### v11

# SPECIAL RELATIVITY

SPECIAL RELATIVITY
Evgeny Saldin Final draft
Applications to astronomy and the accelerator physics
Evgeny Saldin Final draft
New material has been added in the Chapter 4 and Chapter 16. If you have any comments, please send an email to evgueni.saldin@desy.de.
*Preface
Many books have been written on the classical subject of special relativity. However, after years of experience in both relativistic engineering and research, I have come to believe that there is room for a new perspective. This book is not quite like the others, as it aims to shed light on often-overlooked aspects of the subject.
To clarify, this is not a textbook on relativity theory. Instead, it focuses on the nature of special relativistic kinematics, its connection to space and time, and the operational interpretation of coordinate transformations. Every theory contains a number of quantities that can be measured by experiment and expressions that cannot possibly be observed. Whenever we have a theory containing an arbitrary convention, we should examine what parts of the theory depend on the choice of that convention and what parts do not. Unfortunately, this distinction is often overlooked, leading some authors to mistakenly classify certain data as observable when, in reality, they rely on arbitrary choices rather than physical experiments. This oversight results in inconsistencies and paradoxes that should be avoided.
The practical approach adopted in the book should appeal to astronomers, space engineers, accelerator engineers, and more broadly, relativistic engineers. This approach, unusual in the relativistic literature, may be clarified by quoting one of the problems discussed in the text: the new light beam kinematics for rotating frames of references. Since we live on such a rotating (Earth-based) frame of reference, the difference in relativistic kinematics between rotating and non-rotating frames of reference is of great practical as well as theoretical significance. A correct solution of this problem requires the use of relativistic principles even at low velocities since the first-order terms in $(v/c)$, where $v$ is the orbital velocity, play a fundamental role in the non-inertial relativistic kinematics of light propagation.
All the results presented here are derived from the first principles, with every step involving physical reasoning explicitly detailed. To maintain a self-consistent style, auxiliary derivations are provided in the appendices. To encourage readers to develop their own perspectives, each chapter includes a suggested bibliography with relevant remarks. The references list only papers I personally consulted, and while many valuable works remain unmentioned, I apologize for any omission.
I am deeply grateful to my longtime friends Gianluca Geloni and Vitaly Kocharyan for years of discussions on much of the material covered in this book. I also extend my sincere thanks to DESY (Deutsches Electronen-Synchrotron) for the opportunity to work in this fascinating field.
1 Introduction
This book begins with a critical examination of current approaches to special relativity. Traditionally, the theory is framed around Einstein’s two postulates: the principle of relativity and the constancy of the speed of light. However, in a more general geometric approach, the principle of relativity is not a fundamental axiom but rather a consequence of space-time geometry.
We emphasize that the core of special relativity lies in the following fundamental postulate: all physical processes occur in four-dimensional space-time, whose geometry is pseudo-Euclidean. This perspective shapes the presentation of the subject in this book, where the four-dimensional geometric formulation takes precedence over conventional approaches.
The space-time geometric approach accommodates all possible choices of coordinates for reference frames, making the second Einstein postulate—the constancy of the coordinate speed of light—unnecessary in this broader framework. The coordinate speed of light remains isotropic and constant only in Lorentz coordinates, where Einstein’s synchronization of distant clocks and Cartesian space coordinates are applied. Thus, the basic elements of the space-time geometric formulation of the special relativity and the usual Einstein’s formulation, are quite different.
It is important to emphasize that, in practical applications, there are two useful choices for clock synchronization conventions:
(a) Einstein’s convention, which leads to the Lorentz transformations between reference frames.
(b) The absolute time convention, which leads to the Galilean transformations between reference frames.
Absolute time (or simultaneity) can be introduced in special relativity without altering either the logical structure or the convention-independent predictions of the theory. While this choice may seem unconventional within relativity, it is often the most practical when connecting theoretical results to laboratory reality.
It is widely believed that only philosophers of physics discuss the issue of distant clock synchronization. Indeed, a typical physics laboratory does not contain a predefined space-time grid. It is important to recognize that a rule-clock structure exists only in our minds, and the use of hypothetical clocks in special relativity is a necessary prerequisite for applying dynamics and electrodynamics in a coordinate-based framework. This situation often leads physicists to assume that the theory of relativity can be applied to physical processes without a detailed understanding of the clock synchronization procedure. However, many problems in special relativity can only be adequately addressed using a non-standard absolute time synchronization approach. A distinctive feature of this book is its exploration of absolute time coordinatization—specifically, the use of Galilean transformations—within the framework of special relativity.
Chapter 3 presents an ”operational interpretation” of Lorentz and absolute time coordinatizations, making it arguably the most important chapter of this book. Today, asserting the validity of Galilean transformations is considered a ”shocking heresy,” conflicting with the prevailing relativistic intuition and the widely accepted interpretation of special relativity among physicists. The distinction between absolute time synchronization and Einstein’s time synchronization, from an operational perspective, will be a significant revelation for any expert in special relativity. To our knowledge, neither the operational interpretation of absolute time synchronization nor this key distinction has been explored elsewhere in the literature.
We begin by examining the phenomenon of the aberration of light. Light, as a special case of electromagnetic waves, is governed by electrodynamics. It is well known that electrodynamics fully complies with the principles of relativity and, therefore, must accurately describe the properties of light—an inherently relativistic entity.
In Chapters 4 and 5, we undertake a critical reexamination of the existing theory of the aberration of light. Even at the first-order approximation in $v/c$, the phenomenon remains complex, and the literature contains numerous incorrect results. To analyze optical phenomena involving the relative motion of two (or more) light sources, it becomes essential to employ electrodynamics within an absolute time framework. When the sources are independent, introducing Lorentz coordinates poses no issue, as each source can be assigned an individual coordinate system with its own set of clocks. However, the situation changes when a secondary source interacts with the light emitted by a primary source. A key challenge from the perspective of special relativity is that, in this case, a single inertial frame cannot accommodate a common Lorentz coordinatization for both sources. This fundamental difficulty necessitates a deeper examination of how electrodynamics should be applied in such scenarios.
It is widely assumed that when a mirror moves tangentially to its surface, the law of reflection remains the same as for a stationary mirror. However, this common textbook assertion is incorrect. The first fundamental error in textbook explanations is conceptual. Observers typically describe the light beam from a stationary emitter using the Minkowski metric. Textbooks assume that the same metric applies to a moving mirror, incorrectly implying that both the emitter and mirror share a common Lorentz coordinatization within a single inertial frame. This assumption is flawed. The second error arises from the use of an idealized plane wave and an infinite plane mirror in the context of tangential motion. In this scenario, ”mirror motion” is not a physically observable effect. If the mirror were truly infinite, the problem would be time-independent. Only the motion of a finite mirror has physical significance. Experimental methods for measuring aberration increments rely on light signals and thus do not directly measure phase velocity (i.e., frequency and wavevector) but rather the group velocity. What most textbooks overlook is that energy transport is well-defined only when the apertures of both the source and the mirror are explicitly specified.
Consider a point source and an observer, along with their respective measuring devices, all at rest in an inertial laboratory frame. We show that when a finite-aperture mirror moves tangentially at a constant velocity and interacts with radiation in the far field—where the point source effectively produces a plane wave at the mirror—the energy transport of the reflected light experiences a measurable deviation. This effect has important practical implications. According to Babinet’s principle, our theory of light aberration predicts that an analogous deviation occurs for light transmitted through a hole in a moving opaque screen, or equivalently, through the open end of a moving telescope barrel.
Questions related to transmission through a transversely moving pupil detection system (e.g., the moving end of a telescope barrel) often lead to serious misunderstandings. These arise from an inadequate grasp of several complex aspects of statistical optics. Modern explanations suggest that the phenomenon of aberration of light can be interpreted using a ray model. A particularly interesting case is the transmission of light through a moving telescope tube. The rays of light from a star impinge on the telescope tube, and according to the prevailing literature, they do not interact with its sides.
One might naively assume—following textbook reasoning—that the domain of ray optics applies to all spatially incoherent radiation. However, this is a misconception. A completely incoherent spatial source (such as an incandescent lamp or a star) consists of elementary, statistically independent point sources with different offsets. The characteristic dimension of each elementary statistically independent source is approximately $\lambda$, where $\lambda$ is the wavelength of visible radiation. The radiation field generated by such a source can be considered a linear superposition of fields from individually incoherent sources. Each elementary source effectively produces a plane wave in front of the pupil detection system. Moreover, any measuring instrument inevitably influences the detected radiation due to the unavoidable diffraction of a plane wave by an aperture. It is important to note that a linear superposition of radiation fields from elementary point sources preserves fundamental single-point source characteristics, including their independence from source motion.
Chapter 6 explores astronomical applications. Stellar aberration is often considered one of the simplest phenomena in astronomical observations. However, despite its apparent simplicity, it remains one of the most intricate effects of special relativity. A common belief is that stellar aberration depends on the relative velocity between the source (star) and the observer. However, observations clearly indicate that aberration is independent of the relative motion between stars and telescopes on Earth. The lack of symmetry, between the cases when either the source or telescope is moving is shown clearly on the basis of the separation of binary stars. The relative motion of these stars with respect to each other (and hence, with respect to the Earth) is never followed by any aberration, although the motion of these stars is, sometimes, much faster than that of the Earth around the Sun. It should be stressed that it is the telescope and not the star that must change its velocity (relative to the fixed stars) to cause aberration.
The contradiction is so evident that some physicists argue stellar aberration conflicts with the special theory of relativity. Currently, no explanation accounts for why observational data on stellar aberration align with a moving Earth, yet the symmetric case—where the star has relative transverse motion—fails to produce compatible predictions. We show that the absence of widely separated binary stars does not necessitate a fundamental revision of physical principles. Instead, it highlights the need to interpret stellar aberration within a space-time geometric framework.
The phenomenon of the aberration of light is often interpreted, within the corpuscular model of light, as analogous to the oblique fall of raindrops observed by a moving observer. This classical kinematic approach has been employed in astronomy for nearly three centuries to compute stellar aberration. This book presents a relativistic theory of the aberration of light in rotating frames. To explore the effects of relativistic modifications to stellar aberration, we begin with the classical aberration increment, given by $\theta_{a}=v/V$ , where $v$ is the observer’s velocity and $V$ is the velocity of the particles. It is assumed that $v\ll V$. According to the conventional approach, by neglecting terms of order $v^{2}/c^{2}$ in comparison to 1, the stellar aberration increment simplifies to the elementary formula $\theta_{a}=v/c$. This result is particularly notable, as the study of stellar aberration traditionally relies on classical kinematics, specifically the Galilean vectorial law of velocity addition. We conclude that the standard analysis of stellar aberration fails to account for the fundamental distinction between the velocity of light and the velocity of raindrops.
A satisfactory treatment of relativistic modifications should be based on two key relativistic parameters. For an observer on Earth, the velocity relative to the solar reference frame is approximately 30 km/s, corresponding to Earth’s motion around the Sun. In the theory of stellar aberration, we typically consider the small expansion parameter $v/c\simeq 10^{-4}$, neglecting terms of the order of $v^{2}/c^{2}$. However, in addition to this parameter, we must also account for the relativistic quantity $V/c$. For light, $V/c=1$, and according to the special theory of relativity, stellar aberration cannot be treated purely within the framework of classical mechanics. Light is always an ”ultra-relativistic” phenomenon, regardless of how small the ratio $v/c$ may be. Since physical processes occur within the metric structure of space-time, as dictated by the special theory of relativity, the geodesic interval must be used to accurately describe the phenomenon of stellar aberration.
The challenges associated with Earth-based measurements are resolved by recognizing a fundamental asymmetry between Earth-based and Sun-based observers—specifically, the acceleration of the traveling Earth-based observer relative to the fixed stars. We demonstrate that explaining the aberration of light in a rotating frame of reference does not require modifying special relativity or invoking general relativity. Instead, a rigorous application of special relativity is sufficient.
All phenomena in non-inertial reference frames should be analyzed within the space-time geometric framework using the metric tensor. Employing the Langevin metric in a rotating frame of reference allows us to account for all Earth-based experiments. A correct treatment of this problem in a rotating frame necessitates the use of the metric tensor, even for first-order effects in $v/c$. This is because the cross term in the Langevin metric—which represents the first-order deviation of the metric tensor from its Minkowski form—plays a crucial role in the non-inertial kinematics of light beam propagation. The historical progression of studies on optical effects in rotating frames is quite unusual. The Sagnac effect was discovered in 1913 and later described by Langevin in 1921. It is intriguing that the Langevin metric has never been previously applied to stellar aberration theory.
We demonstrate that the aberration of light is a complex phenomenon that must be categorized into various types based on their origin. These categories depend on the cause of the aberration—whether it arises from a change in the observer’s velocity (relative to the fixed stars) or the velocity of the light source. Furthermore, aberration can be further subdivided based on the physical influence of the optical instrument on the measurement.
A theory of stellar aberration within the Earth-based frame of reference must account for two key observations: (1) the annual apparent motion of fixed stars relative to their positions, and (2) the absence of apparent aberration in rotating binary systems (they exhibit aberrations not different from other stars). We present a theory that addresses both of these phenomena while also providing new insights. All Earth-based experiments can be explained by considering the effect of the measuring instrument (i.e., the inevitable physical influence of the telescope on the measurement) and the acceleration of the Earth-bound observer relative to the fixed stars.
In Chapter 10, we analyze the potential for using Earth-based sources to confirm predictions of the relativistic aberration of light theory. We introduce a simple scaling model for stellar aberration, emphasizing that the motion of stars relative to Earth is not accompanied by aberration. Notably, an aberration shift occurs even when a star moves at the same velocity as Earth. We derive a condition for optical similarity between the aberration of light from a distant star moving with Earth’s velocity and that from an Earth-based incoherent source. The proposed method for measuring the aberration angle leverages Earth-based sources and offers a significant advantage: Earth’s rotation should induce a corresponding shift in the observed image. This aberration shift depends solely on $v_{\perp}$, the component of Earth’s orbital velocity perpendicular to its rotation axis. The apparent position of the source image is thus always a little displaced in the direction of the Earth’s motion around the Sun at that moment, and hence describes a small elliptical figure during the annular revolution of the Earth around its axis. In principle, observations could be recorded within a single day.
Special relativity is generally considered a reciprocal theory. However, the aberration of light in an accelerated frame reveals a fundamental asymmetry between accelerated and inertial observers. Notably, without referencing anything external to the Earth-based accelerated frame, one could determine Earth’s velocity relative to the Sun-based frame by measuring the aberration of an Earth-based incoherent light source. While no such experiment has been conducted, astronomical observations confirm what the outcome would be.
Many people who learn special relativity in the usual way find this argument unsettling. The reasoning behind the claim that aberration shift must be symmetrical typically proceeds as follows:
1. A fundamental principle of special relativity states that the metric contains all the necessary information about the physics of a given situation as described in the chosen coordinates.
2. The direction of an observer’s acceleration is encoded in the cross term of the Langevin metric.
3. It is always possible to choose a coordinate system in which the metric of an accelerated frame becomes diagonal. This follows from the pseudo-Euclidean geometry of space-time.
4. Maxwell’s equations hold in all inertial frames, ensuring that light propagates at the same velocity $c$ in each of them. This eliminates any privileged frame, implying that the concept of absolute motion has no physical meaning.
In the argument above, steps (2) and (3) are correct, but step (1), and consequently step (4), are incorrect. Step (3) asserts that all inertial frames are equivalent with respect to physical laws, but not necessarily with respect to physical facts. At first glance, the diagonalization of the Langevin metric might seem to establish symmetry between inertial frames, but where does the asymmetry arise?
The electrodynamics equations alone do not provide a complete description of the physical situation. To solve them, one must also specify the initial conditions. The time after diagonalization is readily obtained by introducing the time offset factor. This time shift introduces a rotation of the plane of simultaneity, which in turn leads to a rotation of the plane wavefront in the accelerated frame. Consequently, after metric diagonalization, the information about the observer’s acceleration is no longer contained in the metric but is instead encoded in the initial conditions—specifically, in the orientation of the radiation wavefront.
One can conclude that not everything is relative in relativity, as the theory also encompasses certain absolute features. Many people assume that, because time and distance in the Lorentz coordinatization have direct physical significance, there must be an underlying physical (dynamical) reason for wavefront rotation. Some believe that the dynamics governed by physical fields are merely concealed within the language of pseudo-Euclidean geometry. However, the paradoxical asymmetry between inertial and accelerated reference frames is best explained through a dynamical perspective. The resolution of this asymmetry paradox identifies the inertial (pseudo-gravitational) force within the accelerated system as the cause of the asymmetry. By applying the principle of equivalence, one can address non-inertial kinematic problems using dynamical methods. In this context, wavefront rotation, which arises when transforming from an inertial frame to an accelerated frame (relative to the fixed stars), can be understood as the effect of a pseudo-gravitational potential gradient during the acceleration process.
Several key points arise from the above results. One interesting question is why we are not discussing the aberration of light emitted by a laser. When considering light aberration, it is essential to distinguish between incoherent and laser sources. Naively, one might expect Earth’s motion around the Sun to introduce an aberration significant enough to affect precise laser-based observations. However, it is important to emphasize that laser light aberration cannot be measured. Inside a laser resonator, the electromagnetic wave travels back and forth, reflecting between mirrors. This repeated reflection effectively cancels out any asymmetry over successive round trips, preventing the manifestation of aberration. A crucial distinction exists between laser and incoherent sources. The radiation from an incoherent source depends on initial conditions, influenced by a mix of spatial and temporal factors—an effect akin to pseudo-gravity experienced by an accelerated observer. In contrast, in an optical resonator, the (3-space) boundary conditions dictate the direction of energy propagation.
Many books claim that no experiment can determine which observer remains ”at rest” during acceleration because, according to the principle of relativity, only relative motion has physical significance. From this, they conclude that any observed asymmetry would contradict the principle of relativity. However, this argument is incorrect. There is no conflict between the fundamental structure of special relativity and the aberration of light phenomena. The principle of special relativity, as stated in Einstein’s writings and every textbook on the subject, asserts that the laws of physics are the same in all inertial frames. The principle of special relativity states that velocity is irrelevant to physical laws, but not to everything.
It is widely assumed that covariant equations must yield covariant solutions. This notion, commonly found in textbooks, is incorrect. For example, from a mathematical perspective, the Lorentz covariance of electrodynamics equations does not necessarily ensure covariant solutions. Here, we emphasize a crucial point: the principle of special relativity, which states that all physical equations must remain invariant under Lorentz transformations (i.e., treating the pseudo-Euclidean geometry of space-time as an axiom of the theory), does not inherently impose reciprocal symmetry in nature.
Textbook authors mistakenly assume that all inertial frames are equivalent. Relativistic (or reciprocal) symmetry is commonly associated with symmetry under Lorentz boosts. However, it is well known that Lorentz boosts alone do not form a group—the composition of two boosts in different directions results in a combination of a pure boost and a spatial (Wigner) rotation. From a mathematical perspective, symmetry between inertial frames requires that the relevant transformations form a group. Since Lorentz boosts alone do not satisfy this criterion, true symmetry between inertial frames is absent. Inertial frames are defined by motion at a constant velocity relative to the fixed stars. A fundamental distinction exists between an accelerated inertial frame (relative to the fixed stars) and one without an acceleration history. In the case of the aberration of light, this difference is closely linked to the Wigner rotation effect.
It is important to note that all well-known tests of special relativity rely on second-order optical (interference) measurements, with the Michelson-Morley experiment being the most prominent example. In such experiments, light from a source is split into two beams, which later recombine to produce an interference pattern. It is not possible to detect the orbital velocity using second-order (interference) experiments.
Phase, as a 4D invariant, is merely a number—it remains unchanged regardless of the chosen inertial frame or coordinate system. This invariance stems from the fundamental geometry of space-time. Consequently, in interference experiments, neither the Doppler effect nor the aberration of light exist as independent, well-defined physical phenomena; only phase has physical significance.
This reasoning explains why special relativity predicts a null fringe shift in Earth-based frames. A careful analysis of all second-order optical experiments within an inertial frame confirms that observed phenomena remain unaffected by uniform motion relative to the fixed stars. While the standard formulation of relativity incorporates the principle of velocity irrelevance, this principle strictly applies in the limiting case of optical second-order measurements.
Beyond interference-based experiments, several other second-order tests exist that do not involve light interference. This raises the question of whether it is possible to determine the motion of an initial inertial frame from within an accelerated frame using time dilation effects. We demonstrate that the slowing of a physical clock accelerated relative to the fixed stars is independent of the reference frame in which this effect is measured. In other words, time dilation is an absolute effect, not a relative one—a fundamental property of pseudo-Euclidean space-time.
Now, let us return to the topic of metric diagonalization. In textbooks, time dilation is typically derived from the diagonal (Minkowski) metric. After diagonalization, these textbooks assert that coordinates in inertial frames should be related through the Lorentz transformation. At first glance, the problem appears completely symmetrical, suggesting that time dilation is a relative effect. However, something is evidently amiss. In Chapter 12, we explore the connection between time dilation and the Langevin metric.
Perhaps the only way to demonstrate that time dilation is absolute is through a mathematical perspective. Mathematically, the argument for the absoluteness of time dilation follows this reasoning: The metric tensor must be a continuous quantity. The Langevin metric is obtained by smoothly transitioning between the metric tensors of an accelerated frame and an inertial frame. To achieve this smooth transition, it is sufficient to relate the coordinates and time of the accelerated observer to those of the inertial observer using a Galilean boost. Under these conditions, the inertial frame’s metric naturally evolves into the Langevin metric of the accelerated frame.
However, after diagonalization, the metric tensor in the accelerated frame must abruptly shift from the Langevin metric to the Minkowski (diagonal) form. This discontinuity leads to a critical conclusion: when applying the Lorentz transformation, if the transformation fails to preserve the continuity of the metric tensor, then symmetry between inertial frames does not hold. Thus, any change in velocity or acceleration—relative to the fixed stars—has an absolute significance.
The formalism of relativistic physics is fundamentally tied to the concept of an absolute space-time structure. This structure is inherently linked to an initial inertial frame, which, in turn, serves as an absolute rest frame. Consequently, uniform motion is not relative, and a particle’s absolute velocity can be determined within this initial frame. The key distinction between the initial inertial frame and Newton’s concept of absolute space lies in the lessons of special relativity, which introduced a unified space-time model. In this framework, Newton’s separate notions of absolute space and absolute time are merged into a single, cohesive space-time continuum.
The proper time interval can be associated with various physical phenomena, including particle lifetimes, atomic transition periods, and nuclear half-life. Until now, all tests of time dilation using atomic clocks in an Earth-based frame have involved circular path geometries. For instance, atomic clocks flown on commercial jets were compared with a reference clock upon returning to the laboratory, ensuring that both were at rest relative to the Earth.
According to special relativity, any time dilation experiment investigating the effects of orbital velocity should use a linear (one-way) path geometry. This is typically observed in experiments involving the decay of rapidly moving unstable particles. However, current experimental techniques lack the sensitivity required to detect the influence of orbital velocity on the average distance traveled by a decaying particle.
In contrast, we describe (see Chapter 10) a simple, first-order optical experiment accessible to most undergraduate physics departments. This experiment illustrates key principles in the foundations of special relativity. Specifically, it demonstrates that the Earth’s speed relative to a Sun-based reference frame could be determined using measurements of aberration from an Earth-based incoherent light source.
In Chapter 16, we describe the particle dynamics method in the Lorentz lab frame, using lab time $t$ as the evolution parameter. The study of relativistic particle motion in a constant magnetic field, as traditionally approached in accelerator engineering, remains closely tied to classical (Newtonian) kinematics. Specifically, it relies on the Galilean vectorial law for velocity addition. This approach, which is non-covariant, implicitly assumes an absolute time coordinatization—though this remains hidden within the formalism.
The absolute time synchronization convention is self-evident, which is why it is rarely discussed in accelerator physics. Standard textbooks suggest that when only a single frame is considered, a detailed understanding of relativity is unnecessary. In this conventional approach, the only modification to classical mechanics is the introduction of relativistic mass, thereby maintaining a distinction between time and space.
From the lab frame perspective, conventional particle tracking treats a trajectory as the result of successive Galilean boosts that follow the motion of an accelerated particle. The preference for a non-covariant approach within the framework of dynamics stems from its simplicity and practicality. By avoiding the intermixing of spatial coordinates and time, the (3+1)-dimensional non-covariant tracking method remains intuitive, transparent, and well-suited to laboratory conditions.
It can be demonstrated that this approach poses no fundamental difficulties in mechanics or electrodynamics. It is both effective and consistent, as the choice of transformation does not alter the underlying physical reality. What matters is that once a particular convention is adopted, it must be applied consistently across both dynamics and electrodynamics.
A common mistake in accelerator physics arises from an incorrect algorithm for solving electromagnetic field equations. When using Maxwell’s equations, only the covariant formulation of the dynamics equations (i.e., in Lorentz coordinates) ensures the correct coupling between Maxwell’s equations and particle trajectories in the lab frame.
Interestingly, employing the conventional coupling of Maxwell’s equations with a corrected Newtonian equation to calculate radiation from a moving source does not always lead to errors. In cases where the source moves rectilinearly along with the emitted light beam, both covariant and non-covariant approaches yield identical trajectories, making Maxwell’s equations consistent with conventional particle tracking. However, this method was mistakenly extended to non-collinear geometries, where it breaks down.
To better understand this issue, we first examine the reasoning presented in textbooks. It is generally assumed that, in Lorentz coordinates, the magnetic field can only alter an electron’s direction of motion, not its speed. However, there is a counterargument to this widely accepted derivation of velocity composition.
Consider the composition of two velocities that do not lie along a straight line. Conventional particle tracking, following Newtonian mechanics, treats velocity as an ordinary vector and combines them geometrically using the parallelogram method. However, this approach fails to account for relativistic effects and leads to inconsistencies in non-collinear motion.
Only by solving the equations of motion in their fully covariant form can one correctly capture the coupling between Maxwell’s equations and particle trajectories in the lab frame. Solving problems related to covariant particle tracking presents two key challenges. First, conventional particle tracking within a single inertial frame is typically based on classical Newtonian mechanics, which does not incorporate Lorentz transformations. Second, there is considerable confusion regarding the combination of non-parallel velocities in special relativity. The standard textbook treatment of relativistic velocity transformation implicitly assumes that the $(x^{\prime},y^{\prime},z^{\prime})$ axes of a moving observer remain parallel to the $(x,y,z)$ axes of the lab frame. In other words, it presumes that both observers share a common three-dimensional space—a misconception. Textbooks often overlook the fact that in Lorentz coordinatization, two observers following different trajectories experience different three-spaces. Special relativity lacks an absolute notion of simultaneity, leading to a fundamental mixing of spatial and temporal coordinates. For one observer, spatial measurements inherently include a contribution from the time coordinate as perceived by another observer. Consequently, there is no well-defined concept of an instantaneous three-space.
At first glance, textbooks on special relativity typically derive the relativistic velocity addition formula analytically from the combination of Lorentz boosts. Many authors argue that the principle of relativity requires the velocity addition law to obey a group composition structure. However, the non-associativity of Einstein’s velocity addition for non-collinear velocities is often overlooked, despite being demonstrable through straightforward algebra. Consequently, constructing a fully covariant trajectory using only Einstein’s velocity addition is impossible—additional structure is needed.
We emphasize that the addition of non-collinear velocities in Lorentz-coordinatized spacetime is governed by the Wigner rotation. A key consequence of the non-commutativity of non-collinear Lorentz boosts is the lack of a shared, global ”ordinary” space. It is crucial to recognize that Einstein’s velocity addition does not directly follow from the full structure of Lorentz transformations. Textbook derivations of transverse (perpendicular) velocity addition often invoke the relation $dt^{\prime}=dt/\gamma$, yet fail to account for the Wigner rotation, which introduces a coupling between transverse position and time.
Only by solving the equations of motion within a fully covariant framework—one that includes the Wigner rotation in the treatment of non-collinear velocities—can the coupling between Maxwell’s equations and particle trajectories in the laboratory frame be properly described.
In Chapter 16, we explore which aspects of dynamical theory depend on the choice of synchronization convention and which remain independent. Our analysis highlights the distinction between the concepts of path and trajectory. A path is a purely geometric notion, representing the spatial course of a particle without reference to time. In contrast, a trajectory provides more information, as it specifies not only the particle’s spatial position but also the corresponding time instant. The path has a precise, objective meaning. For example, in a magnetic field, the curvature radius of the path—and consequently, the three-momentum—is convention-invariant and thus objectively defined. However, the trajectory $\vec{x}(t)$ is convention-dependent, reflecting the inherent conventionality of velocity, and therefore lacks an exact objective meaning. While dynamical theory involves the concept of a particle’s trajectory, direct verification of it is unnecessary. Instead, the trajectory serves as a theoretical tool in the analysis of electrodynamics problems.
In Chapter 18, we present a critical reexamination of radiation theory. To evaluate radiation fields generated by external sources, we must determine the particle velocity $\vec{v}$ and position $\vec{x}$ as functions of the lab-frame time $t$. As discussed earlier, Maxwell’s equations should be solved in the lab frame, incorporating the current and charge density produced by particles moving along their covariant trajectories. For arbitrary values of $v/c$, covariant calculations of radiation processes become highly complex. However, in certain cases, significant simplifications are possible. One such case is a non-relativistic radiation setup. The non-relativistic assumption enables the dipole approximation, which is of great practical importance. When considering only the dipole component of radiation, we disregard all details of the electron trajectory. Consequently, dipole radiation is insensitive to differences between covariant and non-covariant particle trajectories. However, this is merely the first and most practically relevant term in the expansion. To compute higher-order corrections beyond the dipole approximation, detailed knowledge of the electron trajectory is required. In such cases, relying on the non-covariant approach is insufficient, and the covariant trajectory must be used instead.
It is generally assumed that Maxwell’s equations and the corrected form of Newton’s second law can explain all experiments conducted within a single inertial (laboratory) frame. According to standard textbooks, if only one frame is involved, knowledge of relativity is deemed unnecessary. Many physicists, having learned special relativity from textbooks, might say, “For those who want just enough understanding to solve problems, special relativity merely modifies Newton’s laws by introducing a correction factor to mass.” However, this is a misconception. The velocity of light transforms like that of a particle, and one cannot simply apply classical kinematics to mechanics while using Einstein’s kinematics for electrodynamics. We highlight errors in the standard approach to coupling fields and particles through a relatively simple example, ensuring that the core physical principles remain clear and unobscured by unnecessary mathematical complexity. We hope this example will draw greater attention to the crucial role of Wigner rotation in the transformation of non-collinear velocities.
Accelerator physicists studying the application of relativity theory to synchrotron radiation often face a fundamental issue: the distinction between covariant and non-covariant particle trajectories has never been fully understood. Consequently, the contribution of relativistic kinematic effects to synchrotron radiation has been overlooked. Traditionally, accelerator physics has relied on Newtonian kinematics, which is incompatible with Maxwell’s equations. This raises an important question: if storage rings are designed without accounting for relativistic kinematics, how do they function? In reality, electron dynamics in storage rings are significantly influenced by radiation emission. We address this question in detail in Chapter 19, focusing on spontaneous synchrotron radiation from bending magnets and undulators.
Just as the non-relativistic asymptote simplifies calculations, the ultrarelativistic asymptote does the same for covariant analysis, as it naturally leads to the paraxial approximation. The motion of electrons in a bending magnet exhibits cylindrical symmetry, which gives rise to several remarkable effects. We demonstrate that covariant particle tracking is unnecessary for deriving bending magnet radiation. However, there is one case where the conventional approach fails: the covariant framework predicts a nonzero redshift of the critical frequency due to perturbations of electron motion along the magnetic field.
A direct laboratory test of synchrotron radiation theory could highlight the incompatibility between standard relativistic electrodynamics—based on Maxwell’s equations—and non-covariant particle tracking. Despite decades of measurements, synchrotron radiation theory remains experimentally unverified in certain aspects. We explore the potential of synchrotron radiation sources to test the predictions of a revised synchrotron radiation theory, proposing experiments that have never been conducted at existing facilities.
The theory of relativity demonstrates that the relativity of simultaneity—a fundamental effect of relativistic kinematics—is closely linked to extended relativistic objects. Until the 21st century, no macroscopic objects were known to travel at relativistic speeds, and it was widely believed that only microscopic particles in experiments could reach velocities close to the speed of light. However, the 2010s saw rapid advancements in laser light sources in the X-ray wavelength range. An X-ray free electron laser (XFEL) is an example where improvements in accelerator technology make it possible to develop ultrarelativistic macroscopic objects with an internal structure (modulated electron bunches), and the relativistic kinematics plays an essential role in their description.
In Chapter 20, we critically reexamine existing XFEL theory, particularly for readers with limited knowledge of accelerator physics. We focus on the production of coherent undulator radiation by a modulated ultrarelativistic electron beam. Fortunately, understanding this process does not require prior knowledge of undulator radiation theory from Chapter 19, as it can be explained in a straightforward manner. Relativistic kinematics plays a fundamental role in XFEL physics, particularly through the rotation of the modulation wavefront, which, in the ultrarelativistic approximation, is closely linked to the relativity of simultaneity. When particle trajectories are analyzed in a Lorentz reference frame—an inertial frame where time coordinates are assigned using the Einstein synchronization procedure—relativistic kinematic effects, such as the relativity of simultaneity, must be considered. In the ultrarelativistic limit, the orientation of the modulation wavefront, i.e. the orientation of the plane of simultaneity, is always perpendicular to the electron beam velocity when the evolution of the beam’s evolution is described using Lorentz coordinates.
It is important to note that Maxwell’s equations are valid only in Lorentz reference frames. Einstein’s time order must be consistently applied in both dynamics and electrodynamics. Maxwell’s equations dictate that the wavefront of a laser beam is always orthogonal to its propagation direction. At this point, we emphasize that, according to relativity, a modulated electron beam in the ultrarelativistic asymptotic limit exhibits the same kinematics as a laser beam in Lorentz coordinates. Experiments show that this prediction is, in fact, true.
The theory of relativity, formulated within a space-time framework based on pseudo-Euclidean geometry, has been developed for over a century. Recently, it has experienced a rapid expansion in practical applications, particularly in XFEL physics.
2 A Critical Survey of Present Approaches to Special Relativity
2.1 What is Special Relativity?
The laws of physics are invariant under Lorentz transformations. This invariance is a restrictive principle—it does not, by itself, determine the precise form of the dynamics involved. Understanding the postulates of special relativity is analogous to grasping the principle of energy conservation: initially, we accept it as a foundational principle, and later, we seek microscopic interpretations that are consistent with it. For any system governed by energy conservation, there should exist a more fundamental theory that provides insight into the detailed physical processes. Naturally, this deeper theory must also obey the principle of energy conservation.
The conservation of energy is a powerful tool, allowing us to analyze systems even in the absence of a complete fundamental theory. A similar methodological perspective applies to the postulates of special relativity. For instance, suppose we do not understand the microscopic reasons for muon decay, but we do know the decay law in the muon’s rest frame. This decay law, in such a case, serves as a phenomenological model. Its relativistic generalization to any Lorentz frame allows us to predict, for example, the average distance a muon travels before decaying.
Consider the case of muon decay. Suppose we lack a detailed understanding of why a muon disintegrates but have empirical knowledge of its decay law in its rest frame. This law serves as a phenomenological model. By applying relativistic principles, we can extend this law to an arbitrary Lorentz frame, enabling us to predict the muon’s average travel distance. Specifically, if we apply a Lorentz transformation to the decay law, we find that after a distance $\gamma v\tau_{0}$, the muon population in the lab frame is reduced to half of its original value. Here, $\gamma$ is the relativistic factor, and $\tau_{0}$ is the proper lifetime. This result can be interpreted as an increase in the effective lifetime of the muon from $\tau_{0}$ to $\gamma\tau_{0}$ in the lab frame due to time dilation.
However, special relativity is not a complete theory. Constructive (or microscopic) theories, such as electrodynamics or quantum field theory, offer a deeper understanding of physical processes than purely restrictive frameworks like special relativity. Relativistic kinematics serves primarily as a descriptive tool for interpreting how matter fields appear to behave from the perspective of different observers. In principle, any relativistic quantity can be derived directly from the underlying theories of matter, without explicitly invoking relativity.
For example, muons exhibit relativistic behavior in motion because the field forces governing their decay obey Lorentz-covariant quantum field equations. In the ”microscopic” view of relativistic phenomena, the Lorentz covariance of the fundamental laws of physics remains, much like energy conservation, an empirical and unexplained fact—but ultimately, all explanations must terminate at some foundational level.
2.2 Different Approaches to Special Relativity
In literature, three main approaches to special relativity are discussed: Einstein’s approach, the standard covariant approach, and the space-time geometric approach.
Einstein’s formulation is based on two fundamental postulates: the principle of relativity and the constancy of the speed of light.
The standard covariant formulation describes special relativity in terms of pseudo-Euclidean space-time geometry and the invariance of the space-time interval $ds$. However, this interpretation is often limited to cases where the metric is strictly diagonal. Assuming a diagonal metric implicitly imposes Lorentz coordinates, ensuring that different inertial frames are related by Lorentz transformations.
The space-time geometric approach, on the other hand, places primary emphasis on the geometry of space-time. In this framework, space-time is assumed to have a pseudo-Euclidean structure, and only four-tensor quantities are considered physically meaningful. Unlike Einstein’s formulation, where the principle of relativity is postulated, in this approach, it naturally follows from the geometric properties of space-time.
Since the space-time geometric approach accommodates all possible coordinate choices within a given reference frame, Einstein’s second postulate—asserting the constancy of the coordinate speed of light—does not generally hold. The coordinate speed of light remains isotropic and constant only in Lorentz coordinates, where Einstein’s synchronization of distant clocks and Cartesian space coordinates are used.
2.2.1 The Usual Einstein’s Approach
Traditionally, the special theory of relativity is founded on two key postulates:
1. Principle of relativity. The laws of nature are the same (or take the same form) in all inertial frames
2. Constancy of the speed of light. Light propagates with constant velocity $c$ independently of the direction of propagation, and of the velocity of its source.
However, the constancy of the speed of light in all inertial frames is not a fundamental statement of relativity. The core principle of special relativity is the Lorentz covariance of all fundamental physical laws. Contrary to common textbook presentations, the second postulate is not an independent physical assumption but rather a convention that cannot be experimentally tested.
By assuming the constancy of the speed of light in all inertial frames, we implicitly adopt Lorentz coordinates and assume that different inertial frames are related by Lorentz transformations. In this restrictive interpretation of relativity, only Lorentz transformations are used to map event coordinates between inertial observers.
2.2.2 The Usual Covariant Approach
In the standard covariant approach, special relativity is formulated as a theory of space-time with pseudo-Euclidean geometry. Physical quantities are represented by tensors in four-dimensional space-time, making them covariant, and the laws of physics are expressed as four-tensor equations to ensure manifest covariance.
To develop space-time geometry, a metric or measure $ds$ of space-time intervals must be introduced. The choice of measure determines the nature of the geometry. In this framework, an event is mathematically represented by a point in space-time, called a world-point, while the evolution of a particle is depicted as a curve in space-time, known as a world-line. If $ds$ represents the infinitesimal displacement along a particle’s world-line, then
$ds^{2}=c^{2}dT^{2}-dX^{2}-dY^{2}-dZ^{2}~{},$ (1)
where we have selected a special type of coordinate system (a Lorentz coordinate system), defined by the requirement that Eq. (1) holds.
To simplify our writing we will use, instead of variables $T,X,Y,Z$, variables $X^{0}=cT,~{}X^{1}=X,~{}X^{2}=Y,~{}X^{3}=Z$. Then, by adopting the tensor notation, Eq. (1) becomes $ds^{2}=\eta_{ij}dX^{i}dX^{j}$, where Einstein summation is understood. Here $\eta_{ij}$ are the Cartesian components of the metric tensor and by definition, in any Lorentz system, they are given by $\eta_{ij}=\mathrm{diag}[1,-1,-1,-1]$, which is the metric canonical, diagonal form. As a consequence of the space-time geometry, Lorentz coordinate systems are connected by Lorentz transformations.
Physical quantities are represented by space-time geometric (tensor) entities. When a basis is introduced, a tensor as a geometric object consists of both its components and the basis itself. In the conventional covariant approach, one typically considers only the components of tensors in a Lorentz coordinate system—i.e., under the assumption that the basis four-vectors are orthogonal. Consequently, physical laws are expressed exclusively through four-tensor equations written in component form.
However, in this approach, the definition of a tensor relies on the transformation properties of its components. For instance, in standard covariant formalism, the electromagnetic ”tensor” $F^{\mu\nu}$ is not actually a full tensor; rather, it consists only of component values implicitly taken in an orthogonal basis. These components are coordinate-dependent quantities and do not fully describe the physical entity, as they lack explicit reference to the space-time basis itself. This limitation is not an issue when restricting transformations to those between orthogonal bases, such as Lorentz transformations, which are assumed to correctly map event coordinates. However, according to the conventional covariant approach, transformations from a standard orthogonal basis to a non-orthogonal one—such as Galilean transformations—are considered ”incorrect.”
2.2.3 The Space-Time Geometric Approach
We emphasize the significant freedom in choosing a coordinate system for Minkowski space-time. The space-time continuum, defined by the interval in Eq. (1), can be described using arbitrary coordinates, not just Lorentz coordinates. When transitioning to an arbitrary coordinate system, the geometry of four-dimensional space-time remains unchanged. In special relativity, there are no restrictions on the choice of coordinate system: the spatial coordinates $x^{1},x^{2},x^{3}$ can be any parameters defining the position of particles, while the time coordinate $x^{0}$ can correspond to any arbitrary clock.
The components of the metric tensor in the coordinate system $x^{i}$ can be determined by performing the transformation from the Lorentz coordinates $X^{i}$ to the arbitrary variables $x^{j}$, which are fixed as $X^{i}=f^{i}(x^{j})$. One then obtains
$ds^{2}=\eta_{ij}dX^{i}dX^{j}=\eta_{ij}\frac{\partial X^{i}}{\partial x^{k}}\frac{\partial X^{j}}{\partial x^{m}}dx^{k}dx^{m}=g_{km}dx^{k}dx^{m}~{},$ (2)
In textbooks and monographs, special relativity is typically presented in terms of an interval $ds$ expressed in the Minkowski form of Eq. (1), whereas Eq. (2) is often associated with general relativity. However, in the space-time geometric approach, special relativity is fundamentally a theory of four-dimensional space-time with pseudo-Euclidean geometry. In this formulation, the space-time continuum can be described equivalently in any coordinate system, provided that $ds$ remains invariant. In contrast, the conventional presentation of special relativity also relies on the invariance of $ds$ but interprets it in a more restricted sense, assuming a strictly diagonal metric.
Textbook presentations of the special theory of relativity typically follow Einstein’s approach or, more generally, the standard covariant approach, which, as discussed above, deals only with the components of 4-tensors in specific (orthogonal) Lorentz bases. However, these presentations often overlook the fact that transitioning to arbitrary coordinates does not alter the underlying geometry of space-time. As a result, a common misconception among experts is that transforming from an orthogonal Lorentz basis to a nonorthogonal one is incorrect, while a Lorentz transformation—between two orthogonal Lorentz bases—is valid. This belief is unfounded. Physics can be described in any coordinate system; changes in coordinates merely affect how the components of 4-tensors are expressed, without altering the tensors themselves. Although Einstein synchronization (i.e., the choice of Lorentz coordinates) is preferred by physicists for its simplicity and symmetry, it is no more ”physical” than any other coordinate system. In this book, we will explore an especially unconventional choice: the absolute time coordinate system.
2.3 The Myth About the Incorrectness of Galilean Transformations
The role of Galilean transformations within the framework of relativity requires careful reconsideration. Many physicists continue to view Galilean transformations—specifically, transformations from an orthogonal Lorentz basis to a non-orthogonal one—as outdated and incorrect when applied to spatial coordinates and time. A common argument against their validity is that they fail to preserve the form-invariance of Maxwell’s equations under a change of inertial frame. This reasoning is frequently found in well-known textbooks and monographs  [^1] It is widely believed that Galilean transformations, being pre-relativistic, are incompatible with the special theory of relativity. For instance, Bohm [1] states: “… the Galilean law of addition of velocities implies that the speed of light should vary with the speed of the observing equipment. Since this predicted variation is contrary to fact, the Galilean transformations evidently cannot be correct.” Similar assertions appear in more recent pedagogical literature. Drake and Purvis [2], for example, write: “One of the great insights of relativity theory was the realization that Galilean transformations are incorrect. The proper way to translate space-time measurements between inertial frames is through Lorentz transformations.”.
However, the assertion that Galilean transformations are inherently incorrect is a misconception. Special relativity is fundamentally a theory of four-dimensional space-time with a pseudo-Euclidean geometry. From this perspective, the principle of relativity naturally emerges as a geometric property of space-time, which can be described using arbitrary coordinate systems  [^2] The mathematical claim that transitioning to arbitrary coordinates leaves the geometry of space-time unchanged is considered erroneous in standard textbooks. As L. Landau and E. Lifshitz note [3]: ”This formula is called the Galilean transformation. It is easy to verify that this transformation, as was to be expected, does not satisfy the requirements of the theory of relativity; it does not leave the interval between events invariant.” This misconception is attributed to a failure to distinguish between the convention-dependent and convention-invariant aspects of the theory. In pseudo-Euclidean geometry, the interval between events remains invariant under arbitrary coordinate transformations.. Thus, contrary to the standard presentation in most textbooks, Galilean transformations are in fact compatible with the principle of relativity, even though they modify the explicit form of Maxwell’s equations.
2.4 The Non-Relativistic Limit of Lorentz Transformations
It is commonly assumed that the Lorentz transformation reduces to the Galilean transformation in the non-relativistic limit. However, we argue that this widely accepted textbook claim is incorrect and misleading. Kinematics inherently involves a comparative study between two coordinate systems, requiring the assignment of time coordinates to both. The method of clock synchronization determines these time coordinates, and different synchronization conventions lead to different time assignments.
In Minkowski space-time, choosing a synchronization convention is equivalent to selecting a specific coordinate system within an inertial frame of reference. Practical considerations may favor one coordinate system over another in different contexts. In relativistic engineering, we typically choose between the absolute time coordinate and the Lorentz time coordinate. Importantly, the space-time continuum can be equally well described in both frameworks. This implies that, for arbitrary particle speeds, Galilean coordinate transformations effectively characterize the change in reference frame from a laboratory inertial observer to a comoving inertial observer, even within the framework of relativity.
Now, let us examine the non-relativistic limit. For $v/c$ small enough that terms of order $v^{2}/c^{2}$ can be neglected, the Lorentz transformation simplifies to $x^{\prime}=x-vt$, $t^{\prime}=t-xv/c^{2}$. This infinitesimal Lorentz transformation differs from the infinitesimal Galilean transformation $x^{\prime}=x-vt$, $t^{\prime}=t$. The difference lies in the additional term $xv/c^{2}$ in the Lorentz transformation for time, which is a first-order correction.
We wish to emphasize the following point: An infinitesimal Lorentz transformation differs from the Galilean transformation only by incorporating the relativity of simultaneity. This is the sole relativistic effect that appears in the first order of $v/c$. Other higher-order effects, such as Lorentz-Fitzgerald contraction and time dilation, arise from higher-order terms and can be derived mathematically by iterating the infinitesimal transformation  [^3] There is a common misconception that the Lorentz transformation reduces to the Galilean transformation in the non-relativistic limit. For example, French [4] states, ”The reduction of $t^{\prime}=\gamma(t-vx/c^{2})$ to the Galilean relation $t^{\prime}=t$ requires $x\ll ct$ as well as $v/c\ll 1$”. Similar statements are found in recent textbooks, such as Rafelski [5], who asserts: ”A wealth of daily experience shows that the Galilean coordinate transformation (GT) is correct in the nonrelativistic limit in which the speed of light is so large that it plays no physical role. Any coordinate transformation replacing the GT must also agree with this experience, and thus must contain the GT in the nonrelativistic limit.” This conclusion is mathematically flawed. As Baierlein [6] points out: ”If the Lorentz transformation for infinitesimal $v/c$ were to reduce to the Galilean transformation, then the iterative process could never generate a finite Lorentz transformation that is radically different from the Galilean transformation..
The primary distinction between Lorentz coordinatization and absolute time coordinatization lies in the transformation laws that relate coordinates and time between relatively moving systems. It is incorrect to assert, as some textbooks do, that the expression $t^{\prime}=\gamma(t-vx/c^{2})$ reduces to the Galilean relation $t^{\prime}=t$ in the non-relativistic limit. Such a claim would imply that infinitesimal Lorentz transformations become identical to infinitesimal Galilean transformations in this limit, which is not the case. The essence of Lorentz (or Galilean) transformations is captured in their infinitesimal form: relativistic kinematic effects cannot be derived simply by iterating infinitesimal Galilean transformations through a mathematical procedure.
2.5 The Myth about the Constancy of the Speed of Light
It is widely believed that experiments demonstrate the independence of the speed of light in a vacuum from both the source and the observer  [^4] Many textbook authors still treat conventional quantities as if they are empirically measurable. For instance, Cristodoulides [7] states: ”The fact that the Galilean transformation does not preserve Maxwell’s equations has already been mentioned […] On the other hand, experiments show that the speed of light in a vacuum is independent of the source or observer.”. This claim is commonly presented in textbooks but is, in fact, incorrect. The constancy of the speed of light is not an empirical fact but rather a consequence of the chosen synchronization convention, making it impossible to test experimentally  [^5] Since we can empirically access only the round-trip average speed of light, any claims about the magnitude and isotropy of the one-way speed of light depend on the chosen time coordinate system. Such values vary with the synchronization scheme used. As Anderson, Vetharaniam, and Stedman [8] state: ”No experiment, then, is a ’one-way’ experiment. An empirical test of any property of the one-way speed of light is not possible. Such quantities as the one-way speed of light are irreducibly conventional in nature, and recognizing this aspect is to recognize a profound feature of nature.”.
To measure the one-way speed of light, one must first synchronize an infinite set of clocks distributed across space to enable time measurements. However, an inherent circularity arises when clocks are synchronized under the assumption that the one-way speed of light is $c$. If the synchronization process is based on this assumption, then any measurement using these clocks will necessarily yield $c$, as the measurement framework has been defined accordingly.
Thus, the one-way speed of light is a matter of convention rather than a physical quantity with empirical meaning. In contrast, the two-way (or round-trip) speed of light has physical significance because it is directly measurable without reliance on clock synchronization. Round-trip experiments involve observing simultaneity (or lack thereof) at a single spatial location, making them independent of any synchronization convention. All well-established methods for measuring the speed of light are, in fact, round-trip measurements. A prime example is the Michelson-Morley experiment, which employs an interferometer to compare light beams in a two-way manner.
Apparent paradoxes emerge when Galilean transformations are applied within electrodynamics. Many sources claim that Galilean velocity transformations are inconsistent with the electron-theoretic explanations of optical phenomena such as refraction and reflection. This notion is widespread in standard literature. Pauli, for instance, states [9]: ”It is essential that the spherical waves emitted by the dipoles in the body interfere with the incident wave. If we consider the body at rest while the light source moves relative to it, the waves emitted by the dipoles will have a velocity different from that of the incident wave, making interference impossible.”
However, such claims stem from a misunderstanding. It is often asserted that when a light source moves relative to a stationary medium (like glass), the resulting wave emitted by the oscillating electrons in the glass cannot interfere properly with the incident wave due to differing velocities. This is incorrect. Regardless of the light’s velocity, the frequency of the incident wave determines the oscillation frequency of the electrons in the medium. These electrons then re-emit light of the same frequency. Therefore, both the incident and re-emitted waves maintain the same frequency and can interfere. The velocity difference merely results in a spatially varying phase difference, not a breakdown in interference.
2.6 Convention-Dependent Aspects of the Theory
Consider the motion of a charged particle in a given magnetic field. According to the theory of relativity, the trajectory $\vec{x}(t)$ of the particle in the lab frame depends on the chosen synchronization convention for clocks in that frame. Whenever a theory involves an arbitrary convention, it is essential to examine which aspects of the theory depend on that choice and which do not. The former are convention-dependent, while the latter are convention-invariant. Clearly, physically meaningful measurement results must be convention-invariant.
Now, consider the motion of two charged particles in a given magnetic field, which controls their trajectories. Suppose there are two apertures located at points $A$ and $A^{\prime}$. From the solution of the equations of motion, we conclude that the first particle passes through aperture $A$ while the second passes through aperture $A^{\prime}$ simultaneously. The occurrence of these two events—particles passing through $A$ and $A^{\prime}$—has an exact objective meaning, making them convention-invariant. However, the simultaneity of these events is convention-dependent and lacks an exact objective meaning. It is important to emphasize that, consistent with the conventionality of simultaneity, the value of the particle’s speed is also a matter of convention and does not possess a definite objective meaning.
To determine which aspects of dynamical theory depend on a chosen convention and which do not, we examine the distinction between the notions of path and trajectory. Consider the motion of a particle in three-dimensional space, described by the vector-valued function $\vec{x}(t)$ The particle follows a prescribed curve, or path, as it moves. This motion along the path is parameterized by $l(t)$, where $l$ represents a certain parameter—specifically, in our case of interest, the arc length.
Now, if we take the origin of a Cartesian coordinate system and connect it to a point on the path, the resulting vector is the position vector $\vec{x}(l)$  [^6] For a more detailed discussion on the distinction between path and trajectory, see [10].. The key difference between trajectory $\vec{x}(t)$ and path $\vec{x}(l)$ is fundamental: the path has an exact, objective meaning—it is independent of any convention. In contrast, and consistent with the inherent conventionality of velocity, the trajectory $\vec{x}(t)$ is convention-dependent and lacks an exact objective interpretation.
To illustrate this distinction, consider experiments in accelerator physics. Suppose we wish to measure a particle’s momentum. A uniform magnetic field can be used to construct a momentum analyzer for high-energy charged particles. Importantly, this method of determining momentum is convention-independent. The curvature radius of the path within the magnetic field—and consequently, the three-momentum—has an objective, convention-invariant meaning. While dynamical theory includes the concept of a particle’s trajectory, this quantity is not directly measured but instead serves as a tool in the analysis of electrodynamics problems.
2.7 Relativistic Time Dilation and Length Contraction
Experts in the theory of relativity often mistakenly conflate the fundamental properties of Minkowski space-time with the specific forms that certain convention-dependent quantities take under standard Lorentz coordinatization. These quantities, commonly referred to as ”relativistic kinematic effects,” include time dilation, length contraction, and Einstein’s velocity addition. There is a widespread misconception that such effects have direct physical significance, rather than being artifacts of the chosen synchronization convention.
Relativistic kinematic effects are inherently coordinate-dependent and lack absolute physical meaning. As noted by Leubner, Auflinger, and Krumm [11], many physicists incorrectly equate the standard formulations of these effects with relativity itself. Their study demonstrates that adopting a different synchronization convention can lead to significant changes in the appearance of these quantities.
For instance, under Lorentz coordinatization, one observes time dilation. However, in a framework based on absolute time coordinatization, relativistic kinematic effects disappear—no time dilation is observed. Despite such variations, all coordinate-independent quantities, such as the particle path $\vec{x}(l)$ and momentum $\vec{p}$, remain unaffected by changes in clock synchronization.
It is a misconception to believe that physics must rely solely on concepts that are directly measurable. Special relativity, when formulated in Lorentz coordinates, includes constructs such as time dilation and length contraction that are not directly observable. Nevertheless, these constructs are essential for making theoretical predictions. The strength of a theory lies in its predictive power, not in the observability of every individual element.
This idea can be further illustrated. Both the velocity of an electron and the speed of light are convention-dependent quantities. In absolute-time coordinatization, the speed of light differs from the electrodynamic constant $c$, as Maxwell’s equations are not invariant under Galilean transformations. Conversely, in Lorentz coordinatization, the speed of light equals $c$, while the electron’s velocity is changed due to relativistic effects. Yet the dimensionless ratio of the electron’s velocity to the speed of light remains convention-independent. This ratio, unaffected by changes in clock synchronization or rhythm, forms the basis for physically meaningful predictions in electrodynamics—for example, in the analysis of synchrotron radiation emitted by bending magnets.
2.8 Relativistic Particle Dynamics
The accelerated motion of a relativistic charged particle is governed by a covariant equation of motion under the influence of the four-force in the Lorentz lab frame. The particle’s trajectory, $\vec{x}_{cov}(t)$, as observed in this frame, results from a sequence of infinitesimal Lorentz transformations. Importantly, the lab frame time $t$ in the equation of motion cannot be treated as independent of the spatial coordinates. This dependence arises because Lorentz transformations inherently mix space and time coordinates, with relativistic kinematic effects manifesting as a consequence of the relativity of simultaneity.
Let us consider the conventional approach to particle tracking. It is generally accepted that, to describe the dynamics of relativistic particles in an inertial laboratory reference frame, one only needs to account for the relativistic dependence of momentum on velocity. The treatment of relativistic particle dynamics is based on a modified form of Newton’s second law. In a given lab frame, there is an electric field $\vec{E}$ and magnetic field $\vec{B}$. They push on a particle in accordance with
$\frac{d\vec{p}}{dt}=e\left(\vec{E}+\frac{\vec{v}}{c}\times\vec{B}\right)~{},$ (3)
$\vec{p}=m\vec{v}\left(1-\frac{v^{2}}{c^{2}}\right)^{-1/2}~{},$ (4)
where here the particle’s mass, charge, and velocity are denoted by $m$, $e$, and $\vec{v}$ respectively. The Lorentz force law, together with measurements of the acceleration components of test particles, serves to define the components of the electric and magnetic fields. Once these field components are determined from the acceleration of test particles, they can be used to predict the accelerations of other particles.
This solution to the dynamics problem in the lab frame does not rely on Lorentz transformations. Conventional particle tracking treats space-time in a non-relativistic manner, as a (3+1) manifold. In this approach, the only modification to classical mechanics is the introduction of relativistic mass, while time remains distinct from space. There is no mixing of spatial and temporal coordinates  [^7] Clarification of the true content of the non-covariant theory can be found in various advanced textbooks. To quote e.g. Ferrarese and Bini [12]: ”… within a single inertial frame, the time is an absolute quantity in special relativity also. As a consequence, if no more than one frame is involved, one would not expect differences between classical and relativistic kinematics. But in the relativistic context, there are differences in the transformation laws of the various relative quantities (of kinematics or dynamics), when passing from one reference frame to another.” These authors emphasize the unique role of a ”single inertial frame.” This term suggests that a defining characteristic of non-covariant theory is the absence of relativistic kinematics in describing particle motion. A similar viewpoint is expressed by Friedman [15]: ”Within any single inertial frame, things look precisely the same as in Newtonian kinematics: there is an enduring Euclidean three-space, a global time $t$, and the law of motion.”.
Many fascinating phenomena arise when charges move under the influence of electromagnetic fields, often in highly complex situations. However, here we focus on a simpler case: the accelerated motion of particles in a constant magnetic field. In a non-covariant treatment, a magnetic field can change only the direction of a particle’s motion, not its speed (or mass). The study of relativistic particle motion in a constant magnetic field—common in accelerator physics—appears similar to its non-relativistic counterpart in Newtonian dynamics and kinematics. The trajectory of a particle, $\vec{x}(t)$, derived from the corrected form of Newton’s second law, does not account for relativistic kinematic effects, as it relies on the Galilean vector addition of velocities and disregards the relativity of simultaneity.
Now, let us address the important problem of velocity addition in relativity. It is evident that the velocity of light transforms just as a particle’s velocity does. We cannot use one set of kinematics for a particle’s velocity and another for the velocity of light. Consider a scenario where a light source in the lab frame is accelerated from rest to a velocity $u$. Suppose that, in the Lorentz inertial frame where the light source is at rest, an observer sees light propagating with velocity $v=c$. How will this appear to an observer in the lab frame? According to the relativistic velocity addition formula, the answer remains $v=c$. Maxwell’s equations retain their form under Lorentz transformations. However, these transformations lead to non-Galilean velocity transformation rules. Consequently, special relativity dictates that if Maxwell’s equations are to hold in the lab frame, particle trajectories must incorporate relativistic kinematic effects. That is, Maxwell’s equations can be consistently applied in the lab frame only if particle trajectories are described as the result of successive infinitesimal Lorentz transformations.
The absence of relativistic kinematic effects is a fundamental flaw in conventional (3+1) non-covariant particle tracking, making it clearly inconsistent with Maxwell’s electrodynamics. The relativity of simultaneity—i.e., the interdependence of spatial positions and time—is entirely absent from this framework. Consequently, the approach of coupling fields and particles within a ”single inertial frame” is not just flawed but fundamentally and profoundly incorrect.
2.9 Commonly Used Method of Coupling Fields and Particles
The electrodynamics problem is generally considered solvable within a ”single inertial frame” framework, without requiring Lorentz transformations. Standard derivations assume that Maxwell’s equations, along with a corrected version of Newton’s second law, can fully account for all experiments conducted within a single inertial frame, such as the laboratory reference frame.
Going to the electrodynamics problem, the differential form of Maxwell’s equations describing electromagnetic phenomena in the same inertial lab frame (in cgs units) is given by the following expressions:
$\vec{\nabla}\cdot\vec{E}=4\pi\rho~{},$ (5)
$\vec{\nabla}\cdot\vec{B}=0~{},$ (6)
×$\vec{\nabla}\times\vec{E}=-\frac{1}{c}\frac{\partial\vec{B}}{\partial t}~{},$ (7)
×$\vec{\nabla}\times\vec{B}=\frac{4\pi}{c}\vec{j}+\frac{1}{c}\frac{\partial\vec{E}}{\partial t}~{}.$ (8)
Here the charge density $\rho$ and current density $\vec{j}$ are written as
$\rho(\vec{x},t)=\sum_{n}e_{n}\delta(\vec{x}-\vec{x}_{n}(t))~{},$ (9)
$\vec{j}(\vec{x},t)=\sum_{n}e_{n}\vec{v}_{n}(t)\delta(\vec{x}-\vec{x}_{n}(t))~{},$ (10)
where $\delta(\vec{x}-\vec{x}_{n}(t))$ is three-dimensional delta function, $m_{n},e_{n},\vec{x}_{n}(t)$, and $\vec{v}_{n}=d\vec{x}_{n}(t)/dt$ denote the mass, charge, position, and the velocity of the $n$th particle, respectively. To evaluate radiation fields arising from external sources Eq. (8) we need the velocity $\vec{v}_{n}$ and the position $\vec{x}_{n}$ as a function of lab frame time $t$. It is generally accepted by the physics community that the equation of motion, which describes how coordinates of the particle carrying the charge change with time $t$, is described by corrected Newton’s second law Eq. (4).
This coupling of Maxwell’s equations with the corrected Newtonian equation is widely recognized as a valuable approach in accelerator physics, particularly in analytical and numerical studies of radiation properties. This method of treating relativistic dynamics and electrodynamics often leads accelerator physicists to assume that designing particle accelerators is possible without a deep understanding of the theory of relativity.
However, a common misconception in accelerator physics arises from the distinction between the trajectories $\vec{x}(t)$ and $\vec{x}_{cov}(t)$. To examine this issue from the perspective of electrodynamics of relativistically moving charges, consider the evaluation of fields generated by external sources. This requires knowledge of their velocity and position as functions of the lab-frame time $t$. Now, suppose one aims to compute the properties of radiation. Given our previous discussion, an important question emerges: Should one solve Maxwell’s equations in the lab frame using the current and charge density associated with a particle moving along the non-covariant trajectory $\vec{x}(t)$? We argue that the answer is no.
The conventional approach—solving Maxwell’s equations in the lab frame with currents and charge densities based on non-covariant trajectories—is widely regarded as relativistically correct. However, this method contradicts the fundamental principles of special relativity. Remarkably, this critical issue has largely gone unnoticed in the physics community. The only correct way to ensure consistency between Maxwell’s equations and particle trajectories in the lab frame is to solve the equations of motion in fully covariant form.
3 Space-Time and Its Coordinatization
3.1 Choosing a Coordinate System in an Inertial Frame
Let us explore an ”operational interpretation” of Lorentz and absolute time coordinatizations. Every physical phenomenon occurs in both space and time, and a concrete means of representing these dimensions is a frame of reference. A single space-time continuum can be described using different coordinate-time grids, corresponding to various frames of reference. Even the simplest space-time coordinate systems require careful definition.
Clocks play a crucial role in tracking a particle’s motion through a coordinate-time grid. The general approach to determining a particle’s motion is as follows: at any instant, a particle possesses a well-defined velocity, $\vec{v}$, as measured in a laboratory frame of reference. But how is this velocity determined? It is found once coordinates are assigned in the lab frame and then measured over appropriate time intervals along the particle’s trajectory. However, measuring a time interval between events occurring at different spatial points poses a fundamental challenge. To do so—and thereby measure a particle’s velocity within a single inertial lab frame—it is first necessary to synchronize distant clocks. The concept of synchronization is central to understanding special relativity.
There are multiple methods to synchronize distant clocks [^8] According to the thesis of the conventionality of simultaneity [13, 14, 15, 16, 17], simultaneity between distant events is a matter of convention, as it can be legitimately defined in different ways within any given inertial reference frame. Moeller [13], for example, states: ”All methods for the regulation of clocks meet with the same fundamental difficulty. The concept of simultaneity between two events in different places has no exact objective meaning at all since we cannot specify any experimental method by which this simultaneity could be ascertained. The same is therefore true also for the concept of velocity.”. The choice of a synchronization convention is ultimately a matter of selecting a coordinate system within an inertial frame of reference in Minkowski space-time.
The space-time continuum can be described using arbitrary coordinates. However, altering these coordinates does not change the underlying geometry of four-dimensional space-time. In special relativity, we are therefore unrestricted in our choice of a coordinate system. By leveraging the geometric structure of Minkowski space-time, one can define a class of inertial frames and adopt a Lorentz frame with orthonormal basis vectors. Within this chosen Lorentz frame, Einstein’s synchronization procedure—which relies on the constancy of the speed of light in all inertial frames—is applied, along with Cartesian spatial coordinates.
3.2 The Inertial Frame Where a Light Source is at Rest
Let us provide an operational interpretation of Lorentz coordinatization. The fundamental laws of electrodynamics are governed by Maxwell’s equations, which state that light propagates at a constant velocity $c$ in all directions. This isotropy arises naturally from the structure of Maxwell’s theory, which lacks any intrinsic directional dependence. It is often asserted that Maxwell’s equations, in their original form, are strictly valid within inertial frames. However, these equations can only be expressed in coordinate form once a space-time coordinate system has been properly defined.
Assigning Lorentz coordinates to the laboratory frame when a light source is in motion presents a complex challenge. To approach this problem systematically, we begin with the simpler case: assigning space-time coordinates to an inertial frame in which the light source remains at rest.
Our goal is to provide a practical and operational procedure for this assignment. The most natural synchronization method involves first bringing all ideal clocks to the same spatial location, where they can be synchronized. Once synchronized, the clocks are then slowly transported to their designated positions—a process known as slow clock transport  [^9] Eddington, in a text first published in 1923, was apparently the first to describe a synchronization procedure based on the slow transport of clocks [18]. Further details can be found in the review [8] (see also [19])..
Maxwell’s equations in their usual form are valid in any inertial frame where the sources are at rest, and the time coordinate is assigned using the slow clock transport procedure. The same principles apply when charged particles move in a non-relativistic manner. In particular, oscillating charged particles emit radiation. In the non-relativistic case, where the velocities of the oscillating charges are much smaller than $c$, the emitted radiation is predominantly dipole in nature and is well-described by Maxwell’s equations in their standard form.
Let us now examine in more detail how the dipole radiation term arises. The retardation time in the integrands of the radiation field amplitude expression can be neglected when the charge’s trajectory changes insignificantly during this time. The conditions for this approximation can be determined as follows. Denote by $a$ the characteristic size of the system. The retardation time is then of the order $a/c$. To ensure that the charge distribution remains nearly unchanged during this period, it is required that $a\ll\lambda$, where $\lambda$ is the radiation wavelength. This condition implies that the system’s dimensions must be small relative to the radiation wavelength. Expressing this constraint differently, we obtain $v\ll c$, where $v$ represents the characteristic velocity of the charges. Since dipole radiation theory accounts only for the dipole component of the emitted radiation, it disregards detailed information about the electron trajectory. Consequently, it is not surprising that the fields predicted by dipole radiation theory closely resemble those obtained from instantaneous approximations.
The theory of relativity provides an alternative method for clock synchronization, based on the fundamental principle that the speed of light remains constant in all inertial frames. While this is often regarded as a postulate, as we have seen, it is actually a matter of convention. The synchronization process that follows adheres to the standard Einstein synchronization procedure. Consider a dipole radiation source. When the dipole source is at rest, the governing field equations are given by Maxwell’s equations. In dipole radiation theory, we introduce a small expansion parameter, $v/c\ll 1$, and neglect terms of order $v/c$. This effectively means that the theory employs a zero-order nonrelativistic approximation. Einstein synchronization is defined using light signals emitted by the dipole source at rest, under the assumption that light propagates isotropically with velocity $c$ in all directions. Applying the Einstein synchronization procedure in the rest frame of the dipole source allows us to establish a Lorentz coordinate system.
Slow transport synchronization is equivalent to Einstein’s synchronization in an inertial system where the dipole light source is at rest  [^10] Many textbook authors still attribute a reality status to the one-way speed of light. To quote Hrasko [20]: ”It is sometimes claimed that Einstein’s synchronization of distant clocks $A$ and $B$ is circular. … This argument, however, is fallacious. It is true that a one-way measurement of light velocity can be performed only if clocks at the endpoints are correctly synchronized. However, since they need not show the correct coordinate time, they can be synchronized without light signals by transporting them symmetrically from a common site. … As we see, the described thought experiment is capable of proving the constancy of the speed of light if it is true or disproving it if it is false. It therefore provides a solid logical foundation for Einstein’s synchronization prescription.” This reasoning, however, is incorrect. Slow clock transport synchronization is equivalent to Einstein’s synchronization only in an inertial system where the light source is at rest.. In other words, suppose we have two sets of synchronized clocks spaced along the $x$-axis. Suppose that one set of clocks is synchronized by using the slow clock transport procedure and the other by light signals. If we would ride together with any clock in either set, we could see that it has the same time as the adjacent clocks, with which its reading is compared. This is because in our case of interest when the light source is at rest, field equations are the usual Maxwell’s equations and Einstein synchronization is defined in terms of light signals emitted by a source at rest assuming that light propagates with the same velocity $c$ in all directions. Applying either synchronization method in the rest frame establishes a Lorentz coordinate system. In this system, the metric of the light source takes the Minkowski form. We now turn to the topic of electromagnetic waves. In free space, the electric field $\vec{E}$ of an electromagnetic wave emitted by a stationary dipole source satisfies the wave equation: $\Box^{2}\vec{E}=\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial(ct)^{2}=0$. This result is closely related to the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$.
3.3 Motion of a Light Source Relative to an Inertial Frame
We now consider the case in which a light source, initially at rest in the lab frame, is accelerated to a velocity $v$ along the $x$-axis. In other words, we examine an active (physical) boost. When we say that the source undergoes acceleration while the inertial observer (equipped with measuring devices) does not, we mean acceleration relative to the fixed stars. Any acceleration relative to the fixed stars—that is, any active boost of velocity—has an absolute meaning.
In the case of an active boost, we analyze the motion of the same physical system over time, as observed from a single reference frame. A key question is whether the lab’s clock synchronization method depends on the motion of the light source. The answer to this is a matter of convention. The simplest synchronization method involves maintaining the same set of uniformly synchronized clocks used when the light source was at rest—essentially preserving clock transport synchronization (or Einstein synchronization, which is defined using light signals from a stationary dipole source). This choice is the most practical from a laboratory perspective, as it preserves simultaneity and follows the absolute time (or absolute simultaneity) convention.
Absolute simultaneity can be incorporated into special relativity without altering either the logical structure or the (convention-independent) predictions of the theory. We begin with the metric as the true measure of space-time intervals for an inertial observer with coordinates $(t,x,y,z)$. Applying a Galilean boost, we transform the spatial coordinate as $x\to x-vt$ while keeping time unchanged. Substituting this into the Minkowski metric, $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ yields
$ds^{2}=c^{2}(1-v^{2}/c^{2})dt^{2}+2vdxdt-dx^{2}-dy^{2}-dz^{2}~{}.$ (11)
This metric describes the electrodynamics of the moving light source from the viewpoint of the inertial observer measurements. From the structure of the equation, it is evident that the measurement of the electromagnetic field configuration—expressed in terms of the coordinates $x$ and $t$ of a measuring device at rest in the inertial frame—yields the same result as that obtained at the spatial position $x-vt$ at time $t=0$.
To our knowledge, the metric in Eq. 6 is presented here for the first time. This result has not been previously reported in the literature. The agreement between this theoretical framework and empirical observations lends sup- port to our revised formulation of special relativity incorporating absolute simultaneity (see Section 4.14 for further details). By contrast, existing literature often assumes—incorrectly—that an inertial observer and an accelerated light source share the same Minkowski metric, treating this assumption as self-evident.
To understand the physical interpretation of Eq. 11, it is important to recognize that space-time points can be labeled in infinitely many ways using different coordinate systems. These systems are related through coordinate transformations. A new frame of reference can always be introduced by relabeling coordinates and describing physical phenomena using the new labels—a process known as a passive (coordinate) transformation. For instance, in a laboratory frame, one can define a comoving coordinate system to analyze radiation from a moving source in terms of the new (comoving) coordinates. In this system, fields are expressed as functions of the independent variables $x^{\prime},y^{\prime},z^{\prime},t^{\prime}$, which are related to the original coordinates $x,y,z,t$ via a passive Galilean transformation. Specifically, after the passive transformation, the Cartesian coordinates of the source transform as $x^{\prime}=x-vt,~{}y^{\prime}=y,~{}z^{\prime}=z$. This transformation completes with the invariance of simultaneity, $t^{\prime}=t$. The transformation of time and spatial coordinates of any event has the form of a Galilean transformation. One must take into account that $(t,x,y,z)$ are coordinates of the rule clock structure at rest in the inertial frame.
The equivalence of the active and passive perspectives arises because shifting a system in one direction is equivalent to shifting the coordinate system in the opposite direction by the same amount. However, there is no a priori guarantee that the physical world must behave this way. For example, Newtonian mechanics is not invariant under passive rotations, as demonstrated by Newton’s bucket experiment. The empirical principle states that within a single inertial frame, every passive Galilean boost is postulated to correspond to an active Galilean boost.
According to this equivalence of passive and active boosts, Maxwell’s equations remain valid in the comoving frame. In this frame, the fields are expressed as functions of the independent variables $x^{\prime},y^{\prime},z^{\prime}$, and $t^{\prime}$. The electric field $\vec{E}^{\prime}$ of an electromagnetic wave satisfies the equation $\Box^{\prime 2}\vec{E}^{\prime}=\nabla^{\prime 2}\vec{E}^{\prime}-\partial^{2}\vec{E}^{\prime}/\partial(ct^{\prime})^{2}=0$. However, the variables $x^{\prime},y^{\prime},z^{\prime},t^{\prime}$ can be expressed in terms of the independent variables $x,y,z,t$ using a Galilean transformation, so that fields can be written in terms of $x,y,z,t$. From the Galilean transformation $x^{\prime}=x-vt,~{}y^{\prime}=y,~{}z^{\prime}=z,~{}t^{\prime}=t$, after partial differentiation, one obtains $\partial/{\partial t}=\partial/{\partial t^{\prime}}-v\partial/{\partial x^{\prime}}$, $\partial/{\partial x}=\partial/{\partial x^{\prime}}$. Hence the wave equation transforms into
$\Box^{2}\vec{E}=\left(1-\frac{v^{2}}{c^{2}}\right)\frac{\partial^{2}\vec{E}}{\partial x^{2}}-2\left(\frac{v}{c}\right)\frac{\partial^{2}\vec{E}}{\partial ct\partial x}+\frac{\partial^{2}\vec{E}}{\partial y^{2}}+\frac{\partial^{2}\vec{E}}{\partial z^{2}}-\frac{1}{c^{2}}\frac{\partial^{2}\vec{E}}{\partial t^{2}}=0~{},$ (12)
This last result is closely related to the metric equation Eq. 11.
The solution of this equation $F[x-(c+v)t]+G[x+(-c+v)t]$ is the sum of two arbitrary functions, one of argument $x-(c+v)t$ and the other of argument $x+(-c+v)t$. Here we obtained the solution for waves that move in the $x$ direction by supposing that the field does not depend on $y$ and $z$. The first term represents a wave traveling forward in the positive $x$ direction, and the second term represents a wave traveling backward in the negative direction.
We conclude that the speed of light emitted by a moving source, as measured in the laboratory frame $(t,x,y,z)$, depends on the relative velocity $v$ between the source and the observer. In other words, in this scenario, the speed of light follows the Galilean law of velocity addition. Specifically, the coordinate velocity of light along the $x$-axis is given by $dx/dt=c+v$ in the positive direction and $dx/dt=-c+v$ in the negative direction. This apparent deviation from the electrodynamics constant $c$ arises because the clocks are synchronized according to an absolute time convention. Once such a convention is chosen, it must be consistently applied to both mechanics and electrodynamics. Under this absolute time coordinatization, the velocity of light transforms analogously to that of a particle, as expected.
3.4 Discussion
So far, we have examined the Galilean transformation of the electrodynamics equations. Now, we highlight a crucial distinction between fundamental and phenomenological theories. For instance, Newton’s equation in a co-moving frame can be regarded as a phenomenological law. It does not provide a microscopic interpretation of a particle’s inertial mass; instead, rest mass is introduced in an ad hoc manner. This suggests that the particle may possess internal variables that remain unknown. Any phenomenological law valid in the Lorentz rest frame can be embedded in four-dimensional space-time only through Lorentz coordinatization. Now, let us turn our attention to electrodynamics. It is essential to emphasize that electrodynamic laws do not rely on any hidden internal variables or underlying mechanisms. Electromagnetic fields are fundamental, and electrodynamics fully complies with the principles of relativity. As a result, any synchronization scheme can be used to describe electromagnetic field dynamics.
The flexibility in describing physical systems allows us to choose the most suitable representation for each problem. A new frame of reference can always be introduced by relabeling coordinates, enabling the discussion of physical phenomena through passive transformations. In contrast, an active transformation involves the actual motion of a physical system—where its characteristics change due to internal or external interactions. Here, we are concerned with the evolution of the same physical system over time, analyzed from the perspective of a single reference frame.
Let us now explore the correspondence between passive and active transformations, drawing a comparison between special relativity and Newtonian mechanics. In Newton’s non-relativistic theory, the formalism is grounded in an absolute space structure. However, there is relativity concerning uniform motion, meaning we cannot determine a particle’s absolute velocity. Passive Galilean boosts are dynamical symmetries of Newtonian mechanics. The equivalence between active and passive Galilean boosts arises from the fact that moving a system in one direction is equivalent to shifting the coordinate system in the opposite direction by the same amount, as illustrated by Galileo’s famous ship experiment. This symmetry, known as the Galilean group, exists because the equations of motion in Newtonian mechanics do not explicitly depend on velocity. The internal dynamics of the system remain unchanged under a Galilean boost. However, Newtonian mechanics is not invariant under passive rotations. The form of the equations of motion changes due to the introduction of pseudo-forces. As a general discussion of the correspondence between passive and active transformations, we refer the reader to the paper [21]. Newton’s bucket argument contrasts two situations: one where a vessel of water is at rest, and another where the vessel is gradually rotated to a state of uniform angular velocity. The argument against the equivalence of active and passive rotations is based on the observation that, while there is no relative motion between the water and the bucket in both cases, the shape of the water differs. There is no relativity of accelerations in Newtonian mechanics, and we can determine the absolute acceleration of a particle in an inertial frame. Absolute acceleration refers to acceleration relative to the fixed stars.
Let us now examine the correspondence between active and passive transformations in special relativity. Within a single inertial frame, kinematical symmetry corresponds to dynamical symmetry, and the symmetry principle (i.e., the equivalence of the active and passive Galilean boosts) holds. It is important to note that a passive Galilean boost, when considered within a single inertial frame, is merely a different parametrization of the observations made by the inertial observer. Under the passive transformation ($x\to x-vt$, $t\to t$), the motion of fixed stars relative to the observer and their devices remains unchanged. Note that $(t,x,y,z)$ represent the coordinates of the clock structure at rest in the inertial frame. In Chapter 5, we will address the problem of light aberration in accelerated systems using the framework of special relativity. This issue can be adequately tackled only by adopting an approach that employs absolute time coordinatization. To express the laws of electrodynamics in an accelerated frame, we define the metric of the accelerated frame using the inverse Galilean boost. It is crucial to highlight that, when applying an inverse Galilean boost in velocity, we are accounting for the effects of interaction in terms of the acceleration of motion relative to the fixed stars.
This implies that the passive Galilean transformation within a single inertial frame is fundamentally distinct from the (passive) inverse Galilean transformation, $x\to x_{n}+vt$, $t\to t_{n}$. This transformation describes the acceleration of an observer (with their measuring instruments) relative to the fixed stars. It’s important to note that $(t_{n},x_{n},y_{n},z_{n})$ are the coordinates of the rule-clock structure at rest in the accelerated frame. The transition of the coordinates of the light source does not signify a mere change in representation but reflects a genuine change in the absolute motion of the light source (and rule-clock structure) with respect to the fixed stars. It is important to recognize that in the actual process of transforming to an accelerated frame (relative to the fixed stars), the observer will experience inertial forces, which are not accounted for in standard treatments. We argue that the reciprocity of the inertial frames, which is considered in textbooks as relativistically correct, is at odds with special relativity. The arguments concerning the relativity of motion in our case of interest cannot be applied, since the inertial and accelerated reference systems are not equitable.
According to special relativity, there is a significant difference between an accelerated inertial frame (relative to the fixed stars) and an inertial frame that has not undergone any acceleration, with this distinction being intimately tied to inertial forces. While the duration of the acceleration period has a negligible effect on anisotropy in the accelerated frame, the overall impact of inertial forces fundamentally determines the problem. A useful way to conceptualize the asymmetry between inertial and accelerated frames is to view it as the result of pseudo-gravity experienced by the accelerated observer. A uniformly accelerating frame can be treated as an inertial frame supplemented by a uniform pseudo-gravitational field. This concept is at the heart of the equivalence principle, which allows us to apply it to solve non-inertial kinematics problems using dynamics methods.
There is an alternative approach to solving the problem. The formalism of relativistic physics assumes an absolute space-time structure, where there is no relativity of uniform motion, allowing us to determine the system’s absolute velocity in the initial inertial frame. From a space-time geometric perspective, the solution involves recognizing that the metric tensor must be a continuous quantity. When an accelerated system transitions to constant velocity, we can smoothly adjust the metric tensor. The (non-diagonal) Langevin metric arises by matching the metric tensors of the accelerated frame and the initial inertial frame. This perspective may be unsettling for those who have learned relativity from standard textbooks. According to conventional teachings, the problem is entirely symmetric, and both frames are considered equivalent. The widespread belief is that the coordinates of these frames are related by Lorentz transformations, and the metric in both frames is diagonal. However, upon diagonalizing the metric, we find that in the accelerated frame, the metric tensor must abruptly shift from the Langevin form to the Minkowski form. This reveals that, contrary to the textbook understanding, Lorentz transformations do not preserve the continuity of the metric tensor ( see Chapter 5 for further details).
3.5 A Clock Resynchronization Procedure
After the Galilean boost in the case of the absolute time coordinatization we can see that the homogeneous wave equation for the field in the lab frame has nearly but not quite the usual, standard form that takes when there is no acceleration. The main difference consists in the crossed term $\partial^{2}/\partial t\partial x$, which complicates the solution of the equation. To get around this difficulty, we observe that simplification is always possible. The trick needed here is to further make a change of the time variable. Suppose that we describe the effect of radiation by working only up to the first order in $v/c$. In this approximation, it is always possible to introduce a time transformation $t\to t+xv/c^{2}$. After this time shift we obtain the usual wave equation.
Let us now consider the case of an arbitrary velocity along the $x$-axis. We have, then, a general method for finding a solution of the electrodynamics problem in the case of absolute time coordinatization. The new independent variables $(x_{L},t_{L})$ can be expressed in terms of the old independent variables $(x,t)$:
$ct_{L}=t\sqrt{1-v^{2}/c^{2}}+vx/[c^{2}\sqrt{1-v^{2}/c^{2}}]~{},~{}x_{L}=x/\sqrt{1-v^{2}/c^{2}}~{},$ (13)
Since the change variables completed by the Galilean transformation is mathematically equivalent to the Lorentz transformation, it follows that transforming to new variables $x_{L},t_{L}$ leads to the usual Maxwell’s equations. In particular, the wave equation Eq. (12) transforms into
$\Box_{L}^{2}E=\nabla_{L}^{2}E-\partial^{2}E/\partial(ct_{L})^{2}=0~{}.$ (14)
In the new variables, the velocity of light emitted by a moving source is constant in all directions, and equal to the electrodynamics constant $c$.
The combined effect of the Galilean boost and variable changes results in the Lorentz transformation under absolute time coordinatization in the lab frame. However, in this context, the transformation should be regarded merely as a mathematical tool that simplifies solving the electrodynamics problem while maintaining absolute time synchronization.
This raises an interesting question: Is it necessary to transform the solution back into the original variables? We argue that the variable changes introduced above have no intrinsic significance; their meaning is purely conventional. Notably, the time shift $t+xv/c^{2}$ is directly related to the issue of clock synchronization. Moreover, the final rescaling of time and spatial coordinates remains physically unobservable. Crucially, the convention-independent results remain unchanged in the new variables. Consequently, there is no need to revert the solution to its original form.
Consider two light sources labeled ”1” and ”2”. Suppose that in the lab frame the velocities of ”1” and ”2” are $\vec{v}_{1}$, $\vec{v}_{2}$ and $\vec{v}_{1}\neq\vec{v}_{2}$. The question now arises of how to assign a time coordinate to the lab reference frame. We have two choices: an absolute time coordinate or a Lorentz time coordinate. The most natural choice, in terms of connecting to laboratory reality, is absolute time synchronization. In this framework, simultaneity is absolute, meaning that a single set of synchronized clocks is sufficient for both sources in the lab frame. However, Maxwell’s equations are not form-invariant under Galilean transformations, meaning their mathematical form differs in the lab frame. Using absolute time synchronization leads to more complex field equations, which also vary for each source. To introduce Lorentz coordinates, we must adopt a different approach. The only viable method in this scenario is to assign an individual coordinate system—i.e., an independent set of synchronized clocks—to each source.
Now, consider a second secondary source, such as a tangentially moving mirror, interacting with a light beam emitted by
a stationary source. A key challenge in special relativity is that both sources cannot share a common Lorentz frame within a single inertial system. The electrodynamic equations must be identical for both the incident and reflected waves. Under Einstein’s synchronization, the time for the moving mirror can be determined by introducing a time shift of $xv/c^{2}$ to the first order of approximation. In this synchronization scheme, reflection is described using Maxwell’s equations. This time shift effectively rotates the radiation wavefront by an angle $v/c$ for both sources. Now, consider the electrodynamics of the stationary source. In absolute-time coordinatization, an observer describes the light beam from the stationary emitter using the diagonal wave equation. However, the time shift introduces a new crossed term in the wave equation, leading to a non-diagonal form. Further details on this issue are provided in the next chapter (see Section 4.6 for further details).
3.6 A Moving Light Source: Peculiarities of Collinear Geometry
So far, we have considered both covariant and non-covariant approaches to solving the problem of radiation emitted by a moving source in an inertial frame. Now, let us examine the acceleration of a dipole source in the laboratory inertial frame, where it accelerates to a velocity $v$ along the $x$-axis. This raises the question of how to assign synchronization in the lab frame after the source has undergone acceleration. Before acceleration, we adopted a Lorentz coordinate system. However, if synchronization in the lab frame remains unchanged after the acceleration, the electrodynamics of moving charges becomes significantly more complex. In this case, the transformation of time and spatial coordinates does not retain the standard Lorentz form. Maxwell’s equations remain valid in the lab inertial frame only when Lorentz coordinates are properly assigned. To preserve Lorentz coordinatization after acceleration, it is necessary to redefine the space-time coordinate system by introducing new variables $(x_{L},y_{L},z_{L},t_{L})$, as given in Eq. (13). In this new coordinate system, the velocity of light emitted by the moving source remains equal to the electrodynamic constant $c$ in all directions.
The aberration of light and the Doppler effect serve as practical examples to illustrate the distinction between covariant and non-covariant approaches. The most effective way to calculate radiation is by applying the Lorentz transformation formulas between inertial reference frames. Specifically, consider a radiation source at rest in the frame $S^{\prime}$, emitting dipole radiation with frequency $\omega_{0}$. In the laboratory frame $S$, where the source moves with velocity $v$ and the radiation propagates along the direction of motion, we observe the so-called radial Doppler effect. This effect is governed by the well-known formula: $\omega=\omega_{0}\sqrt{1-v^{2}/c^{2}}/(1-v/c)$. The effect of the factor $\sqrt{1-v^{2}/c^{2}}$ can be summarized in the following statement: on the moving object, time is flowing slower than expected (time dilation).
It is worth noting that using the conventional coupling of Maxwell’s equations and the corrected Newton’s equation to calculate radiation from a moving source does not necessarily lead to errors. For rectilinear motion of both the source and the emitted light beam, non-covariant and covariant approaches yield identical trajectories, ensuring compatibility between Maxwell’s equations and conventional particle tracking. However, this method was incorrectly extended to non-collinear geometries.
This raises a reasonable question: why does the same method produce incorrect results when the source moves perpendicular to the direction of its emitted radiation? The key distinction in collinear geometry—where the source moves along the same line as the radiated beam—is that the velocity is perpendicular to the plane of the radiation wavefront (i.e., the plane of simultaneity). Consequently, for collinear motion, the plane of simultaneity in absolute time coordinatization retains the same orientation in Lorentz coordinatization.
In the transverse case, the motion is perpendicular to this direction, and the source undergoes acceleration along the radiation wavefront (i.e., the initial plane of phase simultaneity). Since the orientation of the radiation wavefront is no longer absolute, it depends on the chosen clock synchronization convention.
To analyze this, let us first consider a Lorentz transformation. Specifically, we use a Lorentz boost to describe the uniform translational motion of the light source in the lab frame. It is evident that wavefront phases, which are simultaneous in $S^{\prime}$ but occur at different $x^{\prime}$-locations, are not simultaneous in $S$. When applying a Lorentz boost, a time transformation is introduced: $t^{\prime}=t-xv/c^{2}$. The effect of this transformation is equivalent to a rotation of the radiation wavefront in the lab frame by an angle of approximately $v/c$ in the first-order approximation.
According to Maxwell’s electrodynamics, radiation is always emitted in a direction normal to the radiation wavefront. Consequently, the radiated light beam propagates at an angle $v/c$, leading to the phenomenon known as the aberration of light. This remarkable effect will be the focus of our discussion in the next chapter. In the conventional (3+1) approach, simultaneity is considered absolute, and there is no mixing of spatial coordinates and time when sources change their velocities in an inertial frame. Thus, it appears that the conventional (3+1) approach cannot account for the geometric phenomenon of aberration. By applying Maxwell’s equations alongside an absolute time transformation, neither the wavefront plane nor the direction of energy transport undergoes deflection.
Let us now return to the conventional coupling between Maxwell’s equations and particle trajectories within the ”single lab frame.” According to the theory of special relativity, the standard description of optical phenomena related to the motion of a light source remains valid when radiation propagates along the direction of velocity. For instance, in a collinear geometry, the conventional coupling of fields and particles can provide a useful framework for analyzing the Doppler effect.
[…]
The relativistic motion of an electron is described by $d\vec{p}/dt=e\vec{v}\times\vec{H}/c$ with $\vec{p}=m\vec{v}/\sqrt{1-v^{2}/c^{2}}$. This equation can be written as $d\vec{v}/dt=\vec{v}\times\vec{\omega}_{L}$ where $\vec{\omega}_{L}=e\vec{H}\sqrt{1-v^{2}/c^{2}}/(mc)=\vec{\omega}_{L0}\sqrt{1-v^{2}/c^{2}}$ is the relativistic Larmor frequency. When the source is accelerated, the speed of electrons is increased, and therefore the mass is also increased and the electron is heavier. The emissivity presents a spectrum in which the frequency is given by $\omega=\omega_{L0}\sqrt{1-v^{2}/c^{2}}/(1-v/c)$. The frequency is multiplied by the relativistic Larmor frequency accounting at the same time through the denominator for the Doppler effect caused by the motion parallel to the magnetic field.
We can thus conclude that the frequency of an electron’s oscillations decreases when the source moves with velocity $v$. At first glance, the mechanism of the source described may not seem to involve relativistic time dilation. However, it is implicitly present in the assumption that the mass of a moving object corresponds to its relativistic mass, given by $m_{r}=m/\sqrt{1-v^{2}/c^{2}}$.
According to the covariant approach, various relativistic kinematic effects relevant to the dipole radiation setup emerge in successive orders of approximation. The first-order kinematic term $v/c$ plays a crucial role only in the description of dipole radiation in the perpendicular geometry. In the case of collinear geometry, however, the motion of the dipole source affects only the second-order kinematic terms $(v/c)^{2}$, in accordance with the theory of relativity.
In the non-covariant approach, solving the dynamics problem in the lab frame does not involve Lorentz transformations. This means that, within the lab frame, particle motion appears exactly as predicted by classical mechanics, which assumes an absolute time framework. This absolute-time approach is well-suited for explaining experimental outcomes in collinear geometry. However, we argue that this method of solving Maxwell’s equations in the lab frame is not applicable in the transverse case.
The use of electrodynamics in absolute-time coordinatization becomes essential when adopting a non-covariant (3+1)-dimensional approach (i.e., ”old kinematics”) to relativistic particle dynamics. Conventional particle tracking relies on the lab time $t$ as the independent variable, thereby excluding relativistic kinematic effects from the description. This approach is based on the assumption of absolute time synchronization in the lab frame. If one intends to use corrected Newtonian equations—where time serves as the independent variable—the electrodynamic equations must be formulated in a non-covariant manner (i.e., within the absolute-time coordinatization). Only in this framework does the coupling between electrodynamic equations and particle trajectories remain consistent within a single reference frame.
4 Aberration of Light in an Inertial Frame of Reference
The effect of light aberration in an inertial frame of reference is typically understood as a change in the direction of light propagation attributed to moving light sources. However, even at the first order in $v/c$, describing this phenomenon rigorously is far from straightforward. Many textbooks contain incorrect results, reflecting the complexity of the topic. To minimize mathematical complexity, we will restrict our discussion to first-order effects. This chapter focuses on the physical influence of optical instruments on the measurement of light aberration. First, we will demonstrate that when a finite-aperture mirror moves parallel to itself and a plane wave of light falls normally onto it, the reflected light exhibits aberration—manifested as a deviation in energy transport. The next problem, of greater practical significance, involves the complementary case of transmission through an aperture in a moving opaque screen, as described by Babinet’s principle. A particularly relevant scenario is the transmission of light through the moving end of a telescope barrel, which we will examine in detail.
4.1 The ”Plane Wave” Emitter
We will analyze the effect in a simple case to build an intuitive understanding. The explanation of aberration is based on modeling a single plane-wave emitter. As a simple model of a plane-wave emitter, we use a two-dimensional array of identical coherent elementary sources (dipoles), uniformly distributed on a given ($x-y$) plane $P$. We take the elementary sources to start radiating waves simultaneously with respect to the lab reference frame where the plane $P$ is at rest. Therefore we have a plane full of sources, oscillating together, with their motion in the plane and all having the same amplitude and phase. Let us assume that the elementary sources are oscillating at frequency $\omega$. By reducing the spacing between adjacent sources to a value much smaller than the radiation wavelength $\lambda=2\pi c/\omega$, we can approximate this two-dimensional arrangement as an ideal plane-wave emitter.
The concept of an (infinitely extended) plane wave is widely used in physics due to its well-behaved analytical properties as a solution to Maxwell’s equations. However, it is not physically realizable, as such a wave would carry infinite energy. In any real-world scenario, one must consider a finite source aperture. It is implicitly assumed that the detector measuring the radiation direction is sensitive to energy propagation and that its aperture is sufficiently large compared to the radiation beam size. Ultimately, what is commonly referred to as aberration is, in reality, a deviation in the direction of energy transport.
4.2 A Moving Emitter: Single Frame Description
Let us consider the case of an emitter lying in the $(x-y)$ plane of an inertial laboratory frame, which is accelerated from rest to a velocity $v$ along the $x$-axis. Suppose an observer at rest in this inertial frame measures the direction of energy transport.
There are two possible approaches to analyzing the aberration of light emitted by a single moving source. The first is the covariant approach, commonly used in the literature, which explains the aberration effect by applying a Lorentz boost to determine how the direction of a light beam depends on the velocity of the source relative to the inertial frame. The second is a non-covariant approach, which employs a ”single-frame” description without explicit reference to Lorentz transformations.
Both approaches, whether formulated using Einstein’s synchronization convention or an absolute time synchronization scheme, yield the same result in the case of a single moving emitter. Thus, the choice between them is a matter of pragmatism rather than fundamental necessity.
In this chapter, we present both approaches, beginning with the non-covariant one. It is important to emphasize that the (3+1) non-covariant formulation in relativistic electrodynamics is entirely valid and provides a consistent description. The aberration of light can be treated within a single-inertial-frame framework, without invoking Lorentz transformations. In this context, time remains an absolute quantity in special relativity. But what does ”absolute time” mean? It signifies that simultaneity is preserved within the inertial frame, and there is no intermixing of spatial positions and time coordinates when sources change velocity within this frame. A distinctive feature of this non-covariant approach is the absence of relativistic kinematics in the description of light aberration. For a tangentially moving emitter (relative to its surface), the direction of energy transport of the emitted light undergoes a deviation. Within the ”single-frame” approach, this effect arises due to the Doppler effect, which induces angular frequency dispersion of the emitted light waves (see Fig. 1, left).
Figure 1: Illustration of a plane wave emitter, initially at rest in an inertial frame, accelerating to velocity $v$ along the $x$-axis. The effect of aberration is analyzed using a Galilean boost (left) and a Lorentz boost (right).
Let us examine the electrodynamics of a moving source. To explain the phenomenon of radiation in our case of interest, we apply an (active) Galilean boost to describe the uniform translational motion of the source in the inertial lab frame. Since Maxwell’s equations are not invariant under Galilean transformations, their form is not preserved. The additional terms introduced in the field equations as a result of the Galilean transformation lead to the prediction of the Doppler effect.
The current approach to the moving emitter problem employs the Fourier transform method. In the analysis of linear systems, it is often useful to break down a complex input into simpler components, determine the system’s response to each elementary function, and then superimpose these responses to obtain the total outcome. Fourier analysis serves as a fundamental tool for this decomposition  [^11] For a broader discussion on Fourier transform methods in spatial filtering theory or Abbe diffraction theory, we recommend referring to [22]..
[…]
An emitter with a finite aperture is a kind of active medium that breaks up the radiated beam into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which an active medium can be resolved. Let us assume that the current density of the elementary dipole sources varies according to the law $j_{pol}=g(K_{\perp})\sin{({K}_{\perp}x)}$. From this, we conclude that the active medium of the emmiter is sinusoidally space-modulated.
Starting from the Galilean transformation and applying partial differentiation, we obtain the wave equation Eq.(12). Our goal is to demonstrate how the additional terms introduced in the field equations due to the Galilean transformation lead to the prediction of the Doppler effect. Consider as a possible solution a radiated plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$. With a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ with the wavenumber vector $\vec{k}$ and the frequency $\omega$ equation Eq.(12) becomes: $(1-v^{2}/c^{2})k_{x}^{2}+2vk_{x}\omega/c^{2}+k_{z}^{2}-\omega^{2}/c^{2}=0$.
The wavevector $\vec{k}$ is determined by the initial conditions. Specifically, we set $k_{x}=K_{\perp}$, where $K_{\perp}$ corresponds to the wavenumber of the sinusoidally space-modulated dipole density, and $k_{z}^{2}=\omega_{i}^{2}/c^{2}$, with $\omega_{i}$ denoting the frequency of the emitted radiation prior to acceleration.
Given these initial conditions, we treat $\vec{k}$ as the independent variable and express the frequency as a function of $k_{x}$: $\omega=\omega_{i}+\Delta\omega(k_{x})$, where $\Delta\omega(k_{x})$ represents the frequency shift due to motion. Substituting this relation into the dispersion equation yields a condition linking the wavenumber and the frequency shift: $\Delta\omega=K_{\perp}v$.
[…]
An alternative physical perspective on the Doppler effect for a light beam emitted by a moving source leads to the same conclusion as the one presented above. The phase of the wave at the world point $(\vec{r},t)$ is invariant at a change of the reference system. This invariance holds regardless of the specific coordinate transformation used to describe the change in the reference system. Therefore, the phase $\vec{k}\cdot\vec{r}-\omega t$ must be invariant of the Galilean transformation. Consequently, $\vec{k}^{\prime}\cdot\vec{r}^{\prime}-\omega^{\prime}t^{\prime}=\vec{k}\cdot\vec{r}-\omega t$. Substituting the Galilean transformation formulae $x^{\prime}=x-vt,t^{\prime}=t$ into the phase equality formula we obtain $\Delta\omega=K_{\perp}v$. This frequency shift matches the result obtained directly from the dispersion relation, as expected.
Let us first remind the reader that the velocity of waves is typically defined in terms of the phase difference between oscillations observed at two different points in a free plane wave. This concept is primarily used to compute interference fringes, which visually highlight phase differences. In a plane wave, the phase velocity is given by the ratio $\omega/k$. In addition to phase velocity, another type of velocity—known as the group velocity or energy propagation velocity—can be defined by considering the propagation of a disturbance, such as a change in amplitude, superimposed on a wave train. A simple example of a wave group is formed by the superposition of two waves with frequencies and wave numbers $\omega_{1}=\omega+\Delta\omega,~{}k_{1}=k+\Delta k$ and $\omega_{2}=\omega-\Delta\omega,~{}k_{2}=k-\Delta k$. This combination represents a carrier wave with frequency $\omega$ and a modulation with frequency $\Delta\omega$. The resulting wave can be described as a succession of moving beats (or groups), where the carrier wave travels with velocity $\omega/k$, and the group velocity is given by $v_{g}=\Delta\omega/\Delta k$, which, in the limit, becomes $v_{g}=d\omega/dk$.
Many textbooks on electromagnetic theory discuss the phenomenon of light aberration in the context of plane waves. However, using a plane wave model to explain light aberration leads to an incorrect understanding. When an infinite sinusoidal wave propagates, it maintains a uniform average energy density throughout space. The question then arises: does this energy remain stationary, or does it propagate through space? It is impossible to determine this with certainty based on the plane wave model. All experimental methods used to measure the aberration of light rely on light signals, which measure not the phase velocity, but the signal velocity. In our case of interest, this signal velocity coincides with the group velocity.
In this example, plane waves with different wavenumber vectors propagate from the moving emitter, each with its own frequency. The relation $\Delta\omega/\Delta k_{x}=v$ holds for each scattered wave, regardless of the sign or magnitude of the radiated angle. Specifically, in this case, $\Delta\omega$ represents the Doppler shift, given by $\Delta\omega=\vec{K}_{\perp}\cdot\vec{v}$, and $\Delta k_{x}$ is simply the transverse component of the radiated wavenumber vector, $\Delta k_{x}=K_{\perp}$. These equations indicate that a light beam with a finite transverse size propagates along the $x$-direction with a group velocity $d\omega/dk_{x}=v$.
4.3 A Moving Emitter: Lorentz Transformations-Based Explanation
Another clear way to explain the aberration of light from a moving source is by using a Lorentz boost, which accounts for the uniform translational motion of the source in the laboratory frame. Maxwell’s equations are invariant under Lorentz transformations; however, the Lorentz boost introduces a time coordinate transformation of the form $t^{\prime}=t-xv/c^{2}$, which modifies the timing of events across space.
As a result, features such as the maximum of the electric field reach different positions in the lab frame at different times. For example, the field maximum at a position $x$ along the $x$-axis appears delayed compared to the maximum at $x=0$, with a time shift given by $\Delta t=xv/c^{2}$.
This spatially dependent time delay effectively causes a tilt in the phase front of the radiation. To first order in $v/c$, this corresponds to a rotation of the phase front by an angle approximately equal to $v/c$.
According to Maxwell’s electrodynamics, coherent radiation is always emitted perpendicular to the radiation phase front, as Maxwell’s equations exhibit no intrinsic anisotropy. Thus, when the uniform motion of the source is analyzed via Lorentz transformations, the aberration of light can be understood in the framework of relativistic kinematics. In this description, an additional phase chirp, $d\phi/dx=k_{x}=v\omega_{i}/c^{2}$, emerges, causing the array of identical elementary sources in the moving emitter to acquire different phases. Consequently, the plane wavefront undergoes a rotation under the Lorentz transformation, resulting in a radiated light beam propagating at an angle $v/c$. This yields the well-known phenomenon of aberration of light (Fig. 1, right).
We now consider the group velocity of the radiated beam. For a plane wave of the form $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$, the dispersion equation in Maxwell’s electrodynamics simplifies to $k_{z}^{2}+k_{x}^{2}-\omega^{2}/c^{2}=0$. Using the initial conditions and the Lorentz transformation, we obtain the following expression: $\omega=(\omega_{i}+vK_{\perp})$, $k_{z}=\omega_{i}/c$, $k_{x}=(v\omega_{i}/c^{2}+K_{\perp})$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated current density, $\omega_{i}$ is the frequency of the emitter radiation before the acceleration. Substituting these expressions into the dispersion equation, we verify that it is indeed satisfied, as expected.
As a consequence of the Doppler effect in Lorentz coordinatization, we observe an angular frequency dispersion in the light waves emitted from a moving source with a finite aperture. The Doppler shift, $\Delta\omega$, of a radiated light wave (in the first order approximation) is given by $\Delta\omega=\vec{K}_{\perp}\cdot\vec{v}$, where $K_{\perp}$ is the transverse component of the radiated wavenumber vector. This equation implies that a light beam with a finite transverse extent propagating along the $x$-direction has a group velocity given by $d\omega/dk_{x}=v$.
It is worth examining the significance of the fact that two different approaches—covariant and non-covariant—lead to the same group velocity. The key insight is that both methods accurately describe the same physical reality. Since the group velocity has an objective, convention-independent meaning, both approaches must necessarily yield the same physical results.
The distinction between absolute time coordinatization and Lorentz coordinatization is particularly intriguing. In Chapter 3, we explored how one can transform from absolute time coordinatization to Lorentz coordinatization. Specifically, manipulations involving the rule-clock structure in the lab frame can be understood as a simple change of the time variable, given by the transformation $t\to t+xv/c^{2}$. When combined with the Galilean transformation, this variable change ultimately leads to the Lorentz transformation in the context of absolute time coordinatization in an inertial frame. However, this change of variables has no intrinsic physical significance. Notably, there is a direct link between this time shift and the issue of clock synchronization. Importantly, the convention-independent results of calculations remain unchanged in the transformed variables. Consequently, there is no need to convert the results of electrodynamics problem-solving back into the original (3+1) variables.
From a pedagogical perspective, studying electrodynamics through variable transformations provides valuable insight. It is worth noting that the lack of a dynamical explanation for wavefront rotation in Lorentz coordinatization has been a point of concern for some physicists. However, as discussed in Chapter 3, a useful way to conceptualize wavefront rotation is to view it as a consequence of transitioning to a new time variable within the framework of Galilean (”single-frame”) electrodynamics.
4.4 Reflection from a Mirror Moving Parallel to Its Surface
Up to this point, the previous two sections have not introduced any fundamentally new results. As discussed earlier, when the sources are independent, adopting Lorentz coordinates poses no difficulty—each source can be assigned its own coordinate system, complete with an independent set of clocks. However, the situation becomes more complex when a secondary source interacts with light emitted by a primary source. In such cases, it becomes necessary to apply electrodynamics within an absolute time framework.
With the results from Chapter 3 now at our disposal, we are in a position to examine the intriguing phenomenon of radiation from moving sources.
We begin by analyzing the intensity of light reflected when a stationary source illuminates a moving mirror. It is commonly assumed that if the mirror moves tangentially to its surface, the law of reflection remains unaffected—just as it does for a stationary mirror (see Fig. 2). According to this view, a monochromatic plane wave incident normally on a small aperture of a moving mirror is reflected as an oblique beam. In other words, it is assumed that the energy transport velocity coincides with the phase velocity. This assumption, frequently found in textbooks (see, for example, [9] and [23]), is, however, incorrect.
To examine this claim, we analyze the reasoning found in standard texts 121212 Many sources claim that no aberration occurs when light reflects off a mirror moving parallel to its surface. For example, Sommerfeld [23] states: ”For a mirror moving tangentially to its surface, the law of reflection that holds for a stationary mirror is preserved.”. Textbooks typically analyze reflection using two Lorentz reference frames. The fixed (lab) frame $S$ is at rest relative to the plane-wave emitter, while the moving frame $S^{\prime}$ travels at velocity $v$, in which the mirror is at rest. Both frames use a Cartesian coordinate system with the $x-y$ plane tangent to the mirror’s surface. The $x$-axis aligns with the mirror’s velocity $v$. For simplicity, we consider light incident from the $z$-direction in the lab frame. The incident wave is described by its four-dimensional wave vector, where the time-like component represents the angular frequency $\omega$, and the space-like components define the propagation direction. In the lab frame $(t,x,y,z)$, the wave vector has components $k_{1}=(\omega,0,0,-\omega/c)$, with the negative sign indicating propagation toward the mirror (see Fig. 3a). Our goal is to determine the wave vector of the reflected beam.
Figure 2: A mirror moving parallel to itself. According to standard textbooks treatments, the energy transport of the reflected light beam remains undeviated.
Figure 3: The aberration of light explained using relativistic kinematics, formulated in terms of the wavenumber four-vector. Reflection geometry as observed from (a) the laboratory frame and (b) an inertial frame co-moving with the mirror.
The argument that light reflected from a tangentially moving mirror experiences no aberration proceeds as follows. It is most straightforward to analyze the reflection in the mirror’s rest frame (Fig. 3b), where the surface is stationary and the standard laws of optical reflection apply. We consider aberration effects only up to the first order in $v/c$. In this frame, an observer at rest with respect to the mirror perceives the incoming wave vector as $k_{1}^{\prime}=(\omega,-v\omega/c^{2},0,-\omega/c)$. The effect of reflection is to reverse the sign of the $z^{\prime}$ component of the wave vector, $k^{\prime}_{2}=(\omega,-v\omega/c^{2},0,\omega/c)$. We now obtain the reflected wave vector in the lab frame by applying the inverse Lorentz transformation: $k_{2}=(\omega,0,0,\omega/c)$. This vector represents a light beam traveling away from the mirror, having the same frequency as the incoming beam. This confirms that the reflection follows the usual laws of geometrical optics and that the beam undergoes no aberration.
There are two fundamental errors in solving problems related to the reflection of light from a mirror moving tangentially to its surface. The first error arises from applying the concept of a plane wave and an infinite plane mirror to the case of tangential motion. This assumption leads to an inherent contradiction: if the mirror were truly infinite, the problem would not be time-dependent. This hidden assumption—namely, the absence of time dependence—eliminates aberration, which is unsurprising. However, only the motion of a finite mirror is physically meaningful. From an electrodynamics perspective, only the velocity of a finite mirror has physical significance. Textbook treatments often overlook the crucial interaction between light and the moving edges of the mirror. A well-defined energy transport problem requires specifying both the source and mirror apertures—an essential detail frequently neglected in conventional explanations.
[…]
The second error in the textbook argument is conceptual. Consider an observer at rest in the inertial frame who measures the direction of energy transport. This observer describes the light beam emitted from a stationary source using the Minkowski metric. According to textbooks, the metric of the moving mirror is assumed to be the same Minkowski metric as that of the stationary emitter.
The textbook authors incorrectly assume that a common Lorentz time coordinate axis can be assigned to both the emitter and the mirror. This is a misconception. A fundamental question arises: how should synchronization be established in the laboratory frame after the mirror has undergone acceleration?
A distinctive feature of this problem in the context of relativistic kinematics is that the emitter (along with the observer and measuring devices) remains at rest in the lab’s inertial frame, while the mirror moves at a constant velocity relative to the lab frame, interacting with the radiated light beam. How can we address the issue of the emitter-mirror relative velocity? One approach is to assign a Lorentz time coordinate to describe the emitter’s radiation. However, for the moving mirror, this time coordinate would act as an absolute time reference, and its motion would be described by a Galilean transformation in this framework. Without altering synchronization in the lab frame, a common set of synchronized clocks for both the mirror and the emitter can only be established under absolute time coordinatization—that is, when simultaneity is absolute.
Suppose we re-synchronize the clocks in the lab frame to define the Lorentz time coordinate for the boosted mirror. In this framework, the reflection of light is described using Maxwell’s equations. However, this time transformation merely results in a rotation of the radiation phase front of the incoming plane wave by an angle $v/c$. Consequently, in this new coordinatization, Maxwell’s equations no longer directly apply to describing the emitter’s radiation. Notably, a key feature of this coordinatization is that the energy transport velocity differs from the phase velocity of the incoming light beam.
This problem in special relativity can be effectively analyzed using an absolute time coordinatization approach. We demonstrate that when a mirror with a finite aperture moves tangentially while a plane wave of light falls normally onto it, the energy transport of the reflected light deviates. In this case, the aberration of light arises directly from the time dependence of the mirror’s position, necessitated by its finite aperture.
Notably, in absolute time coordinatization, the wave equation explicitly depends on the velocity vector. Consequently, the solution naturally involves light beams of different frequencies, incorporating the Doppler effect. For a tangentially moving mirror with a finite aperture, angular frequency dispersion—an effect of first order in $v/c$—cannot be neglected.
We have already emphasized that Chapter 3 is arguably the most important part of this book. To our knowledge, the operational interpretation of absolute time coordinatization, as well as the distinction between absolute time synchronization and Einstein’s time synchronization from an operational perspective, has not been explicitly addressed in the literature. Applying the Galilean boost, we substitute $x\to x-vt$ while keeping time unchanged in the Minkowski metric, $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$, leading to the metric given in Eq. 11. This metric characterizes the electrodynamics of the moving mirror as observed from an inertial reference frame, representing the measurements from the perspective of an inertial observer.
There is an intuitively plausible reason why textbook authors conclude that light reflected from a tangentially moving mirror experiences no aberration. A comparison between conventional aberration theory for light and sound theory can provide insight. In Section 4.15, we will consider a sound emitter at rest in the atmospheric frame and examine a mirror moving tangentially along its surface with velocity $v$. According to sound theory, the reflected beam’s energy transport does not deviate because the group velocity equals the phase velocity. When a plane wave falls the mirror normally, it produces an oblique sound beam. The (diagonal) wave equation for the moving mirror in the atmospheric frame is identical to that of an emitter at rest. Textbooks often draw important parallels between sound and light aberration in the first-order approximation  [^13] We should make one additional remark regarding two common errors when solving problems related to reflection from a mirror moving tangentially to its surface. These errors are not independent. Textbooks typically assume that the metric of the moving mirror is diagonal, similar to the metric of a ’plane wave’ emitter at rest, and that the group velocity equals the phase velocity. As a result, regardless of the mirror’s size, there is no deviation in the energy transport of the reflected light beam.. This, however, is a misconception. There is a fundamental difference between the propagation of light and sound. A proper treatment of light aberration must be based on the pseudo-Euclidean geometry of space-time.
4.5 Solving the Emitter-Mirror Problem in (3+1) Space-Time
The aberration of light problem can be addressed within a ”single inertial frame” framework, without invoking Lorentz transformations. When an object is illuminated by a monochromatic, spatially coherent source, a particularly simple method exists for calculating the reflected intensity. This approach, rooted in Fourier transform techniques from spatial filtering theory—specifically, Abbe diffraction theory—offers conceptual elegance.
In this context, Abbe’s method treats the mirror as a diffraction grating that decomposes an incident plane wave into multiple diffracted beams, each representing a Fourier component of the reflected power. Since a finite-aperture mirror is a non-periodic object, it produces an infinite continuum of diffracted beams.
A basic example of a diffraction grating is illustrated in Fig. 4. Assuming the grating’s reflectance follows the function $R=g(K_{\perp})\sin{(\vec{K}_{\perp}\cdot\vec{r})}$, the reflectance is sinusoidally modulated in space. It is important to note that the permanent reflectance distribution grating discussed here is only our mathematical model and we do not need to discuss how it can be created.
Figure 4: Bragg diffraction grating at normal incidence. The reflectance is sinusoidally space-modulated.
The wave vectors $\vec{k}_{i}$ shown in Fig.4 represent the propagation vector of the incident plane wave, which is assumed to be directed perpendicularly to the surface. The vectors $\vec{k_{s}}^{(+)}$ and $\vec{k_{s}}^{(-)}$ indicate the scattered light. The Bragg condition $\vec{k}_{s}=\vec{k}_{i}\pm\vec{K}_{\perp}$, describes the relationship of the incident and scattered waves. Since the scattering occurs from a sinusoidal grating rather than a set of discrete planes (grooves), the first-order maximum dominates.
We assume that the grating vector $\vec{K}_{\perp}$ lies parallel to the surface of the grating, while the incident wave is normal to the surface, as illustrated in Fig. 4. Under the Bragg condition for small angles, the scattering angle is approximately given by $\theta\approx K_{\perp}/k_{i}$.
Let us analyze the reflection of light from a mirror moving parallel to its surface, relying solely on relativistic kinematics. After applying a Galilean boost in the case of absolute time coordinatization, we employed the metric given by Eq. 11, which describes the electrodynamics of a moving light source as observed from an inertial frame.
From the structure of Eq. 11, it becomes evident that the electromagnetic field configuration — expressed in terms of the coordinates $x$ and $t$ of a detector at rest in the inertial frame—coincides with the field configuration at position $x-vt$ at time $t=0$.
Using a Fourier approach, we showed that the problem could be reduced to the case of reflection from a tangentially moving grating. We recall that with a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ with the wavenumber vector $\vec{k}$ and the frequency $\omega$ dispersion equation in the absolute time coordinatization becomes: $(1-v^{2}/c^{2})k_{x}^{2}+2vk_{x}\omega/c^{2}+k_{z}^{2}-\omega^{2}/c^{2}=0$.
Within a single inertial frame, the equivalence between active and passive Galilean boosts implies that translating the system in one direction is indistinguishable from shifting the coordinate system in the opposite direction by the same amount. According to this equivalence, Maxwell’s equations remain valid in the comoving coordinate system. The wavenumber vector of the diffraced wave in the comoving coordinates is fixed by initial conditions. In fact, $k_{z}=\omega_{i}/c$, $k_{x}=K_{\perp}$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated reflectance. Using the initial conditions and the Galilean transformation back to the lab frame, we obtain the following expression $\omega=\omega_{i}+K_{\perp}v$, $k_{z}=\omega_{i}/c$, $k_{x}=K_{\perp}$. Substituting there expressions into dispersion equation, we verify that it is satisfied, as expected.
We find the requirement that the wavenumber $K_{\perp}$ and the frequency change $\Delta\omega$ are related by $\Delta\omega=K_{\perp}v$. This phenomenon is interpreted as a Doppler effect, as the reflection occurs from a moving, rather than stationary, system. As we consider aberration effects only to first order in $v/c$, and under the assumption $cK_{\perp}/\omega_{i}\ll v/c$, no second-order terms appear in the expression for $\Delta\omega$.
A crucial aspect of describing a ”single inertial frame” cannot be overstated. If the light source is at rest while the mirror is in motion, it follows that the electrodynamic equations must remain identical for all electromagnetic waves. In other words, the dispersion equation in absolute time coordinatization must be consistently applied to both incoming and scattered waves (Fig. 4). In our previous discussion on absolute time coordinatization, we established that an emitter at rest must still be governed by Maxwell’s electrodynamics. In this framework, the dispersion equation simplifies to $k_{i}^{2}-\omega_{i}^{2}=0$. From the initial conditions, we derive $\vec{k}_{i}=\vec{e}_{z}k_{z}$ and $\omega_{i}=ck_{z}$. However, an apparent contradiction arises—one that is resolved through a geometric analysis of light reflection. A key feature of this geometry is that even after applying a Galilean transformation along the $x$-axis, the dispersion equation in absolute time coordinatization retains its diagonal form: $k_{z}^{2}-\omega^{2}=0$ for the incident wave. To clarify the fundamental physical principles, we examined a simple case.
[…]
Figure 5: Tangentially moving mirror with a small aperture at normal incidence. The effect of aberration is analyzed using absolute time coordinatization.
One of the key conclusions of the preceding discussion is a striking prediction in the theory of the aberration of light—specifically, the deviation of energy transport for light reflected from a mirror moving tangentially to its surface. When a plane wave (in absolute time coordinatization) falls such a mirror perpendicularly, the reflected beam exhibits a deviation in energy transport (see Fig. 5). This phenomenon can be understood as a straightforward consequence of the Doppler effect [^14] The Doppler effect serves merely as an intermediate mathematical description; when the full calculation is performed, only the geometric (aberration) effect remains..
A common misconception in the literature asserts that the direction of propagation in reflection from a tangentially moving mirror is not determined by the normal to the wavefront [^15] A widespread misunderstanding stems from assuming that the orientation of radiation wavefronts has objective significance. As Norton [24] notes: ”One might try to escape the problem by supposing that the direction of propagation is not always given by the normal to the wavefront. We might identify the direction of propagation with the direction of energy propagation, supposing the latter to transform differently from the wave normal under Galilean transformation. Whatever may be the merits of such proposals, they are unavailable to some trying to implement a principle of relativity. If the direction of propagation of a plane wave is normal to the wavefronts in one inertial frame, then that must be true in all inertial frames.”. This misunderstanding arises from a failure to distinguish between convention-dependent and convention-invariant aspects of the theory. The direction of energy transport is an objective, convention-invariant quantity. In contrast, the phase front orientation—corresponding to an observer’s plane of simultaneity—lacks objective meaning, as no experimental method can precisely determine this orientation due to the finite speed of light.
Since the phase front orientation has no physical reality within the angular range $v/c$, one might ask: why must we account for its exact orientation in electrodynamics calculations? The answer lies in the fact that, when analyzing radiation beam evolution within a single inertial frame, the phase front orientation appears unchanged. While this invariance lacks objective physical meaning, it serves as a useful tool in solving electrodynamics problems. A useful analogy can be drawn with gauge transformations in Maxwell’s electrodynamics.
4.6 A Moving Mirror: Lorentz Transformations-Based Explanation
To maintain a Lorentz coordinate system for a moving mirror in the lab frame, clock resynchronization is necessary. This entails a time shift of the form given in Eq. 13, namely $t\to t+xv/c^{2}$, which is valid to first order in $v/c$. This time shift corresponds to an effective rotation of the phase front of the incoming radiation by an angle of $v/c$. Within this framework, the mirror’s reflection is analyzed using Maxwell’s equations.
For a plane wave of the form $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$, the dispersion equation in Maxwell’s electrodynamics simplifies to $k_{z}^{2}+k_{x}^{2}-\omega^{2}/c^{2}=0$.
In the Lorentz-synchronized coordinates, the electric field of the incident wave transforms to: $\exp[i(-\omega_{i}z/c+v\omega_{i}x/c^{2}-\omega_{i}t)]$. Using a Fourier decomposition, the problem reduces to that of wave reflection from a grating moving tangentially with velocity $v$. By invoking the equivalence between passive and active Lorentz transformations, we find that in the lab frame the diffracted wave components are given by: $\omega=\omega_{i}+K_{\perp}v$, $k_{z}=\omega_{i}/c$, $k_{x}=v\omega_{i}/c^{2}+K_{\perp}$. Substituting these expressions into the dispersion relation confirms that it remains satisfied, as expected.
As a result of the Doppler effect under Lorentz coordinatization, the reflection from a finite-sized moving mirror exhibits angular frequency dispersion. This implies that the reflected light beam has a group velocity along the $x$-direction given by: $d\omega/dk_{x}=v$, consistent with the prediction from Maxwell’s equations that the group velocity equals the phase velocity.
It is also insightful to analyze the electrodynamics of an emitter at rest. A rotation of the $x$-axis without changing the orientation of the $t$-axis is described by a Galilean transformation $x^{\prime}=x-vt$, $t^{\prime}=t$. Conversely, a rotation of the $t$-axis without changing the orientation of the $x$-axis is described by the transformation $x^{\prime}=x$, $t^{\prime}=t+xv/c^{2}$. To first-order, this transformation mathematically is similar to the Galilean transformation $x^{\prime}=x+vt$, $t^{\prime}=t$.
This time transformation introduces a new crossed term in the wave equation which describe emitter radiation, predicting a group velocity $d\omega/dk_{x}=-v$. However, since there is an angle $v/c$ between the wave vector $\vec{k}$ and the $z$-axis, and considering the initial condition, the radiated light beam from the stationary source propagates along the $x$-axis with a group velocity $d\omega/dk_{x}=0$, as expected. This outcome is unsurprising—convention-independent physical results remain unchanged under the new variables.
4.7 Discussion
Physicists investigating the textbook approach to the aberration of light often find themselves puzzled by an apparent inconsistency. When analyzing the reflection of light from an infinite mirror moving normally to its surface, the standard textbook reasoning—based on two Lorentz reference frames—correctly predicts the variation in light frequency upon reflection. Relativistic kinematics successfully describes this phenomenon. However, a natural question arises: why does the same method yield incorrect results when the mirror moves tangentially to its surface?
The key to understanding this discrepancy lies in the nature of the motion. In the case of a normally moving mirror, the velocity is perpendicular to the plane of oscillating elementary sources (dipoles). As a result, the plane of simultaneity in absolute time coordinates retains the same orientation in Lorentz coordinates. For collinear motion, relativistic kinematics affects the system only at the second order.
Now, let us consider frequency measurement within the framework of special relativity. Suppose we use a Fabry-Perot interferometer or a grating spectrometer. In such cases, the measured frequency corresponds to the wavelength of the standing wave. This principle applies universally to all frequency measurements, which inherently obey the relationship between light frequency and interference patterns. Importantly, while the deviation of energy transport direction is a geometric effect, interference patterns themselves are independent of the chosen metric. In other words, in space-time geometry, phase remains a four-dimensional invariant. For collinear motion, both the Minkowski metric and the metric given by Eq. 11 yield identical predictions (see Chapter 13 for further details).
Another crucial aspect to consider is the size of the mirror. In the case of an infinite mirror moving normally to its surface, the motion of the mirror is a physically observable effect. Even for an infinite mirror, the problem remains time-dependent, and its velocity carries physical significance. This makes the infinite mirror a useful conceptual tool for analyzing the Doppler effect in such a scenario.
4.8 Large Aperture Mirror
In this chapter, we focus on a specific range of problem parameters where the effects of emitter edges are neglected. Although this treatment of aberration in light theory is an approximation, it holds significant practical importance. We also consider the case where the transverse size of the emitter is much smaller than that of the moving mirror. It is important to note that this scenario does not occur in stellar aberration measurements. However, analyzing this case allows us to gain a clearer understanding of the underlying physical principles. Notably, it can be shown that in this situation, energy transport deviation is absent (see Fig. 6). To address the moving mirror problem, we employ the Fourier transform method as a direct approach.
Figure 6: Tangentially moving mirror with a large aperture at normal incidence. When a light beam is incident normally on the mirror, the reflected beam undergoes no deviation in energy transport.
4.9 Analysis of Transmission through a Hole in an Opaque Screen
Previously, we demonstrated that when a finite-aperture mirror moves tangentially to its surface and a plane wave of light is incident normally on the mirror, an aberration (deviation in energy transport) occurs in the reflected light. In this section, we consider a more practically significant problem: a screen moving with velocity $v$ along its surface in the lab frame. It is commonly assumed that light passing through a hole in such a moving opaque screen does not experience any deviation in energy transport (see Fig. 7).
However, a common mistake in relativistic optics arises from aberration effects associated with a tangentially moving screen containing a hole. We analyze this system using a Fourier transform method similar to the one employed earlier. The screen with a hole acts as a diffraction grating, decomposing the incident plane wave into multiple diffracted plane-wave components. Each of these components corresponds to a Fourier mode of the transmitted light beam.
The gratings discussed thus far modulate the amplitude of an incident plane wave through a periodic reflection function. However, our analysis can be readily extended to gratings that modulate the amplitude of incident light using a periodic transmission function. Consider a grating whose transmittance varies according to the function $T=g(K_{\perp})\sin{(\vec{K}_{\perp}\cdot\vec{r})}$, as illustrated in Fig. 8. In this case, the transmittance exhibits a sinusoidal spatial modulation. Notably, all previously derived equations remain valid for the forward-scattered beams.
Figure 7: Aberration of light in an inertial frame of reference. A tangentially moving screen with a hole allows the passage of light. According to textbooks, a monochromatic plane wave incident normally on the screen produces a transmitted oblique beam.
Figure 8: Bragg diffraction grating at normal incidence. The transmittance is sinusoidally modulated in space.
Figure 9: Aberration of light in an inertial frame of reference. A tangentially moving screen with a hole allows a normally incident monochromatic plane wave to pass through, generating a transmitted light beam. The explanation of the aberration effect relies on absolute time coordinatization.
Our approach reveals a remarkable prediction of the theory of aberration of light concerning the deviation of energy transport for light passing through a hole in a moving screen. Specifically, when a screen with a hole moves tangentially while a plane wave of light falls normally on it, the transmitted light exhibits a deviation in energy transport (see Fig. 9).
4.10 Spatiotemporal Transformation of the Transmitted Light Beam
Suppose the transmitted light pulse propagates in the $x-z$ plane. Our focus is on the space-time intensity distribution within this plane. Spatiotemporal coupling naturally emerges in the transmitted radiation behind the screen due to the angular-frequency dispersion introduced during the transmission process. In most cases, the emitted light beam can be accurately represented as the product of independent spatial and temporal factors. However, this assumption breaks down when the light must pass through an aperture in a moving opaque screen, as the transmission process disrupts the simple separation of space and time dependencies.
We begin by expressing the field of an emitted pulse as $E=b_{i}(x)\exp[i\omega_{i}(z/c-t)]$. The initial amplitude distribution, $b_{i}(x)$, in front of the moving screen is the optical replica of the emitter’s aperture. The electric field of the transmitted pulse can be represented in the reciprocal domain as $\bar{E}(\Delta k_{x},\Delta\omega)=\bar{E}(K_{\perp},K_{\perp}d\omega/dk_{x})$, which corresponds to the Fourier component of a beam with angular frequency dispersion, where $d\omega/dk_{x}=v$. Taking the inverse Fourier transform, we obtain $E=b(x-vt)\exp[i\omega_{i}(z/c-t)]$. This represents the field immediately behind the moving screen. The beam profile, $b(x-vt)$, is the optical replica of the moving aperture. Now, consider a screen positioned at rest at a distance $l$ behind the moving aperture. For simplicity, we assume that the Fresnel number is large, $N_{F}=D_{h}^{2}\omega/(cl)\gg 1$, allowing us to neglect diffraction effects. Here, $D_{h}$ is the characteristic aperture size. From this, we conclude that the light spot on the observer’s screen moves with the same velocity $v$ as the moving screen. Additionally, we observe that a particular quantity remains unchanged: the combination $x-vt$ is the same for both the spot and the aperture. What does this imply? A careful examination reveals that events occur simultaneously at two separate locations along the $z$-axis  [^16] Spatiotemporal coupling is typically discussed in the literature in the context of ultrashort laser pulse propagation through a grating monochromator. In this scenario, in addition to considering phase fronts, one must also account for planes of constant intensity—known as pulse fronts. In a grating monochromator, the different spectral components of the outgoing pulse propagate in different directions. The electric field of a pulse with angular dispersion can be expressed in the Fourier domain $[\Delta k_{x},\Delta\omega]$ as $\bar{E}(\Delta k_{x}-p\Delta\omega,\Delta\omega)$. Its inverse Fourier transform to the space-time domain $[x,t]$ results in $E(x,t+px)$, representing a pulse with a pulse-front tilt. In our case of interest, we consider light diffracted by a tangentially moving set of gratings, where the Doppler effect induces frequency dispersion given by $d\omega/dk_{x}=v$. Consequently, the spatiotemporal coupling due to light transmission through a hole in a moving screen differs significantly from the conventional pulse-front tilt distortion..
4.11 Applicability of Ray Optics
Today, it is commonly suggested that the phenomenon of light aberration can be explained using a corpuscular model of light. According to this model, light particles (corpuscles) fall a moving screen normally, generating an oblique light beam as shown in Fig. 10. However, this argument remains incorrect and continues to persist. A satisfactory treatment of light aberration should be based on the electromagnetic wave theory.”
Figure 10: A tangentially moving screen with a hole in it. Light corpuscles fall normally on the screen and, according to the literature, produce an oblique light beam.
Some experts argue that the applicability of the corpuscular model in the theory of light should be reinterpreted as the applicability of ray optics.  [^17] To quote Brillouin [25]: ”Fig. 5 explains the situation assuming a simplified device consisting of a parallel plate moving with uniform velocity $v$ in the horizontal direction. Monochromatic light falls normally on the plate and generates an oblique ray.” This oblique effect is demonstrated in Fig. 10.. Let us examine what happens, according to ray optics, in our case of interest. When light rays fall normally on a moving screen, they generate an oblique ray beam, as illustrated in Fig. 10. In this scenario, treating light rays as tiny particles leads to a well-known effect.
Next, we discuss the region of applicability of ray optics. The use of ray optics in the theory of light aberration presents complexities. One might intuitively expect, based on textbook reasoning, that ray optics should apply to any spatially incoherent radiation. However, this assumption leads to incorrect results. Specifically, a completely incoherent source (such as an incandescent lamp or a star) consists of elementary point sources that are statistically independent and have varying offsets.
The radiation field produced by a completely incoherent source can be understood as a linear superposition of the fields of individual elementary point sources. Each elementary source generates an effective plane wave in front of an aperture. Consequently, the transmission process inevitably involves the diffraction of a plane wave through the aperture. Notably, any linear superposition of radiation fields from elementary sources retains fundamental single-source characteristics, such as deviations in the energy transport direction. This reasoning explains why ray optics is not applicable to the aberration theory of light originating from (spatially) completely incoherent sources.
To illustrate the applicability of ray optics in light aberration theory, we consider a specific class of spatially incoherent light beams. A method can be proposed to generate a ray beam from primary sources using an array of randomly phased lasers. Such a planar source produces rays that fall normally on the moving screen (within the laser Rayleigh range) and generate an oblique transmitted ray beam, as shown in Fig. 10. Intuitively, a tangentially moving screen with a hole acts as a switcher for the lasers. A luminous spot moving at velocity $v$ can, in principle, be created more simply—so to speak, ”manually.” By arranging laser-like sources along the $x$-axis and activating them sequentially from left to right with a given time lag, a luminous spot can be made to move at any desired velocity, even exceeding $c$. However, it is crucial to note that in this process, no information is transmitted along the $x$-axis, as each source radiates independently.
4.12 Moving Large Aperture Emitter
Let us now consider a ”plane-wave” emitter that starts at rest in the lab inertial frame and is then accelerated to a velocity $v$ along the $x$-axis. An emitter with a finite aperture acts as an active medium, breaking the radiated beam into multiple diffracted plane-wave components. Each of these beams corresponds to a Fourier component of the emitted radiation. As discussed earlier in this chapter, the energy transport of coherent light from a tangentially moving emitter deviates from the expected path—a well-known phenomenon related to the aberration of light in an inertial frame of reference (Fig. 1).
In our case of interest, the screen remains at rest in the lab frame, while the emitter moves tangentially to its surface at a constant velocity. For simplicity, we assume that the moving ”plane-wave” emitter is significantly larger than the hole in the screen. Now, consider an observer at rest relative to the screen, measuring the direction of energy transport (Fig. 11).
Figure 11: Aberration of light in an inertial frame of reference. A large-aperture plane-wave emitter moves tangentially to its surface, while the screen remains at rest. In this scenario, no aberration of light occur.
If the screen is at rest while the light source is in motion, it is evident that the equations of electrodynamics must remain identical for all electromagnetic waves. In our previous discussion, we established that a stationary screen must be described by Maxwell’s electrodynamics. Consequently, the dispersion equation in the Lorentz coordinatization must be consistently applied to both incoming and scattered waves. In Maxwell’s electrodynamics, the dispersion equation is given by: $k_{x}^{2}+k^{2}_{y}+k^{2}_{z}-\omega^{2}=0$. A key characteristic of the discussed geometry is that, even after applying a Galilean transformation along the $x$-axis, the dispersion equation in the absolute time coordinatization retains the same diagonal form for a normally incident wave: $k^{2}_{z}-\omega^{2}=0$. A closer examination of the physics reveals that, in the inertial laboratory frame where the screen is stationary, the problem corresponds to a steady-state transmission. Since the Doppler effect is absent, the transmitted beam travels vertically, having lost its horizontal group velocity component. As a result, the transmission behaves as depicted in Fig. 11. It is important to emphasize the following point: when light passes through a small aperture, the resulting beam undergoes diffraction, perturbing its fields. This diffraction effect removes any information about the emitter’s motion, making the transmitted light independent of its original source movement.
4.13 A Point Source in an Inertial Frame of Reference
Above, we considered a single moving ”plane wave” emitter in an inertial frame of reference. When analyzing the aberration of light in such a frame, two types of sources are particularly useful to consider:
(a) A ”plane wave”-like emitter
(b) A point-like source (or, more generally, a spatially incoherent source)
The diffraction of a source field is typically categorized into Fresnel (near-zone) and Fraunhofer (far-zone) diffraction. In Fraunhofer diffraction, the wave phase is assumed to vary linearly across the detector aperture, which occurs, for example, when a plane wave impinges on the aperture at an angle to the optical axis. In contrast, Fresnel diffraction replaces this linear phase variation with a quadratic one. In the far zone, both types of sources effectively produce a plane wave in front of the pupil detection system. This implies that, in the Fraunhofer zone, the instrument always exerts a physical influence on the measurement of light aberration.
[…]
Figure 12: Aberration of light in an inertial frame of reference. In the Fraunhofer diffraction region, a point source produces in front of a hole aperture effectively a plane wave. In this case, the aberration of light phenomenon does not occur.
Figure 13: Diagram of a modified source of the ”plane wave” emitter type. A point source of light is placed in the front focal plane of a lens.
One of the key conclusions of the discussion above (see Fig. 11) is that the aberration of a moving point source is absent in an inertial frame (Fig. 12). This leads to several remarkable effects, stemming from the fact that the transmitted light beam does not carry information about the motion of the point source. In fact, this principle is central to the binary star paradox discussed in Chapter 6.
To further explore the aberration of light in an inertial frame, it is useful to consider a slightly modified source—a ”plane wave” emitter. We adopt the model described in Section 4.2 because it is both relatively simple and sufficiently general to serve as a prototype for understanding aberration phenomena. While no one has performed all the ”thought experiments” exactly as described here, the outcomes can be confidently predicted based on the laws of special relativity, which are themselves grounded in experimental evidence.
Figure 13 illustrates the modified ”plane wave” emitter setup. Though it appears more complex at first glance, it is practically feasible to construct. The setup consists of a point source of light positioned in the front focal plane of a lens. The source’s dimensions are assumed to be on the order of $\lambda$, where $\lambda$ is the wavelength of optical radiation.
4.14 Experimental Test: Reflection from a Moving Grating
Textbook treatments suggest there is strong theoretical evidence for the absence of an aberration effect when light reflects from a mirror moving transversely (parallel to its surface). Although this result was established long ago by Pauli [9] and Sommerfeld [23], a direct experimental verification remains absent. This raises a natural question: why has it not been possible to experimentally detect the aberration effect in this configuration? Consider a mirror moving tangentially to its surface, assuming the mirror size is smaller than the transverse size of the incoming light beam. Notably, we analyze a setup where the aberration angle $\theta_{a}=v/c$ is significantly larger than the diffraction divergence $\lambda/D_{m}$ of the reflected radiation. Let us estimate the parameters for the Earth-based setup using tangentially moving mirror and an optical laser. For instance with a light wavelength of $\lambda=0.3~{}\mu\mathrm{m}$ and a mirror size $D_{m}$ of 10 cm, the mirror speed must satisfy $v\gg 1000~{}\mathrm{m/s}$. Clearly, current experimental capabilities are insufficient to meet these demanding conditions. Fortunately, there exists a more practical alternative to this otherwise formidable experimental challenge.
Previously, we analyzed the reflection of light from a mirror moving parallel to its surface, relying solely on relativistic kinematics. After applying a Galilean boost in the case of absolute time coordinatization, we employed the metric given by Eq. 6, which describes the electrodynamics of a moving light source as observed from an inertial frame. Using a Fourier approach, we showed that the problem could be reduced to the case of reflection from a tangentially moving grating. From the resulting dispersion relation, we derived a Doppler shift in the diffraction maxima due to the grating’s motion.
The consistency between our theoretical framework and experimental ob- servations—such as the measured frequency shift by a tangentially moving grating [26]—provides further support for the validity of our approach.
In contrast, special relativity textbooks commonly conclude that no Doppler effect occurs for light reflected from a tangentially moving grating. This argument is typically made for a tangentially moving mirror but applies equally to gratings. The issue stems from the widely accepted assumption that the metric of a moving grating (or mirror) is the same as the Minkowski metric of the stationary emitter.
We now present a simple explanation of our central result. The clearest understanding arises when we accept, as an experimental fact, the presence of a Doppler frequency shift in radiation diffracted from a moving grating.
Within our relativistically consistent framework—based on the equivalence of active and passive boosts within a single inertial frame—the diffracted radiation must satisfy the dispersion relation: $(1-v^{2}/c^{2})k_{x}^{2}+2vk_{x}\omega/c^{2}+k_{z}^{2}-\omega^{2}/c^{2}=0$. The frequency shift $\Delta\omega=K_{\perp}v$ is exactly compensated by the cross term $2vk_{x}\omega/c^{2}$, ensuring that the dispersion relation remains valid.
In contrast, conventional treatments in special relativity textbooks describe diffraction from a moving grating using Maxwell’s equations, leading to the dispersion relation: $k_{z}^{2}+k_{x}^{2}-\omega^{2}/c^{2}=0$. This equation is isotropic and independent of velocity. As a result, it contradicts experimental observations: the Doppler frequency shift cannot be accounted for or compensated.
However, experimental evidence clearly demonstrates that the frequency of the reflected radiation depends on the tangential velocity of the grating. Doppler frequency shifts can be detected with high precision using advanced spectroscopic techniques, even at very low grating velocities. For instance, the authors of [26] used a Mach-Zehnder-type interferometer to measure these shifts.
This striking discrepancy suggests that the conclusions reached by Pauli [9] and Sommerfeld [23] are inconsistent with the principles of special relativity.
4.15 Differences between the Light and Sound Aberration
Having completed our discussion on the aberration of light in an initial inertial frame, we now turn to optical phenomena in an accelerated frame. However, before proceeding, we will first explore optical-acoustic analogies. Sound is a periodic motion of air caused by vibrations of a source. The effect of sound aberration is best understood using a model of a single plane-wave emitter. Specifically, we consider a two-dimensional array of identical, coherent elementary sources uniformly distributed on a given $(x-y)$ plane $P$, all emitting waves simultaneously. We analyze a scenario in which a finite-aperture mirror moves tangentially to its surface. For simplicity, we assume the mirror’s size is small compared to the plane-wave emitter. An observer, at rest in the atmosphere frame, describes the emitted sound beam using the diagonal wave equation: $\nabla^{2}f-\partial^{2}f/\partial(v_{s}t)^{2}=0$, where $v_{s}$ is the speed of sound.
The emitter radiates a plane wavefront in the vertical $z$-direction, and the moving mirror obeys the same wave equation as the stationary emitter. The amplitude of the beam emitted by the oscillating sources can be understood as a superposition of spherical wavelets, as dictated by Huygens’ principle. The wavefront at any given instant is the envelope of these wavelets. Since the diagonal wave equation exhibits no intrinsic anisotropy, the energy transport of the reflected sound beam remains unchanged. A plane wave incident normally on the finite-aperture mirror generates a reflected oblique beam. Notably, in the atmosphere frame, the phase and group velocities remain equal.
Textbooks generally state that, to first order in $v/c$, there is no fundamental difference between the aberration of light and sound. The geometry of reflection from a moving mirror is identical (Fig. 2). Now, consider a screen moving tangentially to its surface with velocity $v$ in the atmosphere frame. According to sound theory, energy transport for a sound beam transmitted through a hole in a moving screen remains undeviated. Consequently, it is commonly assumed that the energy transport for light transmitted through a hole in a moving opaque screen follows the same geometric behavior (Fig. 7). In the next chapter, we will continue our discussion on the aberration of sound.
The conventional treatment of the aberration of light often leads physicists to draw misleading parallels between the aberration of sound and light in the first-order approximation. This is largely due to the fact that the wave equation governing a moving mirror (or screen) retains the same diagonal form as the wave equation of an emitter at rest. However, in sound theory, we operate within Newtonian space and time rather than the pseudo-Euclidean geometry of space-time. This distinction is crucial: the standard analysis of light aberration within a single inertial frame fails to account for the fundamental difference between the velocity of light and that of sound.
A rigorous relativistic treatment must be grounded in a space-time geometric approach. The Minkowski metric remains valid in an inertial frame where the observer is at rest, with time coordinates assigned through the slow clock transport method or Einstein synchronization, which is defined via light signals emitted by a dipole source at rest. In this Lorentz coordinate system, the Minkowski metric accurately describes the electrodynamics of a stationary light source from the perspective of an inertial observer. Now, consider the case where a mirror in an inertial frame accelerates from rest to velocity $v$ along the $x$-axis. Synchronization is maintained using the same set of synchronized clocks, preserving the Minkowski metric for both the inertial observer and the stationary light source. In this scenario, the mirror undergoes a Galilean boost, represented by the transformation $x\to x-vt$, while keeping time unchanged in the observer’s Minkowski metric. This results in Eq. 11, which describes the electrodynamics of the moving mirror from the inertial observer’s perspective. In contrast, prior literature has often made the incorrect assumption that an inertial observer (i.e., an observer at rest in an inertial frame) and an accelerated mirror within the same frame share a common Minkowski metric—an assumption long treated as self-evident. Consequently, the widely accepted approach to special relativity in the first-order approximation effectively reduces to a classical theory akin to sound propagation.
5 Aberration of Light: Non-inertial Frame of Reference
In this chapter, we revisit the problem of light transmission in non-inertial frames, focusing on the aberration of light. Using the Langevin metric in general relativity, we derive the aberration for a pulse of light traveling in an accelerated system.
The aberration problem is fundamentally resolved by recognizing the inherent asymmetry between inertial and non-inertial observers. This asymmetry, akin to the well-known Sagnac effect, arises from the absolute nature of acceleration. As Langevin noted in 1921, ”any change of velocity, or any acceleration, has an absolute meaning” [31]. Mathematically, however, there is no distinction between calculations in general relativity and special relativity when space-time curvature is absent.
The usual arguments for the relativity of motion do not apply here, as inertial and non-inertial reference frames are not equivalent. The asymmetry paradox is typically resolved by recognizing acceleration as the defining factor. Although the duration of acceleration has a negligible effect on anisotropy within the accelerated frame, the acceleration itself fundamentally determines the problem. Interestingly, the lack of a dynamical explanation for this asymmetry in special relativity has puzzled some physicists. A useful perspective is to interpret the asymmetry as a consequence of the pseudo-gravitational effects experienced by an accelerated observer. From a pedagogical standpoint, treating the accelerated frame via the equivalence principle provides valuable insight, a topic we will explore further in Chapter 12.
A natural question arises: how can different observers—such as those based on Earth and the Sun—determine which one experienced acceleration? The surprising answer is that this determination can be made solely through observations of the ”fixed stars.” In principle, acceleration is defined relative to these distant celestial objects, implying an implicit absolute reference frame.
5.1 Absolute Time Coordinatization in Accelerated Systems
We analyze the problem of light aberration in accelerated systems within the framework of special relativity. Specifically, we demonstrate that explaining optical effects in a rotating frame of reference does not require modifying special relativity or invoking general relativity. Instead, a strict adherence to special relativity suffices. To express electrodynamics laws in an accelerated frame, one must take an additional step by defining the metric of the non-inertial frame. Importantly, the metric tensor must remain a continuous function, leading to the concept of smoothly tailoring the metric tensor.
This problem in special relativity can be effectively addressed using an approach based on absolute time coordinatization. At the end of this section, we outline the mathematical extension of special relativity to accelerating systems. However, in some cases, direct physical arguments can provide quicker insights while remaining consistent with rigorous mathematical derivations.
Suppose the initial inertial system $S$ is at rest with respect to the fixed stars, and a system $S_{n}$ (along with an observer and their measuring instruments) is accelerated from rest in $S$ to a velocity $\vec{v}$ along the $x$-axis. In accelerated systems, the concept of maintaining absolute simultaneity remains logically consistent with the smooth tailoring of the metric tensor. The synchronization method involves preserving, without modification, the same set of uniformly synchronized clocks used when $S_{n}$ was initially at rest. It is well known that during acceleration (relative to the fixed stars), Einstein’s clock synchronization procedure cannot be applied. Consequently, by the time $S_{n}$ transitions to uniform motion, the metric in the accelerated reference frame will have acquired a non-diagonal form.
We begin by considering the metric as the true measure of space-time interval for an accelerated observer $S_{n}$ with coordinates $(t_{n},x_{n},y_{n},z_{n})$. To transform to the coordinates of an inertial observer $S$ moving with velocity $-v$ relative to $S_{n}$, we apply the inverse Galilean transformation $x=x_{n}+vt$, while leaving the time coordinate unchanged, $t_{n}=t$. Substituting this into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$, we obtain the transformed metric:
$ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}~{}.$ (15)
This metric describes the electrodynamics of the accelerated light source with the viewpoint of the accelerated observer measurements as viewing this of the inertial observer. Inspecting Eq.(15) we can find the components of the metric tensor $g_{\mu\nu}$ in the coordinate system of $S_{n}$. We obtain $g_{00}=1-v^{2}/c^{2}$, $g_{01}=-v/c$, $g_{11}=-1$. Note that the metric in Eq. (15) is not diagonal, since, $g_{01}\neq 0$, and this implies that time is not orthogonal to space. This result represents the Langevin metric, derived by matching the metric tensors of the accelerated frame and the initial inertial frame. Considering the Langevin metric, we conclude that the length of a physical rod $dx$ and the proper time interval $dt$ in the inertial frame will coincide with the coordinate distance and coordinate time in the accelerated frame, i.e., $dx_{n}=dx$, $dt_{n}=dt$. (For further details, see Chapter 12).
Let us now extend the mathematical framework of the special theory of relativity to include accelerated motion. To keep the mathematical complexity minimal, we assume that the reference frame $S_{n}$ starts at $t=0$ and moves with a constant acceleration $a>0$ until $t=T$. Our goal is to determine an expression for velocity from the perspective of an inertial (non-accelerated) observer.
In relativistic mechanics, uniformly accelerated motion satisfies the equation $d(\gamma\vec{v})/dt=\vec{f}/m=\vec{a}$ which integrates to $\gamma\vec{v}=\vec{a}t$. This leads to the velocity expression $dx/dt=at/\sqrt{1+a^{2}t^{2}/c^{2}}$. We emphasize that $dx(t)/dt$ represents the velocity of the accelerating frame as observed from the inertial reference frame.
Applying the inverse Galilean transformation, we have $dt_{n}=dt$, $dx=dx_{n}+atdt/\sqrt{1+a^{2}t^{2}/c^{2}}$. With this transformation, the metric in the uniformly accelerating observer’s frame takes the form:
$ds^{2}=c^{2}dt_{n}^{2}(1+a^{2}t_{n}^{2}/c^{2})^{-1}-2at_{n}dt_{n}dx_{n}(1+a^{2}t_{n}^{2}/c^{2})^{-1/2}-dx_{n}^{2}~{}.$ (16)
Since the metric tensor of space-time must be a continuous quantity, the coordinates in the accelerated and inertial stages must be matched accordingly. At $t=0$, the metric naturally coincides with that of the $S$ frame, giving $g_{00}=1$. During acceleration, we adopted an absolute time coordinatization. To ensure continuity of the metric tensor as the reference frame transitions from non-inertial to inertial motion, we employ the metric given in Eq. (16). Thus, at $t=T$, the metric of the non-inertial frame must smoothly transition into the Langevin metric. This can be achieved by relating the coordinates and time of the accelerated observer, $(x_{n},t_{n})$, to those of the inertial observer, $(x,t)$, via the Galilean boost: $x_{n}=x-vt$, $t_{n}=t$, where $v=aT/\sqrt{1+a^{2}T^{2}/c^{2}}$.
For coordinate times $T<t_{n}$, the reference frame of the accelerated system follows the metric given by Eq.(15). Importantly, $v$ represents the velocity of the frame $S_{n}$ as observed by an inertial observer during the second (inertial) stage of $S_{n}$ motion. It is evident that at $t=t_{n}=T$, the metric of the non-inertial frame continuously transitions into that of the (formerly accelerated) inertial frame. For a broader discussion on the mathematical extension of special relativity to accelerated motion, we refer the reader to [16].
Textbooks describe inertial frames as moving at constant velocity relative to the fixed stars, with different inertial frames connected by Lorentz transformations. Consequently, one might expect the coordinates and time in frames $S$ and $S_{n}$ (during the inertial stage) to be related by a Lorentz transformation. However, during the inertial segment of the trajectory, the metric in both frames remains diagonal in Lorentz coordinates. Thus, at $t_{n}=T$, the metric tensor of the accelerated frame, according to textbooks, must undergo an abrupt change from $g_{00}=(1+a^{2}T^{2}/c^{2})^{-1}$ to $g_{00}=1$.
5.2 Asymmetry Between Inertial and Accelerated Frames
The velocity of light emitted by a source at rest in the coordinate system $(t,x,y,z)$ of the inertial frame $S$ is $c$. The Minkowski metric, given by Eq. (1), implies a symmetry in the one-way speed of light in this frame. However, in the coordinate system $(t_{n},x_{n},y_{n},z_{n})$ of an accelerated frame $S_{n}$, the speed of light emitted by a source at rest in $S_{n}$ differ from $c$ This is because $(t_{n},x_{n},y_{n},z_{n})$ is related to $(t,x,y,z)$ via a Galilean transformation. This is readily verified if one recalls that the velocity of light in the reference system $S$ is equal to $c$. If $ds$ is the infinitesimal displacement along the world line of a ray of light, then $ds^{2}=0$ and we obtain $c^{2}=(dx/dt)^{2}$. In the accelerated reference system, since $x_{n}=x-vt$ and $t=t_{n}$, this expression takes the form $c^{2}=(dx_{n}/dt_{n}+v)^{2}$, as seen by setting $ds^{2}=0$ in Eq. (15). Consequently, in the accelerated coordinate system $(ct_{n},x_{n})$, the velocity of light parallel to the x-axis, is $dx_{n}/dt_{n}=c-v$ in the positive direction, and $dx_{n}/dt_{n}=-c-v$ in the negative direction, as expected.
We have identified a fundamental asymmetry between inertial and accelerated frames: Maxwell’s equations do not hold from the perspective of an observer at rest with respect to an accelerated frame $S_{n}$. The metric given by Eq. (15), which corresponds to the accelerated frame $S_{n}$, predicts an asymmetry in the one-way speed of light along the direction of relative velocity. Accelerations (with respect to the fixed stars) influence the propagation of light. In an accelerated system $S_{n}$, the velocity of light emitted by a source at rest must be modified by the acceleration—either added to or subtracted from—resulting in different speeds in opposite directions. In contrast, in an inertial frame $S$, the velocity of light emitted by a source at rest remains $c$.
One crucial aspect of describing an accelerated reference frame cannot be overstated: the metric applies to physical laws, not to physical facts. In other words, it is always possible to choose a set of variables in which the metric of the accelerated source remains diagonal. We interpret the Langevin metric as indicating that the laws of electrodynamics, when expressed in an accelerated frame, take the form of anisotropic field equations. Here, ”physical facts” refer specifically to the aberration of light emitted by a single ”plane wave” source within the accelerated frame. To account for this effect, the electrodynamics equation must be integrated using the appropriate initial conditions for the radiation wavefront.
After the boost we can see that acceleration does not affect the wavefront orientation. In fact, the variables $(t,x,y,z)$ can be expressed in terms of the variables $(t_{n},x_{n},y_{n},z_{n})$ by means of Galilean transformation $x_{n}=x-vt$, $t_{n}=t$. This transformation ensures that, following acceleration, the emitted light beam’s wavefront remains perpendicular to the vertical direction $z_{n}$, as illustrated in Fig. 14. We will explore this topic further in Section 5.5.
Figure 14: Aberration of light in an accelerated frame of reference $S_{n}$. The orientation of radiation wavefronts and the anisotropy of the speed of light are depicted using absolute time coordinatization ($t_{n}=t$) for the screen.
Figure 15: Aberration of light in an accelerated frame of reference $S_{n}$. The screen at rest within the accelerated frame. The aberration of transmitted light is independent of the source speed.
It is useful to begin with an overview of some key results. Consider an observer in the accelerated frame $S_{n}$ performing an aberration measurement. To predict to the outcome, the observer must use the non-diagonal metric given in Eq. (15). Due to the inherent asymmetry between inertial and accelerated frames, an intriguing prediction arises in the theory of light aberration. Specifically, if an opaque screen with a hole is initially at rest relative to the fixed stars and then set into motion, the apparent angular position of a ”plane-wave” emitter observed through the aperture in the accelerated frame will shift by an angle $-v/c$. This effect is illustrated in Fig. 14. Notably, the source of this asymmetry is not the relative acceleration between the two frames but rather the difference in their individual acceleration histories with respect to the fixed stars.
In Chapter 4, we analyzed the emitter-screen problem in the reference frame $S$, which remains at rest (i.e., without any history of acceleration) relative to the fixed stars. We found that, in the case of a small aperture hole, no aberration arises from the emitter, regardless of its motion. It is important to emphasize that the aberration of the light beam transmitted through the small aperture hole also remains independent of the emitter’s motion in the accelerated frame $S_{n}$ (see Fig. 15). This is because the cross-term in the metric (Eq. 15), which induces aberration in this particular case, depends solely on the velocity change of the accelerated frame relative to the initial inertial frame.
5.3 Electrodynamic Explanation of Light Aberration
It is well known that electrodynamics fully complies with the principles of relativity. For instance, we can derive the velocity of light in an accelerated frame without relying on the metric equation (15). In an inertial frame, electromagnetic fields are expressed as functions of the independent variables $x,y,z$, and $t$. According to special relativity, Maxwell’s equations remain valid in any Lorentz reference frame, and the electric field $\vec{E}$ of an electromagnetic wave satisfies the wave equation: $\Box^{2}\vec{E}=\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial(ct)^{2}=0$. By applying a Galilean transformation, the variables $x,y,z,t$ can be expressed in terms of the coordinates $x_{n},y_{n},z_{n},t_{n}$ in an accelerated frame. Consequently, the electromagnetic field of a wave emitted by a source at rest in the accelerated frame can be rewritten in terms of these transformed variables. As a result, the wave equation transforms into
$\Box^{2}\vec{E}_{n}=\left(1-\frac{v^{2}}{c^{2}}\right)\frac{\partial^{2}\vec{E}_{n}}{\partial x_{n}^{2}}+2\left(\frac{v}{c}\right)\frac{\partial^{2}\vec{E}_{n}}{\partial t_{n}\partial x_{n}}+\frac{\partial^{2}\vec{E}_{n}}{\partial y_{n}^{2}}+\frac{\partial^{2}\vec{E}_{n}}{\partial z_{n}^{2}}-\frac{1}{c^{2}}\frac{\partial^{2}\vec{E}_{n}}{\partial t_{n}^{2}}=0~{},$ (17)
where coordinates and time are transformed according to a Galilean transformation. Acceleration affects the propagation of light. The coordinate velocity of light parallel to the $x$-axis is given by $dx/dt=c-v$ in the positive direction, and $dx/dt=-c-v$ in the negative direction.
We now consider the transmission through the hole in the opaque screen in the accelerated frame $S_{n}$, Fig. 15. We describe the aberration of light based on the observations made by an observer in the same accelerated frame as the screen. We have already discussed the explanation of the aberration based on electrodynamics. Our earlier discussion is really about as far as anyone would normally need to go with the subject, but we are going to do it again. One reason is that one should know how to deal with what happens to waves in the accelerated frame from the viewpoint of electrodynamics equations. The screen containing a hole is a kind of diffraction grating that breaks up the radiated beam into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which a transmittance can be resolved. Let us assume that the transmittance of the grating varies according to the law $T=g(K_{\perp})\sin{({K}_{\perp}x)}$.
Acceleration influences the field equations, transforming the wave equation into Eq. (17). Consider a transmitted plane wave of the form $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$. With a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ with the wavenumber vector $\vec{k}$ and the frequency $\omega$ equation Eq.(17) becomes: $(1-v^{2}/c^{2})k_{x}^{2}-2vk_{x}\omega/c^{2}+k_{z}^{2}-\omega^{2}/c^{2}=0$. The wavenumber vector of the transmitted plane wave is fixed by initial conditions. Suppose the wavefront of the emitted light beam is initially perpendicular to the vertical direction before acceleration. Then, we have $k_{z}=-\omega_{i}/c$, $k_{x}=K_{\perp}$, where $K_{\perp}$ is the wavenumber accosiated with space-modulated transmittance.
[…]
In our example, plane waves with different wavenumber vectors propagate away from the screen at different frequencies. Each transmitted wave satisfies the equation $\Delta\omega/\Delta k_{x}=-v$, regardless of the sign or magnitude of the transmission angle. This implies that a transmitted light beam with a finite transverse size moves along the $x_{n}$-direction with a group velocity of $d\omega/dk_{x}=-v$ (see Fig. 15).
Now, let us address an apparent paradox. We derived the electrodynamics equations in an accelerated frame using the Galilean transformation of Maxwell’s equations. Intuitively, one might expect everything to be at rest in the accelerated frame. However, special relativity dictates the presence of a time derivative, $\partial/\partial t_{n}=v\partial/\partial x_{n}$, since $\partial/\partial x_{n}$ is nonzero. This situation differs fundamentally from electrodynamics in an inertial frame. We will explore this topic further in Chapter 9.
5.4 A Simple Explanation of Light Aberration in Accelerated Frames
Previously, we analyzed the aberration of light in an accelerated reference frame. Calculating this effect in a non-inertial frame is inherently complex. To tackle this challenge, we employed a metric tensor to derive the electrodynamics equations within the accelerated frame, using the Fourier transform method. A key consequence of the cross term in the wave equation is the emergence of a group velocity, which explains the aberration. Now, we present a straightforward explanation of light aberration in an accelerated frame using electrodynamics. To determine the aberration increment, we apply Galilean transformation laws to the electric and magnetic fields.
We have an expression for the energy flow vector of the electromagnetic field. This vector, $\vec{S}=c\vec{E}\times\vec{B}/(4\pi)$, is called ”Poynting’s vector”. It tells us the rate at which the field energy moves around in space. To observe the effect of light aberration within an accelerated frame, it is essential to establish a coordinate system with a well-defined reference direction. We define the reference directions as those perpendicular to the direction of motion in both the inertial and accelerated frames. For simplicity, we assume that the motion of the aberrated light beam remains confined to a single plane, with its angular position described by a single angle. The analysis is much simplified if we treat separately the case of a radiated beam with its $\vec{E}$-vector parallel to the ”plane of incidence” (that is $xz$-plane) and the case of a radiated beam with the $\vec{E}$-vector perpendicular to the $xz$-plane. We will carry through the analysis for an incoming beam polarized perpendicular to the plane of incidence, but the principle is the same for both. So we take that $\vec{E}$, has only $y$- component. For a magnetic field, we get $\vec{B}=\vec{e}_{z}\times\vec{E}$, where $\vec{e}_{z}$ is the direction of Poynting vector $\vec{S}$ in the inertial frame.
Above, we demonstrated that the metric tensor must be matched in both the accelerational and inertial segments of the trajectory. To achieve this, it is sufficient to relate the coordinates and time of the accelerated observer to those of the inertial observer using a Galilean boost. The electrodynamics equations in an accelerated frame can be obtained by applying a Galilean transformation (with velocity $-v$) to Maxwell’s equations. The electromagnetic fields $\vec{E}_{n}$ and $\vec{B}_{n}$ observed in the accelerated frame differ from the fields $\vec{E}$ and $\vec{B}$ in the inertial frame prior to the active Galilean boost. The transformation of the fields under the Galilean boost is given by $\vec{E}_{n}=\vec{E}-\vec{v}\times\vec{B}/c$, $\vec{B}_{n}=\vec{B}-\vec{v}\times\vec{E}/c$. This expression accounts only for first-order effects in $v/c$. Here, the velocity vector of the accelerated frame relative to the inertial frame is $\vec{v}=v\vec{e}_{x}$. If the electric field $\vec{E}$ is perpendicular to the $xz$-plane, the magnetic field in the inertial frame has only an $x$-component and we have immediately $\vec{E}_{n}=\vec{E}=\vec{e}_{y}E$, $\vec{B}_{n}=(\vec{e}_{z}\times\vec{e}_{y})E-(\vec{e}_{x}\times\vec{e}_{y})vE/c=-\vec{e}_{x}E-\vec{e}_{z}vE/c$. This results in a Poynting vector, $c\vec{E}_{n}\times\vec{B}_{n}/(4\pi)=[-(\vec{e}_{y}\times\vec{e}_{x})-(\vec{e}_{y}\times\vec{e}_{z})v/c]cE^{2}/(4\pi)=[\vec{e}_{z}-\vec{e}_{x}(v/c)]cE^{2}/(4\pi)$. Thus, the radiated beam propagates at an angle $-v/c$ with respect to the $z_{n}$-axis, leading to the phenomenon of aberration of light in the accelerated frame.
We find that the energy transport direction remains the same whether analyzed using electrodynamics equations in an accelerated frame or the Galilean transformation law for electromagnetic fields. In the first approach, the equations show that a transmitted light beam with a finite transverse size propagates along the $x$-direction with a group velocity given by $d\omega/dk_{x}=-v$. In the second, light aberration is interpreted as a change in the direction of the Poynting vector. Both methods yield the same result. In this section, we demonstrate that the group velocity direction and the Poynting vector direction are fundamentally equivalent—two sides of the same coin.
This is a good point to make a general remark about the physical reality of the energy flow, represented by the Poynting vector. In any practical aberration measurement, one must account for a finite radiation beam size. The direction of energy transport—and thus the Poynting vector—is well-defined and convention-invariant. However, a common misconception in electrodynamics arises when considering energy transport in plane waves. In such cases, determining the exact direction of energy flow is fundamentally impossible.
5.5 Clock Resynchronization in Accelerated Systems
It should be noted, however, that there is an alternative satisfactory approach to explaining the effect of light aberration in the accelerated frame $S_{n}$. This explanation utilizes a clock re-synchronization procedure. When the system $S_{n}$ begins to move with constant velocity, the standard procedure for Einstein’s clock synchronization can be applied. This synchronization method is based on the assumption that light signals emitted by a source at rest propagate with the same velocity $c$ in all directions. By using this synchronization procedure, we can select a Lorentz coordinate system for the screen. In this framework, the transmission through the aperture is described using Maxwell’s equations. The interval in the accelerated reference frame $S_{n}$ will then take the diagonal form given by Eq. (1) for the transmitted light beam.
Figure 16: Aberration of light in the accelerated frame of reference $S_{n}$. In this frame, the plane wavefront undergoes rotation following the re-synchronization of the accelerated clock. According to Maxwell’s electrodynamics, the transmitted light beam propagates at an angle of $-v/c$, resulting in the aberration of light.
The time coordinate $t^{\prime}_{n}$ in the $S_{n}$ frame, under Einstein synchronization, is obtained by introducing a time offset of the form given in Eq. 13, where the velocity $v$ is replaced with $-v$: specifically $-x_{n}v/c^{2}$. Applying this offset yields the first-order approximation $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$.
This time adjustment effectively rotates the plane of simultaneity by an angle $-v/c$. Consequently, after re-synchronization, the radiation wavefront undergoes rotation in the accelerated frame, as illustrated in Fig. 16.
The new time coordinate in the accelerated frame is interpreted in accordance with Maxwell’s equations, ensuring their applicability to light propagation. As a result, the transmitted light beam (passing through the aperture) propagates at an angle $-v/c$, demonstrating the phenomenon of the aberration of light. Both approaches yield the same outcome. The choice between these clock synchronization conventions is ultimately pragmatic. While changing the (four-dimensional) coordinate system does not introduce new physical phenomena, it provides a more consistent and coherent description of existing ones.
We restricted our analysis to the diagonalization procedure for nonrelativistic velocities. However, the method can be readily generalized for an arbitrary parameter $v/c$. A full diagonalization can be achieved using the transformation: $t^{\prime}_{n}=t_{n}\sqrt{1-v^{2}/c^{2}}-(vx_{n}/c^{2})/\sqrt{1-v^{2}/c^{2}}$, $x^{\prime}_{n}=x_{n}/\sqrt{1-v^{2}/c^{2}}$. Here, the physical time $t^{\prime}_{n}$ determines the flow of time in a physical process in the accelerated frame $S_{n}$, while $dx^{\prime}_{n}$ denotes the physical distance between two points of three-dimensional space.
5.6 Composition of Motions in Non-Inertial Frames
Until now, we have considered only the case of nonrelativistic motion. We now extend our analysis to arbitrary velocities. As discussed earlier, the metric tensor must be a continuous quantity. The Langevin metric is derived by matching the metric tensors of the accelerated and inertial frames. For a smooth transition of the metric tensor, it is sufficient to relate the coordinates and time of the accelerated observer to those of the inertial observer using the inverse Galilean boost. Under this transformation, the inertial frame metric transitions into the Langevin metric of the accelerated frame. Next, we consider the transformation of the direction of light propagation. The group velocity of light transforms similarly to that of a particle, following the Galileo velocity addition theorem. Let us assume that the accelerated frame moves with velocity $v$, then the angle of aberration $\theta_{a}$ is such that   $\tan\theta_{a}=-v/c$. If $v/c\ll 1$ the angle of aberration is small, being approximated by $\theta_{a}=-v/c$.
It is always possible to choose such variables, in which the metric of an accelerated source will be diagonal. In the Lorentz coordinatization, light propagates with constant velocity $c$ independently of the direction of propagation, and of the velocity of its source. If we look at the transformation $t^{\prime}_{n}=t_{n}\sqrt{1-v^{2}/c^{2}}-(vx_{n}/c^{2})/\sqrt{1-v^{2}/c^{2}}$, $x^{\prime}_{n}=x_{n}/\sqrt{1-v^{2}/c^{2}}$, we see that velocity along the $x_{n}$-axis is the same in the Lorentz coordinatization too: $dx_{n}^{\prime}/dt_{n}^{\prime}=(dx_{n}/dt_{n})/[1-v^{2}/c^{2}-v(dx_{n}/dt_{n})]$. Substituting $dx_{n}/dt_{n}=-v$ into this equation confirms that $dx_{n}^{\prime}/dt_{n}^{\prime}=-v$. The group velocity of the light beam remains $-v$, but when we use the new velocity of light this becomes $\sin\theta_{a}=-v/c$. This represents a correction to the classical aberration of light law within Lorentzian coordinates. Now, suppose an accelerated observer measures the aberration of light. Since time and distance in Lorentzian coordinates retain direct physical significance, the observer finds that the aberration angle satisfies the relation $\sin\theta_{a}=-v/c$.
5.7 Discussion
Let us now return to the topic of metric diagonalization and deepen our understanding of the relationship between the Lorentz coordinate systems $(x,y,z,t)$ and $(x^{\prime}_{n},y^{\prime}_{n},z^{\prime}_{n},t^{\prime}_{n})$. We have already noted that the old coordinates $(x_{n},y_{n},z_{n},t_{n})$ are found by matching the accelerated frame and inertial frame metrics; what does mismatching of the coordinates $(x,y,z,t)$ and $(x^{\prime}_{n},y^{\prime}_{n},z^{\prime}_{n},t^{\prime}_{n})$ mean, in terms of measurements made by the accelerated observer? In both reference frames, the metric remains diagonal, and, according to standard textbooks, the coordinates should be related by the Lorentz transformation. At first glance, the problem appears entirely symmetrical, suggesting that both frames should be equivalent. However, as we demonstrated earlier, there is a fundamental distinction between an accelerated inertial frame and an inertial frame without a history of acceleration.
Where does the asymmetry originate? The electrodynamics equation must be integrated with initial conditions. According to Maxwell’s electrodynamics, coherent radiation is always emitted in the direction normal to the radiation wavefront. In the inertial frame, the wavefront of the emitted light beam is perpendicular to the vertical $z$-axis. However, when considering Lorentz coordinatization, the wavefront orientation is affected in the accelerated frame $S_{n}$. According to relativistic kinematics, the extra phase chirp $d\phi/dx^{\prime}_{n}=k_{x}=-v\omega_{i}/c^{2}$ is introduced as a consequence of this, the plane wavefront rotates after the acceleration. Then, the radiated light beam is propagated at the angle $-v/c$ with respect to the $z_{n}$-axis, yielding the phenomenon of the aberration of light in the accelerated frame $S_{n}$.
This phenomenon is purely kinematic and involves no forces. Many may find this counterintuitive. We explore the resolution of this paradox in detail in Chapter 12, using a dynamical approach to explain the asymmetry between inertial and accelerated reference frames. Without proof, we state the key result: the apparent asymmetry arises from the presence of inertial (pseudo-gravitational) forces in the accelerated frame $S_{n}$. The principle of equivalence allows non-inertial kinematic problems to be addressed using dynamical methods.
Wavefront rotation, which occurs when transitioning from an inertial to an accelerated frame (relative to the fixed stars), can be attributed to the pseudo-gravitational potential gradient during acceleration. A clock’s rate depends on the local pseudo-gravitational potential, leading to an accumulated time difference between spatially separated clocks.
Suppose (for simplicity) that there is a constant acceleration $g$ along the $x_{n}$-axis during the time $T=v/g$. The pseudo-gravitational acceleration is simply equal to the gradient of scalar potential: $g=-\partial\phi/\partial x_{n}$. A clock at a higher gravitational potential (aligned with the acceleration direction) runs faster. The pseudo-gravitational potential difference between two points along the $x_{n}$-axis is given by $\Delta\phi=\phi_{1}-\phi_{2}=(\partial\phi/\partial x_{n})[x_{n}(1)-x_{n}(2)]$. Once the fixed stars move with constant velocity, the potential gradient in $S_{n}$ becomes zero. The accumulated time difference between two spatially separated clocks is $t_{n}(1)-t_{n}(2)=-g[x_{n}(1)-x_{n}(2)]v/(c^{2}|g|)=v[x_{n}(1)-x_{n}(2)]/c^{2}$. This time shift effectively rotates the plane of simultaneity by an angle $-v/c$.
5.8 A Single Moving Emitter in an Accelerated Frame
In Chapter 4, we examined the single-emitter problem within the initial inertial frame of reference, $S$. Our analysis revealed a deviation in the energy transport of light emitted by an accelerated source, which corresponds to a well-known phenomenon: the aberration of light from a tangentially moving emitter in an inertial frame of reference (see Fig. 1). We demonstrated that this aberration can be analyzed using two distinct approaches: the covariant and the non-covariant formulations.
We now extend our analysis to the aberration of light emitted by a ”plane wave” source moving within an accelerated reference frame. In this case, two different methods of coordinatization are possible:
(a) Consider a system $S_{n}$, initially at rest in the inertial frame $S$, which undergoes acceleration along the $x$-axis until it reaches a velocity $v$. The simplest synchronization method, based on absolute time, involves maintaining the same set of uniformly synchronized clocks as when $S_{n}$ was at rest. However, acceleration relative to the fixed stars influences the propagation of light. As a result, by the time $S_{n}$ attains a constant velocity, the velocity of light emitted from a source at rest in the accelerated frame $S_{n}$ will exhibit anisotropy along the $x_{n}$-axis (see Eq. (15)).
Figure 17: Aberration of light in an accelerated frame of reference $S_{n}$. The emitter remains at rest relative to the fixed stars. The wavefront orientation and group velocity are shown using Einstein synchronization.
(b) Once the system $S_{n}$ begins moving with a constant velocity, Einstein’s standard clock synchronization procedure can be applied. This synchronization method is defined using light signals emitted by a source at rest in the accelerated frame, under the assumption that light propagates isotropically with velocity $c$. By employing this Einstein synchronization procedure, we describe the light emitted from the source at rest (in the accelerated frame) using the standard Maxwell equations.
We now consider the case where the emitter is at rest in the initial inertial frame $S$ (i.e., at rest relative to the fixed stars), while the observer, who is at rest in the accelerated frame $S_{n}$, performs the measurement of energy transport direction. It is important to emphasize that the aberration of light emitted by a single source also depends on the motion of the emitter within the accelerated frame $S_{n}$. Due to the fundamental asymmetry between inertial and accelerated frames, the theory of light aberration predicts a remarkable effect. Specifically, if the emitter remains at rest relative to the fixed stars while the observer transitions from rest to motion with respect to the fixed stars, the apparent angular position of the ”plane-wave” emitter, as seen in the accelerated frame, would jump by an angle of $-2v/c$. This aberration effect in the frame $S_{n}$ is illustrated in Fig. 17.
Let us describe the radiation emitted by a source moving with velocity $-v$ in an accelerated frame, as observed by an accelerated observer. The acceleration relative to the accelerated frame influences the propagation of light. Assuming an absolute time coordinatization, we apply a Galilean boost by substituting $x_{n}\to x_{n}+vt_{n}$ while keeping the time coordinate unchanged. Applying this transformation to the Langevin metric yields
$ds^{2}=c^{2}(1-4v^{2}/c^{2})dt_{n}^{2}-4vdx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}~{}.$ (18)
This metric characterizes the electrodynamics of a stationary light source in an inertial frame from the perspective of measurements made by an accelerated observer. From the formula, it is evident that the measurement of the electromagnetic field configuration in the moving system—expressed in terms of the coordinate $x_{n}$ and time $t_{n}$ of a measuring device at rest in the accelerated frame—yields the same result as that of the configuration at the point $x_{n}+vt_{n}$ and time $t_{n}=0$. A similar situation arises in the inertial frame. In Chapter 3, we analyzed the electrodynamics of a moving emitter using the metric given in Eq. (11). Within a single accelerated frame, during the inertial phase, the symmetry principle—namely, the equivalence of active and passive Galilean boosts—is upheld.
Let us now examine the fundamental asymmetry between inertial and accelerated frames in absolute time coordinatization. An inertial observer describes the electrodynamics of a stationary light source using the Minkowski metric. By substituting $x=x_{n}+vt$ and $t_{n}=t$ into the Minkowski metric, we obtain the Langevin metric (Eq. 15) in the comoving frame.
A natural question arises: what is the metric of the inertial observer from the perspective of the accelerated observer? To address this, we transform the coordinates $(t_{n},x_{n},y_{n},z_{n})$, which correspond to the accelerated observer $S_{n}$, using the inverse Galilean transformation. Specifically, we substitute $x_{n}=x-vt$ while keeping time unchanged ( $t_{n}=t$) into the Langevin metric (Eq. 15). This transformation yields the Minkowski metric, $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ which correctly describes the electrodynamics of a stationary light source in the inertial frame. Importantly, this transformation highlights how an inertial observer perceives the measurements of an accelerated observer. In particular, it illustrates the time dilation effect experienced by a physical clock at rest in the accelerated frame. Crucially, the slowdown of the accelerated clock is independent of the reference frame in which this effect is observed (see Chapter 12 for further details).
Another approach to solving the problem involves the Langevin metric diagonalization procedure. In this framework, the new coordinate system in the accelerated frame is interpreted such that an observer perceives a diagonal metric: $ds^{2}=c^{2}(dt^{\prime}_{n})^{2}-(dx_{n}^{\prime})^{2}-(dy_{n}^{\prime})^{2}-(dz_{n}^{\prime})^{2}$ for an emitter at rest in the accelerated frame, and $ds^{2}=c^{2}(1-v^{2}/c^{2})(dt^{\prime}_{n})^{2}-2vdx_{n}^{\prime}dt_{n}^{\prime}-(dx^{\prime}_{n})^{2}-(dy^{\prime}_{n})^{2}-(dz^{\prime}_{n})^{2}$ for a moving emitter. This reveals a symmetry (or reciprocity) in the metrics—and consequently in the electrodynamic equations—between the accelerated and inertial frames. This symmetry arises naturally from the pseudo-Euclidean geometry of space-time. The only difference between this metric and Eq. (11) is the sign of the velocity. After diagonalization, it becomes evident that any apparent asymmetry stems from the initial conditions. Notably, the plane radiation wavefront of a moving emitter (regardless of its velocity) undergoes a rotation by an angle of $-v/c$ in the accelerated frame as a result of the metric diagonalization.
To predict the outcome of aberration measurements, one can alternatively analyze the metric of the moving emitter directly within the absolute-time coordinate system, given by Eq. (18). According to electrodynamics, the group velocity $-2v$ emerges as a direct consequence of the cross term $-4vdxdt$.
Now, let us consider the most general case in which the system $S_{n}$, initially at rest in the inertial frame $S$, is accelerated to velocity $v$ along the $x$-axis. Simultaneously, an emitter in the accelerated frame $S_{n}$ is accelerated from rest to velocity $u$ along the $x_{n}$-axis. Suppose an observer in the accelerated frame $S_{n}$ performs an aberration measurement. To describe the radiation emitted by a source moving with velocity $u$ in the accelerated frame, we must account for the electrodynamics in terms of the accelerated observer. This can be done by employing the metric $ds^{2}=c^{2}[1-(v-u)^{2}/c^{2}]dt_{n}^{2}-2(v-u)dx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$. A closer examination of the physics reveals that the aberration increment is related to the system parameters by the expression $\theta_{a}=-v/c+u/c$. The key point is that the cross term in the metric, which induces anisotropy in the accelerated frame, depends solely on the velocity $v$ of the accelerated frame relative to the initial inertial frame.
Figure 18: Aberration of light in an accelerated frame of reference. The point source is at rest within the accelerated frame, while the wavefront orientation is shown in absolute time coordinatization for the screen.
Figure 19: Aberration of light in an accelerated frame of reference. The point source remains at rest relative to the fixed stars.
5.9 A Point Source in an Accelerated Frame
Now, let us return to observations of an accelerated observer. A key characteristic of point-like sources is that radiation emitted at a given instant forms a spherical wave around the source, and any measuring instrument inevitably influences the observed radiation (see Fig. 18 - Fig. 19). The aberration of a point source is independent of the source’s speed and is attributed solely to the observer’s motion relative to the fixed stars. The cross term in metric Eq. (15) introduces anisotropy in the accelerated frame, altering the direction of transmitted radiation. As a result, a point source exhibits the same aberration pattern as fixed stars, meaning its apparent position shifts by an angular displacement identical to that of the fixed stars.
In conventional aberration theory, an intriguing puzzle arises concerning stellar aberration. Some binary star systems undergo velocity changes on timescales ranging from days to years. At certain times, the components of such systems can have significantly different velocities relative to Earth. However, it is well established that both components always exhibit the same aberration angle. Rotating binary systems follow the same aberration pattern as all fixed stars, appearing to shift annually by the same universal aberration angle. This observation strongly supports our theoretical prediction for the aberration of light from a point source in a rotating frame. In the next chapter, we explore this experimental test in greater detail.
5.10 Aberration of a Tilted Incoming Plane Wave
We now return to the scenario depicted in Fig. 18 – Fig. 19. A point source located in the far zone effectively generates a plane wave in front of the aperture. Above, we considered only the case where the point source is positioned on the optical axis (which is parallel to the $z_{n}$-axis), producing a normally incident plane wave. The fundamental electrodynamic equations must hold for all electromagnetic waves. In other words, the dispersion equation in the accelerated frame should be consistently applied to both the incoming and scattered waves. In Maxwell’s electrodynamics, the dispersion relation simplifies to $k_{x}^{2}+k_{y}^{2}+k_{z}^{2}-\omega^{2}=0$. A notable feature of this case is that, even after a Galilean transformation along the $x$-axis, the dispersion relation in the accelerated frame retains the same diagonal form, $k_{z}^{2}-\omega_{i}^{2}=0$, for a normally incident wave.
Now, let us examine the behavior of an incoming wave with a phase gradient along $x_{n}$. In other words, there is an angle $\alpha$ between the wavevector $\vec{k}$ and the $z_{n}$-axis. For small angles $\alpha\simeq v/c$, we can approximate $\alpha\simeq k_{x}/k_{z}$. Our goal is to express the parameters of the incoming wave in terms of the source offset $x_{o}$ relative to the optical axis. The wave tilt angle can be written as $\alpha\simeq x_{o}/z_{s}$, where $z_{s}$ is the distance from the source to the screen. Next, we analyze the aberration of light emitted by a stationary point source in an accelerated frame. The wavevector of the incoming plane wave is determined by the initial conditions. Assuming that the emitted wavefront is initially tilted by an angle $\alpha$, we obtain $k_{x}=\alpha k=\alpha\omega_{i}/c$, $k_{z}=\sqrt{\omega_{i}^{2}/c^{2}-k_{x}^{2}}$. Since acceleration modifies the field equations, we substitute the plane wave solution $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ into Eq. (17), yielding: $(1-v^{2}/c^{2})k_{x}^{2}-2vk_{x}\omega/c^{2}+k_{z}^{2}-\omega^{2}/c^{2}=0$. A screen with a hole acts as a diffraction grating, splitting the incident plane wave into multiple diffracted components. The wavenumber of the transmitted plane wave along the $x_{n}$-direction is $k_{x}=\alpha k+K_{\perp}$, where $K_{\perp}$ is the wavenumber associated with the sinusoidally modulated transmittance. The dispersion equation then simplifies to $\omega_{i}\Delta\omega=-\omega_{i}K_{\perp}v+c^{2}\alpha kK_{\perp}$. This result indicates that the transmitted light beam propagates along $x_{n}$ with group velocity $d\omega/dk_{x}=-v+c\alpha$. Thus, the apparent angular position of the point source is related to the problem parameters by $\theta=-v/c+\alpha$  [^18] A similar effect arises if the screen moves with velocity $v$ along the $x$-axis in an inertial frame. In Chapter 4, we considered the special case of a normally incident plane wave, but this approach generalizes to tilted wavefronts as well. The deviation in energy transport for light passing through a hole in the moving screen leads to an aberration increment given by $\theta=v/c+\alpha$..
5.11 Non-Reciprocity in the Aberration of Light Theory
In this section, we continue our discussion on the aberration of light emitted by a single moving source. Now, let us consider a scenario with two identical emitters. Suppose the first emitter remains at rest in an inertial frame, while the second emitter is initially at rest but then accelerates to a velocity $v$ along the $x$-axis. An observer, who is at rest in the inertial frame, measures the direction of energy transport. Let us assume that, at the outset, the velocity component of the emitted light along the $x$-axis is zero. The question then arises: How does the light beam from the moving emitter appear? For the inertial observer, the angular displacement of the emitted radiation is given by $\theta_{a}=v/c$. This effect, illustrated in Fig. 20, is a well-known example of the phenomenon of aberration of light.
Figure 20: The aberration of light from stationary and moving sources.
Figure 21: Aberration of light when an accelerated observer redirects an accelerated emitter. After the redirection, an inertial observer measures an angular displacement of $\theta_{a}=2v/c$.
Let us analyze the scenario in which an accelerated observer redirects an accelerated emitter. We consider the specific case where, after redirection, the (group) velocity component of the light beam along the $x_{n}$-axis becomes zero. Following this redirection, an inertial observer would measure an angular displacement of $\theta_{a}=2v/c$. Now, suppose the accelerated observer also conducts an aberration measurement. As shown in Fig. 21, the aberration increment is related to the system’s parameters by the expression $\theta_{a}=-2v/c$. .
Figure 22: Illustration of the accelerated source redirection process. The source acceleration information is encoded in the point source offset relative to the optical axis, which is aligned with the $z_{n}$ axis. Notably, the wavefront orientation lacks a strict objective definition. Consequently, for simplicity, the effect of aberration is explained using two different coordinatizations: absolute time (left) and Lorentz coordinates (right).
Let us revisit the redirection procedure for an accelerated source. A key question arises: where is the information about the source’s acceleration recorded when redirecting light from an accelerated emitter? To understand the redirection of a light source in an accelerated system, we must examine the source’s internal mechanics and observe what occurs during the redirection process. Figure 13 illustrates a schematic of a ”plane wave” emitter. In electrodynamics, redirection can be achieved by introducing an offset between the point source and the optical axis (which is aligned with the $z_{n}$- axis), as shown in Figure 22. This offset encodes the information about the source’s acceleration.
5.12 Two Metrics
In special relativity, an important theorem states that applying a Galilean boost yields the metric of an accelerated emitter in the initial inertial frame, given by Eq. 11. In this framework, a non-accelerated light source is described using the Minkowski (diagonal) metric. To transform the coordinates $(t,x,y,z)$ of an inertial observer, we use the equivalence of active and passive pictures. By substituting $x\to x-vt,t\to t$ into the Minkowski metric, we obtain Eq. 11, which describes the electrodynamics of a moving emitter from the perspective of an inertial observer.
On the other hand, the Langevin metric describes the electrodynamics of an accelerated emitter from the viewpoint of an accelerated observer. A key principle is that the metric tensor must remain continuous. This problem in special relativity can be effectively addressed using an absolute time coordinatization, which relies on Galilean boosts. To transform the inertial observer’s coordinates to the accelerated frame, we apply the inverse Galilean transformation: substituting $x\to x_{n}+vt,t\to t_{n}$ into the Minkowski metric yields the Langevin metric, Eq. (15). Notably, the Langevin metric Eq.15 of the accelerated emitter in the accelerated frame is the same as metric Eq.11 of the accelerated emitter in the inertial frame. The difference is only that the sign of velocity $v$. These two metrics are two sides of the same coin. Ultimately, we describe the electrodynamics of the same accelerated emitter—accelerated with respect to the fixed stars—from both an inertial and an accelerated observer’s viewpoint.
Notably, this theorem is absent from conventional formulations of special relativity. Here, for the first time, we present the metric given in Eq. 11. It is surprising that this result has not previously been reported, especially considering that the Langevin metric—derived by matching metric tensors across frames using Galilean boosts—is already widely used in applications involving optical phenomena in rotating frames.
In the problem under discussion, a puzzle remains. We introduce a comoving coordinate system in the initial inertial frame and analyze radiation from the moving source in terms of the new (comoving) coordinate labels. After performing a passive transformation, the source coordinates transform as $x^{\prime}=x-vt,t^{\prime}=t$. According to the equivalence of active and passive transformations within a single inertial frame, Maxwell’s equations remain valid in the comoving coordinate system. Consequently, the electric field $\vec{E}^{\prime}$ of an electromagnetic wave satisfies the equation $\Box^{\prime 2}\vec{E}^{\prime}=\nabla^{\prime 2}\vec{E}^{\prime}-\partial^{2}\vec{E}^{\prime}/\partial(ct^{\prime})^{2}=0$. This implies that the Minkowski metric, $ds^{2}=c^{2}(dt^{\prime})^{2}-(dx^{\prime})^{2}-(dy^{\prime})^{2}-(dz^{\prime})^{2}$, governs the electrodynamics of the accelerated light source in the comoving coordinate system. However, this raises a fundamental question: How can the Minkowski metric applies in the comoving coordinatization in the inertial frame while the Langevin metric applies in the accelerated frame? This apparent discrepancy requires further explanation.
The difference between measurements made in the two reference frames is particularly intriguing. To analyze this, we first examine the relationship between passive and active transformations. It is important to note that a passive Galilean boost within a single inertial frame, given by $x^{\prime}=x-vt,t^{\prime}=t$, merely represents an alternative parametrization of the inertial observer’s measurements. Eq. (11) describes these measurements. In the case of active transformations within a single inertial frame, the motion of fixed stars relative to the inertial observer and their instruments remains unchanged. This establishes the equivalence between active and passive Galilean boosts. By contrast, the Langevin metric describes the measurements of an accelerated observer. Under the inverse Galilean boost, the motion of fixed stars relative to the space-time grid $(t_{n},x_{n},y_{n},z_{n})$ of the accelerated observer is altered.
On the other hand, the standard textbook treatment of Lorentz transformations implicitly assumes that the $x_{n}y_{n}z_{n}$-axes of the accelerated observer remain parallel to the $xyz$-axes of the inertial observer. In other words, it presumes that both observers share a common three-dimensional space. This assumption, often taken for granted, suggests that the simultaneous acceleration of a rigid reference frame has direct physical significance. However, this is a misconception. The conventional derivation of the aberration of light effect overlooks the conventionality of distant simultaneity. There is an inherent uncertainty in the relative position along the $z_{n}$-axis in the $x$-direction, amounting to $z_{n}v/c$, due to the ambiguity in the timing of acceleration. Consequently, we are dealing with space-like separated events. The orientation of the $z_{n}$-axis lacks an exact objective meaning because of the relativity of simultaneity. The claim that the $z_{n}$-axis orientation with respect to the $z$-axis of the inertial observer has objective meaning to within a certain accuracy can be visualized by the picture of $z_{n}$ axis in the vertical orientation with approximate angle uncertainty (blurring) given by $\Delta\theta=v/c$. This inclination angle between the inertial $z$ and accelerated $z_{n}$ reference directions has no exact objective meaning at all since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this displacement could be ascertained. We will explore this issue further in Chapter 7.
5.13 Ether Theory and the Aberration of Light Effect
The framework of classical physics is rooted in the concept of absolute space, often linked to the hypothetical ”ether”—a substance once thought to provide an absolute frame of rest. In this section, we will show that the initial inertial frame discussed earlier is not physically equivalent to absolute space or the ether. Instead, it represents an abstract idealization: the concept of absolute space-time.
To deepen our analysis, we revisit the classical ether theory in its original form, which posits the existence of a stationary ether that determines the speed of light. Since the ether is presumed immobile, any frame moving relative to it would experience anisotropy in light propagation. A non-inertial frame, accelerating with respect to the ether, would thus detect an ”ether wind” when measuring the speed of light.
We begin by examining the influence of the ether wind on light speed. According to the ether hypothesis, the velocity of light observed from an accelerated frame would be $c+u$ for a beam propagating in the same direction as the accelerated system and $-c+u$ for a beam propagating in the opposite direction, where $u=-v$ represents the ether velocity in the accelerated reference frame. The aberration effect arises as a consequence of the ether wind’s influence on the speed of light (Fig. 23).
Figure 23: Aberration of light in an accelerated frame. According to the ether theory, the ether wind induces anisotropy in a frame accelerated relative to the ether, which causes a change in the direction of radiation.
5.13.1 Lorentz’s Theory of Corresponding States
In 1895, Hendrik Lorentz derived a first-order version of the so-called ”theorem of corresponding states” to provide a general explanation for the null results of first-order optical interference experiments. This theorem establishes a relationship between pairs of electromagnetic field configurations: one in a system at rest in the ether and the other in uniform motion through it. Lorentz demonstrated that if a particular field configuration is allowed by the laws of electrodynamics, its corresponding configuration in the moving system must also be valid. The theorem works as follows. The mathematical description of the configuration in the moving system in terms of ”local time” (instead of ”real time” ) is the same as the description of its corresponding state in the system at rest in terms of real time. Lorentz showed that two field configurations related to one another in many of their observable properties, in particular, they give rise to identical interference patterns. Thus, it is not possible to detect the orbital velocity using interference (e.g. Michelson-Morley) experiments  [^19] For discussion of the relationship between the generally accepted way of looking at special relativity and Lorentz’s ether theory we suggest reading the paper [32]..
We now examine the application of Lorentz’s pre-relativistic theory to the aberration of light in a non-inertial frame of reference. Previously, we analyzed the aberration of light in accelerated systems within the framework of special relativity, considering only first-order effects. Since light is inherently a relativistic phenomenon, its behavior in non-inertial frames requires careful treatment.
In this chapter, we have already discussed the asymmetry between inertial and accelerated frames. A key observation was that the metric tensor must remain continuous. In special relativity, this issue is best addressed through an absolute time coordinatization approach. Maxwell’s equations hold in the initial inertial frame, where the optical system (e.g., a light source and lens, as shown in Fig. 13) is at rest, and time coordinates are assigned via slow clock transport. To ensure continuity, the metric of a non-inertial frame associated with an accelerating optical system (relative to the initial inertial frame) must transition smoothly into the Langevin metric. This requires relating the accelerated observer’s coordinates $(x_{n},y_{n},z_{n},t_{n})$ to those of the inertial observer $(x,y,z,t)$ using the inverse Galilean transformation: $x_{n}=x-vt,y_{n}=y,z_{n}=z$, $t_{n}=t$. Under this transformation, the wave equation takes the form of Eq. 17, where the coordinates and time adhere to Galilean transformation rules. The Langevin metric, along with the initial conditions, then governs optical effects in the accelerated frame.
In the preceding discussion, we analyzed the aberration of light in an accelerated frame using the Fourier transform method. The fundamental asymmetry arises from the electrodynamics equation (Eq. 17). One consequence of the cross-term in the wave equation is the emergence of group velocity, which contributes to the aberration. However, an alternative approach provides an equally satisfactory explanation. This involves clock resynchronization, wherein Einstein’s synchronization method is used to define a Lorentz coordinate system for the optical setup at rest in the accelerated frame. In this framework, Maxwell’s equations describe the field configuration, and the resynchronized time $t^{\prime}_{n}$ is given by $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$ in the first-order approximation. After resynchronization, the asymmetry shifts to the initial condition: the plane wavefront of the emitter undergoes a rotation by an angle $-v/c$ in the accelerated frame.
At first glance, Lorentz’s treatment of an optical setup in an inertial frame moving through the ether appears similar to our approach to such a setup in an accelerated frame (relative to the fixed stars). We now analyze this similarity. One key observation is that our initial inertial frame—one without a history of acceleration relative to the fixed stars—can be interpreted as what Lorentz considered a frame at rest in the ether. In Lorentz’s theory, and more broadly in pre-special relativity physics, the relationship between an observer at rest (using coordinates $(x,y,z,t)$)) and an observer bound to a moving body (using $(x_{n},y_{n},z_{n},t)$ ) was governed by the Galilean transformation: $x_{n}=x-vt,y_{n}=y,z_{n}=z,t_{n}=t$. In Lorentz’s framework, a Galilean boost is empirically equivalent to a smooth adjustment of the metric tensor. Under this transformation, the metric of the initial inertial frame transitions into the Langevin metric of the accelerated frame. Of course, Lorentz did not conceptualize his theory in this way, as Minkowski spacetime was not yet available to him.
Due to this Galilean boost, the homogeneous wave equation for the field in the moving frame (Eq. 17) lacks a standard form. The primary complication arises from the presence of a cross-term, which makes solving the equation more challenging. To circumvent this difficulty, Lorentz observed that simplification is always possible. In solving the wave equation, he introduced the variable $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$, which he called ”local time.” In Lorentz’s theory, this time shift was a mathematical device used to transform the electrodynamics equations into Maxwell’s form.
At first sight, Lorentz’s theorem of corresponding states for first-order phenomena seems sufficient to provide an exact treatment of light aberration in an accelerated frame. One might naively expect that Lorentz’s theory would predict, in full generality, that observations of light aberration from an earth-based incoherent source should reveal Earth’s orbital motion (Fig. 22). However, this leads to an important question: Why did Lorentz conclude, based on his theorem of corresponding states, that no first-order effects of Earth’s orbital motion could be detected?
5.13.2 Analogy with the Sound Theory
A comparison between Lorentz ether theory and the theory of sound may be insightful in this context. In Section 4.15, we analyzed the problem of a sound emitter in the atmosphere frame, specifically considering a screen moving along its surface with velocity $v$. According to the theory of sound, energy transport remains unaffected when sound passes through a hole in the moving screen. When a plane wave falls the screen normally, it generates a transmitted oblique sound beam, as illustrated in Fig. 7. In this scenario, the group velocity matches the phase velocity. The diagonal wave equation, $\nabla^{2}f-\partial^{2}f/\partial(v_{s}t)^{2}=0$, is always valid in the atmosphere frame. However, this equation is not applicable from the perspective of an observer at rest in an accelerated frame. Accelerations relative to the atmosphere frame influence the propagation of sound, leading to a transformation of the wave equation into
$\left(1-\frac{v^{2}}{v_{s}^{2}}\right)\frac{\partial^{2}f}{\partial x_{n}^{2}}+2\left(\frac{v}{v_{s}}\right)\frac{\partial^{2}f}{\partial t\partial x_{n}}+\frac{\partial^{2}f}{\partial y_{n}^{2}}+\frac{\partial^{2}f}{\partial z_{n}^{2}}-\frac{1}{c^{2}}\frac{\partial^{2}f}{\partial t^{2}}=0~{},$ (19)
where coordinates and time are transformed according to the Galilean transformation $x_{n}=x-vt$, $t_{n}=t$. To describe sound propagation, the wave equation must be integrated with the initial conditions of the sound wavefront. Notably, acceleration does not influence the wavefront’s orientation, meaning the emitted sound beam remains perpendicular to the $z_{n}$ axis. Since the air remains stationary in the atmospheric frame, any reference frame moving relative to it experiences anisotropy in sound propagation. As a result, an observer in such a moving frame perceives an ”air wind” when measuring the speed of sound. According to the wave equation (Eq. 19), the sound velocity observed from an accelerated reference frame is $v_{s}-v$ for a beam propagating in the same direction as the accelerated system and $-v_{s}-v$ for a beam propagating in the opposite direction. Here, $-v$ represents the wind velocity in the accelerated frame. This effect, known as aberration, arises due to the influence of the perceived air wind on the speed of sound (Fig. 23).
Let us now analyze the aberration of sound radiated by a single ”plane-wave” emitter in the atmospheric frame. The wavefront radiated by the plane-wave emitter at a given instant time is constructed as an envelope of all the spherical wavelets from the point sources on the emitter that have radiated until this instant. It is not difficult to see that when the emitter is moving at a constant velocity $v$ along the $x$ axis, the wavefront propagates at a speed $v_{s}$ at an angle $\theta=v/v_{s}$ from the vertical. This phenomenon arises due to the motion of the emitter, which causes constructive interference to occur when the projection of the phase velocity onto the $x$-axis matches the velocity of the emitter. According to classical kinematics, this results in an additional phase chirp given by $d\phi/dx=k_{x}=vk/v_{s}=v\omega_{s}/v_{s}^{2}$. Consequently, the plane wavefront undergoes rotation as the emitter accelerates. As a result, the radiated sound beam propagates at an angle $v/v_{s}$, manifesting as the aberration of sound in the atmospheric frame—analogous to the aberration of light.
Now, consider the perspective of an accelerated observer. The wave equation must be integrated with the initial condition set by the sound wavefront. A Galilean boost does not affect the orientation of the wavefront, meaning that the wavefront of the emitted sound beam remains tilted at an angle $v/v_{s}$ relative to the $z_{n}$-axis. The wave vector of the radiated plane wave is determined by this initial condition, leading to $k_{x}=vk/v_{s}=v\omega_{s}/v_{s}^{2}$. Furthermore, the non-diagonal wave equation (Eq. 19) implies that the radiated sound beam moves along the $x_{n}$-axis with a group velocity given by $d\omega/dk_{x}=-v+v_{s}^{2}k_{x}/\omega_{s}=0$. For further details, refer to Section 5.10.
Thus, we derive the same addition theorem for sound beam velocities as for particle velocities in classical physics. According to Lorentz, the wave equation for a moving light (or sound) emitter in the ether (or atmospheric) frame is identical to that of an emitter at rest. Lorentz concluded that Earth’s orbital motion could not be determined using aberration measurements from Earth-based sources. Consequently, ether and sound theories were merely variations of the same classical framework. However, this analysis overlooks a fundamental distinction: the velocity of light differs inherently from the velocity of sound.
When Lorentz formulated his pre-relativistic theory he assumed a space-time background with a well-defined geometrical structure: Newtonian space-time. A close look at the physics of this subject shows that there is a real difference between the inertial frame without an accelerational history relative to the fixed stars and the Lorentz classical ether at rest. The main difference between the initial inertial frame and the absolute space of Newton is that special relativity taught us to think in terms of a unified space-time model. In special relativity, the absolute space and the absolute time of Newton are fused into absolute space-time.
6 Stellar Aberration
6.1 The Corpuscular Model of Light and Stellar Aberration
It is commonly understood that the aberration of light can be interpreted through the corpuscular model of light, drawing an analogy to how a moving observer perceives the oblique fall of raindrops. This classical kinematic approach to the calculation of stellar aberration has been employed in astronomy for nearly three centuries  [^20] The phenomenon of the annual apparent motion of celestial objects, known as stellar aberration, was first discovered by Bradley in 1727, who also explained it using the corpuscular model of light [33]..
However, in the 20th century, it was discovered that Newton’s dynamical laws for light were incorrect, necessitating the introduction of electromagnetic wave theory to rectify these errors. As is well-known in textbooks, the physical basis of stellar aberration arises from the finite speed of light and the change in its direction when observed from a different reference frame. This can be explained as a consequence of the velocity addition formula applied to a light beam when the observer shifts between reference frames.
For an observer on Earth, the velocity relative to the solar reference frame is approximately $v=30~{}\mathrm{km/s}$, corresponding to Earth’s motion around the Sun. To describe the effect of stellar aberration, it is sufficient to consider only the first-order approximation, $v/c=10^{-4}$. The conventional treatment of stellar aberration remains deeply connected to classical (Newtonian) kinematics, relying on the erroneous assumption that observers on Earth and the Sun share a common 3-space.
6.2 The Wave Theory of Stellar Aberration
Most authors treat light propagating through the telescope barrel as a stream of photons rather than as a wave.  [^21] A common misconception is that the rays of light coming from a star simply fall onto the telescope tube without interacting with its sides. As French [4] states: ”Regarding light as composed of a rain of photons, we can easily calculate the change in the apparent direction of a distant object such as a star.” Similar views are also found in recent textbooks. For instance, Rafelski [5] writes: ”We consider a light ray originating from a distant star… The following discussion addresses the observation of well-focused light rays, not (spherical) plane wave light… This view of the experimental situation is accurate because the light emitted by a star consists of an incoherent flux of photons produced in independent atomic processes.” This conclusion, however, is incorrect. We emphasize that the design of a telescope’s optical system is based on classical diffraction theory, which stems directly from classical electromagnetic theory. The resolution of a telescope is fundamentally limited by the diffraction of light waves, and the telescope always affects the measured radiation due to the unavoidable diffraction of starlight by the telescope aperture. In a well-corrected optical system with a circular pupil, the size of the Airy disk (i.e., the size of the image of a point source) is inversely proportional to the pupil’s diameter.
By questioning the validity of this standard reasoning, we argue that a proper treatment of stellar aberration should rely on coherent wave optics. It is easy to show that light from a distant star is approximately coherent over a circular area whose diameter, in practical terms, is far larger than the diameter of the telescope. Thus, the telescope samples only a tiny portion of the star’s coherent light, meaning the waveforms can be approximated as (flat) plane waves.  [^22] While stars are considered incoherent sources, the mutual intensity function produced by an incoherent source is fully described by the Van Cittert-Zernike theorem [34]. Any star can be considered as very distant from the Sun, and telescopes typically operate in the far zone of such sources. For instance, consider Sirius, one of the closest stars to Earth. The coherent area of light from Sirius has a diameter of about 6 meters, a correlation observed by Brown and Twiss in 1956 [35].
Let’s verify whether this assertion holds true. The spatial coherence of a light beam generally refers to the coherence between two points in the field illuminated by the light source. To understand spatial coherence more clearly, we can refer to Young’s two-pinhole experiment (see Fig. 24). In its simplest form, the degree of coherence between two points is described by the contrast of the interference fringes obtained when these points are treated as secondary sources. Consider a source $S$ illuminating two pinholes, $S1$ and $S2$, as shown in Fig. 24. If the source is perfectly incoherent, no interference fringes can be observed by placing two pinholes in the plane of the source. However, it has been demonstrated that when the two pinholes are positioned sufficiently far from the incoherent source, interference fringes with good contrast can be observed. It is often stated that the spatial coherence of light beams increases with distance ”simply through propagation.” It would be valuable to find an elementary explanation that allows us to visualize the physical process behind this phenomenon.
Consider a quasi-monochromatic wave incident on an aperture in an opaque screen, as depicted in Fig. 24. In general, this wave may exhibit partial coherence. As the wave propagates through space, its detailed structure evolves, and similarly, the spatial coherence structure also changes. In this context, the transverse coherence function is said to ”propagate.” Given the spatial coherence at the aperture, our goal is to determine the spatial coherence on the observation screen, which is located at a distance $z$ beyond the aperture. Stellar radiation is inherently stochastic, and for any starlight beam, there exists a characteristic linear dimension, $\Delta r$, which defines the scale of spatially random fluctuations. Fig. 25 illustrates the spiky pattern that appears on the aperture in an opaque screen. When $\Delta r\ll d$, where $d$ is the aperture size, the radiation beyond the aperture remains partially coherent, as shown in Fig. 25. In this case, $\Delta r$ corresponds to the typical linear dimension of the spikes.
Figure 24: Young’s interferometer.
Figure 25: Geometry for propagation of spatial coherence.
Figure 26: Reciprocal width relation of Fourier transform pair. Free space basically acts as a Fourier transformation. This means that the field in the far zone is, phase factor and proportionality factor aside, the spatial Fourier transform of the field at the source plane.
[…]
A star is an incoherent source, characterized by spatial fluctuations on the scale of $\lambdabar\Delta r\sim\lambdabar$. Such a source radiates in all directions. Additionally, the angular spectrum of the starlight contains spikes with a typical width of $(kd)^{-1}$, where $d$ represents the star’s diameter. In astronomical observations using a telescope, each spike observed on the Earth’s surface has a finite thickness, approximately equal to the spatial coherence length of the starlight. Therefore, the coherent area of the observed starlight on the Earth’s surface can be estimated to be of the order of $z/(kd)$, where $z$ is the distance from the star to the Earth. The star chosen by Bradley for this observation was Draconis, one of the nearest stars. For this case, the diameter of the coherent area on the Earth’s surface is roughly estimated to be around 100 meters.
6.3 Stellar aberration in the Context of Special Relativity
It is widely accepted that the theory of relativity aligns with Bradley’s discovery of stellar aberration, interpreting it as a consequence of an observer’s motion relative to a light source 232323 In his paper on relativity, Einstein derived the aberration formula by considering the velocity $v$ as the relative velocity of the star-Earth system. This interpretation has been adopted by numerous textbook authors. As Moeller [13] states: ”This phenomenon, known as aberration, was observed by Bradley, who noticed that stars appear to undergo an annual collective motion in the sky. This apparent motion arises because the observed direction of a light ray from a star depends on the Earth’s velocity relative to the star.”.
The asymmetry between the cases where either the star or the Earth-based telescope is in motion becomes evident when considering the separation of binary stars. Spectroscopic binaries exhibit velocities exceeding that of Earth’s orbit around the Sun. These stars revolve around their common center of mass within just a few days—a period during which Earth’s motion remains practically constant. If the changing velocities of the binary components were comparable to Earth’s orbital velocity, their separation should be easily detectable. However, this is not observed (Fig. 27-Fig. 28)  [^24] In 1950, Ives [36] was the first to highlight a significant challenge that binary star observations pose to the theory of relativity. He argued that the idea of aberration being purely a function of relative motion is contradicted by the existence of spectroscopic binaries with velocities comparable to Earth’s orbital velocity. Yet, these binaries exhibit no aberration effects different from other stars. For instance, the spectroscopic binary Mizar A has well-established orbital parameters. If aberration were solely due to relative velocity, its angular separation should be observable at 1’10”. However, empirical measurements show a value below 0.01”, which is clearly inconsistent with the perspective presented in standard textbooks [37, 38]..
Figure 27: Aberration shift inferred from astronomical observations. The plane wavefronts of starlight entering the telescope are focused by the lens to form a diffraction spot in the focal plane. Two scenarios are considered, each with different velocities of an Earth-based telescope and a star. In the first scenario (left), the star remains stationary relative to the fixed stars. In the second scenario (right), the star moves at the same velocity as the Earth.
Figure 28: Predicted aberration shift according to the conventional theory of stellar aberration.
In Chapter 4, we critically reexamined the textbook claim that wavefronts and raindrops experience the same aberration. Using the theory of relativity, we demonstrated that when a transversely moving mirror reflects a normally incident plane wave of light, the energy transport deviates. This effect arises because, in the absolute time coordinatization, the wave equation depends on the velocity vector. As a result, the energy transport velocity differs from the phase velocity. According to Babinet’s principle, this remarkable prediction should also apply to light transmitted through a hole in a moving opaque screen or, consequently, through the moving open end of a telescope barrel.
The binary star paradox is resolved by recognizing that when light passes through the end of a telescope barrel, diffraction perturbs its fields. As a result, the light beam no longer carries information about the star’s motion relative to the fixed stars. If the telescope were at rest relative to the fixed stars and the observed star began moving, its apparent position in the telescope would not suddenly shift by any angle.
However, a different challenge arises when explaining Earth-based observations—specifically, the apparent shift in the positions of fixed stars as the Earth’s telescope changes its motion relative to them. It is crucial to emphasize that aberration is caused by changes in the telescope’s velocity, not the star’s.
The resolution of this issue lies in the fundamental asymmetry between Earth-based and Sun-based observers—namely, the acceleration of the Earth-based observer relative to the fixed stars. Some experts, such as Selleri, have explicitly recognized this, noting that aberration can be explained in terms of variations in the Earth’s absolute velocity due to its orbital motion, while the star–Earth relative velocity is irrelevant. Thus, planetary acceleration plays a key role in the phenomenon [38]. We derive stellar aberration using the Langevin metric. Although stellar aberration exhibits asymmetry, it does not contradict special relativity. This is because the heliocentric (Sun-based) reference frame is inertial, whereas the geocentric (Earth-based) reference frame is non-inertial.
In the discussion above, we demonstrated that when a small hole is present in an opaque screen at rest relative to the fixed stars, and a point source moves tangentially to the screen, there is no aberration—no deviation in the transport of energy—for light passing through the hole. The transmission appears as illustrated in Fig. 12. The absence of any effect from the motion of the light source relative to the fixed stars in this setup suggests a fundamental issue for stellar aberration theory in the heliocentric frame of reference.
Now, consider an observer at rest with respect to the telescope who measures the direction of energy transport. A careful analysis reveals that in the heliocentric frame, where the telescope is at rest, we are dealing with a steady-state transmission problem. How does the transmitted light beam behave in this case? The beam appears to travel along the telescope’s axis because it has lost its horizontal group velocity component. This situation corresponds to a telescope aligned perpendicularly to the phase front of incoming starlight—pointed directly at the star. If the star’s motion is parallel to the phase front (i.e., perpendicular to the telescope axis), then starlight entering the telescope’s aperture will propagate through its full length without deviation.
A key conclusion from this discussion is that stellar aberration is absent in this configuration. Specifically, binary star components remain unresolved, meaning their velocities do not affect the aberration phenomenon.
The analogy between the obliquity of raindrops and stellar aberration is incorrect. A complete description of all Earth-based experimental observations of stellar aberration is possible only through the theory of relativity and wave optics. Our theory predicts stellar aberration effects that are in full agreement with Bradley’s results (Fig. 29).
Due to the asymmetry between inertial and rotating frames, the theory of light aberration makes a striking prediction: if a telescope is at rest relative to the Earth while the Earth rotates relative to the fixed stars, the observed direction of a star from Earth differs from that seen by a hypothetical observer at rest with respect to the Sun. Specifically, the apparent angle is smaller than the actual angle. The difference between the actual and apparent angles, $\theta_{a}$, is given by the relation: $\theta_{a}=v/c$, where $v$ is the Earth’s orbital velocity around the Sun. This effect arises from the anisotropy introduced by the cross-term in the metric equation (Eq. 15), which influences the direction of radiation (aberration) in the rotating frame. The transmission through the telescope aperture in this frame is illustrated in Fig. 18 - Fig. 19.
Figure 29: The direction of a star as seen from the earth.
In the geocentric frame of reference, stellar aberration is independent of the star’s speed and originates solely from the observer’s motion relative to the Sun. According to our interpretation, there are multiple types of aberration, with stellar aberration in the Earth-based frame being just one of them. We have demonstrated that all Earth-based experiments can be explained by two factors: the effect of the measuring instrument (i.e., the physical influence of the telescope on the measurement) and the acceleration of the Earth-based observer relative to the fixed stars.
6.4 Defining a Physical Coordinate System in Space
The aberration of light is a geometric phenomenon. To detect its effects within an Earth-based frame, a well-defined coordinate system with a reference direction is necessary.
A conventional approach to the aberration of light relies on an implicit assumption of a reference direction, though this assumption often goes unrecognized. Traditionally, the physical interpretation of aberration is based on measurements using rods that remain at rest in the observer’s frame. This local frame of reference is so intuitively accepted that it is rarely questioned or explicitly discussed in aberration theory.
However, a more fundamental spatial description exists beyond the reliance on measuring rods. In an Earth-based frame, the natural reference axis is defined by the gravitational field vector. The plumb line, pointing toward the Earth’s center, establishes the nadir, forming the most fundamental local coordinate system. For instance, James Bradley used a vertically mounted telescope to study stellar aberration. He selected the star Draconis because it transited almost exactly at the zenith. The traditional plumb line provided a sufficiently accurate zenith reference for observing stellar aberration.
6.5 Limits of Applicability
The applicability of our stellar aberration theory is broader than one might expect. Previously, for simplicity, we assumed that a star image is represented by a point spread function in the image plane of a telescope—in other words, that the input signal is effectively a plane wave.
Now, suppose a telescope can resolve details in the star image. A complete understanding of the object-image relationship requires accounting for diffraction effects. Diffraction convolves the ideal image with the Fraunhofer diffraction pattern of the telescope pupil.
A star is a spatially incoherent source  [^25] In statistical optics, stellar emission is described classically at the level of Maxwell’s equations, while the emitting medium—an ensemble of atoms—is treated using quantum mechanics. A star consists of elementary point sources, each statistically independent, with characteristic dimensions on the order of the radiation wavelength, $\lambda$. Within an elementary source volume $(\lambda^{3})$, a vast number of atoms contribute to emission. Semi-classical theory models these atoms as coherent, interacting radiating dipoles. The induced macroscopic dipole moment in an elementary source produces classical electromagnetic radiation. In statistical optics, electromagnetic fields are treated classically until they interact with the atoms of a photosensitive material, where their interaction is quantized. This avoids the need to quantize the electromagnetic field itself—only the field-matter interaction is quantized. The photodetector converts the continuous cycle-averaged classical intensity into a sequence of discrete photocounts.. Because a star is composed of statistically independent elementary point sources at different spatial offsets, each elementary source effectively produces a plane wave at the telescope pupil. An offset in an elementary source tilts the far-zone field. The total radiation field from a completely incoherent source is thus a linear superposition of the fields from these individual point sources. Since the image of each elementary source is a point spread function, the telescope inevitably influences the measurement of any completely incoherent source.
Importantly, any linear superposition of radiation fields from elementary point sources preserves single-source characteristics, such as independence from source motion. This reasoning confirms that our stellar aberration theory remains valid for imaging any completely incoherent source.
7 The Concept of Ordinary Space in Special Relativity
7.1 Inertial frame view of Observations of a Non-Inertial Observer
In this chapter, we introduce a novel approach to the theory of light aberration in non-inertial reference frames, providing an alternative perspective on this complex issue. First, we highlight a fundamental principle: the laws of physics in initial inertial frame must account for all physical phenomena, including observations made by non-inertial observers. It is commonly believed that phenomena in non-inertial (e.g., rotating) reference frames should be analyzed solely within the geometric space-time framework using the Langevin metric. However, we argue that physical phenomena in non-inertial frames can also be effectively described within an inertial frame using standard Einsteinian special relativity, leveraging well-established relativistic kinematic effects.
For example, when observed from an inertial laboratory frame, the interpretation of the Sagnac effect is straightforward. The phase difference between counter-propagating waves, attributed to the Sagnac effect, can be derived from the relativistic velocity addition formula. In other words, within the rotating frame, as seen from the inertial lab frame, the Sagnac effect manifests as a purely kinematic phenomenon governed by Einstein’s velocity addition in special relativity. This analysis assumes that the speed of light in the lab frame is $c$, which, according to special relativity, remains independent of the motion of the source. In contrast, calculating the Sagnac effect within a rotating reference frame is more complex. In such cases, textbook authors often employ a metric tensor—specifically, the Langevin metric—within a four-dimensional Minkowski spacetime to determine the propagation time difference between counter-propagating waves. Could we not analyze the aberration of light in the same manner? Here, too, we apply the Langevin metric in a non-inertial reference frame.
Let us first apply the classical kinematic method to compute the aberration of light effect in a non-inertial frame, as observed from an inertial frame. Textbook authors derive the same velocity addition theorem for light beams as for particles in classical physics. This result is particularly striking, as the study of light aberration is deeply rooted in classical kinematics. When terms of order $v^{2}/c^{2}$ are neglected, the Galilean vectorial velocity addition law is used. However, applying classical kinematics to aberration calculations leads to a significant error. According to our approach, a remarkable prediction emerges regarding light aberration in an accelerated frame (see Fig. 20). To correctly predict the aberration measurement, the accelerated observer must use the Langevin metric. The cross-term in this metric induces anisotropy in the accelerated frame, which alters the radiation direction. In contrast, relying solely on velocity addition (whether Galilean or Einsteinian) suggests that acceleration does not break the symmetry between the accelerated and initial inertial frames.
The error in the last argument arises from the incorrect assumption that an inertial observer and an accelerated observer share a common 3-space. According to special relativity, an inertial observer and an accelerated observer (relative to the fixed stars) have distinct 3-spaces. The classical kinematics method used to compute the aberration of light effect is flawed because it treats the speed of light as a classical velocity, disregarding relativistic effects. In the classical approach, no fundamental distinction is made between the aberration of light and the aberration of raindrops.
It is particularly interesting to observe that geometric effects in our ordinary space are closely linked to the relativity of simultaneity. In Fig. 16, the transmitted light beam propagates at an angle of $-v/c$, resulting in the phenomenon of light aberration. This naturally raises the question: relative to what does a light beam propagate within an accelerated frame when it undergoes an angular displacement of $-v/c$? Consider an observer in the accelerated frame measuring the direction of the light beam. If the plane wavefront of the transmitted beam is focused by a lens, it will form a diffraction spot in the focal plane along the optical axis. Determining the direction of the optical axis with respect to the frame’s axes is equivalent to measuring the angular displacement. To detect the aberration of light within the accelerated frame, it is essential to establish a coordinate system with a well-defined reference direction. This necessitates a detailed examination of the method by which coordinates are assigned—a process that inherently involves a physical procedure. Ultimately, this method must provide a consistent framework for defining coordinates in both inertial and accelerated reference frames.
In ordinary space, an accelerated frame moves relative to an inertial frame along the line of motion, and conversely, the inertial frame moves relative to the accelerated frame along the line of motion. The angle between the observer’s coordinate system axis and the line of motion is a fundamental geometric parameter in ordinary space. By using the line of motion as the reference axis, the accelerated observer can define a second reference axis. To achieve this, we must provide a practical and operational method for assigning an axis perpendicular to the motion line axis. One possible approach is to use a light beam as a reference direction. We define the second reference direction as follows: Suppose the aberration direction within the accelerated frame is determined relative to the fixed direction of a light beam emitted from a ”plane-wave” source at rest in the accelerated frame. In other words, the coordinate system is established using the electromagnetic axis and the line of motion. For simplicity, we assume that the motion of the aberrated light beam lies in the same plane, with its angular position described by a single angle.
When the accelerated system transitions to uniform motion at a constant velocity, Einstein’s standard clock synchronization procedure can be applied. This synchronization method relies on light signals emitted by a source at rest, under the assumption that light propagates isotropically with constant speed $c$. In this framework, the accelerated observer analyzes the light beam from the stationary emitter using the standard Maxwell equations. According to classical electrodynamics, light is emitted perpendicular to the wavefront of the radiation. We consider a specific configuration in which the component of the reference light beam’s velocity along the $x_{n}$-axis lies parallel to the wavefront of the reference light beam. The aberration of the light beam transmitted through an aperture can be characterized by the angle between the directions of the reference and transmitted beams, which quantifies the aberration increment.
Figure 30: Inertial frame view of observations of the accelerated observer.
Let us examine the relationship between reference directions in the inertial frame and those in the accelerated frame. In the context of our study on the aberration of light, this approach relies on the use of local reference light sources. To define reference directions perpendicular to the motion in both frames, we employ two local light sources: one remains at rest in the accelerated frame, while the other is stationary in the inertial frame. In essence, the reference electromagnetic axis in each frame is established by an individual light beam. However, reference axes can also be defined using more practical methods than light beams. For example, in an Earth-based frame, the reference axis may be aligned with the gravitational field vector, such as the standard direction of a plumb line. The theory of relativity is founded on the equivalence of all local physical frames of reference. Consequently, when the aberration angle arises, it manifests identically across all local reference systems.
Now, let us return to the observations of an accelerated observer as viewed from an inertial frame. We begin with the non-covariant approach. The aberration of light problem can be addressed within a ”single inertial frame” description, without invoking Lorentz transformations. According to our approach, when a transversely moving screen with a hole is placed in the path of a normally incident plane wave of light, the transmitted light undergoes a deviation in energy transport (see Fig. 9). However, an alternative and equally valid explanation for the aberration of light from a moving source relies on Lorentz coordinatization. This explanation describes the uniform translational motion of the light source within the inertial frame.
In Chapter 3, we discussed how absolute time coordinatization can be transformed into Lorentz coordinatization. Within the inertial frame, this transformation can be understood as a change of the time variable according to $t\to t+xv/c^{2}$. The combined effect of the Galilean transformation and this variable change results in the Lorentz transformation when using absolute time coordinatization in the inertial frame. Consequently, in this framework, the speed of light remains independent of the motion of the source.
Simultaneity of events is relative for both inertial and accelerated frames. Events occurring at different locations along the direction of relative motion cannot be simultaneous in both frames when using Lorentz coordinatization—an idea analogous to Einstein’s train-embankment thought experiment. On one hand, the wave equation remains invariant under Lorentz transformations. On the other hand, employing Lorentz coordinatization introduces a time transformation, $t\to t+xv/c^{2}$, which results in a rotation of the radiation phase front by an angle $v/c$.
From the preceding discussion, it follows that in an accelerated system, as viewed by an inertial observer, the electromagnetic axis assigned by the accelerated observer will not be parallel to the $z$-axis of the inertial frame (see Fig. 30). We conclude that, in the accelerated frame, the transmitted light beam propagates along the $z_{n}$-axis of the accelerated frame with an angular displacement of $-v/c$.
We determine the reference directions in both the inertial and accelerated frames using two local light sources, leading to intriguing consequences. First, both the accelerated and inertial observers can directly measure the angular displacement of a transmitted light beam relative to the reference light beam in the inertial frame. Consider an observer in the inertial frame measuring the direction of the transmitted light relative to a reference light beam from an emitter at rest in that frame. The inertial observer finds that the angular displacement is positive and equal to $\theta_{a}=v/c$.
Now, we examine the angular displacement of the transmitted light relative to the reference beam within the accelerated frame. On one hand, the angular displacement of the transmitted light with respect to the reference beam from an emitter at rest in the accelerated frame is $-v/c$. On the other hand, the angular displacement of the inertial reference light beam is $-2v/c$ (see Fig. 17). This implies that the accelerated observer can directly measure the angular displacement of the transmitted light with respect to the inertial frame’s reference light beam. Specifically, the relative angular displacement is $\theta_{a}(\mathrm{transmitted~{}beam})-\theta_{a}(\mathrm{inertial~{}reference~{}beam})=-v/c+2v/c=v/c$. This result is consistent with the measurements of the inertial observer, as expected  [^26] We have been discussing the electromagnetic axis. If we instead determine the reference directions in both frames using two relativistic electron sources, an important distinction arises. Unlike a reference light beam, a reference relativistic electron beam would not allow for an exact measurement of the transmitted light’s angular displacement within our experimental setup..
7.2 The Non-Existence of Instantaneous Three-Dimensional Space
We derived the results for a non-inertial observer’s observations using Lorentz transformations. Let us follow the standard textbook assumption that the linear motion of observers remains the same in both the accelerated and inertial frames. At first glance, if the reference axes are orthogonal to the direction of motion, they should remain parallel to each other. However, due to the angular displacement $v/c$, this leads to an apparent paradox.
Figure 31: The inertial observer sees the rotation of the moving axes with respect to the lab frame axes, which is closely associated with the relativity of simultaneity.
The key to resolving this paradox, as noted above, lies in recognizing that the theory of special relativity eliminates the concept of absolute simultaneity. In the case of the relativity of simultaneity, space and time become intertwined—an observer’s spatial measurements inherently include a subtle influence of time as perceived by another observer.
A closer examination reveals that the commonly accepted assumption of a shared $(x)$ line motion among observers is based on the mistaken belief that they share a common 3-space. This is a misconception. The correct approach is to use the light beam as the reference direction. We have already established that accelerated observer can directly measure the angular displacement between the inertial $(z)$ and accelerated $(z_{n})$ electromagnetic reference directions, and this measurement aligns with the inertial observer’s perspective. In other words, the displacement between electromagnetic axes is a physical reality.
From the inertial observer’s viewpoint, the moving frame’s axes rotate relative to the inertial frame axes, as illustrated in Fig. 31. This rotation is a relativistic kinematic effect. One can directly verify that the accelerated frame’s axes are rotated by an angle $\theta_{w}=v/c$ (to first order in $v/c$) relative to the inertial frame axes (see the next section for further details).
7.3 Acceleration of a Rigid Body in Special Relativity
Now, let us explore further consequences of the relativity of simultaneity. The concept of rigid motion, as understood in Newtonian mechanics, cannot be directly carried over to the special theory of relativity. First, let us study the relativistic kinematics of rotation of the moving frame axes with respect to the inertial frame axes. Consider two emitter-detector setups in an initial inertial frame. Suppose one setup is accelerated from rest to velocity $v$ along the $x$-axis, as shown in Fig. 32. The explanation of aberration relies on absolute time coordinatization  [^27] An alternative explanation involves a clock re-synchronization procedure in the initial inertial frame. Under Einstein’s synchronization, the time for the moving source is obtained by introducing an offset factor $xv/c^{2}$ in the first-order approximation. This time shift results in a rotation of the radiation wavefront by an angle $v/c$ for both sources. In this synchronization scheme, the radiation of the moving source is described using Maxwell’s equations.. A key distinction between relativistic and Newtonian kinematics is that, in the inertial frame, the accelerated emitter-detector setup undergoes a shearing motion. While counterintuitive, this effect is not paradoxical—it arises naturally from the relativity of simultaneity and the geometric structure of spacetime in special relativity.
We have demonstrated that the aberration of light in an accelerated frame, as observed from an inertial frame, can be derived from the Lorentz transformations. In other words, the aberration effect in the accelerated frame, viewed from the initial inertial frame, is a kinematic consequence of special relativity (Fig. 33). It is shown that the relativity of simultaneity is responsible for aberrations to first order in $v/c$. Furthermore, we identified a fundamental asymmetry between inertial and accelerated frames: acceleration (relative to the inertial frame) influences the propagation of light in the accelerated frame. Specifically, in an accelerated system, an accelerated detector does not receive light emitted by an accelerated source (Fig. 33). In contrast, a detector at rest in the inertial frame continues to receive light (Fig. 32, left).
Figure 32: Aberration of light in an inertial frame of reference. The orientation of the wavefront lacks an exact objective meaning. For simplicity, only the wavefront orientation in the absolute time coordinatization of the moving source is depicted here.
Figure 33: Aberration of light in an accelerated frame of reference. In an accelerated system, the accelerated detector does not receive light radiated by the accelerated emitter.
Figure 34: Aberration of light in an inertial frame of reference, as predicted by conventional theory.
Figure 35: Aberration of light in an accelerated frame of reference, as predicted by conventional theory. Acceleration does not disrupt the motional symmetry between the accelerated and inertial reference frames.
Common textbook presentations of special relativity typically follow the (3+1) approach, which relies on the implicit assumption that distant clocks are synchronized according to the absolute simultaneity convention. According to these textbooks, when a reference frame at rest is set into motion, all points in the three-dimensional reference grid are assumed to move simultaneously. The standard derivation of the aberration of light effect is based on the hidden assumption that the $(x_{x},y_{n},z_{n})$ axes of the accelerated frame remain parallel to the $(x,y,z)$ axes of the inertial frame. In other words, it is incorrectly presumed that both accelerated and inertial observers share a common three-dimensional space (see Fig. 34 - Fig. 35). According to conventional theory, acceleration does not disrupt the motional symmetry between the accelerated and inertial reference frames. However, our findings (Fig. 32 - Fig. 33) contradict the textbook predictions. A critical flaw in the widely accepted derivation of the aberration of light effect is the omission of Wigner rotation, which, in our analysis, is fundamentally linked to the relativity of simultaneity. In the next chapter, we will explore the theory of Wigner rotation in detail.
7.4 Discussion
Let us return to the observations of an inertial observer. According to the special theory of relativity, an orthogonal emitter-detector setup undergoes a shearing motion during acceleration, as illustrated in Fig. 32. This effect is purely kinematic and does not involve any forces.
How can this be? The answer is not straightforward, but one way to approach the issue is as follows: Our ordinary intuition often fails when dealing with two space-like separated events. The commonly accepted assumption of rigidity in an accelerated frame is based on the belief that the simultaneous acceleration of the reference space grid has a direct physical meaning. However, this is a misconception. When considering two distant events, we must account for the conventionality of distant simultaneity over a time interval of $l/c$, where $l$ is their spatial separation. The position of the accelerated emitter along the $x$-axis relative to the accelerated detector lacks an exact objective meaning. Due to the finite speed of light, no experimental method can precisely determine this position. Instead, there is an inherent uncertainty (blurring) in the relative position along the $x$-direction, quantified as $lv/c$, which arises from uncertainties in the timing of acceleration.
A relevant example involves a light clock—a rod with mirrors at each end. Textbooks typically describe how, from the perspective of an inertial observer, the light inside the orthogonal light clock follows a zigzag path. However, this explanation implicitly assumes that the $(x_{n},y_{n},z_{n})$ coordinate axes of the moving observer remain parallel to the $(x,y,z)$ axes of the inertial observer. This assumption is incorrect. To properly analyze the behavior of light clocks, we must consider both the phase and initialization of the system. The term ”phase” refers to the zigzag pattern of the light path, which cannot be directly measured in the lab frame. Moreover, the relative position of the accelerated bottom mirror with respect to the accelerated top mirror also lacks an exact objective meaning. Here, too, an uncertainty of order $lv/c$ arises, where $l$ is the rod length.
For completeness, we must also analyze the angular displacement of the $x_{n}$-axis. So far, we have considered only non-relativistic speeds, employing a first-order relativistic approximation. It follows that for the displacement of the detector along the $x$-axis in the $z$-direction, we obtain: $\Delta z\simeq\theta_{w}Lv/c\simeq Lv^{2}/c^{2}$. Thus, special relativity indicates that in the limit $v^{2}/c^{2}\to 0$, the kinematic shearing motion (during acceleration) and the Wigner rotation motion become indistinguishable.
8 Kinematics of Wigner Rotation
The focus of this chapter is the Wigner rotation  [^28] It is well known that the composition of non-collinear Lorentz boosts does not result in a simple boost but rather in a Lorentz transformation that includes both a boost and a spatial rotation. This rotation is known as the Wigner rotation [39, 40]. It is sometimes referred to as the Thomas rotation (see, e.g., [13, 41]).. We previously introduced the Wigner rotation in the last chapter, where our discussion primarily centered on the aberration of light. However, the Wigner rotation is also closely related to the aberration of particles. As a relativistic kinematic effect, the Wigner rotation describes how the coordinate axes of a reference frame, moving along a curvilinear trajectory, rotate with respect to the axes of an inertial Lorentz frame. The primary objective of this chapter is to derive an expression for this effect. Throughout, we have aimed to minimize mathematical complexity while maintaining clarity.
8.1 Composition of Lorentz Boosts
Let us now consider a relativistic particle accelerating in an initial inertial frame and analyze its evolution within Lorentz coordinate systems. Since the particle’s permanent rest frame is not inertial, we introduce an infinite sequence of comoving frames to address this difficulty. At each instant, the rest frame is a Lorentz frame centered on the particle, moving alongside it. As the particle’s velocity changes at an infinitesimally later instant, a new Lorentz frame, centered on the particle and moving at the updated velocity, is adopted.
Let us denote the three inertial frames by $K,K^{\prime},K"$. The initial inertial frame is $K$, while $K^{\prime}$ is the rest frame of the particle, which has a velocity $\vec{v}=\vec{v}(\tau)$ relative to $K$, and $K"$ represents the particle’s rest frame at the next instant of proper time $\tau+d\tau$, moving with an infinitesimal velocity $d\vec{v^{\prime}}$ relative to $K^{\prime}$. All three inertial frames are assumed to be Lorentz frames. To clarify velocity measurements in different inertial frames: an observer in $K$ measures the velocity of frame $K^{\prime}$ as $\vec{v}$, while an observer in $K^{\prime}$ measures the velocity of frame $K"$ as $d\vec{v^{\prime}}$.
Now, let us examine the transformation of velocity in the theory of relativity. For a rectilinear motion along the $z$ axis, the velocity transformation follows the equation: $v_{z}(\tau+d\tau)=(dv^{\prime}_{z}+v_{z})/(1+v_{z}dv^{\prime}_{z}/c^{2})$. As it happens with the composition of Galilean boosts, collinear Lorentz boosts commute. This implies that the result of successive collinear Lorentz boosts is independent of which transformation applies first.
Unlike the case of Lorentz boosts in collinear directions, Lorentz boosts in different directions do not commute. While the successive application of two Galilean boosts results in a single Galilean boost, and the successive application of two rotations yields another rotation, the composition of two non-collinear Lorentz boosts does not produce a single Lorentz boost.
We assume that frame $K^{\prime}$ moves with velocity $v_{z}$ along the $x$-axis relative to $K$ and in frame $K^{\prime}$, the particle’s acceleration is perpendicular to the trajectory of the initial inertial frame. Suppose that frame $K"$ moves relative to $K^{\prime}$ with nonrelativistic velocity $dv^{\prime}_{x}$.
From the perspective of an inertial observer, time dilation occurs in the Lorentz frame moving with velocity $\vec{v}_{z}$ relative to the initial inertial frame, given by $dt/\gamma=d\tau$. Due to this time dilation, the velocity increment in the initial inertial frame $dv_{x}$ corresponds to a velocity $dv^{\prime}_{x}=\gamma dv_{x}$ in frames $K^{\prime}$.
8.2 Wigner Rotation
Numerous incorrect expressions for the Wigner rotation appear in the literature, underscoring the need for a careful and transparent treatment of this topic. Rather than deriving all transformation matrices for four-vector components, we adopt a geometric perspective that offers clear and intuitive insight into the phenomenon.
The mathematics involved is notably straightforward, relying only on basic algebra. The key challenge lies in bridging the conceptual gap that arises from the fact that different Lorentz frames cannot be described within a shared three-dimensional space.
[…]
Several key insights can be drawn regarding the theory of Wigner rotation. Notably, the expression for the Wigner rotation angle in the proper frame, $d\vec{\Phi}_{K"}$ can be reformulated as $d\vec{\Phi}_{K"}=(1-1/\gamma)\vec{v}_{K"}\times d\vec{v}_{K"}/v_{K"}^{2}$, where $\vec{v}_{K"}=-\vec{v}$ is the velocity of the original frame relative to the proper frame, and $d\vec{v}_{K"}=-\gamma d\vec{v}$ is its differential change. This alternative form emphasizes the expression of the Wigner rotation in terms of quantities measured in the proper frame.
Transforming this expression back to the lab frame $K$ yields $d\vec{\Phi}_{K}=(1-1/\gamma)\vec{v}\times d\vec{v}/v^{2}$ representing the Wigner rotation angle of the proper frame’s axes as observed in the lab frame. According to the principles of special relativity, this quantity must remain form-invariant under Lorentz transformations  [^29] The correct expression for the Thomas (Wigner) rotation was first obtained by V. Ritus [42]. In deriving expressions for the Thomas (Wigner) rotation, the majority of authors (see e.g. [41]) was supposedly guided by the incorrect expression for Thomas (Wigner) rotation from Moeller’s monograph [13]. The expression obtained by Moeller is given by $\vec{\delta\Phi}=(1-\gamma)\vec{v}\times d\vec{v}/v^{2}=(1-\gamma)\vec{\delta\theta}$ (and subsequently expression for Thomas precession $\Omega_{\mathrm{T}}=(1-\gamma)\omega_{0}$). It should be noted that, in his monograph, Moeller stated several times that this expression is valid in the lab Lorentz frame. Clearly, this expression and correct result differ both in sign and magnitude. An analysis of why Moeller obtained an incorrect expression for the Thomas (Wigner) rotation in the lab frame is the focus of Ritus paper [43]. As analyzed in Ritus’s work [43], Moeller’s error is not computational but rather conceptual. The correct expression, derived decades earlier by several researchers, remained largely unnoticed amid the proliferation of inaccurate results, as discussed in the review [44].. We note that owing to the relativistic effect of time dilation in the reference frame that moves to the lab frame, the Wigner rotation angle in the proper frame is always $\gamma$ time higher than in the lab frame  [^30] We note that in 1986, M. Stranberg obtained an expression for the Thomas (Wigner) rotation correct both in the lab inertial frame and the comoving reference frame [45]. His paper is one of the few that explicitly states that the angle of Thomas (Wigner) rotation in the comoving frame is $\gamma$ times higher than in the lab frame..
We derived the exact relation $d\vec{\Phi}=(1-1/\gamma)\vec{v}\times d\vec{v}/v^{2}$ using only rudimentary knowledge of special relativity. In standard textbooks on relativity, the spatial rotation that arises from the composition of two non-parallel Lorentz boosts is typically introduced through an algebraic approach. This reliance on algebra, without sufficient geometrical interpretation, is one reason why many authors arrive at an incorrect expression for the Wigner rotation. These treatments often describe the rotation of a moving object without addressing its geometrical significance, leading to serious difficulties in both the interpretation of the computations and the physical meaning of the results.
Figure 36: Illustration of the Wigner rotation interpreted with respect to the axes of the proper frame.
8.3 Measuring Wigner rotation
The Wigner rotation is a relativistic kinematic effect that arises from the geometry of pseudo-Euclidean space-time. It manifests as a rotation of the spatial coordinate axes of a particle’s proper frame when the particle follows a curvilinear trajectory in the Lorentz lab frame. In this context, we derive the expression for the infinitesimal rotation resulting from the acceleration of a relativistic particle:
$d\vec{\Phi}=(1-1/\gamma)\vec{v}\times d\vec{v}/v^{2}=\left(1-\frac{1}{\gamma}\right)\vec{\delta\theta}~{}.$ (20)
Here $d\vec{v}$ represents the infinitesimal change in velocity due to acceleration, $d\Phi$ is the Wigner rotation angle describing the rotation of the particle’s proper frame relative to the Lorentz lab frame, and $d\theta$ is the infinitesimal orbital angle. A natural question arises: how can this rotation be observed or measured? Since a moving coordinate system changes its position over time, determining its orientation becomes nontrivial. To explore the physical meaning and possible experimental interpretation of this rotation, we illustrate the problem of defining the orientation of a moving frame using a simple example. The Wigner rotation, when interpreted in the context of the original inertial frame, is intimately related to length contraction  [^31] The length of a moving object depends not only on the structure of space-time but also on our choice of synchronization convention. As with the orientation of radiation phase fronts, such quantities do not possess an exact objective meaning. Under the standard clock synchronization, they are typically regarded as ”relativistic kinematic effects.” The Wigner rotation similarly arises as a coordinate effect, lacking an absolute, observer-independent significance..
Figure 37: Illustration of the Wigner rotation interpreted with respect to the axes of the lab frame.
Suppose that, after performing a Lorentz boost along the $z$-axis with velocity $v_{z}$, an observer in the lab frame rotates the coordinate system by an angle $\vec{\delta\theta}=\vec{v_{z}}\times d\vec{v_{x}}/v_{z}^{2}$, such that the new $x_{r}$-axis is orthogonal to the vector $\vec{v_{z}}+d\vec{v_{x}}$, and the $z_{r}$-axis is aligned with it. Now, consider a rod oriented along the $x$-axis in the comoving frame. The motion takes place in the $(x,z)$-plane, and the rod is initially perpendicular to the velocity $\vec{v_{z}}$. After rotating the lab frame axes, the projection of the rod onto the $z_{r}$-axis becomes simply $l\delta\theta$, where $l$ is the length of the rod in the comoving frame, and also in the lab frame following the first boost along the $z$-axis. After a second infinitesimal Lorentz boost with velocity $d\vec{v_{x}}$, this projection undergoes Lorentz contraction and becomes $l\delta\theta/\gamma$ (Fig. 37). Assume now that the lab observer fixes the orientation of the comoving frame axes. In the ultrarelativistic limit $\gamma\to\infty$, the comoving frame axes become parallel to the rotated lab frame axes $(x_{r},z_{r})$, and thus the projection of the rod onto the $z_{r}$-axis vanishes. However, for an arbitrary (finite) velocity, the comoving frame axes are not aligned with the rotated lab frame axes. Due to the contracted projection, the angular deviation is $-\vec{\delta\theta}/\gamma$. Consequently, one can directly verify that the comoving frame axes are rotated relative to the original lab frame axes $(x,z)$ by an angle $\delta\theta-\delta\theta/\gamma$, which is precisely the Wigner rotation angle, as given by Eq.(20). This example illustrates how the Wigner rotation arises naturally from the composition of Lorentz transformations.
8.4 Discussion
Numerous incorrect expressions for the Wigner (Thomas) rotation can be found in the literature. Here, we examine the underlying reasons behind these errors. In 1959, Bargmann, Michel, and Telegdi (BMT) formulated a consistent relativistic theory for spin dynamics as observed in the laboratory frame, which was later confirmed experimentally [46]. A common misconception is that the BMT equation inherently includes the standard (but incorrect) expression for the Wigner rotation in the Lorentz lab frame. Many physicists, having learned about Wigner (Thomas) rotation from well-known textbooks, argue: ”The highly precise measurements of the electron’s magnetic-moment anomaly, conducted on relativistic electrons, rely on the BMT equation—where the Wigner rotation is an integral part—thus serving as an experimental confirmation of the standard expression for the Wigner rotation.” This belief, however, is both widespread and misleading. It is important to clarify that the results in the Bargmann-Michel-Telegdi paper were derived using a semiclassical approximation of the Dirac equation. The Wigner rotation was not explicitly considered, as the Dirac equation allows for a complete description of spin dynamics without separating it into Larmor and Wigner components.
It should be noted that the application of Wigner (Thomas) rotation theory in relativistic spin dynamics is complex. Bargmann, Michel, and Telegdi began their analysis in the particle’s rest frame, deriving the equation of motion for angular momentum. They then generalized it to the Lorentz lab frame before transforming it back to the rest frame. This raises an interesting question: Why is it convenient to transform this equation to the rest frame at that instant? In experimental practice, working with the proper spin is preferred due to the clear physical interpretation of the three-dimensional spin pseudo-vector $\vec{s}$. Unlike the three-dimensional momentum $\vec{p}$, which has well-defined components in every reference frame, angular momentum is specific to a single frame and does not transform. Thus, any meaningful statement about spin implicitly refers to the particle’s instantaneous rest frame. For example, when we say that a particle’s spin in the lab frame makes an angle $\phi$ with its velocity, we actually mean that in the particle’s rest frame, the spin vector forms the same angle with the direction of the lab frame’s motion.
Physical phenomena in non-inertial frames can still be analyzed within the framework of standard special relativity by applying well-established relativistic kinematic effects. In the context of spin dynamics, this eliminates the need for measurements using non-inertial devices.
The conventional approach to explaining spin dynamics in the lab frame is quite unusual. As previously noted, the laws of physics in the lab frame account for all physical phenomena, including observations made by non-inertial observers. In particular, spin orientation measurements with respect to the lab frame axes can be determined within the proper frame. From the lab frame perspective, the interpretation of spin rotation experiments is straightforward and can be derived from the predictions of the proper observer regarding the measurements made by the lab observer. Thus, the spin rotation measurements by a polarimeter in the lab frame is interpreted from the viewpoint of the proper observer as viewing of the lab observer. This inherent complexity has led textbook authors to derive incorrect expressions for Wigner rotation—an issue that will be discussed further in Chapter 17.
Additionally, it is crucial to examine why errors in Wigner rotation theory went undetected for so long. Wigner (or Thomas) rotation is typically presented within the context of spin kinematics as a peculiar effect of special relativity. This limited perspective likely explains why, until recently, researchers failed to recognize the significant discrepancies between different works.
The Wigner rotation is a fundamental relativistic phenomenon, on par with time dilation and length contraction. Special relativity demonstrates that the aberration effect in the rotating Earth-based frame, as observed from the inertial Sun-based frame (in Lorentz coordinatization), is a kinematic consequence of the theory. In this context, the shift in aberration images within the Earth-based frame, when viewed from the Sun-based frame, is governed by the Wigner rotation. According to Wigner rotation theory, observers following different trajectories experience different 3-spaces. However, previous literature has incorrectly assumed that an Earth-based observer and an inertial (e.g., Sun-based) observer share a common 3-space—an assumption that has long been taken for granted.
9 Aberration of Relativistic Particles
We now examine the phenomenon of aberration for relativistic particles.
Consider a particle source initially at rest in an inertial frame, emitting particles that travel with a constant velocity $V$ along the $z$-axis. Suppose both the observer (along with their instruments) and the particle source are uniformly accelerated from rest to a velocity $v$ along the $x$-axis, under the condition that $v\ll V$. Our goal is to understand how the aberration of the particle beam is perceived by the observer in this accelerated frame.
To approach this question, we begin by summarizing the key result.
The central question is: How should the aberration of a particle beam be described when the source is at rest in the accelerated frame? To predict the outcome of such a measurement, the observer must refer to the Langevin metric, given in Eq. (15). Within this framework, the aberration angle $\theta_{a}$ is related to the system parameters by the expression: $\theta_{a}=-(1-1/\gamma)v/V$, where $\gamma=1/\sqrt{1-V^{2}/c^{2}}$. is the Lorentz factor of the particles. This relativistic expression smoothly reduces to the classical result $\theta_{a}=0$ in the limit $V^{2}/c^{2}\to 0$.
It is crucial to note that a correct description of aberration in the accelerated frame requires incorporating the full metric tensor, even to first order. This is because the cross term in the Langevin metric—representing the leading-order deviation from the Minkowski metric—plays an essential role in capturing the non-inertial kinematics of relativistic particle motion.
In the ultrarelativistic limit $V/c\to 1$, the aberration increment approaches the simple and intuitive form: $\theta_{a}=-v/c$.
9.1 Explanation of Electron Aberration Based on Electrodynamics
The aberration of particle trajectories can be effectively addressed using dynamical arguments. Electromagnetic forces, which determine the behavior of an emitted electron beam, are influenced by the acceleration of the source in such a way that they cause a deviation in the direction of electron transport. To analyze these effects, we can derive the electromagnetic fields in an accelerated frame through a Galilean transformation.
For readers already familiar with special relativity from standard textbooks, it may be advantageous to approach the aberration of particles—both theoretical and experimental—using a microscopic perspective. From this standpoint, one can calculate relativistic quantities directly from the fundamental theories of matter, without explicitly invoking relativistic kinematics. The aberration of particles associated with the transformation from an inertial to an accelerated frame can be understood as the result of a force acting on the system. Importantly, this issue is unrelated to the concept of the ether and is fully contained within the framework of special relativity. Let us consider a specific case in which an electron gun, serving as a source of relativistic electrons, is initially at rest in an inertial frame and then accelerated to a velocity $v$ along the $x$-axis. The electromagnetic forces governing the properties of the emitted electron beam are modified by this acceleration, resulting in a change in the beam’s transport direction. The underlying mechanism behind the change in electron momentum $\vec{p}$, when both the source and the observer (along with their measuring instruments) are at rest in the accelerated frame, can be fully explained within the framework of electrodynamics.
We can derive the electrodynamics equations in an accelerated reference frame by applying a Galilean transformation. When analyzing Maxwell’s equations under such a transformation, a key question arises: What are the transformation laws for the electromagnetic fields $\vec{E}$ and $\vec{B}$?
The electric and magnetic fields $\vec{E}_{n}$ and $\vec{B}_{n}$ observed in the accelerated frame—where the particle source is at rest—differ from the fields $\vec{E}$ and $\vec{B}$ observed in the original inertial frame, prior to the Galilean boost. Our goal is to determine the relationship between these sets of fields.
To proceed, we transform the coordinates $(t,x,y,z)$ of the inertial observer $S$, moving with velocity $-v$ relative to the accelerated observer $S_{n}$, using the Galilean transformation: $x_{n}=x-vt$, $t_{n}=t$. Substituting these into the Minkowski metric, $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$, we obtain the so-called Langevin metric in the accelerated frame Eq. (15): $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$. This transformed metric allows us to study Maxwell’s equations in a non-inertial frame using the Galilean transformation with velocity $-v$. In the specific case where $\vec{B}=0$, the transformed fields to first order in $v/c$ are: $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$. Here, $\vec{v}$ is the velocity of the source in the original inertial frame. This result captures the leading-order relativistic correction to the magnetic field observed in the accelerated frame.
Figure 38: Electron aberration observed when both the electron source and the observer are at rest in an accelerated frame.
[…]
The aberration of particles resulting from the transformation between an inertial and an accelerated coordinate system can be interpreted as a manifestation of the Lorentz force. The electromagnetic forces that govern the behavior of an emitted electron beam are influenced by acceleration relative to the fixed stars, leading to a deviation in the beam’s transport direction, as illustrated in Fig. 38.
9.2 Magnetic Field Measurements in an Accelerated Electron Source
We now turn to an apparent paradox. Special relativity introduces concepts that may seem counterintuitive at first. Yet it remains a rigorously mathematical theory, and it leads to a surprising conclusion: when we consider an accelerated electron source, the electromagnetic fields in the accelerated frame are not static—the problem becomes time-dependent. This is a direct implication of special relativity.
Specifically, if both the electron source and the observer are at rest in the accelerated frame, one might intuitively expect that everything should remain unchanged over time—that is, the electric field should not vary. However, special relativity tells us otherwise. The time derivative of the electric field, $\partial\vec{E}/\partial t_{n}$, is nonzero because of the spatial variation of the field, $\partial\vec{E}/\partial x_{n}\neq 0$, and the relation $\partial\vec{E}/\partial t_{n}=v\partial\vec{E}/\partial x_{n}$. This seemingly paradoxical result is simply a manifestation of the four-dimensional formalism of relativity. It carries no physical contradiction or special consequence—just a shift in perspective due to the geometry of spacetime.
To clarify this, it may help to draw an analogy with a familiar concept in classical electromagnetism: the use of gauge transformations. For example, when we work in the Coulomb gauge, the scalar potential appears to propagate instantaneously. However, this is merely a mathematical artifact of the gauge choice. In a complete and self-consistent calculation, all unphysical instantaneous effects cancel out, preserving causality and agreement with physical observation.
We derived the electrodynamics equations in an accelerated frame by applying the inverse Galilean transformation (with velocity $-v$) of Maxwell’s equations. Under a Galilean transformation, and assuming $\vec{B}=0$ in the original frame, the transformed fields become $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$. The electric field $\vec{E}_{n}$ and the induced magnetic field $\vec{B}_{n}$ are both static, meaning they do not vary with time—the system is entirely time-independent. Many physicists familiar with standard electrodynamics might argue, based on textbook knowledge, that a static magnetic field must arise from electric currents, which in turn require moving charges. Since no charges are in motion in this case, the presence of a magnetic field appears paradoxical. This represents a fundamentally new kind of situation, distinct from traditional electrodynamics in an initial inertial frame.
How can that be? Even when the electron source and the observer are at rest in the accelerated frame, the time derivative $\partial/\partial t_{n}$ is generally nonzero. This can be readily demonstrated by transforming from the coordinates $(t,x,y,z)$ of an inertial observer $S$—moving with velocity $-v$ relative to the non-inertial observer $S_{n}$—to the coordinates $(t_{n},x_{n},y_{n},z_{n})$ of $S_{n}$. Using a Galilean transformation, the inertial coordinates can be related to the non-inertial ones as follows: $x_{n}=x-vt,~{}y_{n}=y,~{}z_{n}=z,~{}t_{n}=t$. This allows all fields to be re-expressed in terms of the variables $x_{n},y_{n},z_{n},t_{n}$. Taking partial derivatives accordingly, we find: $\partial/{\partial t_{n}}=\partial/{\partial t}+v\partial/{\partial x}$, $\partial/{\partial x_{n}}=\partial/{\partial x}$. Therefore, if the fields are time-independent in the inertial frame (i.e., $\partial/\partial t=0$), it follows that: $\partial/\partial t_{n}=v\partial/\partial x_{n}$. Thus, $\partial/\partial t_{n}$ is not necessarily zero, even though the source and observer are at rest in the non-inertial frame.
A natural starting point is to assume that the time derivative $\partial/\partial t_{n}$ originates from a Galilean boost. To generalize, we now consider the case of Lorentzian coordinatization. When the reference frame $S_{n}$ begins moving at a constant velocity, Einstein’s standard procedure for clock synchronization can be applied. In this context, the time coordinate $t^{(L)}_{n}$ in the frame $S_{n}$—as defined by Einstein synchronization—is obtained by introducing an offset term $x_{n}v/c^{2}$ and applying the first-order approximation: $t^{(L)}_{n}=t_{n}-x_{n}v/c^{2}$. This transformation allows us to reinterpret the new time coordinate in the accelerated frame in such a way that Maxwell’s equations remain valid, particularly in describing the aberration of electrons. Taking partial derivatives with respect to the transformed coordinates yields: $\partial/{\partial t^{(L)}_{n}}=v\partial/{\partial x_{n}}$, $\partial/{\partial x^{(L)}_{n}}=\partial/{\partial x_{n}}+(v^{2}/c^{2})\partial/{\partial t_{n}}$. Hence, to first order in $v/c$, we find: $\partial/\partial t^{(L)}_{n}=v\partial/\partial x^{(L)}_{n}$ neglecting terms of order $v^{2}/c^{2}$. It is important to note that changing the (four-dimensional) coordinate system does not give rise to new physical phenomena. The underlying physics remains invariant under such transformations.
When the situation is described as we have done here, the paradox seems to disappear entirely; it becomes quite natural that, in general, space and time can no longer be treated as separate entities. The argument for a paradox typically goes like this: If we could measure the magnetic field $\vec{B}_{n}$ using an ordinary magnetometer, then a paradox would indeed arise. However, such a measurement is not possible. As discussed earlier, the magnetic field $\vec{B}_{n}$ induces a change in the transverse momentum of test particles, described by $\theta_{a}=p_{\perp}/p=(1-1/\gamma)v/V$. For non-relativistic particle velocities, applying a binomial expansion yields $\vec{\theta}_{a}=-[V^{2}/(2c^{2})]\vec{V}\times\vec{v}/V^{2}$, indicating that the momentum perturbation is a relativistic effect. According to special relativity, as $V^{2}/c^{2}\to 0$, the change in particle momentum vanishes. Thus, a classical magnetometer using test particles would detect no trace of the magnetic field. Moreover, it is not just that this particular classical magnetometer cannot detect $\vec{B}_{n}$; if the theory of relativity holds, any classical magnetometer—regardless of its operating principle—would similarly fail to reveal any evidence of the field’s effect.
To understand magnetic field measurements in an accelerated electron source, we begin by examining how a classical magnetometer behaves when observed from an accelerated frame. Since a full treatment is complex, we consider a simplified, conventional type of magnetometer. One practical approach to observing nuclear magnetism is through the phenomenon of nuclear magnetic resonance (NMR), where a proton resonance apparatus can serve effectively as a proton resonance magnetometer. We present a general analysis of this classical type of magnetometer, independent of its experimental precision. In the inertial lab frame, a current loop possessing a magnetic moment $\vec{m}$ acquires an electric dipole moment when it moves with velocity $\vec{v}$ relative to the lab. This dipole moment is given by: $\vec{d}=\vec{v}\times\vec{m}/c$. Since the magnetic properties of materials arise from atomic-scale current loops, a magnetized body undergoing acceleration relative to the lab frame should exhibit an induced electric polarization in that frame. As a result, the potential energy of a moving particle with intrinsic magnetic moment $\vec{m}$, in the presence of a magnetic field $\vec{B}$ and electric field $\vec{E}$, is given by: $U_{p}=-\vec{m}\cdot\vec{B}-\vec{d}\cdot\vec{E}$. To analyze this in an accelerated frame, we apply the Galilean transformation (with velocity $-\vec{v}$) to Maxwell’s equations. Under this transformation, assuming an initial zero electric dipole moment ($\vec{d}=0$), the magnetic and electric dipole moments in the accelerated frame become: $\vec{m}_{n}=\vec{m}$, $\vec{d}_{n}=-\vec{v}\times\vec{m}/c$. From vector algebra, we use the identity $(\vec{a}\times\vec{b})\cdot\vec{c}=\vec{a}\cdot(\vec{b}\times\vec{c})$ to evaluate terms in the potential energy expression: $(-\vec{v}\times\vec{E})\cdot\vec{m}=-\vec{v}\cdot(\vec{E}\times\vec{m})$, $(-\vec{v}\times\vec{m})\cdot\vec{E}=-\vec{v}\cdot(\vec{m}\times\vec{E})$. Thus, the additional contribution to the potential energy evaluates to zero, as expected. Consequently, a conventional magnetometer cannot detect the magnetic field $\vec{B}_{n}$ in the accelerated frame.
9.3 Aberration of Particles from a Moving Source
Let us examine what occurs when the electron gun—serving as the source of a relativistic electron beam—is accelerated from rest to a velocity $v$ along the $x$-axis in the initial inertial frame (see Fig. 39). In other words, we consider an active (physical) Lorentz boost. In an active boost, we are observing the evolution of the same physical system over time, analyzed from the perspective of a single reference frame. The simplest method of clock synchronization in this context is to retain the same set of uniformly synchronized clocks used when the particle source was at rest. In other words, we continue to use the synchronization method based on clock transport or Einstein synchronization, which relies on light signals emitted by a stationary dipole source. This approach is typically the most convenient, as it allows for a smooth and consistent definition of the metric tensor. It also preserves simultaneity and follows the convention of absolute time—or absolute simultaneity.
Figure 39: Electron aberration from a moving source in an inertial frame. The electron gun moves perpendicular to the optical axis.
We can determine the electric field $\vec{E}$ and magnetic field $\vec{B}$ observed in the inertial frame where the source particle is moving with velocity $\vec{v}$ by applying the Galilean transformation. When analyzing the equations of electrodynamics under such a transformation, a natural question arises: What are the transformation laws for the electromagnetic fields? The equivalence between the active and passive viewpoints stems from the fact that moving a physical system in one direction is equivalent to moving the coordinate system in the opposite direction by the same amount. Consequently, the fields in the inertial frame before an active Galilean boost are identical to the fields $\vec{E}^{\prime}$ and $\vec{B}^{\prime}$ in the comoving coordinate system. Under a Galilean transformation, and assuming $\vec{B}^{\prime}=0$ in the comoving coordinate system, the transformed fields in the lab frame are given by: $\vec{E}=\vec{E^{\prime}}$, $\vec{B}=\vec{v}\times\vec{E^{\prime}}/c$. This relation captures only first-order effects in $v/c$, where $\vec{v}$ is the velocity of the source relative to the lab frame.
[…]
We analyzed the effect of particle aberration by considering terms only up to first order in $v/c$. A basic explanation of this effect is well-known: in the inertial laboratory frame, particle aberration can be readily understood through the Galilean transformation of velocities between reference frames. It is important to emphasize that the dynamical reasoning presented here clarifies the physical meaning of particle aberration in the inertial frame. The kinematics effects are only an interpretation of the behavior of the electromagnetic fields.
Let us now return to the magnetic field measurements in the accelerated electron source, as shown in Fig. 39. In the lab frame, the electron source is in motion and its evolution is analyzed from the perspective of an inertial reference system. A key feature of this situation is that the velocity of the source carries real physical significance. Indeed, the problem is explicitly time-dependent. For a moving electron source, special relativity tells us that a magnetic field arises according to $\vec{B}=\vec{v}\times\vec{E}/c$, leading to a change in the transverse momentum of test particles. This change can be quantified as $\theta_{a}=p_{\perp}/p=v/V$, a result that remains valid even in the non-relativistic limit as $V^{2}/c^{2}\to 0$. This brings us to an interesting question: what does an inertial observer measure when using a conventional magnetometer? Given that charges in the source are moving, the resulting current produces a magnetic field described by $\vec{B}=\vec{v}\times\vec{E}/c$. This field is indeed detectable with a standard magnetometer. To understand this more concretely, consider the operation of an NMR magnetometer. When at rest in the lab frame, it measures the additional potential energy experienced by a proton with magnetic moment $\vec{m}$ in the field $\vec{B}$, given by $-\vec{m}\cdot\vec{B}\neq 0$. We come to the conclusion that the magnetic field generated by a moving source in an inertial frame is a real field in the sense we have defined it.
9.4 Inertial Frame View of Observations by a Non-Inertial Observer
Earlier, we demonstrated that the aberration of light observed in an accelerated frame, when viewed from the initial inertial frame, is a well-known kinematic effect described by the special theory of relativity. Can we not interpret the aberration of particles in an accelerated frame in a similar fashion? As we shall see, the effect of particle aberration in non-inertial frames can indeed be understood within an inertial frame using the framework of standard special relativity—specifically, by employing the theory of Wigner rotation.
The Wigner rotation is a relativistic kinematic effect in which the coordinate axes of a particle’s proper reference frame appear to rotate relative to an initial (Lorentz) inertial frame. This rotation arises due to successive Lorentz boosts in different directions. The Wigner rotation angle in the inertial frame can be expressed in vector form as $\vec{\theta_{w}}=(1-1/\gamma)\vec{V}\times\vec{v}/V^{2}$, where $\vec{V}$ is the vector of particle velocity in the initial Lorentz frame before acceleration, $\vec{v}$ is the vector of small velocity change ($v\ll V$) due to acceleration, $\gamma$ is the Lorentz factor associated with $V$, and $\theta_{w}$ is the Wigner rotation angle of the proper frame axes as observed from the (Lorentz) inertial frame. We now examine how this formula applies in the specific case where both the observer and the particle source are accelerated to a velocity $v$ in the inertial frame.
To find the rotation magnitude in the stated problem, we introduce a composition of Lorentz boosts. Let $S$ be the initial inertial frame of reference, $S_{n}$ an accelerated frame with velocity $\vec{v}=\vec{e}_{x}v$ relative to the inertial frame, and $S^{\prime}$ is a frame which moves relative to the $S_{n}$ with velocity $\vec{V}$ along the $z_{n}$-axis. Two sequential boosts from the inertial frame $S$ to $S_{n}$ and then to $S^{\prime}$ are equivalent to the boost from $S$ to $S^{\prime}$ and the subsequent rotation. The frame $S^{\prime}$ is rotated through the angle $\theta_{w}$ with respect to the inertial frame $S$. Now we must be careful about the direction of rotation. There is a good mnemonic rule. The rule says that the direction of the Wigner rotation in the Lorentz lab frame is the same as the direction of the velocity rotation in the Lorentz lab frame. We can easily understand that the coordinate axes of the frame $(x^{\prime},y^{\prime},z^{\prime})$ will be parallel to the coordinate axes of the accelerated frame $(x_{n},y_{n},z_{n})$. As viewed from the inertial frame, the coordinate axes of the accelerated frame $(x_{n},y_{n},z_{n})$ are also rotated through the angle $\theta_{w}$ with respect to the coordinate axes of the inertial frame $(x,y,z)$. Our results show that we cannot remain with the framework of parallel axes $z_{n},z$ when considering the inertial frame view of particle beam observations of the accelerated (up to the velocity $v$) observer in perpendicular geometry.
Consider an electron gun–detector setup initially at rest in an inertial frame of reference. Suppose this system is then uniformly accelerated along the $x$-axis until it attains a velocity $v$. Now, from the perspective of an observer who remains at rest in the original inertial frame, the direction of the electron beam’s propagation appears to shift. This observer measures an angular displacement of the beam given by $\theta_{a}=v/V$, where $V$ is the velocity of the particles emitted by the gun. This effect, known as particle aberration, is a well-established phenomenon. A key distinction between relativistic and Newtonian kinematics emerges when we examine the behavior of the emitter–detector system under acceleration. From the inertial frame’s viewpoint, the accelerated setup undergoes a shearing motion: the detector no longer remains aligned with the electron beam after acceleration. For instance, if the gun–detector assembly was originally aligned along the $z_{n}$-axis, the detector fails to intercept the beam once the acceleration ends. In the accelerated frame, this misalignment is reflected in the angular deviation, and we obtain an expression for the aberration increment: $\theta_{a}=-\theta_{w}=-(1-1/\gamma)v/V$. In the special case of light, where $V\to c$, this expression simplifies to: $\theta_{a}=-v/c$.
The formula $\vec{\theta}_{a}=-(1-1/\gamma)\vec{V}\times\vec{v}/V^{2}$ matches the result previously derived for observations made by a non-inertial observer within the framework of electrodynamics. According to special relativity, the aberration of particle trajectories observed in an accelerated frame—when viewed from an initial inertial frame (using Lorentz coordinates)—is a purely kinematic effect. In such cases, the apparent shift in particle directions (i.e., aberration) within the accelerated frame is governed by the Wigner rotation. The theory of Wigner rotation reveals that observers following different worldlines possess distinct notions of 3-space. This insight challenges a widespread but incorrect assumption in earlier literature: that a non-inertial observer (e.g., Earth-based) and an inertial observer (e.g., Sun-based) share a common 3-space. It is important to emphasize that the dynamical arguments presented above clarify the physical meaning of the Wigner rotation in a rotating frame. The relativistic kinematic effects observed in the rotating frame—when interpreted from the perspective of an inertial observer—are simply manifestations of how electromagnetic fields transform under relativistic motion.
10 Earth-Based Setups to Detect the Aberration Phenomena
10.1 The Potential of Earth-Based Electron Microscopes
Above, we discussed the new relativistic particle kinematics in rotating frames of reference. Since we live on a rotating Earth, the distinction between relativistic kinematics in rotating and non-rotating frames holds both theoretical and practical importance. We propose an aberration-of-particle experiment using an Earth-based particle source. The results of such experiments can be interpreted within the framework of special relativity. However, obtaining an accurate solution in the Earth-based frame requires the use of the metric tensor, even for first-order approximations.
We propose an experiment using commercial 200-kV scanning transmission electron microscope (STEM) as a particle source. One could use a vertically oriented optical setup. We suppose that the electron beam is imaged by a lens to a spot that lies in the image plane. Measurement of the spot shift with respect to the optical axis (which is parallel to the zenith-nadir axis) is equivalent to the determination of the angular displacement. Due to the high stability, the aberration increment could be observed  [^32] The dedicated STEM with high stability was made [47]. It was developed based on a 200-kV commercial instrument. The electron source was installed in the anti-seismic room. Temperature fluctuations in the room were within 0.2 ${}^{o}C$. The magnetic field in the room was less than $0.2mG$ near the source column. To prevent thermal drift, a shield box was placed over the high-voltage generator. The column was covered with rubber sheets to reduce the temperature fluctuations. The image drift measured during a period of 61 minutes was $\simeq 6~{}\mathrm{nm}$. The drift speed and direction during this period were approximately constant. The average drift speed was 0.12 nm/min. It is substantially smaller than that conventional commercial microscope, in which the drift speed is about 2 nm/min..
Consider the inertial sun-based reference system. In this system, there is the earth which rotates around the sun with orbital velocity $v$. Developing into powers of $v/c$ we can classify effects for velocities $v\ll c$ as of the first order, of the second order, or higher orders. Clearly, in the case of the rotating (around the sun) earth-based frame, we consider the small expansion parameter $v/c\simeq 10^{-4}$ neglecting terms of order of $v^{2}/c^{2}$.
Suppose that an earth-based observer performs an aberration of particle measurement. The aberration angle varies with a one-year temporal period. An approximate formula to express the aberration angle can be found by using the Langevin metric in the earth-based frame or by using the Wigner rotation theory in the sun-based frame. In the aberration of particles, the aberration angle is the apparent angular deviation of the position of the particle source relative to the real location of the source. The reference axis in the earth-based frame can be formed by a plumb line, which is the most fundamental local earth-based coordinate system.
When measured with an earth-based particle source, the aberration angle is $\vec{\theta}_{a}=-\vec{\theta}_{w}=-(1-1/\gamma)\vec{V}\times\vec{v}/V^{2}$, where $\vec{v}$ is the orbital velocity of the earth relative to the sun, $\vec{V}$ is the velocity of the accelerated electron, and $\gamma=1/\sqrt{1-V^{2}/c^{2}}$ is the particle relativistic factor. This means that the angle of Wigner rotation depends on $V^{2}/c^{2}$. It should be noted that the presented formula can be considered only as a first-order approximation in $v/c$. In the theory of (infinitesimal) Wigner rotation, we must also consider the relativistic parameter $V/c$. If a particle motion velocity is non-relativistic, the binomial expansion yields $\vec{\theta}_{a}=-[V^{2}/(2c^{2})]\vec{V}\times\vec{v}/V^{2}$.
Let us calculate the parameters for the STEM setup. The Wigner rotation angle parameter (constant of aberration) defines the scale of the aberration increment and is equal to $\theta_{w}=(1-1/1/\gamma)v/V\simeq 40$ $\mu\mathrm{rad}$. Here $\gamma\simeq 1.39$, $V/c=0.7$, $v/c\simeq 100$ $\mu\mathrm{rad}$. Earth moves around the sun, with a consequent change of the direction of $\vec{v}$. Therefore the angle of aberration changes. The total change of the order of $40\mu\mathrm{rad}$. It is oscillatory with a period of one year.
Figure 40: Definitions related observer’s position on the earth.
A specific aspect of our case study needs further investigation. The proposed method of measuring the angle of aberration involves the use of earth-based sources and has a big advantage. The rotation of the earth on its axis should produce a corresponding shift of the image. Obviously, it is important that observation should be recorded in the shortest possible time. In principle, records could be taken over a period of one day. The aberration shift depends only upon the value of $v_{\perp}$, the component of the orbital velocity perpendicular to the earth’s rotation axis 333333 The constant of diurnal aberration defines the scale of the aberration increment and is equal to $\theta_{d}=(1-1/1/\gamma)v_{e}/V\simeq 0.6$ $\mu\mathrm{rad}$. Here $\gamma\simeq 1.39$, $V/c=0.7$, $v_{e}/c\simeq 1,55$ $\mu\mathrm{rad}$ where $v_{e}=0.45$ km/s is the linear velocity of the earth rotation at equator.. In practice, the measurement procedure is complicated by a number of factors, Fig. 40. A correction has to be made for the observer’s position on the earth’s surface and the obliquity of the ecliptic $\delta=23^{o}$. The image appears to move in an ellipse. The shape of the aberration ellipse depends on the observer’s latitude $\phi$. For an observer on the equator, the ellipse degenerates into a line segment, and for an observer at the pole of the earth, the ellipse is a circle. At the value $l=2$ mm of the focal length  [^34] Electron-optical system of the STEM contains an electron gun and several magnetic lenses, stacked vertically to form a lens column. The illumination system of the instrument comprises the electron gun, together with two condenser lenses that focus the electrons onto the spacemen. Electrons entering the lens column appear to come from a virtual source. The first condenser lens (C1) is a strong magnetic lens, with a focal length $f_{1}$ of typically 2 mm. Using a virtual electron source as its object, C1 produces a real image. Because the lens is located $D_{1}=15$ cm below the object (virtual source), its object distance is 15 cm $\gg f_{1}$ and the image distance $\simeq f_{1}$. The second condenser lens (C2) is a weak magnetic lens ($f_{2}\simeq 10$ cm) that provides little magnification but allows the diameter of illumination at the specimen to vary continuously over a wide range. The distance between lens centers is typically 10 cm. Spacemen is located $D_{3}=25$ cm below the C2., the major axis of the aberration ellipse of the order of 80 ($\cos\delta)$ nm. The value of the minor axis depends on the observer’s position and is given by 80 ($\cos\delta\sin\phi))$ nm. When the latitude is $\phi=35^{o}$, the value of the scale of image shift per hour depends on the local time and varies from a value of about 10 nm/60 min to about 20 nm/60 min through the day  [^35] Experimental results (see Fig. 44 p.167 in [47]) shows that the image shift is quite close to the theoretical prediction $\simeq 10~{}\mathrm{nm}$. It is important to note that we used the typical focal length (2 mm) in our estimations. The details of the optical setup may influence the image shift. .
10.2 A Thought Experiment with a Point-Like Source
We now turn to a simple scaling model for stellar aberration, illustrated in Fig. 41. Consider a scale transformation involving both the linear dimensions of the source and the distance between the source and the observer. This approach allows us to derive a condition for optical similarity between the aberration of light from a distant star and that from an Earth-based source. Such a transformation can assist in the preliminary design of an Earth-based experimental setup, where an incandescent lamp is used in place of a star.
Scaling model for stellar aberration using an Earth-based setup with an incandescent lamp. Based on astronomical observations, we know what to expect even without invoking special relativity. Remarkably, by measuring the aberration of light alone—without referencing anything external to Earth—it is possible to determine the Earth’s orbital speed around the Sun.
Figure 41: Scaling model for stellar aberration using an Earth-based setup with an incandescent lamp. Based on astronomical observations, we know what to expect even without invoking special relativity.
A star ( or an incandescent lamp) is a system of elementary statistically independent point sources. Point-source field diffraction can be divided into categories - the Fresnel (near-zone) diffraction and Fraunhofer ( or telescopic) diffraction. If the Fraunhofer (or equivalently, far zone) approximation $z>D^{2}/\lambda$ is satisfied, then the quadratic phase factor is approximately unity over the detection aperture. Here $D$ is the aperture width, $\lambda$ is the wavelength of the radiation, and $z$ is the distance between the source and observer. In the far zone, a point source produces in front of pupil detection effectively a plane wave. The plane wavefronts of starlight entering the telescope are imaged by the lens to a diffraction spot that lies in the focal plane. It is worth noting that we consider the divergence of the transmitted radiation, $\lambda/D$, that is relatively small compared to the aberration angle, $\theta_{a}=v/c$.
An incandescent lamp is a completely incoherent source. The character of the mutual intensity function produced by an incoherent source is fully described by the Van Cittert-Zernike theorem [34]. In our case of interest, imaging system is situated in the far zone $D^{2}/(\lambda z)<1$. It is the source linear dimension $d$ (source diameter) that determines the coherent area of the observed radiation $(z\lambda/d)^{2}$. The condition for neglecting the transverse size of the source, $d$, formulated as the requirement that the change of the correlation function along the aperture be less than unity, that is $d<D$ 363636 In all cases of practical interest, telescopes are situated in the far zone of the distant stars. If the far zone approximation $z>d$ is satisfied, then the Van Cittert-Zernike theorem takes its simplest form. For example, the coherence area of the light emitted by the circular incoherent source is $A_{c}=4\lambda^{2}z^{2}/(\pi d^{2})$. The minimal coherent area of light observed on the earth’s surface from the nearest star has a diameter of about $\lambda z/d\simeq$ 10 m. This diameter is larger than a typical telescope aperture $D$. In this situation, a star can be considered as a point source and a star image is a point spread function in the image plane of a telescope. Our scaling model is based on the far zone approximation. We have only to change the distance and the source diameter. Assuming small parameter conditions $Dd/(z\lambda)<1$, $D^{2}/(z\lambda)<1$, the scaling approach yields the following requirements: $D^{2}/(z\lambda)<1$, $d<D$,.
Let us estimate the parameters for an Earth-based setup using an incandescent lamp. We aim to describe a series of thought experiments—hypothetical, yet physically plausible. At optical frequencies, the conditions required for the Fraunhofer approximation can be quite stringent. For instance, with a wavelength of $\lambda=0.3$ $\mu$m and an aperture width of $D=3$ cm, the observation distance must satisfy $z>3000$ m. Obviously, the required conditions are met in a number of important problems. No one has ever done such an experiment, but we know what would happen from the astronomical observations.
10.3 Practical Setup Using an Incoherent Optical Source
It should be noted that Fraunhofer diffraction patterns can be observed at distances much closer than implied by equation $z>D^{2}/\lambda$ provided the aperture is illuminated by a spherical wave converging towards the observer, or if a positive lens is properly situated between the observer and the aperture.
The operational principle of the Earth-based optical setup designed to detect light aberrations is schematically illustrated in Fig. 42. This configuration follows the well-known two-lens imaging scheme, which enables magnification by adjusting the focal length of the second lens. However, for simplicity, we will assume equal focal lengths for both lenses (i.e., 1:1 imaging) in the following discussion. The two-lens setup is commonly used in image processing applications, as it provides more flexibility for modifying images compared to a single-lens system. In the context of this two-lens arrangement, we now turn to the relatively straightforward task of characterizing the point source radiation in the image plane, assuming that the lenses depicted in Fig. 42 are identical.
To perform an aberration of light measurement, the source linear dimension $d$ should be radically decreased. We assume that the order of magnitude of the dimensions of the ”point” source is about $\lambda$, where $\lambda$ is the optical radiation wavelength. Within an elementary source volume ($\lambda^{3}$) there is an enormous number of atoms. In optics, the emission from such a ”point” source is described classically using Maxwell’s equations, while the emitting medium (considered as an ensemble of atoms) is treated quantum mechanically. Thanks to modern lithography techniques, it is relatively straightforward to create an unresolved point source at optical wavelengths, allowing for the generation of sufficiently bright radiation  [^37] It may be possible to use the masked spatially incoherent UV light source. One could use the light source onto a (micron diameter) hole of a specially designed mask. The mask should be placed at the distance $l<<d$, where $d$ is the transverse size of the spatially incoherent UV light source..
Consider a diffracting aperture that is circular, and let the radius of the aperture be $a$. The physical properties of the lens can be combined in a single number $f$ called the focal length. We assume that the image is produced by a diffraction-limited optical system (i.e. a system that is free from aberrations). Once the wavelength is fixed, the resolution only depends on the numerical aperture $NA=a/f\ll 1$ of the system. The response of the system at point $(x_{i},y_{i})$ of the image plane to a $\delta$ function input at coordinates $(x_{o},y_{o})$ of the object plane is called the point spread function of the system. In our case of interest, one obtains the following amplitude point spread function: $A(x_{i},y_{i})=J_{1}(\alpha)/\alpha$, which is the diffraction pattern of a circular aperture, where $\alpha=2\pi ar_{i}/(f\lambda)$, $x_{0}=0,y_{0}=0$, $r_{i}=(x_{i}^{2}+y_{i}^{2})^{1/2}$. The first zero of the Airy’s pattern is at $r_{i}=0.6\lambda/NA$.
Let us estimate the parameters for the optical setup shown in Fig.42. Suppose $a=1.5$ cm and $\lambda=0.3$ $\mu$m. We will use $NA=0.1$. The resolution analysis is reduced to the theoretical framework of standard imaging theory. One can take advantage of well-known resolution criteria like Rayleigh criteria i.e. $20.6\lambda/NA=$ 2 $\mu$m. One could use a vertically oriented optical setup. The constant of aberration defines the scale of the aberration increment, and is equal to $v/c=100$ $\mu$rad. Measurement of the spot shift with respect to the optical axis (which is parallel to the nadir direction) is equivalent to the determination of the angular displacement. As already pointed out, the proposed method of measuring the angle of aberration involves the use of earth-based sources and has a big advantage. The rotation of the earth on its axis should produce a corresponding shift of the image. In principle, records could be taken over a period of one day. The image appears to move in an ellipse. The major axis of the ellipse is about 14 $\mu$m. The value of the minor axis depends on the observer’s position and is given by $14(\sin\phi)$ $\mu$m  [^38] The small image size—approximately 2 $\mu$m in the numerical case considered—places significant limitations on the use of a CCD as a detector. To address this, one promising alternative is photon detection using a microchannel plate photomultiplier (MCP-PMT). A feasible approach involves using a commercially available precision 2D translation stage with a resolution of 0.5 $\mu$m to control the position of a submicron aperture in an opaque screen. For instance, the aperture diameter could be set to 1 $\mu$m or smaller. The screen is mounted horizontally on the translation stage, which is positioned in the image plane. The MCP-PMT can then detect photons transmitted through the aperture. This setup offers the advantage of compactness, with the overall height of the optical configuration being approximately one meter..
Figure 42: Earth-based optical setup for detecting the aberration of light. The setup employs a two-lens system to form an image using an incoherent light source at optical frequencies.
We derived our results under the assumption that the dimensions of the source are on the order of the wavelength, $\lambda$. It is important to note that the resolution of the proposed optical techniques is not fundamentally limited by the size of the incoherent source. The radiation field generated by an incoherent source can be modeled as a linear superposition of the fields from individual elementary point sources. The image of each elementary point source corresponds to a point spread function (PSF). An optical system illuminated with incoherent light can be fully described by a convolution integral that relates the object distribution to the resulting image distribution. In this framework, the optical system is characterized by its impulse response, namely, the point image. The rounding or softening of the source’s edges is a characteristic effect of the convolution (”smoothing”) process.
It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single-point source characteristics like oscillatory image shift within a period of one day. This argument gives the reason why our technique can be extended to allow for the reconstruction of the angle of aberration from a measurement of image shift of a finite size, incoherent source. The edge object has become useful in aberration shift detection. An alternative application of a ”point” source is an incoherent source with (physically) sharp edges. The edge width of the source is strictly related to the practical resolution achievable. For our particular example, the edge width should be about $1\mu$m.
11 The Aberration of Light from a Laser Source
11.1 Theory of a Laser with an Optical Resonator
Suppose an Earth-based observer attempts to measure the aberration of light using a laser as the light source. It is important to emphasize that the Earth’s orbital motion cannot be detected using a laser in this manner. To understand why, consider a simplified scenario in which the laser resonator consists of plane mirrors. For oscillation to occur within the resonator, the total power loss due to diffraction and reflection must be less than the power gained as light passes through the active medium. Diffraction losses, therefore, play a crucial role—not only in determining the threshold for oscillation but also in shaping the distribution of energy within the interferometer during operation. The aim of this study is to examine how diffraction affects the electromagnetic field in a Fabry–Pérot interferometer operating in free space. The results of this analysis are equally applicable to gas lasers, as long as the interferometer is immersed in the active medium.
The electric field $\vec{E}$ of an electromagnetic wave in a passive resonator satisfies the equation $\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial(ct)^{2}=0$. For simplicity, we shall assume that the field in the resonator is circularly polarized. Such an assumption does not violate the generality of the analysis because the polarization degeneracy takes place in the system under study. The electric field vector in the plane resonator is presented as an oscillation superposition with different longitudinal wave numbers, $E_{x}+iE_{y}=\sum_{m}\widetilde{E}_{m}(x,y,t)\exp(-i\omega_{m}t)\sin(K_{m}z+\delta)$, where $K_{m}=m\pi/l-i/(nl)$, $\omega_{m}=m\pi c/l$, $\delta=i/n$, $n$ is the (complex) refractive index of the mirror matter ($|n|\gg 1)$, $l$ is the distance between the mirrors, and $m$ is an integer number ($m\gg 1$). It is evident that this expression satisfies Leontovich’s boundary conduction  [^39] An example of introducing approximate boundary conditions is presented here. These conditions, originally derived by Leontovich, apply to surfaces of materials with a high refractive index modulus. A more comprehensive theoretical treatment can be found in the monograph [48]. on the mirror surface when $z=0$ and $z=l$. We assume the field change per one resonator pass to be small i.e. $|\partial\widetilde{E}_{m}/\partial t|l/c\ll|\widetilde{E}_{m}|$, which means that the laser frequency is close to the natural frequencies $\omega_{m}$ of the longitudinal resonator modes. Let us find a solution for the field amplitude $\widetilde{E}_{m}$ in the form $\widetilde{E}_{m}=\Phi(x,y)\exp(\Lambda t)$. Substituting the ansatz in the wave equation we get $\nabla_{\perp}\Phi+2i\omega_{m}[\Lambda/c^{2}+2/(nl)]\Phi=0$.
In a plane resonator, the diffraction effects at the edges of the mirrors can be taken into account if one considers the space between the mirrors as a waveguide, and uses diffraction theory at the open end of the waveguide. If almost an integer number of half waves is laid along the length between the mirrors then at the open resonator end we may put boundary conditions on the amplitude function $\Phi(x,y)$, equivalent by its effect to the open waveguide end action  [^40] The analytical method is based on the introduction of specific, complex boundary conditions known as impedance boundary conditions of the resonance type [49]. : $\Phi+(1+i)\beta_{0}\sqrt{cl/(4\omega_{m})}\partial\Phi/\partial\eta=0$, where $\eta$ is the direction of the normal to the imaginary side surface of the resonator, and the parameter $\beta_{0}=0.824$  [^41] Interestingly, $\beta_{0}$ is related to one of the most renowned mathematical functions—the Riemann zeta function $\zeta$. Specifically, it is given by $-\zeta(1/2)/\sqrt{\pi}$; see [49].. The boundary conditions do not depend on the polarization which means that polarization degeneracy occurs in the system under study. Therefore, the problem concerning the open plane resonator excitation using equivalent boundary conditions can be reduced to the more usual classical formulation of the problem of closed resonator excitation.
[…]
The rigorous results of the three-dimensional theory of a plane Fabry-Perot resonator demonstrate that the radiation within the resonator can be represented as a set of discrete modes. Each mode is characterized by a specific decay rate (decrement) and an eigenfunction that describes the field amplitude distribution along the mirror surface. For the eigenmode associated with the transverse wave number $k_{\nu j}$, the corresponding eigenvalue is given by: $\Lambda=-2c/(nl)-i\mu_{\nu j}^{2}(1-2\Delta)/(2R^{2})$. Due to the exponential dependence on $\mu_{\nu j}^{2}$, only the lower order modes tend to survive in the passive open resonator. We denote a normal mode supported by circular plane mirrors as a $TEM_{\nu j}$ mode, with $\nu$ denoting the order of angular variations ( the angular variation are sinusoidal in form) and $j$ denoting the order of radial variation. The fundamental mode, $TEM_{00}$, is identified as the dominant mode for resonators with circular plane mirrors.
11.2 Absent of Light Aberration from an Earth-Based Laser
We have shown that acceleration influences the field equations. Within the resonator, the electric field of an electromagnetic wave satisfies the wave equation, which transforms into Eq. (17) under a Galilean transformation of space and time coordinates. At first glance, measuring the diffraction losses of the dominant mode may seem equivalent to determining the angular displacement. However, it is important to recognize that the difference between the Langevin and Minkowski metrics does not affect the parameters of a laser incorporating an optical resonator.
In our earlier discussion of the Langevin metric, we demonstrated that applying a Galilean transformation to the wave equation yields a more complex, anisotropic form—namely, Eq. (17). This raises a fundamental question: how can we solve this form of the electrodynamics equation, particularly in the presence of boundary conditions? The key lies in recognizing that the underlying electrodynamics theory for optical resonators remains Lorentz-covariant. To address the anisotropic terms and simplify the field equations, we employ a mathematical transformation. By introducing a suitable change of variables, these anisotropic terms can be eliminated from the transformed equations, effectively restoring the Minkowski metric in the accelerated frame.
It is important to stress that this transformation is purely a mathematical convenience, not a reflection of any underlying physical reality. The variables used carry no intrinsic physical meaning; they serve only to streamline the problem-solving process. Crucially, the physical content of the theory remains unchanged—results expressed in the transformed coordinates are entirely equivalent to those in the original frame. Changing the four-dimensional coordinate system introduces no new physical effects.
A comprehensive analysis of experiments performed in an accelerated frame confirms that optical phenomena within a resonator are unaffected by acceleration relative to the fixed stars. This invariance arises because all radiation measurements in such systems are based on standing-wave phenomena—specifically, round-trip or two-beam interference methods.
Let us now consider the aberration of light emitted from a laser resonator at rest in an accelerated frame. Upon analyzing the data, the accelerated observer detects no shift in the direction of energy propagation, as illustrated in Fig. 43. This absence of aberration indicates that the cross-term in the Langevin metric, which would normally cause such an effect, is effectively averaged out over the complete round-trip trajectory of the radiation within the resonator.
Figure 43: Light aberration in an accelerated frame of reference, assuming the laser operates in a steady-state regime.
Figure 44: Light aberration from a plane-wave emitter in an accelerated frame of reference.
The distinction between a ”plane wave” emitter and a laser source leads to a striking prediction in the theory of light aberration. Specifically, if a ”plane wave” emitter initially at rest with respect to the fixed stars begins to move, the apparent angular position of the emitter—as observed from the accelerated frame—undergoes a sudden shift by an angle of $-v/c$ (see Fig. 44). Understanding this phenomenon requires the framework of special relativity; classical theory is insufficient.
But how can such a difference between a ”plane wave” emitter and a laser source arise? To answer this, we must consider: What happens to an accelerated optical resonator? In previous discussions of optical resonators, we applied electrodynamic boundary conditions in 3-space, which define the direction of energy propagation along the optical axis—perpendicular to the mirror surfaces. According to special relativity, acceleration does not alter the orientation of the mirrors relative to the axes of the accelerated frame (i.e., the 3-space of the observer). Thus, an observer co-moving with an accelerating laser source will find that the direction of energy transport (along the optical axis) remains unchanged relative to the observer’s own frame, regardless of the motion with respect to the fixed stars (see Fig. 43).
Special relativity tells us that both accelerated and inertial observers share the same absolute space-time but differ in their respective metrics, and therefore in their associated 3-spaces. In the case of an optical resonator, the direction of energy propagation is dictated by boundary conditions within 3-space. In contrast, the emission from an incoherent source—such as a ”plane wave” emitter—is influenced by initial conditions. For an accelerated observer, these conditions involve a nontrivial mixture of space and time coordinates, a consequence of the pseudo-gravitational field experienced during acceleration. The solution lies in using the absolute time coordinate. After acceleration, the wavefront of the light emitted by the incoherent source is perpendicular to the direction $z_{n}$. Applying this initial condition, together with the Langevin metric in an accelerating reference frame, allows us to fully explain the observed aberration of light from a ”plane wave” emitter (see Fig. 44).
11.3 A Moving Laser Source in an Inertial Frame of Reference
Let us consider the radiation emitted by an accelerated laser source in the inertial frame. The light aberration effect can be described within the framework of standard special relativity, utilizing the relativity of simultaneity. When a laser moves transversely (i.e., parallel to the mirrors), there is a deviation in the energy transport of the light emitted from the optical resonator. In the first-order approximation in $v/c$ the inertial observer would observe an angular displacement of $\theta_{a}=2v/c$, Fig. 45.
When discussing light aberration, it is important to distinguish between the effect from a ”plane wave” emitter and that from the laser source. Consider a ”plane wave” emitter that is accelerated from rest to velocity $v$ in a direction perpendicular to its optical axis  [^42] It is assumed that after acceleration, the laser moves at constant velocity and operates in a steady-state regime.. Suppose an observer, at rest with respect to the inertial frame of reference, measures the direction of light transport. How does the light beam from the moving ”plane wave” emitter appear to this observer? The inertial observer would find an angular displacement of $\theta_{a}=v/c$, as illustrated in Fig. 46.
Figure 45: Light aberration caused by a moving laser source in the initial inertial frame of reference. The laser resonator moves parallel to the mirrors, and it is assumed that the accelerated laser operates in a steady-state regime
Figure 46: Light aberration from a moving ”plane wave” emitter in the initial inertial frame of reference.
The first point to make regarding measurements in an inertial frame is that there is an intuitively plausible way to understand the aberration of light from a moving ”plane wave” emitter. An elementary explanation of this effect is well-known and can be understood in terms of the transformation of velocities between different reference frames. The aberration of light in the initial inertial frame can also be readily explained within the framework of the corpuscular theory of light. This is plausible when considering that a light signal represents a certain amount of electromagnetic energy. Like mass, energy is conserved, so a light signal shares several characteristics with material particles. Consequently, we should expect the group velocities of light signals to follow the same addition theorem as particle velocities. A more detailed treatment based on the wave theory of light confirms this expectation.
In the case of a moving laser source, intuition suggests that the aberration increment would be given by $\theta_{a}=v/c$. However, special relativity introduces an additional factor of 2. The standard derivation of the aberration of light effect does not account for the mixing of positions and time in an accelerated frame. It incorrectly assumes that accelerated and inertial observers share a common 3-space. In contrast, from the perspective of the initial inertial frame, the coordinate axes of the accelerated frame are rotated by an angle $v/c$ relative to those of the inertial frame.
12 Special Relativity and Reciprocal Symmetry
12.1 Pseudo-Gravitation Fields and the Langevin Metric
This chapter explores the nature of special relativistic effects. The dynamics, governed by field equations, are often obscured within the framework of relativistic kinematics. The metric structure of space-time is best understood as an interpretation of how dynamical matter fields are perceived by different observers. A notable concern in the field is the absence of a dynamical explanation for the Langevin metric within special relativity, which has unsettled some physicists. The asymmetry between inertial and accelerated frames can be interpreted as a manifestation of pseudo-gravity experienced by an accelerated observer.
The Langevin metric, which governs the transformation from an inertial coordinate system to an accelerated one (relative to the fixed stars), can be understood as arising from the action of a specific force—an issue typically addressed within the framework of general relativity. However, it would be erroneous to suggest that this provides a direct explanation for the effects discussed in earlier chapters. A central aim of this chapter is to show that non-inertial frames in pseudo-Euclidean space-time can be effectively described using the language of general relativity.
Suppose that the reference frame $S$ is at rest with respect to the fixed stars. Let $(ct,x,y,z)$ represent the Lorentz coordinates in the inertial frame $S$, while $S_{n}$ is a system initially at rest relative to $S$ that undergoes uniform acceleration along the $x$-axis, reaching a velocity $v$. The inverse Galilean transformation $x\to x_{n}+vt$, $t\to t_{n}$ defines a new reference frame, $S_{n}$, which is uniformly accelerated in the $x$-direction, attaining a velocity $v$ relative to $S$. After the acceleration, the frame $S$ is, reciprocally, moving along the $x_{n}$-axis with velocity $-v$ relative to $S_{n}$.
When the system $S_{n}$ begins moving with constant velocity, it becomes an inertial frame of reference. At first glance, there appears to be symmetry (or reciprocity) between the inertial frame $S$ and the inertial frame $S_{n}$. According to the principle of relativity (also known as the principle of reciprocity), an observer accelerated relative to the fixed stars should measure the Minkowski metric as $ds^{2}=c^{2}dt_{n}^{2}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$. This metric is reciprocal to the one measured by the inertial observer: $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$. So, where does the asymmetry arise? The analysis of the equivalence principle reveals that no additional complexities are required to explain the asymmetry between an inertial frame and an accelerated frame (relative to the fixed stars). The resolution of this paradox lies in identifying a pseudo-gravitational potential within the system $S_{n}$, which acts as the source of the asymmetry.
A uniformly accelerating frame can be treated as if it were an inertial frame, with the addition of a uniform pseudo-gravitational field. By ”pseudo-gravitational field,” we refer to an apparent field (not a true gravitational field) that exerts a force on all objects proportional to their mass. ”Uniform” means that the force experienced by each object is constant, regardless of its position. This concept lies at the heart of the equivalence principle. The principle of equivalence allows us to apply dynamical methods to solve problems involving non-inertial reference frames.
To minimize the mathematical complexity of the discussion, we will restrict our analysis to terms up to second order in $v/c$. Consider a reference frame $S$, in which each point undergoes constant acceleration $g=-|g|$ in the negative $x_{n}$-direction for a duration $T=v/|g|$. In this setup, the system $S$ is effectively in free fall within a pseudo-gravitational field. It becomes evident that the pseudo-gravitational acceleration is given by the gradient of a scalar potential, i.e., $g=-\partial\phi/\partial x_{n}$. A straightforward calculation shows that at time $t=T$, the coordinate transforms as $x_{n}=gT^{2}/2+x$ at $t=T$. Once the system $S$ transitions to uniform motion (constant velocity), the pseudo-gravitational field vanishes. To determine the pseudo-gravitational potential difference between the two frames, we compute $\phi=(\partial\phi/\partial x_{n})gT^{2}/2=-v^{2}/2$. Here, we have assumed that $\phi=0$ at $x_{n}=x$.
The analysis of an accelerated frame using the equivalence principle does not involve a real gravitational field and therefore does not require the full framework of general relativity. Nevertheless, the insights provided by general relativity about real gravitational fields apply in a limited sense to pseudo-gravitational fields as well. One key concept we rely on is that time runs more slowly as one moves deeper into the potential well of a pseudo-force field. This effect can be used to our advantage when analyzing the metric in a non-inertial frame.
The equivalence principle implies that gravity can alter the frequency of an electromagnetic wave and cause clocks to run at different rates. Specifically, the frequency of light increases with the absolute value of the gravitational potential. If a pulse of light is emitted at a point where the gravitational potential is $\phi_{1}$ with frequency $\omega_{1}$, then upon reaching a point where the potential is $\phi_{2}$, its frequency—measured in proper time at that location—will be: $\omega_{2}=\omega_{1}[1+(\phi_{1}-\phi_{2})/c^{2}]$.
Frequency is inversely proportional to the rate of local proper time; therefore, the gravitational frequency shift can be reformulated as a time dilation effect. The rate at which a coordinate clock ticks depends on the gravitational potential at its location. The relationship between coordinate time and proper (physical) time is given by: $dt^{(p)}=dt_{n}(1+\phi/c^{2})$. This implies that physical time elapses more slowly in regions with lower gravitational potential—i.e., where the absolute value of the potential is higher. If two identical clocks are placed in a gravitational field for some duration, the clock that remains deeper in the field will lag behind the other, appearing to run slower.
Let us now consider the situation where a reference frame, initially located at a point with gravitational potential $\phi_{1}$, undergoes free fall toward a region of lower potential $\phi_{2}$. Importantly, during fall, the frame retains the gravitational potential of its starting point, $\phi_{1}$. It is worth exploring what it means for the potential to remain ”unchanged” in this context.
Let us first examine the behavior of a clock in free fall. From the perspective of an observer located at $\phi_{2}$, the clock in the freely falling frame appears to tick faster, since it remains associated with the higher potential $\phi_{1}$, regardless of its instantaneous position. Radiation from a standard source within the free-falling frame will therefore appear blue-shifted relative to radiation from a standard source at rest at $\phi_{2}$ 434343 To understand the gravitational blue-shift of a light source in free fall, one must analyze the source’s emission mechanism and its behavior while in motion. The frequency of dipole radiation is proportional to the electron mass. In a gravitational field, mass varies with gravitational potential, according to the relation $m=m_{0}/\sqrt{1+2\phi/c^{2}}$. As a result, the electron mass is greater at points with lower gravitational potential..
A similar time dilation effect occurs in accelerated frames. For a clock at rest in a uniformly accelerated frame, the relation between proper time and coordinate time is also: $dt^{(p)}=dt_{n}(1+\phi/c^{2})=dt_{n}[1-v^{2}/(2c^{2})]$, where $dt_{n}$ is coordinate time in the accelerated frame $S_{n}$, but corresponds to proper time in the inertial frame $S$. From the viewpoint of the accelerated observer, clocks in the inertial frame appear to run faster due to their association with a higher effective (pseudo) potential.
According to the equivalence principle, the consistent aging of the inertial observer compared to the accelerated one arises naturally. The slower passage of time in the pseudo-gravitational field explains this outcome directly, requiring only basic algebraic reasoning.
Furthermore, it is important to note that the phenomenon of the relativity of simultaneity can be understood from a dynamical perspective. Specifically, the time-offset relationship between two spatially separated clocks can be interpreted as the accumulated time difference resulting from the pseudo-gravitational effects experienced by an accelerated observer. This interpretation implies that a clock positioned at a higher gravitational potential—i.e., farther along the direction of acceleration—runs faster than one at a lower potential. The pseudo-gravitational potential difference between two points along the $x_{n}$-axis is given by $\Delta\phi=\phi_{1}-\phi_{2}=(\partial\phi/\partial x_{n})[x_{n}(1)-x_{n}(2)]$. When the system $S$ transitions to uniform motion with constant velocity, the potential gradient in the comoving system $S_{n}$ becomes zero. The accumulated time difference between two spatially separated clocks is then $t^{(p)}(1)-t^{(p)}(2)=-g[x_{n}(1)-x_{n}(2)]v/(c^{2}|g|)=v[x_{n}(1)-x_{n}(2)]/c^{2}$.
The next aspect we must examine is the origin of the Lorentz deformation. Specifically, we observe that this deformation—referring to changes in relative position along the $x_{n}$ direction—is closely connected to shifts in the moments of acceleration. By employing the time-offset relation and integrating with respect to an infinitesimal change $dv$ we arrive at the expression $dl_{n}=dx_{n}[1+v^{2}/(2c^{2})]$, which characterizes the spatial geometry within the accelerated frame. This equation relates the length $dl=dx_{n}$ of the rod in the inertial frame $S$ to the corresponding length $dl_{n}$ in the accelerated frame $S_{n}$. Our analysis, which relies on standard measuring rods in the accelerated frame, demonstrates that the coordinate distance $dx_{n}$ acquires an effective length of $dx_{n}[1+v^{2}/(2c^{2})]$. This is a purely kinematic effect, arising without the involvement of any forces. We therefore conclude that the various relativistic effects are not independent phenomena, but are instead interconnected and can be comprehensively understood through a generalized principle of equivalence.
We now have all the quantities we set out to determine. Let us bring them together into the expression for the new space-time variables. The new independent variables $(x_{n}^{\mathrm{(p)}},t_{n}^{\mathrm{(p)}})$ can be expressed in terms of the original independent variables $(x_{n},t_{n})$ as follows: $ct_{n}^{\mathrm{(p)}}=t_{n}[1-v^{2}/(2c^{2})]-(vx_{n}/c^{2})[1+v^{2}/(2c^{2})]$, $x_{n}^{\mathrm{(p)}}=x_{n}[1+v^{2}/(2c^{2})]$. Above we made a simplification in calculating time and space transformations by considering only low velocities. Despite this approximation, the physical principles underlying the space-time transformation were clearly illustrated. The next natural question is: What is the form of the transformation for arbitrary velocity? The answer is straightforward: $t_{n}^{\mathrm{(p)}}=t_{n}\sqrt{1-v^{2}/c^{2}}-(vx_{n}/c^{2})/\sqrt{1-v^{2}/c^{2}}$, $x_{n}^{\mathrm{(p)}}=x_{n}/\sqrt{1-v^{2}/c^{2}}$. The quantity $dt_{n}^{\mathrm{(p)}}$ represents physical time, which is independent on the choice of coordinate time. It is important to note that $t_{n}$ in the expression for $t_{n}^{\mathrm{(p)}}$ corresponds to the coordinate time in the accelerated frame. Physical time governs the actual flow of events in physical processes. Likewise, the quantity $|dx_{n}^{\mathrm{(p)}}|$ represents the physical distance between two points in three-dimensional space. The space-time interval, when expressed in terms of physically measurable differentials, becomes a diagonal quantity. These measurable differentials, $dt_{n}^{\mathrm{(p)}}$ and $dx_{n}^{\mathrm{(p)}}$, are linear combinations of the coordinate time differential $dt_{n}$ and the spatial differential $dx_{n}$, with coefficients determined by components of the metric tensor. In the new physical coordinates, the space-time interval takes the form: $ds^{2}=(cdt_{n}^{\mathrm{(p)}})^{2}$ $(dx_{n}^{\mathrm{(p)}})^{2}$ $=(1-v^{2}/c^{2})c^{2}dt^{2}_{n}-2vdx_{n}dt_{n}-dx_{n}^{2}$.
The derivation of the Langevin metric in an accelerated frame—based on the concept of a pseudo-gravitational field—highlights a subtle but crucial distinction: the metric determines the form of physical laws, not the physical phenomena themselves. We interpret the Langevin metric to mean that the laws of electrodynamics, in the accelerated frame, take the form given by Eq. (17). By physical facts in this context, we refer to observable phenomena such as the aberration of light emitted by a single plane-wave source in the accelerated frame. To fully describe this aberration effect, the electrodynamics equation must be solved with an appropriate initial condition for the radiation wavefront. After applying an inverse Galilean boost—where the coordinates transform as $x_{n}=x-vt,\quad t_{n}=t$ —the wavefront is found to be perpendicular to the vertical direction $z_{n}$, as depicted in Fig. 15. Together, the Langevin metric and this initial condition account for the observed aberration of light in the accelerated frame.
12.2 The Equivalence Principle and Gravitational Physics
We should make one further remark regarding the equivalence principle. [^44] For a general discussion of the equivalence principle, we recommend reading [50]. It is a common belief that the equivalence principle forms the foundation of general relativity. However, this view is not entirely accurate. The gravitational field is not merely a universal force acting identically on all bodies. According to Einstein’s insight, gravity is intimately connected to the structure of space-time itself. To grasp this idea, we must first understand the role of the Riemann curvature tensor, which, in a precise sense, is the gravitational field. Only when this tensor vanishes can we say that no gravitational field exists. This observation alone suffices to demonstrate that the equivalence principle, as traditionally formulated, does not universally apply. The equivalence principle holds for mechanical processes but fails when applied to electrodynamical phenomena. This distinction implies that it is indeed possible to determine, through internal measurements, whether a system is in an inertial frame or in free fall within a homogeneous gravitational field. For example, in such a field, a free electric charge undergoes uniform acceleration relative to an inertial frame and, as a result, emits radiation. This radiation occurs regardless of whether the field is strong or weak. Conversely, in a uniformly accelerated frame (absent gravity), the free charge remains unaccelerated relative to the inertial frame and, according to classical electrodynamics, does not radiate. Hence, the equivalence principle is explicitly violated in this context.
In the context of special and general relativity, we can examine earth-based experiments where pseudo-gravity and real gravity are intricately combined. With this foundation, we now need to describe the gravitational fields near the Sun, arising from the Sun’s own mass. Consider the simplest case of a transformation from the initial inertial frame, in which spacetime is Minkowskian, to the Earth-based frame. Newton’s theory of gravity is known to hold when gravitational fields are weak enough that velocities do not approach the speed of light. Specifically, this means the gravitational potential energy of a particle is much smaller than its rest mass energy. In such scenarios, general relativity reduces to Newtonian gravity. For weak gravitational fields, the Newtonian potential $\phi$ fully determines the metric, which takes the form: $ds^{2}=(1+2\phi/c^{2})c^{2}dt^{2}-(1-2\phi/c^{2})(dx^{2}+dy^{2}+dz^{2})$. Here, $\phi$ is chosen to be negative so that, at large distances from the mass $M$, $\phi=-GM/r$. The condition for a weak field is that $|m\phi|<<mc^{2}$, which ensures that $|\phi/c^{2}|<<1$. This metric is only accurate to the first order in $\phi/c^{2}$. At large distances from the source of mass, we say that spacetime asymptotically becomes flat.
For a rotating reference frame, such as the Earth revolving around the Sun, the Langevin metric can be applied: $ds^{2}=(1-v^{2}/c^{2})d(ct)^{2}-2(\vec{v}\cdot d\vec{r})dt-(d\vec{r})^{2}$, where $\vec{v}$ is the velocity of the Earth, and $(d\vec{r})^{2}$ represents the square of the coordinate distance. The lack of equivalence between equations in an accelerated reference frame and those in an inertial frame with a static gravitational field becomes clear. In the former case, the non-diagonal metric coefficients depend on time, whereas in the latter, the metric remains diagonal and time-independent. This discrepancy explains why a charge at rest in an accelerated frame will radiate, while a charge at rest in a static gravitational field will not. This difference demonstrates why we cannot conclusively deduce the validity of the equivalence principle from observational data alone.
Now, returning to the perspective of an Earth-based observer, an additional expression for the invariant interval is required: $ds^{2}=(1+2\phi/c^{2}-v^{2}/c^{2})d(ct)^{2}-2(\vec{v}\cdot d\vec{r})dt-(1-2\phi/c^{2})(d\vec{r})^{2}$. At the order of approximation used, this result is simply the superposition of the Langevin metric and the Sun-centered metric in the weak gravitational field of the Sun. Both the Earth’s orbital speed and the Sun’s gravitational field contribute to time dilation and spatial contraction. However, the effects are small, making the Langevin metric an excellent approximation for the metric in the rotating Earth-based reference frame.
12.3 Time Dilation
It is commonly believed that special relativity is a reciprocal theory. However, an instructive counterexample to this reciprocity is provided by the time dilation effect. Consider the passage of time in two inertial reference frames. One frame, denoted by $S$, is considered to be at rest relative to the fixed stars. The second frame, $S_{n}$, undergoes acceleration relative to $S$ until it reaches a velocity $v$ along the $x$-axis. As previously discussed, the transformation from an inertial to an accelerated frame was apparently first introduced by Langevin. A central mathematical approach in the analysis of time dilation is the smooth matching—or tailoring—of the metric tensor. The transformation to the accelerated frame, with coordinates denoted by subscript ”n”, takes a Galilean form: $t_{n}=t$, $x_{n}=x-vt$ Following this transformation, the metric in the inertial frame transitions continuously into the Langevin metric of the accelerated inertial frame $S_{n}$ (see Chapter 5 for details).
In these new coordinates, the invariant interval given by Eq. (1) becomes Eq. (15). It is important to note that although coordinate clocks are fixed in the accelerated frame, they register the same time as clocks in the inertial frame. However, due to the presence of the cross-term $dt_{n}dx_{n}$, the clocks in the accelerated frame are not synchronized in the conventional (Einstein) sense—that is, by exchanging light signals between clocks and accounting for round-trip travel time.
Suppose we have two identical physical clocks located at the same point in an inertial reference frame $S$, and let their readings coincide at the initial moment $t=0$. The first clock remains at rest in frame $S$, while the second clock, at $t=0$, begins to move with constant velocity $v$ along the $x$ axis. Now, consider the situation from the perspective of the accelerated reference frame $S_{n}$ in which the second clock is always at rest. The two frames are not equivalent, as $S_{n}$ is accelerated with respect to the fixed stars. In this accelerated frame, the first clock moves with velocity $dx_{n}/dt_{n}=-v$. Taking into account the Langevin metric Eq.(15): $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}$, we find that the proper time interval for the first (moving) clock is $d\tau=ds/c=dt_{n}$. This implies that the time shown by the first physical clock (which is moving in the accelerated frame) coincides with the coordinate time $t_{n}$ in that frame. Since the second clock is at rest in the accelerated frame, its proper time evolves according to $d\tau_{n}=\sqrt{1-v^{2}/c^{2}}dt_{n}$. This result demonstrates the time dilation experienced by a physical clock at rest in the accelerated frame when compared to clocks in the inertial frame. Importantly, we observe that the slowdown of the accelerated clock does not depend on the choice of reference frame in which the effect is measured.
Let us now consider a different and very practical scenario. Suppose we split the accelerated frame into two parts. Frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ both accelerate relative to the inertial frame $S$, reaching the same speed $v$, but in opposite directions. From the perspective of the inertial frame $S$, the situation is perfectly symmetrical. Therefore, no difference in time dilation should exist between physical clocks at rest in $S_{n}^{(1)}$ and $S_{n}^{(2)}$, as compared to a clock at rest in $S$. The following simple analysis confirms this expectation. Assume we have two physical clocks: the first is at rest in frame $S_{n}^{(1)}$, while the second is at rest in frame $S_{n}^{(2)}$. From the point of view of $S_{n}^{(1)}$, the second clock appears to move with velocity $dx_{n}/dt_{n}=-2v$. Using the Langevin metric, $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}$, we calculate the proper time interval for the second clock as: $d\tau_{2}=ds/c=\sqrt{1-v^{2}/c^{2}}dt_{n}$. This result shows that the time measured by the first clock, $d\tau_{1}$, is equal to the proper time $d\tau_{2}$ of the second clock. This outcome may appear paradoxical at first. After all, $S_{n}^{(1)}$ and $S_{n}^{(2)}$ are moving relative to each other at a non-zero speed. Yet we observe an interesting and important fact: The mere presence of relative velocity does not necessarily lead to time dilation.
As discussed above, non-inertial frames can be effectively described using the framework of general relativity. The time dilation observed when transitioning from an inertial to an accelerated frame can be interpreted as the consequence of a specific force acting within the system. A resolution to the symmetry paradox reveals that a pseudo-gravitational field, present within the systems $S_{n}^{(1)}$ and $S_{n}^{(2)}$, serves as the agent responsible for the time dilation effect. Clocks measuring proper time and sharing identical histories of acceleration run at the same rates and produce consistent measurements, even when at rest in distinct inertial frames.
Let us now consider a different scenario. Suppose that frame $S_{n}$ accelerates relative to the inertial frame $S$ until it reaches a velocity $v$. Then, two new frames, $S_{n}^{(1)}$ and $S_{n}^{(2)}$, each accelerate relative to $S_{n}$ to the same speed $v$, but in opposite directions. From the perspective of $S_{n}$, this setup is not symmetric, and the resulting time dilation experienced in the accelerated frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ differs. In frame $S_{n}$, frame $S_{n}^{(1)}$ moves with velocity $-v$, so the proper time in this frame is simply $d\tau_{n}^{(1)}=dt=dt_{n}$. Meanwhile, frame $S_{n}^{(2)}$ moves with velocity $v$ relative to $S_{n}$. Using the Langevin metric, $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}$, we find that the proper time in frame $S_{n}^{(2)}$ is given by $d\tau_{n}^{(2)}=\sqrt{1-4v^{2}/c^{2}}dt_{n}$. This analysis has been carried out from the point of view of an observer at rest in frame $S_{n}$. It is natural to ask: how would the situation appear to an observer in the inertial frame $S$? From the perspective of frame $S$ the frame $S_{n}^{(2)}$ moves with velocity $dx/dt=2v$. Applying the Minkowski metric, we again obtain: $d\tau_{n}^{(2)}=\sqrt{1-4v^{2}/c^{2}}dt$, which is consistent with our previous result. It is worth noting that the mathematics involved in relating proper times is remarkably straightforward in this case. This simplicity arises because we are using Galilean boosts and the Galilean addition of velocities.
12.4 The Light-Clock and Observations by a Non-Inertial Observer
To understand how a clock slows down in an accelerated system, we need to examine the internal workings of the clock and observe what happens as it moves. We’ll explore this effect using a simplified example—a device known as a light clock. This clock consists of a meter stick with mirrors at both ends; a ”tick” corresponds to a pulse of light bouncing between the mirrors. Now, let us consider two configurations of the light clock: one with its length parallel to the direction of motion and the other oriented orthogonally. According to the principle of relativity, all clocks—regardless of their construction—must measure time consistently in all inertial frames. We will demonstrate that, regardless of the orientation of the light clock relative to its motion, the elapsed time measured remains the same. Two light clocks illustrating these configurations are shown in Fig. 47.
Figure 47: Parallel and orthogonal light-clock setups.
Now, let us examine the workings of a physical light clock under acceleration and observe what happens to it in such a scenario. We analyze the behavior of the clock from the perspective of an observer co-moving with the accelerated frame in which the clock resides. To begin, we configure the light clock so that the light beam travels between two mirrors aligned parallel to the direction of motion relative to the fixed stars. Within the context of the Langevin metric, the speed of light emitted by an accelerated source is no longer equal to $c$. Considering an infinitesimal displacement $ds$ along the worldline of a light beam, and using the condition $ds^{2}=0$, we derive from Eq. (15) that: $c^{2}=(dx_{n}/dt_{n}+v)^{2}$. This implies that, in the accelerated coordinate system $(ct_{n},x_{n})$, the velocity of light emitted in the direction parallel to the $x_{n}$-axis is: $dx_{n}/dt_{n}=c-v$ in the positive direction, and $dx_{n}/dt_{n}=-c-v$ in the negative direction. Let $l_{n}$ represent the (as yet undetermined) length of the rod or meter stick in this frame. The time it takes for light to travel from the left mirror to the right mirror is then: $t_{1}=l_{n}/(c-v)$, and for the return path, from the right mirror to the left: $t_{2}=l_{n}/(c+v)$. Therefore, the total time interval between the emission and reception of the light signal is: $\Delta t=t_{1}+t_{2}=2l_{n}/[c(1-v^{2}/c^{2})]$.
To begin, we must determine the length of the physical measuring rods. For measuring distances between fixed points in the accelerated frame, we use physical rods identical to those employed in the inertial frame, but now considered at rest relative to the accelerated frame. Regarding spatial distances, the guiding interpretive principle is that $\sqrt{-ds^{2}}$ represents the length of an infinitesimal rod whose endpoints are simultaneous according to standard simultaneity in the rod’s rest frame. Using the Langevin metric Eq.(15), we find that for $dt_{n}=0$, it follows that $dx_{n}=dx$. This implies that the length of a physical rod at rest in the inertial frame matches the coordinate distance in the accelerated frame. However, when we apply this principle to rods at rest in the accelerated frame, we face a complication: the condition $dt_{n}=0$ does not necessarily correspond to standard simultaneity in the accelerated frame. As a result, the length of a rod at rest in the accelerated frame generally differs from that of a rod at rest in the inertial frame.
The spatial geometry in an accelerated frame is described by the differential spatial line element defined within it. While spatial geometry possesses a degree of conventionality due to coordinate dependence, the proper spatial line element, denoted $dl_{n}$, holds particular significance, as it defines a coordinate-independent spatial geometry in the accelerated frame. The relation $dl_{n}^{2}=(-g_{\alpha\beta}+g_{0\alpha}g_{0\beta}/g_{00})dx_{\alpha}dx_{\beta}$ establishes the connection between the three-dimensional spatial line element and the four-dimensional spacetime metric. It can be shown that $dl_{n}$ remains invariant under coordinate transformations within the same reference frame. In contrast, the spatial line element $dl$, obtained by setting $dt_{n}=0$ in Eq.(15) with $dl^{2}=dx_{n}^{2}$, differs from $dl_{n}$.
Using Eq. (15), we obtain the relation $dl_{n}^{2}=dx_{n}^{2}/(1-v^{2}/c^{2})$, which defines the spatial geometry in the accelerated frame. This equation is asymmetric with respect to the lengths $dx_{n}=dl$ and $dl_{n}$, as it connects the physical length $dl$ of a rod measured in the inertial frame $S$ with the physical length $dl_{n}$ in the accelerated frame $S_{n}$. Our analysis, based on the use of standard measuring rods within the accelerated frame to probe its geometrical properties, shows that a coordinate distance $dx_{n}$ corresponds to a physical length of $dx_{n}/\sqrt{1-v^{2}/c^{2}}$. Thus, we can directly determine the asymmetry in the length of the meter stick from the metric. This can be interpreted as an effect of acceleration with respect to the fixed stars, suggesting that space in an accelerated frame is inherently non-Euclidean.
It is important to emphasize that the Langevin metric characterizes the measurements of an accelerated observer as interpreted by an inertial observer who employs the simplest form of clock synchronization—namely, maintaining the same set of synchronized clocks used when the light clock was at rest. This method preserves Einstein synchronization, established via light signals emitted from a stationary dipole source. Within this framework of absolute simultaneity, the inertial observer perceives the accelerated observer as measuring the speed of light along the $x$-axis as $c-v$ in the positive direction and $-c-v$ in the negative direction. Furthermore, the physical meter stick (with mirrors mounted on it), as measured by the accelerated observer, appears compressed along the direction of motion relative to the fixed stars.
It is crucial to note that the Langevin metric, in fact, reflects measurements performed by an inertial observer using the Minkowski metric, subsequently re-expressed via the transformation $x\to x_{n}+vt$, $t\to t_{n}$. The coordinate length $dx_{n}$ in the accelerated frame corresponds to a length measured by the inertial observer in the moving (accelerated) frame under the assumption of absolute time synchronization in the inertial frame. The length of a moving object thus depends not only on the spacetime structure but also on the synchronization convention adopted. Consequently, it is unsurprising that the measured length of a moving physical rod differs from its proper length when at rest.
As a result, the time interval between the emission and reception of the light signal in the parallel light clock within the accelerated frame is given by: $\Delta t_{n}=(2l/c)/\sqrt{1-v^{2}/c^{2}}$. Since $t_{n}=t$ and $\Delta t=2l/c$ represent the time indicated by a clock at rest in the inertial frame, this implies that physical clocks at rest in the accelerated frame tick more slowly compared to those in the original inertial frame (see the Section 13.1 for further details).
Let us consider a light clock that is accelerated to a velocity $v$, perpendicular to the direction of the optical pulse. In other words, we orient the light clock in such a way that the effects of length contraction—such as those observed in the case of a meter stick aligned with the direction of motion—do not arise. We analyze this orthogonal light clock from the perspective of an observer co-moving with the accelerated clock
If $ds$ is the infinitesimal displacement along the worldline of a light beam, then $ds^{2}=0$. Applying Eq. (15), we find that $(dz_{n}/dt_{n})^{2}=c^{2}(1-v^{2}/c^{2})$. Hence, the time interval between the emission and reception of the light signal is $\Delta t_{n}=2l/[c\sqrt{1-v^{2}/c^{2}}]$. Since $t_{n}=t$ and $\Delta t=2l/c$ represent the time indicated by a physical clock at rest in the inertial frame, this result shows that clocks at rest in the accelerated frame run slower compared to those in the original inertial frame.
12.5 Metrics in Accelerated Frames
Let us examine the implications of assuming that the inertial frame has no history of acceleration and that the space-time metric, as perceived by an observer at rest, is diagonal. We begin by considering two emitters within this inertial frame: the first is stationary, while the second moves with velocity $v$. Applying a Galilean boost yields the metric of the accelerated emitter in the inertial frame, as given in Eq. (11). In this coordinate system, the non-accelerated light source is described by the standard diagonal metric: $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$.
Let us now return to the discussion of the accelerated frame. In this frame, we observe a reciprocal velocity: the second emitter is at rest, while the first emitter moves with velocity $-v$. Now what about reciprocity in physics? According to conventional theory, acceleration does not inherently break the symmetry of motion between an accelerated frame and an inertial one. Consequently, an observer in the accelerated frame should still obtain a diagonal metric for the second emitter, which is at rest relative to this frame. By the principle of reciprocity, the metric given in Eq. (15) should also describe the aberration of light as observed for the first emitter, which appears to be moving in the accelerated frame. In fact, the metric of the moving emitter in the accelerated frame matches that of the moving emitter in the inertial frame, as given by Eq. (11). Reciprocity implies that the only difference lies in the sign of the velocity $v$. The principle of reciprocity states that all inertial frames are equivalent. However, this relativity principle does not hold and the Lorentz covariance of all physical equations does not dictate the reciprocal symmetry of nature.
We will examine the asymmetry between inertial and accelerated reference frames. A space-time geometrical analysis of the accelerated frame reveals a fundamental non-reciprocity between it and the initial inertial frame. Specifically, within the accelerated frame, we derive the Langevin metric, Eq. (15), associated with the second emitter. This metric governs the behavior of electrodynamics in the accelerated frame. Our analysis highlights a crucial asymmetry: Maxwell’s equations, which hold in inertial frames, are no longer valid from the perspective of an observer at rest in the accelerated frame. This breakdown reflects a deeper distinction between the two types of frames. Importantly, because we adopt an absolute time coordinatization, the initial conditions—such as the orientation of radiation wavefronts—remain identical in both frames  [^45] Many physicists contend that the universality of Maxwell’s equations across inertial frames rules out the existence of any privileged frame. As Dieks observes [51], “As we already mentioned, it is a basic principle of the special theory of relativity that the line element $ds$ supplies all information about the physics of the situation, as described in the given coordinates.” However, this widely held view is misleading. The metric alone does not fully specify the physical situation. To solve electrodynamics problems, one must also provide appropriate initial conditions..
We now turn to an apparent paradox. The principle that all inertial frames are equivalent is often supported by the following argument: To maintain Lorentz coordinates in an accelerated frame, one must perform a diagonalization of the metric tensor. In both the original and the accelerated frames, the metric appears diagonal, and standard treatments suggest that the coordinates in these frames should be related by a Lorentz transformation. At first glance, the situation appears symmetric—both frames seem equivalent. However, a crucial asymmetry emerges when we consider the physical distinction between an inertial frame and one that has undergone acceleration. Specifically, we should expect the metric tensor to reflect this difference. After the diagonalization procedure, the metric tensor in the accelerated frame must transition discontinuously—from a form such as $g_{00}=(1-v^{2}/c^{2})<1,g_{01}=-v/c$ to the standard inertial form $g_{00}=1,g_{01}=0$. This discontinuity implies that the Lorentz transformation does not preserve the smooth structure of the metric across frames. As a result, the symmetry between inertial frames breaks down in this context. This reveals a subtle yet significant distinction: while the mathematical form of the metric may appear similar in both frames, the physical histories of the frames—particularly the presence or absence of acceleration—lead to observable and meaningful differences. In this way, the accelerated frame retains a kind of “memory” of its non-inertial past, breaking the naive equivalence suggested by symmetry arguments.
13 Relativistic Measurements
13.1 The Concept of Time in Special Relativity
The theory of relativity reveals that our intuitive understanding of time does not align with its true nature. To clarify the concept of time, we adopt the principle that ”each physical quantity is defined by the method used to measure it.” Let us explore how this principle applies. Consider the case of length. We have a standard unit of length, and we can measure the length of any object accordingly. However, this measurement refers to objective physical length only when the object is at rest. For moving objects—such as a rod in motion—length becomes a convention-dependent quantity. It relies on the chosen method of synchronization and lacks an exact, objective meaning. With this in mind, we can now turn to the problem of defining time within the framework of special relativity.
Here is a simple example that illustrates the concept clearly. Suppose we know the law governing muon decay in its rest frame. When we apply a Lorentz transformation to this decay law, we find that, in the laboratory frame, the characteristic lifetime of the particle appears to increase from $\tau_{0}$ to $\gamma\tau_{0}$. This result can be interpreted as follows: after traveling a distance of $\gamma v\tau_{0}$, the number of muons observed in the lab frame is reduced to half of the original population. In this context, the time measured corresponds to a ”length,” as defined by the earlier metric discussion. This leads to an intriguing question: could this interpretation apply to all time measurements within the framework of special relativity? We argue that the answer is yes.
Our next example concerns frequency measurements. Consider a Fabry-Perot interferometer. In this setup, the frequency of light is effectively measured through the ”length” of the standing wave it produces. Frequency measurements in such cases inherently follow the principles of interference. Another example that illustrates the connection between frequency and interference is the grating spectrometer. Here, the frequency is inferred from the position of the light spot along the dispersion direction. Although their designs differ, there is no fundamental physical distinction between a diffraction grating and an interferometer—both rely on the interference of incident and reflected light waves to perform frequency measurements.
Let us now explore how to compare the clock rates between the initial inertial frame $S$ and the accelerated frame $S_{n}$. It is generally understood that when comparing clock rates between two frames $S$ and $S_{n}$ in relative motion, one cannot simply compare the readings of a single clock in each frame. This is because clocks from different frames coincide at the same spatial point only once. Therefore, in at least one of the frames, multiple clocks—assumed to be synchronized—must be used for meaningful comparison.
We have already pointed out that, within the framework of special relativity, all time measurements ultimately reflect what is essentially a measurement of ”length.” When it comes to measuring the length of objects at rest, we are dealing with an objective physical quantity—one that is independent of any particular convention. Since our empirical access is limited to such length measurements, it is difficult to accept the standard textbook claim that one can only compare the reading of a single clock in one frame with the readings of multiple synchronized clocks in another. Let us now apply our operational interpretation of time measurements to the example of a light clock, in order to see how this approach functions in practice.
Let us consider a specific thought experiment. We examine a situation in which time marks can be imprinted on a moving object (a screen) by a clock. To begin, we describe an orthogonal light clock from the perspective of an observer in the same accelerated frame as the clock. The screen is at rest in the initial inertial frame. Suppose the top mirror of the light clock is semitransparent, allowing time marks to be recorded on the (moving) screen, thereby measuring the time intervals between clock ”clicks.” In the case of the Langevin metric the speed of light emitted transversely by the accelerated source is given by $c\sqrt{1-v^{2}/c^{2}}$. Accordingly, the time interval between the emission and return of the light signal is $\Delta t_{n}=2l/[c\sqrt{1-v^{2}/c^{2}}]$. During this interval, the object (i.e., the screen in the initial inertial frame) falls a distance $\Delta x_{n}=\Delta x=2lv/[c\sqrt{1-v^{2}/c^{2}}]$, which corresponds to the spacing between consecutive time marks.
Taking into account the Langevin metric (Eq. 15), we find that at $dt_{n}=dt=0$, it follows that $dx_{n}=dx$, i.e. the length of the physical rod in the inertial frame coincidence with the coordinate distance in the accelerated frame. Since $t_{n}=t$ and $\Delta t=2l/c$ represent the time indicated by a physical clock at rest in the initial inertial frame, it follows that clocks at rest in the accelerated frame run slower compared to those in the inertial frame. In fact, the rate of the accelerated clock as measured by the inertial observer corresponds to the spacing between time marks on the screen in the initial inertial frame.
Second, we describe the orthogonal light clock from the perspective of an observer who is initially at rest in the same inertial frame as the clock. According to this observer, the time interval between the emission and reception of the light signal is given by $\Delta t=2l/c$. During this interval, the screen in the accelerated frame falls a distance $\Delta x=\Delta x_{n}=2lv/c$, which corresponds to the spacing between the time marks.
An analysis using standard measuring rods in an accelerated frame to examine its geometrical properties reveals that the coordinate distance $dx_{n}$ corresponds to a physical length of $dx_{n}/\sqrt{1-v^{2}/c^{2}}$. This relationship directly reflects the asymmetry in the length of the meter stick, as derived from the Langevin metric. This result indicates that the physical spacing between time marks, as measured in the accelerated frame, experiences compression in the direction of motion relative to the fixed stars. Since $\Delta x_{n}=\Delta x$, and $\Delta x=2vl/c$ represents the physical time rate shown by a clock at rest in the inertial frame, it follows that the accelerated observer perceives a shortening of the mark spacing, given by $\Delta l_{n}=2lv\sqrt{1-v^{2}/c^{2}}]/c$. Here $\Delta l_{n}$ denotes the physical spacing of time marks in the accelerated frame. This thought experiment clearly demonstrates the absence of reciprocity in time dilation.
13.2 Interference Phenomena
We now turn our attention to a specific aspect of the phenomenon of interference. To explore this, let’s examine the processes of reflection and refraction within the framework of special relativity. At this point, we are prepared to analyze what happens when light emitted by a moving source passes through a stationary medium, such as glass. The key task is to calculate the interference between the incident and reflected radiation. At first glance, one might assume that calculating both refraction and reflection in a single inertial frame would require accounting for differences in the metrics, and consequently the electrodynamics equations, for a moving emitter and a stationary glass. This suggests that the possibility of calculating interference effects is not immediately apparent.
Figure 48: The light emitted by a moving source reflects off a stationary mirror in a collinear arrangement. Interference occurs between the incident radiation and the radiation resulting from the material.
If the mirror is stationary and the emitter is in motion, it is evident that the equations governing electrodynamics must be the same for all electromagnetic waves. In other words, a consistent metric must be applied to both the incident and scattered waves. However, this raises a paradox: how can the electrodynamics equations be identical for both the incident and reflected waves, given that they are governed by two different metrics?
In a previous discussion, we addressed the issue of two distinct metrics in the context of light aberration. In Chapter 4, we focused on the first-order approximation and highlighted that the apparent paradox is resolved. One aspect of the aberration setup can be easily understood based on concepts we have already explored. It is advantageous to begin with the simpler case of a normally incident plane wave, and only later in Section 5.10 do we introduce the theory that includes tilted incoming plane waves. This approach is beneficial because the unique geometry of light aberration ensures that, even after applying a Galilean transformation, the electrodynamics equations in an absolute time coordinate system will retain Maxwell’s form for a normally incident plane wave. Now, let us examine the situation in a collinear geometry, as illustrated in Figure 48.
We begin by summarizing the central idea. The apparent paradox is resolved by recognizing that both the Minkowski metric and the metric given by Eq. (11) yield identical predictions. The key point is that all methods used to measure interference—specifically, those involving standing waves—are effectively round-trip measurements. In contrast, the deviation in the direction of energy transport arises from a geometric effect. Since empirical measurements grant access only to the round-trip (two-way) average speed of light, the one-way speed of light remains a matter of convention without intrinsic physical meaning. The two-way speed of light, however, is directly measurable and therefore possesses physical significance.
We now turn to an important aspect of interference phenomena. It is commonly believed that, in a setup involving a moving light source and a stationary glass medium, the incident wave and the wave scattered by the dipoles of the glass cannot interfere—as required by the electron theory of dispersion—because they travel at different velocities in an absolute time framework. This notion, however, is incorrect. Physically, an incident wave of a given frequency will excite electrons in the glass to oscillate at that same frequency, regardless of the wave’s velocity. These oscillating electrons then re-emit radiation at the same frequency. Thus, the incident and scattered waves at any given location share the same frequency and can interfere. The difference in their velocities manifests as a relative phase shift that varies spatially. This phase variation influences both the velocity and amplitude envelope of the resulting wave formed by the superposition of the incident and scattered components.
The following simple analysis supports these ideas. We consider the case where the metric is non-diagonal, as given in Eq. 11. Let the incoming wave be represented by $\exp[i(\omega t-kx)]$, where the phase velocity is $\omega/k=c+v$. Similarly, let the scattered (outgoing) wave be represented by $\exp[i(\omega t+k^{\prime}x+\phi)]$, having the same amplitude and frequency, but a different velocity $\omega/k^{\prime}=c-v$, and a phase offset $\phi$. The superposition of these two waves is given by: $\exp i(\omega t-kx)+\exp i(\omega t+k^{\prime}x+\phi)=2[\cos[(k+k^{\prime})x/2+\phi/2]]\exp i[\omega t-(k-k^{\prime})x/2+\phi/2]$. Here, the cosine term acts as an amplitude envelope that is stationary in space. Its periodicity is inversely proportional to the difference in the wave numbers $k$ and $k^{\prime}$ of the two components. This expression can be further simplified by substituting the wave numbers in terms of $\omega$, $c$, and $v$, leading to: $2[\cos[\omega x/[c(1-v^{2}/c^{2})]+\phi/2]]\exp i[\omega t-xv\omega/[c^{2}(1-v^{2}/c^{2})]+\phi/2]$.
Consider a source at rest emitting waves with a natural frequency $\omega_{0}$. In the laboratory frame, after applying a Galilean transformation, the velocity of the incoming wave becomes $c+v$. Given this, the frequency observed in the lab frame is: $\omega=\omega_{0}/[1-v/(c+v)]=\omega_{0}(1+v/c)$. This frequency shift corresponds to the well-known Doppler effect. Incorporating this into the expression for the superposition of two waves, we obtain: $2[\cos[\omega_{0}x/[c(1-v/c)]+\phi/2]]\exp i[\omega_{0}(1+v/c)t-xv\omega_{0}/[c^{2}(1-v/c)]+\phi/2]$.
Suppose an observer in the laboratory conducts a standing wave measurement. It is important to analyze which aspects of the measured data depend on the chosen synchronization convention and which remain invariant. We emphasize that time oscillation lacks intrinsic physical meaning—its interpretation arises solely from a chosen convention. In particular, consider the time shift $xv\omega_{0}(1+v/c)/c^{2}$ appearing in the phase term $\exp i[\omega_{0}(1+v/c)t-xv\omega_{0}/[c^{2}(1-v/c)]+\phi/2]$. This shift is closely tied to the issue of synchronizing distant clocks. Furthermore, from a physical standpoint, the absolute scale of time—or equivalently, frequency—cannot be directly identified.
Suppose the laboratory observer measures the wavelength of a standing wave. The relation $dl_{i}^{2}=(-g_{11}+g_{01}g_{01}/g_{00})dx^{2}$ describes the connection between the spatial line element and the metric. Using Eq. (11), we find $dl_{i}^{2}=dx^{2}/(1-v^{2}/c^{2})$, which defines the spatial geometry in the context of the Langevin metric. This result reflects the Lorentz contraction of measuring rods. When analyzing the measured data, the laboratory observer finds that the wave number is given by $\sqrt{1-v^{2}/c^{2}}\omega_{0}/[c(1-v/c)]=\omega_{0}\sqrt{1+v/c}/[c\sqrt{1-v/c}]$. Notably, this is the same factor one obtains under the assumption of a diagonal metric. In other words, the laboratory observer will measure the same two-way speed of light regardless of the choice of metric representation.
The reason a laboratory observer measures the same wavelength lies in the underlying principle of interference. All interference phenomena are governed by a fundamental law known as the relativistic invariance of phase. This principle states that a specific quantity—the phase—remains unchanged under transformations between reference frames. The invariance of phase is a powerful tool for analysis, as it allows us to draw meaningful conclusions (as will be illustrated in the next section) without requiring detailed knowledge of the full set of transformation formulas. However, a puzzle remains in the scenario under discussion.
Let us consider a dipole source that accelerates in the laboratory inertial frame to a velocity $v$ along the $x$-axis. After acceleration, assuming synchronization in the lab frame remains unchanged, we encounter a complex situation involving the dynamics and electrodynamics of moving charges.
In conventional (non-covariant) particle tracking, the particle’s motion can be modeled as a sequence of infinitesimal Galilean transformations, with a corrected form of Newton’s second law applied at each step. Interestingly, this approach does not invoke Lorentz transformations. Within the inertial lab frame, the particle’s motion appears exactly as classical mechanics would predict—following absolute time, without invoking the relativity of simultaneity. That is, time and position remain distinct and do not mix in this description.
In contrast, Maxwell’s equations are valid in the inertial frame only when Lorentz coordinates are used. This raises a fundamental question: How can classical mechanics (with Galilean kinematics) be used alongside electrodynamics governed by Einstein’s kinematics? This apparent inconsistency calls for explanation.
The resolution lies in the nature of the coupling between particles and fields in collinear geometries. In such configurations—where both the particle motion and the emitted radiation are aligned along the same axis—the standard Maxwell equations and the corrected Newtonian dynamics are sufficient to explain interference phenomena within a single inertial frame.
In this context, the dynamical evolution of the particle can be alternatively interpreted as a sequence of infinitesimal Lorentz transformations, rather than Galilean ones. But how is this possible? As discussed in Section 3.6, the key insight lies in the fact that Galilean boosts form a commutative group—just as collinear Lorentz boosts do. This commutative property plays a critical role in the analysis.
In the case of collinear geometry, where the particle moves along the same line as the emitted radiation, the velocity vector is perpendicular to the radiation’s wavefront (or the plane of simultaneity). In this configuration, the motion of the source influences only higher-order kinematic terms—those proportional to powers of $v/c$.
At first glance, Newtonian particle dynamics in collinear geometry appears to be entirely non-relativistic. However, relativistic effects do manifest, albeit subtly, through the implicit assumption that the mass of the moving particle corresponds to its relativistic mass.
14 A Preferred Inertial Frame and Second Order Experiments
14.1 Second-Order Optical Experiments and Special Relativity
An illustrative example of time dilation is observed in the resonance absorption of gamma rays in Mössbauer rotating disk experiments. When measuring the resonance absorption for the same relative velocity between the source and the observer, a blue shift is detected when the source is positioned at the center of the disk, while a red shift appears when the source is located at the rim. This result reveals an asymmetry in the transverse Doppler effect [^46] The first experimental confirmation of time dilation via resonance absorption of gamma rays was reported by Champeney, Isaak, and Khan in 1965 [52].. Notably, when both the source and observer are situated at the rim, no frequency shift is observed. In such a configuration, a pulse emitted by one accelerated observer on the rotating disk, which appears redshifted at the center due to time dilation, is received by a diametrically opposite observer on the rim without any frequency change. This occurs despite both observers moving at the same speed $v$ in opposite directions in the laboratory frame. Their mutual time dilation effects effectively cancel the frequency shift in this scenario.
At first glance, the transverse Doppler effect in an accelerated frame reveals a fundamental asymmetry between accelerated and inertial observers. This leads to the question: is it experimentally possible to determine the state of motion of an inertial frame $S$ from the perspective of an accelerated frame $S_{n}$, using the second-order Doppler redshift and blueshift of moving sources? Specifically, can this effect be used to determine the motion of the Earth relative to a Sun-centered inertial frame?
Let us examine the resonance absorption of gamma rays in a Mössbauer rotating disk experiment. Consider a setup where the absorber is located at the center of a disk, which remains at rest relative to the Earth-based frame, while the emitter is positioned on the edge of the disk. The emitter thus moves in a circular path around the absorber with a constant tangential velocity $\vec{w}$. To keep the mathematical complexity manageable, we restrict our analysis to terms up to second order in $w/c$ and $v/c$, where $v$ is the orbital velocity of the Earth. The frequency of the emitter changes with the velocity $\vec{w}$ in a manner as it suggests itself from the transverse Doppler effect. Indeed, the inner frequency of the emitter depends on its velocity $\vec{w}$, so that (using Langevin metric in the Earth-based frame) $\omega(t_{n})=\omega_{0}\sqrt{1-|\vec{v}+\vec{w}|^{2}}$. Expanding this expression and neglecting higher-order terms, we obtain $\omega(t_{n})=\omega_{0}[1-v^{2}/(2c^{2})-w^{2}/(2c^{2})-\vec{v}\cdot\vec{w}(t_{n})/c^{2}]$. Since the physical clock is at rest in the earth-based frame, the reading of its proper time is $d\tau=(1-v^{2}/2)dt_{n}$. We therefore expect that radiation emitted with a constant intrinsic frequency $\omega_{0}$ will arrive at the absorber with a time-dependent frequency, given by $\omega(\tau)=\omega_{0}[1-w^{2}/(2c^{2})-\vec{v}\cdot\vec{w}(\tau)/c^{2}]$. Because the emitter moves in a circular path, the scalar product $\vec{v}\cdot\vec{w}(\tau)$ varies periodically, leading to periodic fluctuations in the observed frequency at the disk center. Consequently, the resonance absorption measured at the absorber will also vary periodically with the emitter’s position along the disk’s circumference.
In the actual experiment, no effect on absorption was observed when the disk was set into rotation. This negative result can be explained by considering that only a portion of the total frequency fluctuation is initially taken into account. A more refined analysis reveals the necessity of including the frequency shift caused by the so-called radial Doppler effect. We analyze a configuration based on perpendicular geometry, where, at first glance, the radial Doppler effect would appear to be absent. However, as discussed in previous chapters, the aberration of light in the Earth-based reference frame implies that the radial component of the emitter’s velocity is given by $\vec{w}\cdot\vec{v}/c$. As a result, the emitted radiation reaching the center of the disk exhibits a periodically varying frequency shift $\delta\omega=\omega_{0}\vec{v}\cdot\vec{w}(\tau)/c^{2}$, which exactly matches the periodic variation in the internal resonance frequency. Thus, the two effects — the intrinsic frequency variation and the radial Doppler shift — cancel each other out, leading to no net observable effect.
The second-order Doppler redshift observed in moving atoms is a key prediction of special relativity. The first experimental verification of the transverse Doppler effect was conducted in 1938 by Ives and Stilwell [53]. In their experiment, ionized hydrogen molecules were accelerated in a cathode tube to energies of up to 28 keV. The frequencies of light emitted in both the forward (parallel) and backward (antiparallel) directions relative to the motion of the particles were measured using a stationary spectrograph. Let us now analyze this experiment. We discuss a setup based on a parallel geometry. Radiation emitted by the moving atoms, exhibiting Doppler-shifted spectral lines, was observed alongside radiation from stationary atoms present in the same working volume, which produced an unshifted reference line. By including this reference, Ives and Stilwell effectively circumvented the challenging task of directly measuring the asymmetry between the red- and blue-shifted lines. Instead, they compared both to the stationary reference line, simplifying the analysis.
At first glance, the Earth’s orbital motion relative to the Sun can be examined through this type of experiment. [^47] In the Ives and Stilwell experiment, the effect arising from the Earth’s orbital motion relative to the Sun was within the limits of experimental accuracy. This limitation is due to the high velocity of the emitter ($w\gg v$). In this discussion, we present a general analysis of such experiments, independent of specific experimental limitations. Specifically, if the internal frequency of an atom depends on its velocity $\vec{w}$, then (using the Langevin metric in the Earth-based frame) we have $\omega(t_{n})=\omega_{0}\sqrt{1-|\vec{v}+\vec{w}|^{2}}$, which implies that the frequency of radiation recorded by a spectrograph should, in principle, depend on the Earth’s orbital velocity $\vec{v}$. In rotating disk experiments, the observable effect due to the transverse Doppler shift is counteracted by another relativistic effect. For the case of perpendicular geometry, we have previously explained the experimental results by accounting for both the aberration of light and the radial (i.e., classical) Doppler effect. In contrast, in the collinear geometry, the aberration of light is absent, and only the radial Doppler effect needs to be considered. By analyzing the geometry of this configuration, we find that the observed frequency (i.e., the frequency of the radiation reaching the detector) is given by $\omega=\omega_{0}[1-v^{2}/(2c^{2})-w^{2}/(2c^{2})+\vec{v}\cdot\vec{w}(t_{n})/c^{2}]/[1-w/(c+\vec{w}\cdot\vec{v}/w)]$, where, according to the Langevin metric, the coordinate velocity of light in the Earth-based frame is $c^{*}=c+\vec{w}\cdot\vec{v}/w$. Thus, just as in the case of the rotating disk experiment, we find that in measurements of the second-order Doppler redshift for moving atoms, the two effects caused by the Earth’s orbital motion effectively cancel each other. As a result, no net observable effect remains.
Such cancellations are typically understood to arise from deep underlying principles. Given the series of negative results—like the one obtained by Michelson and Morley—it seems reasonable to assume that the consistent failure of various optical experiments to detect the Earth’s orbital velocity $v$ is no coincidence. This has led to the widely accepted view that a fundamental law of nature prevents the determination of $v$ through laboratory-based experiments. However, in the present case, no such profound implication appears to be at play. This outcome reflects a specific limitation of second-order optical experiments. As discussed in earlier chapters, we already know that it is possible to detect the Earth’s orbital velocity—without referencing anything external—by using the phenomenon of light aberration.
14.2 Features of Second Order Optical Experiments
The principle of relativity has been experimentally confirmed, though only to a limited extent. In second-order experiments, the effects caused by Earth’s orbital motion cancel each other out, leaving no observable impact. This suggests that second-order experiments inherently prevent us from determining Earth’s velocity relative to the Sun through laboratory measurements. Let us examine this claim more closely.
The key effects under consideration involve the interference of two light beams. A light source is split into two beams using a suitable apparatus, and their overlap produces an interference pattern. Since both beams originate from the same source, second-order interference experiments cannot detect Earth’s orbital velocity. This is because phase is a four-dimensional invariant—independent of the chosen inertial frame and coordinate system—due to the fundamental geometry of space-time.
This discussion shows that special relativity correctly predicts the observed null fringe shift in Earth-based frames. One of the most well-known interference experiments that demonstrates this is the Michelson-Morley experiment. Next, we turn to the second-order Doppler shift in moving atoms. The frequencies of emitted light were measured using a grating spectrograph. Notably, there is no fundamental distinction between a diffraction grating and an interferometer in this context.
An intriguing example of this effect is the measurement of resonance absorption of gamma rays in a Mössbauer rotating disk experiment. To illustrate this effect, consider a source positioned at a great distance from a thin absorber plate. We aim to determine the resulting field at a distant point on the opposite side of the absorber. According to electrodynamics, the total electric field is the sum of contributions from the external source and from the fields generated by charges within the absorber. The absorber consists of atoms with nuclei, and when the external field acts on these nuclei, it induces motion in their charges. Moving charges generate additional fields, acting as new radiators. Consequently, the two plane waves that interfere propagate in the same direction, leading to interference between the incident and coherently forward-scattered waves. Our model of the nuclear oscillator includes a damping force. With damping taken into account, the refractive index becomes complex, resulting in destructive interference. Thus, this too is an interference phenomenon.
14.3 Wigner Rotation and the Trouton-Noble Experiment
Next, we will explore additional second-order experiments that are not directly related to light propagation but can be understood through general dynamical principles. One of the most significant of these is the Trouton-Noble experiment [54]. The Trouton-Noble experiment aims to detect the rotational motion of a charged parallel-plate capacitor suspended at rest in the Earth’s frame. This motion is used as a means to measure the Earth’s velocity relative to the Sun (i.e., relative to the fixed stars). The fundamental concept behind the experiment can be described as follows.
Consider two opposite point charges $+e$ and $-e$ in the Earth-based frame; the radius-vector pointing from $-e$ to $+e$ be denoted by $\vec{r}_{n}$. If the charges are at rest the force acting upon $+e$ can be written $\vec{F}_{n}=e\vec{E}=-e^{2}\vec{r}_{n}/r^{3}_{n}$. As the force acts in the direction $\vec{r}_{n}$ the moment of force produced by the pair of charges vanishes, $\vec{M}_{n}=\vec{r}_{n}\times\vec{F}_{n}=0$.
If the pair of charges is made to move with a velocity $\vec{v}$ in the Sun-centered frame then the positive charge will be under the action of the Coulomb attraction of $-e$ and also under the influence of the magnetic field $\vec{B}=e(\vec{v}\times\vec{r})/(cr^{3})$. Thus the total force acting upon $e$ is given by $\vec{F}+e^{2}(\vec{v}\times(\vec{v}\times\vec{r}))/(c^{2}r^{3})$. Since $\vec{v}\times(\vec{v}\times\vec{r})=\vec{v}(\vec{v}\cdot\vec{r})-v^{2}\vec{r}$ we find that the moment of force produced by the pair of charges is equal to $\vec{M}=e^{2}(\vec{v}\cdot\vec{r})(\vec{r}\times\vec{v})/(c^{2}r^{3})$. Denoting the angle between $\vec{v}$ and $\vec{r}$ by $\theta$, we find for the absolute value of moment of force $M=e^{2}v^{2}\sin{\theta}\cos{\theta}/(c^{2}r)$.
In the actual experiment, a charged condenser was suspended from an elastic string. It was positioned such that the angle $\theta=\pi/4$; that is, the line perpendicular to the surface of the condenser plates formed an angle of $\pi/4$ with the assumed direction of the Earth’s orbital velocity, denoted by the vector $\vec{v}$. The experiment yielded negative results—no effect proportional to $v^{2}/c^{2}$ was observed, contrary to expectations.
The absence of torque on the capacitor in the Earth-based frame is fully consistent with special relativity and can be explained in a straightforward manner. The electrodynamic equations in the Earth-based frame can be derived by applying a Galilean transformation (with velocity $-v$) to Maxwell’s equations. Assuming the magnetic field $\vec{B}=0$ in the original frame, the transformed fields are given by $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$, where $\vec{v}$ is the velocity of the Earth relative to the Sun-based frame. Consequently, the induced magnetic field in the Earth-based frame does not exert any torque on the capacitor, which is at rest in that frame.
The presence of a magnetic field gives rise to a magnetic force, which, in turn, exerts a torque on a charge in the sun-based inertial frame. This leads to an apparent paradox: the mechanical equations governing force and torque appear to differ between inertial frames. Specifically, a torque—and thus a time-dependent change in three-dimensional angular momentum—exists in one inertial frame but not in another that is in relative motion  [^48] In [55], Jackson discusses the apparent paradox that different mechanical equations for force and torque govern the motion of a charged particle in different inertial frames. A similar paradox arises in the standard explanation of the Trouton–Noble experiment. In the usual approach [55], this paradox stems from the hidden assumption that all inertial frames share a common three-dimensional space. However, we demonstrate that a relativistically accurate description of the Trouton–Noble experiment can be achieved only by taking into account the Wigner rotation effect.. We argue that the root of this paradox lies in a flawed assumption: that an Earth-based observer and a Sun-based observer share a common three-dimensional space. According to special relativity, observers following different worldlines (i.e., with different trajectories through spacetime) possess different three-dimensional hypersurfaces of simultaneity—different 3-spaces. The standard treatment of the Trouton-Noble experiment implicitly assumes that the coordinate axes $(x_{n},y_{n},z_{n})$ of the moving Earth-based observer are parallel to the axes $(x,y,z)$ of the stationary Sun-based observer. We demonstrate that a consistent explanation, fully aligned with special relativity, emerges only when the equations of motion are expressed in covariant form. This approach properly accounts for the relativistic transformation of coordinate systems and resolves the apparent contradiction.
The key to resolving the paradox commonly associated with explanations of the Trouton–Noble experiment lies in recognizing that spatial measurements made by an Earth-based observer involve a mixing of space and time coordinates, as viewed from the Sun-based (inertial) frame. While it may seem intuitive to assume that the angle between the velocity vector $\vec{v}$ and the radius vector $\vec{r}_{n}$ in the Earth frame should match the angle between $\vec{v}$ and $\vec{r}$ in the Sun frame, this is not actually the case. The apparent rotation of the radius vector about the velocity vector is intrinsically linked to the phenomenon of length contraction. Consider a rod aligned along the $x$-axis in the Earth-based frame. Under a Lorentz boost along the $x$-axis with velocity $v$, the rod appears contracted in the Sun-based frame to a length of $l_{n}/\gamma$, where $l_{n}$ is the rod’s proper length in the Earth frame and $\gamma$ is the Lorentz factor. The relationship between the angle $\theta_{n}$ in the Earth frame and $\theta$ in the Sun frame is given by $\gamma\tan\theta_{n}=\tan\theta$. Assuming $v^{2}/c^{2}\ll 1$, this leads to the expression $\Delta\theta/\cos^{2}\theta=-[v^{2}/(2c^{2})]\tan\theta$, where $\Delta\theta=\theta_{n}-\theta$ represents the rotation angle. This angle is further related to the system parameters by $\Delta\theta=-[v^{2}/(2c^{2})]\sin\theta\cos\theta$. Consequently, the torque (moment of force) exerted by a pair of charges in the Earth-based frame is found to be $M_{n}=e^{2}v^{2}\sin\theta\cos\theta/(2c^{2}r)$.
But we’re still not done! We need to apply an additional correction to the moment of force in the Earth-based frame. A more careful analysis reveals that we must account for the change in the perpendicular component of the velocity, $v_{\perp}=v\sin\theta$, due to the Wigner rotation. Since the force acts along the direction of $\vec{r}$, there exists a component of the charge’s acceleration that is perpendicular to the velocity $\vec{v}$. This component results in a rotation of the Earth-based (proper) frame axes as seen from the Sun-based frame. The rate at which the Wigner rotation occurs is given by: $d\vec{\Phi}/dt=[v^{2}/(2c^{2})](\vec{v}/v^{2})\times d\vec{v}/dt$. The direction of this rotation of the Earth-based frame (in the Sun-based frame) is the same as the direction of the rotation of the velocity vector in the Sun-based frame. We already have all the expressions we need. For the rate of change of the angle $\theta$, we find: $d\theta/dt=[v^{2}/(2c^{2})]e^{2}\sin\theta/(mvr^{2})$, where $m$ is the mass of the charge. Thus, the second correction to the moment of force is: $mrdv_{\perp}/dt=mv\cos\theta d\theta/dt=e^{2}v^{2}\sin\theta\cos\theta/(2c^{2}r)$. As a result, the net moment of force produced by a pair of charges in the Earth-based frame vanishes.
14.4 Time Dilation in Moving Atomic Clock
The next topic we need to address is the time dilation effect. We have already discussed this subject in Chapter 12. Here’s a brief summary of our findings:
1. We established that the slowing down of a physical clock that is accelerated with respect to the fixed stars does not depend on the choice of the reference frame in which the effect is measured. In other words, the time dilation effect is absolute, not relative.
2. Due to the asymmetry between the initial inertial and accelerated frames, special relativity makes a striking prediction. Suppose frame $S$ is at rest relative to the fixed stars, and frame $S_{n}$ is accelerated with respect to $S$ until it reaches a speed $v$. Now consider two clocks in frame $S_{n}$ both symmetrically accelerated to speed $v$, but moving in opposite directions. Surprisingly, the time dilation experienced by these two clocks in $S_{n}$ is not the same. This discrepancy arises because the clocks have different acceleration histories relative to the fixed stars.
A natural question arises: is it possible to experimentally determine the state of motion of the inertial frame $S$ from the perspective of the accelerated frame $S_{n}$, using the time dilation effect observed in a moving (atomic) reference clock? At first glance, it may seem that the Earth’s orbital motion could be detected through time dilation experiments. While this is theoretically possible, a more detailed analysis is required to properly address the question.
The proper time interval can be associated with phenomena such as particle lifetimes, atomic transition periods, and nuclear half-lives. The key point, however, is that this interval is entirely determined by the motion of a clock within its initial inertial frame.
In 1971, Hafele and Keating conducted a landmark experiment in which four atomic clocks were synchronized with a laboratory reference clock and then flown around the Earth aboard commercial aircraft—two traveling eastward and two westward  [^49] This experiment provided the first experimental confirmation of time dilation as predicted by special relativity, using atomic clocks in circular motion. An important follow-up came from the Global Positioning System (GPS), which also involves relativistic effects. GPS satellites, numbering around 80, orbit the Earth at velocities of approximately 4 km/s. The system demonstrates that satellite clocks tick more slowly due to their relative motion, with a measured discrepancy of about 7.1 microseconds per day. Interestingly, while there is a noticeable time difference between the clocks on Earth and those in orbit, no asymmetries are observed among the satellites themselves. This suggests that relative velocity alone does not account for time dilation. Furthermore, the experiment revealed no effect of Earth’s orbital velocity on clock rates, a result attributed to the fact that both the satellites and Earth-based clocks share a common orbital path around the Sun, rendering orbital effects unobservable in this context [56, 57].. Upon return, the airborne clocks were compared with the laboratory reference clock. The results showed that the clocks carried on the planes had run slower than those left stationary on Earth. The observed discrepancies matched the predictions of special relativity, confirming the role of relative acceleration in time dilation. Notably, the experiment found no measurable influence of Earth’s orbital motion on clock rates, reinforcing the idea that time dilation effects are tied to local motion relative to the chosen inertial frame.
[…]
15 The Principle of Relativity and Modern Cosmology
In the preceding discussion, we explored the fundamental concepts necessary for understanding the phenomenon of the aberration of light. At first glance, the source-observer asymmetry associated with stellar aberration appears to contradict the principle of relativity. This principle, first introduced by Poincaré—who coined the term—states that an observer in uniform motion relative to the fixed stars cannot, by any internal measurement, detect this motion, provided they do not observe external celestial objects.
Now, let us examine the stellar aberration experiment more closely. One might argue that in the case of stellar aberration, the key distinction is that the accelerated (Earth-based) observer does, in fact, observe external celestial objects—the stars. However, the motion of the stars relative to Earth is never accompanied by any aberration. The aberration shift, as inferred from astronomical observations, remains even when a star moves with the same velocity as Earth. Thus, when considering Earth-based observations of stellar aberration, we cannot rely solely on the notion of ”looking outside to the stars.”
In Chapter 10, we introduced a simple scaling model for stellar aberration and derived a condition for optical similarity between the aberration of light from a distant star and that from an incoherent Earth-based source. This analysis revealed that even without observing external celestial objects, it is possible to determine Earth’s orbital velocity around the Sun through aberration of light measurements.
There is no fundamental conflict between the structure of special relativity and the aberration of light phenomena. As discussed in previous chapters, special relativity does not require the principle of the irrelevancy of velocity formulated earlier. The principle of special relativity applies to physical laws, not to physical facts. It asserts that the same laws must hold in all inertial frames, which we interpret to mean that these laws should be expressed by equations that retain the same form in all inertial frames.
In special relativity, the transformations preserving this form-invariance between inertial frames are the Lorentz transformations  [^50] In the special theory of relativity the form-invariant transformations between inertial frames are Lorentz transformations [58]. For a general discussion about the Principle of Relativity we suggest reading the paper [59]. . According to the special principle of relativity, all inertial frames are equivalent concerning physical laws, but not necessarily concerning physical facts.
The existence of absolute acceleration in nature implies the existence of absolute velocity.
A closer examination of the problem of time dilation in an accelerated frame reveals the fundamental asymmetry between an accelerated frame and an initial inertial frame (i.e., one without a history of acceleration). The proper time of any particle moving arbitrarily within the initial inertial frame always runs slower than the physical time of that frame. This is a fundamental characteristic of pseudo-Euclidean space-time.
Uniform motion is not relative in this framework, allowing us to determine a particle’s absolute velocity within the initial inertial frame. The formalism of classical physics relies on the structure of absolute space, which, in turn, must be associated with some underlying ”substance”—the ether—which serves as an absolute rest frame.
Could space-time itself be the new ether? It shares similar properties, notably providing an ”absolute rest reference frame.” The key distinction between the initial inertial frame and Newtonian absolute space lies in the shift introduced by special relativity: we now conceive of a unified space-time model, without an inherent separation between space and time.
While the initial inertial frame is not physically identical to absolute space, it represents the concept of absolute space-time.
The principle of relativity has a long history in physics. In Newtonian mechanics, physical laws remain invariant under Galilean transformations. However, from a mathematical perspective, the Lorentz transformation is fundamentally different from the Galilean transformation. Notably, Lorentz boosts alone do not form a group, whereas the set of Galilean boosts does. This reflects the essence of Galilean relativity: all inertial frames are equivalent, meaning that no one frame is preferred over another. The underlying reason for this symmetry is that the equations of motion in Newtonian mechanics do not explicitly depend on velocity, ensuring that a system’s internal dynamics remain unchanged under a Galilean boost.
Historically, Galileo’s relativity principle was mistakenly assumed to apply universally across all of physics. However, the principle of relativity does not hold across the entire domain of Lorentz-covariant physical laws, and Lorentz covariance itself is not a fundamental symmetry of nature. Since the space-time metric tensor is a continuous quantity, an accelerated inertial frame is connected to its initial state by a Galilean transformation rather than a Lorentz transformation. This distinction is particularly evident in electrodynamics: Galilean transformations do not preserve the form-invariance of Maxwell’s equations under a change of inertial frames. The reason is that electrodynamic equations explicitly depend on velocity, leading to the breakdown of Galilean relativity in this context.
According to special relativity, a fundamental difference exists between an accelerated inertial frame (relative to the fixed stars) and an inertial frame without an acceleration history. This distinction is closely linked to inertial forces and forms the core idea of the equivalence principle. Understanding the relationship between inertia and gravity remains an open question in fundamental physics.
Now we want to discuss a serious difficulty associated with the ideas of modern cosmology. The difficulty we speak of is associated with the concept of source-observer asymmetry when applied to the motion relative to the cosmic microwave background (CMB). One may say that perhaps there is no use worrying about these difficulties since there are so many things about the universe that we still don’t understand. We live in a universe governed by the unknown. Dark energy and dark matter together make up nearly 96 percent of the universe’s total energy-mass, yet their true nature remains a mystery. However, it is essential to recognize that the foundational assumptions of modern cosmology are only approximations. As our understanding deepens, these assumptions will inevitably evolve.
We will not delve into the principles of modern cosmology at this time but will assume their validity and focus on discussing some of their implications.
According to the cosmological principle, the universe should appear isotropic — lacking any preferred direction — to a comoving observer with no peculiar motion relative to the cosmic fluid of the expanding universe. However, if such an observer has a peculiar motion, it can introduce a dipole anisotropy in the observed properties of certain objects. This anisotropy can, in turn, be used to infer the observer’s peculiar velocity.
A well-known example of this phenomenon is the anisotropy observed in the cosmic microwave background (CMB) radiation. The CMB dipole is almost universally attributed to kinematic effects, specifically the Doppler shift caused by the relative motion between the Earth (the observer) and the reference frame in which the CMB appears nearly isotropic. By subtracting this dipole component, the CMB rest frame is established as the fundamental reference frame of the universe.
The observed dipole indicates that the solar system is moving at 370 km/s relative to the observed universe in the direction of galactic longitude $l=264^{o}$ and latitude $b=48^{O}$. This is quite far from the galactic rotation direction of 250 km/s towards $l=90^{o}$ and $b=0$. The motion relative to the cosmic microwave background results from the sum of many components of velocity due to the gravitational attraction of various mass concentrations. The existence of clusters and super clusters of galaxies and our motion is a natural consequence of the large-scale organization of matter. The peculiar velocity consists of the five vector contributions: the motion of the earth around the sun (30 km/s), the hypothetical circular motion around our galaxy (250 km/s), the motion of our galaxy in the local group, and the motion of the local group with respect to the cosmic microwave background. The origin of the velocity of the local group is still uncertain and has been under discussion over the past two decades. This peculiar velocity is believed to be generated by the spatial inhomogeneities of mass (mainly dark matter) distribution in nearly large-scale structures. The velocity of the local group with respect to the cosmic microwave background is estimated to be 500 km/s.
Now let us see how aberration of light varies with the speed relative to the rest frame of the universe. Consider first the textbook explanation. It is very easy to understand how the aberration of light effect comes about from the point of view of the conventional aberration of light theory. The existence of an aberration of the line of sight by motion can be recognized by considering an observer in a car driving through a rainstorm: raindrops falling vertically appear to be oblique. An important application of the Galilean velocity transformation law is provided by stellar aberration, the change in the apparent direction of a star caused by the Earth’s motion around the Sun. The apparent positions of all fixed stars are thus always a little displaced in the direction of the earth’s motion at that moment, and hence describe a small elliptical figure during the annual revolution of the Earth around the Sun. What about the motion relative to the CMB rest frame? According to textbooks, if the Earth’s motion were uniform, the aberration effect would be undetectable since the ”true” direction of the star is unknown. Indeed, who can say where a given star should be? On the other hand, the true direction of an Earth-based source is known. But, according to conventional theory, the aberration of light phenomenon does not exist in an aberration of light experiments using an Earth-based light source.
We are now in a position to understand the nature of our difficulty with the motion relative to the CMB. It is believed that the dipole anisotropy is produced by the Doppler effect due to acceleration (during the billions of years) of the solar system with respect to the rest frame of the universe. This acceleration is believed to be generated by the spatial inhomogeneities of (Dark matter) mass distribution in nearly large-scale structures. A correct solution of the aberration problem in the Earth-based frame requires the use of metric tensor even in first-order experiments since the crossed term in the Langevin metric plays a fundamental role in the non-inertial kinematics of a light ( or relativistic particle ) beam produced by the Earth-based source. According to the relativistic theory of aberration presented in this book, the proper rotation of the Earth on its axis should produce a corresponding shift of the image. The aberration shift will depend also upon the value of $v_{\perp}^{CMB}$, the component of the solar system velocity (relative to the CMB) perpendicular to the Earth’s rotation axis. The rotation of the Earth produces aberration in an amount larger enough to be taken into account in precise observation work using an electron microscope as an Earth-based particle source. Experimental results show (see Chapter 10) that the image shift is quite close to the theoretical prediction for the (30 km/s) orbital velocity and clearly indicate that the signal associated with motion (370 km/s) relative to the CMB does not exist. The simplest explanation is that the CMB dipole might be of non-kinematical origin.
More recent observations by astronomers also cast doubt on the CMB dipole, being the ultimate representative of the solar peculiar velocity. In recent years observations have emerged hinting at an anisotropic universe. The discovery of the preferred direction in the universe was serendipitous. Until very recently the velocity of the solar system in the rest frame of the universe is inferred from CMB temperature dipole anisotropy. Obviously, an independent measurement of this velocity is needed to fully establish the kinematical origin of the dipole. Another such quantity that could be employed to look for departures from isotropy is the angular distribution of distant radio sources in the sky. This could provide an independent check on the interpretation of CMB dipole. The radio data clearly indicate that significantly larger dipoles exist in the rest frame of the radio galaxies. While the velocity of the solar system inferred from the CMB temperature dipole anisotropy is 370 km/s, the radio dipole measurements find the speed of motion to be around 1000 km/s (i.e. to be around three times larger than that of CMB). From the Hubble diagram of quasars, the motion of the solar system is derived, which is out to be the largest value ever found, $8\cdot 10^{3}$ km/s  [^51] Recent observations of the cosmic dipole anisotropies is revolutionized our understanding of the universe. The various dipoles, including CMB dipole, all pointing along the same direction, suggest a preferred direction in the universe, raising thereby uncomfortable questions about the cosmological principle, the basis of the standard model in modern cosmology [60, 61, 62]..
On the other hand, a common direction for all these dipoles, determined from completely independent surveys by different groups employing different techniques, indicates that these dipoles are not resulting from some systematics in the observations or the data analysis, but could instead suggest an inherent anisotropy. This is totally unexpected in a standard model of the universe. We have a preferred direction, aligned with the CMB dipole, in the universe. That is, going to the CMB rest frame, we see an anisotropic background. There is a difficulty with such some sort of an ”axis” of the universe which, in turn, would be against the cosmological principle. Three independent dipole vectors pointing along the same particular direction could imply an anisotropic universe, violating the cosmological principle, a cornerstone of modern cosmology.
Let us be conservative and say that there are two kinds of the Earth velocity. The total velocity with respect to the initial (i.e. privilege ) inertial frame could be the sum of the orbital velocity and the velocity of the Sun. The velocity of the Sun consists the two contributions: the circular motion around our galaxy and the motion of our galaxy. In Earth-based experiments where we measure the Earth’s velocity with respect to the privileged frame by seeing the aberration of light (or particles), we are measuring the recorded (in the accelerational history) velocity. So the velocity with respect to the privilege frame consists of two contributions: a recorded motion plus an unrecorded motion. We know that there is definitely a recorded (orbital) motion. It is therefore impossible to get all the earth’s velocity to be recorded in the accelerational history in the way we hoped. Clearly something else has to be added.
Let’s take a closer look at why the Earth moves relative to the initial inertial frame. For the Earth to be in motion, a force must have acted upon it at some point. To deepen our understanding, it is useful to examine the origins of these forces in greater detail. We must say immediately that orbital motion is the result of non-gravitational forces. It is well known that a full 98 percent of all the angular momentum in the solar system is concentrated in the planets, yet a staggering 99.8 percent of all the mass in our solar system is in our sun. Perhaps the first scientifically respectable theory of the origin of the solar system was given by Hoyle (1960) [63]. He invoked the action of the magnetic field to transfer angular momentum from the central body, the Sun, to the ejected matter which eventually formed the planets. In contrast, the unrecorded motion around our galactic and the motion of our galactic is the result of the action of gravity.
We began by discussing the gravitational interaction between the Earth and the Sun. Our confidence in the theory of gravity is so strong that we use it to describe the forces acting between entire galaxies. However, we may be making too great an extrapolation, applying our limited understanding of gravity to cosmic scales. Perhaps the entire difficulty is that a modification of gravity may be responsible for unrecorded velocities. Given the many unresolved mysteries surrounding dark matter, dark energy, and cosmic anisotropy, a deeper exploration of fundamental physics is essential to uncovering the true nature of the universe.
16 Relativistic Particle Dynamics
16.1 Manifestly Covariant Formulation
The equations of dynamics can be formulated as tensor equations within Minkowski space-time. Upon selecting a coordinate system, these equations may be expressed in terms of components rather than abstract geometric objects. Leveraging the geometric structure of Minkowski space-time, one can identify the class of inertial frames and, for any such frame, adopt a Lorentz frame with orthonormal basis vectors. In any Lorentz coordinate system, the law of motion takes the form:
$m\frac{d^{2}x_{\mu}}{d\tau^{2}}=eF^{\mu\nu}\frac{dx_{\nu}}{d\tau}~{},$ (21)
where here the particle’s mass and charge are denoted by $m$ and $e$ respectively. The electromagnetic field is described by a second-rank, antisymmetric tensor with components $F^{\mu\nu}$. The coordinate-independent proper time $\tau$ is a parameter describing the evolution of the physical system under the relativistic laws of motion, Eq. (21).
The covariant equation of motion for a relativistic charged particle, subject to the four-force $K_{\mu}=eF^{\mu\nu}dx_{\nu}/d\tau$, as given in Eq. (21), represents a relativistic extension of Newton’s second law. While the classical form $md\vec{v}/dt=\vec{f}$ remain valid in the particle’s instantaneous comoving Lorentz frame, the relativistic formulation embeds these three spatial equations into the four-dimensional Minkowski spacetime  [^52] The notion of embedding relies on the fact that Newton’s second law can always be applied in any Lorentz frame in which the particle is instantaneously at rest [64]. That is, if a comoving Lorentz frame is defined at a given instant, the particle’s evolution can be accurately predicted in this frame over an infinitesimal time interval. Geometrically, Newton’s law holds rigorously on a hyperplane orthogonal to the particle’s world line. As the particle moves, the hyperplane—and its normal vector $u_{\mu}$—tilts accordingly along the world line. To formalize this embedding, one introduces a projection operator $\hat{P}_{\perp}$ that continuously projects vectors in Minkowski space onto hyperplanes orthogonal to the world line. This operator is given by $(\hat{P}_{\perp})_{\mu\nu}=\eta_{\mu\nu}-u_{\mu}u_{\nu}/u^{2}$ [64]. The presence of $\hat{P}_{\perp}$ highlights that only three of the four components of the covariant equation are independent. .
The immediate generalization of $md\vec{v}/dt=\vec{f}$ to an arbitrary Lorentz frame is given by Eq. (21), as can be verified by reducing the equation to the particle’s rest frame. In Lorentz coordinates, the four-velocity $u^{\mu}=dx^{\mu}/d\tau$ satisfies the kinematic constraint $u^{\mu}u_{\mu}=c^{2}$. Due to this condition, the four-dimensional equation of motion, Eq. (21), contains only three independent components. Using the explicit form of the Lorentz force, one can show that Eq. (21) automatically preserves the constraint $u^{\mu}u_{\mu}=c^{2}$, as required. To demonstrate this, we take the scalar product of both sides of the equation of motion with $u_{\mu}$. Exploiting the antisymmetry of the electromagnetic field tensor, $F^{\mu\nu}=-F^{\nu\mu}$, we find $u_{\mu}du^{\mu}/d\tau=eF^{\mu\nu}u_{\mu}u_{\nu}=0$. This result implies that the quantity $Y=(u^{2}-c^{2})$ is conserved, since $dY/d\tau=0$.
16.2 Conventional Particle Tracking
Having expressed the equation of motion in 4-vector form, Eq. (21), and determined the components of the 4-force, we have, on the one hand, ensured compliance with the principle of relativity and, on the other hand, obtained a complete set of four equations describing particle motion. This represents a covariant, relativistic generalization of Newton’s three-dimensional equation of motion, formulated with the particle’s proper time as the evolution parameter.
We now aim to describe the motion of a particle in the Lorentz laboratory frame, using the lab time $t$ as the evolution parameter. To begin, we consider the standard treatment found in textbooks. Let us determine the spatial components (i.e., the first three components) of the four-force. For this purpose, we examine the spatial part of the dynamical equation, Eq. (21): $\vec{Q}=(dt/d\tau)d(m\gamma\vec{v})/dt=\gamma d(m\gamma\vec{v})/dt$. The prefactor $\gamma$ appears due to the change in the evolution parameter from the proper time $\tau$—natural for covariant formulations—to the lab frame time $t$, which is necessary for introducing the conventional three-force vector $\vec{f}$. In this context, the spatial part of the four-force is related to the three-force by $\vec{Q}=\gamma\vec{f}$. Explicitly, the relativistic form of the three-force is given by
$\frac{d}{dt}\left(\frac{m\vec{v}}{\sqrt{1-v^{2}/c^{2}}}\right)=e\left(\vec{E}+\frac{\vec{v}}{c}\times\vec{B}\right)~{}.$ (22)
The time component of the four-force equation is
$\frac{d}{dt}\left(\frac{mc^{2}}{\sqrt{1-v^{2}/c^{2}}}\right)=e\vec{E}\cdot\vec{v}~{}.$ (23)
The evolution of the particle is governed by these four equations, together with the invariant constraint
$\mathcal{E}^{2}/c^{2}-|\vec{p}|^{2}=mc^{2}~{}.$ (24)
According to the non-covariant (3+1) approach we seek the initial value solution to these equations. Using the explicit expression for Lorentz force we find that the three equations Eq.(22) automatically imply the constraint Eq.(24), once this is satisfied initially at $t=0$. In this (3+1) formalism, the four-dimensional equations of motion are decomposed into three spatial and one temporal component, effectively eliminating any mixing between space and time in the dynamical equation Eq.(21). This approach to relativistic particle dynamics relies on the use of three independent equations of motion Eq.(22) for three independent coordinates and velocities, ”independent” meaning that equation Eq.(23) (and constraint Eq.(24)) are automatically satisfied.
One might expect that the particle’s trajectory in the lab frame, denoted as $\vec{x}(t)$ and derived from the previous reasoning, should correspond to the covariant trajectory $\vec{x}_{cov}(t)$. However, identifying them in this way leads to paradoxical results. Specifically, the trajectory $\vec{x}(t)$ fails to account for relativistic kinematic effects. In the non-covariant (3+1) approach, the dynamics in the lab frame are formulated without reference to Lorentz transformations. As a result, the motion of a particle in a constant magnetic field, when described from the lab frame, appears identical to its Newtonian counterpart—relativistic effects are absent in this formulation. In conventional particle tracking, the lab-frame trajectory $\vec{x}(t)$ can be interpreted as arising from a sequence of Galilean boosts, which follow the motion of the particle as it accelerates under the influence of the magnetic field. The standard Galilean velocity addition rule is applied to determine the boosts at each instant, tracing the particle’s progression along its curved path.
The old kinematics is especially surprising, because we are based on the use of the covariant approach. Where does it come from? The previous commonly accepted derivation of the equations for the particle motion in the three-dimensional space from the covariant equation Eq.(21) includes one delicate point. In Eq.(22) and Eq.(23) the restriction $\vec{p}=m\vec{v}/\sqrt{1-v^{2}/c^{2}}$ has already been imposed. One might well wonder why because in the accepted covariant approach, the solution of the dynamics problem for the momentum in the lab frame makes no reference to the three-dimensional velocity. Equation Eq.(21) tells us that the force is the rate of change of the momentum $\vec{p}$, but does not tell us how momentum varies with speed. The four-velocity cannot be decomposed into $u=(c\gamma,\vec{v}\gamma)$ when we deal with a particle accelerating along a curved trajectory in the Lorentz lab frame.
Actually, the decomposition $u=(c\gamma,\vec{v}\gamma)$ comes from the relation $u_{\mu}=dx_{\mu}/d\tau=\gamma dx_{\mu}/dt=(c\gamma,\vec{v}\gamma)$. In other words, the presentation of the time component as the relation $d\tau=dt/\gamma$ between the $\tau$ and lab time $t$ is based on the hidden assumption that the type of clock synchronization, which provides the time coordinate $t$ in the lab frame, is based on the use of the absolute time convention.
It is important to stress at this point that the situation when only one clock in the comoving frame is involved in dynamics cannot be realized. The Newton law can be written down in the proper frame only when a space-time coordinate system $(x^{\prime},y^{\prime},z^{\prime},\tau)$ has been specified. The type of clock synchronization which provides the time coordinate in the Newton equation has never been discussed in textbooks. It is clear that without an answer to the question about the method of distant clock synchronization used, not only the concept of acceleration but also the dynamics law has no physical meaning. A proper frame approach to relativistic particle dynamics is forcefully based on a definite (Einstein) coordinatization assumption. After this, the dynamics theory states that the equation of motion in the proper frame is $md^{2}\vec{x^{\prime}}/d\tau^{2}=\vec{f}^{\prime}$. It should be stressed that in the case of velocity increment $\Delta\vec{v^{\prime}}=\Delta\vec{x^{\prime}}/\Delta\tau$, we also deal with the distant events. It can be said with some abuse of language that a 3-velocity vector is always a ”spatially extended” object. Let us return to our consideration of the relation $d\tau=dt/\gamma$ between the $\tau$ and lab time $t$. The calculation carried out in the case of a spatially extended object shows that the temporal coincidence of two distant events has absolute character: $\Delta\tau=0$ implies $\Delta t=0$.
It should be noted that usually the notation ”$\tau$” arose from the proper time. Now let us recall the standard concept of the object’s proper time. Let a point-like object move uniformly and rectilinearly relative to the lab frame $K$. The proper frame $K^{\prime}$ can be fixed to the moving object. The object is at rest in this frame, so that events happening with this object are registered by one clock. This clock counts the proper time at the point where the object is located. As we already know, the proper time on moving object flows slower than the time $t$. This phenomenon was called time dilation. The proper time can also be introduced for a point particle moving with acceleration. This standard interpretation of the proper time $\tau$ has nothing to do with dynamics and one may wonder where this contradiction existing between the name and the content of time in dynamic law comes from.
To avoid being overly abstract for too long, we have introduced a concrete example. Consider the case where the velocity increment is not aligned with the direction of uniform motion. For instance, imagine a particle moving ”upward” with a velocity increment of $\Delta y^{\prime}/\Delta\tau$ in the frame $S_{n}$, while the frame $S_{n}$ itself is moving ”horizontally” with velocity $v$ in the lab frame. According to standard textbook treatments, one would apply the relation $d\tau=dt/\gamma$, leading to the familiar result: $\Delta v_{y}=\Delta v^{\prime}_{y}/\gamma$. However, this approach can lead to paradoxical outcomes. Specifically, it implies a trajectory that does not appropriately mix spatial and temporal components. One of the key conclusions drawn from the discussion in Chapter 7 is that special relativity does not allow for an absolute notion of simultaneity. Consequently, there is no well-defined concept of an instantaneous three-dimensional space. The standard expression for the transformation of a velocity increment, $\Delta v_{y}=\Delta v^{\prime}_{y}/\gamma$, rests on a hidden assumption—that the $y^{\prime}$-axis and $y$-axis remain parallel. In reality, the Lorentz transformation induces a Wigner rotation, which causes these axes to become oblique. This insight immediately reveals a fundamental distinction between non-covariant and covariant descriptions of point particle trajectories. (We will explore velocity transformations in more detail in Section 16.5.)
16.3 Incorrect Expansion of the Relation $d\tau=dt/\gamma$ for Arbitrary Motion
Textbook authors are fundamentally mistaken in their widespread belief about the general validity of the standard momentum–velocity relation. According to the theory of relativity, the equation $\vec{p}_{cov}=m\vec{v}_{cov}/\sqrt{1-v^{2}_{cov}/c^{2}}$ does not hold for particles moving along curved trajectories in the Lorentz lab frame. Many experts who have studied relativity through conventional textbooks may find this claim counterintuitive or even unsettling at first.
But how can such an unconventional momentum–velocity relation arise? We know that the components of the four-momentum $p_{\mu}=(E/c,\vec{p})$ transform between Lorentz frames in the same way as the components of the four-position vector $x=(x_{0},\vec{x})$. However, when translating from four-vector notation back to the familiar three-dimensional velocity vector $\vec{v}$, defined via $\vec{v}=d\vec{x}/dx_{0}$, subtleties emerge. Surprises are to be expected in this transition, especially when dealing with non-rectilinear motion.
It is well established that for rectilinear accelerated motion, the standard momentum–velocity relation does hold. In such cases, the combination of the usual relation and the covariant velocity transformation law (Einstein’s velocity addition) is consistent with the covariant momentum transformation. Both the non-covariant and covariant approaches yield the same trajectory in the Lorentz lab frame.
We can see why by examining the transformation of the 3-velocity in the theory of relativity. For a rectilinear motion, this transformation is performed as $v=(v^{\prime}+V)/(1+v^{\prime}V/c^{2})$. The relativistic factor $1/\sqrt{1-v^{2}/c^{2}}$ is given by: $1/\sqrt{1-v^{2}/c^{2}}=(1+v^{\prime}V/c^{2})/(\sqrt{1-v^{\prime 2}/c^{2}}\sqrt{1-V^{2}/c^{2}})$. The new momentum is then simply $mv$ times the above expression. But we want to express the new momentum in terms of the primed momentum and energy, and we note that $p=(p^{\prime}+\mathcal{E}^{\prime}V/c^{2})/\sqrt{1-V^{2}/c^{2}}$. Thus, for a rectilinear motion, the combination of Einstein’s addition law for parallel velocities and the usual momentum-velocity relation is consistent with the covariant momentum transformation.
It is well known that collinear Lorentz boosts, like their Galilean counterparts, form a commutative group. This implies that the result of applying successive Lorentz boosts in the same direction is independent of the order in which they are applied. However, Lorentz boosts in different directions do not commute and do not form a group. In contrast, Galilean boosts always form a group, regardless of direction.
A key relativistic effect arises when composing non-collinear Lorentz boosts: the result is not simply another boost but a combination of a boost and a spatial rotation, known as the Wigner rotation. This phenomenon has no analogue in non-covariant physics. One of its consequences is the emergence of a non-trivial relationship between momentum and velocity: $\vec{p}_{cov}\neq m\vec{v}_{cov}/\sqrt{1-v^{2}_{cov}/c^{2}}$, which also lacks a non-covariant counterpart. Relativity further reveals that this unusual momentum-velocity relationship is tied to motion along curved trajectories. In such cases, the concept of a single, shared ”ordinary space” breaks down, leading to a fundamental distinction between covariant and non-covariant particle trajectories.
16.4 Standard Integration of the 4D Covariant Equation of Motion
Attempts to solve the dynamical equation Eq. (21) in a manifestly covariant form can be found in the literature. However, the resulting trajectory does not account for relativistic kinematic effects. Consequently, it cannot be identified with $\vec{x}_{cov}(t)$, even though it may initially appear to follow a covariant prescription.
We begin by analyzing the textbook’s treatment of the integration of the covariant equation of motion. Consider, for example, the motion of a charged particle in a given electromagnetic field. Of particular practical importance is the simplest case: a uniform electromagnetic field, meaning that the field tensor $F^{\mu\nu}$ remains constant throughout the entire space-time region of interest. Specifically, we focus on the motion of a particle in a constant, homogeneous magnetic field, represented by the tensor components $F^{\mu\nu}=B(e^{\mu}_{2}e^{\nu}_{3}-e^{\nu}_{2}e^{\mu}_{3})$ where $e^{\mu}_{2}$ and $e^{\mu}_{3}$ are orthonormal space like basis vectors $e^{2}_{2}=e^{2}_{3}=-1$, $e_{2}\cdot e_{3}=0$. In the lab frame of reference where $e^{\mu}_{0}$ is taken as the time axis, and $e^{\mu}_{2}$ and $e^{\mu}_{3}$ are space vectors the field is indeed purely magnetic, of magnitude $B$ and parallel to the $e_{1}$ axis. Let us set the initial four-velocity $u^{\mu}(0)=\gamma ce^{\mu}_{0}+\gamma ve^{\mu}_{2}$, where $v$ is the initial particle’s velocity relative to the lab observer along the axis $e_{2}$ at the instant $\tau=0$, and $\gamma=1/\sqrt{1-v^{2}/c^{2}}$. The components of the equation of motion are then $du^{(0)}/d\tau=0$, $du^{(1)}/d\tau=0$, $du^{(2)}/d\tau=-eBu^{(3)}/(mc)$, $du^{(3)}/d\tau=eBu^{(2)}/(mc)$. We seek the initial value solution to these equations as done in the existing literature  [^53] Textbook authors, such as those in [5, 64, 65], conclude from the covariant equations that a charged particle in a constant magnetic field moves in a uniform circular trajectory. This analysis does not account for relativistic kinematics, and instead, the Galilean vector law of velocity addition is applied. The non-relativistic kinematics are derived from the relation $d\tau=dt/\gamma$. It is only after making this substitution for $d\tau$ hat the authors obtain the familiar formula for the non-covariant trajectory.. A distinctive aspect of relativistic dynamics is the presence of constraints. The evolution of the particle is subject to $mdu^{\mu}/d\tau=eF^{\mu\nu}u_{\nu}$, but also to the constraint $u^{2}=c^{2}$. However, such a condition can be weakened requiring its validity at certain values of $\tau$ only, let us say initially, at $\tau=0$. Therefore, if $Y(\tau)=u^{2}-c^{2}$ vanishes initially, i.e. $Y(0)=0$, then $Y(\tau)=0$ at any $\tau$. In other words, the differential Lorentz-force equation implies the constraint $u^{2}=c^{2}$ once this is satisfied initially. Integrating with respect to the proper time we have $u^{\mu}(\tau)=\gamma e^{\mu}_{0}+\gamma v[e^{\mu}_{2}\cos(\omega\tau)+e^{\mu}_{3}\sin(\omega\tau)]$ where $\omega=eB/(mc)$. We see that $\gamma$ is constant with time, meaning that the energy of a charged particle moving in a constant magnetic field is constant. After two successive integrations we have $X^{\mu}(\tau)=X^{\mu}(0)+\gamma c\tau e^{\mu}_{0}+R[e^{\mu}_{2}\sin(\omega\tau)-e^{\mu}_{3}\cos(\omega\tau)]$ where $R=\gamma v/\omega$. This enables us to find the time dependence $[0,X^{(2)}(t),X^{(3)}(t)]$ of the particle’s position since $t/\gamma=\tau$. From this solution, we conclude that the motion of a charged particle in a constant magnetic field is indeed uniform circular motion, consistent with classical expectations, yet derived within the fully covariant relativistic framework.
One might expect that the particle’s trajectory in the lab frame, derived from the previous reasoning $[0,X^{(2)}(t),X^{(3)}(t)]$, should be identified with $\vec{x}_{cov}(t)$. However, the trajectory $[0,X^{(2)}(t),X^{(3)}(t)]$ does not include relativistic kinematics effects. We have found that the standard integration of the four-dimensional covariant equation of motion, as shown in Eq. (21), yields a particle trajectory that is mathematically identical to the one derived from Newtonian kinematics and dynamics. Specifically, the trajectory of the electron fails to incorporate relativistic effects, and instead, the Galilean law of velocity addition is applied.
It seems that we must have made a mistake. Notice that we are using textbooks to solve covariant equations of motion. While we did not make a computational error in our integrations, we did make a conceptual one. First, we should clarify that there is no issue with the initial integration of Eq. (21) with respect to proper time $\tau$ from the initial conditions. This yields the four-momentum, which has a well-defined, objective meaning—it is convention-invariant. The key point to recognize is that the concept of velocity is only introduced in the second step of the integration. However, in the accepted covariant approach, the solution to the dynamics problem for momentum in the lab frame makes no reference to three-dimensional velocity. The initial condition we used is $u^{\mu}(0)=\gamma ce^{\mu}_{0}+\gamma ve^{\mu}_{2}$, which includes $\gamma c$ and $\gamma v$, not as velocities but as the time and space components of the initial four-momentum. The three-dimensional trajectory and velocity, which are convention-dependent, emerge only after the second integration step. So, where does the old kinematics come from? The second integration is performed using the relation $d\tau=dt/\gamma$. It is only after substituting for $d\tau$ that we recover the conventional (non-covariant) expression for the trajectory of an electron in a constant magnetic field.
We should expect results similar to those obtained in the case of the (3+1) non-covariant particle tracking. In fact, based on the structure of the four components of the equation of motion (Eq. 21), we can derive an equivalent mathematical formulation of the dynamical problem. As previously mentioned, the evolution of the particle in the lab frame is subject to a constraint. This implies that the dynamics are described by only three independent equations of motion. It is easy to see from the initial set of four equations, $du^{(0)}/d\tau=0$, $du^{(1)}/d\tau=0$, $du^{(2)}/d\tau=-eBu^{(3)}/(mc)$, $du^{(3)}/d\tau=eBu^{(2)}/(mc)$, that the presentation of the time component simply as the relation $d\tau=dt/\gamma$ between proper time and coordinate time is just a simple parametrization that yields the corrected Newton’s equation Eq.(22) as another equivalent form of these four equations in terms of absolute time $t$ instead of proper time of the particle. This approach to solving the dynamical equations from the initial conditions relies on three independent spatial coordinates and velocities, without constraints, and is closely tied to classical kinematics. The presentation of the time component simply as the relation $d\tau=dt/\gamma$ between proper time and coordinate time is based on the hidden assumption that the type of clock synchronization, which provides the time coordinate $t$ in the lab frame, is based on the use of the absolute time convention.
16.5 Convention-Invariant Particle Tracking
So far, we have described the motion of a particle in three-dimensional space using the vector-valued function $\vec{x}(t)$. This function defines a prescribed curve (or path) along which the particle moves. The motion along this path is parameterized by $l(t)$, where $l$ represents a particular parameter—specifically, the arc length in the case of interest.
It is important to distinguish between the concepts of path and trajectory. A trajectory provides a complete description of a particle’s motion by specifying its position as a function of time. In contrast, a path is a purely geometric construct, independent of time. Paths can take various forms, such as straight lines, circular arcs, or helical curves.
If we connect the origin of a Cartesian coordinate system to a point on the path, the resulting vector is called the position vector, denoted by $\vec{x}(l)$. The derivative of this vector with respect to arc length, $d\vec{x}(l)/dl$, yields the tangent vector to the curve at that point. The direction of this tangent vector is determined by the orientation of the curve as traced by increasing arc length.
As established in Chapter 2, the path $\vec{x}(l)$ has an exact, objective meaning—it is invariant under changes in convention. Similarly, the components of the four-momentum vector $mu=(\mathcal{E}/c,\vec{p})$ are also convention-invariant. In contrast, the trajectory $\vec{x}(t)$—as measured in a laboratory frame—is convention-dependent, reflecting the conventionality inherent in the definition of velocity. As such, it lacks the same level of objective meaning as the path $\vec{x}(l)$.
We now describe how to determine the position vector $\vec{x}(l)_{cov}$ in the convention-invariant particle tracking approach. Consider a particle moving in a uniform magnetic field, with no electric field present. Using Eq. (21), we obtain:
$\frac{d\vec{p}}{d\tau}=\frac{e}{mc}~{}\vec{p}\times\vec{B},~{}~{}\frac{d\mathcal{E}}{d\tau}=0~{}~{}.$ (25)
Since $d\mathcal{E}/d\tau=0$ and the constraint $\mathcal{E}^{2}/c^{2}-|\vec{p}|^{2}=m^{2}c^{2}$ holds, it follows that $dp/d\tau=0$, where $p=|\vec{p}|=m|d\vec{x}_{cov}|/d\tau$. The unit vector $\vec{p}/p$ satisfies: $\vec{p}/p=d\vec{x}_{cov}/|d\vec{x}_{cov}|=d\vec{x}_{cov}/dl$, where $|d\vec{x}_{cov}|=dl$ is the differential path length.
From this we derive:
$\frac{d^{2}\vec{x}_{cov}}{dl^{2}}=\frac{d\vec{x}_{cov}}{dl}\times\left(\frac{e\vec{B}}{pc}\right)~{}.$ (26)
These equations are identical to those obtained using the non-covariant particle tracking approach. Consequently, $\vec{x}(l)_{cov}=\vec{x}(l)$, as expected. Both formulations describe the same physical reality. Since the curvature radius of the trajectory in a magnetic field—and therefore the three-momentum—has an objective (i.e., convention-invariant) meaning, both approaches yield the same physical results.
16.6 Lorentz-Covariant Particle Tracking
In conventional particle tracking, the trajectory of a particle, denoted by $\vec{x}(t)$, is often described from the laboratory frame as the result of successive Galilean boosts. These boosts account for the motion of a particle accelerating in a constant magnetic field. The classical Galilean rule for the addition of velocities is used to determine the appropriate boost at each instant, effectively tracking the particle’s motion along its curved trajectory.
However, we cannot rely on Newtonian kinematics for mechanics while simultaneously applying Einsteinian kinematics to electrodynamics. A consistent and accurate description requires solving the dynamical equations in their covariant form. Only this approach ensures the correct coupling between Maxwell’s equations and the particle trajectories observed in the laboratory frame. In this covariant framework, the trajectory $\vec{x}_{cov}(t)$ is understood, from the lab frame, as the result of successive Lorentz transformations. This introduces relativistic kinematic effects that are absent in the classical treatment.
Let us apply our algorithm for reconstructing $\vec{x}_{cov}(t)$ to a few examples to observe its performance in practice. To begin, we revisit the standard textbook derivation of velocity transformations between reference frames. Assume a frame $K^{\prime}$ moves relative to the $K$ system with velocity $V$ along the $x$ axis. Let $\vec{v}=d\vec{x}/dt,$ be the vector of the particle velocity in the $K$ system and $\vec{v^{\prime}}=d\vec{x^{\prime}}/dt^{\prime}$ the velocity vector of the same particle in the $K^{\prime}$ system. From Lorentz transformation we have $dx=\gamma(dx^{\prime}+Vdt^{\prime})$, $dy=dy^{\prime}$, $dz=dz^{\prime}$, $dt=\gamma(dt^{\prime}+Vdx^{\prime}/c^{2})$, where $\gamma=1/\sqrt{1-V^{2}/c^{2}}$. Dividing the spatial components by the time component yields the relativistic velocity transformation formulas: $v_{x}=(v_{x}^{\prime}+V)/(1+Vv^{\prime}_{x}/c^{2})$, $v_{y}=v^{\prime}_{y}/[\gamma(1+Vv^{\prime}_{x}/c^{2})]$, $v_{z}=v^{\prime}_{z}/[\gamma(1+Vv^{\prime}_{x}/c^{2})]$. These expressions describe the relativistic law of composition of velocities and form the basis of velocity transformations in special relativity, as presented in standard texts.
We begin with the simple case of relativistically combining perpendicular velocities. While there is no length contraction in directions perpendicular to motion, time dilation still affects the observed velocities. As a result, in the frame $K$, the particle’s velocity components are $v_{x}=V$, $v_{y}=v_{y}^{\prime}/\gamma$. It is important to clarify what is meant by ”perpendicular” in this context. Two velocity vectors can be meaningfully said to be perpendicular only if they are defined within the same reference frame. Therefore, when textbook authors assert that $\vec{e}_{y^{\prime}}v_{y}^{\prime}$ and $\vec{e}_{x}V$ are perpendicular, what they actually rely on—often implicitly—is the assumption that the spatial axes $(x^{\prime},y^{\prime},z^{\prime})$ and $(x,y,z)$ are parallel. In other words, they assume that the observers in frames $K$ and $K^{\prime}$ share a common 3-space. This is a subtle but important misconception.
One of the consequences of the non-commutativity of non-collinear Lorentz boosts is the absence of a shared, common notion of ordinary space. Suppose that in the Lorentz frame $K^{\prime}$, a traveler observes a particle moving along the $y^{\prime}$ axis. In other words, the particle has a velocity component $v^{\prime}_{y}$ in this frame, while the frame $K^{\prime}$ itself is moving with velocity $V$ along the $x$ axis relative to another Lorentz frame $K$. To an observer at rest in frame $K$, the moving frame $K^{\prime}$ appears rotated with respect to $K$. Let us consider the simple case where the boost velocity $V$ approaches the speed of light (i.e., the ultrarelativistic limit), while the transverse velocity $v^{\prime}_{y}$ remains small. This means we are taking the limit where $\gamma\gg 1\gg v^{\prime}_{y}/c$, treating $v^{\prime}_{y}/c$ as a small parameter and keeping terms up to second order. In this regime, the axes of the $K^{\prime}$ frame appear rotated with respect to those of $K$ by an angle $\theta_{w}=v^{\prime}_{y}/(\gamma c)$. Notably, this rotation is in the same direction as the rotation of the particle’s velocity vector in $K$  [^54] To determine the rotation angle in the given problem, we consider a composition of Lorentz boosts. Two successive boosts—first from frame $K$ to $K^{\prime}$, and then an infinitesimal boost to the particle’s proper frame—are equivalent to a single boost from $K$ directly to the proper frame, followed by a rotation. The proper frame is rotated by an angle $d\theta_{w}=dv^{\prime}_{y}/(\gamma c)$ relative to frame $K$. The coordinate axes of the proper frame remain parallel to those of frame $K^{\prime}$. From the perspective of frame $K$, the coordinate axes of frame $K^{\prime}$ are also rotated by the angle $d\theta_{w}$ with respect to those of frame $K$. We now consider the physical consequences of this rotation. First, note that the transverse velocity component $v^{\prime}_{y}$ acquires a projection onto the $x$-axis when viewed from frame $K$. As a result, the component of the particle’s velocity along the $x$-axis in frame $K$ is slightly reduced from $V$.
Our goal is to calculate the decrement in the horizontal velocity in frame $K$. This relativistic correction to the horizontal velocity appears only at second order in $v^{\prime}_{y}/(\gamma c)$. The particle’s trajectory, as viewed from $K$, can be considered the result of a sequence of infinitesimal Lorentz transformations. By integrating over infinitesimal transverse velocity increments $dv^{\prime}_{y}$, we find that the corrected horizontal velocity is $v_{x}=V-(v^{\prime}_{y})^{2}/(2\gamma^{2}c)=V-v_{y}^{2}/(2c)$, where $v_{y}$ is the transverse component of the particle’s velocity as seen in frame $K$. So that the Lorentz transformation effectively rotates the particle’s velocity vector $\vec{V}$ by an angle $v_{y}/V=v_{y}/c$ in ultrarelativistic asymptotic.
This result stands in contrast with the textbook prediction, which states that the total speed of the particle in frame $K$ increases from $V$ to $V+v_{y}^{2}/(2c)$. Our result contradicts this claim. The standard derivation of velocity addition relies on the relation $dt^{\prime}=dt/\gamma$, but neglects the mixing of transverse spatial coordinates with time in Lorentz transformations. This omission leads to an incorrect conclusion about the magnitude of the resulting velocity.
16.7 Relativistic Velocity Composition and the Wigner Rotation
Solving problems related to covariant particle tracking presents two major challenges.
The first challenge arises from the conventional approach to particle tracking within a single inertial frame, which is based entirely on classical Newtonian mechanics, without incorporating Lorentz transformations. Standard textbooks assert that when only one frame is involved, relativity is unnecessary. In this framework, the only relativistic modification is the introduction of relativistic mass, while space and time remain distinct, with no mixing of spatial and temporal components. Surprisingly, this crucial aspect has received little attention within the physics community.
The second challenge stems from confusion surrounding the composition of non-parallel velocities in special relativity. In Newtonian mechanics, velocity addition follows the parallelogram rule, treating velocities as ordinary vectors. However, in relativistic kinematics, velocity composition must be derived from the combination of Lorentz boosts. A well-known issue is that Lorentz boosts alone do not form a group, which raises a fundamental concern: if a set of transformations lacks group structure, the resulting theory is mathematically incomplete.
Textbook authors often assume that the principle of relativity demands that Einstein’s velocity addition law obeys group composition. However, the non-associativity of non-collinear velocity addition is not widely recognized. Many physicists, having learned relativity from standard textbooks, might find this claim surprising, as Einstein’s velocity addition is often assumed to be associative. Yet a straightforward algebraic exercise demonstrates its non-associativity. As a result, constructing a covariant trajectory, $\vec{x}_{cov}(t)$, using only Lorentz boosts is impossible—additional elements are required.
The root cause of this issue lies in the presence of Wigner rotation. In fact, the breakdown of commutativity and associativity in Einstein’s velocity addition law naturally manifests as Wigner rotation, offering deeper insight into the structure of relativistic transformations. The next section will explore the relativistic transformation of velocities in greater detail.
16.8 Mathematical Analysis of Relativistic Velocity Composition
Above, we derived specific quantitative results for the relativistic combination of velocities using only elementary concepts from special relativity. However, some readers may find a more mathematically rigorous explanation clearer.
As a consequence of the space-time geometry, Lorentz coordinate systems are connected by Lorentz transformations. Consider a four-vector, $x=(x_{0},\vec{x})$, with squared-length $x^{2}=\eta_{ij}x^{i}x^{j}=(x_{0})^{2}-|\vec{x}|^{2}$ where $\eta_{ij}=\mathrm{diag}[1,-1,-1,-1]$ is the Minkowski space-time metric. Since Lorentz transformations preserve the space-time interval, they satisfy $x^{\prime}=\Lambda x$ such that $\eta_{ij}=\Lambda^{m}_{i}\Lambda^{n}_{j}\eta_{mn}$. This equation defines the most general 4 × 4 Lorentz transformation matrix $\Lambda$. The set of all such matrices forms a symmetry-group of the Minkowski metric under matrix multiplication, denoted by $O(1,3)$. In particular, $O(1,3)$ is a Lie group, appropriately called the Lorentz group.
It is widely accepted that relationship between inertial frames are completely determined by their relative velocities and relative orientations. Consequently, the generic Lorentz transformation denoted as $\Lambda(\vec{v},\vec{\theta})$, is parameterized by the velocity $\vec{v}$ and the rotation parameter $\vec{\theta}$, representing a simultaneous boost and rotation. Here, $\vec{v}$ is a vector in three-dimensional Euclidean space, wile $\vec{\theta}$ belongs to the group $O(3)$, which consists of all possible spatial rotations in three-dimensions. This parameterization of the Lorentz group is commonly found in textbooks.
Now, we turn to a significant challenge associated with this parametrization of the Lorentz group. A comparison with the Galilean group may provide insight. It is well known that the Galilean transformation $G(\vec{v},\vec{\theta})$, which preserves the group composition law, is isomorphic to the semi-direct product group $R^{3}\times O(3)$, where $R^{3}$ represents the group of Galilean boosts and $O(3)$ denotes the rotation group.
In contrast, the composition of Lorentz transformations is more intricate. The key distinction between the Galilean and Lorentz groups lies in their structure: the Lorentz group is simple, meaning it does not naturally decompose into a direct product of rotations and boosts. Unlike in the Galilean case, $O(3)$ is not a normal subgroup of the Lorentz group $O(1,3)$, and the set of Lorentz boosts does not even form a group. The theory of the Galilean transformation composition law is straightforward and well understood. In contrast, the analogous theory of the Lorentz transformation composition law is unsatisfactory.
We conclude that the standard analysis of the relationship between inertial frames overlooks the fundamental difference between Newtonian space and time and the pseudo-Euclidean geometry of space-time. In classical mechanics, the components of the velocity three-vector transform under a Galilean transformation in the same way as the components of the position three-vector $\vec{x}$. This is because, according to classical kinematics, all inertial frames share the same ordinary three-dimensional space. In contrast, in special relativity, different inertial frames have distinct three-dimensional spaces within the framework of space-time. Therefore, unexpected results may arise in special relativity when transitioning from the four-vector formalism back to the three-dimensional velocity vector $\vec{v}$, which is expressed in terms of the four-vector components $(x_{0},\vec{x})$ as $\vec{v}=d\vec{x}/dx_{0}$.
Most texts on special relativity present the relativistic velocity addition formula only for parallel velocities. In this simplified case, Einstein’s velocity addition is both commutative and associative, making it a commutative group operation. However, in the general case, Einstein’s velocity addition $\oplus$ does not form a group.
The relativistic reciprocity principle states that if an inertial frame $K^{\prime}$ moves with velocity $\vec{v}$ relative to another inertial frame $K$, then, reciprocally, the velocity of $K$ relative to $K^{\prime}$ is $-\vec{v}$. This principle suggests that the inverse of a velocity $\vec{v}$ is simply its negation: $\vec{v}^{-1}=-\vec{v}$, analogous to the Galilean addition of velocities. For composite velocities, $\vec{u}\oplus\vec{v}$, reciprocity implies $(\vec{u}\oplus\vec{v})^{-1}=-(\vec{u}\oplus\vec{v})=(-\vec{v})\oplus(-\vec{u})=-(\vec{v}\oplus\vec{u})$. However, since Einstein’s velocity addition $\oplus$ is generally non-commutative, this leads to a contradiction, known as the reciprocity paradox. This paradox is closely related to the Mocanu paradox, as discussed by Mocanu [66] and was latter resolved by Ungar [67] who demonstrated the crucial role of the Wigner (Thomas) rotation in the transformation. Furthermore, Ungar [67] also showed that Einstein’s velocity addition is also non-associative meaning that $(\vec{w}\oplus(\vec{v}\oplus\vec{u}))\neq((\vec{w}\oplus\vec{v})\oplus\vec{u})$. Being non-associative, the Einstein’s addition $\oplus$ is not a group operation. Non-associativity is paradoxical: for a system of four bodies the $\oplus$ of three non-parallel velocities gives two distinct velocities between two bodies.
Most textbooks present the Lorentz transformation as being generated by a boost vector $\vec{v}$, interpreted as the unique reciprocal velocity between two inertial frames. However, we emphasize that the reciprocal property $\vec{v}^{-1}=-\vec{v}$ is inconsistent with Einstein’s two postulates of special relativity  [^55] For a more in-depth discussion on the relationship between the reciprocity axiom and Einstein’s two postulates of special relativity, we recommend the chapter ”How Do You Add Relative Velocities?” by Oziewicz in [69].. The Lorentz isometry group $\Lambda$ acts on four-dimensional spacetime, and all subgroups of the Lorentz group operate on four-vectors in spacetime. Consequently, there is no concept of common (instantaneous) spatial frame within the framework of the Lorentz group. The Lorentz transformation $\Lambda$ does not naturally decompose into a unique combination of a rotation and a boost. Any such decomposition would imply the existence of a preferred frame, contradicting the principle of relativity.
In Minkowski space, where spacetime follows pseudo-Euclidean geometry, equivalence of all inertial frames is not possible. This suggests the existence of an explicit, preferred (initial) inertial reference frame within special relativity, challenging the conventional interpretation. The relationship between inertial frames are not determined solely by their relative velocities 565656 In particular, the presence of non-zero relative velocity cannot, by itself, be a cause of time dilation (see Section 12.3 for more detail)..
The inverse of a velocity vector, $\vec{v}^{-1}=-\vec{v}$, is often presented in textbooks as a self-evident fact requiring no further discussion. However, this assumption contradicts a fundamental axiom of relativity theory—namely, the axiom of pseudo-Euclidean spacetime geometry.
For example, in his contemporary textbook, Barut defines Lorentz boost as (see Eq. (1.28) and the accompanying discussion in [6]):
$L_{\vec{v}}=(L_{-\vec{v}})^{-1}=\left(\begin{array}[]{cccc}\gamma&-\beta\gamma&0&0\\ -\beta\gamma&\gamma&0&0\\ 0&0&1&0\\ 0&0&0&1\end{array}\right).$
It is evident that all textbooks, including [6], define the Lorentz boost as a 4x4 basis-dependent matrix. Next we shall discuss some physical applications of the transformation $L_{\vec{v}}$.
Consider a coordinate system $K_{1}$ moving with respect to $K_{0}$ with a velocity $\vec{v}_{01}$ and a frame $K_{2}$ moving with respect to $K_{1}$ with a velocity $\vec{v}_{12}$. It is important to note that $\vec{v}_{01}$ represents the velocity of frame $K_{1}$ as measured by an observer at rest in $K_{0}$, while $\vec{v}_{12}$ represents the velocity of $K_{2}$ as measured by an observer in $K_{1}$. The question then arises: what is the Lorentz transformation from $K_{0}$ to $K_{2}$? Let the corresponding Lorentz transformations be denoted by $L_{1}$, $L_{2}$ and $L$, respectively. Then, according to standard textbooks conventions, we have $L=L_{2}L_{1}$.
Now, consider the case, where $\vec{v}_{01}$ and $\vec{v}_{12}$ are perpendicular. The velocity $\vec{v}_{02}$ of frame $K_{2}$ as observed from $K_{0}$ is given by $\vec{v}_{02}=\vec{v}_{01}+\vec{v}_{12}\sqrt{1-v_{01}^{2}/c^{2}}$. According to reciprocity principle, accepted in all textbooks, an observer in $K_{2}$ sees frame $K_{1}$ moving with velocity $-\vec{v}_{12}$, while an observer in $K_{1}$ sees frame $K_{0}$ moving with velocity $-\vec{v}_{01}$. Similarly, applying Einstein’s velocity addition, we find that the velocity $\vec{v}_{20}$ of frame $K_{0}$ as observed from $K_{2}$ is: $\vec{v}_{20}=-\vec{v}_{12}-\vec{v}_{01}\sqrt{1-v_{12}^{2}/c^{2}}$. According to textbooks, $\vec{v}_{20}$ should be reciprocal of $\vec{v}_{02}$, but they point in different directions. The angle between $\vec{v}_{02}$ and $\vec{v}_{20}$ is the Wigner rotation angle. This reciprocity discrepancy leads to the Mocanu contradiction and highlights the non-associativity of Einstein’s velocity addition.
While composite velocities in Galilean relativity fully determine the relationships between inertial frames, their relativistic counterparts do not. In addition to the relativistic velocity composition between inertial frames, one must also account for the associated Wigner (Thomas) rotations.
In his papers and monographs—written at an advanced mathematical level — Ungar attempted to resolve the non-associativity paradox of Einstein’s velocity addition $\oplus$ by refining the composition law (see [68] and references therein). The improved composition law replaces $\vec{u}\oplus\vec{v}$ with $[\vec{u}\oplus\vec{v},\mathrm{W}(\vec{u},\vec{v})]$, where $\mathrm{W}(\vec{u},\vec{v})$ represents the Wigner (Thomas) rotation generated by $\vec{u}$ and $\vec{v}$. Notably, the binary operation $\oplus$ remains the standard Einstein velocity addition.
However, this approach is unsatisfactory even from a mathematical standpoint for a fundamental reason. As demonstrated in Section 16.6, the Wigner rotation is not merely an additive correction but is intrinsically woven into the structure of relativistic velocity addition itself. Consider the asymptotic case of two perpendicular velocities: $\vec{v}_{01}$ approaching the speed of light and $\vec{v}_{12}$ being small. In this limit, $\gamma=(1-v_{01}^{2}/c^{2})^{-1/2}\gg 1\gg v_{12}/c$. The frame $K_{1}$ with the moving frame $K_{2}$ seems to rotate about frame $K_{0}$. We want now to look at the consequence of this Wigner rotation. In the Section 16.6 we found that in Lorentz coordinatization, the velocity $v_{01}$ of frame $K_{2}$ in frame $K_{0}$, after a boost with velocity $\vec{v}_{12}$, undergoes a simple rotation by an angle $v_{12}/(\gamma v_{01})=v_{12}/(\gamma c)$ without any change in its magnitude. In contrast, Ungar’s approach—based on the standard Einstein velocity addition within an improved composition law—suggests that the total speed of $K_{2}$ in the Lorentz frame $K_{0}$ increases from $v_{01}$ to $v_{01}+v_{12}^{2}/(2\gamma^{2}c)$.
In all special relativity textbooks, the principle of reciprocity states that the inverse of velocity addition must be identical to that in Galilean absolute time. Consequently, Einstein’s velocity addition is based on the relation $dt^{\prime}=dt/\gamma$, which implicitly assumes an absolute time convention. However, it does not account for the interplay between transverse positions and time, which is inherent in Lorentz boost matrices. In reality, every Lorentz boost represents a transformation between reference frames within a Lorentzian framework (i.e., under Lorentz coordinatization). Thus, we conclude that the commonly accepted derivation of velocity composition does not follow from the composition of Lorentz boosts, $L=L_{2}L_{1}$.
Under Lorentz coordinatization, moving frames inherently possess different simultaneity relations. As a result, the velocity $\vec{v}$ and its inverse $\vec{v}^{-1}$ are each tangent to a different instantaneous space and these two spaces are not parallel. Consequently, the inverse velocity cannot be truly reciprocal. The reciprocity of frames contradicts relativity of simultaneity. In Enstein’s velocity addition all velocities $\vec{v}_{01}$, $\vec{v}_{12}$, $\vec{v}_{02}$, are implicitly assumed to lie within the same preferred 3-space as seen by a preferred observer. In other words, the reciprocity of relative velocities holds only under an absolute time framework.
The standard textbook presentation of the relativistic transformations implicitly assumes that observers in $K_{0}$, $K_{1}$, and $K_{2}$ inertial frames share a common 3-space. It is often assumed that the acceleration of a rigid reference frame grid has a direct physical meaning. However, this is a misconception. Different inertial frames have distinct ordinary spaces and implicit assumption of velocity reciprocity contradicts Einstein’s two postulates.
The concepts of observer independence and Lorentz covariance are fundamentally distinct. We conclude that there is a striking conflict between the Principle of Relativity (i.e., the equivalence of all inertial frames) and the axiom of pseudo-Euclidean spacetime geometry. In other words, only one of these principles can hold. Based on experimental evidence, we adopt the Lorentz covariance principle  [^57] However, an alternative perspective exists in the literature. Oziewicz, for instance, favors the Principle of Relativity [69]..
This is a good point to make some general remarks about Einstein’s velocity addition. It is important to recognize that it does not directly follow from the Lorentz transformations. While Lorentz boosts intertwine space and time coordinates, Einstein’s velocity addition relies on the assumption of a shared three-dimensional space across all inertial frames.
Despite this, Einstein’s velocity addition is widely used in relativistic dynamics and electrodynamics, raising a reasonable question: How can this be? After all, we know that this form of velocity addition does not constitute a mathematical group—it is non-associative, and without associativity, the development of a consistent and complete theory would seem impossible.
Earlier, we demonstrated that the main issue lies in the definition of relative velocity. If we postulate, as Einstein did, the principle of relativistic reciprocity, we inevitably encounter the non-associativity of velocity addition. Nevertheless, one should recognize the existence of a preferred (initial) inertial frame without an accelerational history relative to the fixed stars. Based on experimental evidence, we adopt the Sun-centered frame as the preferred inertial frame (see Chapter 15 for further discussion). Once a preferred frame is established, there is one—and only one—well-defined relative velocity between any pair of inertial frames.
Let us apply this approach to an example and observe how it performs in practice. Suppose an observer in the initial inertial frame $K_{0}$ establishes a coordinate system using Einstein’s synchronization procedure, based on light signals emitted by a stationary dipole source, under the assumption that light propagates isotropically at speed $c$. Now consider two additional moving inertial frames, $K_{1}$ and $K_{2}$, whose motions are analyzed by the observer in $K_{0}$. The simplest synchronization method involves maintaining the same set of synchronized clocks, thus preserving simultaneity and adhering to an absolute-time coordinatization in $K_{0}$.
We assume that system $K_{1}$ moves with velocity $v_{x}$ along the $x$-axis relative to $K_{0}$, and $K_{2}$ moves with velocity $v_{y}$ along the $y$-axis relative to $K_{1}$. It is important to note that $v_{x}$ and $v_{y}$ represent the velocities as measured by the observer in $K_{0}$.
Applying Einstein’s velocity addition, the observer in $K_{0}$ concludes that an observer in $K_{1}$ sees frame $K_{2}$ moving with a velocity $v_{y}/\sqrt{1-v_{x}^{2}/c^{2}}$. According to Einstein, the velocity of frame $K_{2}$ as observed directly from $K_{0}$ is given by $\vec{v}_{02}=v_{x}\vec{e}_{x}+v_{y}\vec{e}_{y}$ consistent with Galilean velocity addition within the initial inertial frame.
Now, consider a different perspective. The observer in $K_{0}$ sees the relative velocity of $K_{2}$ with respect to $K_{1}$ as $-v_{y}$. Similarly, an observer in $K_{2}$ sees $K_{1}$ moving with velocity $-v_{y}/\sqrt{1-v_{02}^{2}/c^{2}}$. According to Einstein’s velocity addition, the velocity of $K_{1}$ as observed from $K_{0}$ is $\vec{v}_{01}=v_{x}\vec{e}_{x}$ as expected.
We thus conclude that Einstein’s velocity addition is consistent with Galilean velocity addition when viewed from the initial inertial frame. In this scheme, reciprocity is absent, and all paradoxes are eliminated. Indeed, there is no reciprocity in the relative velocities between $K_{1}$ and $K_{2}$, because time dilation is absolute, not relative, in the presence of a preferred inertial frame.
Now, let us consider the Earth-based frame. For an observer on Earth, the velocity relative to the initial inertial frame is approximately $v_{orb}\simeq 30~{}\mathrm{km/s}$, allowing us to neglect terms of order $v_{orb}^{2}/c^{2}$. As previously demonstrated, current experimental techniques usually lack the sensitivity required to detect the influence of orbital velocity on the relativistic particle dynamics. Consequently, in accelerator physics, the Earth-based frame can be treated with good accuracy as the initial inertial frame.
It is important to emphasize that Einstein’s velocity addition is fully compatible with conventional particle tracking within a single inertial frame, which itself relies on the absolute time convention. Crucially, Einstein’s velocity addition inherently incorporates relativistic time dilation. Similarly, conventional particle tracking also accounts for relativistic time dilation, albeit implicitly, through the use of relativistic mass for moving particles.
This approach presents no fundamental challenges in mechanics. The preference for a non-covariant formulation within the framework of dynamics arises from its simplicity and practicality. What is essential is the consistent application of the chosen convention across both dynamics and electrodynamics. In an absolute-time coordinatization, the equations of electrodynamics are not isotropic. This poses a challenge, as electrodynamics involves partial differential equations, whereas dynamics deals only with ordinary differential equations. Only by adopting a kinematic framework based on Lorentz coordinatization can we ensure the correct coupling between the isotropic Maxwell equations and particle trajectories in the laboratory frame.
Let us now come to our conclusion: Einstein’s velocity addition requires the existence of a preferred inertial frame. Therefore, Einstein’s velocity addition is consistent with the concept of an initial inertial frame and stands in contradiction to the Principle of Relativity, namely the equivalence of all inertial frames.
17 Relativistic Spin Dynamics
We aim to demonstrate that while studying relativistic particle dynamics, we simultaneously explored a wide range of related topics. As an example, consider the motion of spin in a given field. The equations governing spin dynamics can be formulated as tensor equations in Minkowski space-time. Here, we focus on the case of a particle with a magnetic moment $\vec{\mu}$ in a microscopically homogeneous electromagnetic field. Notably, the torque influences only the spin, while the force affects only the momentum. Consequently, the overall motion of the system in any reference frame is determined solely by its charge, independent of its magnetic dipole moment—a topic addressed in the previous chapter. Our current focus is on the dynamics of spin motion.
17.1 Magnetic Dipole at Rest in an Electromagnetic Field
Let us first consider the spin precession for a nonrelativistic charged particle. The proportionality of magnetic moment $\vec{\mu}$ and angular momentum $\vec{s}$ has been confirmed in many ”gyromagnetic” experiments on many different systems. The constant of proportionality is one of the parameters charactering a particular system. It is normally specified by giving the gyromagnetic ratio or $g$ factor, defined by $\vec{\mu}=ge\vec{s}/(2mc)$. This formula says that the magnetic moment is parallel to the angular momentum and can have any magnitude. For an electron $g$ is very nearly 2.
Suppose that a particle is at rest in an external magnetic field $\vec{B}_{R}$. The equation of motion for the angular momentum in its rest frame is $d\vec{s}/d\tau=\vec{\mu}\times\vec{B}_{R}=eg\vec{s}\times\vec{B}_{R}/(2mc)=\vec{\omega}_{s}\times\vec{s}$. In other words, the spin precesses around the direction of the magnetic field with the frequency $\omega_{s}=-eg\vec{B}_{R}/(2mc)$. In the same nonrelativistic limit the velocity processes around the direction of $\vec{B}_{R}$ with the frequency $\omega_{p}=-(e/mc)\vec{B}_{R}$: $d\vec{v}/d\tau=(e/mc)\vec{v}\times\vec{B}_{R}$. Thus, for $g=2$ spin and velocity precess with the same frequency, so that the angle between them is conserved.
17.2 Derivation of the Covariant (BMT) Equation of Spin Motion
In deriving the equation for spin motion, we rely on the well-established dynamics in the rest frame and known relativistic transformation laws. Since spin is inherently defined in a specific frame (the rest frame), constructing expressions with well-defined transformation properties requires introducing a four-quantity associated with spin. A natural choice is the four-(pseudo)vector $S$, defined such that in the rest frame, its spatial components correspond to the spin components, while its time-like component is zero. We refer to $S$ as the four-spin; when normalized by its invariant length, it becomes the polarization four-vector. Since it is space-like, its spatial part never vanishes in any frame.
Let the spin of the particle be represented in the rest frame by $\vec{s}$. The four-vector $S^{\alpha}$ is by definition required to be purely spatial at time $\tau$ in an instantaneous Lorentz rest frame $R(\tau)$ of the particle and to coincide at this time with the spin $\vec{s}(\tau)$ of the particle; that is $S_{R}^{\alpha}(\tau)=(0,\vec{S}_{R}(\tau))=(0,\vec{s}(\tau))$. At a later instant $\tau+\Delta\tau$ in an instantaneous inertial rest frame $R(\tau+\Delta\tau)$, we have similarly $S_{R}^{\alpha}(\tau+\Delta\tau)=(0,\vec{S}_{R}(\tau+\Delta\tau))=(0,\vec{s}(\tau+\Delta\tau))$.
The BMT equation [46] is the manifestly covariant equation of motion for a four-vector spin $S^{\alpha}$ in an electromagnetic field $F^{\alpha\beta}$:
$\frac{dS^{\alpha}}{d\tau}=\frac{ge}{2mc}\left[F^{\alpha\beta}S_{\beta}+\frac{1}{c^{2}}u^{\alpha}\left(S_{\lambda}F^{\lambda\mu}u_{\mu}\right)\right]-\frac{1}{c^{2}}u^{\alpha}\left(S_{\lambda}\frac{du^{\lambda}}{d\tau}\right)~{},$ (27)
where $u_{\mu}=dx_{\mu}/d\tau$ is the four-dimensional particle velocity vector. With Eq.(21), one has
$\frac{dS^{\alpha}}{d\tau}=\frac{e}{mc}\left[\frac{g}{2}F^{\alpha\beta}S_{\beta}+\frac{g-2}{2c^{2}}u^{\alpha}\left(S_{\lambda}F^{\lambda\mu}u_{\mu}\right)\right]~{},$ (28)
The BMT equation is valid in any Lorentz frame and, together with the covariant force law, consistently describes the motion of a charged particle with spin and magnetic moment. The covariant equation of spin motion for a relativistic particle subjected to the four-force $Q^{\mu}=eF^{\mu\nu}u_{\nu}$ (as expressed in the Lorentz lab frame, Eq. (27)) serves as a relativistic generalization of the classical spin dynamics equation in the particle’s rest frame. This generalization embeds the three Larmor spin precession equations into the four-dimensional structure of Minkowski space-time.
In Lorentz coordinates, there exists a kinematic constraint $S^{\mu}u_{\mu}=0$, which expresses the orthogonality between the four-spin and the four-velocity vectors. Due to this condition, the four-dimensional dynamical equation, Eq. (27), effectively contains only three independent equations of motion. By substituting the explicit form of the Lorentz force, one finds that Eq. (27) inherently satisfies the constraint $S^{\mu}u_{\mu}=0$, as required. To demonstrate this, note that in any Lorentz frame, the time component of the spin satisfies $S^{0}=\vec{S}\cdot\vec{v}$. Although $S^{0}$ vanishes in the particle’s rest frame, its derivative with respect to proper time, $dS^{0}/d\tau$, does not necessarily vanish. In fact, the condition $d(S^{\mu}u_{\mu})/d\tau=0$ leads directly to $dS^{0}/d\tau=\vec{S}\cdot d\vec{v}/d\tau$. The immediate generalization of $d\vec{S}/d\tau=eg\vec{S}\times\vec{B}_{R}/(2mc)$ and $dS^{0}/d\tau=\vec{S}\cdot d\vec{v}/d\tau$ to arbitrary Lorentz frames is Eq.(27) as can be checked by reducing to the rest frame. A methodological parallel can thus be drawn between this formulation and the relativistic generalization of Newton’s second law.
To fully grasp the significance of embedding the spin dynamics law within Minkowski space-time, it is important to recall that the spin dynamics equation, as presented earlier in the Lorentz comoving frame, is characterized as a phenomenological law. Its formulation does not provide a microscopic interpretation of a particle’s magnetic moment. In other words, the spin dynamics law is generally accepted as phenomenological, with the magnetic moment introduced in an ad hoc manner. The coordinate system in which the classical equations of motion for a particle’s angular momentum are valid is referred to as the Lorentz rest frame. The relativistic generalization of the three-dimensional equation $d\vec{s}/d\tau=eg\vec{s}\times\vec{B}_{R}/(2mc)$ to an arbitrary Lorentz frame enables accurate predictions of spin behavior across different reference frames.
17.3 Changing Spin Variables
When Bargmann, Michel, and Telegdi first formulated the correct laws of spin dynamics, they derived a manifestly covariant equation in Minkowski space-time, Eq. (28), describing the motion of the four-spin $S^{\mu}$. Their approach closely resembled the four-tensor equations already established in relativistic particle dynamics. However, Eq. (28) is more intricate than it might initially appear. It consists of a system of coupled differential equations, making direct solutions challenging—even in highly symmetric cases, such as a uniform magnetic field. The question of how to solve this four-tensor equation is particularly intriguing. In relativistic spin dynamics, a specific and highly convenient method is employed to tackle this problem.
To apply Eq.(28) to specific problems it is convenient to introduce a three vector $\vec{s}~{}$ by the equation $\vec{s}=\vec{S}+S^{0}\vec{p}c/(\mathcal{E}+mc^{2})$. With the help of this relation, one can work out the equation of motion for $\vec{s}~{}$. Let us restrict our treatment of spinning particle dynamics to purely transverse magnetic fields. This means that the magnetic field vector $\vec{B}$ is oriented normally to the particle line motion. In this practically important case one has, after somewhat lengthy calculations [46]:
$\frac{d\vec{s}}{d\tau}=\vec{\Omega}\times\vec{s}=-\frac{e}{2mc}\left[\left(g-2+\frac{2}{\gamma}\right)\gamma\vec{B}\right]\times\vec{s}~{},$ (29)
What must be recognized is that in the accepted covariant approach (indeed, Eq.(28) is manifestly covariant), the solution of the dynamics problem for the spin in the lab frame makes no reference to the three-dimensional velocity. In fact, the Eq.(29) includes relativistic factor $\gamma$ and vector $\vec{v}/c$, which are actually notations: $\gamma=\mathcal{E}/(mc^{2})$, $\vec{v}/c=\vec{p}c/\mathcal{E}$. All quantities $\mathcal{E},\vec{p},\vec{B}$ are defined in the lab frame and possess exact, objective meaning—that is, they are independent of any choice of convention. The evolution parameter $\tau$ is likewise measured in the lab frame and has a similarly objective interpretation. For example, it is straightforward to show that $d\tau=mdl/|\vec{p}|$, where $dl$ denotes the differential path length.
We now have the equation in a form that is convenient for solving. Consider a charged, spinning particle moving through a bending magnet of length $dl$ in the lab frame. The orbital deflection angle of the particle is given by $d\theta=-eBdl/(|\vec{p}|c)$. Additionally, the proper time interval is $d\tau=mdl/|\vec{p}|$. From Eq. (29), we can express the spin rotation angle relative to the lab frame axes as $\Omega d\tau$, which becomes: $\Omega d\tau=[(g/2-1)\gamma d\theta+d\theta]$.
The spin vector $\vec{s}$ directly represents the spin as perceived in a comoving system. In the lab frame, if we say that a particle’s spin makes an angle $\phi$ with its velocity, this means that in the particle’s rest frame, the spin makes the same angle with the direction of motion of the lab frame. This leads us to conclude that the conventional approach used to describe spin dynamics in the lab frame is rather unconventional. Specifically, the measurement of spin rotation in the lab frame is interpreted, from the proper observer’s perspective, as viewing this of the lab observer.
17.4 An Alternative Approach to BMT Theory
Above we described the BMT equation, Eq.(29), in the standard manner. It uses a spin quantity defined in the proper frame but observed with respect to the lab frame axes. That means that we know the orientation of the proper spin with respect to the lab coordinate system which is moving with velocity $-\vec{v}$ and acceleration $-\gamma d\vec{v}/d\tau$ in the proper frame. Let’s look at what the equation Eq.(29) says in a little more detail. It will be more convenient if we rewrite this equation as
$d\vec{s}=\vec{\Omega}d\tau\times\vec{s}=-eg\gamma\vec{B}d\tau/(2mc)\times\vec{s}+e(\gamma-1)\vec{B}d\tau/(mc)\times\vec{s}~{}.$ (30)
Now let’s see how we can write the right-hand side of Eq.(30). The first term is that we would expect the spin rotation due to a torque with respect to the proper frame axes $d\vec{\phi}_{L}=-eg\gamma\vec{B}d\tau/(2mc)=(g/2)\gamma d\vec{\theta}$. Here $d\vec{\theta}=-eBdl/(|\vec{p}|c)$ is the angle of the velocity rotation in the lab frame. It has also been made evident by our analysis in Chapter 8 that the angle of rotation $d\vec{\phi}_{W}=-e(\gamma-1)\vec{B}d\tau/(mc)=(\gamma-1)d\vec{\theta}$ corresponds to the Wigner rotation of the lab frame axes with respect to the proper frame axes. With these definitions, we have $d\vec{s}=\vec{\Omega}d\tau\times\vec{s}=d\vec{\phi}_{L}\times\vec{s}-d\vec{\phi}_{W}\times\vec{s}$.
We now present a new approach to the BMT theory, offering an alternative method for addressing this complex problem. It is important to note that $d\vec{\phi}_{L}$ and $d\vec{\phi}_{W}$ represent rotations with respect to the axes of the proper frame. However, our primary goal is to determine the spin motion relative to the laboratory frame axes. In doing so, we must pay close attention to the signs of the rotations, as they play a critical role in the analysis.
A helpful mnemonic can be used to remember the signs of various rotational effects. It consists of three parts:
1. The direction of velocity rotation in the proper frame is the same as the direction of velocity rotation in the lab frame.
2. The direction of lab frame rotation as observed in the proper frame is the same as the direction of velocity rotation in the proper frame.
3. When $g>0$ (which is the case for an electron, where $g$ is positive and approximately equal to 2), the direction of spin rotation due to a torque in the proper frame matches the direction of velocity rotation in the proper frame.
We now turn to the question of determining the proper spin rotation relative to the lab frame axes. This is straightforward: the relevant rotation angle is given by the difference $d\vec{\phi}_{L}-d\vec{\phi}_{W}$. With this, we begin to grasp the fundamental framework of spin dynamics. It becomes clear why the Wigner rotation of the lab frame axes, as seen from the proper frame, must be accounted for when analyzing spin dynamics relative to the lab frame.
Why is the new derivation of the BMT equation so simple? The key reason is that the decomposition of the particle’s spin motion relative to the lab frame axes into dynamic and kinematic components can only be properly realized in the particle’s proper frame. In this frame, there is no need to consider a relativistic ”generalization” of the phenomenological classical equation of motion for angular momentum. This approach allows the spin dynamics problem to be separated into two distinct parts: the trivial (Larmor) dynamic problem and the kinematic problem of the Wigner rotation of the lab frame axes in the proper frame.
17.5 Spin Tracking
By expressing the spin motion equation in four-vector form, Eq. (28), and determining the components of the four-force, we have not only ensured compliance with the principle of relativity but also obtained the four-component formulation of the spin motion equation. This represents a covariant relativistic generalization of the conventional three-dimensional equation for the motion of a magnetic moment, where the particle’s proper time serves as the evolution parameter. Next, we aim to describe the spin motion relative to the Lorentz lab frame, using the lab time $t$ as the evolution parameter.
When going from the proper time $\tau$ to the lab time $t$, the frequency of spin precession with respect to the lab frame can be obtained using the well-known formula $d\tau=dt/\gamma$. We then find
$\frac{d\vec{s}}{dt}=\vec{\varpi}\times\vec{s}=-\frac{e}{2mc}\left[\left(g-2+\frac{2}{\gamma}\right)\vec{B}\right]\times\vec{s}~{}.$ (31)
The frequency of spin precession can be written in the form $\varpi=\omega_{0}[1+\gamma(g/2-1)]$, where $\omega_{0}$ is the particle revolution frequency.
The old kinematics comes from the relation $d\tau=dt/\gamma$. The presentation of the time component simply as the relation $d\tau=dt/\gamma$ between proper time and coordinate time is based on the hidden assumption that the type of clock synchronization that provides the time coordinate $t$ in the lab frame is based on the use of the absolute time convention.
In the previous chapter, we saw that the particle path $\vec{x}(l)$ has an exact objective meaning i.e. it is convention-invariant. The spin orientation $\vec{s}$ at each point of the particle path $\vec{x}(l)$ has also exact objective meaning. In contrast to this, and consistently with the conventionality of the three-velocity, the function $\vec{s}(t)$ describing the spinning particle in the lab frame has no exact objective meaning.
We now describe how to determine the spin orientation along the path $\vec{s}(l)$ in convention-invariant spin tracking. Starting with the covariant equation (Eq. 27), we obtain Eq. (Eq. 29). By using the relation $d\tau=mdl/|\vec{p}|$, we arrive at the convention-invariant equation of spin motion:
$\frac{d\vec{s}}{dl}=-\frac{e\mathcal{E}}{m|\vec{p}|c^{3}}\left[\left(\frac{g}{2}-1+\frac{mc^{2}}{\mathcal{E}}\right)\vec{B}\right]\times\vec{s}=\left[\left(\frac{g}{2}-1\right)\frac{\mathcal{E}}{mc^{2}}+1\right]\frac{d\vec{\theta}}{dl}\times\vec{s}~{},$ (32)
where $l$ is the path length, used as the evolution parameter. These three equations precisely correspond to the equations for components of the proper spin vector derived from the non-covariant spin tracking equation (Eq. 31).
It is important to note that there are two distinct approaches—covariant and non-covariant—that yield the same spin orientation $\vec{s}(l)$ along the path. Both approaches describe the same physical reality correctly, meaning that the orientation of the proper spin $\vec{s}$ at any point along the particle’s path in the magnetic field is objective and convention-invariant.
18 Relativity and Electrodynamics
The differential form of Maxwell’s equations describing electromagnetic phenomena in the Lorentz lab frame is given by Eq.(8). Now let us use these equations to discuss the phenomena called radiation. To evaluate radiation fields arising from external sources in Eq. (8), we need to know the velocity $\vec{v}$ and the position $\vec{x}$ as a function of the lab frame time $t$. As discussed above, it is generally accepted that one should solve the Maxwell’s equations in the lab frame with current and charge density created by particles moving along non-covariant trajectory like $\vec{x}(t)$. The trajectory $\vec{x}(t)$, which follows from the solution of the corrected Newton’s second law Eq. (4) under the absolute time convention, does not include, however, relativistic kinematics effects.
We now turn to the motion of a particle in a given magnetic field. Maxwell’s equations, in their standard form, are valid only within Lorentz reference frames. According to the proper coupling between fields and particles, Eq. 10 implies the following expressions for the charge and current densities:
$\rho(\vec{x},t)=e\delta(\vec{x}-\vec{x}_{cov}(t))~{},$ (33)
$\vec{j}(\vec{x},t)=e\vec{v}_{cov}(t)\delta(\vec{x}-\vec{x}_{cov}(t))~{},$ (34)
where $\vec{v}_{cov}=d\vec{x}_{cov}/dt$. Here, the covariant trajectory of the particle, as observed in the laboratory frame, results from a sequence of Lorentz transformations.
[…]
18.1 The Dipole Approximation
The distinction between covariant and non-covariant particle trajectories was never fully understood. As a result, physicists failed to recognize the contribution of relativistic kinematic effects to radiation. This oversight naturally raises an important question: why did this error in radiation theory remain undetected for so long?
To address this question, we will analyze the subject more mathematically. For an arbitrary velocity parameter $v/c$, performing covariant calculations of the radiation process is highly challenging. However, in certain cases, significant simplifications arise. One such case is the non-relativistic radiation regime. The non-relativistic asymptotic limit offers a fundamental simplification for covariant calculations. This is because the non-relativistic assumption justifies the dipole approximation, which is of great practical importance. When considering only the dipole component of radiation, all details of the electron trajectory are effectively ignored. Consequently, dipole radiation is completely insensitive to the distinction between covariant and non-covariant particle trajectories  [^59] Similarly, the ultrarelativistic asymptotic limit also simplifies covariant calculations. This is due to the paraxial approximation, which naturally follows from the ultrarelativistic assumption. More details on this effect will be discussed in the next chapter..
[…]
Suppose we are interested in the radiation generated by an electron and observed far away from it. In this case, it is possible to find a relatively simple expression for the electric field. We indicate the electron velocity in units of $c$ with $\vec{\beta}$, the electron trajectory in three dimensions with $\vec{r}(t)$ and the observation position with $\vec{r}_{0}$. Finally, we introduce the unit vector
$\vec{n}=\frac{\vec{r}_{0}-\vec{r}(t)}{|\vec{r}_{0}-\vec{r}(t)|}$ (35)
pointing from the retarded position of the electron to the observer. In the far zone, by definition, the unit vector $\vec{n}$ is nearly constant in time. If the position of the observer is far away enough from the charge, one can make the expansion $\left|\vec{r}_{0}-\vec{r}(t)\right|=r_{0}-\vec{n}\cdot\vec{r}(t)$.
We then obtain the following approximate expression for the the radiation field in the space-frequency domain:
[…]
where $\omega$ is the frequency, $(-e)$ is the negative electron charge and we make use of Gaussian units. A different constant of proportionality in Eq.(LABEL:revwied) and the well-known textbooks is to be ascribed to the use of different units and the definition of the Fourier transform.
Let us now examine in greater detail how the dipole radiation term arises. In the integrands of the expression for the radiation field amplitude, Eq. (LABEL:revwied), the time argument $\vec{r}(t)\cdot(\vec{n}/c)$ can be neglected if the trajectory of the charge changes insignificantly during this interval. Determining the conditions under which this approximation holds is straightforward. Let $a$ represent the characteristic size of the system. Then, the time delay can be estimated as $\vec{r}(t)\cdot(\vec{n}/c)\sim a/c$. To ensure that the charge distribution remains essentially unchanged over this duration, it is necessary that $a\ll\lambda$, where $\lambda$ is the wavelength of the emitted radiation. This condition can also be expressed in an alternative form: $v\ll c$, where $v$ denotes the characteristic velocity of the charges in the system.
We consider the radiation corresponding to the zeroth-order term in the expansion of Eq. (LABEL:revwied) in powers of $\vec{r}(t)\cdot(\vec{n}/c)$. In this approximation, all details of the electron trajectory $\vec{r}(t)$ are neglected. This is the essence of the dipole approximation, where the spatial scale of the electron’s motion is assumed to be much smaller than the wavelength of the emitted radiation. Under this condition, Eq. (LABEL:revwied) yields fields that closely resemble those predicted by instantaneous, non-retarded theories. Therefore, it is appropriate to use a non-covariant approach when analyzing dipole radiation.
However, this is only the first and most practically significant term. The remaining terms indicate that there are higher-order corrections to the dipole radiation approximation. Calculating these corrections requires detailed knowledge of the electron’s trajectory. In particular, to determine the correction to dipole radiation, we must use the covariant form of the trajectory rather than relying on a non-covariant approach. Nevertheless, corrections to multipole radiation are generally expected to be small. For instance, the covariant correction to quadrupole radiation is typically regarded as a contribution of higher order than the quadrupole term itself.
18.2 An Illustrative Example
It is beneficial to illustrate errors in standard coupling fields and particles in accelerator and plasma physics using a relatively simple example, where the core physical concepts remain clear and unobscured by unnecessary mathematical complexities. This example is primarily intended for readers with limited knowledge of accelerator and synchrotron radiation physics. Fortunately, the error in standard coupling fields and particle interactions can be explained in a straightforward manner.
An electron kicker setup serves as a practical example for illustrating the distinction between covariant and non-covariant trajectories. Consider a simple case where an ultrarelativistic electron, moving with velocity $v$ along the $z$-axis in the inertial (lab) frame, is subjected to a weak dipole magnetic field directed along the $x$-axis. For simplicity, we assume that the kick angle is small compared to $1/\gamma$, where $\gamma=1/\sqrt{1-v^{2}/c^{2}}$ is the Lorentz factor. This corresponds to the limit $\gamma\gg 1\gg\gamma v_{y}/v$. We begin with non-covariant particle tracking calculations. The electron’s trajectory, derived from the corrected form of Newton’s second law under the absolute time convention, does not account for relativistic effects. Consequently, as in classical Newtonian kinematics, the Galilean vector addition of velocities is employed. Non-covariant particle dynamics predicts that the electron’s direction changes as a result of the kick, while its speed remains constant (see Fig. 49). According to this model, the magnetic field $B\vec{e}_{x}$ affects only the direction of motion and not the magnitude of the velocity. After the kick, the components of the beam’s velocity are $(0,v_{y},v_{z})$, where $v_{z}=\sqrt{v^{2}-v_{y}^{2}}$. In the ultrarelativistic limit, where $\gamma\gg 1$, we can use a second-order approximation to obtain $v_{z}=v[1-v_{y}^{2}/(2v^{2})]=v[1-v_{y}^{2}/(2c^{2})]$.
Figure 49: Illustration of the difference between covariant and non-covariant trajectories. The setup shows the motion of a relativistic electron accelerated by a kicker field, assuming $\gamma\gg 1\gg\gamma v_{y}/v$ for simplicity. In the non-covariant particle tracking framework, the magnetic field $B\vec{e}_{x}$ affects only the direction of the electron’s motion without altering its speed.
In contrast, covariant particle tracking—which relies on Lorentz coordinates—yields different results for the electron’s velocity. To explore this, consider a sequence of passive Lorentz transformations used to track the motion of a relativistic electron accelerated by a kicker field. Let $S$ be the lab frame and $S^{\prime}$ a frame comoving with the electron at velocity $\vec{v}$ relative to $S$. Upstream of the kicker, the electron is at rest in the $S^{\prime}$ frame. This implies that $S^{\prime}$ is related to $S$ by a Lorentz boost $L(\vec{v})$, with $\vec{v}$ aligned along the $z$ axis. The boost transforms a four-vector even $X$ in space-time according to $X^{\prime}=L(\vec{v})X$. Now, let us analyze the particle’s dynamics from the perspective of the $S^{\prime}$ frame. In this frame, the electron remains at rest, while the kicker moves toward it with velocity $-\vec{v}$. Due to its motion, the kicker’s magnetic field generates an electric field perpendicular to it. As the kicker interacts with the stationary electron in $S^{\prime}$, the particle experiences a combination of perpendicular electric and magnetic fields.
We consider a small expansion parameter, $\gamma v_{y}/c\ll 1$, and retain terms up to second order—specifically, terms of order $(\gamma v_{y}/c)^{2}$—while neglecting higher-order contributions such as $(\gamma v_{y}/c)^{3}$. This corresponds to employing the second-order kick angle approximation. It is straightforward to show that the acceleration in crossed fields leads to a particle velocity $v^{\prime}_{y}=\gamma v_{y}$ along the $y$-axis and $v^{\prime}_{z}=-v(\gamma v_{y}/c)^{2}/2$ along the $z$-axis. Within this second-order approximation, relativistic corrections to velocity composition—typically arising at third order—do not appear.
[…]
According to standard textbooks, in relativistic kinematics, the composition of perpendicular velocities is analyzed using Einstein’s velocity addition theorem. We previously discussed the law of velocity composition in Section 15.5, where it was shown that, within the Lorentz coordinatization, velocity addition is governed by the Wigner rotation. In the ultrarelativistic approximation, a simple result emerges: $v_{tot}=v_{z}=v[1-v_{y}^{2}/(2c^{2})]$. This implies that a Lorentz transformation induces a rotation of the particle’s velocity $v_{z}$ by an angle approximately equal to $v_{y}/c$ (Fig. 50). This result highlights that the apparent contradiction is rooted in the presence of the Wigner rotation. As discussed in earlier chapters, the relation $dt^{\prime}=dt/\gamma$ is not valid when addressing the relativistic composition of perpendicular velocities within the Lorentz framework.
It is important to highlight another effective approach to covariant electron tracking. We now revisit the motion of a relativistic electron accelerated by the kicker field, analyzing it from the perspective of an inertial lab observer without changing reference frames. Specifically, we consider an electron in the lab frame undergoing acceleration due to the kicker field—an example of an active boost. In this scenario, we track the evolution of the system over time within a single reference frame.
The simplest synchronization method involves maintaining the same set of uniformly synchronized clocks as in the pre-kick state, without modification. This approach, based on the absolute time (or absolute simultaneity) convention, preserves simultaneity. The standard Galilean velocity addition rule is then applied to track the electron’s motion instant by instant as it moves through a constant magnetic field.
In Chapter 3, we discussed how absolute time coordinatization can be transformed into Lorentz coordinatization. By combining Galilean transformations with variable changes, we ultimately derive the Lorentz transformation within the framework of absolute time coordinatization in the lab frame. To further illustrate this, we now examine how distant lab clocks resynchronize during the electron’s acceleration. This analysis establishes a direct link between the electron’s speed reduction after the kick (in Lorentz coordinates) and time dilation. As previously noted, the Lorentz coordinate system is purely a conceptual construct. Hypothetical clocks are required only for applying Maxwell’s equations in synchrotron radiation calculations.
Suppose that, upstream of the kicker, we choose a Lorentz coordinate system in the lab frame. Immediately after the electron enters the magnetic field, its velocity changes by an infinitesimal amount $d\vec{v}$ along the $y$-axis. At this initial stage, Eq. (22) allows us to express the differential $d\vec{v}$ in terms of the differential $dt$ within the Lorentz coordinate system defined upstream of the kicker. If clock synchronization is fixed, this corresponds to adopting the absolute time convention. To maintain Lorentz coordinates in the lab frame—as discussed previously—we must perform a clock resynchronization by introducing an infinitesimal time shift. The simplest case arises when the kick angle $\theta_{k}$ is very small, allowing us to work up to second-order terms in $(\theta_{k}\gamma)^{2}$. This restriction greatly simplifies the calculations for two reasons. First, relativistic corrections due to the composition of non-collinear velocity increments only appear at order $(\gamma\theta_{k})^{3}$ and can therefore be neglected. Second, time dilation effects also enter only at higher orders. Thus, Eq. (22) enables us to express the small velocity change $\Delta\vec{v}$ due to the kick in the initial Lorentz coordinate system and defer clock resynchronization until after the kicker. Consequently, the post-kick motion can be described as a composition of two Lorentz boosts in the perpendicular $y$ and $z$ directions. The first boost imparts a velocity $v\theta_{k}\vec{e}_{y}$ along the $y$-axis, while the second applies an additional velocity $-(v\theta_{k}^{2}/2)\vec{e}_{z}$ along the $z$ axis. The second-order approximation ensures that the two boosts commute.
Figure 50: Trajectory of a relativistic electron accelerated by a kicker field. To describe the kicker’s influence on the electron’s motion using a Lorentz transformation, it is necessary to resynchronize the clock by introducing a time shift and applying a time scale adjustment.
To maintain a Lorentz coordinate system in the lab frame after the electron receives a transverse kick, a resynchronization of clocks is required. In this context, operations involving the rule clock structure in the lab frame can be interpreted as a change of variables governed by the transformation in Eq. 13: $y_{L}=\gamma_{y}y$, $t_{L}=(t/\gamma_{y}+\gamma_{y}yv_{y}/c^{2})$, where $\gamma_{y}=1+v_{y}^{2}/(2c^{2})$ and the electron’s displacement is given by $y=v_{y}t$.
This transformation effectively rescales the time variable—adjusting the synchronization of all clocks from $t$ to $\gamma_{y}t$, with $\gamma_{y}\simeq 1+\theta_{k}^{2}/2$. As a result, it becomes evident that the electron’s downstream speed is no longer independent of its transverse motion in the magnetic field (see Fig. 50). Although no second-order relativistic correction appears in the transverse ($y$-axis) velocity component, the longitudinal component is modified. Specifically, the longitudinal velocity $v_{z}$ changes to $v_{z}/\gamma_{y}$ with $v_{z}=v(1-\theta_{k}^{2}/2)$, leading to $v_{z}/\gamma_{y}=v(1-\theta_{k}^{2})$. As a result, the total electron speed in the lab frame, after clock resynchronization downstream of the kicker, decreases from $v$ to $v(1-\theta_{k}^{2}/2)$.
This analysis is performed in the limit of a small kick angle $\gamma v_{y}/c\ll 1$. Nevertheless, even in this simplified case and for a single electron, we are able to clearly demonstrate the distinction between covariant and non-covariant particle trajectories.
18.3 Electron Motion Accelerated by a Kicker in a Bending Magnet
In our relativistic, yet non-covariant, analysis of electron motion in a magnetic field, the electron maintains the same velocity—and therefore the same relativistic factor $\gamma$—both upstream and downstream of the kicker. Suppose the electron then enters a bending magnet, i.e., a region with a uniform magnetic field directed along the $y$-axis (see Fig. 51). The resulting motion within the bending magnet closely resembles that predicted by non-relativistic dynamics, with the only difference being the presence of the relativistic factor $\gamma$ in the expression for the cyclotron frequency: $\omega_{c}=eB/(m\gamma)$. The curvature radius $R$ of the trajectory follows from the relation $v_{\perp}/R=\omega_{c}$, where $v_{\perp}=v(1-\theta_{k}^{2}/2)$ is the component of the velocity perpendicular to the magnetic field $\vec{B}=B\vec{e}_{y}$. According to non-covariant particle tracking, the correction to the radius $R$ after the kick is of order $\theta_{k}^{2}$.
Figure 51: Geometry of radiation production in a bending magnet. In non-covariant particle tracking, the kicker’s magnetic field $B\vec{e}_{x}$ can change the direction of the electron’s motion but does not affect its speed.
At first glance, one might expect that in covariant particle tracking, the total speed of an electron in the lab frame—after passing through the kicker—would decrease from $v$ to $v(1-\theta_{k}^{2}/2)$. This, in turn, would suggest a corresponding decrease in the magnitude of the three-momentum $|\vec{p}|$ from $m\gamma v$ to $m\gamma v(1-\gamma^{2}\theta_{k}^{2}/2)$, based on our approximation. However, such a change in momentum would imply a correction to the bending radius $R$ on the order of $\gamma^{2}\theta_{k}^{2}$, leading to a stark contradiction with the radius calculated using noncovariant tracking. Since the curvature radius in the bending magnet has an objective, convention-invariant meaning, this apparent discrepancy poses a paradox. The resolution lies in recognizing that, in Lorentz coordinates, the momentum three-vector $\vec{p}$ transforms as the spatial part of the four-momentum $p_{\mu}$ under Lorentz boosts. To analyze the motion of the relativistic electron affected by the kicker, we consider a composition of Lorentz boosts that follow the particle’s trajectory. Under such a transformation, the longitudinal component of the momentum remains effectively unchanged—accurate to order $\theta_{k}^{2}$.
Let us verify the validity of this assertion. The four-momentum is given by $p_{\mu}=[\mathcal{E}/c,\vec{p}]$. Consider the Lorentz frame $S^{\prime}$, which is at rest with respect to the electron upstream of the kicker. In the special case where the electron is initially at rest in this frame, its four-momentum is $p_{\mu}^{\prime}=[mc,\vec{0}]$. We now focus on the dynamics in frame $S^{\prime}$. Acceleration due to the crossed electric and magnetic fields of the kicker induces a velocity component $v_{y}^{\prime}=\gamma v_{y}$ along the $y$-axis, and a second-order longitudinal component $v_{z}^{\prime}=-v(\gamma v_{y})^{2}/2$ along the $z$-axis. After the interaction with the kicker, the four-momentum in $S^{\prime}$ becomes: $p_{\mu}^{\prime}=[mc+mv_{y}^{\prime 2}/(2c),0,mv_{y}^{\prime},mv_{z}^{\prime}]$, where the expression is evaluated to second order in $(\gamma v_{y}/c)^{2}$, consistent with our approximation scheme. We observe that the transverse acceleration results in an increase in the time-like component of the four-momentum—that is, the energy of the electron. Specifically, the energy increases from $mc^{2}$ to $mc^{2}+m(\gamma v_{y})^{2}/2$. Recall that frame $S^{\prime}$ is related to the lab frame $S$ via a Lorentz boost. Under a boost to a frame moving with velocity $\vec{v}=-v\vec{e}_{z}$, the longitudinal component of the momentum (which is normal to the magnetic field of the bending magnet) transforms as: $p_{z}=\gamma(p_{z}^{\prime}+vp_{0}^{\prime}/c)=\gamma mv$. This shows that, within our approximation, the momentum component along the $z$-axis remains unchanged. Furthermore, applying the Lorentz transformation to the time component yields: $p_{0}=\gamma(p^{\prime}_{0}+vp^{\prime}_{z})=\gamma mc$, consistent with our earlier result.
It is intriguing to examine the implications of having two distinct approaches—covariant and noncovariant—that yield the same particle three-momentum. The key insight is that both frameworks accurately describe the same physical reality. The curvature radius of the trajectory in a magnetic field, and consequently the three-momentum, possesses an objective meaning, making it invariant under different conventions.
18.4 Redshift of the Synchrotron Radiation Critical Frequency
Next, we explore the intriguing problem of synchrotron radiation emission in a bending magnet, both with and without the influence of a preceding kick. We will focus on the physical interpretation without delving into computational details (see the next chapter for a more thorough analysis). Consider the setup illustrated in Fig. 51. An ultrarelativistic electron is initially moving along the $z$-axis in the laboratory frame. Before entering a uniform magnetic field directed along the $y$-axis (i.e., a bending magnet), the electron receives a small transverse kick from a weak dipole field oriented along the $x$-axis. An electron undergoing acceleration along a curved trajectory emits electromagnetic radiation. At relativistic speeds, this radiation is concentrated into a narrow cone, tangent to the electron’s path. Furthermore, the radiation amplitude is significantly enhanced in this direction—a phenomenon known as Doppler boosting. Synchrotron radiation arises specifically when a relativistic electron is accelerated by a bending magnet. In the ultrarelativistic limit, the key features of the emitted spectrum can be intuitively understood to depend primarily on the small difference between the speed of the electron and the speed of light.
We now turn to the radiation emitted by an ultrarelativistic electron traversing a bending magnet. Consider the case in which the electron is moving toward the observer (Fig. 52). The electromagnetic source follows a trajectory $\vec{x}(t^{\prime})$ over time, where $t^{\prime}$ denotes the emission time of the radiation. However, because electromagnetic signals propagate at a finite speed (the speed of light), a signal emitted at time $t^{\prime}$ from position $\vec{x}(t^{\prime})$ reaches the observer at a later time $t$. Consequently, the observer perceives the motion of the source as a function of $t$, not $t^{\prime}$. It is important to note that the prime notation here indicates retarded time and should not be confused with primes denoting quantities in a Lorentz-transformed frame, as used in subsequent sections.
Let the coordinates of the electron be denoted by $(x,y,z)$, where the $z$-axis is aligned with the direction of observation. We continue to assume that the detector is located far from the radiation source. At a given moment $t^{\prime}$, the electron’s position components are $x(t^{\prime})$, $y(t^{\prime})$, and $z(t^{\prime})$. The distance to the observer is approximately $R(t^{\prime})=R_{0}-z(t^{\prime})$. If we denote the observation time by $t$, then $t^{\prime}$ is not equal to $t$; rather, it is delayed by the time it takes light to travel from the electron to the observer. Disregarding a constant delay—equivalent to shifting the time origin—we find the relation $ct=ct^{\prime}-z(t^{\prime})$. Our goal is to express $x$ as a function of the observation time $t$, rather than the emission time $t^{\prime}$. Assuming $c-v\ll c$, we can use the approximation: $dt/dt^{\prime}=(c-v\cos\theta)/c\simeq(1-v/c+\theta^{2}/2)\simeq(1/2)(1/\gamma^{2}+\theta^{2})$, where $\theta$ is the observation angle (Fig. 52).
[…]
Figure 52: Geometry for synchrotron radiation from a bending magnet.
Figure 53: Formation length for the bend.
It is evident from the preceding discussion that, according to conventional synchrotron radiation theory, introducing a kick to the electron’s motion results merely in a rigid rotation of the angular distribution of the emitted radiation, aligning it with the new direction of motion. This conclusion is intuitive: the electron retains its speed after the kick and, due to the Doppler effect, emits radiation along the new trajectory.
However, a more rigorous treatment—based on the proper coupling between fields and particles—reveals a noteworthy prediction of synchrotron radiation theory for the setup described above. Specifically, there is a red shift in the critical frequency of the synchrotron radiation observed in the kicked direction. To demonstrate this effect, we turn to a covariant formulation that explicitly incorporates Lorentz transformations. When the kick is applied, covariant particle tracking predicts a nonzero red shift in the critical frequency. This shift arises because, in Lorentz coordinates, the electron’s velocity decreases from $v$ to $v-v\theta_{k}^{2}/2$, while the speed of light remains unchanged at the electrodynamic constant $c$. The resulting red shift in the critical frequency $\omega_{cr}\simeq R/(c\gamma^{3})$ can be approximated by the relation $\Delta\omega_{cr}/\omega_{cr}\simeq-\gamma^{2}v_{y}^{2}/c^{2}=-\gamma^{2}\theta_{k}^{2}$. Here, we observe a second-order correction in $\theta_{k}^{2}$ that is, however, multiplied by a large factor $\gamma^{2}$.
It is worth noting, however, that there exists an alternative and equally satisfactory explanation for the redshift. This interpretation employs a non-covariant framework, which considers non-covariant particle trajectories and applies Galilean transformations to the electromagnetic field equations. In this approach, the electron velocity remains unchanged according to non-covariant particle tracking. However, Maxwell’s equations are not invariant under Galilean transformations, leading to a change in the speed of light. Specifically, the velocity of light increases from $c$ (before the kick) to $c(1+\theta_{k}^{2}/2)$ (after the kick). This variation arises because, under the absolute time convention, clocks are not resynchronized after the kick. As a result, the speed of light differs from the electrodynamical constant $c$. Nevertheless, the ratio of the electron velocity to the speed of light remains invariant across different synchronization conventions—that is, it does not depend on the method of distant clock synchronization or the rate of the clocks. Our calculations demonstrate that both covariant and non-covariant treatments—when the correct coupling between fields and particles is used—yield the same prediction for the redshift. This prediction is convention-invariant and depends solely on the dimensionless parameter $v_{coord}/c_{coord}=v(1-\theta^{2}_{k}/2)/c=v/[c(1+\theta^{2}_{k}/2)]$, where $v_{coord}$ is the coordinate velocity of the electron, and $c_{coord}$ is the coordinate velocity of light.
To validate the predictions of our synchrotron radiation theory, we propose an experimental test using third-generation synchrotron radiation sources. Synchrotron radiation from bending magnets spans a broad frequency range; however, employing narrow-bandwidth sources offers a more promising approach for studying redshifts in the radiation spectrum. Narrower bandwidths enhance the sensitivity of output intensity to redshift effects and reduce the demands on beam kicker strength and photon beamline aperture. Undulators, which generate quasi-monochromatic synchrotron radiation, are particularly well-suited for this purpose. By forcing the electron beam into a periodic undulating path, they induce interference effects that significantly narrow the emitted bandwidth. Since undulators typically consist of many periods, the resulting bandwidth scales inversely with the number of periods. Consequently, insertion devices at third-generation facilities provide an effective means to boost sensitivity to redshift signals, even at small kick angles, $\theta_{k}<1/\gamma$. In essence, we argue that, despite decades of experimental work, synchrotron radiation theory remains insufficiently confirmed, and a more precise test is both timely and necessary.
18.5 Discussion
We now wish to emphasize that when an electron is accelerated in the lab frame by a kicker, this acceleration is encoded in its covariant trajectory. A compelling question arises: Where, physically, is the information about the electron’s acceleration located when using absolute time coordinatization?
[…]
This apparent paradox—of an unchanged velocity despite acceleration—can be explained through dynamical reasoning involving the electron’s non-covariant trajectory. Without entering into a full derivation, we summarize the key insight as follows:
For a rapidly moving electron, the transverse electric and magnetic fields are nearly equal in magnitude and orthogonal to each other—indistinguishable from the fields of an actual radiation beam. On a microscopic level, when the electron passes through a kicker, its electromagnetic fields are perturbed and now carry imprints of that acceleration.
Under the framework of classical (Galilean) kinematics, the orientation of the virtual radiation phase front remains unchanged. However, since Maxwell’s equations are not invariant under Galilean transformations—as discussed throughout this article—the adoption of such kinematics necessitates the use of anisotropic field equations. Consequently, although the phase front remains planar, the direction of propagation is no longer perpendicular to it. In other words, the radiation beam’s direction of motion diverges from the normal to its phase front.
Within the absolute time synchronization convention, electrodynamics predicts that the virtual radiation beam propagates along the direction of the kick, with a phase front tilt $\theta_{k}$. This tilt is the crucial clue to resolving the paradox. The key insight is this: The information about the electron’s acceleration is not lost—it is embedded in the perturbation of the electron’s self-fields. Mathematically, this information is encoded in the tilt of the phase front of the virtual radiation beam.
Let us now briefly consider some additional, intriguing cases. Previously, we examined a simple scenario in which the kick angle is small. In such cases, when an electron passes through a weak kicker, it emits no appreciable synchrotron radiation. More precisely, the radiation is indistinguishable from the electron’s intrinsic electromagnetic fields. We can therefore assume that the electron remains in a “field-dressed” state — still accompanied by its usual virtual radiation fields — after passing through the weak kicker.
Now, let us turn to a different situation. Suppose an ultrarelativistic electron in the laboratory frame is deflected by a dipole magnetic field with a kick angle much larger than $1/\gamma$, This setup corresponds to a typical synchrotron radiation scenario.
At the exit of the strong kicker, the electron emerges effectively “naked” or “field-free,” meaning it is no longer accompanied by its usual virtual field structure. A process then begins in which the “field-dressed” state is gradually re-established — that is, the characteristic electromagnetic fields associated with a fast-moving charge begin to reform — over a certain formation length in the downstream straight section.
At this point, information about the electron’s recent acceleration is carried away by the emitted synchrotron radiation. However, the electron’s original self-fields and its acceleration history are effectively erased in the strong kick. As a result, the electromagnetic fields surrounding the electron downstream of the kicker no longer encode this history. In other words, the electron’s covariant trajectory beyond the kicker contains no imprint of the preceding acceleration. We will revisit this topic in more detail in the next chapter.
19 Synchrotron Radiation
19.1 Introductory Remarks
Accelerator physics has traditionally been framed within the context of classical (Newtonian) kinematics, which is fundamentally incompatible with Maxwell’s equations. Here, we aim to revisit this perspective by applying modern insights from dynamics and electrodynamics to explore a central question in greater detail: Why did the error in synchrotron radiation theory remain undetected for so long?
The electromagnetic radiation phenomena under consideration are inherently complex. In a general setup, covariant calculations of the radiation process can be highly intricate. However, certain configurations allow for significant simplifications. One such case is the synchrotron radiation setup. Much like the non-relativistic limit, the ultrarelativistic limit introduces essential simplifications into covariant calculations. This is because the ultrarelativistic regime justifies the paraxial approximation: the formation length of the radiation is much longer than its wavelength, and as a result, radiation is emitted within a narrow cone of angles on the order $1/\gamma$ or smaller. Consequently, the small-angle approximation becomes valid. In this regime, the transverse velocity is small compared to the speed of light, allowing us to use a second-order relativistic approximation for the transverse motion. Rather than using a small total velocity parameter $(v/c)$ as in the non-relativistic case, we now deal with a small transverse velocity parameter $(v_{\perp}/c)$. The next step is to analyze the longitudinal motion using the same approach. Notably, in synchrotron radiation, the longitudinal dynamics are particularly straightforward. When expanding transformations up to second order in $(v_{\perp}/c)^{2}$, relativistic corrections to the longitudinal motion are absent within this approximation.
In the covariant framework, relativistic kinematic effects associated with synchrotron radiation appear as successive corrections at different orders of approximation:
$\cdot$ First order $(v_{\perp}/c)$: Effects such as the relativity of simultaneity and the Wigner rotation arise. In the ultrarelativistic limit, the Wigner rotation manifests already at first order and is a direct consequence of the relativity of simultaneity.
$\cdot$ Second order $(v_{\perp}/c)^{2}$: Effects such as time dilation and the relativistic correction to the velocity addition law emerge. Relativistic correction in the law of composition of velocities, which already appears in the second order, results directly from the time dilation.
The first-order kinematic term $(v_{\perp}/c)$ plays a significant role primarily in the description of coherent radiation from a modulated electron beam. In a storage ring, however, the longitudinal positions of electrons within a bunch are essentially uncorrelated. As a result, the radiation emitted by different electrons is also uncorrelated. In this case, the total radiated power is simply the sum of the individual contributions—intensities are added rather than electric fields. For incoherent synchrotron radiation, where one considers the motion of a single ultrarelativistic electron in a constant magnetic field, relativistic effects influence only the second-order kinematic terms, specifically $(v_{\perp}/c)^{2}$.
19.2 Paraxial Approximation of the Radiation Field
The general method to derive the frequency spectrum is to transform the electric field from the time domain to the frequency domain by the use of the Fourier transform. First, let us rewrite Eq. (LABEL:revwied) as follows
[…]
Eq. (LABEL:revwied) and Eq. (LABEL:rrevwied) are equivalent but include different integrands. This is no mistake, as different integrands can lead to the same integral (see Appendix A1).
We call $z_{0}$ the observation distance along the optical axis of the system, while $(x_{0},y_{0})$ fixes the transverse position of the observer. Using the complex notation, in this and in the following sections we assume that the temporal dependence of fields with a certain frequency is of the form:
$\vec{E}\sim\vec{\bar{E}}(z_{0},x_{0},y_{0},\omega)\exp(-i\omega t)~{}.$ (40)
With this choice for the temporal dependence we can describe a plane wave traveling along the positive $z$-axis with
$\vec{E}=\vec{E}_{a}\exp\left(\frac{i\omega}{c}z_{0}-i\omega t\right)~{}.$ (41)
In the following, we will always assume that the ultra-relativistic approximation is satisfied, which is the case for SR setups. As a consequence, the paraxial approximation applies too. The paraxial approximation implies a slowly varying envelope of the field with respect to the wavelength. It is therefore convenient to introduce the slowly varying envelope of the transverse field components as
$\vec{\widetilde{E}}(z_{0},x_{0},y_{0},\omega)=\vec{\bar{E}}(z_{0},x_{0},y_{0},\omega)\exp{\left(-i\omega z_{0}/c\right)}~{}.$ (42)
We will now replace all vectors by their components to obtain directional dependency of the synchrotron radiation. The emission angle $\theta=\sqrt{\theta_{x}^{2}+\theta_{y}^{2}}$ is taking with respect to the $z$-axis. Here $\theta_{x}=x_{0}/z_{0}$ is the observation angle projected onto the $x-z$ plane, $\theta_{y}=y_{0}/z_{0}$ is the observation angle projected onto the $y-z$ plane. The components of the unit vector $\vec{n}$ can be approximated by $n_{z}=1-\theta_{x}^{2}/2-\theta_{y}^{2}/2$, $n_{x}=\theta_{x}$, $n_{y}=\theta_{y}$, so $\vec{n}\cdot(\vec{r}_{0}-\vec{r})=(z_{0}-z)(1-\theta_{x}^{2}/2-\theta_{y}^{2}/2)-x\theta_{x}-y\theta_{y}$. We consider the motion in a static magnetic field. According to conventional particle tracking the magnitude of the velocity is constant and is equal $v=ds/dt=\mathrm{const.}$, where $s(z)$ is the longitudinal coordinate along the path. The transverse components of the envelope of the field in Eq. (LABEL:rrevwied) in the far zone and paraxial approximation finally becomes
[…]
where the total phase $\Phi_{T}$ is
$\Phi_{T}=\omega\left[{s(z^{\prime})\over{v}}-{z^{\prime}\over{c}}\right]$ (44)
$+\frac{\omega}{2c}\left[z_{0}(\theta_{x}^{2}+\theta_{y}^{2})-2\theta_{x}x(z^{\prime})-2\theta_{y}y(z^{\prime})+z^{\prime}(\theta_{x}^{2}+\theta_{y}^{2})\right]~{}.$ (45)
Here $v_{x}(z^{\prime})$ and $v_{y}(z^{\prime})$ are the horizontal and the vertical components of the transverse velocity of the electron, $x(z^{\prime})$ and $y(z^{\prime})$ specify the transverse position of the electron as a function of the longitudinal position, $\vec{e}_{x}$ and $\vec{e}_{y}$ are unit vectors along the transverse coordinate axis.
19.3 Undulator Radiation
Traditionally, courses on synchrotron radiation theory begin by following the historical development of the field, starting with radiation from bending magnets. In contrast, we will begin this chapter by exploring a more advanced topic: the theory of synchrotron radiation from undulator setups, where covariant calculations of the radiation process become notably straightforward.
To generate specific characteristics of synchrotron radiation, special insertion devices known as undulators are often employed. In such setups, the resonance approximation—which is always applicable—significantly simplifies the theoretical treatment. This approximation complements, rather than replaces, the paraxial one, and leverages another large parameter: the number of undulator periods $N_{w}\gg 1$.
The frequency of the radiation emitted by a particle traversing an undulator can be derived by analyzing the interference between radiation produced at successive undulator periods (see Fig. 54). This frequency is subject to Doppler shifting, and the shortest wavelength is observed along the undulator axis. Under the resonance approximation, radiation at this shortest wavelength is emitted within an angular cone much narrower than $1/\gamma$ (see Fig. 55). This sharply defines the relevant observation angles. Outside the diffraction angle, the intensity falls to zero with an accuracy $1/N_{w}$.
This leads us to an important question: Why did the error in insertion device theory remain undetected for so long? We address this in detail below, focusing specifically on radiation within the central cone (see Fig. 55). In this approximation, the electron trajectory is effectively neglected. Regardless of the strength of the undulator (i.e., the undulator parameter), the amplitude of the electron’s oscillation remains much smaller than the diffraction-limited size of the radiation at the undulator exit—again, a consequence of $N_{w}\gg 1$. As a result, undulator radiation theory, in this context, produces fields closely resembling those predicted by dipole-like instantaneous emission models. For practical purposes, especially when describing radiation into the central cone, the conventional (non-covariant) approach has therefore been considered adequate.
However, there is a specific scenario in which the conventional theory breaks down. The covariant formulation predicts a non-zero red shift in the resonance frequency when the electron motion experiences perturbations—specifically, transverse kicks with respect to the longitudinal axis. Experimental observations support this correction for spontaneous undulator emission, validating the covariant approach in such cases.
19.3.1 Conventional Theory
Equation (43) provides a general framework for characterizing the far-field radiation from an electron following an arbitrary trajectory. In this section, we present a straightforward derivation of the frequency-domain representation of the radiated field produced by an electron traversing a planar undulator. The magnetic field along the undulator axis is given by
$\vec{B}(z)=\vec{e}_{y}B_{w}\cos(k_{w}z)~{},$ (46)
Here $k_{w}=2\pi/\lambda_{w}$, and $\lambda_{w}$ is the undulator period. The Lorentz force is used to derive the equation of motion of the electron in the presence of a magnetic field. Integration of this equation gives
$v_{x}(z)=-{c\theta_{s}}\sin(k_{w}z)=-\frac{c\theta_{s}}{2i}\left[\exp(ik_{w}z)-\exp(-ik_{w}z)\right]~{}.$ (47)
Here $\theta_{s}=K/\gamma$, where $K$ is the deflection parameter defined as
$K=\frac{e\lambda_{w}B_{w}}{2\pi mc^{2}}~{},$ (48)
$m$ being the electron mass at rest and $B_{w}$ being the maximal magnetic field of the undulator on the axis.
In this case, the electron path is given by
$x(z)=r_{w}\cos(k_{w}z)~{},$ (49)
where $r_{w}=\theta_{s}/k_{w}$ is the oscillation amplitude.
We write the undulator length as $L=N_{w}\lambda_{w}$, where $N_{w}$ is the number of undulator periods. With the help of Eq. (43) we obtain an expression, valid in the far zone:
[…]
Here
$\Phi_{T}=\left({\omega\over{2c\bar{\gamma}_{z}^{2}}}+{\omega\theta^{2}\over{2c}}\right)z^{\prime}-{K\theta_{x}\over{\gamma}}{\omega\over{k_{w}c}}\cos(k_{w}z^{\prime})-{K^{2}\over{8\gamma^{2}}}{\omega\over{k_{w}c}}\sin(2k_{w}z^{\prime})~{},$ (52)
where the average longitudinal Lorentz factor $\bar{\gamma}_{z}$ is defined as
$\bar{\gamma}_{z}=\frac{\gamma}{\sqrt{1+K^{2}/2}}~{}.$ (54)
The choice of the integration limits in Eq. (51) implies that the reference system has its origin in the center of the undulator.
Figure 54: Constructive interference of radiation from the successive poles
Figure 55: Geometry for radiation production from an undulator
Typically, calculating the intensity distribution using Eq. (51) alone is insufficient, as it neglects additional contributions—both interfering and non-interfering—from other segments of the electron’s trajectory. To obtain the total field in a given setup, one must account for the full electron path and include these extra terms alongside Eq. (51). However, there are specific scenarios where the contribution from Eq. (51) dominates over the others. In such cases, Eq. (51) can be considered to have an independent physical significance.
One such situation arises when the resonance approximation is applicable. This approximation does not replace the paraxial approximation, which is based on the condition $\gamma^{2}\gg 1$, but is instead used in conjunction with it. It exploits another typically large parameter—the number of undulator periods, $N_{w}\gg 1$. Under these conditions, the integral over $dz^{\prime}$ in Eq. (51) simplifies significantly. This simplification occurs regardless of the frequency of interest, due to the long integration range compared to the characteristic scale of the undulator period.
A well-known expression for the angular distribution of the first harmonic field in the far zone (see Appendix A2 for a detailed derivation) can be derived from Eq. (51). This expression is axisymmetric and can therefore be represented as a function of a single observation angle, $\theta$, where $\theta^{2}=\theta_{x}^{2}+\theta_{y}^{2}$. The resulting distribution for the slowly varying envelope of the electric field is given by:
[…]
Here $\omega=\omega_{r}+\Delta\omega$, $C=k_{w}\Delta\omega/\omega_{r}$ and
$\omega_{r}=2k_{w}c\bar{\gamma}_{z}^{2}~{},$ (57)
is the fundamental resonance frequency. Finally $A_{JJ}$ is defined as
$A_{JJ}=J_{0}\left(\frac{K^{2}}{4+2K^{2}}\right)-J_{1}\left(\frac{K^{2}}{4+2K^{2}}\right)~{},$ (58)
$J_{n}$ being the $n$-th order Bessel function of the first kind. The integration over longitudinal coordinate can be carried out leading to the well-known final result:
${\vec{\widetilde{E}}}(z_{0},\vec{\theta})=-\frac{K\omega eL}{2c^{2}z_{0}\gamma}A_{JJ}\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\mathrm{sinc}\left[\frac{L}{2}\left(C+\frac{\omega\theta^{2}}{2c}\right)\right]\vec{e}_{x}~{},$ (59)
where $\mathrm{sinc}(\cdot)\equiv\sin(\cdot)/(\cdot)$. Therefore, the field is horizontally polarized and azimuthal symmetric. Eq. (60) describes a field with a spherical wavefront centered in the middle of the undulator.
19.3.2 Why Did the Error in Insertion Device Theory Remain Undetected so Long?
We have seen that, in full generality, the expression for the undulator field in the far zone under the ultrarelativistic (i.e., paraxial) approximation can be written as Eq. (141). Within the resonance approximation ($N_{w}\gg 1$), and for frequencies near the first harmonic, this expression simplifies to the well-known form given by Eq. (60), where the field is horizontally polarized and exhibits azimuthal symmetry. The angular divergence of the radiation is significantly smaller than the characteristic angle $1/\bar{\gamma}_{z}$.
This narrow divergence arises from the resonance nature of undulator radiation, mathematically expressed through the sinc-like term $\sin(\cdot)/(\cdot)$. To characterize the angular width of the radiation peak around the forward direction ($\theta=0$), we introduce a small angular displacement $\Delta\theta$. By identifying the first zero of the $\sin(\cdot)/(\cdot)$ function at $C=0$, we can define the natural angular width of the first harmonic radiation, denoted $\theta_{c}$. The resulting cone of radiation with aperture $\theta_{c}$ is commonly referred to as the central cone. It can be shown that the angular width satisfies $\theta_{c}^{2}=1/(2N_{w}\bar{\gamma}_{z}^{2})\ll 1/\bar{\gamma}_{z}^{2}$.
We now aim to determine the characteristic transverse size of the radiation field at the exit of the undulator. Radiation emitted by the magnetic poles interferes coherently along the undulator axis, with constructive interference occurring within an angle of approximately $\sqrt{c/(\omega L_{w})}$. This corresponds to a transverse interference size at the undulator exit on the order of $\sqrt{cL_{w}/\omega}$. Meanwhile, the amplitude of the electron’s transverse oscillation is given by $r_{w}=c\theta_{s}/k_{w}=cK/(\gamma k_{w})$. Comparing this with the interference size, we find $r_{w}^{2}/(cL_{w}/\omega)=K^{2}\omega/(L_{w}k_{w}^{2}\gamma^{2})=K^{2}/[\pi N_{w}(1+K^{2}/2)]\ll 1$, where we used the relation $\gamma^{2}=(1+K^{2}/2)\bar{\gamma}_{z}^{2}$. This inequality holds regardless of the value of $K$, since $N_{w}\gg 1$. Therefore, the electron’s oscillation amplitude is always much smaller than the diffraction-limited size of the radiation field at the undulator exit.
We analyze the radiation associated with the first-order term in the expansion of Eq. (147) in powers of the small parameter $[K\theta_{x}\omega/(\gamma k_{w}c)]$ (see Eq. (150)). However, this approximation neglects the term $K\theta_{x}\omega\cos(k_{w}z^{\prime})/(\gamma k_{w}c)$, and as a result, all information about the transverse electron trajectory in the phase factor of Eq. (53) is lost. Under this approximation, the scale of the electron’s orbit is much smaller than the radiation’s diffraction size, and Eq. (60) yields fields that closely match those predicted by dipole radiation theory. Therefore, we consider the non-covariant approach sufficient for describing transverse electron motion in this context.
Several important points can be made regarding the result discussed above. As previously explained, by considering only the radiation emitted within the central cone, we overlook critical information about the electron’s transverse motion. To present a complete analysis, we must also account for the effects of acceleration along the $z$-direction—that is, along the undulator axis. We assume that the transverse velocity $v_{\perp}(z)$ is small compared to the speed of light $c$, and we introduce $v_{\perp}/c$ as a small expansion parameter. In this framework, we neglect terms of order $(v_{\perp}/c)^{3}$, corresponding to a second-order relativistic approximation for the transverse dynamics.
It is worth noting that, under the ultrarelativistic approximation, analyzing the longitudinal motion is considerably simpler than analyzing the transverse motion. In a constant magnetic field, the electron acquires a transverse velocity $v_{\perp}$, and its longitudinal velocity is reduced by $\Delta v_{z}=-v(v_{\perp}/c)^{2}/2$, where $v$ is the total velocity.
When we apply the relevant transformations consistent with a second-order approximation in $(v_{\perp}/c)^{2}$, no relativistic correction to the longitudinal motion arises. Consequently, within this level of approximation, the distinction between covariant and non-covariant constrained electron trajectories has no impact on the undulator radiation observed in the central cone.
19.3.3 Influence of the Kick According to Conventional Theory
Equation (60) can be extended to describe a particle with an initial offset $\vec{l}$ and deflection angle $\vec{\eta}$ relative to the longitudinal axis, under the assumption that the magnetic field in the undulator is independent of the particle’s transverse position (see Appendix A3). Although this generalization can be derived directly from Eq. (161), it is often more efficient to apply certain intuitive geometrical arguments that are consistent with a rigorous mathematical treatment.
Consider first the effect of a transverse offset $\vec{l}$ with respect to the longitudinal axis $z$. Since the particle experiences the same magnetic field, the far-field radiation pattern is simply shifted by $\vec{l}$. As a result, Eq. (60) can be generalized by replacing the transverse observation coordinate $\vec{r}_{0}$ with $\vec{r}_{0}-\vec{l}$. Equivalently, the observation angle $\vec{\theta}=\vec{r}_{0}/z_{0}$ should be substituted with $\vec{\theta}-\vec{l}/z_{0}$, leading to:
$\widetilde{{E}}\left(z_{0},\vec{l},\vec{\theta}\right)$ $=$ $-\frac{K\omega eL}{2c^{2}z_{0}\gamma}A_{JJ}\exp\left[i\frac{\omega z_{0}}{2c}\left|\vec{\theta}-\frac{\vec{l}}{z_{0}}\right|^{2}\right]\mathrm{sinc}\left[\frac{\omega L\left|\vec{\theta}-\left({\vec{l}}/{z_{0}}\right)\right|^{2}}{4c}\right]~{}.$ (61)
Let us now examine the effect of a deflection angle $\vec{\eta}$. Assuming the magnetic field experienced by the electron remains independent of its transverse position, the electron’s trajectory continues to follow a sinusoidal path. However, the effective undulator period is modified and becomes $\lambda_{w}/\cos(\eta)\simeq(1+\eta^{2}/2)\lambda_{w}$, due to the projection of the undulator axis onto the new trajectory. This leads to a relative red shift in the resonant wavelength given by $\Delta\lambda/\lambda\sim\eta^{2}/2$.
In typical scenarios of interest, the deflection angle can be approximated as $\eta\sim 1/\gamma$. Consequently, the relative red shift scales as $\Delta\lambda/\lambda\sim 1/\gamma^{2}$. This should be compared with the intrinsic relative bandwidth of the resonance, which is approximately $\Delta\lambda/\lambda\sim 1/N_{w}$, where $N_{w}$ denotes the number of undulator periods. For instance, if $\gamma>10^{3}$, the red shift induced by the deflection angle becomes negligible in all practical situations.
Therefore, introducing a deflection angle effectively corresponds to a rigid rotation of the entire system. When performing this rotation, it is important to note that the phase factor in Eq. (61)—which represents a spherical wavefront emanating from $z=0$—remains invariant under rotation. In contrast, the argument of the $\mathrm{sinc}(\cdot)$ function in Eq. (61) is altered by the rotation, as the point $(z_{0},0,0)$ is mapped to $(z_{0},-\eta_{x}z_{0},-\eta_{y}z_{0})$. As a result, after the rotation, Eq. (61) transforms accordingly:
$\widetilde{{E}}\left(z_{0},\vec{\eta},\vec{l},\vec{\theta}\right)=-\frac{K\omega eLA_{JJ}}{2c^{2}z_{0}\gamma}\exp\left[i\frac{\omega z_{0}}{2c}\left|\vec{\theta}-\frac{\vec{l}}{z_{0}}\right|^{2}\right]\mathrm{sinc}\left[\frac{\omega L\left|\vec{\theta}-\left({\vec{l}}/{z_{0}}\right)-\vec{\eta}\right|^{2}}{4c}\right]$ (62)
Finally, in the far-zone regime, we can take the limit $l/z_{0}\ll 1$, which justifies neglecting the term ${\vec{l}}/{z_{0}}$ in the argument of the $\mathrm{sinc}(\cdot)$ function, as well as the quadratic phase term $\omega l^{2}/(2cz_{0})$. As a result, Eq. (63) simplifies further, yielding a generalized form of Eq. (60) in its final expression:
$\widetilde{{E}}\left(z_{0},\vec{\eta},\vec{l},\vec{\theta}\right)$ $=$ $-\frac{K\omega eLA_{JJ}}{2c^{2}z_{0}\gamma}\exp\left[i\frac{\omega}{c}\left(\frac{z_{0}\theta^{2}}{2}-\vec{\theta}\cdot\vec{l}\right)\right]\mathrm{sinc}\left[\frac{\omega L\left|\vec{\theta}-\vec{\eta}\right|^{2}}{4c}\right]~{}.$ (64)
It is evident from the above that, within the framework of conventional synchrotron radiation theory, considering the radiation from a single electron at a detuning $C$ from resonance, the effect of a kick is merely a rigid rotation of the angular distribution in the direction of the new electron trajectory. This is intuitively reasonable, since after the kick, the electron maintains its velocity and emits radiation in the new, kicked direction due to the Doppler effect. Following this rotation, Eq. (60) transforms into Eq. (64).
19.3.4 Influence of the Kick According to Correct Coupling of Fields and Particles
According to the correct coupling between fields and particles, undulator radiation theory makes a striking prediction regarding radiation from a single electron, both with and without a transverse kick. Specifically, when a kick is introduced, the theory predicts a red shift in the resonance wavelength of the undulator radiation along the direction of the electron’s velocity.
To demonstrate this effect, we consider a covariant treatment that explicitly employs Lorentz transformations. Within this framework, introducing a kick leads to a non-zero red shift in the resonance frequency. This shift arises because, in Lorentz coordinates, the electron’s longitudinal velocity is reduced—from $v$ to $v-v\theta_{k}^{2}/2$while the speed of light remains constant at its electrodynamical value, $c$.
Consequently, the expression given in Eq. (168) requires correction: it should use the modified electron velocity $v-v\theta_{k}^{2}/2$ rather than the unperturbed velocity $v$. The resulting shift in the total phase $\Phi_{T}$, appearing under the integral in Eq. (161), is given by $\Delta\Phi_{T}=\omega\theta_{k}^{2}z^{\prime}/(2c)$, where we have used the ultrarelativistic approximation.
Now, consider the case where the electron moves along a straight, constrained trajectory parallel to the undulator axis, without any kick. The corresponding radiation field in the far zone is described by Eq. (60). Referring back to Eq. (64), we find that the conventional undulator radiation theory provides the following expression for the radiation field after the kick:
${\vec{\widetilde{E}}}=-\frac{K\omega eL}{2c^{2}z_{0}\gamma}A_{JJ}\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\mathrm{sinc}\left[\frac{L}{2}\left(C+\frac{\omega\left|\vec{\theta}-\vec{\theta}_{k}\right|^{2}}{2c}\right)\right]\vec{e}_{x}~{}.$ (65)
The covariant equations indicate that, when the kick is applied, the resulting radiation field is described by the following formula:
${\vec{\widetilde{E}}}=-\frac{K\omega eL}{2c^{2}z_{0}\gamma}A_{JJ}\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\mathrm{sinc}\left[\frac{L}{2}\left(C+\frac{\omega\theta_{k}^{2}}{2c}+\frac{\omega\left|\vec{\theta}-\vec{\theta}_{k}\right|^{2}}{2c}\right)\right]\vec{e}_{x}~{},$ (67)
This formula is nearly, but not exactly, of the same form as Eq. (66). The key difference lies in the additional term $\omega\theta_{k}^{2}/(2c)$ appearing in the argument of the $\mathrm{sinc}$ function. Special attention should be paid to the shift in resonance frequency between the undulator radiation setups with and without a kick.
Recalling the definition of the detuning parameter, $C=k_{w}\Delta\omega/\omega_{r}$, we can express the redshift in resonance frequency as $\Delta\omega/\omega_{r}=-\omega_{r}\theta_{k}^{2}/(2k_{w}c)$. Alternatively, this redshift can also be written in terms of the relativistic factor as $\Delta\omega/\omega_{r}=-\gamma^{2}\theta_{k}^{2}/(1+K^{2}/2)$. This reveals a second-order correction in $\theta_{k}^{2}$, which, however, is amplified by the large factor $\bar{\gamma}_{z}^{2}$.
We are now prepared to more generally examine how the field expression is modified by the introduction of a kick. Suppose that, in the absence of a kick, the electron moves along a trajectory that makes an angle $\vec{\eta}$ with respect to the undulator axis. The corresponding field is given by Eq. (64). Let $\vec{\theta}_{k}$ denote the kick angle of the electron relative to its initial trajectory. According to the conventional approach, the expression for the field after the kick becomes:
${\vec{\widetilde{E}}}=-\frac{K\omega eL}{2c^{2}z_{0}\gamma}A_{JJ}\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\mathrm{sinc}\left[\frac{L}{2}\left(C+\frac{\omega\left|\vec{\theta}-\vec{\eta}-\vec{\theta}_{k}\right|^{2}}{2c}\right)\right]\vec{e}_{x}~{}.$ (69)
In contrast, the covariant approach gives
${\vec{\widetilde{E}}}=-\frac{K\omega eL}{2c^{2}z_{0}\gamma}A_{JJ}\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\mathrm{sinc}\left[\frac{L}{2}\left(C+\frac{\omega\theta_{k}^{2}}{2c}+\frac{\omega\left|\vec{\theta}-\vec{\eta}-\vec{\theta}_{k}\right|^{2}}{2c}\right)\right]\vec{e}_{x}~{},$ (71)
This brings us to an intriguing situation. According to conventional theory, the resonance wavelength depends solely on the observation angle relative to the electron’s velocity direction. However, Eq. (70) shows that for any kick angle $\vec{\theta}_{k}$ and for any angle $\vec{\eta}$ between the undulator axis and the initial electron velocity, the radiation emitted along the velocity direction experiences no redshift. This highlights a critical distinction between the conventional and covariant formulations.
In contrast, the result from the covariant approach—Eq. (72)—explicitly depends on the magnitude of the kick angle $\theta_{k}$. In this case, radiation along the velocity direction exhibits a redshift only when the kick angle is nonzero.
We are thus led to an important conclusion: when the electron is accelerated in the lab frame upstream of the undulator, the covariant trajectory retains this information.
19.3.5 Experimental Test of SR Theory in 3rd Generation Light Source
One way to highlight the incompatibility between the standard approach to relativistic electrodynamics, which typically involves Maxwell’s equations, and the particle trajectories derived from non-covariant particle tracking is through a direct laboratory test of synchrotron radiation theory. Essentially, we are asserting that, despite extensive measurements accumulated over several decades, synchrotron radiation theory has yet to be definitively confirmed by experimental evidence.
Let us now explore the potential of synchrotron radiation sources to validate the predictions of a revised synchrotron radiation theory. In modern synchrotron radiation sources, the electron beam emittance is sufficiently small that one can disregard the effects of finite beam size and angular divergence, particularly in the soft X-ray wavelength range. This makes it feasible to model the synchrotron radiation source using the approximation of a filament electron beam, thereby enabling the use of analytical models for single electron synchrotron radiation fields.
Figure 56: Illustration of the proposed experimental setup to test synchrotron radiation theory using a third-generation light source. Top: Electron beam without a kick. Bottom: Electron beam kicked by an angle $\eta$. In both configurations, the X-ray pulse is filtered through a monochromator, and the total energy is recorded by a detector as a function of undulator detuning.
The basic setup for the test experiment is illustrated in Fig. 56. The soft X-ray undulator beamline should be tuned to a minimum photon energy, typically corresponding to the ”water window” wavelength range. The radiation pulse passes through a monochromator filter $F$, and its energy is then measured by the detector. Precise monochromatization of the undulator radiation is not required in this case; a monochromator line width of $\Delta\omega/\omega\simeq 10^{-3}$ is sufficient.
To conduct the proposed test experiment, it is essential to control the beam kick, for example, using a corrector magnet. Without the kick, the maximum pulse energy detected will align with the monochromator line tuned to resonance. However, when the kick is applied, conventional synchrotron radiation theory predicts no redshift in the resonance wavelength. In contrast, one of the immediate consequences of the corrected theory is the occurrence of a non-zero redshift.
The proposed experimental procedure is relatively simple, as it involves relative measurements in the electron beam’s velocity direction, both with and without the transverse kick. This measurement is crucial because conventional theory predicts the absence of redshift, a result that, to our knowledge, has never been observed at synchrotron radiation facilities. However, an XFEL-based experiment confirms our correction for spontaneous undulator emission—further details can be found in the next chapter.
19.4 Synchrotron Radiation from Bending Magnets
Consider a relativistic electron moving along a circular orbit. In the standard treatment, the observer is located in a vertical plane tangent to the electron’s circular trajectory at the origin, positioned at an angle $\theta$ above the orbital plane. In this configuration, the $z$-axis is not fixed but varies with the observer’s position. Notably, the electron’s motion exhibits cylindrical symmetry about the vertical axis through the center of the orbit. Owing to this symmetry, it is sufficient—when calculating the spectral and angular distribution of emitted photons—to consider an observer in this specific geometric arrangement, without resorting to a more general configuration. It is important to emphasize the distinction between the general geometric setup used here and the more restrictive arrangement often assumed in standard synchrotron radiation treatments of bending magnet radiation.
To proceed, we begin with a broad overview of the main results. The goal is to solve the electrodynamics problem using Maxwell’s equations in their conventional form. Since the particle undergoes relativistic acceleration, its dynamics must be analyzed within the framework of special relativity. However, defining Lorentz coordinates for the laboratory frame becomes nontrivial in the presence of acceleration. The only consistent method to introduce Lorentz coordinates in this context is to employ individual coordinate systems—so-called ”ruler-clock” structures—at each point along the electron’s trajectory.
Several notable effects emerge from the cylindrical symmetry and the use of the paraxial approximation. In our analysis, we demonstrate that the derivation of bending magnet radiation does not require covariant particle tracking. Nonetheless, there is one key instance in which conventional theory fails: the covariant approach predicts a nonzero redshift of the critical frequency when perturbations influence the electron’s motion along the magnetic field—that is, when the trajectory deviates from the nominal orbit.
19.4.1 Conventional Theory
Consider a single relativistic electron moving along a circular orbit and an observer. In conventional treatments, the horizontal observation angle $\theta_{x}$ is typically assumed to be zero. This simplification arises because most textbooks focus on calculating the intensity radiated by a single electron in the far field—specifically, the square modulus of the field amplitude—without addressing more complex scenarios such as source imaging.
Equation (43) can be used to calculate the far-zone radiation field emitted by a relativistic electron moving along a circular arc. Assuming a geometry with a fixed $z$-coordinate, the transverse position of the electron can be expressed as a function of the curvilinear abscissa $s$ as follows:
$\vec{r}(s)=-R\left(1-\cos(s/R)\right)\vec{e_{x}}$ (73)
and
$z(s)=R\sin(s/R)$ (74)
where $R$ is the bending radius.
Since the integral in Eq. (43) is performed along $z$ we should invert $z(s)$ in Eq. (74) and find the explicit dependence $s(z)$:
$s(z)=R\arcsin(z/R)\simeq z+{z^{3}\over{6R^{2}}}$ (75)
so that
$\vec{r}(z)=-{z^{2}\over{2R}}\vec{e_{x}}~{},$ (76)
where the expansion in Eq. (75) and Eq. (76) is justified, once again, in the framework of the paraxial approximation.
With Eq. (43) we obtain the radiation field amplitude in the far zone:
[…]
where
$\Phi_{T}=\omega\left[\left({\theta_{x}^{2}+\theta_{y}^{2}\over{2c}}z_{0}\right)+\left({1\over{2\gamma^{2}c}}+{\theta_{x}^{2}+\theta_{y}^{2}\over{2c}}\right)z^{\prime}\right.$ (78)
$\left.+\left({\theta_{x}\over{2Rc}}\right)z^{\prime 2}+\left(1\over{6R^{2}c}\right)z^{\prime 3}\right]~{}.$ (79)
One can easily reorganize the terms in Eq. (79) to obtain
$\Phi_{T}=\omega\left[\left({\theta_{x}^{2}+\theta_{y}^{2}\over{2c}}z_{0}\right)-{R\theta_{x}\over{2c}}\left({1\over{\gamma^{2}}}+{\theta_{x}^{2}\over{3}}+\theta_{y}^{2}\right)\right.$ (80)
$\left.+\left({{1\over{\gamma^{2}}}+\theta_{y}^{2}}\right){\left(z^{\prime}+R\theta_{x}\right)\over{2c}}+{\left(z^{\prime}+R\theta_{x}\right)^{3}\over{6R^{2}c}}\right]~{}.$ (81)
With redefinition of $z^{\prime}$ as $z^{\prime}+R\theta_{x}$ under integral we obtain the final result:
[…]
×$\times\exp\left\{{i\omega\left[{z^{\prime}\over{2\gamma^{2}c}}\left(1+\gamma^{2}\theta_{y}^{2}\right)+{z^{\prime 3}\over{6R^{2}c}}\right]}\right\}~{},$ (83)
where
$\Phi_{s}={\omega z_{0}\over{2c}}\left(\theta_{x}^{2}+\theta_{y}^{2}\right)$ (84)
and
$\Phi_{0}=-{\omega R\theta_{x}\over{2c}}\left({1\over{\gamma^{2}}}+{\theta_{x}^{2}\over{3}}+\theta_{y}^{2}\right)~{}.$ (85)
In standard treatments of bending magnet radiation, the phase term $\exp(i\Phi_{0})$ is absent. The horizontal observation angle $\theta_{x}$ is always equal to zero.
19.4.2 Why Did the Error in Synchrotron Radiation Remain Undetected so Long?
Our case of interest involves an ultrarelativistic electron undergoing circular acceleration. As previously noted, conventional (non-covariant) particle tracking describes the dynamical evolution in the laboratory frame using the absolute time convention. In this framework, simultaneity is treated as absolute, requiring only a single set of synchronized clocks in the lab frame to describe the electron’s accelerated motion. However, adopting the absolute time convention leads to significantly more complex field equations, which vary with the particle’s velocity—that is, they differ at each point along the electron’s trajectory. This complexity is precisely why the covariant approach is preferred when treating both dynamics and electrodynamics.
We begin by considering an electron moving along a circular trajectory that lies in the $(x,z)$-plane and is tangent to the $z$-axis. Due to the cylindrical symmetry of the problem, it is not necessary to consider a general observation point to calculate the spectral and angular distributions of the emitted photons. Instead, we assume the observer is located in the vertical plane tangent to the trajectory at the origin. Within the ultrarelativistic (paraxial) approximation, we perform transformations accurate up to order $v^{2}_{x}/c^{2}$, which significantly simplifies the calculations. This second-order approximation allows us to neglect higher-order terms while retaining essential relativistic effects.
In the lab frame, manipulations involving the rule-clock structure can be interpreted as a change of variables governed by the transformation in Eq. (13): $x_{L}=\gamma_{x}x$, $t_{L}=(t/\gamma_{x}+\gamma_{x}xv_{x}/c^{2})$, where $\gamma_{x}=1+v_{x}^{2}/(2c^{2})$ reflects the second-order approximation.
This combination of a Galilean transformation and a change of variables effectively results in a transverse Lorentz transformation. Since this modified Galilean approach is mathematically equivalent to a Lorentz transformation, applying these variable changes naturally leads to the correct form of Maxwell’s equations.
To retain Lorentz coordinates in the laboratory frame, as previously discussed, it suffices to apply a time transformation given by $\Delta t=t_{L}-t=-[v_{x}^{2}/(2c^{2})]t+xv_{x}/c^{2}$. At this order of expansion, the relativistic correction to the particle’s offset $x$ does not appear; such corrections arise only at order $v_{x}^{3}/c^{3}$. Therefore, for our case of interest, $x_{L}=x$.
Although this time transformation was initially derived for a specific scenario, the result generalizes to any transverse velocity direction. In general, we can write: $\Delta t=t_{L}-t=-[|\vec{v}_{\perp}|^{2}/(2c^{2})]z^{\prime}/c+\vec{r}_{\perp}\cdot\vec{v}_{\perp}/c^{2}$. To complete our analysis, we now consider the relativistic correction to longitudinal motion. We emphasize again that when evaluating the transformations up to second order in $(v_{\perp}/c)^{2}$, no correction to the longitudinal motion appears at this level of approximation.
In summary, we have demonstrated a covariant approach applicable to arbitrary trajectories—providing a general framework for analyzing the system directly in the space-frequency domain under the paraxial approximation.
Let us now explore how to apply the covariant method to a specific case. We will use our understanding of the relativistically correct approach for calculating synchrotron radiation emission to determine the photon angular-spectral density distributions from a bending magnet.
In the ultrarelativistic limit, the electron undergoes uniform transverse acceleration, given by $a=v^{2}/R\simeq c^{2}/R$. Under this approximation, the electron’s velocity and displacement in the transverse direction can be expressed as: $v_{x}=at=az^{\prime}/v=az^{\prime}/c$, $x=at^{2}/2=az^{\prime 2}/(2c^{2})$. With these expressions, we now have all the quantities needed to evaluate the relativistic time shift: $\Delta t=t_{L}-t=-a^{2}z^{\prime 3}/(2c^{5})+a^{2}z^{\prime 3}/(2c^{5})=0$. There is no time difference! This means that covariant particle tracking is not required to derive the radiation from a bending magnet.
Such a precise cancellation often hints at a deeper underlying principle. However, in this case, it appears to be a mere coincidence—there is no evident profound significance.
This cancellation is not surprising when one analyzes the general expression for the radiation field from a bending magnet in the far zone, given by Eq. (83).
In our previous discussion of undulator radiation, we learned that the relativistic correction appears only when the transverse electron trajectory is included in the total phase $\Phi_{T}$ under the integral Eq.(43). Referring back to Eq.(45) for the phase factor $\Phi_{T}$, we see that the term which depends on the transverse position of the electron can be written as $\exp i(\omega/c)[\theta_{x}x(z^{\prime})+\theta_{y}y(z^{\prime})]$. We conclude that the observation angle $\theta_{x}$ in the total phase factor under the integral must be related to the contribution of the transverse electron trajectory.
Now, examining Eq. (83), we observe that the phase factor includes only the $\theta_{y}$ component of the observation angle. This implies that the transverse constrained motion of the electron, $x(z^{\prime})$ in the bending magnet does not contribute to synchrotron radiation. Therefore, it is justified to use a non-covariant approach when analyzing the constrained electron motion along the nominal orbit in the $(x,z)$-plane.
We emphasize that the cancellation of the relativistic time shift and the independence of the Fraunhofer propagator (more precisely, the paraxial approximation of the Green’s function for the inhomogeneous Helmholtz equation in the space-frequency domain) with respect to the observation angle $\theta_{x}$ in the far zone are two aspects of the same phenomenon. Both arise from the cylindrical symmetry inherent in the motion of an electron along a circular arc.
Due to this symmetry, calculating the spectral and angular photon distributions in the far field does not require placing the observer at a general position. Instead, it suffices to consider an observer located in the vertical plane tangent to the circular trajectory at the origin. In this configuration, the observation angle $\theta_{x}=0$, while $\theta_{y}$ corresponds to elevation above the orbital plane. In other words, within this specific geometry, the $z$-axis is effectively defined by the observer’s position.
However, this approach offers limited insight into the near-field radiation, where the Fresnel propagator must be used. In the near zone, the radiation field is sensitive to the detailed, constrained motion of the electron. Far-field arguments, though, are useful in demonstrating that synchrotron radiation from bending magnets is unaffected by differences between non-covariant and covariant electron trajectories. Te near-zone field can be calculated from the knowledge of the far-zone field. This is made possible through the use of the space-frequency domain and by exploiting the paraxial approximation [83]. The cancellation of the relativistic time shift ensures consistent results in both the far and near zones, as expected  [^62] It is important to emphasize that the inverse field problem—based solely on far-field data—cannot be solved without invoking the paraxial approximation..
19.4.3 Influence of the Kick According to Conventional Theory
Up to this point, we have examined the case of an electron moving along a circular trajectory confined to the $(x,z)$-plane and tangent to the $z$-axis. The resulting phase difference in the emitted fields is influenced by both the observer’s position and the specifics of the electron’s motion. We now shift focus to the radiation emitted by a single electron traversing a bending magnet, allowing for arbitrary angular deflection and spatial offset relative to the nominal orbit.
An approximate representation of the electron’s path—given by Eqs. (155) and (158) in Appendix A3—can be used to characterize the emitted field for a general trajectory. By applying Eqs. (75) and (76), we derive an approximate expression for $s(z)$:
$s(z)=z+{z^{3}\over{6R^{2}}}+{z^{2}\eta_{x}\over{2R}}+{z\eta_{x}^{2}\over{2}}+{z\eta_{y}^{2}\over{2}}$ (86)
so that
$\vec{v}_{\perp}(z)=\left(-{vz\over{R}}+v\eta_{x}\right)\vec{e_{x}}+\left(v\eta_{y}\right)\vec{e_{y}}~{}$ (87)
and
$\vec{r}(z)=\left(-{z^{2}\over{2R}}+\eta_{x}z+l_{x}\right)\vec{e_{x}}+\left(\eta_{y}z+l_{y}\right)\vec{e_{y}}~{}.$ (88)
It is clear that the offsets $l_{x}$ and $l_{y}$ are consistently subtracted from $x_{0}$ and $y_{0}$, respectively. This implies that a shift in the particle trajectory on the vertical plane is equivalent to an opposite shift of the observer. Taking this into account, we define the modified angles $\bar{\theta}_{x}=\theta_{x}-l_{x}/z_{0}$ and $\bar{\theta}_{y}=\theta_{y}-l_{y}/z_{0}$ in order to obtain
[…]
and
$\Phi_{T}=\omega\left({\bar{\theta}_{x}^{2}+\bar{\theta}_{y}^{2}\over{2c}}z_{0}\right)+{\omega\over{2c}}\left({1\over{\gamma^{2}}}+\left(\bar{\theta}_{x}-\eta_{x}\right)^{2}+\left(\bar{\theta}_{y}-\eta_{y}\right)^{2}\right)z^{\prime}$ (90)
$+\left({\omega(\bar{\theta}_{x}-\eta_{x})\over{2Rc}}\right)z^{\prime 2}+\left(\omega\over{6R^{2}c}\right)z^{\prime 3}~{}.$ (91)
One can easily reorganize the terms in Eq. (91) to obtain
$\Phi_{T}=\omega\left({\bar{\theta}_{x}^{2}+\bar{\theta}_{y}^{2}\over{2c}}z_{0}\right)-{\omega R(\bar{\theta}_{x}-\eta_{x})\over{2c}}$ (92)
×$\times\left({1\over{\gamma^{2}}}+(\bar{\theta}_{y}-\eta_{y})^{2}+{(\bar{\theta}_{x}-\eta_{x})^{2}\over{3}}\right)$ (93)
$+\left({{1\over{\gamma^{2}}}+(\bar{\theta}_{y}-\eta_{y})^{2}}\right){\omega\left(z^{\prime}+R(\bar{\theta}_{x}-\eta_{x})\right)\over{2c}}$ (94)
$+{\omega\left(z^{\prime}+R(\bar{\theta}_{x}-\eta_{x})\right)^{3}\over{6R^{2}c}}~{}.$ (95)
Redefinition of $z^{\prime}$ as $z^{\prime}+R(\bar{\theta}_{x}-\eta_{x})$ gives the result
[…]
×$\times\exp\left\{{i\omega\left[{z^{\prime}\over{2\gamma^{2}c}}\left(1+\gamma^{2}(\bar{\theta}_{y}-\eta_{y})^{2}\right)+{z^{\prime 3}\over{6R^{2}c}}\right]}\right\}~{},$ (97)
where
$\Phi_{s}={\omega z_{0}\over{2c}}\left(\bar{\theta}_{x}^{2}+\bar{\theta}_{y}^{2}\right)$ (98)
and
$\Phi_{0}=-{\omega R(\bar{\theta}_{x}-\eta_{x})\over{2c}}\left({1\over{\gamma^{2}}}+(\bar{\theta}_{y}-\eta_{y})^{2}+{(\bar{\theta}_{x}-\eta_{x})^{2}\over{3}}\right)~{}.$ (99)
In the far zone we can neglect terms in $l_{x}/z_{0}$ and $l_{y}/z_{0}$, which leads to
[…]
×$\times\exp\left\{{i\omega\left[{z^{\prime}\over{2\gamma^{2}c}}\left(1+\gamma^{2}\left(\theta_{y}-\eta_{y}\right)^{2}\right)+{z^{\prime 3}\over{6R^{2}c}}\right]}\right\}~{},$ (101)
where
$\Phi_{s}={\omega z_{0}\over{2c}}\left(\theta_{x}^{2}+\theta_{y}^{2}\right)$ (102)
and
$\Phi_{o}\simeq-{\omega R({\theta_{x}}-\eta_{x})\over{2c}}\left({1\over{\gamma^{2}}}+(\theta_{y}-\eta_{y})^{2}+{(\theta_{x}-\eta_{x})^{2}\over{3}}\right)-{\omega\over{c}}(l_{x}\theta_{x}+l_{y}\theta_{y})~{}.$ (103)
It is evident from the discussion above that the field distribution in the far zone depends solely on the observation angle relative to the direction of the electron’s velocity.
According to the conventional (though incorrect) approach to coupling fields and particles, radiation theory predicts the behavior of bending magnet radiation from a single electron, both with and without an applied kick. Specifically, when a kick is introduced, the angular distribution in the far zone is predicted to undergo a rigid rotation.
19.4.4 Influence of the Kick According to Correct Coupling of Fields and Particles
Let us now discuss the covariant treatment, which explicitly employs Lorentz transformations. Consider the radiation from a bending magnet, emitted by a single electron that receives a transverse kick with respect to the nominal orbit in the $(x,z)$-plane. Due to the kick, there is an additional translation along the $y$-axis with a constant velocity $v_{y}=v\theta_{k}$. The corresponding offset of the electron can be written as: $y=\theta_{k}z^{\prime}$. We now incorporate the velocity and offset into the relativistic time shift: $\Delta t=t_{L}-t=-\theta_{k}^{2}z^{\prime}/(2c)+\theta_{k}^{2}z^{\prime}/c=\theta_{k}^{2}z^{\prime}/(2c)$. Thus, the resulting shift in the total phase accumulated along the path is given by: $\Delta\Phi_{T}=\omega\theta_{k}^{2}z^{\prime}/(2c)$. This result is consistent with our redshift calculation in the undulator case when a kick is introduced, as expected.
We would like to make a historical note: the distinction between covariant and non-covariant particle trajectories was not fully understood in the past. As a result, accelerator physicists did not recognize that relativistic kinematic effects contribute to synchrotron radiation.
This raises an important question: how can storage rings function effectively under such assumptions?
The key point is that in this context, the dynamics of the electron beam are primarily governed by synchrotron radiation emitted from bending magnets. Due to the cylindrical symmetry of the system, both covariant and non-covariant descriptions of electron motion along a circular arc produce similar synchrotron radiation characteristics—with one important exception.
The covariant framework predicts a small redshift in the critical frequency of the emitted radiation when the electron motion experiences perturbations in the vertical direction. However, since synchrotron radiation from bending magnets spans a broad frequency range, the overall output intensity is largely insensitive to this redshift.
19.5 Problem-Solving: Multiple Trajectory Kicks
We now consider the case of $n$ arbitrarily spaced kickers, each introducing a distinct rotation angle. Our goal is to explore how covariant particle tracking can be applied in this scenario and to gain insight into the resulting dynamics—particularly when an undulator is placed downstream of the kicker configuration.
Formally, calculating the radiation from the undulator requires accounting for all trajectory perturbations experienced by the electron since its generation. This may seem surprising at first, but it aligns with the general expression for the radiation field from a single electron, as given by Eq. (LABEL:revwied).
It is important to note that, in principle, the entire history of the electron - from $t^{\prime}=-\infty$ to $t^{\prime}=\infty$must be considered, since the integral in Eq. (LABEL:revwied) extends over this range. However, the physical interpretation of this statement depends on the specific scenario. In practice, the integration should be restricted to the time intervals during which the electron contributes significantly to the radiation field.
Ultimately, it is electrodynamics itself that determines which portions of the particle’s trajectory are relevant for calculating the undulator radiation and which may be neglected. The most crucial general statement in this context is that the trajectory must be computed using covariant methods—provided one wishes to employ the standard formulation of Maxwell’s equations.
Let us consider the ultrarelativistic assumption $1/\gamma^{2}\ll 1$, which is commonly valid in typical synchrotron radiation setups. In general, introducing a small parameter into a physical theory leads to significant simplifications. Specifically, the ultrarelativistic approximation corresponds to a paraxial regime, under which Eq. (LABEL:revwied) simplifies to Eq. (43)
Let us now consider a scenario in which the deflection angle of the first bending magnet upstream of the undulator is much larger than $1/\gamma$. In other words, we examine a standard synchrotron radiation setup, where an electron enters the system through a bending magnet, traverses a straight section, passes through an undulator, continues along another straight section, and exits via a second bending magnet.
Although the integration in Eq.(43) is performed from $\infty$- ∞ to $\infty$, the only (edge) part of the trajectory into the bending magnets contributing to the integral is of the order of the radiation formation length $L_{f}$.
Mathematically, this is reflected in the behavior of $\Phi_{T}(z^{\prime})$ in Eq.(43), which exhibits increasingly rapid oscillations as $z^{\prime}$ exceeds the formation length. At the critical wavelength, the formation length is approximately $R/\gamma$, where $R$ is the bending radius. This corresponds to an angular interval along the orbit of roughly $\Delta\theta\simeq 1/\gamma$.
Typically, the critical wavelength of the radiation from a bending magnet in a synchrotron radiation source is about 0.1 nm. The formation length in this case is only a few millimeters.
Note that for ultrarelativistic systems, the formation length is generally much longer than the radiation wavelength. This seemingly counterintuitive result arises because, in ultrarelativistic systems, it is not possible to localize radiation sources within a macroscopic portion of the trajectory.
The formation length can be interpreted as the longitudinal size of a single electron source. It is meaningless to specify the exact location where electromagnetic signals are emitted within the formation length. As a result, in the context of the radiative process within a bending magnet, we cannot distinguish between radiation emitted at point $A$ and radiation emitted at point $B$ if the distance between these two points is shorter than the formation length $L_{f}$.
[…]
Intuitively, bending magnets can be seen as ”switchers” for the ultrarelativistic electron trajectory. While we focus on bending magnets here, other configurations could also serve as switchers, provided they share a common characteristic: the switching process must depend exponentially on the distance from the starting point. In this context, a characteristic length $d_{s}$ can be associated with any switcher.
[…]
Figure 57: Standard undulator radiation setup. As the electron traverses a bending magnet, it emits synchrotron radiation, effectively shedding the Frank-Ginzburg fields associated with its relativistic motion. At the exit of the bending magnet, the electron is effectively “naked,” having lost its field dressing. A new field configuration forms within a characteristic *formation length* at the beginning of the straight section downstream. For third-generation synchrotron radiation sources, the transverse size of this field-dressed ultrarelativistic electron is typically on the order of a few microns..
[…]
The electron’s trajectory can be divided into two fundamentally distinct segments: before and after the switcher. When the electron is accelerated upstream of the switcher in the lab frame, information about this acceleration is embedded in the first segment of its covariant trajectory. However, this pre-acceleration history—along with the associated electromagnetic fields of the ultrarelativistic electron—is effectively ”washed out” during the switching process. As a result, the electron enters the straight section in a ”naked” state, devoid of its previous field configuration.
We begin by describing the field formation process along the straight section downstream of the switcher, using a covariant framework. The first step is to synchronize distant clocks within the lab frame, employing the standard Einstein synchronization procedure. Within this Lorentzian lab frame, we assume the electron follows a rectilinear trajectory at constant velocity $v$; this serves as our initial condition. From here, the evolution of the electromagnetic field can be analyzed using Maxwell’s equations.
When one analyzes the process of ”field-dressed” electron formation from the viewpoint of the noncovariant approach, one assumes the same initial conditions (rectilinear trajectory with velocity $v$) for the electron motion. Then one solves the electrodynamics problem of field formation by using the usual Maxwell’s equations. We already mentioned that the type of clock synchronization that results in the time coordinate $t$ in an electron trajectory $\vec{x}(t)$ is never discussed in accelerator physics. However, we know that the usual Maxwell’s equations are only valid in the Lorentz frame. The noncovariant approach is obviously based on a definite synchronization assumption, but this is actually a hidden assumption. In our case of interest, within the lab frame, the Lorentz coordinates are then automatically enforced. So one should not be surprised to find that in this simple case of rectilinear motion, there is no difference between covariant and noncovariant calculations of the initial conditions at the undulator entrance.
Due to the nature of undulator radiation, the radiation field within the central cone can be accurately calculated using the instantaneous (dipole-like) approximation. This justifies the use of a noncovariant approach to describe the constrained electron motion through the undulator.
In conclusion, for a standard synchrotron radiation setup, it does not matter whether one adopts the covariant approach with Einstein synchronization or the noncovariant approach based on absolute time. Both yield the same results for the radiation field within the central cone.
Let us now examine the behavior of a weak dipole magnet (commonly referred to as a kicker) installed in the straight section upstream of the undulator. This kicker is characterized by a small kick angle such that $(\gamma\theta_{k})^{2}\ll 1$. What should we expect in terms of undulator radiation under these conditions?
At first glance, the setup appears similar to the switcher configuration: the electron trajectory is again divided into two segments—before and after the kicker. However, there is a crucial difference. In this case, electrodynamics dictates that both segments of the trajectory must be considered when calculating the resulting undulator radiation.
As the electron passes through the kicker, there is no significant synchrotron radiation. More precisely, any emitted radiation is indistinguishable from the electron’s self-fields. As a result, the virtual radiation fields are not washed out as they are in the switcher scenario. The electron emerges from the kicker still ”field-dressed,” but with a perturbed electromagnetic field that now carries information about the acceleration experienced relative to an inertial frame.
According to conventional theory—rooted in Newtonian kinematics—the Galilean vector law of velocity addition is applied. Within this non-covariant framework, the electron’s direction changes after the kick, but its speed remains constant. In contrast, covariant particle tracking—based on Lorentz coordinates—leads to a different prediction: the electron’s speed is slightly reduced, from $v$ to $v(1-\theta_{k}^{2}/2)$. This discrepancy arises because the covariant approach properly accounts for the relativistic addition of non-parallel velocities.
In standard electrodynamics, the usual algorithm involves solving Maxwell’s equations using particle trajectories derived from non-covariant tracking. According to this method, the undulator radiation emitted along the post-kick velocity direction exhibits no redshift in resonance frequency, regardless of the kick angle $\theta_{k}$.
However, when the coupling between fields and particles is treated correctly within the covariant framework, a notable prediction emerges. Specifically, synchrotron radiation theory indicates a redshift in the resonance frequency of the undulator radiation in the kicked direction. This redshift is given by: $\Delta\omega_{r}/\omega_{r}=-\gamma^{2}\theta_{k}^{2}/(1+K^{2}/2)$. This result underscores the importance of using covariant dynamics in accurately modeling radiation phenomena in relativistic beamlines.
19.6 Helical Trajectories and Synchrotron Radiation
The observation of a redshift in bending magnet radiation inherently suggests a similar issue within the conventional theory of cyclotron radiation. In the ultrarelativistic regime, well-established analytical expressions describe the spectral and angular distribution of radiation emitted by an electron moving in a uniform magnetic field, where the motion has a non-relativistic component parallel to the field and an ultrarelativistic component perpendicular to it.
According to the conventional framework—just as in the case of bending magnet radiation—the angular and spectral distribution of the emitted radiation depends on the total velocity of the particle, a consequence of the Doppler effect. In contrast, the covariant approach predicts a redshift of the critical frequency that arises specifically when there are perturbations in the electron’s motion along the direction of the magnetic field.
It is important to note that cyclotron-synchrotron radiation is a fundamental emission process in both plasma physics and astrophysics. Therefore, the corrections we propose have significant implications extending well beyond synchrotron radiation facilities.
19.6.1 Existing Theory
Let us now examine the relativistic cyclotron radiation in greater detail. Here, we present only the final results and discuss their connection to the conventional synchrotron radiation theory associated with bending magnets. In the case of uniform translational motion with non-relativistic velocity along the direction of the magnetic field (see Fig. 58), a widely accepted expression in astrophysics describes the angular and spectral distributions of radiation emitted by an ultra-relativistic electron following a helical trajectory [^63] The angular and spectral distributions of radiation from an ultra-relativistic electron on a helical orbit were derived in [73, 74]. These results are now standard examples in the literature (see, e.g., [75]) and are not elaborated upon here.
${\vec{\widetilde{E}}}(\chi,\alpha)$$\sim\Bigg{\{}\vec{e}_{x}\left[(\xi^{2}+\psi^{2})K_{2/3}\left(\frac{\omega}{2\omega_{c}}\left(1+\frac{\psi^{2}}{\xi^{2}}\right)^{3/2}\right)\right]$ (105)
$-i\vec{e}_{y}\left[(\xi^{2}+\psi^{2})^{1/2}\psi K_{1/3}\left(\frac{\omega}{2\omega_{c}}\left(1+\frac{\psi^{2}}{\xi^{2}}\right)^{3/2}\right)\right]\Bigg{\}}~{},$
where $K_{1/3}$ and $K_{2/3}$ are the modified Bessel functions, $\xi=1/\gamma$, $\psi=\chi-\alpha$, ($\chi$ is the angle between $\vec{v}$ and $\vec{B}$ and $\alpha$ that between $\vec{n}$ and $\vec{B}$); the angle $\psi$ is the angular distance between the direction of the electron velocity $\vec{v}$ and the direction of observation $\vec{n}$. Here the $\omega_{c}$ is defined by $3eB\gamma^{2}/(2mc)$.
Figure 58: Geometry for radiation production from helical motion.
We have already discussed the radiation emitted by an ultrarelativistic electron following a helical trajectory in the previous section. Equation (101) presents the result we derived earlier for synchrotron radiation from a single electron undergoing an angular deflection relative to the nominal orbit. At first glance, Eq. (105) appears different from Eq. (101). However, the two become equivalent upon introducing the small deflection angle $\eta_{y}=\pi/2-\chi$ and the observation angle $\theta_{y}=\pi/2-\alpha$, under the assumption that the observer lies in the vertical plane tangent to the trajectory (i.e., $\theta_{x}=\eta_{x}=0$). The integrals in Eq. (101) can then be expressed in terms of modified Bessel functions:
[…]
Then, making the necessary variable changes, the formula reduces to Eq.( 105).
19.6.2 Methodology of Solving Problems Involving Boosts
The derivation leading to Eq. (105) is rather involved, so it is useful to present an independent and more intuitive argument. A particularly simple way to analyze radiation from ultrarelativistic helical motion leverages special relativity and requires minimal calculation.
In the case of uniform translational motion, radiation analysis can often be simplified by identifying a reference frame in which the problem is already solved—such as the frame where the particle undergoes circular motion—and then transforming the result back to the original frame.
The reference system $S^{\prime}$ in which the electron moves in circular motion can be transformed to a reference system $S$ in which the electron proceeds following a helical trajectory. Eq. ( 105) holds, indeed, in the frame $S$ for a particle whose velocity is $(v_{x},v_{y},v_{z})=(v_{0}\sin\chi\sin\phi,v_{0}\cos\chi,v_{0}\sin\chi\cos\phi)$. The Lorentz transformation, which leads to the value $v_{y}=v_{0}\cos\chi$ for the $y$-component of the velocity yields $(v_{x},v_{y},v_{z})=(v^{\prime}\sin\phi^{\prime}/\gamma_{y},v_{y},v^{\prime}\cos\phi^{\prime}/\gamma_{y})$, where $\gamma_{y}=1/\sqrt{1-v^{2}_{y}/c^{2}}$, $v^{\prime}$ is the velocity of the electron in the frame $S^{\prime}$ and the phase angle $\phi^{\prime}=\phi$ is invariant. This means that, in order to end up in $S$ with a transverse (to the magnetic field direction) velocity $v_{\perp}=v_{0}\sin\chi$, one must start in $S^{\prime}$ with $v^{\prime}=\gamma_{y}v_{0}\sin\chi$. In the ultrarelativistic approximation $\gamma_{\perp}^{2}=1/(1-v^{2}_{\perp}/c^{2})\gg 1$, and one finds the simple result $v_{0}=v^{\prime}$, so that a Lorentz boost with non-relativistic velocity $v_{y}$ leads to a rotation of the particle velocity $\vec{v}_{0}$ of the angle $\eta=\pi/2-\chi\simeq v_{y}/c\ll 1$ (if angle $\eta$ is small and $v_{0}\simeq c$, we would write $\gamma_{y}\sin\chi\simeq 1$). If one transforms the radiation field for a particle in a circular motion in the system $S^{\prime}$, one obtains the result that the effect of a boost amounts to a rigid rotation of the angular-spectral distribution of the radiation emitted by the electron moving with velocity $v_{0}$ on a circle that is, once more, Eq. ( 105)  [^64] A covariant approach to analyzing radiation from helical motion is presented in [76]. It is generally assumed that $\vec{x}(t)=\vec{x}(t)_{cov}$, which is why [76] makes no distinction between the covariant and non-covariant treatments of electron motion along a helix downstream of the kicker setup..
Note: the subscript ”” used here to indicate the velocity in the transverse to the magnetic field direction should not be confused with the ”” referring to an acceleration of the electron in the transverse direction in the proceeding sections.
It follows quite naturally that the covariant analysis of radiation from helical motion, as considered above, is grounded in the use of Lorentz transformations. In other words, within the laboratory frame, Lorentz coordinates are inherently applied. It is assumed that, in this Lorentz lab frame, the electron follows a helical trajectory with velocity $v_{0}$, which serves as an initial condition. In the ultrarelativistic approximation, applying a Lorentz boost along the direction of the magnetic field with a non-relativistic velocity $v_{y}$ transforms the motion into a circular trajectory, still with velocity $v_{0}$. As a result, a single boost along the field direction does not alter the radiation properties.
Let us consider synchrotron radiation from a single electron that has received a transverse kick relative to the nominal orbit in the $(x,z)$-plane. As a result of this kick, the electron experiences an acceleration that induces a velocity component $v_{y}=v\theta_{k}$ along the $y$-axis and a corresponding change $\Delta v_{z}=-v\theta_{k}^{2}/2$ along the $z$-axis. Limiting the analysis to second-order terms simplifies the calculations significantly.
To describe the motion of the particle downstream of the kicker, we can employ a sequence of two commuting, non-collinear Lorentz boosts.
When the kick is applied, covariant particle tracking predicts a modification to the initial conditions at the entrance of the synchrotron radiation setup. If we retain the Lorentz coordinate system of the lab frame downstream of the kicker, we find that the covariant velocity of the particle on the resulting helical orbit decreases from $v$ to $v_{0}=v-v\theta_{k}^{2}/2$. Applying a covariant analysis of radiation for a helical trajectory with this reduced velocity leads to a redshift in the critical wavelength.
It is important to note that the single passive Lorentz boost to the reference frame $S^{\prime}$, discussed above, is merely a mathematical device. One can interpret this transformation as a change of variables, rather than a physical alteration of the system. Under this kinematic transformation, neither the electron’s nor the observer’s motion relative to the fixed stars is affected. True (i.e., physically real) acceleration is a dynamic process and cannot be captured by coordinate transformations alone.
Therefore, we conclude that accelerating an electron with respect to the fixed stars in the lab inertial frame—before it enters the uniform magnetic field—is an absolute acceleration. The effects of this acceleration are intrinsically encoded in the covariant particle trajectory.
19.6.3 On the Advanced ”Paradox” Related to the Coupling Fields and Particles
We now wish to highlight that two distinct sets of initial conditions can result in the same uniform translation along the magnetic field direction in the Lorentz lab frame.
To illustrate this, we first consider an electron moving along a circular trajectory in the $(x,z)$-plane. Next, we rotate the magnetic field vector $\vec{B}$ within the $(y,z)$-plane by a small angle $\theta_{0}$, under the assumption that $(\theta_{0}\gamma)^{2}\ll 1$. In this configuration, the electron moves uniformly along the direction of the magnetic field with a velocity component $v\theta_{0}$. Importantly, this rotation of the magnetic field does not affect the radiation characteristics of the electron in circular motion. This invariance is intuitive: after rotating the bending magnet, the electron retains its velocity, and due to the Doppler effect, it continues to emit radiation along the direction of motion. The change in the curvature radius $R$ due to the rotation is of order $\theta_{0}^{2}$, and thus can be neglected.
We now consider a second scenario. Suppose a kicker is installed in the straight section upstream of the bending magnet, imparting a kick in the $y$-direction with angle $\theta_{k}=\theta_{0}$. When such a kick is applied, a redshift in the critical wavelength of the emitted radiation occurs. This is a consequence of the relativistic velocity addition law: the longitudinal component of the electron’s velocity is reduced from $v$ to $v-v\theta_{0}^{2}/2$ after the kick. The relative redshift of the critical frequency $\omega_{c}$ is given by: $\Delta\omega_{c}/\omega_{c}=-(3/2)\gamma^{2}\theta_{0}^{2}$. Here we observe a second-order correction in $\theta_{0}$, but it is enhanced by the large factor $\gamma^{2}$.
In the covariant framework, the outcome depends on the absolute magnitude of the kick angle $\theta_{0}$. Radiation emitted along the velocity direction experiences a redshift only when the kick angle is nonzero. This result reflects the concept of ”absolute” acceleration—acceleration relative to the fixed stars.
The distinction between these two situations—both ultimately resulting in uniform translation along the direction of the magnetic field—is quite intriguing. This difference arises due to the presence of two distinct Lorentz coordinate systems within the lab frame. When we accelerate the electron upstream of the bending magnet, we effectively change its associated Lorentz frame.
To maintain a consistent Lorentz coordinate system in the lab frame after the electron receives a transverse kick, a clock resynchronization must be performed. Consequently, we should expect the electron’s velocity to be modified. This explains the difference between the two experimental setups: when the electron’s motion upstream of the bending magnet remains unperturbed (relative to the fixed stars), no clock resynchronization is necessary. However, when we do perturb the electron’s motion, resynchronization is required.
Figure 59: Two sets of initial conditions result in the same uniform motion along the magnetic field direction when using absolute time coordinatization in the lab frame. In both setups, the magnitude of the electron’s velocity and the orientation of the velocity vector relative to the magnetic field vector are identical.
This leads us to consider an apparent paradox. The argument is as follows: under absolute time coordinatization in the lab frame, the initial conditions at the entrance to the bending magnet appear to be identical in both cases. Specifically, the electron’s velocity magnitude and its orientation with respect to the magnetic field are the same. Yet, when we accelerate the electron upstream of the magnet, this information is seemingly absent from its noncovariant trajectory.
So, where is the information about this acceleration encoded? Since the electron is traditionally considered a structureless particle, the situation seems paradoxical.
[…]
When the electron is subjected to a transverse kick—such as by a kicker magnet—its self-fields are perturbed, now reflecting information about the acceleration. Under the traditional (Galilean) kinematics, the orientation of the virtual radiation phase front remains unchanged. However, Maxwell’s equations are not invariant under Galilean transformations. As discussed throughout this book, adopting Galilean kinematics leads to anisotropic field equations.
As a result, although the phase front of the virtual beam remains planar, the direction of propagation is no longer perpendicular to the phase front. That is, the motion of the virtual radiation beam and the normal to its phase front diverge. Therefore, within the absolute time synchronization convention, electrodynamics predicts that the virtual radiation beam propagates in the kicked direction with a phase front tilt $\theta_{k}$.
In this way, the information about the electron’s acceleration is embedded in the perturbation of its self-electromagnetic field.
20 Relativity and X-Ray Free Electron Lasers
20.1 Introductory Remarks
In the previous chapter, we examined why the error in radiation theory remained undetected for so long. Within the covariant framework, relativistic kinematic effects in synchrotron radiation emerge in successive orders of approximation. Rather than using the total velocity parameter $(v/c)$ as in non-relativistic cases, we employ the small transverse velocity parameter $(v_{\perp}/c)$.
In our earlier discussion of bending magnet radiation, we found that the motion of a single ultrarelativistic electron in a constant magnetic field, according to relativistic theory, affects kinematic terms only at the second order, i.e., $(v_{\perp}/c)^{2}$ It has been shown that, due to a combination of the ultrarelativistic (or paraxial) approximation and the specific symmetry of conventional synchrotron radiation setups, second-order relativistic kinematic effects cancel out—except for a non-zero redshift in the critical frequency. This redshift arises when the electron experiences perturbations in the direction of the bending magnetic field. However, because synchrotron radiation from bending magnets spans a broad frequency range, the output intensity is largely insensitive to this redshift. As a result, spontaneous synchrotron radiation shows no observable sensitivity to the differences between covariant and non-covariant particle trajectories.
This situation changes in the 21st century with the advent of X-ray Free Electron Lasers (XFELs). In XFELs, first-order kinematic effects $(v_{\perp}/c)$ play a crucial role in the description of radiation. In such cases, the covariant coupling between fields and particles predicts outcomes that sharply contrast with those of conventional treatments.
In this chapter, we critically reexamine existing XFEL theory, with particular focus on coherent undulator radiation from modulated electron beams. The discussion is primarily intended for readers with limited background in accelerator and XFEL physics.
Figure 60: Illustration of a well-known outcome in conventional (non-covariant) particle tracking: a micro-bunched electron beam passes through a weak dipole magnet (kicker), receiving a kick by an angle $\theta_{k}$. The beam’s propagation axis is deflected, while the orientation of the wavefront remains unchanged.
The conventional theory of X-ray Free Electron Lasers (XFELs) typically relies on a hybrid approach: Newtonian kinematics—corrected for relativistic mass—is applied to particle dynamics, while Einsteinian principles underpin electrodynamics. In practice, the relativistic treatment of particle motion often reduces to a modified form of Newton’s second law, with velocity composition still grounded in Galilean transformations.
For rectilinear motion of a modulated electron beam, both covariant and non-covariant approaches yield identical particle trajectories. Consequently, Maxwell’s equations remain consistent with conventional particle tracking in such cases. However, the relativity of simultaneity—which involves the intermixing of spatial and temporal coordinates—leads to fundamental differences between covariant and non-covariant descriptions when the beam follows a curved path. According to the theory of relativity, these discrepancies emerge specifically in the presence of acceleration along curved trajectories.
First-order relativistic effects in XFELs can arise under various conditions, particularly when the electron beam experiences a trajectory kick [^65] Angular kicks are commonly used in XFEL diagnostics and experiments. For instance, standard gain length measurements rely on such kicks. They are also employed in beam-splitting techniques, where polarization components are separated through angular deflection of the modulated beam [77, 78]..
Among these, one of the most illustrative effects involves the generation of coherent undulator radiation by an ultrarelativistic, modulated electron beam that has been deflected by a weak dipole field prior to entering a downstream undulator. This scenario serves as a critical test for the correct coupling between electromagnetic fields and particle dynamics.
To set the stage, it is helpful to begin with a broad overview of key results. Let us now consider the predictions of the standard XFEL theory in the context of non-collinear electron beam motion. A well-established result from conventional particle tracking states that, following a weak transverse kick, the trajectory of the electron beam changes direction while the modulation wavefront retains its original orientation (see Fig. 60). This results in a misalignment between the electron beam direction and the normal to the modulation wavefront—commonly referred to as a wavefront tilt.
In the conventional framework, this wavefront tilt is treated as a physical phenomenon. That is, a transverse kick does not alter the orientation of the modulation wavefront, thereby reducing the efficiency of radiation in the direction of the beam motion  [^66] In standard XFEL operation, the electron beam is guided to remain closely aligned with the undulator axis. Nonetheless, random perturbations in the focusing system can introduce angular deviations or ”kicks.” The resultant misalignment between beam trajectory and wavefront normal has been discussed in the literature. It is traditionally understood that coherent radiation is emitted along the wavefront normal. Consequently, under the conventional field-particle coupling (which we argue is flawed), the mismatch between these directions diminishes radiation efficiency [79]. Studies of trajectory errors in XFEL amplification have shown that undulator magnetic fields must meet stringent tolerances. Interestingly, these tolerances have proven to be more stringent than necessary according to the corrected, four-dimensional (4D) covariant XFEL theory—an insight that helps explain the remarkable success of XFEL technology in recent decades..
Figure 61: Result of covariant particle tracking. In the ultrarelativistic limit, the modulation wavefront—or equivalently, the plane of simultaneity—is always oriented perpendicular to the electron beam’s velocity when the evolution of the modulated beam is described using Lorentz coordinates. According to the theory of relativity, a modulated electron beam in this limit exhibits the same kinematic behavior, in Lorentz coordinates, as a laser beam. Maxwell’s equations dictate that the wavefront of a laser beam is always orthogonal to its direction of propagation.
The covariant approach, applied within the frameworks of both mechanics and electrodynamics, predicts an effect that stands in stark contrast to conventional theory. Specifically, in the ultrarelativistic limit, the modulation wavefront—defined as a plane of simultaneity—is always perpendicular to the velocity of the electron beam (see Fig. 61). Consequently, Maxwell’s equations predict strong coherent undulator radiation emitted in the direction of the kick imparted to the modulated electron beam.
Experimental results confirm this prediction. Observations from XFEL facilities have demonstrated that even the direction of coherent undulator radiation emission lies outside the predictive scope of the conventional theory (see Section 20.6 for further discussion).
It is worth noting that the lack of a dynamical explanation for the readjustment of the modulation wavefront under Lorentz coordinatization has been a point of concern for some XFEL experts. However, we suggest that a useful way to conceptualize this readjustment is to view it as a consequence of transforming to a new time variable within the Galilean (i.e., single-frame) formulation of electrodynamics.
20.2 Modulation Wavefront Orientation
Let us consider a modulated electron beam propagating along the $z$-axis of a Cartesian coordinate system $(x,y,z)$ in the laboratory frame. For example, assume that the modulation wavefront is perpendicular to the beam velocity $v_{z}$. How can we determine this orientation?
As the electron bunch moves, its position changes over time. A natural way to characterize the orientation of the modulation wavefront is to ask: at what time does each electron cross the $y$-axis of the reference frame?
If a synchronization convention has been adopted—i.e., a method for timing distant events—we can use it to define the orientation of the modulation wavefront. Specifically, if electrons corresponding to the region of maximum density cross the $y$-axis simultaneously at a given position $z$, then the wavefront is perpendicular to the $z$-axis. In other words, the modulation wavefront can be identified as a plane of simultaneous events—where the events are the arrivals of electrons at the point of maximum density. In short, it is a plane of simultaneity.
We now examine the case where the electron beam is accelerated in the lab frame to acquire a small transverse velocity component $v_{y}$ along the $y$-axis. This raises an important question: how should synchronization be defined in the lab frame after this acceleration? Prior to the kicker setup, we chose a Lorentz coordinate system for the lab frame  [^67] Conventional XFEL theory adopts the absolute time convention to describe particle dynamics. Notably, employing the standard coupling of Maxwell’s equations with the corrected Newtonian equations to compute radiation from an ultrarelativistic electron beam in an undulator does not inherently introduce inaccuracies. In the case of straight-line motion—whether for the electron beam or the emitted undulator radiation—both covariant and non-covariant approaches produce identical trajectories. This ensures internal consistency between Maxwell’s equations and conventional particle tracking methods. However, this consistency brings to the forefront a critical question: how should synchronization be defined in the laboratory frame for a modulated electron beam upstream of the kicker setup? This is particularly relevant considering that the absolute time coordinatization was originally adopted in alignment with FEL theory. The most straightforward synchronization method employs the same set of synchronized clocks used in conventional particle tracking. This approach defines simultaneity based on light signals emitted from a stationary dipole source, assuming isotropic light propagation at speed $c$ in all directions.. After the beam acquires a small transverse velocity $v_{y}$, maintaining the original synchronization leads to complications in the electrodynamics of moving charges. As a result of such a boost, the transformation of time and spatial coordinates has the form of a Galilean transformation.
To maintain a Lorentz coordinate system in the lab frame following the acceleration, it is necessary to perform a clock resynchronization. Specifically, this involves a time shift of the form given in Eq. 13: $t\to t+yv_{y}/c^{2}$. This adjustment is valid in the first-order approximation, where $v_{y}/c$ is so small that $v_{y}^{2}/c^{2}$ can be neglected and one arrives at the coordinate transformation $y\to y+v_{y}t$, $t\to t+yv_{y}/c^{2}$. This differs from a pure Galilean transformation by the inclusion of the relativity of simultaneity—the only first-order relativistic correction in $v_{y}/c$.
This resynchronization introduces a time shift for electrons located at different transverse positions. For instance, electrons at a transverse position $y$ with maximum density will cross the lab frame’s $y$-axis at a later time compared to electrons at $y=0$. The resulting time shift is given by $\Delta t=yv_{y}/c^{2}$. This time shift results in an effective rotation of the modulation wavefront by an angle: $v_{z}\Delta t/y=v_{z}v_{y}/c^{2}$ in the first order approximation. In ultrarelativistic limits, $v_{z}\simeq c$, and the modulation wavefront rotates exactly as the velocity vector $\vec{v}$.
What does this readjustment of the wavefront mean in terms of measurements? In the framework of absolute time coordination, simultaneity between two events is considered absolute. This absolute nature of temporal coincidence stems from the convention of absolute time synchronization. In this classical kinematics view, the modulation wavefront remains unchanged. However, within the covariant approach, we adopt a criterion for simultaneity based on the invariance of the speed of light. It becomes evident that, due to the motion of electrons along the $y$-axis (i.e., along the plane of simultaneity prior to the boost), with a velocity $v_{y}$, the simultaneity of events is no longer absolute. Instead, it becomes dependent on the kick angle $\theta=v_{y}/c$. This reasoning mirrors the principles of Einstein’s train-embankment thought experiment.
The orientation of the wavefront lacks an exact objective meaning due to the relativity of simultaneity. However, the statement that the wavefront orientation has an objective meaning within a certain accuracy can be illustrated by considering the wavefront in its proper orientation, with the angle’s uncertainty (or “blurring”) given by $\Delta\theta\simeq v_{z}(v_{y}/c^{2})$. This relationship defines the limits within which nonrelativistic theory remains applicable.
For a very nonrelativistic electron beam, where $v_{z}^{2}/c^{2}$ is negligible, the angle of ”blurring” becomes extremely small. In this case, the wavefront tilt angle $\theta=v_{y}/v_{z}$ is almost perfectly sharp, with $\Delta\theta/\theta\simeq v_{z}^{2}/c^{2}\ll 1$. This represents the limiting case of nonrelativistic kinematics. The angle ”blurring” is a characteristic feature of relativistic beam motion.
In the ultrarelativistic limit the wavefront tilt has no exact objective meaning. This is because, due to the finite speed of light, no experimental method exists by which we could measure this tilt with certainty.
20.3 XFEL Radiation Configuration
One of the most fundamental effects that serve as a crucial test for the proper coupling of fields and particles is the production of coherent undulator radiation by a modulated ultrarelativistic electron beam, which is deflected by a weak dipole field before entering a downstream undulator. Our goal is to study the emission of coherent undulator radiation from such a system.
At the heart of an XFEL source is the undulator, which compels electrons to follow curved, periodic trajectories. There are two primary undulator configurations: helical and planar. To understand the basic principles of undulator operation, let us begin by examining the helical undulator.
The magnetic field along the axis of the helical undulator is given by
$\vec{B}_{w}=\vec{e}_{x}B_{w}\cos(k_{w}z)-\vec{e}_{y}B_{w}\sin(k_{w}z)~{},$ (108)
where $k_{w}=2\pi/\lambda_{w}$ is the undulator wavenumber and $\vec{e}_{x,y}$ are unit vectors directed along the $x$ and $y$ axes. We neglected the transverse variation of the magnetic field. It is necessary to mention that in XFEL engineering we deal with very high-quality undulator systems, which have a sufficiently wide good-field-region, so that our studies, which refer to a simple model of undulator field nevertheless yield a correct quantitative description in a large variety of practical problems. The Lorentz force $\vec{F}=-e\vec{v}\times\vec{B}_{w}/c$ is used to derive the equation of motion of electrons with charge $-e$ and mass $m$ in the presence of magnetic field
$m\gamma\frac{dv_{x}}{dt}=\frac{e}{c}v_{z}B_{y}=-\frac{e}{c}v_{z}B_{w}\sin(k_{w}z)~{},$ (109)
$m\gamma\frac{dv_{y}}{dt}=-\frac{e}{c}v_{z}B_{x}=-\frac{e}{c}v_{z}B_{w}\cos(k_{w}z)~{}.$ (110)
Introducing $\tilde{v}=v_{x}+iv_{y}$, $dz=v_{z}dt$ we obtain
$m\gamma\frac{d\tilde{v}}{dz}=-i\frac{e}{c}(B_{x}+iB_{y})=-i\frac{e}{c}B_{w}\exp(-ik_{w}z)~{}.$ (111)
Integration of the latter equation gives
$\frac{\tilde{v}}{c}=\theta_{w}\exp(-k_{w}z)~{},$ (112)
where $\theta_{w}=K/\gamma$ and $K=eB_{w}/(k_{w}mc^{2})$ is the undulator parameter. The explicit expression for the electron velocity in the field of the helical undulator has the form
$\vec{v}=c\theta_{w}[\vec{e}_{x}\cos(k_{w}z)-\vec{e}_{y}\sin(k_{w}z)]~{},$ (113)
This means that the reference electron in the undulator moves along the constrained helical trajectory parallel to the $z$ axis. As a rule, the electron rotation angle $\theta_{w}$ is small and the longitudinal electron velocity $v_{z}$ is close to the velocity of light, $v_{z}=\sqrt{v^{2}-v^{2}_{\perp}}\simeq v(1-\theta_{w}^{2}/2)\simeq c$.
Let us consider a modulated ultrarelativistic electron beam moving along the $z$-axis within the field of a helical undulator. In this study, we make the following assumptions: First, in the absence of any deflection, the electrons follow constrained helical trajectories that are parallel to the $z$-axis. Second, the electron beam density at the entrance of the undulator is given by
$n=n_{0}(\vec{r}_{\perp})[1+a\cos\omega(z/v_{z}-t)]~{},$ (114)
where $a=\mathrm{const.}$ In other words we consider the case in which there are no variations in amplitude and phase of the density modulation in the transverse plane. Under these assumptions, the transverse current density may be written in the form
$\vec{j}_{\perp}=-e\vec{v}_{\perp}(z)n_{0}(\vec{r}_{\perp})[1+a\cos\omega(z/v_{z}-t)]~{}.$ (115)
Even though the measured quantities are real, it is generally more convenient to use complex representation, starting with real $\vec{j}_{\perp}$, one defines the complex transverse current density:
$j_{x}+ij_{y}=-ec\theta_{w}n_{0}(\vec{r}_{\perp})\exp(-ik_{w}z)[1+a\cos\omega(z/v_{z}-t)]~{}.$ (116)
The transverse current density has an angular frequency $\omega$ and two waves traveling in the same direction with variations $\exp i(\omega z/v_{z}-k_{w}z-\omega t)$ and $\exp-i(\omega z/v_{z}+k_{w}z-\omega t)$ will add to give a total current proportional to $\exp(-ik_{w}z)[1+a\cos\omega(z/v_{z}-t)]$. The factor $\exp i(\omega z/v_{z}-k_{w}z-\omega t)$ indicates a fast wave, while the factor $\exp-i(\omega z/v_{z}+k_{w}z-\omega t)$ indicates a slow wave. The use of the word ”fast” (”slow”) here implies a wave with a phase velocity faster (slower) than the beam velocity.
Having defined the sources, we now turn to the electrodynamics problem. Maxwell’s equations can be manipulated in various ways to derive forms that are more suitable for specific applications. For instance, from Maxwell’s equations (Eq. 8), we can derive an equation that depends solely on the electric field vector $\vec{E}$ (in Gaussian units):
$c^{2}\vec{\nabla}\times(\vec{\nabla}\times{\vec{E}})=-\partial^{2}\vec{E}/\partial t^{2}-4\pi\partial\vec{j}/\partial t~{}.$ (117)
With the help of the identity
×(×$\vec{\nabla}\times(\vec{\nabla}\times{\vec{E}})=\vec{\nabla}(\vec{\nabla}\cdot{\vec{E}})-\nabla^{2}\vec{E}$ (118)
and Poisson equation
$\vec{\nabla}\cdot\vec{E}=4\pi\rho$ (119)
we obtain the inhomogeneous wave equation for $\vec{E}$
$c^{2}\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial t^{2}=4\pi c^{2}\vec{\nabla}\rho+4\pi\partial\vec{j}/\partial t~{}.$ (120)
Once the charge and current densities $\rho$ and $\vec{j}$ are specified as a function of time and position, this equation allows one to calculate the electric field $\vec{E}$ at each point of space and time. This nonhomogeneous wave equation thus serves as the complete and accurate formulation for describing radiation. However, we aim to apply it to a simplified scenario in which the second term on the right-hand side—associated with the current density—dominates the contribution to the radiation field.
It is important to recall that our focus is on coherent undulator radiation, which exhibits a divergence much smaller than the angle $1/\gamma$. Under this condition, it can be shown that the gradient term, $4\pi c^{2}\vec{\nabla}\rho$, on the right-hand side of the wave equation becomes negligible. As a result, we simplify the wave equation to the following form:
$c^{2}\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial t^{2}=4\pi\partial\vec{j}_{\perp}/\partial t~{}.$ (121)
We now consider the case in which the phase velocity of the current wave is close to the speed of light. This condition can be satisfied under the resonance condition $\omega/c=\omega/v_{z}-k_{w}$. This is the condition for synchronism between the transverse electromagnetic wave and the fast transverse current wave with the propagation constant $\omega/v_{z}-k_{w}$. When the phase velocity of the current wave matches that of the electromagnetic wave, a spatial resonance can occur between the electromagnetic field and the electrons. This resonance enables a cumulative interaction between the modulated electron beam and the transverse electromagnetic wave, even in free space.
Under this synchronism condition—and provided the undulator has a sufficiently large number of periods—the contributions of all other (non-synchronous) waves become negligible. Thus, it is justified to focus solely on the resonant wave in analyzing the beam–field interaction.
Here follows an explanation of the resonance condition which is elementary in the sense that we can see what is happening physically. The field of electromagnetic wave has only transverse components, so the energy exchange between the electron and electromagnetic wave is due to a transverse component of the electron velocity. For effective energy exchange between the electron and the wave, the scalar product $-e\vec{v}_{\perp}\cdot\vec{E}$ should be kept nearly constant along the whole undulator length. We see that required synchronism $k_{w}+\omega/c-\omega/v_{z}=0$ takes place when the wave advances the electron beam by the wavelength at one undulator period $\lambda_{w}/v_{z}=\lambda/(c-v_{z})$, where $\lambda=2\pi/\omega$ is the radiation wavelength. This tells us that the angle between the transverse velocity of the particle $\vec{v}_{\perp}$ and the vector of the electric field $\vec{E}$ remains nearly constant. Since $v_{z}\simeq c$ this resonance condition may be written as $\lambda\simeq\lambda_{w}/(2\gamma_{z}^{2})=\lambda_{w}(1+K^{2})/(2\gamma^{2})$.
We will employ an adiabatic approximation, which is applicable in all practical XFEL scenarios where the modulation wavelength is much shorter than the electron bunch length $\sigma_{b}$, i.e. $\sigma_{b}\omega/c\gg 1$. Since our focus is on coherent emission near the modulation wavelength, the theory of coherent undulator radiation is most naturally formulated in the space-frequency domain. This is because, in such cases, one is typically interested in the radiation properties at a fixed modulation frequency.
We first apply a temporal Fourier transformation to the inhomogeneous wave equation to obtain the inhomogeneous Helmholtz equation
$c^{2}\nabla^{2}\vec{\bar{E}}+\omega^{2}\vec{\bar{E}}=-4\pi i\omega\vec{\bar{j}}_{\perp}~{},$ (122)
where $\vec{\bar{j}}_{\perp}(\vec{r},\omega)$ is the Fourier transform of the current density $\vec{j}_{\perp}(\vec{r},t)$. The solution can be represented as a weighted superposition of solutions corresponding to a unit point source located at $\vec{r}^{\prime}$. The Green function for the inhomogeneous Helmholtz equation is given by (for unbounded space and outgoing waves)
$4\pi G(\vec{r},\vec{r^{\prime}},\omega)=\frac{1}{|\vec{r}-\vec{r^{\prime}}|}\exp\left[i\frac{\omega}{c}|\vec{r}-\vec{r^{\prime}}|\right]~{},$ (123)
with $|\vec{r}-\vec{r^{\prime}}|=\sqrt{(x^{\prime}-x)^{2}+(y^{\prime}-y)^{2}+(z^{\prime}-z)}^{2}$. With the help of this Green function we can write a formal solution for the field equation as:
[…]
This is simply a mathematical description of the concept of Huygens’ secondary sources and wave propagation, which is, of course, well-known. However, it is worth recalling how this directly follows from Maxwell’s equations. We can consider the amplitude of the radiation emitted by the plane of oscillating electrons as the resultant of radiated spherical wavelets. This is because Maxwell’s theory exhibits no intrinsic anisotropy. The electrons situated on the plane of simultaneity each generate spherical wavelets, which, according to Huygens’ principle, combine to form the resulting radiated wave. If the plane of simultaneity is the $xy$-plane (i.e., the beam modulation wavefront is perpendicular to the $z$-axis), Huygens’ construction reveals that plane wavefronts will be emitted along the $z$-axis.
In summary, according to Maxwell’s electrodynamics, coherent radiation is always emitted in the direction normal to the modulation wavefront. We have already emphasized that Maxwell’s equations are valid only within a Lorentz reference frame, i.e., in an inertial frame where the Einstein synchronization procedure is applied to assign values to the time coordinates. It is crucial to apply Einstein’s time order consistently, both in dynamics and electrodynamics. Our previous discussion naturally leads to the conclusion that Maxwell’s equations in the lab frame are compatible only with covariant trajectories, $\vec{x}_{cov}(t)$, which are calculated using Lorentz coordinates and, therefore, include relativistic kinematic effects.
Let us revisit the modulated electron beam that was kicked transversely with respect to its direction of motion, as discussed earlier. Conventional particle tracking indicates that, while the direction of the electron beam changes after the kick, the orientation of the modulation wavefront remains unchanged. In other words, the direction of the electron’s motion and the normal to the wavefront are not aligned. Therefore, according to the conventional coupling of fields and particles, which we consider incorrect, the coherent undulator radiation produced in the kicked direction downstream in the undulator is expected to be significantly suppressed once the kick angle exceeds the divergence of the output coherent radiation.
To estimate the loss in radiation efficiency in the kicked direction using the conventional coupling of fields and particles, we assume that the spatial profile of the modulation closely follows the electron beam’s transverse distribution, modeled as a Gaussian with standard deviation $\sigma$.
A modulated electron beam in an undulator can be viewed as a sequence of periodically spaced radiating oscillators. These oscillators emit radiation that interferes constructively in the forward direction ($\theta=0$) when they are all in phase, leading to strong on-axis emission.
To understand the angular distribution, consider a triangle formed by a radiation path at a small angle $\theta$, with altitude $r\simeq\theta z$ and base $z$. The diagonal path length $s$ exceeds the base by an amount $\Delta=s-z\simeq z\theta^{2}/2$. When $\Delta$ equals one radiation wavelength, destructive interference occurs, as the phase contributions from different oscillators become uniformly distributed over 00 to $2\pi$.
In the limit of a small electron beam size ($\sigma\to 0$), constructive interference occurs within an angle $\Delta\theta\backsimeq\sqrt{c/(\omega L_{w})}=1/(\sqrt{4\pi N_{w}}\gamma_{z})\ll 1/\gamma$, where $L_{w}=\lambda_{w}N_{w}$ is the undulator length. In the limit for the large size of the electron beam, the angle of coherence is about $\Delta\theta\backsimeq c/(\omega\sigma)$ instead. The boundary between these two asymptotes is for sizes of about $\sigma_{\mathrm{dif}}\backsimeq\sqrt{cL_{w}/\omega}$. The parameter $\omega\sigma^{2}/(cL_{w})$ can be referred to as the electron beam Fresnel number. For XFELs, the transverse beam size $\sigma$ typically exceeds $\sigma_{\mathrm{dif}}$, indicating a large Fresnel number. Consequently, the angular distribution of radiation in the far zone is approximately Gaussian with standard deviation $\sigma_{\mathrm{\theta}}\backsimeq c/(\sqrt{2}\omega\sigma)$.
However, under the conventional framework, a discrepancy arises after the beam is kicked: the direction of electron motion no longer aligns with the modulation wavefront. As a result, the radiation intensity in the kicked direction is suppressed and can be approximated as $I\backsimeq I_{0}\exp[-\theta_{k}^{2}/(2\sigma_{\mathrm{\theta}}^{2})]$, where $I_{0}$ is the on-axis intensity without kick and $\theta_{k}$ is the kick angle. This exponential suppression reflects the misalignment between the wavefront and the new direction of electron motion.
We presented a study of the very idealized situation to illustrate the difference between the conventional and covariant coupling of fields and particles. We solved the dynamics problem of the motion of relativistic electrons in the prescribed force field of a weak kicker magnet by working only up to the order of $\gamma\theta_{k}$. This approximation is of particular theoretical interest because it is relatively simple and at the same time forms the basis for understanding relativistic kinematic effects such as relativity of simultaneity.
Let us discuss the region of validity of our small kick angle approximation $\theta_{k}\gamma\ll 1$. Since in XFELs, the Fresnel number is rather large, we can always consider a kick angle that is relatively large compared to the divergence of the output coherent radiation, and, at the same time, it is relatively small compared to the angle $1/\gamma$. In fact, from $\omega\sigma^{2}/(cL_{w})\gg 1$, with some rearranging, we obtain $\sigma_{\theta}^{2}\simeq c^{2}/(\omega^{2}\sigma^{2})\ll c/(\omega L_{w})$. Then we recall that $\sqrt{c/(\omega L_{w})}=1/(\sqrt{4\pi N_{w}}\gamma_{z})\ll 1/\gamma$. Therefore, the first-order approximation used to analyze the kicker setup in this chapter is of practical significance for XFEL engineering.
One of the goals of this chapter is to demonstrate the experimental predictions that we expect from our corrected radiation theory. To illustrate the essential physical principles clearly, we worked out a simple case. Surprisingly, the first-order approximation used to analyze the kicker setup in this chapter also has important practical applications.
As shown above, our corrected coupling of fields and particles predicts an effect that is in complete contrast to the conventional treatment. Specifically, in the ultrarelativistic limit, the plane of simultaneity—the wavefront orientation of the modulation—is always perpendicular to the electron beam velocity. As a result, we predict strong emission of coherent undulator radiation from the modulated electron beam in the direction of the kick, as shown in Fig. 62.
Figure 62: Experimental setup for testing XFEL theory. According to the covariant approach, in the ultrarelativistic limit, the modulation wavefront is always perpendicular to the velocity of the electron beam. As a result, Maxwell’s equations predict a strong emission of coherent undulator radiation in the direction of the applied kick. Experimental results confirm that this prediction holds true.
XFEL experts have indeed observed an apparent wavefront readjustment due to relativistic kinematics effects, though this conclusion was never drawn. In this book, we are the first to consider the idea that the results of the conventional theory of radiation from relativistically moving charges are inconsistent with special relativity. In previous literature, identification of the trajectories in the source part of the usual Maxwell’s equations with the trajectories calculated by conventional particle tracking in the (”single”) lab frame has always been taken as self-evident.
20.4 Modulation Wavefront Tilt and Maxwell’s Theory
In the existing literature, the theoretical analysis of XFELs driven by electron beams with wavefront tilt is typically based on standard simulation codes and conventional applications of Maxwell’s equations. However, even when using only a kicker setup—without the inclusion of undulator radiation—we can demonstrate that the traditional coupling between fields and particles in XFEL theory is fundamentally flawed.
This conventional XFEL theory relies on the absolute time convention (i.e., old kinematics) for describing particle dynamics. In this work, we present a straightforward demonstration of the inconsistency between standard particle tracking methods and Maxwell’s electrodynamics. Our goal is to show, in a clear and simple manner, that conventional XFEL theory fails to accurately describe the electromagnetic field distribution produced by a fast-moving, modulated electron beam downstream of a kicker.
Under Maxwell’s electrodynamics, the fields associated with a modulated electron beam moving at constant velocity exhibit intriguing behavior as the beam velocity approaches the speed of light. Specifically, in the space-time domain, these fields increasingly resemble those of a laser beam (see Appendix A4). For a rapidly moving, modulated electron beam, the electric and magnetic fields are nearly equal in magnitude, transverse, and mutually perpendicular. In the limit $v\to c$, they become virtually indistinguishable from the fields of a laser beam. According to Maxwell’s equations, the wavefront of such a beam is always orthogonal to its direction of propagation. This is indeed the case for virtual laser-like radiation beam in the region upstream the kicker.
Let us now examine the impact of a transverse kick on the modulation wavefront of the electron beam. Conventional particle tracking predicts that the kick causes a misalignment between the direction of electron motion and the normal to the modulation wavefront, effectively tilting the wavefront.
This leads to an inherent contradiction: within the conventional ”single-frame” approach, the post-kick propagation direction of the radiation beam is no longer perpendicular to its wavefront. In other words, the direction of motion of the virtual radiation beam and the normal to its wavefront diverge. As a result, the beam appears to propagate along the kicked trajectory while maintaining a tilted wavefront—a scenario indistinguishable from real radiation in the ultrarelativistic limit, yet fundamentally inconsistent with Maxwell’s electrodynamics.
The existing literature includes theoretical treatments of XFELs driven by electron beams with tilted modulation wavefronts. These analyses typically rely on Maxwell’s equations and standard simulation tools. However, by isolating the kicker setup—excluding the undulator—we demonstrate that the conventional XFEL framework fails to correctly model the coupling between fields and particles.
This discrepancy underscores a persistent issue in XFEL physics. Initially emerging in the context of coherent undulator radiation from ultrarelativistic, modulated electron beams, the problem has now shifted focus to the tilt of the self-electromagnetic fields of the modulated electron beam.
20.5 Discussion
Finally, we offer some remarks on an alternative perspective for analyzing our complex problem. While we have already obtained results using Lorentz coordinatization, it is instructive to understand why coherent undulator radiation is still emitted in the kicked direction when described within the framework of absolute time synchronization.
The underlying physics is, in fact, quite simple. If we interpret the emission of coherent undulator radiation from a kicked, modulated electron beam in terms of the aberration of light, we can gain a remarkably intuitive understanding of the phenomena occurring in the undulator. Within the framework of Galilean-transformed electrodynamics, a measurement of the coherent radiation will detect emission only along the kicked direction. This is directly analogous to the classical aberration effect—i.e., the apparent deviation in the direction of energy transport—associated with light emitted by a moving source in an inertial frame (see Fig. 1).
There is an intuitively appealing way to grasp the aberration of light in this context. As shown in Fig. 39, the aberration effect in an inertial frame can be readily understood within the corpuscular (particle-like) theory of light. This phenomenon arises naturally from the transformation of velocities between reference frames.
Applying this analogy to a transversely kicked electron beam in an undulator, it is reasonable to expect a similar aberration effect. This becomes especially clear when one considers that a pulse of undulator radiation carries a quantifiable amount of electromagnetic energy. Like mass, energy is a conserved quantity, and in many respects, a coherent X-ray pulse behaves analogously to a stream of particles. Accordingly, we expect that the group velocity of undulator radiation pulses transforms according to the same velocity addition rules that apply to particles in an inertial frame. A more rigorous analysis based on the wave theory of light fully supports this expectation.
Let us now demonstrate how Galilean-transformed electrodynamics predicts a deviation in the direction of energy transport for a radiated X-ray pulse.
A modulated electron beam with finite transverse size, when traversing an undulator, effectively acts as an active medium. This medium diffracts the emitted radiation into multiple plane-wave components, each corresponding to a Fourier component of the spatial modulation of the beam. Consider an ultrarelativistic electron beam propagating along the $z$-axis and kicked transversely along the $y$-axis. The transverse electron density is assumed to vary according to $\rho_{e}=g(K_{\perp})\cos(K_{\perp}y)$, where $K_{\perp}$ is the transverse wavenumber associated with a particular Fourier component.
By applying the Galilean transformation and performing the necessary partial differentiations, one arrives at the modified wave equation, Eq. (12). The additional terms introduced by the Galilean transformation give rise to a predicted Doppler effect. One significant consequence of this effect, in the context of absolute time coordinatization, is an angular frequency dispersion in the light emitted from the kicked, modulated electron beam with finite transverse size.
Specifically, the Doppler shift $\Delta\omega$ of the radiated wave, to first order, is given by $\Delta\omega=K_{\perp}v_{y}$, where $v_{y}$ is the transverse velocity component of the kicked beam. This relation implies that a coherent X-ray beam with finite transverse extent propagates along the $y$-axis with a group velocity given by $d\omega/dk_{y}=v_{y}$.
Thus, according to the corrected coupling between conventional particle tracking and electrodynamics, the direction of X-ray beam propagation is identified with the direction of energy transport, rather than with the orientation of the wavefront. This distinction arises because, under Galilean transformation, the energy transport direction and the wave normal transform differently.
20.6 Experimental Test: Electrodynamics - Dynamics Coupling
The fact that our theory aligns well with experimental observations is clearly illustrated by its agreement with results from an ”X-ray beam-splitting” experiment. This technique separates a circularly polarized XFEL pulse from the linearly polarized XFEL background, thereby maximizing the degree of circular polarization.
In this experiment, it was demonstrated that when a modulated electron beam is kicked by an angle significantly larger than the divergence of the XFEL radiation, the modulation wavefront realigns itself along the new direction of beam motion (see Fig. 62). This realignment is essential to explain the emission of coherent radiation from a short undulator placed downstream of the kicker and oriented along the kicked direction—see, for example, Fig. 14 in [80].
Although these results were unexpected at the time, they had immediate practical implications. The observed ”apparent wavefront readjustment” made it possible to eliminate the unwanted linearly polarized background radiation without any additional hardware—an important breakthrough for XFEL polarization control.
In the existing literature, theoretical treatments of XFELs driven by electron beams with tilted modulation wavefronts have been presented (e.g., [81, 82]), relying on Maxwell’s equations and implemented through standard simulation codes using Maxwell solvers. However, we assert that this approach is fundamentally flawed. In the ultrarelativistic limit relevant to XFELs, a tilted modulation wavefront is inconsistent with Maxwell’s electrodynamics.
Specifically, within the Lorentz lab frame—where Maxwell’s equations are valid and Einstein synchronization is enforced—the concept of a tilted modulation wavefront contradicts the principles of special relativity. When the beam is treated properly in Lorentz coordinates, the modulation wavefront must remain perpendicular to the electron beam velocity. Therefore, the notion of a persistent wavefront tilt under these conditions is not physically meaningful.
It is important to note that the results of the beam-splitting experiment at LCLS support our revised understanding of spontaneous undulator emission [78]. The experiment clearly demonstrated that when a modulated electron beam is kicked by an angle much larger than the divergence of the XFEL radiation, the modulation wavefront realigns itself along the new direction of beam motion. As a result, coherent radiation emitted from an undulator placed downstream of the kicker is observed along the kicked direction, with practically no suppression.
Within the framework of conventional XFEL theory, this leads to a second major puzzle. According to the standard model of undulator radiation, if a modulated electron beam is in perfect resonance before the kick, then after the kick it should remain in resonance along the new velocity direction. However, experimental observations contradict this expectation. Specifically, the data show a redshift in the resonance wavelength following the kick. Maximum radiation power is achieved only when the undulator is detuned to match the reduced longitudinal velocity of the beam after the kick [78].
It is worth emphasizing that any linear superposition of radiation fields from individual electrons preserves the fundamental characteristics of single-particle emission—such as its parametric dependence on undulator settings and polarization. Consider a modulated beam deflected by a weak dipole field before entering a downstream undulator. The resulting radiation field is simply the coherent sum of emissions from individual electrons. Since the observed coherent undulator radiation exhibits a redshift after the kick, it follows that the radiation from each individual electron must also experience a redshift under the same conditions.
This line of reasoning reinforces the conclusion that the results of the beam-splitting experiment in [78] validate our correction to the conventional theory of spontaneous undulator emission.
20.7 Debunking the Modulation Wavefront Tilt Myth
In conventional XFEL theory, wavefront tilt is typically regarded as a physically real effect. However, a common misconception in accelerator physics relates to the interpretation of this tilt. In the ultrarelativistic regime, wavefront tilt does not have an exact, objective meaning. Its value depends on the choice of clock synchronization procedure in the lab frame and, consequently, can be assigned arbitrarily within the interval $(0,\theta_{k})$.
For example, when the evolution of a modulated electron beam is described using Lorentz coordinates, the orientation of the modulation wavefront is always perpendicular to the electron beam velocity—i.e., $\theta_{tilt}=0$. This illustrates that wavefront tilt is a coordinate-dependent artifact rather than a physical observable. As a general principle, no genuine physical effect can depend on an arbitrary constant or an arbitrary function.
Let us consider a specific example. Some papers (see, e.g., [79]) claim that wavefront tilt leads to significant degradation of electron beam modulation in XFELs. To analyze this, suppose the modulation wavefront is initially perpendicular to the beam velocity.
One effect that can influence XFEL performance is the presence of betatron oscillations, which introduce an additional spread in longitudinal velocity. Even for particles with identical energies, differing betatron angles lead to different longitudinal velocities. Thus, beyond the velocity spread caused by energy dispersion, betatron motion introduces another source of longitudinal velocity spread.
To assess the impact of this effect, we consider the total dispersion in longitudinal velocity. The deviation from the nominal longitudinal velocity is given by $\Delta v_{z}=v\Delta\gamma/\gamma^{3}-v\Delta\theta^{2}/2$. The finite angular spread of the electron beam results in a difference in arrival times at a given longitudinal position, which is the well-known normal debunching effect.
From the standpoint of conventional XFEL theory, this time difference is said to be amplified by the kick angle $\theta_{k}$. According to conventional (non-covariant) particle tracking, the wavefront is tilted by an angle $\theta_{tilt}=\theta_{k}$. It is commonly believed that this tilt has physical significance, and that the deviation of the longitudinal velocity component—defined as the component perpendicular to the wavefront under Galilean kinematics—is given by $\Delta v_{z}=-v|\Delta\vec{\theta}+\vec{\theta}_{k}|^{2}/2$. If this description were accurate, the cross term $v\vec{\theta}_{k}\cdot\Delta\vec{\theta}$ would indeed lead to a significant degradation of the modulation amplitude. This phenomenon, referred to as modulation smearing, is proposed as a separate mechanism from normal debunching (see [79]).
However, this interpretation overlooks a critical point. Many experts assume that all forms of debunching are physically meaningful. According to the theory of relativity, while normal debunching is an objective effect, the so-called smearing mechanism is not. The proposed smearing is a coordinate-dependent artifact—it arises solely from the choice of a particular reference frame and synchronization convention in four-dimensional spacetime, and thus lacks physical meaning.
Let us now examine, from a physical perspective, why the smearing mechanism is not valid even within Galilean kinematics. In this classical framework, the cross term $v\vec{\theta}_{k}\cdot\Delta\vec{\theta}$ would appear to cause modulation degradation in the forward direction. However, Galilean-transformed electrodynamics tells us that coherent radiation is observable only in the kicked direction. In that direction, the cross term does not appear in the expression for the deviation of the longitudinal velocity component. It follows naturally that the smearing effect is not a real physical phenomenon.
21 Appendix
21.1 A1. Radiation from Moving Charges
We start with the solution of Maxwell’s equation in the space-time domain, the well-known Lienard-Wiechert expression, and we subsequently apply a Fourier transformation. The Lienard-Wiechert expression for the electric field of a point charge $(-e)$ reads (see, e.g. [41]):
$\vec{{E}}(\vec{r}_{o},t)$ $=$ $-e{\vec{n}-\vec{\beta}\over{\gamma^{2}(1-\vec{n}\cdot\vec{\beta})^{3}|\vec{r}_{o}-\vec{r^{\prime}}|^{2}}}-{e\over{c}}{\vec{n}\times[(\vec{n}-\vec{\beta})\times\dot{\vec{\beta}}]\over{(1-\vec{n}\cdot\vec{\beta})^{3}|\vec{r}_{o}-\vec{r^{\prime}}|}}.$ (125)
$R=|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|$ denotes the displacement vector from the retarded position of the charge to the point where the fields are calculated. Moreover, $\vec{\beta}=\vec{v}/c$, $\vec{\dot{\beta}}=\vec{\dot{v}}/c$, while $\vec{v}$ and $\vec{\dot{v}}$ denote the retarded velocity and acceleration of the electron. Finally, the observation time $t$ is linked with the retarded time $t^{\prime}$ by the retardation condition $R=c(t-t^{\prime})$. As is well-known, Eq. (125) serves as a basis for the decomposition of the electric field into a sum of two quantities. The first term on the right-hand side of Eq. (125) is independent of acceleration, while the second term linearly depends on it. For this reason, the first term is called ”velocity field”, and the second ”acceleration field” [41]. The velocity field differs from the acceleration field in several respects, one of which is the behavior in the limit for a very large distance from the electron. There one finds that the velocity field decreases like $R^{-2}$, while the acceleration field only decreases as $R^{-1}$. Let us apply a Fourier transformation:
[…]
As in Eq. (125) one may formally recognize a velocity and an acceleration term in Eq. (127) as well. Since Eq. (127) follows directly from Eq. (125), that is valid in the time domain, the magnitude of the velocity and acceleration parts in Eq. (127), that include terms in $1/R^{2}$ and $1/R$ respectively, do not depend on the wavelength $\lambda$. It is instructive to take advantage of integration by parts. With the help of
${1\over{c}}{d\over{dt^{\prime}}}|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|=-\vec{n}\cdot\vec{\beta}~{}~{}~{}~{}~{}\mathrm{and}~{}~{}~{}~{}{d\vec{n}\over{dt^{\prime}}}={c\over{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|}}\left[-\vec{\beta}+\vec{n}\left(\vec{n}\cdot\vec{\beta}\right)\right]~{},$ (128)
Eq. (127) can be written as
$\vec{\bar{E}}$ (((( $\vec{r}_{o},\omega)=-{e}\int_{-\infty}^{\infty}dt^{\prime}{\vec{n}\over{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|^{2}}}\exp\left[i\omega\left(t^{\prime}+{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right]$ (130)
[…]
Eq. (LABEL:revtrasfbiss) may now be integrated by parts. When edge terms can be dropped one obtains [83]
[…]
×$\times\exp\left\{i\omega\left(t^{\prime}+{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right\}~{}.$
The only assumption made going from Eq. (127) to Eq. (133) is that edge terms in the integration by parts can be dropped. This assumption can be justified by means of physical arguments in the most general situation accounting for the fact that the integral in $dt^{\prime}$ has to be performed over the entire history of the particle and that at $t^{\prime}=-\infty$ and $t^{\prime}=+\infty$, the electron does not contribute to the field anymore. Let us give a concrete example for an ultra-relativistic electron. Imagine that bending magnets are placed at the beginning and at the end of a given setup, such that they deflect the electron trajectory of an angle much larger than the maximal observation angle of interest for radiation from a bending magnet. This means that the magnets would be longer than the formation length associated with the bends, i.e. $L_{\mathrm{fb}}\sim(c\rho^{2}/\omega)^{1/3}$, where $\rho$ is the bending radius. In this way, intuitively, the magnets act like switches: the first magnet switches the radiation on, and the second switches it off. Then, what precedes the upstream bend and what follows the downstream bend does not contribute to the field detected at the screen position. With this caveat Eq. (133) is completely equivalent to Eq. (127).
The derivation of Eq. (133) is particularly instructive because shows that each term in Eq. (133) is due to a combination of velocity and acceleration terms in Eq. (127). In other words the terms in $1/R$ and in $1/R^{2}$ in Eq. (133) appear as a combination of the terms in $1/R$ (acceleration term) and $1/R^{2}$ (velocity term) in Eq. (127). As a result, one can say that there exist contributions to the radiation from the velocity part in Eq. (127). The presentation in Eq. (133) is more interesting with respect to that in Eq. (127) (although equivalent to it) because of the magnitude of the $1/R^{2}$-term in Eq. (133) can directly be compared with the magnitude of the $1/R$-term inside the integral sign.
The bottom line is that physical sense can be ascribed only to the integral in Eq. (127) or Eq. (133). The integrand is, in fact, an artificial construction. In this regard, it is interesting to note that the integration by parts giving Eq. (133) is not unique. First, we find that [83]
${\vec{n}\times[(\vec{n}-\vec{\beta})\times\vec{\dot{\beta}}]\over{|\vec{r_{o}}-\vec{r^{\prime}}|(1-\vec{n}\cdot\vec{\beta})^{2}}}$ $=$ ${1\over{|\vec{r_{o}}-\vec{r^{\prime}}|}}{d\over{dt^{\prime}}}\left[{\vec{n}\times(\vec{n}\times\vec{\beta})\over{(1-\vec{n}\cdot\vec{\beta})}}\right]$ (135)
$-\left[{\vec{\dot{n}}(\vec{n}\cdot\vec{\beta})+\vec{n}(\vec{\dot{n}}\cdot\vec{\beta})-\vec{\dot{n}}(\vec{n}\cdot\vec{\beta})^{2}-\vec{\beta}(\vec{\dot{n}}\cdot\vec{\beta})\over{|\vec{r_{o}}-\vec{r^{\prime}}|(1-\vec{n}\cdot\vec{\beta})^{2}}}\right].$
Note that Eq. (135) accounts for the fact that $\vec{n}=(\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime}))/|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|$ is not a constant in time. Using Eq. (135) in the integration by parts, we obtain
[…]
×$\times\exp\left\{i\omega\left(t^{\prime}+{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right\}~{}.$
Similarly as before, the edge terms have been dropped. Eq. (127), Eq. (133) and Eq. (137) are equivalent but include different integrands. This is no mistake, as different integrands can lead to the same integral.
If the position of the observer is far away enough from the charge, one can make the expansion $\left|\vec{r}_{0}-\vec{r}(t)\right|=r_{0}-\vec{n}\cdot\vec{r}(t)$. Using Eq. (137), we obtain Eq. (LABEL:revwied).
21.2 A2. Undulator Radiation in the Far Zone
Calculations pertaining undulator radiation are well established see e.g. [84]. In all generality, the field in Eq. (51) can be written as
${\vec{\widetilde{E}}}=\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\frac{i\omega e}{c^{2}z_{0}}$ (138)
×$\times\int_{-L/2}^{L/2}dz^{\prime}\left\{\frac{K}{2i\gamma}\left[\exp\left(2ik_{w}z^{\prime}\right)-1\right]\vec{e}_{x}+\vec{\theta}\exp\left(ik_{w}z^{\prime}\right)\right\}$ (139)
×$\times\exp\left[i\left(C+{\omega\theta^{2}\over{2c}}\right)z^{\prime}-{K\theta_{x}\over{\gamma}}{\omega\over{k_{w}c}}\cos(k_{w}z^{\prime})-{K^{2}\over{8\gamma^{2}}}{\omega\over{k_{w}c}}\sin(2k_{w}z^{\prime})\right]~{}.$ (140)
Here $\omega=\omega_{r}+\Delta\omega$, $C=k_{w}\Delta\omega/\omega_{r}$ and
$\omega_{r}=2k_{w}c\bar{\gamma}_{z}^{2}~{},$ (142)
is the fundamental resonance frequency.
Using the Anger-Jacobi expansion:
$\exp\left[ia\sin(\psi)\right]=\sum_{p=-\infty}^{\infty}J_{p}(a)\exp\left[ip\psi\right]~{},$ (143)
where $J_{p}(\cdot)$ indicates the Bessel function of the first kind of order $p$, to write the integral in Eq. (141) in a different way:
${\vec{\widetilde{E}}}=\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\frac{i\omega e}{c^{2}z_{0}}\sum_{m,n=-\infty}^{\infty}J_{m}(u)J_{n}(v)\exp\left[\frac{i\pi n}{2}\right]$ (144)
×$\times\int_{-L/2}^{L/2}dz^{\prime}\exp\left[i\left(C+{\omega\theta^{2}\over{2c}}\right)z^{\prime}\right]$ (145)
×$\times\left\{\frac{K}{2i\gamma}\left[\exp\left(2ik_{w}z^{\prime}\right)-1\right]\vec{e}_{x}+\vec{\theta}\exp\left(ik_{w}z^{\prime}\right)\right\}\exp\left[i(n+2m)k_{w}z^{\prime}\right]~{},$ (146)
where [^68] Here the parameter $v$ should not be confused with the velocity.
$u=-\frac{K^{2}\omega}{8\gamma^{2}k_{w}c}~{}~{}~{}~{}\mathrm{and}~{}~{}~{}v=-\frac{K\theta_{x}\omega}{\gamma k_{w}c}~{}.$ (148)
Up to now we just re-wrote Eq. (51) in a different way. Eq. (51) and Eq. (147) are equivalent. Of course, definition of $C$ is suited to investigate frequencies around the fundamental harmonic but no approximation is taken besides the paraxial approximation.
Whenever
$C+\frac{\omega\theta^{2}}{{2c}}\ll k_{w}~{},$ (149)
[…]
Within the resonance approximation we further select frequencies such that $|\Delta\omega|/\omega_{r}\ll 1$ (i.e. $|C|\ll k_{w}$). Note that this condition on frequencies automatically selects observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$. In fact, if one considers observation angles outside the range $\theta^{2}\ll 1/\gamma_{z}^{2}$, condition (149) is not fulfilled, and the integrand in Eq. (147) exhibits fast oscillations on the integration scale $L$. As a result, one obtains zero transverse field, $\vec{\widetilde{E}}=0$, with accuracy $1/N_{w}$. Under the constraint imposed by $|\Delta\omega|/\omega_{r}\ll 1$ , independently of the value of $K$ and for observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$, we have
$|v|={K|\theta_{x}|\over{\gamma}}{\omega\over{k_{w}c}}=\left(1+\frac{\Delta\omega}{\omega_{r}}\right)\frac{2\sqrt{2}K}{\sqrt{2+K^{2}}}\bar{\gamma}_{z}|\theta_{x}|\lesssim\bar{\gamma}_{z}|\theta_{x}|\ll 1~{}.$ (150)
[…]
Similar reasonings can be done for frequencies around higher harmonics with a different definition of the detuning parameter $C$. However, around odd harmonics, the before-mentioned expansion, together with the application of the resonance approximation for $N_{w}\gg 1$ (fast oscillating terms in powers of $\exp[ik_{w}z^{\prime}]$ effectively average to zero), yields extra-simplifications.
Here we are dealing specifically with the first harmonic. Therefore, these extra-simplifications apply. We neglect both the term in $\cos(k_{w}z^{\prime})$ in the phase of Eq. (141) and the term in $\vec{\theta}$ in Eq. (141). First, non-negligible terms in the expansion of $J_{n}(v)$ are those for small values of $n$, since $J_{n}(v)\sim v^{n}$, with $|v|\ll 1$. The value $n=0$ gives a non-negligible contribution $J_{0}(v)\sim 1$. Then, since the integration in $dz^{\prime}$ is performed over a large number of undulator periods $N_{w}\gg 1$, all terms of the expansion in Eq. (147) but those for $m=-1$ and $m=0$ average to zero due to resonance approximation. Note that surviving contributions are proportional to $K/\gamma$, and can be traced back to the term in $\vec{e}_{x}$ only, while the term in $\vec{\theta}$ in Eq. (147) averages to zero for $n=0$. Values $n=\pm 1$ already give negligible contributions. In fact, $J_{\pm 1}(v)\sim v$. Then, the term in $\vec{e}_{x}$ in Eq. (147) is $v$ times the term with $n=0$ and is immediately negligible, regardless of the values of $m$. The term in $\vec{\theta}$ would survive averaging when $n=1,~{}m=-1$ and when $n=-1,~{}m=0$. However, it scales as $\vec{\theta}v$. Now, using condition $|\Delta\omega|/\omega_{r}\ll 1$ we see that, for observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$, $|\vec{\theta}|~{}|v|\sim(\sqrt{2}~{}K~{}/\sqrt{2+K^{2}}~{})~{}\bar{\gamma}_{z}\theta^{2}\ll K/\gamma$. Therefore, the term in $\vec{\theta}$ is negligible with respect to the term in $\vec{e}_{x}$ for $n=0$, that scales as $K/\gamma$. All terms corresponding to larger values of $|n|$ are negligible.
Summing up, all terms of the expansion in Eq. (143) but those for $n=0$ and $m=-1$ or $m=0$ give negligible contribution. After definition of
$A_{JJ}=J_{0}\left(\frac{\omega K^{2}}{8k_{w}c\gamma^{2}}\right)-J_{1}\left(\frac{\omega K^{2}}{8k_{w}c\gamma^{2}}\right)~{},$ (151)
that can be calculated at $\omega=\omega_{r}$ since $|C|\ll k_{w}$, we have
[…]
21.3 A3. Approximating the Electron Path
Let us now discuss the case of the radiation from a single electron with an arbitrary angular deflection $\vec{\eta}$ and an arbitrary offset $\vec{l}$ with respect to a reference orbit defined as the path through the origin of the coordinate system, that is $x(0)=y(0)=0$.
If the magnetic field in the setup does not depend on the transverse coordinates, i.e. $B=B(z)$, an initial offset $x(0)=l_{x}$, $y(0)=l_{y}$ shifts the path of an electron of $\vec{l}$. Similarly, an angular deflection $\vec{\eta}=(\eta_{x},\eta_{y})$ at $z=0$ tilts the path without modifying it. Cases when the magnetic field of SR sources include focusing elements (or the natural focusing of insertion devices) are out of the scope of this paper. Assuming further that $\eta_{x}\ll 1$ and $\eta_{y}\ll 1$, which is typically justified for ultrarelativistic electron beams, one obtains the following approximation for the electron path:
$x(z)=x_{r}(z)+\eta_{x}z+l_{x}~{},y(z)=y_{r}(z)+\eta_{y}z+l_{y}~{},$ (154)
where the subscript ‘r’ refers to the reference path. This gives a parametric description of the path of a single electron with offset $\vec{l}$ and deflection $\vec{\eta}$. The curvilinear abscissa on the path can then be written as
[…]
$=s_{r}(z)+\frac{\eta^{2}z}{2}+x_{r}(z)\eta_{x}+y_{r}(z)\eta_{y}~{},$ (158)
[…]
We now substitute Eq. (155) and Eq. (158) into Eq. (43) to obtain:
[…]
×$\times\left[\left({v_{x}(z^{\prime})\over{c}}-(\theta_{x}-\eta_{x})\right){\vec{e_{x}}}+\left({v_{y}(z^{\prime})\over{c}}-(\theta_{y}-\eta_{y})\right){\vec{e_{y}}}\right]~{},$ (160)
where the total phase $\Phi_{T}$ is
$\Phi_{T}=\omega\left[\frac{s_{r}(z^{\prime})}{v}+\frac{\eta^{2}z^{\prime}}{2v}+\frac{1}{v}[x_{r}(z^{\prime})\eta_{x}+y_{r}(z^{\prime})\eta_{y}]-{z^{\prime}\over{c}}\right]$ (162)
$+\frac{\omega}{2c}\left[z_{0}(\theta_{x}^{2}+\theta_{y}^{2})-2\theta_{x}x_{r}(z^{\prime})-2\theta_{x}\eta_{x}z^{\prime}-2\theta_{x}l_{x}\right.$ (163)
$\left.-2\theta_{y}y(z^{\prime})-2\theta_{y}\eta_{y}z^{\prime}-2\theta_{y}l_{y}+z^{\prime}(\theta_{x}^{2}+\theta_{y}^{2})\right]~{},$ (164)
which can be rearranged as
$\Phi_{T}\simeq\omega\left[{s_{r}(z^{\prime})\over{v}}-{z^{\prime}\over{c}}\right]-\frac{\omega}{c}(\theta_{x}l_{x}+\theta_{y}l_{y})$ (165)
$+\frac{\omega}{2c}\left[z_{0}(\theta_{x}^{2}+\theta_{y}^{2})-2(\theta_{x}-\eta_{x})x_{r}(z^{\prime})\right.$ (166)
$\left.-2(\theta_{y}-\eta_{y})y_{r}(z^{\prime})+z^{\prime}\left((\theta_{x}-\eta_{x})^{2}+(\theta_{y}-\eta_{y})^{2}\right)\right]~{}.$ (167)
21.4 A4. Self-Fields of a Modulated Electron Beam
In general, the electrodynamical theory is based on the exploitation, for the ultra-relativistic particles, of the small parameter $1/\gamma^{2}$. By this, Maxwell’s equations are reduced to much simpler equations with the help of paraxial approximation.
Whatever the method used to present results, one needs to solve Maxwell’s equations in unbounded space. We introduce a cartesian coordinate system, where a point in space is identified by a longitudinal coordinate $z$ and transverse position $\vec{r}_{\perp}$. Accounting for electromagnetic sources, i.e. in a region of space where current and charge densities are present, the following equation for the field in the space-frequency domain holds in all generality:
$c^{2}\nabla^{2}\vec{\bar{E}}+\omega^{2}\vec{\bar{E}}=4\pi c^{2}\vec{\nabla}\bar{\rho}-4\pi i\omega\vec{\bar{j}}~{},$ (169)
where $\bar{\rho}(z,\vec{r}_{\perp},\omega)$ and $\vec{\bar{j}}(z,\vec{r}_{\perp},\omega)$ are the Fourier transforms of the charge density $\rho(z,\vec{r}_{\perp},t)$ and of the current density $\vec{j}(z,\vec{r}_{\perp},t)$. Eq. (169) is the well-known Helmholtz equation. Here $\vec{\bar{E}}$ indicates the Fourier transform of the electric field in the space-time domain.
Eq. (169) can be solved with the help of an appropriate Green’s function $G(z-z^{\prime},\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot})$ yielding
[…]
the integration in $d\vec{r^{\prime}}_{\bot}$ being performed over the entire transverse plane. An explicit expression for the Green’s function to be used in Eq. (170) is given by
$G(z-z^{\prime},\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot})$ $=$ $-\frac{\exp\left\{i(\omega/c)\left[\left|\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}\right|^{2}+\left(z-z^{\prime}\right)^{2}\right]^{1/2}\right\}}{4\pi\left[\left|\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}\right|^{2}+\left(z-z^{\prime}\right)^{2}\right]^{1/2}}~{},$ (171)
that automatically includes the proper boundary conditions at infinity.
[…]
$\mathcal{D}\left[\vec{\widetilde{E}}_{\bot}(z,\vec{r}_{\bot},\omega)\right]=\vec{g}(z,\vec{r}_{\bot},\omega)~{},$ (172)
where the differential operator $\mathcal{D}$ is defined by
$\mathcal{D}\equiv\left({\nabla_{\bot}}^{2}+{2i\omega\over{c}}{\partial\over{\partial z}}\right)~{},$ (173)
${\nabla_{\bot}}^{2}$ being the Laplacian operator over transverse cartesian coordinates. Eq. (172) is Maxwell’s equation in paraxial approximation.
[…]
The source-term vector $\vec{g}(z,\vec{r}_{\bot})$ is specified by the trajectory of the source electrons, and can be written in terms of the Fourier transform of the transverse current density, $\vec{\bar{j}}_{\bot}(z,\vec{r}_{\bot},\omega)$, and of the charge density, $\bar{\rho}(z,\vec{r}_{\bot},\omega)$, as
$\vec{g}=$ $-{4\pi}\exp\left[-\frac{i\omega z}{c}\right]\left(\frac{i\omega}{c^{2}}\vec{\bar{j}}_{\bot}-\vec{\nabla}_{\bot}\bar{\rho}\right)~{}.$ (174)
$\vec{\bar{j}}_{\bot}$ and $\bar{\rho}$ are regarded as given data. We will treat $\vec{\bar{j}}_{\bot}$ and $\bar{\rho}$ as macroscopic quantities, without investigating individual electron contributions. In the time domain, we may write the charge density $\rho(\vec{r},t)$ and the current density $\vec{j}(\vec{r},t)$ as
$\rho(\vec{r},t)=\frac{1}{v}\rho_{\bot}(\vec{r}_{\bot})f\left(t-\frac{z}{v}\right)$ (175)
and
$\vec{j}(\vec{r},t)$ $=$ $\frac{1}{v}\vec{v}\rho_{\bot}(\vec{r}_{\bot})f\left(t-\frac{z}{v}\right)~{},$ (176)
where $v$ denote the velocity of the electron. The quantity $\rho_{\bot}$ has the meaning of transverse electron beam distribution, while $f$ is the longitudinal charge density distribution.
In the space-frequency domain, Eq. (175) and Eq. (LABEL:curr) transform to:
$\bar{\rho}(\vec{r}_{\bot},z,\omega)=\rho_{\bot}\left(\vec{r}_{\bot}\right)\bar{f}(\omega)\exp\left[{i\omega z)/v}\right]~{},$ (178)
and
$\vec{\bar{j}}(\vec{r}_{\bot},z,\omega)=\vec{v}\rho_{\bot}\left(\vec{r}_{\bot}\right)\bar{f}(\omega)\exp\left[{i\omega z/v}\right]~{}.$ (179)
It should be remarked that $\bar{\rho}$ and $\vec{\bar{j}}=\bar{\rho}\vec{v}$ satisfy the continuity equation. In other words, one can find $\vec{\nabla}\cdot\vec{\bar{j}}=i\omega\bar{\rho}$.
We find an exact solution of Eq. (172) without any other assumption about the parameters of the problem. A Green’s function for Eq. (172), namely the solution corresponding to the unit point source can be written as (see e.g. [83]):
$G(z-z^{\prime};\vec{r_{\bot}}-\vec{r^{\prime}_{\bot}})$ $=$ $-{1\over{4\pi(z-z^{\prime})}}\exp\left\{i\omega{\mid\vec{r_{\bot}}-\vec{r^{\prime}_{\bot}}\mid^{2}\over{2c(z-z^{\prime})}}\right\}~{},$ (180)
assuming $z-z^{\prime}>0$. When $z-z^{\prime}<0$ the paraxial approximation does not hold, and the paraxial wave equation Eq. (172) should be substituted, in the space-frequency domain, by a more general Helmholtz equation. Yet, the radiation formation length for $z-z^{\prime}<0$ is very short with respect to the case $z-z^{\prime}>0$, i.e. we can neglect contributions from sources located at $z-z^{\prime}<0$.
Since it is assumed that electrons are moving along the $z$-axis, we have $\vec{\bar{j}}_{\perp}=0$. Thus, after integration by parts, we obtain the solution
[…]
×$\times\frac{1}{z-z^{\prime}}{\rho_{\bot}}\left(\vec{r^{\prime}}_{\bot}\right)\left(\frac{\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}}{z-z^{\prime}}\right)~{}.$
Eq. (LABEL:ggeneralfin) describes the field at any position $z$.
[…]
$\vec{\widetilde{E}}_{\bot}(z,\vec{r}_{\bot})$ $=$ $-\frac{i\omega\bar{f}(\omega)}{c}\int d\vec{r^{\prime}}_{\bot}{\rho_{\bot}}\left(\vec{r^{\prime}}_{\bot}\right)\exp\left[\frac{i\omega z}{2c\gamma^{2}}\right]\Bigg{\{}\left[\frac{ic}{\omega}\frac{\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}}{\mid\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}\mid}\cdot\frac{d}{d\left[\mid\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}\mid\right]}\right]$ (185)
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\int_{0}^{\infty}\frac{d\xi}{\xi}\exp\left[+i\omega\frac{\mid\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}\mid^{2}}{2c\xi}-\frac{i\omega\xi}{2c\gamma^{2}}\right]\Bigg{\}}$
We now use the fact that, for any real number $\alpha>0$:
[…]
where $K_{0}$ is the zero order modified Bessel function of the second kind. Using Eq. (186) we can write Eq. (185) as
[…]
where $K_{1}(\cdot)$ is the modified Bessel function of the first order.
[…]
$\vec{\widetilde{E}}_{\bot}(z,\vec{r}_{\bot})=-\frac{2\omega e}{c^{2}\gamma}\exp\left[\frac{i\omega z}{2c\gamma^{2}}\right]\frac{\vec{r}_{\bot}}{r_{\bot}}K_{1}\left(\frac{\omega r_{\bot}}{c\gamma}\right)~{}.$ (189)
[…]

### v12

@@
 Evgeny Saldin Final draft
-New material has been added in the Chapter 4 and Chapter 16. If you have any comments, please send an email to evgueni.saldin@desy.de.
+Current draft was completed by a new Chapter 17. If you have any comments, please send an email to evgueni.saldin@desy.de.
 *Preface
@@
 However, this approach fails to account for relativistic effects and leads to inconsistencies in non-collinear motion.
-Only by solving the equations of motion in their fully covariant form can one correctly capture the coupling between Maxwell’s equations and particle trajectories in the lab frame.
 Solving problems related to covariant particle tracking presents two key challenges.
@@
 Only by solving the equations of motion within a fully covariant framework—one that includes the Wigner rotation in the treatment of non-collinear velocities—can the coupling between Maxwell’s equations and particle trajectories in the laboratory frame be properly described.
-In Chapter 16, we explore which aspects of dynamical theory depend on the choice of synchronization convention and which remain independent.
 Our analysis highlights the distinction between the concepts of path and trajectory.
@@
 Instead, the trajectory serves as a theoretical tool in the analysis of electrodynamics problems.
-In Chapter 18, we present a critical reexamination of radiation theory.
+In Chapter 16, we derived specific quantitative results for relativistic velocity addition using only elementary principles of special relativity.
+However, readers seeking a more mathematically rigorous treatment may prefer Chapter 17, where we present a comprehensive account of the mathematical structure underlying the composition of Lorentz boosts.
+In Chapter 19, we present a critical reexamination of radiation theory.
 To evaluate radiation fields generated by external sources, we must determine the particle velocity $\vec{v}$ and position $\vec{x}$ as functions of the lab-frame time $t$.
@@
 Many physicists, having learned special relativity from textbooks, might say, “For those who want just enough understanding to solve problems, special relativity merely modifies Newton’s laws by introducing a correction factor to mass.” However, this is a misconception.
-The velocity of light transforms like that of a particle, and one cannot simply apply classical kinematics to mechanics while using Einstein’s kinematics for electrodynamics.
+The velocity of light transforms like that of a particle, and one cannot simply apply classical kinematics to mechanics while using relativistic kinematics for electrodynamics.
 We highlight errors in the standard approach to coupling fields and particles through a relatively simple example, ensuring that the core physical principles remain clear and unobscured by unnecessary mathematical complexity.
@@
 In reality, electron dynamics in storage rings are significantly influenced by radiation emission.
-We address this question in detail in Chapter 19, focusing on spontaneous synchrotron radiation from bending magnets and undulators.
+We address this question in detail in Chapter 20, focusing on spontaneous synchrotron radiation from bending magnets and undulators.
 Just as the non-relativistic asymptote simplifies calculations, the ultrarelativistic asymptote does the same for covariant analysis, as it naturally leads to the paraxial approximation.
@@
 An X-ray free electron laser (XFEL) is an example where improvements in accelerator technology make it possible to develop ultrarelativistic macroscopic objects with an internal structure (modulated electron bunches), and the relativistic kinematics plays an essential role in their description.
-In Chapter 20, we critically reexamine existing XFEL theory, particularly for readers with limited knowledge of accelerator physics.
+In Chapter 21, we critically reexamine existing XFEL theory, particularly for readers with limited knowledge of accelerator physics.
 We focus on the production of coherent undulator radiation by a modulated ultrarelativistic electron beam.
-Fortunately, understanding this process does not require prior knowledge of undulator radiation theory from Chapter 19, as it can be explained in a straightforward manner.
+Fortunately, understanding this process does not require prior knowledge of undulator radiation theory from Chapter 20, as it can be explained in a straightforward manner.
 Relativistic kinematics plays a fundamental role in XFEL physics, particularly through the rotation of the modulation wavefront, which, in the ultrarelativistic approximation, is closely linked to the relativity of simultaneity.
@@
 the orientation of the plane of simultaneity, is always perpendicular to the electron beam velocity when the evolution of the beam’s evolution is described using Lorentz coordinates.
-It is important to note that Maxwell’s equations are valid only in Lorentz reference frames.
-Einstein’s time order must be consistently applied in both dynamics and electrodynamics.
-Maxwell’s equations dictate that the wavefront of a laser beam is always orthogonal to its propagation direction.
-At this point, we emphasize that, according to relativity, a modulated electron beam in the ultrarelativistic asymptotic limit exhibits the same kinematics as a laser beam in Lorentz coordinates.
-Experiments show that this prediction is, in fact, true.
+It is important to recognize that Maxwell’s equations are valid only within Lorentz reference frames.
+Consequently, Einstein’s concept of time ordering must be applied consistently in both dynamics and electrodynamics.
+According to Maxwell’s equations, the wavefront of a laser beam is always orthogonal to its direction of propagation.
+Notably, in the framework of special relativity, a modulated electron beam in the ultrarelativistic (asymptotic) limit exhibits the same kinematic behavior as a laser beam when described in Lorentz coordinates.
+Experimental evidence confirms the accuracy of this prediction.
 The theory of relativity, formulated within a space-time framework based on pseudo-Euclidean geometry, has been developed for over a century.
@@
 2.1 What is Special Relativity?
-The laws of physics are invariant under Lorentz transformations.
-This invariance is a restrictive principle—it does not, by itself, determine the precise form of the dynamics involved.
-Understanding the postulates of special relativity is analogous to grasping the principle of energy conservation: initially, we accept it as a foundational principle, and later, we seek microscopic interpretations that are consistent with it.
-For any system governed by energy conservation, there should exist a more fundamental theory that provides insight into the detailed physical processes.
-Naturally, this deeper theory must also obey the principle of energy conservation.
-The conservation of energy is a powerful tool, allowing us to analyze systems even in the absence of a complete fundamental theory.
+Special relativity is founded on the principle that the laws of physics are invariant under Lorentz transformations.
+This principle is inherently restrictive: it does not dictate the specific form of the dynamics involved, but rather constrains any physical theory to be consistent with Lorentz symmetry.
+Understanding the postulates of special relativity is conceptually similar to understanding the principle of energy conservation.
+Initially, we encounter energy conservation as a general rule.
+Later, we seek deeper, microscopic theories—such as classical mechanics or quantum field theory—that not only comply with this principle but also explain the underlying mechanisms.
+These deeper theories must, of course, recover energy conservation as a consequence of their more fundamental laws.
+The utility of the energy conservation principle lies in its ability to guide analysis, even when the full fundamental theory is not yet known.
 A similar methodological perspective applies to the postulates of special relativity.
-For instance, suppose we do not understand the microscopic reasons for muon decay, but we do know the decay law in the muon’s rest frame.
-This decay law, in such a case, serves as a phenomenological model.
-Its relativistic generalization to any Lorentz frame allows us to predict, for example, the average distance a muon travels before decaying.
-Consider the case of muon decay.
-Suppose we lack a detailed understanding of why a muon disintegrates but have empirical knowledge of its decay law in its rest frame.
-This law serves as a phenomenological model.
-By applying relativistic principles, we can extend this law to an arbitrary Lorentz frame, enabling us to predict the muon’s average travel distance.
-Specifically, if we apply a Lorentz transformation to the decay law, we find that after a distance $\gamma v\tau_{0}$, the muon population in the lab frame is reduced to half of its original value.
-Here, $\gamma$ is the relativistic factor, and $\tau_{0}$ is the proper lifetime.
-This result can be interpreted as an increase in the effective lifetime of the muon from $\tau_{0}$ to $\gamma\tau_{0}$ in the lab frame due to time dilation.
-However, special relativity is not a complete theory.
-Constructive (or microscopic) theories, such as electrodynamics or quantum field theory, offer a deeper understanding of physical processes than purely restrictive frameworks like special relativity.
-Relativistic kinematics serves primarily as a descriptive tool for interpreting how matter fields appear to behave from the perspective of different observers.
-In principle, any relativistic quantity can be derived directly from the underlying theories of matter, without explicitly invoking relativity.
-For example, muons exhibit relativistic behavior in motion because the field forces governing their decay obey Lorentz-covariant quantum field equations.
+For example, suppose we do not fully understand the microscopic mechanism behind muon decay.
+However, if we know the decay law in the muon’s rest frame, we can treat this as a phenomenological law.
+By applying the principles of special relativity, we can generalize this decay law to any Lorentz frame and make testable predictions.
+In particular, if we transform the decay law to the laboratory frame using a Lorentz boost, we find that the population of muons is reduced to half its original value after they travel a distance of $\gamma v\tau_{0}$, where $\gamma$ is the Lorentz factor and $\tau_{0}$ is the muon’s proper lifetime.
+This implies that, from the lab’s perspective, the muon’s lifetime appears to be extended to $\gamma\tau_{0}$.
+Such predictions are direct consequences of the relativistic transformation of time intervals.
+Despite its successes, special relativity is not a complete theory.
+It is a restrictive framework—more a set of symmetry constraints than a constructive (microscopic) explanation.
+Constructive theories, such as electrodynamics or quantum field theory, provide a more detailed understanding of physical phenomena.
+Special relativity, in this context, serves primarily as a kinematical interpretation of how physical processes appear from different inertial frames.
+Importantly, all relativistic effects can, in principle, be derived directly from underlying physical laws—without invoking relativity as an independent framework—provided those laws are Lorentz covariant.
+For instance, muons in motion exhibit relativistic behavior because the quantum field equations governing their interactions and decay are Lorentz invariant.
 In the ”microscopic” view of relativistic phenomena, the Lorentz covariance of the fundamental laws of physics remains, much like energy conservation, an empirical and unexplained fact—but ultimately, all explanations must terminate at some foundational level.
@@
 A common argument against their validity is that they fail to preserve the form-invariance of Maxwell’s equations under a change of inertial frame.
-This reasoning is frequently found in well-known textbooks and monographs  [^1] It is widely believed that Galilean transformations, being pre-relativistic, are incompatible with the special theory of relativity.
+This reasoning is frequently found in well-known textbooks and monographs.
+[^1] It is widely believed that Galilean transformations, being pre-relativistic, are incompatible with the special theory of relativity.
 For instance, Bohm [1] states: “… the Galilean law of addition of velocities implies that the speed of light should vary with the speed of the observing equipment.
@@
 Drake and Purvis [2], for example, write: “One of the great insights of relativity theory was the realization that Galilean transformations are incorrect.
-The proper way to translate space-time measurements between inertial frames is through Lorentz transformations.”.
+The proper way to translate space-time measurements between inertial frames is through Lorentz transformations.”
 However, the assertion that Galilean transformations are inherently incorrect is a misconception.
 Special relativity is fundamentally a theory of four-dimensional space-time with a pseudo-Euclidean geometry.
-From this perspective, the principle of relativity naturally emerges as a geometric property of space-time, which can be described using arbitrary coordinate systems  [^2] The mathematical claim that transitioning to arbitrary coordinates leaves the geometry of space-time unchanged is considered erroneous in standard textbooks.
+From this perspective, the principle of relativity naturally emerges as a geometric property of space-time, which can be described using arbitrary coordinate systems.
+[^2] The mathematical claim that transitioning to arbitrary coordinates leaves the geometry of space-time unchanged is considered erroneous in standard textbooks.
 As L.
@@
 It is easy to verify that this transformation, as was to be expected, does not satisfy the requirements of the theory of relativity; it does not leave the interval between events invariant.” This misconception is attributed to a failure to distinguish between the convention-dependent and convention-invariant aspects of the theory.
-In pseudo-Euclidean geometry, the interval between events remains invariant under arbitrary coordinate transformations..
+In pseudo-Euclidean geometry, the interval between events remains invariant under arbitrary coordinate transformations.
 Thus, contrary to the standard presentation in most textbooks, Galilean transformations are in fact compatible with the principle of relativity, even though they modify the explicit form of Maxwell’s equations.
@@
 This is the sole relativistic effect that appears in the first order of $v/c$.
-Other higher-order effects, such as Lorentz-Fitzgerald contraction and time dilation, arise from higher-order terms and can be derived mathematically by iterating the infinitesimal transformation  [^3] There is a common misconception that the Lorentz transformation reduces to the Galilean transformation in the non-relativistic limit.
+Other higher-order effects, such as Lorentz-Fitzgerald contraction and time dilation, arise from higher-order terms and can be derived mathematically by iterating the infinitesimal transformation.
+[^3] There is a common misconception that the Lorentz transformation reduces to the Galilean transformation in the non-relativistic limit.
 For example, French [4] states, ”The reduction of $t^{\prime}=\gamma(t-vx/c^{2})$ to the Galilean relation $t^{\prime}=t$ requires $x\ll ct$ as well as $v/c\ll 1$”.
@@
 Any coordinate transformation replacing the GT must also agree with this experience, and thus must contain the GT in the nonrelativistic limit.” This conclusion is mathematically flawed.
-As Baierlein [6] points out: ”If the Lorentz transformation for infinitesimal $v/c$ were to reduce to the Galilean transformation, then the iterative process could never generate a finite Lorentz transformation that is radically different from the Galilean transformation..
+As Baierlein [6] points out: ”If the Lorentz transformation for infinitesimal $v/c$ were to reduce to the Galilean transformation, then the iterative process could never generate a finite Lorentz transformation that is radically different from the Galilean transformation.”
 The primary distinction between Lorentz coordinatization and absolute time coordinatization lies in the transformation laws that relate coordinates and time between relatively moving systems.
@@
 This motion along the path is parameterized by $l(t)$, where $l$ represents a certain parameter—specifically, in our case of interest, the arc length.
-Now, if we take the origin of a Cartesian coordinate system and connect it to a point on the path, the resulting vector is the position vector $\vec{x}(l)$  [^6] For a more detailed discussion on the distinction between path and trajectory, see [10]..
+Now, if we take the origin of a Cartesian coordinate system and connect it to a point on the path, the resulting vector is the position vector $\vec{x}(l)$.
+[^6] For a more detailed discussion on the distinction between path and trajectory, see [10].
 The key difference between trajectory $\vec{x}(t)$ and path $\vec{x}(l)$ is fundamental: the path has an exact, objective meaning—it is independent of any convention.
@@
 In this approach, the only modification to classical mechanics is the introduction of relativistic mass, while time remains distinct from space.
-There is no mixing of spatial and temporal coordinates  [^7] Clarification of the true content of the non-covariant theory can be found in various advanced textbooks.
+There is no mixing of spatial and temporal coordinates.
+[^7] Clarification of the true content of the non-covariant theory can be found in various advanced textbooks.
 To quote e.g.
@@
 But in the relativistic context, there are differences in the transformation laws of the various relative quantities (of kinematics or dynamics), when passing from one reference frame to another.” These authors emphasize the unique role of a ”single inertial frame.” This term suggests that a defining characteristic of non-covariant theory is the absence of relativistic kinematics in describing particle motion.
-A similar viewpoint is expressed by Friedman [15]: ”Within any single inertial frame, things look precisely the same as in Newtonian kinematics: there is an enduring Euclidean three-space, a global time $t$, and the law of motion.”.
+A similar viewpoint is expressed by Friedman [15]: ”Within any single inertial frame, things look precisely the same as in Newtonian kinematics: there is an enduring Euclidean three-space, a global time $t$, and the law of motion.”
 Many fascinating phenomena arise when charges move under the influence of electromagnetic fields, often in highly complex situations.
@@
 $\vec{\nabla}\cdot\vec{B}=0~,$ (6)
-×$\vec{\nabla}\times\vec{E}=-\frac{1}{c}\frac{\partial\vec{B}}{\partial t}~{},$ (7)
-×$\vec{\nabla}\times\vec{B}=\frac{4\pi}{c}\vec{j}+\frac{1}{c}\frac{\partial\vec{E}}{\partial t}~{}.$ (8)
+$\vec{\nabla}\times\vec{E}=-\frac{1}{c}\frac{\partial\vec{B}}{\partial t}~,$ (7)
+$\vec{\nabla}\times\vec{B}=\frac{4\pi}{c}\vec{j}+\frac{1}{c}\frac{\partial\vec{E}}{\partial t}~.$ (8)
 Here the charge density $\rho$ and current density $\vec{j}$ are written as
@@
 The concept of synchronization is central to understanding special relativity.
-There are multiple methods to synchronize distant clocks [^8] According to the thesis of the conventionality of simultaneity [13, 14, 15, 16, 17], simultaneity between distant events is a matter of convention, as it can be legitimately defined in different ways within any given inertial reference frame.
+There are multiple methods to synchronize distant clocks.
+[^8] According to the thesis of the conventionality of simultaneity [13, 14, 15, 16, 17], simultaneity between distant events is a matter of convention, as it can be legitimately defined in different ways within any given inertial reference frame.
 Moeller [13], for example, states: ”All methods for the regulation of clocks meet with the same fundamental difficulty.
 The concept of simultaneity between two events in different places has no exact objective meaning at all since we cannot specify any experimental method by which this simultaneity could be ascertained.
-The same is therefore true also for the concept of velocity.”.
-The choice of a synchronization convention is ultimately a matter of selecting a coordinate system within an inertial frame of reference in Minkowski space-time.
+The same is therefore true also for the concept of velocity.” The choice of a synchronization convention is ultimately a matter of selecting a coordinate system within an inertial frame of reference in Minkowski space-time.
 The space-time continuum can be described using arbitrary coordinates.
@@
 The most natural synchronization method involves first bringing all ideal clocks to the same spatial location, where they can be synchronized.
-Once synchronized, the clocks are then slowly transported to their designated positions—a process known as slow clock transport  [^9] Eddington, in a text first published in 1923, was apparently the first to describe a synchronization procedure based on the slow transport of clocks [18].
-Further details can be found in the review [8] (see also [19])..
+Once synchronized, the clocks are then slowly transported to their designated positions—a process known as slow clock transport.
+[^9] Eddington, in a text first published in 1923, was apparently the first to describe a synchronization procedure based on the slow transport of clocks [18].
+Further details can be found in the review [8] (see also [19]).
 Maxwell’s equations in their usual form are valid in any inertial frame where the sources are at rest, and the time coordinate is assigned using the slow clock transport procedure.
@@
 Applying the Einstein synchronization procedure in the rest frame of the dipole source allows us to establish a Lorentz coordinate system.
-Slow transport synchronization is equivalent to Einstein’s synchronization in an inertial system where the dipole light source is at rest  [^10] Many textbook authors still attribute a reality status to the one-way speed of light.
+Slow transport synchronization is equivalent to Einstein’s synchronization in an inertial system where the dipole light source is at rest.
+[^10] Many textbook authors still attribute a reality status to the one-way speed of light.
 To quote Hrasko [20]: ”It is sometimes claimed that Einstein’s synchronization of distant clocks $A$ and $B$ is circular.
@@
 It therefore provides a solid logical foundation for Einstein’s synchronization prescription.” This reasoning, however, is incorrect.
-Slow clock transport synchronization is equivalent to Einstein’s synchronization only in an inertial system where the light source is at rest..
+Slow clock transport synchronization is equivalent to Einstein’s synchronization only in an inertial system where the light source is at rest.
 In other words, suppose we have two sets of synchronized clocks spaced along the $x$-axis.
@@
 This metric describes the electrodynamics of the moving light source from the viewpoint of the inertial observer measurements.
-From the structure of the equation, it is evident that the measurement of the electromagnetic field configuration—expressed in terms of the coordinates $x$ and $t$ of a measuring device at rest in the inertial frame—yields the same result as that obtained at the spatial position $x-vt$ at time $t=0$.
+From the structure of the equation, it is evident that the measurement of the electromagnetic field configuration—expressed in terms of the coordinates $x$ and $t$ of a measuring device at rest in the inertial frame—yields the same result as that obtained at the spatial position xvtx-vt at time $t=0$.
 To our knowledge, the metric in Eq.
@@
 This last result is closely related to the metric equation Eq. 11.
-The solution of this equation $F[x-(c+v)t]+G[x+(-c+v)t]$ is the sum of two arbitrary functions, one of argument $x-(c+v)t$ and the other of argument $x+(-c+v)t$.
+The solution of this equation F[x(c+v)t]+G[x+(c+v)t]F[x-(c+v)t]+G[x+(-c+v)t] is the sum of two arbitrary functions, one of argument x(c+v)tx-(c+v)t and the other of argument x+(c+v)tx+(-c+v)t.
 Here we obtained the solution for waves that move in the $x$ direction by supposing that the field does not depend on $y$ and $z$.
@@
 In other words, in this scenario, the speed of light follows the Galilean law of velocity addition.
-Specifically, the coordinate velocity of light along the $x$-axis is given by $dx/dt=c+v$ in the positive direction and $dx/dt=-c+v$ in the negative direction.
+Specifically, the coordinate velocity of light along the $x$-axis is given by $dx/dt=c+v$ in the positive direction and dx/dt=c+vdx/dt=-c+v in the negative direction.
 This apparent deviation from the electrodynamics constant $c$ arises because the clocks are synchronized according to an absolute time convention.
@@
 This reveals that, contrary to the textbook understanding, Lorentz transformations do not preserve the continuity of the metric tensor ( see Chapter 5 for further details).
-3.5 A Clock Resynchronization Procedure
-After the Galilean boost in the case of the absolute time coordinatization we can see that the homogeneous wave equation for the field in the lab frame has nearly but not quite the usual, standard form that takes when there is no acceleration.
-The main difference consists in the crossed term $\partial^{2}/\partial t\partial x$, which complicates the solution of the equation.
-To get around this difficulty, we observe that simplification is always possible.
-The trick needed here is to further make a change of the time variable.
-Suppose that we describe the effect of radiation by working only up to the first order in $v/c$.
-In this approximation, it is always possible to introduce a time transformation $t\to t+xv/c^{2}$.
-After this time shift we obtain the usual wave equation.
+At this point, a reasonable question arises: Why address this topic now?
+Why not wait until Chapter 5?
+The answer is that the subject is subtle and complex, and the best way to understand it is through gradual exposure.
+The first step is to develop an intuitive sense of what should occur in various situations.
+Later, with a more solid grasp of relativistic effects, we will revisit the topic and refine our understanding with greater precision.
+3.5 A Clock Re-Synchronization Procedure
+Most existing studies assume that an observer at rest and a moving source share the same Minkowski metric, thereby neglecting the need for any clock re-synchronization procedure.
+However, this standard approach is flawed, as it fails to account for the principle of equivalence between active and passive boosts within a single inertial frame.
+Our analysis shows that after applying a Galilean boost under an absolute time coordinatization, the homogeneous wave equation for the field in the lab frame appears nearly—but not exactly—in its standard form (valid in the absence of acceleration).
+The key deviation lies in the presence of a mixed derivative term, $\partial^{2}/\partial t,\partial x$, which complicates the equation’s solution.
+To resolve this, we note that a simplification is always possible by introducing a suitable change of the time variable.
+If we limit our analysis to first-order terms in $v/c$, a time transformation of the form $t\to t+xv/c^{2}$ can be applied.
+This time shift effectively eliminates the mixed derivative term, reducing the equation to the standard wave equation form.
 Let us now consider the case of an arbitrary velocity along the $x$-axis.
@@
 Consequently, there is no need to revert the solution to its original form.
-Consider two light sources labeled ”1” and ”2”.
+Consider two independent light sources labeled ”1” and ”2”.
 Suppose that in the lab frame the velocities of ”1” and ”2” are $\vec{v}_{1}$, $\vec{v}_{2}$ and $\vec{v}_{1}\neq\vec{v}_{2}$.
@@
 The only viable method in this scenario is to assign an individual coordinate system—i.e., an independent set of synchronized clocks—to each source.
-Now, consider a second secondary source, such as a tangentially moving mirror, interacting with a light beam emitted by
-a stationary source.
-A key challenge in special relativity is that both sources cannot share a common Lorentz frame within a single inertial system.
-The electrodynamic equations must be identical for both the incident and reflected waves.
-Under Einstein’s synchronization, the time for the moving mirror can be determined by introducing a time shift of $xv/c^{2}$ to the first order of approximation.
-In this synchronization scheme, reflection is described using Maxwell’s equations.
-This time shift effectively rotates the radiation wavefront by an angle $v/c$ for both sources.
-Now, consider the electrodynamics of the stationary source.
-In absolute-time coordinatization, an observer describes the light beam from the stationary emitter using the diagonal wave equation.
-However, the time shift introduces a new crossed term in the wave equation, leading to a non-diagonal form.
-Further details on this issue are provided in the next chapter (see Section 4.6 for further details).
+The situation becomes more complex when a secondary source interacts with light emitted by a primary source.
+A fundamental challenge in special relativity is that both sources cannot, in general, be described using a single shared Lorentz frame within a common inertial reference system.
+This issue, and its implications for electrodynamics, is discussed in more detail in the next chapter.
 3.6 A Moving Light Source: Peculiarities of Collinear Geometry
@@
 4 Aberration of Light in an Inertial Frame of Reference
-The effect of light aberration in an inertial frame of reference is typically understood as a change in the direction of light propagation attributed to moving light sources.
-However, even at the first order in $v/c$, describing this phenomenon rigorously is far from straightforward.
-Many textbooks contain incorrect results, reflecting the complexity of the topic.
-To minimize mathematical complexity, we will restrict our discussion to first-order effects.
-This chapter focuses on the physical influence of optical instruments on the measurement of light aberration.
-First, we will demonstrate that when a finite-aperture mirror moves parallel to itself and a plane wave of light falls normally onto it, the reflected light exhibits aberration—manifested as a deviation in energy transport.
-The next problem, of greater practical significance, involves the complementary case of transmission through an aperture in a moving opaque screen, as described by Babinet’s principle.
-A particularly relevant scenario is the transmission of light through the moving end of a telescope barrel, which we will examine in detail.
+In this chapter, we explore several practical applications of the concepts introduced earlier, using the spacetime metric given in Eq.
+(11).
+Our first topic addresses the phenomenon of light aberration caused by a light source moving perpendicular to the direction of radiation.
+Light aberration in an inertial frame is generally understood as an apparent shift in the direction of light propagation due to the motion of the source.
+While often treated as a first-order effect in $v/c$, a rigorous description is not trivial—even at this level of approximation.
+Many standard textbooks contain inaccurate or oversimplified accounts, underscoring the subtlety of the underlying physics.
+To keep the analysis manageable, we will limit our discussion to first-order effects in $v/c$.
+Our focus will be on the physical influence of optical components on the observed aberration of light.
+We begin by examining the case of a finite-aperture mirror moving parallel to its surface, with an incident plane wave striking it normally.
+We will show that the reflected light exhibits aberration, observable as a deviation in the direction of energy transport.
+We then turn to a more practically significant scenario: the transmission of light through an aperture in a moving opaque screen, invoking Babinet’s principle.
+A particularly important application is the passage of light through the moving entrance of a telescope barrel—a configuration we will analyze in detail.
 4.1 The ”Plane Wave” Emitter
-We will analyze the effect in a simple case to build an intuitive understanding.
-The explanation of aberration is based on modeling a single plane-wave emitter.
-As a simple model of a plane-wave emitter, we use a two-dimensional array of identical coherent elementary sources (dipoles), uniformly distributed on a given ($x-y$) plane $P$.
-We take the elementary sources to start radiating waves simultaneously with respect to the lab reference frame where the plane $P$ is at rest.
-Therefore we have a plane full of sources, oscillating together, with their motion in the plane and all having the same amplitude and phase.
-Let us assume that the elementary sources are oscillating at frequency $\omega$.
-By reducing the spacing between adjacent sources to a value much smaller than the radiation wavelength $\lambda=2\pi c/\omega$, we can approximate this two-dimensional arrangement as an ideal plane-wave emitter.
-The concept of an (infinitely extended) plane wave is widely used in physics due to its well-behaved analytical properties as a solution to Maxwell’s equations.
-However, it is not physically realizable, as such a wave would carry infinite energy.
-In any real-world scenario, one must consider a finite source aperture.
-It is implicitly assumed that the detector measuring the radiation direction is sensitive to energy propagation and that its aperture is sufficiently large compared to the radiation beam size.
-Ultimately, what is commonly referred to as aberration is, in reality, a deviation in the direction of energy transport.
-4.2 A Moving Emitter: Single Frame Description
-Let us consider the case of an emitter lying in the $(x-y)$ plane of an inertial laboratory frame, which is accelerated from rest to a velocity $v$ along the $x$-axis.
+To build intuitive insight into the phenomenon of aberration, we begin by analyzing a simple, idealized case: a single ”plane-wave” emitter.
+We model this emitter as a two-dimensional array of identical, coherent elementary sources (dipoles), uniformly distributed over a fixed (xxyy) plane.
+These sources are assumed to begin radiating simultaneously in the lab frame, where the plane is at rest.
+All dipoles oscillate with the same frequency $\omega$, amplitude, and phase, and their motion is confined to the plane.
+By decreasing the spacing between adjacent sources to a value much smaller than the radiation wavelength, $\lambda=2\pi c/\omega$, this arrangement effectively approximates an ideal plane-wave emitter.
+The resulting wavefronts are planar, and the emitted radiation resembles an ideal plane wave.
+While the concept of an (infinitely extended) plane wave is a useful mathematical abstraction—owing to its simplicity and exact correspondence with solutions to Maxwell’s equations—it is not physically realizable, as such a wave would carry infinite energy.
+In any real scenario, the emitter has a finite aperture.
+It is also assumed that any detector used to measure the radiation direction is sensitive to energy flow (rather than to phase fronts, for instance) and has an aperture large enough compared to the beam size to capture the overall direction of energy propagation.
+What is commonly described as aberration, in this context, is in fact a shift in the direction of energy transport.
+4.2 A Moving Emitter: Galilean Transformations-Based Explanation
+Let us consider the case of an emitter lying in the (xy)(x-y) plane of an inertial laboratory frame, which is accelerated from rest to a velocity $v$ along the $x$-axis.
 Suppose an observer at rest in this inertial frame measures the direction of energy transport.
@@
 The first is the covariant approach, commonly used in the literature, which explains the aberration effect by applying a Lorentz boost to determine how the direction of a light beam depends on the velocity of the source relative to the inertial frame.
-The second is a non-covariant approach, which employs a ”single-frame” description without explicit reference to Lorentz transformations.
+The second is a non-covariant approach, which employs an old kinematics description without explicit reference to Lorentz transformations.
 Both approaches, whether formulated using Einstein’s synchronization convention or an absolute time synchronization scheme, yield the same result in the case of a single moving emitter.
@@
 The effect of aberration is analyzed using a Galilean boost (left) and a Lorentz boost (right).
-Let us examine the electrodynamics of a moving source.
-To explain the phenomenon of radiation in our case of interest, we apply an (active) Galilean boost to describe the uniform translational motion of the source in the inertial lab frame.
-Since Maxwell’s equations are not invariant under Galilean transformations, their form is not preserved.
-The additional terms introduced in the field equations as a result of the Galilean transformation lead to the prediction of the Doppler effect.
 The current approach to the moving emitter problem employs the Fourier transform method.
 In the analysis of linear systems, it is often useful to break down a complex input into simpler components, determine the system’s response to each elementary function, and then superimpose these responses to obtain the total outcome.
-Fourier analysis serves as a fundamental tool for this decomposition  [^11] For a broader discussion on Fourier transform methods in spatial filtering theory or Abbe diffraction theory, we recommend referring to [22]..
-[…]
+Fourier analysis serves as a fundamental tool for this decomposition.
+[^11] For a broader discussion on Fourier transform methods in spatial filtering theory or Abbe diffraction theory, we recommend referring to [22].
+Consider the inverse transform relationship $g(x)=\int_{-\infty}^{\infty}G(K)\exp(iKx)dK$ which express the profile function in terms of its wavenumber spectrum.
+We may regard this expression as a decomposition of the function $g(x)$ into a linear combination (in our case into an integral) of elementary functions, each with specific form $\exp(iKx)$.
+From this, it is clear that the number $G(K)$ is simply a weighting factor that must be applied to the elementary function of wavenumber $K$ to synthesize the desired $g(x)$.
 An emitter with a finite aperture is a kind of active medium that breaks up the radiated beam into a number of diffracted beams of plane waves.
@@
 Let us assume that the current density of the elementary dipole sources varies according to the law $j_{pol}=g(K_{\perp})\sin{({K}_{\perp}x)}$.
-From this, we conclude that the active medium of the emmiter is sinusoidally space-modulated.
+From this, we conclude that the active medium of the emitter is sinusoidally space-modulated.
+Let us examine the electrodynamics of a moving source.
+To explain the phenomenon of radiation in our case of interest, we apply an (active) Galilean boost to describe the uniform translational motion of the source in the inertial lab frame.
+According to equivalence of passive and active boosts, Maxwell’s equations remains valid in the comoving frame.
+The measurement of the electromagnetic field configuration - expressed in terms of coordinates $x$ and $t$ of a measuring device at rest - yields the same result that obtained at the spatial position xvtx-vt at time $t=0$.
+Applying a Galilean boost, we transform the spatial coordinate as $x\to x-vt$, while keeping time unchanged.
+Since Maxwell’s equations are not invariant under Galilean transformations, their form is not preserved.
 Starting from the Galilean transformation and applying partial differentiation, we obtain the wave equation Eq.(12).
@@
 With a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ with the wavenumber vector $\vec{k}$ and the frequency $\omega$ equation Eq.(12) becomes: $(1-v^{2}/c^{2})k_{x}^{2}+2vk_{x}\omega/c^{2}+k_{z}^{2}-\omega^{2}/c^{2}=0$.
-The wavevector $\vec{k}$ is determined by the initial conditions.
-Specifically, we set $k_{x}=K_{\perp}$, where $K_{\perp}$ corresponds to the wavenumber of the sinusoidally space-modulated dipole density, and $k_{z}^{2}=\omega_{i}^{2}/c^{2}$, with $\omega_{i}$ denoting the frequency of the emitted radiation prior to acceleration.
-Given these initial conditions, we treat $\vec{k}$ as the independent variable and express the frequency as a function of $k_{x}$: $\omega=\omega_{i}+\Delta\omega(k_{x})$, where $\Delta\omega(k_{x})$ represents the frequency shift due to motion.
-Substituting this relation into the dispersion equation yields a condition linking the wavenumber and the frequency shift: $\Delta\omega=K_{\perp}v$.
-[…]
-An alternative physical perspective on the Doppler effect for a light beam emitted by a moving source leads to the same conclusion as the one presented above.
-The phase of the wave at the world point $(\vec{r},t)$ is invariant at a change of the reference system.
+The wavevector $\vec{k}$ is determined by the initial conditions before the acceleration and subsequent Galilean transformation.
+Specifically, we set $k^{\prime}_{x}=K_{\perp}$, where $K_{\perp}$ corresponds to the wavenumber of the sinusoidally space-modulated dipole density, and $(k_{z}^{\prime})^{2}=\omega_{i}^{2}/c^{2}-K_{\perp}^{2}$, with $\omega_{i}$ denoting the frequency of the emitted radiation prior to acceleration.
+The phase of the wave at the world point $(\vec{r},t)$ is invariant at a change of the reference system (when the field is zero, everyone measures the field as zero).
 This invariance holds regardless of the specific coordinate transformation used to describe the change in the reference system.
@@
 Consequently, $\vec{k}^{\prime}\cdot\vec{r}^{\prime}-\omega^{\prime}t^{\prime}=\vec{k}\cdot\vec{r}-\omega t$.
-Substituting the Galilean transformation formulae $x^{\prime}=x-vt,t^{\prime}=t$ into the phase equality formula we obtain $\Delta\omega=K_{\perp}v$.
-This frequency shift matches the result obtained directly from the dispersion relation, as expected.
+Substituting the Galilean transformation formulae $x^{\prime}=x-vt,t^{\prime}=t$ into the phase equality formula we obtain $\omega=\omega_{i}+K_{\perp}v$.
+This transformation completes with the invariance of wavenumber vector $k_{x}=K_{\perp}$, $k_{z}^{2}=\omega_{i}^{2}/c^{2}-K_{\perp}^{2}$.
+Substituting these into dispersion relation confirms that it remains satisfied, as expected.
 Let us first remind the reader that the velocity of waves is typically defined in terms of the phase difference between oscillations observed at two different points in a free plane wave.
@@
 These equations indicate that a light beam with a finite transverse size propagates along the $x$-direction with a group velocity $d\omega/dk_{x}=v$.
+[…]
 4.3 A Moving Emitter: Lorentz Transformations-Based Explanation
-Another clear way to explain the aberration of light from a moving source is by using a Lorentz boost, which accounts for the uniform translational motion of the source in the laboratory frame.
-Maxwell’s equations are invariant under Lorentz transformations; however, the Lorentz boost introduces a time coordinate transformation of the form $t^{\prime}=t-xv/c^{2}$, which modifies the timing of events across space.
-As a result, features such as the maximum of the electric field reach different positions in the lab frame at different times.
-For example, the field maximum at a position $x$ along the $x$-axis appears delayed compared to the maximum at $x=0$, with a time shift given by $\Delta t=xv/c^{2}$.
-This spatially dependent time delay effectively causes a tilt in the phase front of the radiation.
+An alternative and insightful approach to understanding the aberration of light from a moving source is through Lorentz coordinatization.
+This method provides a relativistically consistent explanation of the phenomenon within the framework of Maxwell’s equations.
+This explanation utilizes a clock-re-synchronization procedure.
+The time coordinate $t_{L}$ in the lab frame, under Lorentz coordinatization, is obtained by introducing a time offset of the form given in Eq.
+(13).
+Although Maxwell’s equations are invariant under Lorentz transformations, the Lorentz boost introduces a transformation of the time coordinate of the form $t^{\prime}=t-xv/c^{2}$.
+This spatially varying time shift introduces a delay in the emission of radiation across the source plane, effectively tilting the phase front of the emitted wave.
 To first order in $v/c$, this corresponds to a rotation of the phase front by an angle approximately equal to $v/c$.
-According to Maxwell’s electrodynamics, coherent radiation is always emitted perpendicular to the radiation phase front, as Maxwell’s equations exhibit no intrinsic anisotropy.
-Thus, when the uniform motion of the source is analyzed via Lorentz transformations, the aberration of light can be understood in the framework of relativistic kinematics.
-In this description, an additional phase chirp, $d\phi/dx=k_{x}=v\omega_{i}/c^{2}$, emerges, causing the array of identical elementary sources in the moving emitter to acquire different phases.
-Consequently, the plane wavefront undergoes a rotation under the Lorentz transformation, resulting in a radiated light beam propagating at an angle $v/c$.
-This yields the well-known phenomenon of aberration of light (Fig.
+This rotation manifests physically in the fact that features of the wave, such as the electric field maximum, reach different spatial positions at different times in the lab frame.
+For instance, the field maximum at position $x$ along the $x$-axis is delayed relative to the field maximum at $x=0$, with a time shift given by $\Delta t=xv/c^{2}$.
+In accordance with Maxwell’s electrodynamics, coherent radiation is always emitted perpendicular to the phase front.
+Since Maxwell’s equations contain no intrinsic directional bias, the tilt in the wavefront—induced by the Lorentz time transformation—results in a corresponding shift in the direction of energy propagation.
+Within this relativistic description, the moving emitter exhibits a phase chirp across its aperture, given by $d\phi/dx=k_{x}=v\omega_{i}/c^{2}$, which means that the individual elementary sources now oscillate with position-dependent phase.
+As a result, the wavefront undergoes a rotation, and the emitted radiation propagates at an angle $\theta_{a}=v/c$, capturing the essence of the aberration effect (see Fig.
 1, right).
-We now consider the group velocity of the radiated beam.
-For a plane wave of the form $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$, the dispersion equation in Maxwell’s electrodynamics simplifies to $k_{z}^{2}+k_{x}^{2}-\omega^{2}/c^{2}=0$.
+We now examine the group velocity of the emitted beam.
+For a plane wave described by $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$, the dispersion relation in vacuum, derived from Maxwell’s equations, is $k_{z}^{2}+k_{x}^{2}-\omega^{2}/c^{2}=0$.
+In special relativity, the four quantities $(\omega,k_{x},k_{y},k_{z})$ form a four-vector and transform accordingly under Lorentz transformations.
+We will restrict our analysis to terms up to first order in $v/c$.
 Using the initial conditions and the Lorentz transformation, we obtain the following expression: $\omega=(\omega_{i}+vK_{\perp})$, $k_{z}=\omega_{i}/c$, $k_{x}=(v\omega_{i}/c^{2}+K_{\perp})$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated current density, $\omega_{i}$ is the frequency of the emitter radiation before the acceleration.
-Substituting these expressions into the dispersion equation, we verify that it is indeed satisfied, as expected.
+Substituting these into the dispersion relation confirms that it remains satisfied, as expected.
 As a consequence of the Doppler effect in Lorentz coordinatization, we observe an angular frequency dispersion in the light waves emitted from a moving source with a finite aperture.
@@
 This equation implies that a light beam with a finite transverse extent propagating along the $x$-direction has a group velocity given by $d\omega/dk_{x}=v$.
-It is worth examining the significance of the fact that two different approaches—covariant and non-covariant—lead to the same group velocity.
-The key insight is that both methods accurately describe the same physical reality.
-Since the group velocity has an objective, convention-independent meaning, both approaches must necessarily yield the same physical results.
-The distinction between absolute time coordinatization and Lorentz coordinatization is particularly intriguing.
-In Chapter 3, we explored how one can transform from absolute time coordinatization to Lorentz coordinatization.
-Specifically, manipulations involving the rule-clock structure in the lab frame can be understood as a simple change of the time variable, given by the transformation $t\to t+xv/c^{2}$.
-When combined with the Galilean transformation, this variable change ultimately leads to the Lorentz transformation in the context of absolute time coordinatization in an inertial frame.
-However, this change of variables has no intrinsic physical significance.
-Notably, there is a direct link between this time shift and the issue of clock synchronization.
-Importantly, the convention-independent results of calculations remain unchanged in the transformed variables.
-Consequently, there is no need to convert the results of electrodynamics problem-solving back into the original (3+1) variables.
-From a pedagogical perspective, studying electrodynamics through variable transformations provides valuable insight.
-It is worth noting that the lack of a dynamical explanation for wavefront rotation in Lorentz coordinatization has been a point of concern for some physicists.
-However, as discussed in Chapter 3, a useful way to conceptualize wavefront rotation is to view it as a consequence of transitioning to a new time variable within the framework of Galilean (”single-frame”) electrodynamics.
+It is important to recognize the significance of the fact that both covariant (Lorentz-based) and non-covariant (absolute time-based) approaches yield the same group velocity.
+This agreement underscores the physical objectivity of the result: group velocity is a measurable, convention-independent quantity, and must therefore remain invariant regardless of the chosen coordinatization.
+The distinction between absolute time coordinatization and Lorentz coordinatization, as discussed in Chapter 3, is particularly illuminating.
+For non-relativistic velocities, the transition between these two frameworks can be understood as a simple redefinition of the time variable: $t\to t+xv/c^{2}$.
+When combined with the Galilean transformation, this leads naturally to the Lorentz transformation, even within the context of an absolute-time formalism.
+However, this change of variables has no intrinsic physical significance—it represents merely a reparametrization of time, not a real physical effect.
+This coordinate transformation is closely tied to the issue of clock synchronization.
+Crucially, the physical content of electrodynamics remains unchanged under such transformations.
+As a result, once a solution is obtained in the transformed variables, there is no need to revert to the original (3+1) variables; the results remain valid and physically meaningful.
+It is commonly assumed that the theory of relativity can be applied to physical processes without a detailed understanding of the clock synchronization procedure.
+Most textbooks suggest that an operational interpretation of Lorentz coordinatization within a single inertial frame is unnecessary.
+Nevertheless, distinguishing between absolute-time and Lorentz coordinatization (from an operational standpoint) deepens conceptual understanding—even in the simple case of a single moving source.
+While some physicists have expressed concern about the lack of a ”dynamical” explanation for wavefront rotation in Lorentz coordinates, Chapter 3 offers a compelling perspective: the rotation arises naturally from the transformation of the time coordinate, which alters the simultaneity structure of spacetime events.
+Within the absolute-time framework, this provides an intuitive—though coordinate-dependent—explanation of the aberration phenomenon.
 4.4 Reflection from a Mirror Moving Parallel to Its Surface
-Up to this point, the previous two sections have not introduced any fundamentally new results.
-As discussed earlier, when the sources are independent, adopting Lorentz coordinates poses no difficulty—each source can be assigned its own coordinate system, complete with an independent set of clocks.
-However, the situation becomes more complex when a secondary source interacts with light emitted by a primary source.
-In such cases, it becomes necessary to apply electrodynamics within an absolute time framework.
-With the results from Chapter 3 now at our disposal, we are in a position to examine the intriguing phenomenon of radiation from moving sources.
+Until now, the previous two sections have not introduced any fundamentally new results.
+As previously discussed, when the sources are independent, adopting Lorentz coordinates presents no difficulty—each source can be assigned its own coordinate system, along with an independent set of clocks.
+The situation becomes more complex, however, when a secondary source interacts with light emitted by a primary source within a single inertial frame.
+This challenge in special relativity can be effectively addressed using an electrodynamics-based approach framed within an absolute time perspective using metric Eq.
+(11).
+As noted earlier, the metric in Eq.
+(11) has not been previously reported in the literature.
+Existing studies generally assume—incorrectly—that both sources share the same Minkowski metric and are governed by Maxwell’s equations within a single inertial (laboratory) frame.
+This assumption is often treated as self-evident.
+In effect, the conventional approach relies on a hidden premise: that the principle of equivalence between active and passive boosts does not holds within a single inertial frame.
+Our formulation challenges this assumption.
+The agreement between our theoretical framework and experimental observations lends strong support to our revised formulation of special relativity, which incorporates the metric of Eq.
+(11) (see Section 4.14 for further discussion).
+With the results of Chapter 3 now established, we are prepared to examine the intriguing phenomenon of radiation from moving sources.
 We begin by analyzing the intensity of light reflected when a stationary source illuminates a moving mirror.
@@
 This assumption, frequently found in textbooks (see, for example, [9] and [23]), is, however, incorrect.
-To examine this claim, we analyze the reasoning found in standard texts 121212 Many sources claim that no aberration occurs when light reflects off a mirror moving parallel to its surface.
-For example, Sommerfeld [23] states: ”For a mirror moving tangentially to its surface, the law of reflection that holds for a stationary mirror is preserved.”.
-Textbooks typically analyze reflection using two Lorentz reference frames.
+To examine this claim, we analyze the reasoning found in standard texts.
+121212 Many sources claim that no aberration occurs when light reflects off a mirror moving parallel to its surface.
+For example, Sommerfeld [23] states: ”For a mirror moving tangentially to its surface, the law of reflection that holds for a stationary mirror is preserved.” Textbooks typically analyze reflection using two Lorentz reference frames.
 The fixed (lab) frame $S$ is at rest relative to the plane-wave emitter, while the moving frame $S^{\prime}$ travels at velocity $v$, in which the mirror is at rest.
-Both frames use a Cartesian coordinate system with the $x-y$ plane tangent to the mirror’s surface.
+Both frames use a Cartesian coordinate system with the xyx-y plane tangent to the mirror’s surface.
 The $x$-axis aligns with the mirror’s velocity $v$.
@@
 Consequently, the solution naturally involves light beams of different frequencies, incorporating the Doppler effect.
-For a tangentially moving mirror with a finite aperture, angular frequency dispersion—an effect of first order in $v/c$—cannot be neglected.
+For a tangentially moving mirror with a finite aperture, angular frequency dispersion—an effect of first order in v/cv/c—cannot be neglected.
 We have already emphasized that Chapter 3 is arguably the most important part of this book.
@@
 The (diagonal) wave equation for the moving mirror in the atmospheric frame is identical to that of an emitter at rest.
-Textbooks often draw important parallels between sound and light aberration in the first-order approximation  [^13] We should make one additional remark regarding two common errors when solving problems related to reflection from a mirror moving tangentially to its surface.
+Textbooks often draw important parallels between sound and light aberration in the first-order approximation.
+[^13] We should make one additional remark regarding two common errors when solving problems related to reflection from a mirror moving tangentially to its surface.
 These errors are not independent.
 Textbooks typically assume that the metric of the moving mirror is diagonal, similar to the metric of a ’plane wave’ emitter at rest, and that the group velocity equals the phase velocity.
-As a result, regardless of the mirror’s size, there is no deviation in the energy transport of the reflected light beam..
+As a result, regardless of the mirror’s size, there is no deviation in the energy transport of the reflected light beam.
 This, however, is a misconception.
@@
 Under the Bragg condition for small angles, the scattering angle is approximately given by $\theta\approx K_{\perp}/k_{i}$.
-Let us analyze the reflection of light from a mirror moving parallel to its surface, relying solely on relativistic kinematics.
-After applying a Galilean boost in the case of absolute time coordinatization, we employed the metric given by Eq.
-11, which describes the electrodynamics of a moving light source as observed from an inertial frame.
+Let us analyze the reflection of light from a mirror moving tangentially (i.e., parallel to its surface), relying solely on relativistic kinematics.
+To describe this scenario, we adopt the framework of absolute time coordinatization and apply a Galilean boost.
+In this context, we utilize the metric given by Eq.
+11, which characterizes the electrodynamics of a moving light source as observed from an inertial frame.
 From the structure of Eq.
-11, it becomes evident that the electromagnetic field configuration — expressed in terms of the coordinates $x$ and $t$ of a detector at rest in the inertial frame—coincides with the field configuration at position $x-vt$ at time $t=0$.
-Using a Fourier approach, we showed that the problem could be reduced to the case of reflection from a tangentially moving grating.
-We recall that with a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ with the wavenumber vector $\vec{k}$ and the frequency $\omega$ dispersion equation in the absolute time coordinatization becomes: $(1-v^{2}/c^{2})k_{x}^{2}+2vk_{x}\omega/c^{2}+k_{z}^{2}-\omega^{2}/c^{2}=0$.
+11, it becomes evident that the electromagnetic field configuration—expressed in terms of the coordinates $x$ and $t$ of a detector at rest in the inertial frame—coincides with the field configuration at position xvtx-vt at time $t=0$.
+Using a Fourier analysis, we show that the problem can be reduced to that of reflection from a tangentially moving diffraction grating.
+Consider a plane wave of the form $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ where $\vec{k}$ is the wavevector and $\omega$ the angular frequency.
+In the absolute time coordinatization, the dispersion relation takes the form: $(1-v^{2}/c^{2})k_{x}^{2}+2vk_{x}\omega/c^{2}+k_{z}^{2}-\omega^{2}/c^{2}=0$.
 Within a single inertial frame, the equivalence between active and passive Galilean boosts implies that translating the system in one direction is indistinguishable from shifting the coordinate system in the opposite direction by the same amount.
-According to this equivalence, Maxwell’s equations remain valid in the comoving coordinate system.
-The wavenumber vector of the diffraced wave in the comoving coordinates is fixed by initial conditions.
-In fact, $k_{z}=\omega_{i}/c$, $k_{x}=K_{\perp}$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated reflectance.
-Using the initial conditions and the Galilean transformation back to the lab frame, we obtain the following expression $\omega=\omega_{i}+K_{\perp}v$, $k_{z}=\omega_{i}/c$, $k_{x}=K_{\perp}$.
-Substituting there expressions into dispersion equation, we verify that it is satisfied, as expected.
-We find the requirement that the wavenumber $K_{\perp}$ and the frequency change $\Delta\omega$ are related by $\Delta\omega=K_{\perp}v$.
-This phenomenon is interpreted as a Doppler effect, as the reflection occurs from a moving, rather than stationary, system.
-As we consider aberration effects only to first order in $v/c$, and under the assumption $cK_{\perp}/\omega_{i}\ll v/c$, no second-order terms appear in the expression for $\Delta\omega$.
+This equivalence ensures that Maxwell’s equations retain their form in the comoving coordinate system.
+In the comoving frame, the wavevector of the diffracted wave is determined by the initial conditions.
+Specifically, we have: $k_{z}^{2}=\omega_{i}^{2}/c^{2}-k_{x}^{2}$, $k_{x}=K_{\perp}$, where $K_{\perp}$ denotes the spatial frequency of the sinusoidally space-modulated reflectance.
+Transforming back to the lab frame using a Galilean transformation, we find that the frequency shifts according to: $\omega=\omega_{i}+K_{\perp}v$, while the components of the wavevector remain unchanged.
+Substituting these expressions into the dispersion relation yields: $(1-v^{2}/c^{2})K_{\perp}^{2}+2vK_{\perp}(\omega_{i}+K_{\perp}v)/c^{2}+\omega_{i}^{2}/c^{2}-K_{\perp}^{2}-(\omega_{i}+K_{\perp}v)^{2}/c^{2}=0$, which is readily verified to be satisfied, as expected.
+This analysis leads to the conclusion that the frequency shift $\Delta\omega$ and the spatial frequency $K_{\perp}$ are related by: $\Delta\omega=K_{\perp}v$.
+This relation is recognized as a manifestation of the Doppler effect, arising from reflection off a moving (as opposed to stationary) surface.
 A crucial aspect of describing a ”single inertial frame” cannot be overstated.
@@
 To clarify the fundamental physical principles, we examined a simple case.
-[…]
+One of the consequences of the Doppler effect is the angular frequency dispersion of light waves reflected from a moving mirror with a finite aperture.
+If $\vec{n}=\vec{k}/|\vec{k}|$ denotes a unit vector in the direction of the wave normal, and $\vec{v}$ is the mirror velocity vector relative to the lab frame, we get the equation $\omega_{s}=\omega_{i}(1+\vec{n}\cdot\vec{v}/c)=\omega_{i}+(\omega_{i}v/c)\cos\theta$.
+The Doppler effect is responsible for angular frequency dispersion to the first order of $v/c$ even when $\vec{n}\cdot\vec{v}=0$ (i.e when $\cos\theta=0$).
+In fact, $\d\omega_{s}/d\theta=-(\omega_{i}v/c)\sin\theta=-\omega_{i}v/c$ at $\theta=\pi/2$.
+We can rewrite this equation in a different way.
+The differential of the scattered angle is given by $d\theta=-dk_{x}/k_{i}$.
+With the help of this relation and account for that $k_{i}=\omega_{i}/c$ we have $d\omega_{s}/dk_{x}=v$.
 Figure 5: Tangentially moving mirror with a small aperture at normal incidence. The effect of aberration is analyzed using absolute time coordinatization.
@@
 This phenomenon can be understood as a straightforward consequence of the Doppler effect [^14] The Doppler effect serves merely as an intermediate mathematical description; when the full calculation is performed, only the geometric (aberration) effect remains..
-A common misconception in the literature asserts that the direction of propagation in reflection from a tangentially moving mirror is not determined by the normal to the wavefront [^15] A widespread misunderstanding stems from assuming that the orientation of radiation wavefronts has objective significance.
+A common misconception in the literature asserts that the direction of propagation in reflection from a tangentially moving mirror is not determined by the normal to the wavefront.
+[^15] A widespread misunderstanding stems from assuming that the orientation of radiation wavefronts has objective significance.
 As Norton [24] notes: ”One might try to escape the problem by supposing that the direction of propagation is not always given by the normal to the wavefront.
@@
 Whatever may be the merits of such proposals, they are unavailable to some trying to implement a principle of relativity.
-If the direction of propagation of a plane wave is normal to the wavefronts in one inertial frame, then that must be true in all inertial frames.”.
-This misunderstanding arises from a failure to distinguish between convention-dependent and convention-invariant aspects of the theory.
+If the direction of propagation of a plane wave is normal to the wavefronts in one inertial frame, then that must be true in all inertial frames.” This misunderstanding arises from a failure to distinguish between convention-dependent and convention-invariant aspects of the theory.
 The direction of energy transport is an objective, convention-invariant quantity.
@@
 4.6 A Moving Mirror: Lorentz Transformations-Based Explanation
-To maintain a Lorentz coordinate system for a moving mirror in the lab frame, clock resynchronization is necessary.
-This entails a time shift of the form given in Eq.
-13, namely $t\to t+xv/c^{2}$, which is valid to first order in $v/c$.
+Another clear way to explain the aberration of light from a moving mirror is by using a Lorentz boost, which accounts for the uniform translational motion of the mirror in the laboratory frame.
+Maxwell’s equations are invariant under Lorentz transformations; however, the Lorentz boost introduces a space-time coordinates transformation of the form Eq.
+(13).
+To minimize the mathematical complexity of the discussion, we will restrict our analysis to terms up to first order in $v/c$.
+This entails only a time shift of the form $t\to t+xv/c^{2}$.
 This time shift corresponds to an effective rotation of the phase front of the incoming radiation by an angle of $v/c$.
@@
 This implies that the reflected light beam has a group velocity along the $x$-direction given by: $d\omega/dk_{x}=v$, consistent with the prediction from Maxwell’s equations that the group velocity equals the phase velocity.
-It is also insightful to analyze the electrodynamics of an emitter at rest.
+It is also insightful to analyze the electrodynamics of an emitter at rest in the case of Lorentz coordinatization.
 A rotation of the $x$-axis without changing the orientation of the $t$-axis is described by a Galilean transformation $x^{\prime}=x-vt$, $t^{\prime}=t$.
@@
 4.10 Spatiotemporal Transformation of the Transmitted Light Beam
-Suppose the transmitted light pulse propagates in the $x-z$ plane.
+Suppose the transmitted light pulse propagates in the xzx-z plane.
 Our focus is on the space-time intensity distribution within this plane.
@@
 This represents the field immediately behind the moving screen.
-The beam profile, $b(x-vt)$, is the optical replica of the moving aperture.
-Now, consider a screen positioned at rest at a distance $l$ behind the moving aperture.
+The beam profile, b(xvt)b(x-vt), is the optical replica of the moving aperture.
+Now, consider an observer’s screen positioned at rest at a distance $l$ behind the moving aperture.
 For simplicity, we assume that the Fresnel number is large, $N_{F}=D_{h}^{2}\omega/(cl)\gg 1$, allowing us to neglect diffraction effects.
@@
 From this, we conclude that the light spot on the observer’s screen moves with the same velocity $v$ as the moving screen.
-Additionally, we observe that a particular quantity remains unchanged: the combination $x-vt$ is the same for both the spot and the aperture.
+Additionally, we observe that a particular quantity remains unchanged: the combination xvtx-vt is the same for both the spot and the aperture.
 What does this imply?
-A careful examination reveals that events occur simultaneously at two separate locations along the $z$-axis  [^16] Spatiotemporal coupling is typically discussed in the literature in the context of ultrashort laser pulse propagation through a grating monochromator.
+A careful examination reveals that events occur simultaneously at two separate locations along the $z$-axis.
+[^16] Spatiotemporal coupling is typically discussed in the literature in the context of ultrashort laser pulse propagation through a grating monochromator.
 In this scenario, in addition to considering phase fronts, one must also account for planes of constant intensity—known as pulse fronts.
 In a grating monochromator, the different spectral components of the outgoing pulse propagate in different directions.
-The electric field of a pulse with angular dispersion can be expressed in the Fourier domain $[\Delta k_{x},\Delta\omega]$ as $\bar{E}(\Delta k_{x}-p\Delta\omega,\Delta\omega)$.
+The electric field of a pulse with angular dispersion can be expressed in the Fourier domain $\Delta k_{x},\Delta\omega]$ as $\bar{E}(\Delta k_{x}-p\Delta\omega,\Delta\omega)$.
 Its inverse Fourier transform to the space-time domain $[x,t]$ results in $E(x,t+px)$, representing a pulse with a pulse-front tilt.
 In our case of interest, we consider light diffracted by a tangentially moving set of gratings, where the Doppler effect induces frequency dispersion given by $d\omega/dk_{x}=v$.
-Consequently, the spatiotemporal coupling due to light transmission through a hole in a moving screen differs significantly from the conventional pulse-front tilt distortion..
+Consequently, the spatiotemporal coupling due to light transmission through a hole in a moving screen differs significantly from the conventional pulse-front tilt distortion.
 4.11 Applicability of Ray Optics
@@
 Monochromatic light falls normally on the plate and generates an oblique ray.” This oblique effect is demonstrated in Fig.
-10..
+10.
 Let us examine what happens, according to ray optics, in our case of interest.
@@
 (b) A point-like source (or, more generally, a spatially incoherent source)
+[…]
 The diffraction of a source field is typically categorized into Fresnel (near-zone) and Fraunhofer (far-zone) diffraction.
@@
 In the far zone, both types of sources effectively produce a plane wave in front of the pupil detection system.
-This implies that, in the Fraunhofer zone, the instrument always exerts a physical influence on the measurement of light aberration.
-[…]
 Figure 12: Aberration of light in an inertial frame of reference.
@@
 In this case, the aberration of light phenomenon does not occur.
-Figure 13: Diagram of a modified source of the ”plane wave” emitter type. A point source of light is placed in the front focal plane of a lens.
-One of the key conclusions of the discussion above (see Fig.
-11) is that the aberration of a moving point source is absent in an inertial frame (Fig.
-12).
-This leads to several remarkable effects, stemming from the fact that the transmitted light beam does not carry information about the motion of the point source.
-In fact, this principle is central to the binary star paradox discussed in Chapter 6.
-To further explore the aberration of light in an inertial frame, it is useful to consider a slightly modified source—a ”plane wave” emitter.
-We adopt the model described in Section 4.2 because it is both relatively simple and sufficiently general to serve as a prototype for understanding aberration phenomena.
-While no one has performed all the ”thought experiments” exactly as described here, the outcomes can be confidently predicted based on the laws of special relativity, which are themselves grounded in experimental evidence.
-Figure 13 illustrates the modified ”plane wave” emitter setup.
-Though it appears more complex at first glance, it is practically feasible to construct.
-The setup consists of a point source of light positioned in the front focal plane of a lens.
-The source’s dimensions are assumed to be on the order of $\lambda$, where $\lambda$ is the wavelength of optical radiation.
+Let us now consider the case of a moving point source.
+Specifically, we examine a setup where the screen is at rest, while the point source moves tangentially.
+A key takeaway from the earlier discussion is that no aberration is observed for a large-aperture plane wave emitter moving tangentially in an inertial frame (see Fig.
+11).
+Remarkably, the same conclusion holds for a point source.
+This stems from the linearity of Maxwell’s equations: the governing laws of electrodynamics are linear differential equations.
+As discussed previously, the most effective approach to analyzing a moving emitter is via the Fourier transform method.
+A point source can be decomposed into a superposition of spatial Fourier components—each corresponding to a radiated plane wave.
+The current density associated with these components can be modeled as $j_{pol}=j_{0}\sin{({\vec{K}}_{\perp}\cdot\vec{r})}$.
+Each elementary dipole source radiates a plane wave of the form $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$, where the wavevector $\vec{k}$ is determined by the initial spatial modulation.
+Near normal incidence, we set $k_{x}=K_{\perp}$, where $K_{\perp}$ is the transverse spatial frequency corresponding to the modulated dipole density.
+For small angles, the incidence angle is approximately given by $\theta\simeq K_{\perp}/k_{z}\simeq K_{\perp}c/\omega$.
+[…]
+This leads to a remarkable conclusion: the transmitted beam carries no information about the source’s tangential motion.
+This principle lies at the heart of the binary star paradox discussed in Chapter 6.
 4.14 Experimental Test: Reflection from a Moving Grating
@@
 The effect of sound aberration is best understood using a model of a single plane-wave emitter.
-Specifically, we consider a two-dimensional array of identical, coherent elementary sources uniformly distributed on a given $(x-y)$ plane $P$, all emitting waves simultaneously.
+Specifically, we consider a two-dimensional array of identical, coherent elementary sources uniformly distributed on a given (xy)(x-y) plane $P$, all emitting waves simultaneously.
 We analyze a scenario in which a finite-aperture mirror moves tangentially to its surface.
@@
 5.1 Absolute Time Coordinatization in Accelerated Systems
-We analyze the problem of light aberration in accelerated systems within the framework of special relativity.
-Specifically, we demonstrate that explaining optical effects in a rotating frame of reference does not require modifying special relativity or invoking general relativity.
-Instead, a strict adherence to special relativity suffices.
-To express electrodynamics laws in an accelerated frame, one must take an additional step by defining the metric of the non-inertial frame.
-Importantly, the metric tensor must remain a continuous function, leading to the concept of smoothly tailoring the metric tensor.
-This problem in special relativity can be effectively addressed using an approach based on absolute time coordinatization.
-At the end of this section, we outline the mathematical extension of special relativity to accelerating systems.
-However, in some cases, direct physical arguments can provide quicker insights while remaining consistent with rigorous mathematical derivations.
-Suppose the initial inertial system $S$ is at rest with respect to the fixed stars, and a system $S_{n}$ (along with an observer and their measuring instruments) is accelerated from rest in $S$ to a velocity $\vec{v}$ along the $x$-axis.
-In accelerated systems, the concept of maintaining absolute simultaneity remains logically consistent with the smooth tailoring of the metric tensor.
-The synchronization method involves preserving, without modification, the same set of uniformly synchronized clocks used when $S_{n}$ was initially at rest.
-It is well known that during acceleration (relative to the fixed stars), Einstein’s clock synchronization procedure cannot be applied.
-Consequently, by the time $S_{n}$ transitions to uniform motion, the metric in the accelerated reference frame will have acquired a non-diagonal form.
-We begin by considering the metric as the true measure of space-time interval for an accelerated observer $S_{n}$ with coordinates $(t_{n},x_{n},y_{n},z_{n})$.
-To transform to the coordinates of an inertial observer $S$ moving with velocity $-v$ relative to $S_{n}$, we apply the inverse Galilean transformation $x=x_{n}+vt$, while leaving the time coordinate unchanged, $t_{n}=t$.
-Substituting this into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$, we obtain the transformed metric:
+We investigate the phenomenon of light aberration in accelerated systems within the framework of special relativity.
+In particular, we demonstrate that explaining optical effects in rotating frames does not require modifying special relativity or invoking general relativity.
+A rigorous application of special relativity is sufficient.
+However, to express electrodynamics in a non-inertial frame, one must take the additional step of defining the metric of that frame.
+Crucially, the metric tensor must remain a continuous function, leading naturally to the concept of smoothly tailoring the metric.
+This issue in special relativity can be effectively addressed using an approach based on absolute time coordinatization.
+Toward the end of this section, we outline how special relativity may be mathematically extended to describe accelerating systems.
+In some situations, however, direct physical reasoning can yield rapid insights while remaining consistent with formal derivations.
+Consider an inertial frame $S$, assumed to be at rest relative to the fixed stars.
+A second frame, $S_{n}$, along with an observer and measuring instruments, is accelerated from rest in $S$ to a velocity $v$ along the $x$-axis.
+In such accelerated systems, the notion of absolute simultaneity remains logically consistent with the smooth tailoring of the metric tensor.
+Synchronization in this context involves preserving the same uniformly synchronized clocks that were used when $S_{n}$ was initially at rest.
+It is well known that during acceleration relative to the fixed stars, Einstein’s synchronization procedure cannot be directly applied.
+As a result, by the time $S_{n}$ reaches uniform motion, the metric in its reference frame will generally acquire a non-diagonal form.
+We begin by treating the metric as the true measure of spacetime intervals for an accelerated observer in frame $S_{n}$, with coordinates $(t_{n},x_{n},y_{n},z_{n})$.
+To express this in the coordinates of an inertial observer in $S$, moving at velocity v-v relative to $S_{n}$, we apply the inverse Galilean transformation $x=x_{n}+vt$, keeping the time coordinate unchanged: $t_{n}=t$.
+Substituting into the Minkowski metric, $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$, we obtain the transformed (Langevin) metric:
 $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}~.$ (15)
+It is crucial to note that Langevin metric, in fact, reflects measurements performed by an inertial observer using the Minkowski metric, subsequently re-expressed via the transformation $x\to x_{n}+vt$, $t\to t_{n}$.
 This metric describes the electrodynamics of the accelerated light source with the viewpoint of the accelerated observer measurements as viewing this of the inertial observer.
@@
 This result represents the Langevin metric, derived by matching the metric tensors of the accelerated frame and the initial inertial frame.
-Considering the Langevin metric, we conclude that the length of a physical rod $dx$ and the proper time interval $dt$ in the inertial frame will coincide with the coordinate distance and coordinate time in the accelerated frame, i.e., $dx_{n}=dx$, $dt_{n}=dt$.
-(For further details, see Chapter 12).
+To interpret Eq.
+(15) physically, recall that a new frame of reference can always be introduced via a passive coordinate transformation—a mere relabeling of coordinates used to describe physical events.
+For example, an observer in the inertial frame can introduce a comoving coordinate system to analyze radiation from a moving source.
+In this system, fields are expressed as functions of $(t_{n},x_{n},y_{n},z_{n})$, which relate to the original coordinates $(t,x,y,z)$ through a passive Galilean transformation.
+From the structure of the Langevin metric, it is evident that the electromagnetic field measured at position $x_{n}$ and time $t_{n}$ by a device at rest in the accelerated frame will match the field measured by an identical device at rest in the inertial frame at position $x=x_{n}+vt$ and time $t=t_{n}$.
+Substituting this coordinate relationship into the Minkowski metric of the inertial observer leads directly to Eq.
+(15).
+In describing physical phenomena in an accelerated frame, one must distinguish between coordinate quantities and physical quantities.
+The Langevin metric allows us to determine how these relate.
+For instance, the length of a physical rod $dx$ and the proper time interval $dt$ in the inertial frame coincide with the coordinate length and time in the accelerated frame: $dx_{n}=dx$, $dt_{n}=dt$.
+For a more detailed discussion on how to relate coordinate and physical quantities in non-inertial frames, refer to Chapter 12.
 Let us now extend the mathematical framework of the special theory of relativity to include accelerated motion.
@@
 We emphasize that $dx(t)/dt$ represents the velocity of the accelerating frame as observed from the inertial reference frame.
-Applying the inverse Galilean transformation, we have $dt_{n}=dt$, $dx=dx_{n}+atdt/\sqrt{1+a^{2}t^{2}/c^{2}}$.
+Applying the inverse Galilean transformation, we have $dt=dt_{n}$, $dx=dx_{n}+at_{n}dt_{n}/\sqrt{1+a^{2}t_{n}^{2}/c^{2}}$.
 With this transformation, the metric in the uniformly accelerating observer’s frame takes the form:
@@
 If $ds$ is the infinitesimal displacement along the world line of a ray of light, then $ds^{2}=0$ and we obtain $c^{2}=(dx/dt)^{2}$.
-In the accelerated reference system, since $x_{n}=x-vt$ and $t=t_{n}$, this expression takes the form $c^{2}=(dx_{n}/dt_{n}+v)^{2}$, as seen by setting $ds^{2}=0$ in Eq.
+In the accelerated reference system, since $x_{n}=x-vt$ and $t_{n}=t$, this expression takes the form $c^{2}=(dx_{n}/dt_{n}+v)^{2}$, as seen by setting $ds^{2}=0$ in Eq.
 (15).
@@
 This transformation ensures that, following acceleration, the emitted light beam’s wavefront remains perpendicular to the vertical direction $z_{n}$, as illustrated in Fig.
-14.
+13.
 We will explore this topic further in Section 5.5.
@@
 Due to the inherent asymmetry between inertial and accelerated frames, an intriguing prediction arises in the theory of light aberration.
-Specifically, if an opaque screen with a hole is initially at rest relative to the fixed stars and then set into motion, the apparent angular position of a ”plane-wave” emitter observed through the aperture in the accelerated frame will shift by an angle $-v/c$.
+Specifically, if an opaque screen with a hole is initially at rest relative to the fixed stars and then set into motion, the apparent angular position of a ”plane-wave” emitter observed through the aperture in the accelerated frame will shift by an angle v/c-v/c.
 This effect is illustrated in Fig.
-14.
+13.
 Notably, the source of this asymmetry is not the relative acceleration between the two frames but rather the difference in their individual acceleration histories with respect to the fixed stars.
@@
 It is important to emphasize that the aberration of the light beam transmitted through the small aperture hole also remains independent of the emitter’s motion in the accelerated frame $S_{n}$ (see Fig.
-15).
+14).
 This is because the cross-term in the metric (Eq.
@@
 Acceleration affects the propagation of light.
-The coordinate velocity of light parallel to the $x$-axis is given by $dx/dt=c-v$ in the positive direction, and $dx/dt=-c-v$ in the negative direction.
+The coordinate velocity of light parallel to the $x_{n}$-axis is given by $dx_{n}/dt_{n}=c-v$ in the positive direction, and $dx_{n}/dt_{n}=-c-v$ in the negative direction.
 We now consider the transmission through the hole in the opaque screen in the accelerated frame $S_{n}$, Fig.
-15.
+14.
 We describe the aberration of light based on the observations made by an observer in the same accelerated frame as the screen.
@@
 (17).
-Consider a transmitted plane wave of the form $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$.
-With a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ with the wavenumber vector $\vec{k}$ and the frequency $\omega$ equation Eq.(17) becomes: $(1-v^{2}/c^{2})k_{x}^{2}-2vk_{x}\omega/c^{2}+k_{z}^{2}-\omega^{2}/c^{2}=0$.
-The wavenumber vector of the transmitted plane wave is fixed by initial conditions.
-Suppose the wavefront of the emitted light beam is initially perpendicular to the vertical direction before acceleration.
-Then, we have $k_{z}=-\omega_{i}/c$, $k_{x}=K_{\perp}$, where $K_{\perp}$ is the wavenumber accosiated with space-modulated transmittance.
-[…]
+Consider a transmitted plane wave of the form $\exp(i\vec{k}_{n}\cdot\vec{r}_{n}-i\omega_{n}t_{n})$.
+With a plane wave $\exp(i\vec{k}_{n}\cdot\vec{r}_{n}-i\omega_{n}t_{n})$ with the wavenumber vector $\vec{k}_{n}$ and the frequency $\omega_{n}$ equation Eq.(17) becomes: $(1-v^{2}/c^{2})(k_{n})_{x}^{2}-2v(k_{n})_{x}\omega_{n}/c^{2}+(k_{n})_{z}^{2}-\omega_{n}^{2}/c^{2}=0$.
+The wavevector $\vec{k}_{n}$ is determined by the initial conditions before the acceleration.
+Transforming to the accelerated frame using Galilean transformation $x=x_{n}+vt$, $t=t_{n}$ we find that the frequency shifts according to: $\omega_{n}=\omega_{i}-K_{\perp}v$, while the components of wavevector remain unchanged.
+Substituting these into dispersion relation confirms that it remains satisfied, as expected.
 In our example, plane waves with different wavenumber vectors propagate away from the screen at different frequencies.
@@
 This implies that a transmitted light beam with a finite transverse size moves along the $x_{n}$-direction with a group velocity of $d\omega/dk_{x}=-v$ (see Fig.
-15).
+14).
 Now, let us address an apparent paradox.
 We derived the electrodynamics equations in an accelerated frame using the Galilean transformation of Maxwell’s equations.
+From the Galilean transformation $x_{n}=x-vt,~t_{n}=t$, after partial differentiation, one obtains $\partial/{\partial t}=\partial/{\partial t_{n}}-v\partial/{\partial x_{n}}$, $\partial/{\partial x}=\partial/{\partial x_{n}}$.
 Intuitively, one might expect everything to be at rest in the accelerated frame.
@@
 To achieve this, it is sufficient to relate the coordinates and time of the accelerated observer to those of the inertial observer using a Galilean boost.
-The electrodynamics equations in an accelerated frame can be obtained by applying a Galilean transformation (with velocity $-v$) to Maxwell’s equations.
+The electrodynamics equations in an accelerated frame can be obtained by applying a Galilean transformation (with velocity v-v) to Maxwell’s equations.
 The electromagnetic fields $\vec{E}_{n}$ and $\vec{B}_{n}$ observed in the accelerated frame differ from the fields $\vec{E}$ and $\vec{B}$ in the inertial frame prior to the active Galilean boost.
@@
 This results in a Poynting vector, $c\vec{E}_{n}\times\vec{B}_{n}/(4\pi)=[-(\vec{e}_{y}\times\vec{e}_{x})-(\vec{e}_{y}\times\vec{e}_{z})v/c]cE^{2}/(4\pi)=[\vec{e}_{z}-\vec{e}_{x}(v/c)]cE^{2}/(4\pi)$.
-Thus, the radiated beam propagates at an angle $-v/c$ with respect to the $z_{n}$-axis, leading to the phenomenon of aberration of light in the accelerated frame.
+Thus, the radiated beam propagates at an angle v/c-v/c with respect to the $z_{n}$-axis, leading to the phenomenon of aberration of light in the accelerated frame.
 We find that the energy transport direction remains the same whether analyzed using electrodynamics equations in an accelerated frame or the Galilean transformation law for electromagnetic fields.
@@
 In this frame, the plane wavefront undergoes rotation following the re-synchronization of the accelerated clock.
-According to Maxwell’s electrodynamics, the transmitted light beam propagates at an angle of $-v/c$, resulting in the aberration of light.
+According to Maxwell’s electrodynamics, the transmitted light beam propagates at an angle of v/c-v/c, resulting in the aberration of light.
 The time coordinate $t^{\prime}_{n}$ in the $S_{n}$ frame, under Einstein synchronization, is obtained by introducing a time offset of the form given in Eq.
-13, where the velocity $v$ is replaced with $-v$: specifically $-x_{n}v/c^{2}$.
+13, where the velocity $v$ is replaced with v-v: specifically $x_{n}v/c^{2}$.
 Applying this offset yields the first-order approximation $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$.
-This time adjustment effectively rotates the plane of simultaneity by an angle $-v/c$.
+This time adjustment effectively rotates the plane of simultaneity by an angle v/c-v/c.
 Consequently, after re-synchronization, the radiation wavefront undergoes rotation in the accelerated frame, as illustrated in Fig.
-16.
+15.
 The new time coordinate in the accelerated frame is interpreted in accordance with Maxwell’s equations, ensuring their applicability to light propagation.
-As a result, the transmitted light beam (passing through the aperture) propagates at an angle $-v/c$, demonstrating the phenomenon of the aberration of light.
+As a result, the transmitted light beam (passing through the aperture) propagates at an angle v/c-v/c, demonstrating the phenomenon of the aberration of light.
 Both approaches yield the same outcome.
@@
 Substituting $dx_{n}/dt_{n}=-v$ into this equation confirms that $dx_{n}^{\prime}/dt_{n}^{\prime}=-v$.
-The group velocity of the light beam remains $-v$, but when we use the new velocity of light this becomes $\sin\theta_{a}=-v/c$.
+The group velocity of the light beam remains v-v, but when we use the new velocity of light this becomes $\sin\theta_{a}=-v/c$.
 This represents a correction to the classical aberration of light law within Lorentzian coordinates.
@@
 According to relativistic kinematics, the extra phase chirp $d\phi/dx^{\prime}_{n}=k_{x}=-v\omega_{i}/c^{2}$ is introduced as a consequence of this, the plane wavefront rotates after the acceleration.
-Then, the radiated light beam is propagated at the angle $-v/c$ with respect to the $z_{n}$-axis, yielding the phenomenon of the aberration of light in the accelerated frame $S_{n}$.
+Then, the radiated light beam is propagated at the angle v/c-v/c with respect to the $z_{n}$-axis, yielding the phenomenon of the aberration of light in the accelerated frame $S_{n}$.
 This phenomenon is purely kinematic and involves no forces.
@@
 The accumulated time difference between two spatially separated clocks is $t_{n}(1)-t_{n}(2)=-g[x_{n}(1)-x_{n}(2)]v/(c^{2}|g|)=v[x_{n}(1)-x_{n}(2)]/c^{2}$.
-This time shift effectively rotates the plane of simultaneity by an angle $-v/c$.
+This time shift effectively rotates the plane of simultaneity by an angle v/c-v/c.
 5.8 A Single Moving Emitter in an Accelerated Frame
@@
 Due to the fundamental asymmetry between inertial and accelerated frames, the theory of light aberration predicts a remarkable effect.
-Specifically, if the emitter remains at rest relative to the fixed stars while the observer transitions from rest to motion with respect to the fixed stars, the apparent angular position of the ”plane-wave” emitter, as seen in the accelerated frame, would jump by an angle of $-2v/c$.
+Specifically, if the emitter remains at rest relative to the fixed stars while the observer transitions from rest to motion with respect to the fixed stars, the apparent angular position of the ”plane-wave” emitter, as seen in the accelerated frame, would jump by an angle of 2v/c-2v/c.
 This aberration effect in the frame $S_{n}$ is illustrated in Fig.
-17.
-Let us describe the radiation emitted by a source moving with velocity $-v$ in an accelerated frame, as observed by an accelerated observer.
+16.
+Let us describe the radiation emitted by a source moving with velocity v-v in an accelerated frame, as observed by an accelerated observer.
 The acceleration relative to the accelerated frame influences the propagation of light.
@@
 After diagonalization, it becomes evident that any apparent asymmetry stems from the initial conditions.
-Notably, the plane radiation wavefront of a moving emitter (regardless of its velocity) undergoes a rotation by an angle of $-v/c$ in the accelerated frame as a result of the metric diagonalization.
+Notably, the plane radiation wavefront of a moving emitter (regardless of its velocity) undergoes a rotation by an angle of v/c-v/c in the accelerated frame as a result of the metric diagonalization.
+This implies that the radiated light beam has a group velocity along $x$-direction given by $d\omega/dk_{x}=-2v$.
 To predict the outcome of aberration measurements, one can alternatively analyze the metric of the moving emitter directly within the absolute-time coordinate system, given by Eq.
 (18).
-According to electrodynamics, the group velocity $-2v$ emerges as a direct consequence of the cross term $-4vdxdt$.
+According to electrodynamics, the group velocity 2v-2v emerges as a direct consequence of the cross term 4vdxdt-4vdxdt.
 Now, let us consider the most general case in which the system $S_{n}$, initially at rest in the inertial frame $S$, is accelerated to velocity $v$ along the $x$-axis.
@@
 17 - Fig.
-19).
+18).
 The aberration of a point source is independent of the source’s speed and is attributed solely to the observer’s motion relative to the fixed stars.
@@
 17 – Fig.
-19.
+18.
 A point source located in the far zone effectively generates a plane wave in front of the aperture.
@@
 Assuming that the emitted wavefront is initially tilted by an angle $\alpha$, we obtain $k_{x}=\alpha k=\alpha\omega_{i}/c$, $k_{z}=\sqrt{\omega_{i}^{2}/c^{2}-k_{x}^{2}}$.
-Since acceleration modifies the field equations, we substitute the plane wave solution $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ into Eq.
-(17), yielding: $(1-v^{2}/c^{2})k_{x}^{2}-2vk_{x}\omega/c^{2}+k_{z}^{2}-\omega^{2}/c^{2}=0$.
+Since acceleration modifies the field equations, we substitute the plane wave solution $\exp(i\vec{k}_{n}\cdot\vec{r}_{n}-i\omega_{n}t_{n})$ into Eq.
+(17), yielding: $(1-v^{2}/c^{2})(k_{n})_{x}^{2}-2v(k_{n})_{x}\omega_{n}/c^{2}+(k_{n})_{z}^{2}-\omega_{n}^{2}/c^{2}=0$.
 A screen with a hole acts as a diffraction grating, splitting the incident plane wave into multiple diffracted components.
-The wavenumber of the transmitted plane wave along the $x_{n}$-direction is $k_{x}=\alpha k+K_{\perp}$, where $K_{\perp}$ is the wavenumber associated with the sinusoidally modulated transmittance.
+The wavenumber of the transmitted plane wave along the $x_{n}$-direction is $(k_{n})_{x}=\alpha k+K_{\perp}$, where $K_{\perp}$ is the wavenumber associated with the sinusoidally modulated transmittance.
 The dispersion equation then simplifies to $\omega_{i}\Delta\omega=-\omega_{i}K_{\perp}v+c^{2}\alpha kK_{\perp}$.
@@
 This effect, illustrated in Fig.
-20, is a well-known example of the phenomenon of aberration of light.
+19, is a well-known example of the phenomenon of aberration of light.
 Figure 19: The aberration of light from stationary and moving sources.
@@
 As shown in Fig.
-21, the aberration increment is related to the system’s parameters by the expression $\theta_{a}=-2v/c$.
+20, the aberration increment is related to the system’s parameters by the expression $\theta_{a}=-2v/c$.
 .
+Figure 21: Diagram of a modified source of the ”plane wave” emitter type. A point source of light is placed in the front focal plane of a lens.
 Figure 22: Illustration of the accelerated source redirection process.
@@
 Consequently, for simplicity, the effect of aberration is explained using two different coordinatizations: absolute time (left) and Lorentz coordinates (right).
+To further explore the aberration of light in an accelerated frame, it is useful to consider a slightly modified source—a ”plane wave” emitter.
+We adopt the model described in Section 4.2 because it is both relatively simple and sufficiently general to serve as a prototype for understanding aberration phenomena.
+While no one has performed all the ”thought experiments” exactly as described here, the outcomes can be confidently predicted based on the laws of special relativity, which are themselves grounded in experimental evidence.
+Figure 21 illustrates the modified ”plane wave” emitter setup.
+Though it appears more complex at first glance, it is practically feasible to construct.
+The setup consists of a point source of light positioned in the front focal plane of a lens.
+The source’s dimensions are assumed to be on the order of $\lambda$, where $\lambda$ is the wavelength of optical radiation.
 Let us revisit the redirection procedure for an accelerated source.
@@
 11, which describes the electrodynamics of a moving emitter from the perspective of an inertial observer.
-On the other hand, the Langevin metric describes the electrodynamics of an accelerated emitter from the viewpoint of an accelerated observer.
+On the other hand, the Langevin metric describes the electrodynamics of an accelerated emitter from the viewpoint of an accelerated observer measurements.
 A key principle is that the metric tensor must remain continuous.
@@
 These two metrics are two sides of the same coin.
-Ultimately, we describe the electrodynamics of the same accelerated emitter—accelerated with respect to the fixed stars—from both an inertial and an accelerated observer’s viewpoint.
 Notably, this theorem is absent from conventional formulations of special relativity.
@@
 This implies that the Minkowski metric, $ds^{2}=c^{2}(dt^{\prime})^{2}-(dx^{\prime})^{2}-(dy^{\prime})^{2}-(dz^{\prime})^{2}$, governs the electrodynamics of the accelerated light source in the comoving coordinate system.
-However, this raises a fundamental question: How can the Minkowski metric applies in the comoving coordinatization in the inertial frame while the Langevin metric applies in the accelerated frame?
-This apparent discrepancy requires further explanation.
-The difference between measurements made in the two reference frames is particularly intriguing.
-To analyze this, we first examine the relationship between passive and active transformations.
-It is important to note that a passive Galilean boost within a single inertial frame, given by $x^{\prime}=x-vt,t^{\prime}=t$, merely represents an alternative parametrization of the inertial observer’s measurements.
-Eq.
-(11) describes these measurements.
-In the case of active transformations within a single inertial frame, the motion of fixed stars relative to the inertial observer and their instruments remains unchanged.
-This establishes the equivalence between active and passive Galilean boosts.
-By contrast, the Langevin metric describes the measurements of an accelerated observer.
-Under the inverse Galilean boost, the motion of fixed stars relative to the space-time grid $(t_{n},x_{n},y_{n},z_{n})$ of the accelerated observer is altered.
-On the other hand, the standard textbook treatment of Lorentz transformations implicitly assumes that the $x_{n}y_{n}z_{n}$-axes of the accelerated observer remain parallel to the $xyz$-axes of the inertial observer.
+However, a conceptual tension emerges: Why does the Minkowski metric apply in the comoving description (within an inertial frame), while the Langevin metric governs the physics in the truly accelerated frame?
+This apparent inconsistency requires careful examination.
+The key lies in recognizing that inertial and accelerated observers do not share the same three-dimensional space in Minkowski spacetime.
+Although this might seem paradoxical if both observers were assumed to share a common spatial geometry, special relativity resolves the issue: observers following different worldlines inhabit different instantaneous 3-spaces due to the relativity of simultaneity.
+To analyze this, we begin by examining the relationship between passive and active transformations.
+A passive Galilean boost within a single inertial frame, given by $x^{\prime}=x-vt,t^{\prime}=t$, simply re-expresses the results of measurements made by the inertial observer using a new set of coordinates.
+In the case of active transformations within the same inertial frame, the motion of fixed stars relative to the observer and their instruments remains unchanged.
+This establishes the equivalence between active and passive Galilean boosts in inertial frames.
+In contrast, the Langevin metric describes measurements made by an accelerated observer.
+Under an inverse Galilean boost, the motion of fixed stars relative to the accelerated observer’s space-time grid $(t_{n},x_{n},y_{n},z_{n})$ is no longer preserved.
+This highlights the fundamental difference in how motion is perceived in non-inertial frames.
+When observers adopt absolute-time coordinatization, relativistic kinematics becomes remarkably simple.
+In this approach, all asymmetries between the initial inertial and accelerated frames are absorbed into their respective frame metrics.
+Alternatively, consider observers using Lorentz coordinatization.
+Here, the frame metrics become diagonal, and the complexities of special relativity manifest as counterintuitive kinematic effects.
+Notably, in this coordinatization, the inertial and accelerated observers experience different three-dimensional spatial geometries.
+On the other hand, the standard textbook treatment of Lorentz transformations implicitly assumes that the $x_{n}y_{n}z_{n}$-axes of the accelerated observer remain parallel to the xyzxyz-axes of the inertial observer.
 In other words, it presumes that both observers share a common three-dimensional space.
@@
 However, this is a misconception.
-The conventional derivation of the aberration of light effect overlooks the conventionality of distant simultaneity.
-There is an inherent uncertainty in the relative position along the $z_{n}$-axis in the $x$-direction, amounting to $z_{n}v/c$, due to the ambiguity in the timing of acceleration.
-Consequently, we are dealing with space-like separated events.
-The orientation of the $z_{n}$-axis lacks an exact objective meaning because of the relativity of simultaneity.
-The claim that the $z_{n}$-axis orientation with respect to the $z$-axis of the inertial observer has objective meaning to within a certain accuracy can be visualized by the picture of $z_{n}$ axis in the vertical orientation with approximate angle uncertainty (blurring) given by $\Delta\theta=v/c$.
-This inclination angle between the inertial $z$ and accelerated $z_{n}$ reference directions has no exact objective meaning at all since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this displacement could be ascertained.
-We will explore this issue further in Chapter 7.
+For example, conventional derivation of the aberration of light effect overlooks the conventionality of distant simultaneity.
+There exists an inherent ambiguity in the spatial position along the $z_{n}$-axis (in the $x$-direction) due to uncertainties in the timing of acceleration.
+This uncertainty can be quantified as $z_{n}v/c$, revealing that the associated events are space-like separated.
+Consequently, the orientation of the $z_{n}$-axis lacks an exact objective meaning, due to the relativity of simultaneity.
+The idea that the direction of the $z_{n}$-axis relative to the inertial $z$-axis is physically well-defined is, at best, an approximation.
+This relationship can be visualized as a vertical $z_{n}$ axis with an angular uncertainty (or ”blurring”) given by: $\Delta\theta=v/c$.
+This inclination angle has no precisely defined physical meaning, since—given the finite speed of light—there is no experimental method to determine such a spatial displacement.
+Let us illustrate these ideas with the example of a plane-wave emitter moving along the $x$-axis with velocity $v$ in an initial inertial frame.
+Assume that clocks are synchronized according to the Einstein synchronization procedure, using light signals from a stationary dipole source.
+The electrodynamics of the moving source in this inertial frame is described by the metric in Eq.
+11.
+As previously shown, the radiated beam exhibits aberration: due to the motion of the source, the direction of energy propagation is altered.
+However, when we introduce a comoving coordinate system via a passive Galilean boost, this aberration disappears.
+In this comoving system, the light beam propagates along the $z$-axis.
+Crucially, because the boost is performed within a single inertial frame, the comoving coordinate axes $(x^{\prime},y^{\prime},z^{\prime})$ remain parallel to the original inertial axes $(x,y,z)$.
+That is, both systems inhabit the same three-dimensional space.
+Let us now examine the relationship between the coordinate system of the inertial frame and that of the comoving frame.
+To detect aberration in the inertial frame, one must first establish a coordinate system, which requires a physical, operational procedure.
+In this sense, the spatial grid $(x,y,z)$ of the inertial frame represents a physical reality.
+By contrast, the comoving coordinate grid $(x^{\prime},y^{\prime},z^{\prime})$ introduced through the passive Galilean boost is a purely mathematical construct—there is no need to consider how it would be physically realized.
+It is important to emphasize that a passive Galilean boost within a single inertial frame merely provides an alternative parametrization of the same observations made by an inertial observer using real, physical reference axes.
+Now consider the perspective of an accelerated observer.
+As before, to detect any aberration effect within the accelerated frame, a coordinate system must be established.
+This requires the observer to define a practical, operational method for assigning spatial coordinates $(x_{n},y_{n}z_{n})$.
+The accelerated observer finds that the angular displacement of the beam is negative, given by $\theta_{a}=-v/c$.
+This result leads to a key insight: from the viewpoint of the inertial observer, the $z_{n}$-axis of the accelerated observer is not parallel to the inertial $z$-axis.
+In other words, the inertial and accelerated observers inhabit different three-spaces within Minkowski spacetime.
+This subtle but important distinction reflects the breakdown of simultaneity in relativistic physics and anticipates deeper discussions that will follow in Chapter 7.
 5.13 Ether Theory and the Aberration of Light Effect
@@
 We begin by examining the influence of the ether wind on light speed.
-According to the ether hypothesis, the velocity of light observed from an accelerated frame would be $c+u$ for a beam propagating in the same direction as the accelerated system and $-c+u$ for a beam propagating in the opposite direction, where $u=-v$ represents the ether velocity in the accelerated reference frame.
+According to the ether hypothesis, the velocity of light observed from an accelerated frame would be $c+u$ for a beam propagating in the same direction as the accelerated system and c+u-c+u for a beam propagating in the opposite direction, where u=vu=-v represents the ether velocity in the accelerated reference frame.
 The aberration effect arises as a consequence of the ether wind’s influence on the speed of light (Fig.
@@
 Maxwell’s equations hold in the initial inertial frame, where the optical system (e.g., a light source and lens, as shown in Fig.
-13) is at rest, and time coordinates are assigned via slow clock transport.
+21) is at rest, and time coordinates are assigned via slow clock transport.
 To ensure continuity, the metric of a non-inertial frame associated with an accelerating optical system (relative to the initial inertial frame) must transition smoothly into the Langevin metric.
@@
 In this framework, Maxwell’s equations describe the field configuration, and the resynchronized time $t^{\prime}_{n}$ is given by $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$ in the first-order approximation.
-After resynchronization, the asymmetry shifts to the initial condition: the plane wavefront of the emitter undergoes a rotation by an angle $-v/c$ in the accelerated frame.
+After resynchronization, the asymmetry shifts to the initial condition: the plane wavefront of the emitter undergoes a rotation by an angle v/c-v/c in the accelerated frame.
 At first glance, Lorentz’s treatment of an optical setup in an inertial frame moving through the ether appears similar to our approach to such a setup in an accelerated frame (relative to the fixed stars).
@@
 According to the wave equation (Eq.
-19), the sound velocity observed from an accelerated reference frame is $v_{s}-v$ for a beam propagating in the same direction as the accelerated system and $-v_{s}-v$ for a beam propagating in the opposite direction.
-Here, $-v$ represents the wind velocity in the accelerated frame.
+19), the sound velocity observed from an accelerated reference frame is $v_{s}-v$ for a beam propagating in the same direction as the accelerated system and $v_{s}-v$ for a beam propagating in the opposite direction.
+Here, v-v represents the wind velocity in the accelerated frame.
 This effect, known as aberration, arises due to the influence of the perceived air wind on the speed of sound (Fig.
@@
 6.3 Stellar aberration in the Context of Special Relativity
-It is widely accepted that the theory of relativity aligns with Bradley’s discovery of stellar aberration, interpreting it as a consequence of an observer’s motion relative to a light source 232323 In his paper on relativity, Einstein derived the aberration formula by considering the velocity $v$ as the relative velocity of the star-Earth system.
+It is widely accepted that the theory of relativity aligns with Bradley’s discovery of stellar aberration, interpreting it as a consequence of an observer’s motion relative to a light source.
+232323 In his paper on relativity, Einstein derived the aberration formula by considering the velocity $v$ as the relative velocity of the star-Earth system.
 This interpretation has been adopted by numerous textbook authors.
 As Moeller [13] states: ”This phenomenon, known as aberration, was observed by Bradley, who noticed that stars appear to undergo an annual collective motion in the sky.
-This apparent motion arises because the observed direction of a light ray from a star depends on the Earth’s velocity relative to the star.”.
+This apparent motion arises because the observed direction of a light ray from a star depends on the Earth’s velocity relative to the star.”
 The asymmetry between the cases where either the star or the Earth-based telescope is in motion becomes evident when considering the separation of binary stars.
@@
 17 - Fig.
-19.
+18.
 Figure 29: The direction of a star as seen from the earth.
@@
 7.1 Inertial frame view of Observations of a Non-Inertial Observer
-In this chapter, we introduce a novel approach to the theory of light aberration in non-inertial reference frames, providing an alternative perspective on this complex issue.
-First, we highlight a fundamental principle: the laws of physics in initial inertial frame must account for all physical phenomena, including observations made by non-inertial observers.
-It is commonly believed that phenomena in non-inertial (e.g., rotating) reference frames should be analyzed solely within the geometric space-time framework using the Langevin metric.
-However, we argue that physical phenomena in non-inertial frames can also be effectively described within an inertial frame using standard Einsteinian special relativity, leveraging well-established relativistic kinematic effects.
-For example, when observed from an inertial laboratory frame, the interpretation of the Sagnac effect is straightforward.
-The phase difference between counter-propagating waves, attributed to the Sagnac effect, can be derived from the relativistic velocity addition formula.
-In other words, within the rotating frame, as seen from the inertial lab frame, the Sagnac effect manifests as a purely kinematic phenomenon governed by Einstein’s velocity addition in special relativity.
-This analysis assumes that the speed of light in the lab frame is $c$, which, according to special relativity, remains independent of the motion of the source.
-In contrast, calculating the Sagnac effect within a rotating reference frame is more complex.
-In such cases, textbook authors often employ a metric tensor—specifically, the Langevin metric—within a four-dimensional Minkowski spacetime to determine the propagation time difference between counter-propagating waves.
-Could we not analyze the aberration of light in the same manner?
-Here, too, we apply the Langevin metric in a non-inertial reference frame.
+In this chapter, we introduce a novel approach to the theory of light aberration in non-inertial reference frames, offering a fresh perspective on this complex subject.
+We begin by emphasizing a key principle: the laws of physics in an initial inertial frame must fully account for all physical phenomena, including those observed by non-inertial observers.
+This principle was already applied in Chapter 5, where we analyzed phenomena in non-inertial (e.g., rotating) reference frames using the Langevin metric.
+This metric captures the electrodynamics of an accelerated light source from the viewpoint of measurements made by an accelerated observer—but interpreted from the perspective of an inertial observer.
+Our findings there can be summarized as follows:
+1. Inertial Frame Setup
+We assume that the inertial frame has no history of acceleration and that spacetime, as perceived by an observer at rest, is described by the standard diagonal Minkowski metric: $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$.
+Applying a Galilean transformation $x\to x-vt$, $t\to t$ we obtain the metric of an accelerated emitter in the inertial frame, as presented in Eq.
+(11).
+In this coordinate system $(t,x,y,z)$, the devices of the inertial observer remain governed by the standard diagonal metric.
+2. Non-Inertial Frame Transformation
+To describe an accelerated observer with coordinates $(t_{n},x_{n},y_{n},z_{n})$, we apply the inverse Galilean transformation: $x\to x_{n}+vt$, with $t\to t_{n}$ unchanged.
+Substituting this into the Minkowski metric yields the Langevin metric, as shown in Eq.
+(15).
+However, we argue that physical phenomena in accelerated frames can also be effectively analyzed within an inertial frame using the framework of standard Einsteinian special relativity and its well-established kinematic effects.
+For instance, consider the Sagnac effect.
+From the perspective of an inertial laboratory frame, its interpretation becomes straightforward.
+The phase difference between counter-propagating waves arises from the relativistic velocity addition formula.
+That is, as viewed from the lab frame, the Sagnac effect emerges purely as a kinematic consequence of special relativity.
+This analysis assumes the speed of light in the lab frame is $c$, which—according to special relativity—is independent of the source’s motion when using Lorentz coordinates.
+In contrast, analyzing the Sagnac effect within a rotating frame is more intricate.
+In such cases, authors typically employ a spacetime approach using a metric tensor—namely, the Langevin metric—within a four-dimensional Minkowski framework, to compute the propagation time differences between counter-propagating waves.
+This leads us to a natural question: Can we analyze the aberration of light in the same way?
+Indeed, we propose applying the Langevin metric to describe light aberration in non-inertial frames as well.
+That said, this approach requires a clock resynchronization procedure in the inertial frame.
+To maintain Lorentz coordinates for a moving light source, one must diagonalize the metric given in Eq.
+(11), as detailed in Section 3.5.
+The Lorentz coordinates can then be expressed in terms of the original coordinates $(t,x,y,z)$, using the transformation given in Eq.
+(13).
+7.2 Relativity of Simultaneity
 Let us first apply the classical kinematic method to compute the aberration of light effect in a non-inertial frame, as observed from an inertial frame.
@@
 According to our approach, a remarkable prediction emerges regarding light aberration in an accelerated frame (see Fig.
-20).
+19).
 To correctly predict the aberration measurement, the accelerated observer must use the Langevin metric.
@@
 In Fig.
-16, the transmitted light beam propagates at an angle of $-v/c$, resulting in the phenomenon of light aberration.
-This naturally raises the question: relative to what does a light beam propagate within an accelerated frame when it undergoes an angular displacement of $-v/c$?
+15, the transmitted light beam propagates at an angle of v/c-v/c, resulting in the phenomenon of light aberration.
+This naturally raises the question: relative to what does a light beam propagate within an accelerated frame when it undergoes an angular displacement of v/c-v/c?
 Consider an observer in the accelerated frame measuring the direction of the light beam.
@@
 Now, let us return to the observations of an accelerated observer as viewed from an inertial frame.
-We begin with the non-covariant approach.
-The aberration of light problem can be addressed within a ”single inertial frame” description, without invoking Lorentz transformations.
-According to our approach, when a transversely moving screen with a hole is placed in the path of a normally incident plane wave of light, the transmitted light undergoes a deviation in energy transport (see Fig.
-9).
-However, an alternative and equally valid explanation for the aberration of light from a moving source relies on Lorentz coordinatization.
-This explanation describes the uniform translational motion of the light source within the inertial frame.
 In Chapter 3, we discussed how absolute time coordinatization can be transformed into Lorentz coordinatization.
@@
 30).
-We conclude that, in the accelerated frame, the transmitted light beam propagates along the $z_{n}$-axis of the accelerated frame with an angular displacement of $-v/c$.
+We conclude that, in the accelerated frame, the transmitted light beam propagates along the $z_{n}$-axis of the accelerated frame with an angular displacement of v/c-v/c.
 We determine the reference directions in both the inertial and accelerated frames using two local light sources, leading to intriguing consequences.
@@
 Now, we examine the angular displacement of the transmitted light relative to the reference beam within the accelerated frame.
-On one hand, the angular displacement of the transmitted light with respect to the reference beam from an emitter at rest in the accelerated frame is $-v/c$.
-On the other hand, the angular displacement of the inertial reference light beam is $-2v/c$ (see Fig.
-17).
+On one hand, the angular displacement of the transmitted light with respect to the reference beam from an emitter at rest in the accelerated frame is v/c-v/c.
+On the other hand, the angular displacement of the inertial reference light beam is 2v/c-2v/c (see Fig.
+16).
 This implies that the accelerated observer can directly measure the angular displacement of the transmitted light with respect to the inertial frame’s reference light beam.
@@
 Here, too, an uncertainty of order $lv/c$ arises, where $l$ is the rod length.
-For completeness, we must also analyze the angular displacement of the $x_{n}$-axis.
-So far, we have considered only non-relativistic speeds, employing a first-order relativistic approximation.
-It follows that for the displacement of the detector along the $x$-axis in the $z$-direction, we obtain: $\Delta z\simeq\theta_{w}Lv/c\simeq Lv^{2}/c^{2}$.
-Thus, special relativity indicates that in the limit $v^{2}/c^{2}\to 0$, the kinematic shearing motion (during acceleration) and the Wigner rotation motion become indistinguishable.
 8 Kinematics of Wigner Rotation
-The focus of this chapter is the Wigner rotation  [^28] It is well known that the composition of non-collinear Lorentz boosts does not result in a simple boost but rather in a Lorentz transformation that includes both a boost and a spatial rotation.
+The focus of this chapter is the Wigner rotation.
+[^28] It is well known that the composition of non-collinear Lorentz boosts does not result in a simple boost but rather in a Lorentz transformation that includes both a boost and a spatial rotation.
 This rotation is known as the Wigner rotation [39, 40].
-It is sometimes referred to as the Thomas rotation (see, e.g., [13, 41])..
+It is sometimes referred to as the Thomas rotation (see, e.g., [13, 41]).
 We previously introduced the Wigner rotation in the last chapter, where our discussion primarily centered on the aberration of light.
@@
 Let us denote the three inertial frames by $K,K^{\prime},K$".
-The initial inertial frame is $K$, while $K^{\prime}$ is the rest frame of the particle, which has a velocity $\vec{v}=\vec{v}(\tau)$ relative to $K$, and $K"$ represents the particle’s rest frame at the next instant of proper time $\tau+d\tau$, moving with an infinitesimal velocity $d\vec{v^{\prime}}$ relative to $K^{\prime}$.
+The initial inertial frame is $K$, while $K^{\prime}$ is the rest frame of the particle, which has a velocity $\vec{v}=\vec{v}(\tau)$ relative to $K$, and KK" represents the particle’s rest frame at the next instant of proper time $\tau+d\tau$, moving with an infinitesimal velocity $d\vec{v^{\prime}}$ relative to $K^{\prime}$.
 All three inertial frames are assumed to be Lorentz frames.
-To clarify velocity measurements in different inertial frames: an observer in $K$ measures the velocity of frame $K^{\prime}$ as $\vec{v}$, while an observer in $K^{\prime}$ measures the velocity of frame $K"$ as $d\vec{v^{\prime}}$.
+To clarify velocity measurements in different inertial frames: an observer in $K$ measures the velocity of frame $K^{\prime}$ as $\vec{v}$, while an observer in $K^{\prime}$ measures the velocity of frame KK" as $d\vec{v^{\prime}}$.
 Now, let us examine the transformation of velocity in the theory of relativity.
 For a rectilinear motion along the $z$ axis, the velocity transformation follows the equation: $v_{z}(\tau+d\tau)=(dv^{\prime}_{z}+v_{z})/(1+v_{z}dv^{\prime}_{z}/c^{2})$.
-As it happens with the composition of Galilean boosts, collinear Lorentz boosts commute.
-This implies that the result of successive collinear Lorentz boosts is independent of which transformation applies first.
-Unlike the case of Lorentz boosts in collinear directions, Lorentz boosts in different directions do not commute.
-While the successive application of two Galilean boosts results in a single Galilean boost, and the successive application of two rotations yields another rotation, the composition of two non-collinear Lorentz boosts does not produce a single Lorentz boost.
-We assume that frame $K^{\prime}$ moves with velocity $v_{z}$ along the $x$-axis relative to $K$ and in frame $K^{\prime}$, the particle’s acceleration is perpendicular to the trajectory of the initial inertial frame.
-Suppose that frame $K"$ moves relative to $K^{\prime}$ with nonrelativistic velocity $dv^{\prime}_{x}$.
-From the perspective of an inertial observer, time dilation occurs in the Lorentz frame moving with velocity $\vec{v}_{z}$ relative to the initial inertial frame, given by $dt/\gamma=d\tau$.
-Due to this time dilation, the velocity increment in the initial inertial frame $dv_{x}$ corresponds to a velocity $dv^{\prime}_{x}=\gamma dv_{x}$ in frames $K^{\prime}$.
+Just as with Galilean transformations, collinear Lorentz boosts commute—the order in which collinear boosts are applied does not affect the final result.
+This means that the composition of two Lorentz boosts in the same direction yields another Lorentz boost in that direction, with the total velocity given by the formula above.
+However, when Lorentz boosts are applied in non-collinear directions, they no longer commute.
+While the composition of two Galilean boosts or two spatial rotations yields another transformation of the same kind (i.e., a boost or a rotation), the combination of two non-parallel Lorentz boosts results in a transformation that is not purely a boost—it includes a rotation known as the Wigner rotation.
+To illustrate this, suppose frame $K^{\prime}$ moves with velocity $v_{z}$ along the $z$-axis relative to $K$ and in $K^{\prime}$, the particle experiences an infinitesimal acceleration in the $x^{\prime}$-direction, perpendicular to the trajectory of the initial inertial frame $K$.
+That is, frame KK" moves relative to $K^{\prime}$ with a small velocity $dv^{\prime}_{x}$ along the $x^{\prime}$-axis.
+From the perspective of an inertial observer, time dilation occurs in the Lorentz frame moving with velocity $\vec{v}_{z}$ relative to the initial inertial frame, given by $dt/\gamma=d\tau$, where $\gamma=1/\sqrt{1-v_{z}^{2}/c^{2}}$.
+Due to this time dilation, the velocity increment in the initial inertial frame $dv_{x}$ corresponds to a velocity $dv^{\prime}_{x}=\gamma dv_{x}$ in $K^{\prime}$.
 8.2 Wigner Rotation
@@
 The key challenge lies in bridging the conceptual gap that arises from the fact that different Lorentz frames cannot be described within a shared three-dimensional space.
-[…]
-[…]
+Consider a succession of inertial frames $K\to K^{\prime}\to K$".
+From the perspective of the initial inertial frame, an observer in the frame KK" sees frame $K$ moving with velocity $v_{z}$ and velocity increment $\gamma d\vec{v}_{x}$.
+The corresponding rotation of the velocity direction in the KK" frame is given by $d\theta_{K"}=\gamma dv_{x}/v_{z}$ (Fig.
+36).
+In the $K$ frame, the velocity rotation angle is $dv_{x}/v_{z}=d\theta$, with both rotations occurring in the same direction.
+If the proper and lab observers were to share a common three-dimensional space, this discrepancy would lead to a paradox — an apparent inconsistency in the theory.
+However, in Minkowski spacetime, no such paradox arises: the lab’s 3-space is rotated relative to the proper 3-space by an angle $d\Phi_{K"}=\gamma dv_{x}/v_{z}-dv_{x}/v_{z}=(\gamma-1)\d\theta$ in the same direction as the velocity vector’s rotation in the proper frame.
+This angular difference $d\Phi_{K"}=(\gamma-1)\d\theta$ is the Wigner rotation — the rotation of the original frame’s axes as seen from the proper frame KK".
+In vector form, this can be expressed as $d\vec{\Phi}_{K"}=(\gamma-1)\vec{v}\times d\vec{v}/v^{2}$.
 Several key insights can be drawn regarding the theory of Wigner rotation.
@@
 Transforming this expression back to the lab frame $K$ yields $d\vec{\Phi}_{K}=(1-1/\gamma)\vec{v}\times d\vec{v}/v^{2}$ representing the Wigner rotation angle of the proper frame’s axes as observed in the lab frame.
-According to the principles of special relativity, this quantity must remain form-invariant under Lorentz transformations  [^29] The correct expression for the Thomas (Wigner) rotation was first obtained by V.
+According to the principles of special relativity, this quantity must remain form-invariant under Lorentz transformations.
+[^29] The correct expression for the Thomas (Wigner) rotation was first obtained by V.
 Ritus [42].
@@
 As analyzed in Ritus’s work [43], Moeller’s error is not computational but rather conceptual.
-The correct expression, derived decades earlier by several researchers, remained largely unnoticed amid the proliferation of inaccurate results, as discussed in the review [44]..
-We note that owing to the relativistic effect of time dilation in the reference frame that moves to the lab frame, the Wigner rotation angle in the proper frame is always $\gamma$ time higher than in the lab frame  [^30] We note that in 1986, M.
+The correct expression, derived decades earlier by several researchers, remained largely unnoticed amid the proliferation of inaccurate results, as discussed in the review [44].
+We note that owing to the relativistic effect of time dilation in the reference frame that moves to the lab frame, the Wigner rotation angle in the proper frame is always $\gamma$ time higher than in the lab frame.
+[^30] In 1986, M.
 Stranberg obtained an expression for the Thomas (Wigner) rotation correct both in the lab inertial frame and the comoving reference frame [45].
-His paper is one of the few that explicitly states that the angle of Thomas (Wigner) rotation in the comoving frame is $\gamma$ times higher than in the lab frame..
+His paper is one of the few that explicitly states that the angle of Thomas (Wigner) rotation in the comoving frame is $\gamma$ times higher than in the lab frame [44].
 We derived the exact relation $d\vec{\Phi}=(1-1/\gamma)\vec{v}\times d\vec{v}/v^{2}$ using only rudimentary knowledge of special relativity.
@@
 To explore the physical meaning and possible experimental interpretation of this rotation, we illustrate the problem of defining the orientation of a moving frame using a simple example.
-The Wigner rotation, when interpreted in the context of the original inertial frame, is intimately related to length contraction  [^31] The length of a moving object depends not only on the structure of space-time but also on our choice of synchronization convention.
+The Wigner rotation, when interpreted in the context of the original inertial frame, is intimately related to length contraction.
+[^31] The length of a moving object depends not only on the structure of space-time but also on our choice of synchronization convention.
 As with the orientation of radiation phase fronts, such quantities do not possess an exact objective meaning.
-Under the standard clock synchronization, they are typically regarded as ”relativistic kinematic effects.” The Wigner rotation similarly arises as a coordinate effect, lacking an absolute, observer-independent significance..
+Under the standard clock synchronization, they are typically regarded as ”relativistic kinematic effects.” The Wigner rotation similarly arises as a coordinate effect, lacking an absolute, observer-independent significance.
 Figure 37: Illustration of the Wigner rotation interpreted with respect to the axes of the lab frame.
@@
 However, for an arbitrary (finite) velocity, the comoving frame axes are not aligned with the rotated lab frame axes.
-Due to the contracted projection, the angular deviation is $-\vec{\delta\theta}/\gamma$.
+Due to the contracted projection, the angular deviation is $\vec{\delta\theta}/\gamma$.
 Consequently, one can directly verify that the comoving frame axes are rotated relative to the original lab frame axes $(x,z)$ by an angle $\delta\theta-\delta\theta/\gamma$, which is precisely the Wigner rotation angle, as given by Eq.(20).
@@
 Thus, the spin rotation measurements by a polarimeter in the lab frame is interpreted from the viewpoint of the proper observer as viewing of the lab observer.
-This inherent complexity has led textbook authors to derive incorrect expressions for Wigner rotation—an issue that will be discussed further in Chapter 17.
+This inherent complexity has led textbook authors to derive incorrect expressions for Wigner rotation—an issue that will be discussed further in Chapter 18.
 Additionally, it is crucial to examine why errors in Wigner rotation theory went undetected for so long.
@@
 Our goal is to determine the relationship between these sets of fields.
-To proceed, we transform the coordinates $(t,x,y,z)$ of the inertial observer $S$, moving with velocity $-v$ relative to the accelerated observer $S_{n}$, using the Galilean transformation: $x_{n}=x-vt$, $t_{n}=t$.
+To proceed, we transform the coordinates $(t,x,y,z)$ of the inertial observer $S$, moving with velocity v-v relative to the accelerated observer $S_{n}$, using the Galilean transformation: $x_{n}=x-vt$, $t_{n}=t$.
 Substituting these into the Minkowski metric, $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$, we obtain the so-called Langevin metric in the accelerated frame Eq.
 (15): $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$.
-This transformed metric allows us to study Maxwell’s equations in a non-inertial frame using the Galilean transformation with velocity $-v$.
+This transformed metric allows us to study electrodynamics equations in a non-inertial frame using the Galilean transformation with velocity v-v.
 In the specific case where $\vec{B}=0$, the transformed fields to first order in $v/c$ are: $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$.
@@
 Figure 38: Electron aberration observed when both the electron source and the observer are at rest in an accelerated frame.
-[…]
+The accelerated observer experimentally determines the Lorentz force $\vec{F}_{n}$ acting on a charge $e$ moving with velocity $\vec{V}$ in the region of the electron gun, where electric and magnetic fields $\vec{E}_{n}$ and $\vec{B}_{n}$ are present.
+The force is given by: $\vec{F}_{n}=e\vec{E}-e\vec{V}\times\vec{v}\times\vec{E}/c^{2}$.
+In the special case where $\vec{v}\cdot\vec{V}=0$ and $\vec{E}\cdot\vec{V}=EV$, the magnetic term simplifies to: $e\vec{V}\times\vec{v}\times\vec{E}/c^{2}=-\vec{v}(e\vec{V}\cdot\vec{E}/c^{2})$.
+Using the relativistically correct equation of motion, $\vec{F}_{n}=d\vec{p}/dt$, we expect that after a short time interval $\d t$, the emitted electron will acquire a transverse momentum in the accelerated frame given by: $\d\vec{p}_{\perp}=-\vec{v}(e\vec{V}\cdot\vec{E}/c^{2})\d t$.
+The differential change in kinetic energy of the accelerated electrons is: $\d T=\vec{F}\cdot\vec{V}\d t=\vec{F}\cdot\d\vec{s}$.
+Upon integration, this yields: $T=e\int\vec{E}\cdot\d\vec{s}=mc^{2}/\sqrt{1-V^{2}/c^{2}}-mc^{2}$.
+Given our condition $\vec{v}\cdot\vec{V}=0$, the aberration angle increment becomes: $\theta_{a}=p_{\perp}/p$, where the total momentum of the accelerated electron is: $p=mV/\sqrt{1-V^{2}/c^{2}}$.
+Substituting, we find: $\theta_{a}=-(1-1/\gamma)v/V$.
+Expressed in vector form, the aberration increment is: $\vec{\theta}_{a}=-(1-1/\gamma)\vec{V}\times\vec{v}/V^{2}$.
 The aberration of particles resulting from the transformation between an inertial and an accelerated coordinate system can be interpreted as a manifestation of the Lorentz force.
@@
 In a complete and self-consistent calculation, all unphysical instantaneous effects cancel out, preserving causality and agreement with physical observation.
-We derived the electrodynamics equations in an accelerated frame by applying the inverse Galilean transformation (with velocity $-v$) of Maxwell’s equations.
+We derived the electrodynamics equations in an accelerated frame by applying the inverse Galilean transformation (with velocity v-v) of Maxwell’s equations.
 Under a Galilean transformation, and assuming $\vec{B}=0$ in the original frame, the transformed fields become $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$.
@@
 Even when the electron source and the observer are at rest in the accelerated frame, the time derivative $\partial/\partial t_{n}$ is generally nonzero.
-This can be readily demonstrated by transforming from the coordinates $(t,x,y,z)$ of an inertial observer $S$—moving with velocity $-v$ relative to the non-inertial observer $S_{n}$—to the coordinates $(t_{n},x_{n},y_{n},z_{n})$ of $S_{n}$.
+This can be readily demonstrated by transforming from the coordinates $(t,x,y,z)$ of an inertial observer SS—moving with velocity v-v relative to the non-inertial observer $S_{n}$—to the coordinates $(t_{n},x_{n},y_{n},z_{n})$ of $S_{n}$.
 Using a Galilean transformation, the inertial coordinates can be related to the non-inertial ones as follows: $x_{n}=x-vt,~y_{n}=y,~z_{n}=z,~t_{n}=t$.
@@
 As a result, the potential energy of a moving particle with intrinsic magnetic moment $\vec{m}$, in the presence of a magnetic field $\vec{B}$ and electric field $\vec{E}$, is given by: $U_{p}=-\vec{m}\cdot\vec{B}-\vec{d}\cdot\vec{E}$.
-To analyze this in an accelerated frame, we apply the Galilean transformation (with velocity $-\vec{v}$) to Maxwell’s equations.
+To analyze this in an accelerated frame, we apply the Galilean transformation (with velocity $\vec{v}$) to Maxwell’s equations.
 Under this transformation, assuming an initial zero electric dipole moment ($\vec{d}=0$), the magnetic and electric dipole moments in the accelerated frame become: $\vec{m}_{n}=\vec{m}$, $\vec{d}_{n}=-\vec{v}\times\vec{m}/c$.
@@
 This relation captures only first-order effects in $v/c$, where $\vec{v}$ is the velocity of the source relative to the lab frame.
-[…]
+An inertial observer experimentally measures the Lorentz force $\vec{F}$ acting on a charge $e$ moving with velocity $\vec{V}$ in the region of an electron gun, where both electric and magnetic fields, $\vec{E}$ and $\vec{B}$, are present.
+The force is given by: $\vec{F}=e\vec{E^{\prime}}+e\vec{V}\times\vec{v}\times\vec{E^{\prime}}/c^{2}$, where $\vec{E}^{\prime}$ is the electric field in the coordinate system moving with velocity $\vec{v}$ relative to the inertial frame.
+In the special case where $\vec{v}\cdot\vec{V}=0$ and $\vec{E}^{\prime}\cdot\vec{V}=E^{\prime}V$, the magnetic component of the force simplifies to: $e\vec{V}\times\vec{v}\times\vec{E^{\prime}}/c^{2}=\vec{v}(e\vec{V}\cdot\vec{E^{\prime}}/c^{2})$.
+The differential change in the kinetic energy of the accelerated electrons is: $dT=\vec{F}\cdot\vec{V}dt=\vec{F}\cdot d\vec{s}$.
+Integrating this expression yields the total kinetic energy: $T=e\int\vec{E^{\prime}}\cdot d\vec{s}=mc^{2}/\sqrt{1-V^{2}/c^{2}}-mc^{2}$.
+The total transverse momentum $p_{\perp}$ of the accelerated electron beam, to first order in $v/c$, is given by: $p_{\perp}=p_{0}+\Delta p_{\perp}=mv+vT/c^{2}=mv/\sqrt{1-V^{2}/c^{2}}$, where $p_{0}=mv$ is the initial transverse momentum of the emitted electron.
+Since we have chosen $\vec{v}\cdot\vec{V}=0$, the aberration angle $\theta_{a}$ becomes: $\theta_{a}=p_{\perp}/p$, where $p=mV/\sqrt{1-V^{2}/c^{2}}$ is the momentum of the accelerated electron.
+Substituting, we obtain: $\theta_{a}=v/V$.
+This can be written in vector form as: $\vec{\theta}_{a}=\vec{V}\times\vec{v}/V^{2}$.
+It is important to note that this result holds for an arbitrary Lorentz factor $\gamma=1/\sqrt{1-V^{2}/c^{2}}$.
 We analyzed the effect of particle aberration by considering terms only up to first order in $v/c$.
@@
 To understand this more concretely, consider the operation of an NMR magnetometer.
-When at rest in the lab frame, it measures the additional potential energy experienced by a proton with magnetic moment $\vec{m}$ in the field $\vec{B}$, given by $-\vec{m}\cdot\vec{B}\neq 0$.
+When at rest in the lab frame, it measures the additional potential energy experienced by a proton with magnetic moment $\vec{m}$ in the field $\vec{B}$, given by $\vec{m}\cdot\vec{B}\neq 0$.
 We come to the conclusion that the magnetic field generated by a moving source in an inertial frame is a real field in the sense we have defined it.
@@
 The Wigner rotation angle in the inertial frame can be expressed in vector form as $\vec{\theta_{w}}=(1-1/\gamma)\vec{V}\times\vec{v}/V^{2}$, where $\vec{V}$ is the vector of particle velocity in the initial Lorentz frame before acceleration, $\vec{v}$ is the vector of small velocity change ($v\ll V$) due to acceleration, $\gamma$ is the Lorentz factor associated with $V$, and $\theta_{w}$ is the Wigner rotation angle of the proper frame axes as observed from the (Lorentz) inertial frame.
+[^32] This expression for the Wigner rotation is derived in Chapter 8 for the limiting case of an infinitesimal rotation angle, $\delta\theta=v/V$, corresponding to a small change in the particle’s velocity vector.
+The derivation is based on the composition of Lorentz boosts in the form $L=L(\vec{v})L(\vec{V})$.
+In the context of an accelerated electron source, the relevant velocity composition arises from $L=L(\vec{V})L(\vec{v})$.
+A careful analysis shows that both orderings yield the same Wigner rotation expression, due to the infinitesimal nature of the boost $L(\vec{v})$.
 We now examine how this formula applies in the specific case where both the observer and the particle source are accelerated to a velocity $v$ in the inertial frame.
@@
 In the accelerated frame, this misalignment is reflected in the angular deviation, and we obtain an expression for the aberration increment: $\theta_{a}=-\theta_{w}=-(1-1/\gamma)v/V$.
-In the special case of light, where $V\to c$, this expression simplifies to: $\theta_{a}=-v/c$.
+In the special case of light this expression simplifies to: $\theta_{a}=-v/c$.
+[^33] It’s straightforward to verify that the infinitesimal Wigner rotation angle $\theta_{w}=(1-1/\gamma)v/V$ is always less than $v/c$ regardless of the particle velocity $V$.
+Specifically, $\theta_{w}=(v/c)(1-\sqrt{1-\beta^{2}})/\beta$, where $\beta=V/c$ and function $(1-\sqrt{1-\beta^{2}})/\beta$ is strictly less than 1 for $0<\beta<1$.
+As discussed in Chapter 5, there is an inherent uncertainty in the relative position along the $z_{n}$-axis in the $x$-direction, amounting to $z_{n}v/c$, due to the ambiguity in the timing of acceleration.
+This uncertainty arises because the relevant events are space-like separated.
+Consequently, the orientation of the $z_{n}$ axis relative to the inertial frame’s $z$-axis carries an angular uncertainty of $v/c$.
+The Wigner inclination angle $\theta_{w}$ always lies within this uncertainty bound.
 The formula $\vec{\theta}_{a}=-(1-1/\gamma)\vec{V}\times\vec{v}/V^{2}$ matches the result previously derived for observations made by a non-inertial observer within the framework of electrodynamics.
@@
 Measurement of the spot shift with respect to the optical axis (which is parallel to the zenith-nadir axis) is equivalent to the determination of the angular displacement.
-Due to the high stability, the aberration increment could be observed  [^32] The dedicated STEM with high stability was made [47].
-It was developed based on a 200-kV commercial instrument.
+Due to the high stability, the aberration increment could be observed.
+[^34] The dedicated STEM with high stability was made [47] It was developed based on a 200-kV commercial instrument.
 The electron source was installed in the anti-seismic room.
@@
 The average drift speed was 0.12 nm/min.
-It is substantially smaller than that conventional commercial microscope, in which the drift speed is about 2 nm/min..
+It is substantially smaller than that conventional commercial microscope, in which the drift speed is about 2 nm/min.
 Consider the inertial sun-based reference system.
@@
 In principle, records could be taken over a period of one day.
-The aberration shift depends only upon the value of $v_{\perp}$, the component of the orbital velocity perpendicular to the earth’s rotation axis 333333 The constant of diurnal aberration defines the scale of the aberration increment and is equal to $\theta_{d}=(1-1/1/\gamma)v_{e}/V\simeq 0.6$ $\mu\mathrm{rad}$.
+The aberration shift depends only upon the value of $v_{\perp}$, the component of the orbital velocity perpendicular to the earth’s rotation axis 353535 The constant of diurnal aberration defines the scale of the aberration increment and is equal to $\theta_{d}=(1-1/1/\gamma)v_{e}/V\simeq 0.6$ $\mu\mathrm{rad}$.
 Here $\gamma\simeq 1.39$, $V/c=0.7$, $v_{e}/c\simeq 1,55$ $\mu\mathrm{rad}$ where $v_{e}=0.45$ km/s is the linear velocity of the earth rotation at equator..
@@
 For an observer on the equator, the ellipse degenerates into a line segment, and for an observer at the pole of the earth, the ellipse is a circle.
-At the value $l=2$ mm of the focal length  [^34] Electron-optical system of the STEM contains an electron gun and several magnetic lenses, stacked vertically to form a lens column.
+At the value $l=2$ mm of the focal length, the major axis of the aberration ellipse of the order of 80 ($\cos\delta)$ nm.
+[^36] Electron-optical system of the STEM contains an electron gun and several magnetic lenses, stacked vertically to form a lens column.
 The illumination system of the instrument comprises the electron gun, together with two condenser lenses that focus the electrons onto the spacemen.
@@
 The distance between lens centers is typically 10 cm.
-Spacemen is located $D_{3}=25$ cm below the C2., the major axis of the aberration ellipse of the order of 80 ($\cos\delta)$ nm.
+Spacemen is located $D_{3}=25$ cm below the C2.
 The value of the minor axis depends on the observer’s position and is given by 80 ($\cos\delta\sin\phi))$ nm.
-When the latitude is $\phi=35^{o}$, the value of the scale of image shift per hour depends on the local time and varies from a value of about 10 nm/60 min to about 20 nm/60 min through the day  [^35] Experimental results (see Fig.
+When the latitude is $\phi=35^{o}$, the value of the scale of image shift per hour depends on the local time and varies from a value of about 10 nm/60 min to about 20 nm/60 min through the day.
+[^37] Experimental results (see Fig.
 44 p.167 in [47]) shows that the image shift is quite close to the theoretical prediction $\simeq 10~\mathrm{nm}$.
@@
 The details of the optical setup may influence the image shift.
-.
 10.2 A Thought Experiment with a Point-Like Source
@@
 It is the source linear dimension $d$ (source diameter) that determines the coherent area of the observed radiation $(z\lambda/d)^{2}$.
-The condition for neglecting the transverse size of the source, $d$, formulated as the requirement that the change of the correlation function along the aperture be less than unity, that is $d<D$ 363636 In all cases of practical interest, telescopes are situated in the far zone of the distant stars.
+The condition for neglecting the transverse size of the source, $d$, formulated as the requirement that the change of the correlation function along the aperture be less than unity, that is $d<D$.
+383838 In all cases of practical interest, telescopes are situated in the far zone of the distant stars.
 If the far zone approximation $z>d$ is satisfied, then the Van Cittert-Zernike theorem takes its simplest form.
@@
 We have only to change the distance and the source diameter.
-Assuming small parameter conditions $Dd/(z\lambda)<1$, $D^{2}/(z\lambda)<1$, the scaling approach yields the following requirements: $D^{2}/(z\lambda)<1$, $d<D$,.
-Let us estimate the parameters for an Earth-based setup using an incandescent lamp.
-We aim to describe a series of thought experiments—hypothetical, yet physically plausible.
-At optical frequencies, the conditions required for the Fraunhofer approximation can be quite stringent.
-For instance, with a wavelength of $\lambda=0.3$ $\mu$m and an aperture width of $D=3$ cm, the observation distance must satisfy $z>3000$ m.
-Obviously, the required conditions are met in a number of important problems.
-No one has ever done such an experiment, but we know what would happen from the astronomical observations.
+Assuming small parameter conditions $Dd/(z\lambda)<1$, $D^{2}/(z\lambda)<1$, the scaling approach yields the following requirements: $D^{2}/(z\lambda)<1$, $d<D$.
+[…]
 10.3 Practical Setup Using an Incoherent Optical Source
@@
 In optics, the emission from such a ”point” source is described classically using Maxwell’s equations, while the emitting medium (considered as an ensemble of atoms) is treated quantum mechanically.
-Thanks to modern lithography techniques, it is relatively straightforward to create an unresolved point source at optical wavelengths, allowing for the generation of sufficiently bright radiation  [^37] It may be possible to use the masked spatially incoherent UV light source.
+Thanks to modern lithography techniques, it is relatively straightforward to create an unresolved point source at optical wavelengths, allowing for the generation of sufficiently bright radiation.
+[^39] It may be possible to use the masked spatially incoherent UV light source.
 One could use the light source onto a (micron diameter) hole of a specially designed mask.
-The mask should be placed at the distance $l<<d$, where $d$ is the transverse size of the spatially incoherent UV light source..
+The mask should be placed at the distance $l<<d$, where $d$ is the transverse size of the spatially incoherent UV light source.
 Consider a diffracting aperture that is circular, and let the radius of the aperture be $a$.
@@
 The first zero of the Airy’s pattern is at $r_{i}=0.6\lambda/NA$.
-Let us estimate the parameters for the optical setup shown in Fig.42.
-Suppose $a=1.5$ cm and $\lambda=0.3$ $\mu$m.
-We will use $NA=0.1$.
-The resolution analysis is reduced to the theoretical framework of standard imaging theory.
-One can take advantage of well-known resolution criteria like Rayleigh criteria i.e.
-$20.6\lambda/NA=$ 2 $\mu$m.
-One could use a vertically oriented optical setup.
-The constant of aberration defines the scale of the aberration increment, and is equal to $v/c=100$ $\mu$rad.
-Measurement of the spot shift with respect to the optical axis (which is parallel to the nadir direction) is equivalent to the determination of the angular displacement.
-As already pointed out, the proposed method of measuring the angle of aberration involves the use of earth-based sources and has a big advantage.
-The rotation of the earth on its axis should produce a corresponding shift of the image.
-In principle, records could be taken over a period of one day.
-The image appears to move in an ellipse.
-The major axis of the ellipse is about 14 $\mu$m.
-The value of the minor axis depends on the observer’s position and is given by $14(\sin\phi)$ $\mu$m  [^38] The small image size—approximately 2 $\mu$m in the numerical case considered—places significant limitations on the use of a CCD as a detector.
-To address this, one promising alternative is photon detection using a microchannel plate photomultiplier (MCP-PMT).
-A feasible approach involves using a commercially available precision 2D translation stage with a resolution of 0.5 $\mu$m to control the position of a submicron aperture in an opaque screen.
-For instance, the aperture diameter could be set to 1 $\mu$m or smaller.
-The screen is mounted horizontally on the translation stage, which is positioned in the image plane.
-The MCP-PMT can then detect photons transmitted through the aperture.
-This setup offers the advantage of compactness, with the overall height of the optical configuration being approximately one meter..
+[…]
 Figure 42: Earth-based optical setup for detecting the aberration of light. The setup employs a two-lens system to form an image using an incoherent light source at optical frequencies.
@@
 The rounding or softening of the source’s edges is a characteristic effect of the convolution (”smoothing”) process.
-It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single-point source characteristics like oscillatory image shift within a period of one day.
-This argument gives the reason why our technique can be extended to allow for the reconstruction of the angle of aberration from a measurement of image shift of a finite size, incoherent source.
-The edge object has become useful in aberration shift detection.
-An alternative application of a ”point” source is an incoherent source with (physically) sharp edges.
-The edge width of the source is strictly related to the practical resolution achievable.
-For our particular example, the edge width should be about $1\mu$m.
+[…]
 11 The Aberration of Light from a Laser Source
@@
 The electric field vector in the plane resonator is presented as an oscillation superposition with different longitudinal wave numbers, $E_{x}+iE_{y}=\sum_{m}\widetilde{E}_{m}(x,y,t)\exp(-i\omega_{m}t)\sin(K_{m}z+\delta)$, where $K_{m}=m\pi/l-i/(nl)$, $\omega_{m}=m\pi c/l$, $\delta=i/n$, $n$ is the (complex) refractive index of the mirror matter ($|n|\gg 1)$, $l$ is the distance between the mirrors, and $m$ is an integer number ($m\gg 1$).
-It is evident that this expression satisfies Leontovich’s boundary conduction  [^39] An example of introducing approximate boundary conditions is presented here.
+It is evident that this expression satisfies Leontovich’s boundary conduction on the mirror surface when $z=0$ and $z=l$.
+[^41] An example of introducing approximate boundary conditions is presented here.
 These conditions, originally derived by Leontovich, apply to surfaces of materials with a high refractive index modulus.
 A more comprehensive theoretical treatment can be found in the monograph [48].
-on the mirror surface when $z=0$ and $z=l$.
 We assume the field change per one resonator pass to be small i.e.
@@
 In a plane resonator, the diffraction effects at the edges of the mirrors can be taken into account if one considers the space between the mirrors as a waveguide, and uses diffraction theory at the open end of the waveguide.
-If almost an integer number of half waves is laid along the length between the mirrors then at the open resonator end we may put boundary conditions on the amplitude function $\Phi(x,y)$, equivalent by its effect to the open waveguide end action  [^40] The analytical method is based on the introduction of specific, complex boundary conditions known as impedance boundary conditions of the resonance type [49].
-: $\Phi+(1+i)\beta_{0}\sqrt{cl/(4\omega_{m})}\partial\Phi/\partial\eta=0$, where $\eta$ is the direction of the normal to the imaginary side surface of the resonator, and the parameter $\beta_{0}=0.824$  [^41] Interestingly, $\beta_{0}$ is related to one of the most renowned mathematical functions—the Riemann zeta function $\zeta$.
-Specifically, it is given by $-\zeta(1/2)/\sqrt{\pi}$; see [49]..
+If almost an integer number of half waves is laid along the length between the mirrors then at the open resonator end we may put boundary conditions on the amplitude function $\Phi(x,y)$, equivalent by its effect to the open waveguide end action : $\Phi+(1+i)\beta_{0}\sqrt{cl/(4\omega_{m})}\partial\Phi/\partial\eta=0$, where $\eta$ is the direction of the normal to the imaginary side surface of the resonator, and the parameter $\beta_{0}=0.824$.
+[^42] The analytical method is based on the introduction of specific, complex boundary conditions known as impedance boundary conditions of the resonance type [49].
+Interestingly, $\beta_{0}$ is related to one of the most renowned mathematical functions—the Riemann zeta function $\zeta$.
+Specifically, it is given by $\zeta(1/2)/\sqrt{\pi}$; see [49].
 The boundary conditions do not depend on the polarization which means that polarization degeneracy occurs in the system under study.
 Therefore, the problem concerning the open plane resonator excitation using equivalent boundary conditions can be reduced to the more usual classical formulation of the problem of closed resonator excitation.
-[…]
+Let us consider a resonator equipped with plane circular mirrors with radius $R$.
+The distance between the mirrors is equal to $l$.
+The system is azimuthally symmetric relative to the resonator axis which coincides with the $z$ axis of the polar coordinates.
+We shall seek the solution of the wave equation in the following form: $\Phi(\vec{r})=\Phi_{\nu}(r)\cos(\nu\phi),\Phi_{\nu}(r)\sin(\nu\phi)$, where $\nu$ is an integer.The eigenfunctions of the resonator are the solution of the homogeneous equation $r^{2}d^{2}\Phi_{\nu j}/dr^{2}+rd\Phi_{\nu j}/dr+(k^{2}_{\nu j}r^{2}-\nu^{2})\Phi_{\nu j}=0$, satisfying the impedance boundary conditions.
+The eigenfunctions $\Phi_{\nu j}$ are orthogonal and form a complete system.
+In the first approximation for small parameter $M=1/\sqrt{N}$, where $N=\omega_{m}R^{2}/(cl)\gg 1$ can be referred to as the Fresnel number, the functions $\Phi_{\nu j}$ have the form $\Phi_{\nu}(r)=J_{\nu}(k_{\nu j}r)/J_{\nu+1}(\mu_{\nu j})$, where $\mu_{\nu j}$ is thejjth root of the Bessel function of order $\nu$, $k_{\nu j}=\mu_{\nu j}(1-\Delta)/R$, $\Delta=(1+i)\beta_{0}M/2$.
+Each mode for $\nu>0$ is four times degenerate.
+Firstly, for each value of the azimuthal number $\nu>0$ there are two linear independent functions $\propto\cos(\nu\phi)$ and $\propto\sin(\nu\phi)$.
+Secondly, because the boundary conditions do not depend on the polarization, two different field polarizations exist for each linear independent eigenfunction.
 The rigorous results of the three-dimensional theory of a plane Fabry-Perot resonator demonstrate that the radiation within the resonator can be represented as a set of discrete modes.
@@
 The distinction between a ”plane wave” emitter and a laser source leads to a striking prediction in the theory of light aberration.
-Specifically, if a ”plane wave” emitter initially at rest with respect to the fixed stars begins to move, the apparent angular position of the emitter—as observed from the accelerated frame—undergoes a sudden shift by an angle of $-v/c$ (see Fig.
+Specifically, if a ”plane wave” emitter initially at rest with respect to the fixed stars begins to move, the apparent angular position of the emitter—as observed from the accelerated frame—undergoes a sudden shift by an angle of v/c-v/c (see Fig.
 44).
@@
 When discussing light aberration, it is important to distinguish between the effect from a ”plane wave” emitter and that from the laser source.
-Consider a ”plane wave” emitter that is accelerated from rest to velocity $v$ in a direction perpendicular to its optical axis  [^42] It is assumed that after acceleration, the laser moves at constant velocity and operates in a steady-state regime..
+Consider a ”plane wave” emitter that is accelerated from rest to velocity $v$ in a direction perpendicular to its optical axis.
+[^43] It is assumed that after acceleration, the laser moves at constant velocity and operates in a steady-state regime.
 Suppose an observer, at rest with respect to the inertial frame of reference, measures the direction of light transport.
@@
 The inverse Galilean transformation $x\to x_{n}+vt$, $t\to t_{n}$ defines a new reference frame, $S_{n}$, which is uniformly accelerated in the $x$-direction, attaining a velocity $v$ relative to $S$.
-After the acceleration, the frame $S$ is, reciprocally, moving along the $x_{n}$-axis with velocity $-v$ relative to $S_{n}$.
+After the acceleration, the frame $S$ is, reciprocally, moving along the $x_{n}$-axis with velocity v-v relative to $S_{n}$.
 When the system $S_{n}$ begins moving with constant velocity, it becomes an inertial frame of reference.
@@
 So, where does the asymmetry arise?
-The analysis of the equivalence principle reveals that no additional complexities are required to explain the asymmetry between an inertial frame and an accelerated frame (relative to the fixed stars).
+The analysis of the equivalence principle reveals that no additional complexities are required to explain the asymmetry between an inertial frame and an accelerated frame.
 The resolution of this paradox lies in identifying a pseudo-gravitational potential within the system $S_{n}$, which acts as the source of the asymmetry.
@@
 To minimize the mathematical complexity of the discussion, we will restrict our analysis to terms up to second order in $v/c$.
-Consider a reference frame $S$, in which each point undergoes constant acceleration $g=-|g|$ in the negative $x_{n}$-direction for a duration $T=v/|g|$.
-In this setup, the system $S$ is effectively in free fall within a pseudo-gravitational field.
+Consider a reference frame $S$, in which each point undergoes constant acceleration g=|g|g=-|g| in the negative $x_{n}$-direction over a time interval $T=v/|g|$.
+In this configuration, frame $S$ is effectively in free fall within a pseudo-gravitational field experienced in the accelerating frame $S_{n}$.
 It becomes evident that the pseudo-gravitational acceleration is given by the gradient of a scalar potential, i.e., $g=-\partial\phi/\partial x_{n}$.
 A straightforward calculation shows that at time $t=T$, the coordinate transforms as $x_{n}=gT^{2}/2+x$ at $t=T$.
-Once the system $S$ transitions to uniform motion (constant velocity), the pseudo-gravitational field vanishes.
-To determine the pseudo-gravitational potential difference between the two frames, we compute $\phi=(\partial\phi/\partial x_{n})gT^{2}/2=-v^{2}/2$.
-Here, we have assumed that $\phi=0$ at $x_{n}=x$.
+Once the system $S$ transitions to uniform motion (i.e.
+constant velocity), the pseudo-gravitational field vanishes.
+To determine the pseudo-gravitational potential difference between the two frames, we compute $\phi=(\partial\phi/\partial x_{n})gT^{2}/2=-v^{2}/2$, where we have chosen the potential to be zero at $x_{n}=x$, i.e., $\phi=0$ at that point.
 The analysis of an accelerated frame using the equivalence principle does not involve a real gravitational field and therefore does not require the full framework of general relativity.
@@
 From the perspective of an observer located at $\phi_{2}$, the clock in the freely falling frame appears to tick faster, since it remains associated with the higher potential $\phi_{1}$, regardless of its instantaneous position.
-Radiation from a standard source within the free-falling frame will therefore appear blue-shifted relative to radiation from a standard source at rest at $\phi_{2}$ 434343 To understand the gravitational blue-shift of a light source in free fall, one must analyze the source’s emission mechanism and its behavior while in motion.
+Radiation from a standard source within the free-falling frame will therefore appear blue-shifted relative to radiation from a standard source at rest at $\phi_{2}$.
+444444 To understand the gravitational blue-shift of a light source in free fall, one must analyze the source’s emission mechanism and its behavior while in motion.
 The frequency of dipole radiation is proportional to the electron mass.
 In a gravitational field, mass varies with gravitational potential, according to the relation $m=m_{0}/\sqrt{1+2\phi/c^{2}}$.
-As a result, the electron mass is greater at points with lower gravitational potential..
+As a result, the electron mass is greater at points with lower gravitational potential.
 A similar time dilation effect occurs in accelerated frames.
@@
 These measurable differentials, $dt_{n}^{\mathrm{(p)}}$ and $dx_{n}^{\mathrm{(p)}}$, are linear combinations of the coordinate time differential $dt_{n}$ and the spatial differential $dx_{n}$, with coefficients determined by components of the metric tensor.
-In the new physical coordinates, the space-time interval takes the form: $ds^{2}=(cdt_{n}^{\mathrm{(p)}})^{2}$ $(dx_{n}^{\mathrm{(p)}})^{2}$ $=(1-v^{2}/c^{2})c^{2}dt^{2}_{n}-2vdx_{n}dt_{n}-dx_{n}^{2}$.
+In the new physical coordinates, the space-time interval takes the form: $ds^{2}=(cdt_{n}^{\mathrm{(p)}})^{2}$ - $(dx_{n}^{\mathrm{(p)}})^{2}$ $(1-v^{2}/c^{2})c^{2}dt^{2}_{n}-2vdx_{n}dt_{n}-dx_{n}^{2}$.
 The derivation of the Langevin metric in an accelerated frame—based on the concept of a pseudo-gravitational field—highlights a subtle but crucial distinction: the metric determines the form of physical laws, not the physical phenomena themselves.
@@
 After applying an inverse Galilean boost—where the coordinates transform as $x_{n}=x-vt,\quad t_{n}=t$ —the wavefront is found to be perpendicular to the vertical direction $z_{n}$, as depicted in Fig.
-15.
+14.
 Together, the Langevin metric and this initial condition account for the observed aberration of light in the accelerated frame.
 12.2 The Equivalence Principle and Gravitational Physics
-We should make one further remark regarding the equivalence principle.
-[^44] For a general discussion of the equivalence principle, we recommend reading [50].
-It is a common belief that the equivalence principle forms the foundation of general relativity.
-However, this view is not entirely accurate.
-The gravitational field is not merely a universal force acting identically on all bodies.
-According to Einstein’s insight, gravity is intimately connected to the structure of space-time itself.
-To grasp this idea, we must first understand the role of the Riemann curvature tensor, which, in a precise sense, is the gravitational field.
-Only when this tensor vanishes can we say that no gravitational field exists.
-This observation alone suffices to demonstrate that the equivalence principle, as traditionally formulated, does not universally apply.
-The equivalence principle holds for mechanical processes but fails when applied to electrodynamical phenomena.
-This distinction implies that it is indeed possible to determine, through internal measurements, whether a system is in an inertial frame or in free fall within a homogeneous gravitational field.
-For example, in such a field, a free electric charge undergoes uniform acceleration relative to an inertial frame and, as a result, emits radiation.
-This radiation occurs regardless of whether the field is strong or weak.
-Conversely, in a uniformly accelerated frame (absent gravity), the free charge remains unaccelerated relative to the inertial frame and, according to classical electrodynamics, does not radiate.
-Hence, the equivalence principle is explicitly violated in this context.
+We should add a further remark concerning the equivalence principle.
+[^45] For a general discussion of the equivalence principle, we recommend [50].
+It is commonly believed that the equivalence principle lies at the foundation of general relativity.
+However, this perspective is not entirely accurate.
+Gravity, as understood through Einstein’s insight, is not merely a universal force acting identically on all bodies—it is fundamentally tied to the geometry of space-time itself.
+To appreciate this idea, one must consider the role of the Riemann curvature tensor, which, in a precise mathematical sense, represents the gravitational field.
+The vanishing of this tensor is the necessary and sufficient condition for the absence of gravity.
+This fact alone shows that the traditional formulation of the equivalence principle has only limited applicability.
+While the equivalence principle holds for mechanical processes, it breaks down when applied to electrodynamical phenomena.
+This distinction implies that internal measurements can, in principle, distinguish between an inertial frame and a frame in free fall within a homogeneous gravitational field.
+For instance, consider a free electric charge in such a gravitational field.
+Relative to an inertial frame, it experiences uniform acceleration and consequently emits radiation—regardless of whether the gravitational field is weak or strong.
+In contrast, in a uniformly accelerated frame in flat space-time (i.e., in the absence of gravity), a freely falling charge does not accelerate relative to the inertial frame and, according to classical electrodynamics, does not radiate.
+Thus, in this context, the equivalence principle is explicitly violated.
 In the context of special and general relativity, we can examine earth-based experiments where pseudo-gravity and real gravity are intricately combined.
@@
 A central mathematical approach in the analysis of time dilation is the smooth matching—or tailoring—of the metric tensor.
-The transformation to the accelerated frame, with coordinates denoted by subscript ”n”, takes a Galilean form: $t_{n}=t$, $x_{n}=x-vt$ Following this transformation, the metric in the inertial frame transitions continuously into the Langevin metric of the accelerated inertial frame $S_{n}$ (see Chapter 5 for details).
+The transformation to the accelerated frame, with coordinates denoted by subscript ”n”, takes a Galilean form: $t_{n}=t$, $x_{n}=x-vt$ .
+Following this transformation, the metric in the inertial frame transitions continuously into the Langevin metric of the accelerated inertial frame $S_{n}$ (see Chapter 5 for details).
 In these new coordinates, the invariant interval given by Eq.
@@
 From the perspective of $S_{n}$, this setup is not symmetric, and the resulting time dilation experienced in the accelerated frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ differs.
-In frame $S_{n}$, frame $S_{n}^{(1)}$ moves with velocity $-v$, so the proper time in this frame is simply $d\tau_{n}^{(1)}=dt=dt_{n}$.
+In frame $S_{n}$, frame $S_{n}^{(1)}$ moves with velocity v-v, so the proper time in this frame is simply $d\tau_{n}^{(1)}=dt=dt_{n}$.
 Meanwhile, frame $S_{n}^{(2)}$ moves with velocity $v$ relative to $S_{n}$.
@@
 Now, let us consider two configurations of the light clock: one with its length parallel to the direction of motion and the other oriented orthogonally.
-According to the principle of relativity, all clocks—regardless of their construction—must measure time consistently in all inertial frames.
+According to the principle of relativity (i.e.
+pseudo-Euclidean spacetime geometry), all clocks—regardless of their construction—must measure time consistently in all inertial frames.
 We will demonstrate that, regardless of the orientation of the light clock relative to its motion, the elapsed time measured remains the same.
@@
 As a result, the length of a rod at rest in the accelerated frame generally differs from that of a rod at rest in the inertial frame.
-The spatial geometry in an accelerated frame is described by the differential spatial line element defined within it.
-While spatial geometry possesses a degree of conventionality due to coordinate dependence, the proper spatial line element, denoted $dl_{n}$, holds particular significance, as it defines a coordinate-independent spatial geometry in the accelerated frame.
-The relation $dl_{n}^{2}=(-g_{\alpha\beta}+g_{0\alpha}g_{0\beta}/g_{00})dx_{\alpha}dx_{\beta}$ establishes the connection between the three-dimensional spatial line element and the four-dimensional spacetime metric.
-It can be shown that $dl_{n}$ remains invariant under coordinate transformations within the same reference frame.
-In contrast, the spatial line element $dl$, obtained by setting $dt_{n}=0$ in Eq.(15) with $dl^{2}=dx_{n}^{2}$, differs from $dl_{n}$.
+The spatial geometry in an accelerated frame is characterized by the differential spatial line element defined within that frame.
+Although spatial geometry is inherently coordinate-dependent and thus exhibits a degree of conventionality, the proper spatial line element, denoted by $dl_{n}$, holds particular significance.
+It defines a coordinate-independent notion of spatial distance within the accelerated frame.
+By contrast, the spatial line element $dl$, obtained by setting $dt_{n}=0$ in Eq.
+(15) such that $dl^{2}=dx_{n}^{2}$, differs from $dl_{n}$.
+Now, how do physical notions of length and time manifest in the coordinate system $(t_{n},x_{n},y_{n},z_{n})$?
+In an accelerated frame, consider the spacetime interval given by $ds^{2}=c^{2}g_{00}dt^{2}+2cg_{01}dxdt+g_{11}dx^{2}$.
+We can decompose this expression as follows: $ds^{2}=c^{2}[\sqrt{g_{00}}dt+g_{01}dx/(c\sqrt{g_{00}})]^{2}+(g_{11}-g_{01}^{2}/g_{00})dx^{2}$.
+Here, the quantity $dT^{2}=[\sqrt{g_{00}}dt+g_{01}dx/(c\sqrt{g_{00}})]^{2}$ has a time-like character and represents the square of physical time in the accelerated frame, while $dX^{2}=(g_{11}-g_{01}^{2}/g_{00})dx^{2}$ is space-like and corresponds to the square of physical length.
+These quantities are directly tied to the spacetime interval in a manner analogous to the Minkowski metric in special relativity [16].
+They are thus physically meaningful and invariant under coordinate transformations within the accelerated frame.
+The relation $dl_{n}^{2}=(-g_{11}+g_{01}g_{01}/g_{00})dx^{2}$ establishes the connection between the three-dimensional spatial line element and the four-dimensional spacetime metric.
 Using Eq.
@@
 This method preserves Einstein synchronization, established via light signals emitted from a stationary dipole source.
-Within this framework of absolute simultaneity, the inertial observer perceives the accelerated observer as measuring the speed of light along the $x$-axis as $c-v$ in the positive direction and $-c-v$ in the negative direction.
+Within this framework of absolute simultaneity, the inertial observer perceives the accelerated observer as measuring the speed of light along the $x$-axis as cvc-v in the positive direction and cv-c-v in the negative direction.
 Furthermore, the physical meter stick (with mirrors mounted on it), as measured by the accelerated observer, appears compressed along the direction of motion relative to the fixed stars.
@@
 12.5 Metrics in Accelerated Frames
-Let us examine the implications of assuming that the inertial frame has no history of acceleration and that the space-time metric, as perceived by an observer at rest, is diagonal.
-We begin by considering two emitters within this inertial frame: the first is stationary, while the second moves with velocity $v$.
-Applying a Galilean boost yields the metric of the accelerated emitter in the inertial frame, as given in Eq.
+In this section, we continue our exploration of non-reciprocity in special relativity.
+We begin by examining the implications of assuming that the inertial frame has no history of acceleration and that the space-time metric, as perceived by an observer at rest, is diagonal.
+Consider two emitters within this inertial frame: the first remains stationary, while the second moves with velocity $v$.
+A Galilean boost transforms the metric associated with the moving emitter, resulting in the form given in Eq.
 (11).
-In this coordinate system, the non-accelerated light source is described by the standard diagonal metric: $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$.
-Let us now return to the discussion of the accelerated frame.
-In this frame, we observe a reciprocal velocity: the second emitter is at rest, while the first emitter moves with velocity $-v$.
-Now what about reciprocity in physics?
-According to conventional theory, acceleration does not inherently break the symmetry of motion between an accelerated frame and an inertial one.
-Consequently, an observer in the accelerated frame should still obtain a diagonal metric for the second emitter, which is at rest relative to this frame.
-By the principle of reciprocity, the metric given in Eq.
-(15) should also describe the aberration of light as observed for the first emitter, which appears to be moving in the accelerated frame.
-In fact, the metric of the moving emitter in the accelerated frame matches that of the moving emitter in the inertial frame, as given by Eq.
-(11).
-Reciprocity implies that the only difference lies in the sign of the velocity $v$.
+In the coordinates of the inertial frame, the non-accelerated emitter is described by the standard diagonal Minkowski metric: $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$.
+Now, let us shift our attention to the accelerated frame in which the second emitter is at rest, and the first appears to move with velocity v-v.
+This raises a fundamental question: Does the principle of reciprocity still apply in this context?
+Conventional theory holds that acceleration does not intrinsically violate the symmetry of motion between inertial and non-inertial frames.
+Therefore, an observer in the accelerated frame should also describe the metric of the stationary (second) emitter as diagonal.
+By reciprocity, the moving (first) emitter should then be described in this frame by a metric analogous to Eq.
+(11), differing only by the sign of $v$.
+However, this leads to a deeper inquiry into the principle of reciprocity itself.
 The principle of reciprocity states that all inertial frames are equivalent.
 However, this relativity principle does not hold and the Lorentz covariance of all physical equations does not dictate the reciprocal symmetry of nature.
-We will examine the asymmetry between inertial and accelerated reference frames.
-A space-time geometrical analysis of the accelerated frame reveals a fundamental non-reciprocity between it and the initial inertial frame.
-Specifically, within the accelerated frame, we derive the Langevin metric, Eq.
-(15), associated with the second emitter.
-This metric governs the behavior of electrodynamics in the accelerated frame.
-Our analysis highlights a crucial asymmetry: Maxwell’s equations, which hold in inertial frames, are no longer valid from the perspective of an observer at rest in the accelerated frame.
-This breakdown reflects a deeper distinction between the two types of frames.
-Importantly, because we adopt an absolute time coordinatization, the initial conditions—such as the orientation of radiation wavefronts—remain identical in both frames  [^45] Many physicists contend that the universality of Maxwell’s equations across inertial frames rules out the existence of any privileged frame.
-As Dieks observes [51], “As we already mentioned, it is a basic principle of the special theory of relativity that the line element $ds$ supplies all information about the physics of the situation, as described in the given coordinates.” However, this widely held view is misleading.
-The metric alone does not fully specify the physical situation.
-To solve electrodynamics problems, one must also provide appropriate initial conditions..
-We now turn to an apparent paradox.
-The principle that all inertial frames are equivalent is often supported by the following argument: To maintain Lorentz coordinates in an accelerated frame, one must perform a diagonalization of the metric tensor.
-In both the original and the accelerated frames, the metric appears diagonal, and standard treatments suggest that the coordinates in these frames should be related by a Lorentz transformation.
-At first glance, the situation appears symmetric—both frames seem equivalent.
-However, a crucial asymmetry emerges when we consider the physical distinction between an inertial frame and one that has undergone acceleration.
-Specifically, we should expect the metric tensor to reflect this difference.
-After the diagonalization procedure, the metric tensor in the accelerated frame must transition discontinuously—from a form such as $g_{00}=(1-v^{2}/c^{2})<1,g_{01}=-v/c$ to the standard inertial form $g_{00}=1,g_{01}=0$.
-This discontinuity implies that the Lorentz transformation does not preserve the smooth structure of the metric across frames.
-As a result, the symmetry between inertial frames breaks down in this context.
-This reveals a subtle yet significant distinction: while the mathematical form of the metric may appear similar in both frames, the physical histories of the frames—particularly the presence or absence of acceleration—lead to observable and meaningful differences.
-In this way, the accelerated frame retains a kind of “memory” of its non-inertial past, breaking the naive equivalence suggested by symmetry arguments.
+A geometric analysis of space-time in the accelerated frame reveals this asymmetry.
+Specifically, within the accelerated frame, the Langevin metric—given in Eq.
+(15)—emerges as the correct description of the second emitter’s space-time.
+This metric governs electrodynamics in the accelerated frame and departs from the simple diagonal form of inertial frames.
+This departure has significant consequences: Maxwell’s equations, valid in inertial frames, no longer retain their form for an observer at rest in the accelerated frame.
+This breakdown signals a deeper structural difference between inertial and accelerated reference frames.
+Notably, because we adopt an absolute time coordinatization, the initial conditions—such as the orientation of radiation wavefronts—remain identical in both frames.
+We now arrive at a paradox.
+The equivalence of all inertial frames is often supported by the claim that, after an appropriate coordinate transformation (e.g., diagonalization of the metric tensor), both the inertial and accelerated frames yield diagonal metrics.
+This seems to imply symmetry: both frames appear to be related by a Lorentz transformation and should be physically equivalent.
+[^46] Many physicists argue that the universality of Maxwell’s equations across inertial frames rules out the existence of any privileged frame.
+As Dieks notes [51], “As we already mentioned, it is a basic principle of the special theory of relativity that the line element $ds$ supplies all information about the physics of the situation, as described in the given coordinates.”
+However, a key asymmetry becomes evident when we consider the physical history of the frames.
+While the inertial frame remains unaccelerated, the accelerated frame retains a memory of its dynamical past.
+This memory is encoded in the metric structure.
+After diagonalization, the metric tensor in the accelerated frame undergoes a discontinuous shift—from a form such as $g_{00}=(1-v^{2}/c^{2})<1,g_{01}=-v/c$ to the standard inertial values: $g_{00}=1,g_{01}=0$.
+This discontinuity highlights the fact that the Lorentz transformation does not preserve the smoothness of the metric across such transitions.
+Hence, the apparent symmetry between frames breaks down when physical considerations—specifically, the history of acceleration—are taken into account.
+This leads us to a subtle but important conclusion: although the mathematical form of the metric may appear similar in both frames, the physical context—particularly the presence or absence of past acceleration—introduces observable differences.
+The accelerated frame thus retains a “memory” of its non-inertial history, breaking the naive equivalence suggested by purely kinematic symmetry arguments.
 13 Relativistic Measurements
@@
 This result can be interpreted as follows: after traveling a distance of $\gamma v\tau_{0}$, the number of muons observed in the lab frame is reduced to half of the original population.
-In this context, the time measured corresponds to a ”length,” as defined by the earlier metric discussion.
+In this context, the time measured corresponds to a ”length”.
 This leads to an intriguing question: could this interpretation apply to all time measurements within the framework of special relativity?
@@
 Using Eq.
-(11), we find $dl_{i}^{2}=dx^{2}/(1-v^{2}/c^{2})$, which defines the spatial geometry in the context of the Langevin metric.
+(11), we find $dl_{i}^{2}=dx^{2}/(1-v^{2}/c^{2})$, which defines the spatial geometry in the context of the non-diagonal metric.
 This result reflects the Lorentz contraction of measuring rods.
@@
 When measuring the resonance absorption for the same relative velocity between the source and the observer, a blue shift is detected when the source is positioned at the center of the disk, while a red shift appears when the source is located at the rim.
-This result reveals an asymmetry in the transverse Doppler effect [^46] The first experimental confirmation of time dilation via resonance absorption of gamma rays was reported by Champeney, Isaak, and Khan in 1965 [52]..
+This result reveals an asymmetry in the transverse Doppler effect.
+[^47] The first experimental confirmation of time dilation via resonance absorption of gamma rays was reported by Champeney, Isaak, and Khan in 1965 [52].
 Notably, when both the source and observer are situated at the rim, no frequency shift is observed.
@@
 The fundamental concept behind the experiment can be described as follows.
-Consider two opposite point charges $+e$ and $-e$ in the Earth-based frame; the radius-vector pointing from $-e$ to $+e$ be denoted by $\vec{r}_{n}$.
-If the charges are at rest the force acting upon $+e$ can be written $\vec{F}_{n}=e\vec{E}=-e^{2}\vec{r}_{n}/r^{3}_{n}$.
+Consider two opposite point charges +e+e and e-e in the Earth-based frame; the radius-vector pointing from e-e to +e+e be denoted by $\vec{r}_{n}$.
+If the charges are at rest the force acting upon +e+e can be written $\vec{F}_{n}=e\vec{E}=-e^{2}\vec{r}_{n}/r^{3}_{n}$.
 As the force acts in the direction $\vec{r}_{n}$ the moment of force produced by the pair of charges vanishes, $\vec{M}_{n}=\vec{r}_{n}\times\vec{F}_{n}=0$.
-If the pair of charges is made to move with a velocity $\vec{v}$ in the Sun-centered frame then the positive charge will be under the action of the Coulomb attraction of $-e$ and also under the influence of the magnetic field $\vec{B}=e(\vec{v}\times\vec{r})/(cr^{3})$.
+If the pair of charges is made to move with a velocity $\vec{v}$ in the Sun-centered frame then the positive charge will be under the action of the Coulomb attraction of e-e and also under the influence of the magnetic field $\vec{B}=e(\vec{v}\times\vec{r})/(cr^{3})$.
 Thus the total force acting upon $e$ is given by $\vec{F}+e^{2}(\vec{v}\times(\vec{v}\times\vec{r}))/(c^{2}r^{3})$.
@@
 The absence of torque on the capacitor in the Earth-based frame is fully consistent with special relativity and can be explained in a straightforward manner.
-The electrodynamic equations in the Earth-based frame can be derived by applying a Galilean transformation (with velocity $-v$) to Maxwell’s equations.
+The electrodynamic equations in the Earth-based frame can be derived by applying a Galilean transformation (with velocity v-v) to Maxwell’s equations.
 Assuming the magnetic field $\vec{B}=0$ in the original frame, the transformed fields are given by $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$, where $\vec{v}$ is the velocity of the Earth relative to the Sun-based frame.
@@
 This leads to an apparent paradox: the mechanical equations governing force and torque appear to differ between inertial frames.
-Specifically, a torque—and thus a time-dependent change in three-dimensional angular momentum—exists in one inertial frame but not in another that is in relative motion  [^48] In [55], Jackson discusses the apparent paradox that different mechanical equations for force and torque govern the motion of a charged particle in different inertial frames.
+Specifically, a torque—and thus a time-dependent change in three-dimensional angular momentum—exists in one inertial frame but not in another that is in relative motion.
+[^49] In [55], Jackson discusses the apparent paradox that different mechanical equations for force and torque govern the motion of a charged particle in different inertial frames.
 A similar paradox arises in the standard explanation of the Trouton–Noble experiment.
 In the usual approach [55], this paradox stems from the hidden assumption that all inertial frames share a common three-dimensional space.
-However, we demonstrate that a relativistically accurate description of the Trouton–Noble experiment can be achieved only by taking into account the Wigner rotation effect..
+However, we demonstrate that a relativistically accurate description of the Trouton–Noble experiment can be achieved only by taking into account the Wigner rotation effect.
 We argue that the root of this paradox lies in a flawed assumption: that an Earth-based observer and a Sun-based observer share a common three-dimensional space.
@@
 The key point, however, is that this interval is entirely determined by the motion of a clock within its initial inertial frame.
-In 1971, Hafele and Keating conducted a landmark experiment in which four atomic clocks were synchronized with a laboratory reference clock and then flown around the Earth aboard commercial aircraft—two traveling eastward and two westward  [^49] This experiment provided the first experimental confirmation of time dilation as predicted by special relativity, using atomic clocks in circular motion.
+In 1971, Hafele and Keating conducted a landmark experiment in which four atomic clocks were synchronized with a laboratory reference clock and then flown around the Earth aboard commercial aircraft—two traveling eastward and two westward.
+[^50] This experiment provided the first experimental confirmation of time dilation as predicted by special relativity, using atomic clocks in circular motion.
 An important follow-up came from the Global Positioning System (GPS), which also involves relativistic effects.
@@
 This suggests that relative velocity alone does not account for time dilation.
-Furthermore, the experiment revealed no effect of Earth’s orbital velocity on clock rates, a result attributed to the fact that both the satellites and Earth-based clocks share a common orbital path around the Sun, rendering orbital effects unobservable in this context [56, 57]..
+Furthermore, the experiment revealed no effect of Earth’s orbital velocity on clock rates, a result attributed to the fact that both the satellites and Earth-based clocks share a common orbital path around the Sun, rendering orbital effects unobservable in this context [56, 57].
 Upon return, the airborne clocks were compared with the laboratory reference clock.
@@
 Notably, the experiment found no measurable influence of Earth’s orbital motion on clock rates, reinforcing the idea that time dilation effects are tied to local motion relative to the chosen inertial frame.
-[…]
+The null result of the experiment described above can be understood by considering that the setup involves a circular path geometry.
+A more careful analysis reveals that it is essential to account for the fact that the moving clocks are ultimately compared to a reference clock located at the same point and at rest relative to the Earth.
+The moving clock travels around the reference clock along a circular path with velocity $w$.
+We will analyze the effect by expanding terms up to second order in $w/c$ and $v/c$, where $v$ is the Earth’s orbital velocity.
+The inner rate of the moving clock depends on velocity $w$, so that $d\tau=dt_{n}[1-v^{2}/(2c^{2})-w^{2}/(2c^{2})+\vec{v}\cdot\vec{w}(t_{n})/c^{2}]$.
+Since the reference clock is at rest in the Earth-based frame, the reading of its proper time is $d\tau=dt_{n}(1-v^{2}/2)$.
+Analyzing the geometry of the situation, we see that for motion along a closed circular path, the cross term averages out over a complete cycle.
+Specifically, the integral of the mixed term vanishes: $\int\vec{v}\cdot\vec{w}(t_{n})dt_{n}=\int\vec{v}\cdot d\vec{s}=0$.
+Therefore, no observable effect from the Earth’s orbital velocity remains, which explains the null result.
 15 The Principle of Relativity and Modern Cosmology
@@
 It asserts that the same laws must hold in all inertial frames, which we interpret to mean that these laws should be expressed by equations that retain the same form in all inertial frames.
-In special relativity, the transformations preserving this form-invariance between inertial frames are the Lorentz transformations  [^50] In the special theory of relativity the form-invariant transformations between inertial frames are Lorentz transformations [58].
+In special relativity, the transformations preserving this form-invariance between inertial frames are the Lorentz transformations.
+[^51] In the special theory of relativity the form-invariant transformations between inertial frames are Lorentz transformations [58].
 For a general discussion about the Principle of Relativity we suggest reading the paper [59].
-.
 According to the special principle of relativity, all inertial frames are equivalent concerning physical laws, but not necessarily concerning physical facts.
@@
 to be around three times larger than that of CMB).
-From the Hubble diagram of quasars, the motion of the solar system is derived, which is out to be the largest value ever found, $8\cdot 10^{3}$ km/s  [^51] Recent observations of the cosmic dipole anisotropies is revolutionized our understanding of the universe.
-The various dipoles, including CMB dipole, all pointing along the same direction, suggest a preferred direction in the universe, raising thereby uncomfortable questions about the cosmological principle, the basis of the standard model in modern cosmology [60, 61, 62]..
+From the Hubble diagram of quasars, the motion of the solar system is derived, which is out to be the largest value ever found, $8\cdot 10^{3}$ km/s.
+[^52] Recent observations of the cosmic dipole anisotropies is revolutionized our understanding of the universe.
+The various dipoles, including CMB dipole, all pointing along the same direction, suggest a preferred direction in the universe, raising thereby uncomfortable questions about the cosmological principle, the basis of the standard model in modern cosmology [60, 61, 62].
 On the other hand, a common direction for all these dipoles, determined from completely independent surveys by different groups employing different techniques, indicates that these dipoles are not resulting from some systematics in the observations or the data analysis, but could instead suggest an inherent anisotropy.
@@
 (21), represents a relativistic extension of Newton’s second law.
-While the classical form $md\vec{v}/dt=\vec{f}$ remain valid in the particle’s instantaneous comoving Lorentz frame, the relativistic formulation embeds these three spatial equations into the four-dimensional Minkowski spacetime  [^52] The notion of embedding relies on the fact that Newton’s second law can always be applied in any Lorentz frame in which the particle is instantaneously at rest [64].
+While the classical form $md\vec{v}/dt=\vec{f}$ remain valid in the particle’s instantaneous comoving Lorentz frame, the relativistic formulation embeds these three spatial equations into the four-dimensional Minkowski spacetime.
+[^53] The notion of embedding relies on the fact that Newton’s second law can always be applied in any Lorentz frame in which the particle is instantaneously at rest [64].
 That is, if a comoving Lorentz frame is defined at a given instant, the particle’s evolution can be accurately predicted in this frame over an infinitesimal time interval.
@@
 The presence of $\hat{P}_{\perp}$ highlights that only three of the four components of the covariant equation are independent.
-.
 The immediate generalization of $md\vec{v}/dt=\vec{f}$ to an arbitrary Lorentz frame is given by Eq.
@@
 The calculation carried out in the case of a spatially extended object shows that the temporal coincidence of two distant events has absolute character: $\Delta\tau=0$ implies $\Delta t=0$.
-It should be noted that usually the notation ”$\tau$” arose from the proper time.
+It should be noted that usually the notation $\tau$” arose from the proper time.
 Now let us recall the standard concept of the object’s proper time.
@@
 After two successive integrations we have $X^{\mu}(\tau)=X^{\mu}(0)+\gamma c\tau e^{\mu}_{0}+R[e^{\mu}_{2}\sin(\omega\tau)-e^{\mu}_{3}\cos(\omega\tau)]$ where $R=\gamma v/\omega$.
-This enables us to find the time dependence $[0,X^{(2)}(t),X^{(3)}(t)]$ of the particle’s position since $t/\gamma=\tau$.
+This enables us to find the time dependence $0,X^{(2)}(t),X^{(3)}(t)]$ of the particle’s position since $t/\gamma=\tau$.
 From this solution, we conclude that the motion of a charged particle in a constant magnetic field is indeed uniform circular motion, consistent with classical expectations, yet derived within the fully covariant relativistic framework.
-One might expect that the particle’s trajectory in the lab frame, derived from the previous reasoning $[0,X^{(2)}(t),X^{(3)}(t)]$, should be identified with $\vec{x}_{cov}(t)$.
-However, the trajectory $[0,X^{(2)}(t),X^{(3)}(t)]$ does not include relativistic kinematics effects.
+One might expect that the particle’s trajectory in the lab frame, derived from the previous reasoning $0,X^{(2)}(t),X^{(3)}(t)]$, should be identified with $\vec{x}_{cov}(t)$.
+However, the trajectory $0,X^{(2)}(t),X^{(3)}(t)]$ does not include relativistic kinematics effects.
 We have found that the standard integration of the four-dimensional covariant equation of motion, as shown in Eq.
@@
 Since the curvature radius of the trajectory in a magnetic field—and therefore the three-momentum—has an objective (i.e., convention-invariant) meaning, both approaches yield the same physical results.
-16.6 Lorentz-Covariant Particle Tracking
+16.6 Einstein’s Velocities Addition Vs. Covariant Addition
 In conventional particle tracking, the trajectory of a particle, denoted by $\vec{x}(t)$, is often described from the laboratory frame as the result of successive Galilean boosts.
@@
 This is a subtle but important misconception.
+This is a good point to make some general remarks about Einstein’s velocity addition.
+It should be noted that presented above the commonly accepted derivation of velocity composition does not follow from the composition of Lorentz boosts.
+In particular, second boost from $K^{\prime}$ to particle proper frame does not Lorentz boost.
+Einstein’s velocity addition is based on the relation $dt^{\prime}=dt/\gamma$, which implicitly assumes an absolute time convention for second boost.
+However, it does not account for the interplay between transverse positions $y^{\prime}$ and time $t^{\prime}$, which is inherent in Lorentz boost matrices.
+In contrast, every Lorentz boost represents a transformation between reference frames within a Lorentzian framework (i.e., under Lorentz coordinatization including also transverse direction).
+While Lorentz boosts intertwine space and time coordinates, Einstein’s velocity addition relies on the assumption of a shared three-dimensional space across all inertial frames.
+The consequence of this difference is the Wigner rotation accompanying the composition of non-parallel 3D velocities.
 One of the consequences of the non-commutativity of non-collinear Lorentz boosts is the absence of a shared, common notion of ordinary space.
 Suppose that in the Lorentz frame $K^{\prime}$, a traveler observes a particle moving along the $y^{\prime}$ axis.
-In other words, the particle has a velocity component $v^{\prime}_{y}$ in this frame, while the frame $K^{\prime}$ itself is moving with velocity $V$ along the $x$ axis relative to another Lorentz frame $K$.
-To an observer at rest in frame $K$, the moving frame $K^{\prime}$ appears rotated with respect to $K$.
+In other words, the particle has a velocity component $v^{\prime}_{y}$ in this frame, while the frame $K^{\prime}$ itself is moving with velocity $V$ along the $x$ axis relative to initial Lorentz frame $K$.
+The proper frame is rotated by an angle $\theta_{w}=v^{\prime}_{y}/(\gamma c)$ relative to frame $K$.
+The coordinate axes of the proper frame remain parallel to those of frame $K^{\prime}$.
+From the perspective of frame $K$, the coordinate axes of frame $K^{\prime}$ are also rotated by the angle $\theta_{w}$ with respect to those of frame $K$.
 Let us consider the simple case where the boost velocity $V$ approaches the speed of light (i.e., the ultrarelativistic limit), while the transverse velocity $v^{\prime}_{y}$ remains small.
@@
 In this regime, the axes of the $K^{\prime}$ frame appear rotated with respect to those of $K$ by an angle $\theta_{w}=v^{\prime}_{y}/(\gamma c)$.
-Notably, this rotation is in the same direction as the rotation of the particle’s velocity vector in $K$  [^54] To determine the rotation angle in the given problem, we consider a composition of Lorentz boosts.
-Two successive boosts—first from frame $K$ to $K^{\prime}$, and then an infinitesimal boost to the particle’s proper frame—are equivalent to a single boost from $K$ directly to the proper frame, followed by a rotation.
-The proper frame is rotated by an angle $d\theta_{w}=dv^{\prime}_{y}/(\gamma c)$ relative to frame $K$.
-The coordinate axes of the proper frame remain parallel to those of frame $K^{\prime}$.
-From the perspective of frame $K$, the coordinate axes of frame $K^{\prime}$ are also rotated by the angle $d\theta_{w}$ with respect to those of frame $K$.
+Notably, this rotation is in the same direction as the rotation of the particle’s velocity vector in $K$.
 We now consider the physical consequences of this rotation.
@@
 Our result contradicts this claim.
-The standard derivation of velocity addition relies on the relation $dt^{\prime}=dt/\gamma$, but neglects the mixing of transverse spatial coordinates with time in Lorentz transformations.
-This omission leads to an incorrect conclusion about the magnitude of the resulting velocity.
-16.7 Relativistic Velocity Composition and the Wigner Rotation
-Solving problems related to covariant particle tracking presents two major challenges.
-The first challenge arises from the conventional approach to particle tracking within a single inertial frame, which is based entirely on classical Newtonian mechanics, without incorporating Lorentz transformations.
-Standard textbooks assert that when only one frame is involved, relativity is unnecessary.
-In this framework, the only relativistic modification is the introduction of relativistic mass, while space and time remain distinct, with no mixing of spatial and temporal components.
-Surprisingly, this crucial aspect has received little attention within the physics community.
-The second challenge stems from confusion surrounding the composition of non-parallel velocities in special relativity.
-In Newtonian mechanics, velocity addition follows the parallelogram rule, treating velocities as ordinary vectors.
-However, in relativistic kinematics, velocity composition must be derived from the combination of Lorentz boosts.
-A well-known issue is that Lorentz boosts alone do not form a group, which raises a fundamental concern: if a set of transformations lacks group structure, the resulting theory is mathematically incomplete.
-Textbook authors often assume that the principle of relativity demands that Einstein’s velocity addition law obeys group composition.
+The standard derivation of velocity addition neglects the mixing of transverse spatial coordinate $y^{\prime}$ with time $t^{\prime}$ in Lorentz transformations.
+This omission leads to an incorrect conclusion about the magnitude of the resulting velocity in initial inertial frame $K$.
+It is important to highlight another effective approach to covariant particle tracking.
+Here, we revisit the motion of an accelerated relativistic particle from the perspective of an inertial lab observer—without changing reference frames.
+Specifically, we consider a particle undergoing acceleration within the lab frame, representing an example of an active boost.
+In this scenario, we track the system’s evolution entirely within a single (lab) reference frame.
+The simplest synchronization method maintains the same set of uniformly synchronized clocks used before the acceleration, without any adjustments.
+This approach is based on the absolute time (or absolute simultaneity) convention, which preserves simultaneity across space.
+Under this convention, the standard Galilean velocity addition rule can be applied instant by instant to track the particle’s motion as it accelerates.
+As discussed in Chapter 3, absolute time coordinatization can be transformed into Lorentz coordinatization.
+By combining Galilean transformations with appropriate changes of variables, we can derive the Lorentz transformation within the framework of absolute time.
+To further illustrate this idea, we now examine how distant lab clocks resynchronize during the particle’s acceleration.
+This analysis reveals a direct connection between the transformation of particle velocity (in Lorentz coordinates) and time dilation.
+As previously emphasized, the Lorentz coordinate system is a conceptual construct.
+The notion of synchronized clocks in this system is purely hypothetical, used primarily for the application of Maxwell’s equations in radiation calculations.
+Suppose that, before acceleration, we choose a Lorentz coordinate system in the lab frame.
+Immediately after the acceleration, particle velocity changes by an infinitesimal amount $d\vec{v}$ along the $y$-axis.
+If clock synchronization is fixed, this corresponds to adopting the absolute time convention.
+To maintain Lorentz coordinates in the lab frame—as discussed previously—we must perform a clock resynchronization by introducing an infinitesimal time shift.
+The simplest case arises when the $v_{y}$ is very small, allowing us to work up to second-order terms in $v_{y}^{2}/c^{2}$.
+This restriction greatly simplifies the calculations for two reasons.
+First, relativistic corrections due to the composition of non-collinear velocity increments only appear at order $v_{y}^{3}/c^{3}$ and can be neglected.
+Second, time dilation effects also enter only at higher orders.
+To maintain a Lorentz coordinate system in the lab frame after the acceleration, a resynchronization of clocks is required.
+In this context, operations involving the rule clock structure in the lab frame can be interpreted as a change of variables governed by the transformation in Eq.
+13: $y_{L}=\gamma_{y}y$, $t_{L}=(t/\gamma_{y}+\gamma_{y}yv_{y}/c^{2})$, where $\gamma_{y}=1+v_{y}^{2}/(2c^{2})$ and the particle’s displacement is given by $y=v_{y}t$.
+This transformation effectively rescales the time variable—adjusting the synchronization of all clocks from $t$ to $\gamma_{y}t$, with $\gamma_{y}\simeq 1+\theta_{k}^{2}/2$.
+As a result, we see that the particle’s speed after acceleration is unaffected by its transverse motion.
+Although no second-order relativistic correction appears in the transverse ($y$-axis) velocity component, the longitudinal component is modified.
+Specifically, the longitudinal velocity $V$ changes to $V/\gamma_{y}$ with $v_{z}=V-v_{y}^{2}/(2c)$.
+Thus, after clock resynchronization, the total electron speed in the lab frame remains VV—a result that aligns perfectly with the outcome derived earlier using a sequence of Lorentz transformations in changing frames.
+All pieces fit together seamlessly.
+We will explore the topic of relativistic velocity composition in greater detail in the next chapter
+17 Mathematical Analysis of Relativistic Velocity Composition
+In this chapter, we shall explore the subject of relativistic velocity addition more mathematically.
+Textbook authors often assume that the principle of relativity requires Einstein’s velocity addition law to obey group composition.
 However, the non-associativity of non-collinear velocity addition is not widely recognized.
-Many physicists, having learned relativity from standard textbooks, might find this claim surprising, as Einstein’s velocity addition is often assumed to be associative.
-Yet a straightforward algebraic exercise demonstrates its non-associativity.
-As a result, constructing a covariant trajectory, $\vec{x}_{cov}(t)$, using only Lorentz boosts is impossible—additional elements are required.
-The root cause of this issue lies in the presence of Wigner rotation.
+Many physicists, having studied relativity from standard textbooks, might find this claim surprising, since Einstein’s velocity addition is commonly assumed to be associative.
+Yet, a straightforward algebraic exercise reveals its non-associative nature.
+The root cause of this issue lies in the presence of the Wigner rotation.
 In fact, the breakdown of commutativity and associativity in Einstein’s velocity addition law naturally manifests as Wigner rotation, offering deeper insight into the structure of relativistic transformations.
-The next section will explore the relativistic transformation of velocities in greater detail.
-16.8 Mathematical Analysis of Relativistic Velocity Composition
-Above, we derived specific quantitative results for the relativistic combination of velocities using only elementary concepts from special relativity.
-However, some readers may find a more mathematically rigorous explanation clearer.
-As a consequence of the space-time geometry, Lorentz coordinate systems are connected by Lorentz transformations.
+17.1 Parametrization of Lorentz Transformations
+Einstein formally defined the special theory of relativity through two explicit postulates.
+Minkowski, in 1908, recognized that these postulates could be reformulated as a single geometric axiom: spacetime possesses a pseudo-Euclidean geometry.
+As a consequence of this spacetime geometry, Lorentz coordinate systems are related via Lorentz transformations.
 Consider a four-vector, $x=(x_{0},\vec{x})$, with squared-length $x^{2}=\eta_{ij}x^{i}x^{j}=(x_{0})^{2}-|\vec{x}|^{2}$ where $\eta_{ij}=\mathrm{diag}[1,-1,-1,-1]$ is the Minkowski space-time metric.
 Since Lorentz transformations preserve the space-time interval, they satisfy $x^{\prime}=\Lambda x$ such that $\eta_{ij}=\Lambda^{m}_{i}\Lambda^{n}_{j}\eta_{mn}$.
-This equation defines the most general 4 × 4 Lorentz transformation matrix $\Lambda$.
-The set of all such matrices forms a symmetry-group of the Minkowski metric under matrix multiplication, denoted by $O(1,3)$.
-In particular, $O(1,3)$ is a Lie group, appropriately called the Lorentz group.
-It is widely accepted that relationship between inertial frames are completely determined by their relative velocities and relative orientations.
-Consequently, the generic Lorentz transformation denoted as $\Lambda(\vec{v},\vec{\theta})$, is parameterized by the velocity $\vec{v}$ and the rotation parameter $\vec{\theta}$, representing a simultaneous boost and rotation.
-Here, $\vec{v}$ is a vector in three-dimensional Euclidean space, wile $\vec{\theta}$ belongs to the group $O(3)$, which consists of all possible spatial rotations in three-dimensions.
-This parameterization of the Lorentz group is commonly found in textbooks.
-Now, we turn to a significant challenge associated with this parametrization of the Lorentz group.
-A comparison with the Galilean group may provide insight.
-It is well known that the Galilean transformation $G(\vec{v},\vec{\theta})$, which preserves the group composition law, is isomorphic to the semi-direct product group $R^{3}\times O(3)$, where $R^{3}$ represents the group of Galilean boosts and $O(3)$ denotes the rotation group.
-In contrast, the composition of Lorentz transformations is more intricate.
-The key distinction between the Galilean and Lorentz groups lies in their structure: the Lorentz group is simple, meaning it does not naturally decompose into a direct product of rotations and boosts.
-Unlike in the Galilean case, $O(3)$ is not a normal subgroup of the Lorentz group $O(1,3)$, and the set of Lorentz boosts does not even form a group.
-The theory of the Galilean transformation composition law is straightforward and well understood.
-In contrast, the analogous theory of the Lorentz transformation composition law is unsatisfactory.
-We conclude that the standard analysis of the relationship between inertial frames overlooks the fundamental difference between Newtonian space and time and the pseudo-Euclidean geometry of space-time.
-In classical mechanics, the components of the velocity three-vector transform under a Galilean transformation in the same way as the components of the position three-vector $\vec{x}$.
-This is because, according to classical kinematics, all inertial frames share the same ordinary three-dimensional space.
-In contrast, in special relativity, different inertial frames have distinct three-dimensional spaces within the framework of space-time.
-Therefore, unexpected results may arise in special relativity when transitioning from the four-vector formalism back to the three-dimensional velocity vector $\vec{v}$, which is expressed in terms of the four-vector components $(x_{0},\vec{x})$ as $\vec{v}=d\vec{x}/dx_{0}$.
+This condition defines the most general $4\times 4$ Lorentz transformation matrix $\Lambda$.
+The set of all such matrices forms the symmetry group of the Minkowski metric under matrix multiplication, denoted by $O(1,3)$.
+In particular, $O(1,3)$ is a Lie group known as the Lorentz group.
+It is widely accepted that the relationship between inertial frames is fully determined by their relative velocities and orientations.
+Consequently, a general Lorentz transformation, denoted $\Lambda(\vec{v},\vec{\theta})$, is parameterized by a velocity vector $\vec{v}$ and a rotation vector $\vec{\theta}$, representing a combined boost and spatial rotation.
+Here, $\vec{v}$ is a vector in three-dimensional Euclidean space, while $\vec{\theta}$ belongs to the Lie algebra of the rotation group $O(3)$, which consists of all possible spatial rotations in three dimensions.
+This parameterization of the Lorentz group is standard in the literature.
+Now, we turn to a significant conceptual and structural challenge associated with this parametrization of the Lorentz group. A comparison with the Galilean group provides valuable insight.
+It is well known that the Galilean transformation $G(\vec{v},\vec{\theta})$, which satisfies the group composition law, is isomorphic to the semi-direct product group $\mathbb{R}^{3}\rtimes O(3)$, where $\mathbb{R}^{3}$ represents the group of Galilean boosts and $O(3)$ denotes the rotation group.
+In this structure, $O(3)$ acts as a normal subgroup, and the composition of transformations remains straightforward.
+In contrast, the composition of Lorentz transformations is significantly more intricate.
+The essential distinction between the Galilean and Lorentz groups lies in their group-theoretic structure: the Lorentz group is a **simple** Lie group, meaning it does not admit a nontrivial decomposition into a direct or semi-direct product of boosts and rotations.
+In particular, $O(3)$ is not a normal subgroup of the Lorentz group $O(1,3)$.
+As a result, the combined action of Lorentz boosts and spatial rotations does not obey the same algebraic structure as in the Galilean case.
+This leads to an important conclusion: standard textbook analyses of inertial frames often overlook the fundamental geometric difference between Newtonian space and time, and the pseudo-Euclidean geometry of Minkowski spacetime.
+In classical mechanics, the components of the velocity three-vector transform under Galilean transformations in the same way as the components of the position vector $\vec{x}$.
+This correspondence arises because, in Newtonian physics, all inertial observers share a common, absolute three-dimensional space.
+Time is also absolute, so the spatial position and velocity vectors can be consistently defined across all inertial frames.
+However, in special relativity, this intuition fails.
+When transitioning from the four-dimensional spacetime formalism back to the three-dimensional velocity vector $\vec{v}$, defined as $\vec{v}=d\vec{x}/dx^{0}$, subtle but profound issues emerge.
+The Lorentz group $O(1,3)$ acts on four-dimensional spacetime, and all of its subgroups also act on four-vectors, not directly on spatial vectors.
+Therefore, there is no concept of a universal or common instantaneous three-dimensional spatial frame among all inertial observers within the framework of the Lorentz group.
+This absence of a shared simultaneity structure underlies many of the non-intuitive features of relativistic kinematics, including the non-associativity and non-commutativity of velocity addition.
+Every Lorentz transformation that relates inertial observers corresponds to an isometry of the Minkowski metric.
+However, when two reference frames are related by a general Lorentz transformation—i.e., an element of the Lorentz group O(1,3)O(1,3)—the concept of relative velocity between them is not immediately well-defined.
+To extract the relative velocity, one typically attempts to factor the Lorentz transformation into a pure boost (parameterized by a velocity vector) and a spatial rotation.
+However, such a factorization is not unique: there are many ways to choose the rotation subgroup $O(3)$ within $O(1,3)$, and different choices can lead to different definitions of the relative velocity $\vec{v}$.
+Mathematically, the space of all boosts (i.e., all Lorentz transformations modulo spatial rotations) is the coset space $O(1,3)/O(3)$.
+Since $O(3)$ is not a normal subgroup of $O(1,3)$, different embeddings of $O(3)$ yield different decompositions and thus different velocities.
+The commonly used decomposition, often referred to in the literature as the ”standard choice,” implicitly assumes a privileged reference frame—a specific frame in which the rotation subgroup $O(3)$ is defined.
+In the absence of such a preferred frame, the Lorentz transformation between two systems remains well-defined as an isometry of spacetime, but the notion of a unique relative velocity becomes ambiguous.
+To define a single, well-determined velocity, one must fix a specific reference frame as distinguished.
+Physically, this can be modeled by introducing a preferred timelike vector field on spacetime, which selects a particular congruence of worldlines—effectively, a material or ”ether-like” reference frame.
+However, this dependence on a distinguished timelike direction violates the relativity principle, which asserts the equivalence of all inertial frames.
+Thus, within the geometry of pseudo-Euclidean spacetime, the equivalence of all reference systems cannot be maintained.
+17.2 Observer-Independence and Lorentz Covariance
+A standard approach to representing Lorentz transformations is via transformation matrices acting on spacetime four-vectors.
+Any Lorentz transformation can, in principle, be decomposed into a product of a rotation and a boost.
+However, the explicit form of this decomposition depends on the choice of reference frame, which in the matrix formalism is determined by the selection of a distinguished timelike vector, typically denoted by $e_{0}$, in a chosen basis.
+For example, in his widely used textbook, Barut defines a Lorentz boost (see Eq. (1.28) and the accompanying discussion in [66]) as:
+$L_{\vec{v}}=(L_{-\vec{v}})^{-1}=\left(\begin{array}[]{cccc}\gamma&-\beta\gamma&0&0\\ -\beta\gamma&\gamma&0&0\\ 0&0&1&0\\ 0&0&0&1\end{array}\right).$
+where $\beta=v/c$ and $\gamma=1/\sqrt{1-\beta^{2}}$. This boost corresponds to motion along the $x$-axis and is defined relative to a specific basis.
+However, the Lorentz boost is not fully determined by the velocity parameter $\vec{v}$ alone.
+The very notion of velocity presupposes a distinction between space and time—i.e., a $1+3$ decomposition of spacetime—which requires the selection of a specific timelike direction.
+This decomposition identifies the observer through the choice of a basis vector $e_{0}$, which represents the time direction in that observer’s rest frame.
+As made explicit in Barut and other standard references [66], the Lorentz boost is expressed as a $4\times 4$ matrix relative to a given orthonormal basis.
+In the “standard” formulation, the preferred inertial frame is associated with the basis vector $e_{0}=(1,0,0,0)$.
+Rotations then occur within the three-dimensional spatial subspace orthogonal to $e_{0}$, and any boost involving a velocity vector $\vec{v}$ orthogonal to $e_{0}$ implicitly defines the boost with respect to this preferred frame.
+Thus, even though Lorentz transformations are covariant under changes of inertial observers, their explicit matrix representations—and the definition of velocity itself—rely on a choice of observer.
+Decompositions of spacetime based on different choices of the timelike basis vector $e_{0}$ are mutually incomparable, as they correspond to distinct three-dimensional spatial slices for which no canonical identification exists.
+Successive Lorentz boosts are composed through matrix multiplication, assuming both boosts are expressed in the same basis.
+Specifically, if $L_{\vec{v}_{1}}$ is the boost from frame $K_{0}$ to frame $K_{1}$, and $L_{\vec{v}_{2}}$ is the boost from $K_{1}$ to $K_{2}$, then their composition is given by: $L_{\vec{v}_{2}}L_{\vec{v}_{1}}$.
+Here, $\vec{v}_{1}$ is the velocity of frame $K_{1}$ relative to the preferred frame $K_{0}$, and $\vec{v}_{2}$ is the velocity of frame $K_{2}$ relative to $K_{1}$, as observed from $K_{0}$.
+It is crucial to distinguish between observer-independence and Lorentz covariance.
+Although Lorentz covariance refers to the invariance of physical laws under Lorentz transformations, it does not imply that all inertial frames are physically equivalent in a pseudo-Euclidean spacetime.
+The geometry of Minkowski space allows for Lorentz covariance without guaranteeing the equivalence of all inertial observers.
+This subtle distinction opens the door to the possibility of a physically preferred inertial frame, thereby challenging the traditional interpretation of the principle of relativity.
+In this context, relative velocity alone does not fully characterize the relationship between inertial frames.
+[^55] In particular, a non-zero relative velocity does not, by itself, cause time dilation.
+See Section 12.3 for further discussion.
+A deeper tension emerges between the principle of relativity—which asserts the equivalence of all inertial frames—and the geometric structure of Minkowski spacetime.
+In fact, only one of these principles can hold unconditionally.
+[^56] For a more detailed analysis, see “How Do You Add Relative Velocities?” by Oziewicz in [67].
+Oziewicz argues that within pseudo-Euclidean spacetime, the equivalence of all reference frames is not possible.
+Given the strength of experimental evidence supporting Lorentz invariance, we adopt Lorentz covariance as the foundational principle of the theory.
+As discussed in earlier chapters, further mathematical considerations support the view that pseudo-Euclidean geometry does not entail the complete equivalence of all inertial frames.
+One key insight comes from analyzing the continuity of the metric tensor under coordinate transformations.
+The Langevin metric, for instance, arises from a continuous deformation of the Minkowski metric when transitioning between an inertial frame and an accelerated frame.
+This smooth transition can be modeled via a Galilean boost relating the coordinate systems of the inertial and accelerated observers.
+However, restoring the diagonal Minkowski form in the accelerated frame often requires an abrupt coordinate transformation, resulting in a discontinuity in the metric tensor.
+This discontinuity signals a breakdown in the symmetry between inertial frames: if a Lorentz transformation fails to preserve the smoothness of the metric, then the equivalence of the frames involved is compromised.
+This raises a natural and important question: How can observers identify a preferred inertial frame?
+Within the structure of the Lorentz group, relative velocity between frames enters exclusively through the boost transformations.
+However, a Lorentz boost is always defined relative to a specific reference frame—typically one associated with a distinguished timelike vector.
+Special relativity, in its standard formulation, does not specify such a frame; yet physical observations suggest the existence of a unique inertial frame—one not subjected to prior accelerations, such as relative to the fixed stars.
+There is, however, no intrinsic contradiction between the mathematical framework of special relativity and the physical asymmetry introduced by selecting a preferred frame.
+The theory is built on the requirement that physical laws remain invariant under Lorentz transformations.
+But special relativity is not a self-contained theory in the sense of fully determining physical phenomena from dynamical laws alone—initial conditions must also be specified.
+The principle of relativity asserts that the form of physical laws is independent of velocity, but not necessarily that initial conditions must be velocity-independent.
+While Lorentz coordinatization appears to establish symmetry among inertial frames, the asymmetry emerges from the history of acceleration.
+Once a frame has undergone acceleration, information about that acceleration is not encoded in the transformation laws themselves but in the initial conditions from which they operate.
+Consider, for instance, the relativistic aberration of particles emitted from a source at rest in an accelerated frame.
+As discussed in Chapter 9, special relativity predicts that—alongside a static electric field—a magnetic field also arises in the accelerated frame, leading to deviations in the observed trajectory of particles.
+This deviation is a measurable consequence of the frame’s prior acceleration.
+The mathematical framework of special relativity remains entirely self-consistent when formulated relative to an initial inertial frame—one that has not undergone prior acceleration with respect to the distant stars.
+Such a frame functions as a *privileged reference system*.
+As demonstrated in Chapter 7, when the laws of physics are formulated relative to this frame, they accurately account for all observed phenomena, including those reported by accelerated observers.
+In particular, phenomena like the aberration of light in an accelerated frame can be fully derived using standard tools of special relativity—specifically, Lorentz transformations and Wigner rotations—when analyzed from the standpoint of the initial inertial frame.
+17.3 Einstein Velocity Addition and Conventional Particle Tracking
+Composition of relativistic velocities, also named as ”Einstein’s addition of velocities” is somewhat confusing due to fact that this addition is nonassociative.
+This is contrast with associativity of the Lorentz group.
 Most texts on special relativity present the relativistic velocity addition formula only for parallel velocities.
-In this simplified case, Einstein’s velocity addition is both commutative and associative, making it a commutative group operation.
+In this simplified case, Einstein’s velocity addition $(\oplus)$ is both commutative and associative, making it a commutative group operation analogous to the Galilean velocities addition $(+)$.
 However, in the general case, Einstein’s velocity addition $\oplus$ does not form a group.
-The relativistic reciprocity principle states that if an inertial frame $K^{\prime}$ moves with velocity $\vec{v}$ relative to another inertial frame $K$, then, reciprocally, the velocity of $K$ relative to $K^{\prime}$ is $-\vec{v}$.
+The relativistic reciprocity principle states that if an inertial frame $K^{\prime}$ moves with velocity $\vec{v}$ relative to another inertial frame $K$, then, reciprocally, the velocity of $K$ relative to $K^{\prime}$ is $\vec{v}$.
 This principle suggests that the inverse of a velocity $\vec{v}$ is simply its negation: $\vec{v}^{-1}=-\vec{v}$, analogous to the Galilean addition of velocities.
+This is often presented in textbooks as a self-evident fact requiring no further discussion.
 For composite velocities, $\vec{u}\oplus\vec{v}$, reciprocity implies $(\vec{u}\oplus\vec{v})^{-1}=-(\vec{u}\oplus\vec{v})=(-\vec{v})\oplus(-\vec{u})=-(\vec{v}\oplus\vec{u})$.
 However, since Einstein’s velocity addition $\oplus$ is generally non-commutative, this leads to a contradiction, known as the reciprocity paradox.
-This paradox is closely related to the Mocanu paradox, as discussed by Mocanu [66] and was latter resolved by Ungar [67] who demonstrated the crucial role of the Wigner (Thomas) rotation in the transformation.
-Furthermore, Ungar [67] also showed that Einstein’s velocity addition is also non-associative meaning that $(\vec{w}\oplus(\vec{v}\oplus\vec{u}))\neq((\vec{w}\oplus\vec{v})\oplus\vec{u})$.
+This paradox is closely related to the Mocanu paradox, as discussed by Mocanu [68] and was latter resolved by Ungar [69] who demonstrated the crucial role of the Wigner rotation in the transformation.
+[^57] It is sometimes referred to as the Thomas rotation Furthermore, Ungar [69] also discovered that Einstein’s velocity addition is non-associative meaning that $(\vec{w}\oplus(\vec{v}\oplus\vec{u}))\neq((\vec{w}\oplus\vec{v})\oplus\vec{u})$.
 Being non-associative, the Einstein’s addition $\oplus$ is not a group operation.
 Non-associativity is paradoxical: for a system of four bodies the $\oplus$ of three non-parallel velocities gives two distinct velocities between two bodies.
-Most textbooks present the Lorentz transformation as being generated by a boost vector $\vec{v}$, interpreted as the unique reciprocal velocity between two inertial frames.
-However, we emphasize that the reciprocal property $\vec{v}^{-1}=-\vec{v}$ is inconsistent with Einstein’s two postulates of special relativity  [^55] For a more in-depth discussion on the relationship between the reciprocity axiom and Einstein’s two postulates of special relativity, we recommend the chapter ”How Do You Add Relative Velocities?” by Oziewicz in [69]..
-The Lorentz isometry group $\Lambda$ acts on four-dimensional spacetime, and all subgroups of the Lorentz group operate on four-vectors in spacetime.
-Consequently, there is no concept of common (instantaneous) spatial frame within the framework of the Lorentz group.
-The Lorentz transformation $\Lambda$ does not naturally decompose into a unique combination of a rotation and a boost.
-Any such decomposition would imply the existence of a preferred frame, contradicting the principle of relativity.
-In Minkowski space, where spacetime follows pseudo-Euclidean geometry, equivalence of all inertial frames is not possible.
-This suggests the existence of an explicit, preferred (initial) inertial reference frame within special relativity, challenging the conventional interpretation.
-The relationship between inertial frames are not determined solely by their relative velocities 565656 In particular, the presence of non-zero relative velocity cannot, by itself, be a cause of time dilation (see Section 12.3 for more detail)..
-The inverse of a velocity vector, $\vec{v}^{-1}=-\vec{v}$, is often presented in textbooks as a self-evident fact requiring no further discussion.
-However, this assumption contradicts a fundamental axiom of relativity theory—namely, the axiom of pseudo-Euclidean spacetime geometry.
-For example, in his contemporary textbook, Barut defines Lorentz boost as (see Eq. (1.28) and the accompanying discussion in [6]):
-$L_{\vec{v}}=(L_{-\vec{v}})^{-1}=\left(\begin{array}[]{cccc}\gamma&-\beta\gamma&0&0\\ -\beta\gamma&\gamma&0&0\\ 0&0&1&0\\ 0&0&0&1\end{array}\right).$
-It is evident that all textbooks, including [6], define the Lorentz boost as a 4x4 basis-dependent matrix. Next we shall discuss some physical applications of the transformation $L_{\vec{v}}$.
+Let us illustrate non-commutativity of Einstein’s velocity addition using only simple case of relativistically combining perpendicular velocities.
 Consider a coordinate system $K_{1}$ moving with respect to $K_{0}$ with a velocity $\vec{v}_{01}$ and a frame $K_{2}$ moving with respect to $K_{1}$ with a velocity $\vec{v}_{12}$.
 It is important to note that $\vec{v}_{01}$ represents the velocity of frame $K_{1}$ as measured by an observer at rest in $K_{0}$, while $\vec{v}_{12}$ represents the velocity of $K_{2}$ as measured by an observer in $K_{1}$.
-The question then arises: what is the Lorentz transformation from $K_{0}$ to $K_{2}$?
-Let the corresponding Lorentz transformations be denoted by $L_{1}$, $L_{2}$ and $L$, respectively.
-Then, according to standard textbooks conventions, we have $L=L_{2}L_{1}$.
+The question then arises: what is the velocity transformation from $K_{0}$ to $K_{2}$?
 Now, consider the case, where $\vec{v}_{01}$ and $\vec{v}_{12}$ are perpendicular.
+The ”perpendicular” means that velocity $\vec{v}_{12}$ is perpendicular to line motion of frame $K_{0}$ in $K_{1}$.
 The velocity $\vec{v}_{02}$ of frame $K_{2}$ as observed from $K_{0}$ is given by $\vec{v}_{02}=\vec{v}_{01}+\vec{v}_{12}\sqrt{1-v_{01}^{2}/c^{2}}$.
-According to reciprocity principle, accepted in all textbooks, an observer in $K_{2}$ sees frame $K_{1}$ moving with velocity $-\vec{v}_{12}$, while an observer in $K_{1}$ sees frame $K_{0}$ moving with velocity $-\vec{v}_{01}$.
+According to reciprocity principle, accepted in all textbooks, an observer in $K_{2}$ sees frame $K_{1}$ moving with velocity $\vec{v}_{12}$, while an observer in $K_{1}$ sees frame $K_{0}$ moving with velocity $\vec{v}_{01}$.
 Similarly, applying Einstein’s velocity addition, we find that the velocity $\vec{v}_{20}$ of frame $K_{0}$ as observed from $K_{2}$ is: $\vec{v}_{20}=-\vec{v}_{12}-\vec{v}_{01}\sqrt{1-v_{12}^{2}/c^{2}}$.
 According to textbooks, $\vec{v}_{20}$ should be reciprocal of $\vec{v}_{02}$, but they point in different directions.
-The angle between $\vec{v}_{02}$ and $\vec{v}_{20}$ is the Wigner rotation angle.
 This reciprocity discrepancy leads to the Mocanu contradiction and highlights the non-associativity of Einstein’s velocity addition.
-While composite velocities in Galilean relativity fully determine the relationships between inertial frames, their relativistic counterparts do not.
-In addition to the relativistic velocity composition between inertial frames, one must also account for the associated Wigner (Thomas) rotations.
-In his papers and monographs—written at an advanced mathematical level — Ungar attempted to resolve the non-associativity paradox of Einstein’s velocity addition $\oplus$ by refining the composition law (see [68] and references therein).
-The improved composition law replaces $\vec{u}\oplus\vec{v}$ with $[\vec{u}\oplus\vec{v},\mathrm{W}(\vec{u},\vec{v})]$, where $\mathrm{W}(\vec{u},\vec{v})$ represents the Wigner (Thomas) rotation generated by $\vec{u}$ and $\vec{v}$.
-Notably, the binary operation $\oplus$ remains the standard Einstein velocity addition.
-However, this approach is unsatisfactory even from a mathematical standpoint for a fundamental reason.
-As demonstrated in Section 16.6, the Wigner rotation is not merely an additive correction but is intrinsically woven into the structure of relativistic velocity addition itself.
-Consider the asymptotic case of two perpendicular velocities: $\vec{v}_{01}$ approaching the speed of light and $\vec{v}_{12}$ being small.
-In this limit, $\gamma=(1-v_{01}^{2}/c^{2})^{-1/2}\gg 1\gg v_{12}/c$.
-The frame $K_{1}$ with the moving frame $K_{2}$ seems to rotate about frame $K_{0}$.
-We want now to look at the consequence of this Wigner rotation.
-In the Section 16.6 we found that in Lorentz coordinatization, the velocity $v_{01}$ of frame $K_{2}$ in frame $K_{0}$, after a boost with velocity $\vec{v}_{12}$, undergoes a simple rotation by an angle $v_{12}/(\gamma v_{01})=v_{12}/(\gamma c)$ without any change in its magnitude.
-In contrast, Ungar’s approach—based on the standard Einstein velocity addition within an improved composition law—suggests that the total speed of $K_{2}$ in the Lorentz frame $K_{0}$ increases from $v_{01}$ to $v_{01}+v_{12}^{2}/(2\gamma^{2}c)$.
-In all special relativity textbooks, the principle of reciprocity states that the inverse of velocity addition must be identical to that in Galilean absolute time.
-Consequently, Einstein’s velocity addition is based on the relation $dt^{\prime}=dt/\gamma$, which implicitly assumes an absolute time convention.
-However, it does not account for the interplay between transverse positions and time, which is inherent in Lorentz boost matrices.
-In reality, every Lorentz boost represents a transformation between reference frames within a Lorentzian framework (i.e., under Lorentz coordinatization).
-Thus, we conclude that the commonly accepted derivation of velocity composition does not follow from the composition of Lorentz boosts, $L=L_{2}L_{1}$.
 Under Lorentz coordinatization, moving frames inherently possess different simultaneity relations.
@@
 The reciprocity of frames contradicts relativity of simultaneity.
+The commonly accepted derivation of velocity composition does not follow from the composition of Lorentz boosts and based on absolute time convention.
 In Enstein’s velocity addition all velocities $\vec{v}_{01}$, $\vec{v}_{12}$, $\vec{v}_{02}$, are implicitly assumed to lie within the same preferred 3-space as seen by a preferred observer.
-In other words, the reciprocity of relative velocities holds only under an absolute time framework.
-The standard textbook presentation of the relativistic transformations implicitly assumes that observers in $K_{0}$, $K_{1}$, and $K_{2}$ inertial frames share a common 3-space.
 It is often assumed that the acceleration of a rigid reference frame grid has a direct physical meaning.
 However, this is a misconception.
-Different inertial frames have distinct ordinary spaces and implicit assumption of velocity reciprocity contradicts Einstein’s two postulates.
-The concepts of observer independence and Lorentz covariance are fundamentally distinct.
-We conclude that there is a striking conflict between the Principle of Relativity (i.e., the equivalence of all inertial frames) and the axiom of pseudo-Euclidean spacetime geometry.
-In other words, only one of these principles can hold.
-Based on experimental evidence, we adopt the Lorentz covariance principle  [^57] However, an alternative perspective exists in the literature.
-Oziewicz, for instance, favors the Principle of Relativity [69]..
-This is a good point to make some general remarks about Einstein’s velocity addition.
-It is important to recognize that it does not directly follow from the Lorentz transformations.
-While Lorentz boosts intertwine space and time coordinates, Einstein’s velocity addition relies on the assumption of a shared three-dimensional space across all inertial frames.
-Despite this, Einstein’s velocity addition is widely used in relativistic dynamics and electrodynamics, raising a reasonable question: How can this be?
-After all, we know that this form of velocity addition does not constitute a mathematical group—it is non-associative, and without associativity, the development of a consistent and complete theory would seem impossible.
-Earlier, we demonstrated that the main issue lies in the definition of relative velocity.
-If we postulate, as Einstein did, the principle of relativistic reciprocity, we inevitably encounter the non-associativity of velocity addition.
-Nevertheless, one should recognize the existence of a preferred (initial) inertial frame without an accelerational history relative to the fixed stars.
-Based on experimental evidence, we adopt the Sun-centered frame as the preferred inertial frame (see Chapter 15 for further discussion).
-Once a preferred frame is established, there is one—and only one—well-defined relative velocity between any pair of inertial frames.
-Let us apply this approach to an example and observe how it performs in practice.
-Suppose an observer in the initial inertial frame $K_{0}$ establishes a coordinate system using Einstein’s synchronization procedure, based on light signals emitted by a stationary dipole source, under the assumption that light propagates isotropically at speed $c$.
-Now consider two additional moving inertial frames, $K_{1}$ and $K_{2}$, whose motions are analyzed by the observer in $K_{0}$.
-The simplest synchronization method involves maintaining the same set of synchronized clocks, thus preserving simultaneity and adhering to an absolute-time coordinatization in $K_{0}$.
-We assume that system $K_{1}$ moves with velocity $v_{x}$ along the $x$-axis relative to $K_{0}$, and $K_{2}$ moves with velocity $v_{y}$ along the $y$-axis relative to $K_{1}$.
-It is important to note that $v_{x}$ and $v_{y}$ represent the velocities as measured by the observer in $K_{0}$.
-Applying Einstein’s velocity addition, the observer in $K_{0}$ concludes that an observer in $K_{1}$ sees frame $K_{2}$ moving with a velocity $v_{y}/\sqrt{1-v_{x}^{2}/c^{2}}$.
-According to Einstein, the velocity of frame $K_{2}$ as observed directly from $K_{0}$ is given by $\vec{v}_{02}=v_{x}\vec{e}_{x}+v_{y}\vec{e}_{y}$ consistent with Galilean velocity addition within the initial inertial frame.
-Now, consider a different perspective.
-The observer in $K_{0}$ sees the relative velocity of $K_{2}$ with respect to $K_{1}$ as $-v_{y}$.
-Similarly, an observer in $K_{2}$ sees $K_{1}$ moving with velocity $-v_{y}/\sqrt{1-v_{02}^{2}/c^{2}}$.
-According to Einstein’s velocity addition, the velocity of $K_{1}$ as observed from $K_{0}$ is $\vec{v}_{01}=v_{x}\vec{e}_{x}$ as expected.
-We thus conclude that Einstein’s velocity addition is consistent with Galilean velocity addition when viewed from the initial inertial frame.
-In this scheme, reciprocity is absent, and all paradoxes are eliminated.
-Indeed, there is no reciprocity in the relative velocities between $K_{1}$ and $K_{2}$, because time dilation is absolute, not relative, in the presence of a preferred inertial frame.
-Now, let us consider the Earth-based frame.
-For an observer on Earth, the velocity relative to the initial inertial frame is approximately $v_{orb}\simeq 30~{}\mathrm{km/s}$, allowing us to neglect terms of order $v_{orb}^{2}/c^{2}$.
-As previously demonstrated, current experimental techniques usually lack the sensitivity required to detect the influence of orbital velocity on the relativistic particle dynamics.
-Consequently, in accelerator physics, the Earth-based frame can be treated with good accuracy as the initial inertial frame.
-It is important to emphasize that Einstein’s velocity addition is fully compatible with conventional particle tracking within a single inertial frame, which itself relies on the absolute time convention.
+Despite this, Einstein’s velocity addition is widely used in relativistic dynamics raising a reasonable question: How can this be?
+It is important to emphasize that Einstein’s velocity addition is fully compatible with conventional particle tracking within the initial inertial frame, which itself relies on the absolute time convention.
 Crucially, Einstein’s velocity addition inherently incorporates relativistic time dilation.
 Similarly, conventional particle tracking also accounts for relativistic time dilation, albeit implicitly, through the use of relativistic mass for moving particles.
-This approach presents no fundamental challenges in mechanics.
+Non-covariant approach presents no fundamental challenges in mechanics.
 The preference for a non-covariant formulation within the framework of dynamics arises from its simplicity and practicality.
@@
 Only by adopting a kinematic framework based on Lorentz coordinatization can we ensure the correct coupling between the isotropic Maxwell equations and particle trajectories in the laboratory frame.
-Let us now come to our conclusion: Einstein’s velocity addition requires the existence of a preferred inertial frame.
-Therefore, Einstein’s velocity addition is consistent with the concept of an initial inertial frame and stands in contradiction to the Principle of Relativity, namely the equivalence of all inertial frames.
+17.4 Velocity Addition and the Wigner Rotation
+While the coordinate transformation formula for Lorentz boosts and the velocity addition formula may appear formally similar, they exhibit significant differences in structure and properties.
+Specifically, the composition of Lorentz boosts is associative, as it corresponds to the matrix product of Lorentz transformation matrices.
+However, this composition does not form a group under velocity addition alone.
+In contrast, Einstein velocity addition is not associative.
+The consequence of this difference is the Wigner rotation accompanying the composition of non-parallel velocities.
+For completeness, it is worth noting earlier attempts to address the limitations of Einstein velocity composition.
+In a series of mathematically rigorous works, Ungar proposed a refined framework to resolve the apparent paradox of non-associativity in relativistic velocity addition (see [70, 71] and references therein).
+In this refined framework, the composition of velocities is not solely represented by $\vec{u}\oplus\vec{v}$, but rather by the pair $\vec{u}\oplus\vec{v},\mathrm{R}(\vec{u},\vec{v})]$, where $\mathrm{R}(\vec{u},\vec{v})$ denotes the Wigner rotation arising from the non-commutative nature of Lorentz boosts.
+Notably, the binary operation $\oplus$ remains the standard Einstein velocity addition; what changes is the recognition that a rotation must accompany it to fully capture the transformation.
+It should be noted that the Wigner rotation leaves invariant preferred observer.
+A common textbook derivation of the Wigner rotation begins by decomposing the product of two Lorentz boosts, $L(\vec{v})$ and $L(\vec{u})$, as follows: $L(\vec{v})L(\vec{u})=L(\vec{w})R(\vec{u},\vec{v})$, where $R(\vec{u},\vec{v})$ is identified as the Wigner rotation, and the resultant boost $L(\vec{w})$ corresponds to the Einstein velocity addition $\vec{w}=\vec{u}\oplus\vec{v}$.
+However, this standard approach is mathematically incomplete, as it treats the rotation $R$ as a secondary effect rather than an intrinsic component of the velocity composition itself.
+As shown in Section 16.6, the Wigner rotation is not a mere additive correction—it is fundamentally embedded in the structure of relativistic velocity addition.
+Consider, for example, the composition of two perpendicular velocities $\vec{v}_{01}$ and $\vec{v}_{12}$ as discussed earlier.
+Suppose $\vec{v}_{01}$ approaches the speed of light, while $\vec{v}_{12}$ remains small.
+In this regime, the Lorentz factor $\gamma_{01}=(1-v_{01}^{2}/c^{2})^{-1/2}$ becomes very large, i.e., $\gamma_{01}\gg 1\gg v_{12}/c$.
+Treating $v_{12}/c$ as a small parameter and retaining terms up to second order, we observe that the intermediate frame $K_{1}$ appears to rotate with respect to $K_{0}$.
+This effect is a manifestation of the Wigner rotation.
+Ungar’s approach—based on the Einstein velocity addition within an improved composition law—suggests that the total speed of $K_{2}$ in the Lorentz frame $K_{0}$ increases from $v_{01}$ to $v_{01}+v_{12}^{2}/(2\gamma^{2}c)$.
+However, this standard velocities addition overlooks a critical point.
+In the Section 16.6 we found that in Lorentz coordinatization, the total velocity vector of frame $K_{2}$ in $K_{0}$, after a boost with velocity $\vec{v}_{12}$ in $K_{1}$, undergoes a simple rotation by an angle $v_{12}/(\gamma c)$ without any change in its magnitude, $v_{02}=v_{01}$.
+This is intuitively reasonable, since we are using Lorentz coordinatization and in our relativistic asymptotic $v_{01}$ approaching the speed of light.
+Let us now re-examine, from a covariant perspective, why Einstein’s velocity addition formula proves inadequate—even when only two velocities are involved.
+This topic was already introduced in the previous chapter, where we provided a general overview.
+That earlier discussion covers what is typically sufficient for most purposes. However, we will now revisit the topic in greater detail to develop a more complete understanding.
+The boost from frame $K_{1}$ to $K_{2}$, as observed from $K_{0}$, can be understood as a sequence of infinitesimal Lorentz transformations, each involving a small velocity increment $\Delta v_{12}/\gamma_{01}$.
+To provide a complete analysis, we must also account for the Wigner rotation which occurs in the same direction as the rotation of the velocity vector of $K_{2}$ as seen from $K_{0}$.
+The coordinate axes of the frame $K_{2}$ remain parallel to those of frame $K_{1}$.
+From the perspective of frame $K_{0}$, the coordinate axes of frame $K_{1}$ are also rotated by the same angle with respect to those of frame $K_{0}$.
+The infinitesimal velocity increment of $K_{2}$ as observed from $K_{0}$ can be decomposed into two physically meaningful components: 1) a transverse component of $\Delta v_{\perp}\simeq\Delta v_{12}/\gamma_{01}$, perpendicular to the original velocity $\vec{v}_{01}$, and 2) a tangential component of $\theta_{w}\Delta v_{12}/\gamma_{01}\simeq\theta_{w}\Delta v_{\perp}$ which is aligned with the direction of $\vec{v}_{01}$, and arises due to the Wigner rotation.
+Here $\theta_{w}$ denotes the Wigner rotation angle.
+In the ultra-relativistic limit we approximate it as $\theta_{w}\simeq v_{\perp}/c$.
+Notably, analyzing the transverse component is considerably more straightforward than handling the tangential one.
+In fact, to second order in $v_{12}/c$, no relativistic corrections affect the transverse velocity component.
+We now proceed to compute the net effect of accumulating these infinitesimal increments.
+This process is equivalent to integrating over the transverse velocity: $\int\theta_{w}(v_{\perp})dv_{\perp}$.
+Evaluating this from $v_{\perp}=0$ to $v_{\perp}=v_{12}/\gamma_{01}$ gives the total tangential correction due to the Wigner rotation.
+This results in a tangential velocity component of $v_{01}-v_{12}^{2}/(2\gamma_{01}^{2}c)$.
+Thus, to second order in $v_{12}/c$, the total speed of frame $K_{2}$ with respect to $K_{0}$ remains $v_{02}=v_{01}$.
+This analysis is carried out in the limit of a small Wigner rotation angle, $\theta_{w}=v_{12}/(\gamma_{01}c)\ll 1/\gamma_{01}$.
+Even within this simplified regime, we can clearly demonstrate the fundamental difference between covariant velocity addition and Einstein’s velocity addition.
+To properly interpret the physical meaning of covariant velocity addition as presented here, it is essential to recognize that Einstein’s velocity addition is fully compatible with conventional particle tracking methods, which are based on the absolute time convention.
+However, the only rigorous way to maintain consistency between Maxwell’s equations and particle trajectories is to formulate and solve the equations of motion in a fully covariant manner.
+Experimental results support our correction to the conventional theory of relativistic velocity addition; further details are provided in the final two chapters.
+17.5 Lorentz-Covariant Particle Tracking
+In covariant particle tracking, a particle’s trajectory is described from the perspective of an initial inertial frame as a result of successive Lorentz boosts.
+Let us now apply this approach to a concrete example and assess its practical implications.
+Based on experimental evidence, we adopt the Sun-centered frame as the preferred inertial frame (see Chapter 15 for further discussion).
+Once such a preferred frame is established, a unique and well-defined relative velocity exists between any two inertial frames.
+Now, let us consider the Earth-based frame.
+For an observer on Earth, the velocity relative to the initial inertial frame is approximately $v_{orb}\simeq 30~\mathrm{km/s}$.
+As previously demonstrated, current experimental techniques usually lack the sensitivity required to detect the influence of orbital velocity on the relativistic particle dynamics.
+Consequently, in accelerator physics, the Earth-based frame can be treated with good accuracy as the preferred inertial frame.
+Suppose an observer in the initial inertial frame $K_{0}$ establishes a coordinate system using Einstein’s synchronization procedure, which relies on light signals emitted by a dipole source at rest and assumes that light propagates isotropically at speed $c$.
+This procedure enables the construction of a Lorentz coordinate system in $K_{0}$.
+Now, consider two additional inertial frames, $K_{1}$ and $K_{2}$, both moving relative to $K_{0}$.
+The simplest synchronization method is to retain the same set of synchronized clocks across all frames.
+This preserves simultaneity and leads to an absolute-time coordinatization based on $K_{0}$.
+To describe physical phenomena across these frames, the preferred observer uses a metric tensor to derive the (electro)dynamical equations.
+According to the equivalence between passive and active Lorentz boosts within a single inertial frame, the metric in Eq.
+11 characterizes the measurements of the preferred observer.
+Let us now examine velocity composition within the absolute-time coordinatization framework.
+Assume that frame $K_{1}$ moves with velocity $\vec{v}_{01}$ relative to $K_{0}$, and frame $K_{2}$ moves with velocity $\vec{v}_{12}$ relative to $K_{1}$.
+Importantly, both $\vec{v}_{01}$ and $\vec{v}_{12}$ are defined from the perspective of the observer in $K_{0}$.
+[^58] The velocity $v_{12}$ used here denotes the relative velocity as measured by a preferred observer, and should not be confused with the $v_{12}$ in the preceding section, which refers to the relative velocity measured in frame $K_{1}$ when observing this observer from frame $K_{0}$.
+In this setup, velocity transformation is particularly straightforward: under absolute-time coordinatization, Galilean velocity addition applies.
+Thus, the velocity of $K_{2}$ as observed from $K_{0}$ is simply $\vec{v}_{02}=\vec{v}_{12}+\vec{v}_{01}$.
+This observation highlights a key advantage: when the preferred observer adopts the absolute-time coordinatization, relativistic kinematics becomes extremely simple.
+In this formulation, all the complexity of relativistic particle tracking is absorbed into the frame metrics.
+Consider now two electrons, both moving with the same velocity $\vec{v}_{02}$ relative to the preferred frame.
+The first electron is accelerated directly from rest to $\vec{v}_{02}$ within the preferred frame.
+The second reaches the same final velocity through two steps: first accelerated to $\vec{v}_{01}$, then from $\vec{v}_{01}$ to $\vec{v}_{02}$.
+Despite having the same final velocity, the two electrons have different acceleration histories.
+As a result, the electrodynamics of these two electrons will differ due to differences in their respective metrics.
+The contrast between these seemingly equivalent cases is both subtle and intriguing.
+This paradoxical asymmetry between the two setups is most naturally understood from the perspective of electrodynamics.
+Our everyday intuition often breaks down when confronted with the counterintuitive nature of special relativity.
+The dynamics governed by electromagnetic fields are, in this context, encoded in the language of pseudo-Euclidean geometry, which can obscure their physical implications.
+The idea that each electron ”remembers” the acceleration it experienced relative to a preferred frame may seem surprising at first.
+However, this is entirely consistent with our understanding of relativistic electrodynamics.
+Let us consider the case where the velocity $\vec{v}_{01}$ approaches the speed of light, while the velocity $\vec{v}_{12}$ is perpendicular to $\vec{v}_{01}$ and remains small.
+More specifically, we are interested in the limit where $\gamma_{01}\gg 1\gg v_{12}/c\gg 1/\gamma_{01}$.
+In this setup, the first electron is accelerated in parallel geometry, reaching a final velocity $\vec{v}_{02}$.
+The second electron, in contrast, is finally accelerated in a direction perpendicular to the ultrarelativistic velocity $\vec{v}_{01}$.
+It is well known that transverse acceleration in the ultrarelativistic regime results in radiation that is enhanced by a factor of approximately $\gamma_{01}^{2}$ compared to the case of parallel acceleration.
+Thus, in the ultrarelativistic limit, the first electron (undergoing parallel acceleration) emits negligible radiation.
+In contrast, the second electron emits significant synchrotron radiation due to its transverse acceleration.
+From the standpoint of electrodynamics, these two configurations are fundamentally different.
+In particular, the radiation emitted by the second electron carries information about its acceleration history.
+Now let us consider a scenario where the deflection angle $v_{12}/c$ of the second electron is smaller than $1/\gamma_{01}$.
+In this regime, synchrotron radiation is effectively absent.
+This leads to a compelling question: in the absence of radiation, where—within the framework of absolute time coordinatization—is the information about the second electron’s acceleration physically located?
+According to conventional particle tracking, both electrons end up with the same final velocity after acceleration.
+However, the second electron underwent a qualitatively different acceleration process, raising subtle questions about how this information is encoded in the system when no radiation is emitted.
+One might be tempted to assert: “Since an electron is a structureless particle, this situation seems paradoxical.” However, this is not accurate.
+An ultrarelativistic electron cannot be regarded as structureless in the conventional sense.
+The electromagnetic field associated with such an electron exhibits a macroscopic transverse extent.
+The spatial distribution of this virtual radiation field is well described by the Ginzburg–Frank formula [73] (see Appendix A4 for further details).
+For a rapidly moving electron, the transverse components of the electric and magnetic fields are nearly equal in magnitude and orthogonal to one another—effectively indistinguishable from those of a real radiation beam.
+On a microscopic level, when the electron experiences transverse acceleration, its electromagnetic field is perturbed in a way that reflects this acceleration.
+[^59] Certain assumptions are necessary when applying relativistic kinematics in practice.
+If the electron undergoes a sudden acceleration, there is a minimum distance—known as the formation length—required for the development of the (virtual laser-like) self-field of an ultrarelativistic electron.
+Only when the distance traveled after the first acceleration exceeds this formation length can a second acceleration (or ’boost’) be considered kinematically independent.
+If this condition is not met, the two accelerations effectively combine into a single boost from rest to a final velocity $v_{02}$ and the electron’s acceleration history will appear nearly identical from the perspective of relativistic electrodynamics.
+Under the framework of classical (Galilean) kinematics, the orientation of the virtual radiation phase front remains unchanged.
+However, since Maxwell’s equations are not invariant under Galilean transformations—as discussed throughout this book—the adoption of such kinematics necessitates the use of anisotropic field equations.
+Consequently, although the phase front remains planar, the direction of propagation is no longer perpendicular to it.
+In other words, the (virtual) radiation beam’s direction of motion diverges from the normal to its phase front.
+Within the absolute time coordinatization, electrodynamics predicts that the virtual radiation beam of the second electron propagates along the direction of $\vec{v}_{02}$, with a phase front tilt $v_{12}/c$.
+This tilt is the crucial clue to resolving the paradox.
+The key insight is this: The information about the electron’s acceleration is not lost—it is embedded in the perturbation of the electron’s self-fields.
+Mathematically, this information is encoded in the difference between metrics of these two electrons
+This rises a reasonable question: why does there is no virtual phase front perturbation in the situation with first electron acceleration?
+The key distinction is in collinear geometry - the first electron accelerates perpendicular to the plane of the virtual radiation wavefront (i.e., the plane of simultaneity).
+In the case of the boost with velocity $v_{12}$ the second electron undergoes acceleration along the virtual radiation wavefront.
+Absolute time coordinatization in the preferred frame can be transformed into Lorentz coordinatization.
+In the Chapter 3 we consider a general method for metric diagonalization in the case when particle accelerated from rest to velocity $v$ in the preferred frame.
+The change variables according to Eq.
+13 completed by Galilean transformation is mathematically equivalent to the Lorentz transformation with the same velocity.
+As a result, covariant particle tracking in the case of the first electron is straightforward.
+Similar there is no problem with the first boost of the second electron.
+To maintain a Lorentz coordinate system in the preferred frame after the second boost with velocity $v_{12}$, additional changes in rule-clock structure is required as described in Section 16.6.
+Thus, after addition variable change, the total speed of the second electron decries in the Lorentz coordinatization from $v_{02}$ to $v_{01}$.
+This result stands in contrast with the speed of the first electron which remains $v_{02}$.
+In covariant particle tracking, information about difference in acceleration history between two setups is recorded in the covariant velocities difference.
+An equally valid interpretation exists within the non-covariant framework, which tracks particle motion using absolute time and applies Galilean transformations to field equations.
+The velocity of an electron and the speed of light are convention-dependent.
+In absolute-time coordinatization, the speed of light differs from the electrodynamics constant $c$ because Maxwell’s equations are not invariant under Galilean transformations.
+However, the dimensionless ratio of electron velocity to light velocity remains convention-independent, meaning it is unaffected by distant clock synchronization or variations in clock rhythms.
+In electrodynamics, parameters related to synchrotron radiation, which have direct physical significance, depend on this ratio.
+According to non-covariant particle tracking, the second electron’s velocity is equal $v_{02}$.
+However, Maxwell’s equations are not invariant under Galilean transformations, leading to a change in the speed of light.
+Specifically, the velocity of light increases from $c$, before the second boost with velocity $v_{12}$, to $c+v_{12}^{2}/(2c)]$ after the second boost.
+This variation arises because, under the absolute time convention, clocks are not resynchronized after the boost.
+As a result, the speed of light differs from the electrodynamical constant $c$ after the second boost.
+Nevertheless, the ratio of the electron velocity to the speed of light remains invariant across different synchronization conventions—that is, it does not depend on the method of distant clock synchronization or the rate of the clocks.
+Our calculations demonstrate that both covariant and non-covariant treatments—when the correct coupling between fields and particles is used—yield the same prediction for the relativistic electrodynamics phenomena.
+This prediction is convention-invariant and depends solely on the dimensionless parameter $v_{coord}/c_{coord}=[v_{01}+v_{12}^{2}/(2c)]/[c+v_{12}^{2}/(2c)]=v_{01}/c$, where $v_{coord}$ is the coordinate velocity of the second electron, and $c_{coord}$ is the coordinate velocity of light.
+This is illustrated that all relativistic effects can, in principle, be derived directly from underlying physical laws - without invoking relativity as an independent framework.
+Further details are provided in the final two chapters.
+We have shown that covariant particle tracking in the preferred frame can be achieved without switching reference frames.
+Initially, we employed conventional particle tracking—i.e., absolute time coordinatization—within the preferred frame.
+Then, by applying Galilean transformations alongside appropriate redefinitions of space and time variables, we derived the covariant particle trajectory, $\vec{x}(t)_{cov}$.
+From the standpoint of electrodynamics involving relativistically moving charges, this approach ensures consistency between Maxwell’s equations and particle trajectories within the preferred frame.
+There is, however, one important case where tracking requires a change of reference frame: the description of a particle’s angular momentum (spin) relative to the preferred frame.
+In the non-relativistic regime, spin is represented as a three-dimensional pseudovector $\vec{s}$.
+For a relativistically moving particle, by contrast, angular momentum must be described using the antisymmetric four-tensor $S_{\mu\nu}$.
+In experimental practice, however, the spin vector $\vec{s}$ remains the preferred object due to its intuitive physical interpretation as a three-dimensional pseudovector.
+As a result, any meaningful statement about $\vec{s}$ is understood to refer implicitly to the particle’s instantaneous rest frame—that is, its proper frame.
+As previously noted in Chapter 8, spin orientation measurements with respect to the preferred frame axes can be determined within the proper frame.
+In other words, spin rotation is interpreted from the viewpoint of the proper observer as the apparent rotation of the preferred frame axes.
+Successive Lorentz boosts to the proper frame are composed via matrix multiplication, assuming all transformations are expressed in the same (preferred) basis.
+However, under Lorentz coordinatization, moving frames inherently adopt different simultaneity conventions.
+As a result, the instantaneous spatial hypersurfaces of the proper and preferred frames differ and are not parallel.
+From a purely kinematic standpoint, spin rotation within the proper frame is absent.
+However, the preferred frame undergoes relativistic motion relative to the proper frame, which gives rise to the Wigner (Thomas) rotation of the preferred frame axes as seen from the proper frame.
+This rotation is given by $d\vec{\phi}_{W}=(\gamma-1)d\theta$, where $d\theta$ is the orbital deflection angle of the particle in the preferred frame.
+The direction of the preferred frame’s rotation, as observed in the proper frame, aligns with the direction of velocity rotation—both in the preferred and the proper frames.
+Accordingly, the spin vector rotates with respect to the preferred frame axes as $d\vec{s}=-d\vec{\phi}_{W}\times\vec{s}=-(\gamma-1)d\theta\times\vec{s}$.
+In the next chapter, we will continue our discussion of relativistic spin rotation.
 18 Relativistic Spin Dynamics
@@
 18.2 Derivation of the Covariant (BMT) Equation of Spin Motion
-In deriving the equation for spin motion, we rely on the well-established dynamics in the rest frame and known relativistic transformation laws.
-Since spin is inherently defined in a specific frame (the rest frame), constructing expressions with well-defined transformation properties requires introducing a four-quantity associated with spin.
-A natural choice is the four-(pseudo)vector $S$, defined such that in the rest frame, its spatial components correspond to the spin components, while its time-like component is zero.
-We refer to $S$ as the four-spin; when normalized by its invariant length, it becomes the polarization four-vector.
-Since it is space-like, its spatial part never vanishes in any frame.
+To derive the equation governing spin motion, we start from the well-established spin dynamics in the particle’s rest frame and apply known relativistic transformation laws.
+Since spin is inherently defined in the rest frame, constructing covariant expressions requires introducing a four-dimensional quantity associated with spin that transforms appropriately under Lorentz transformations.
+There are two general ways to generalize the spin vector $\vec{s}$—and, by extension, the Larmor equation $d\vec{s}/d\tau=\vec{\mu}\times\vec{B}_{R}$—to arbitrary frames: one approach uses an axial four-vector, while the other employs an antisymmetric rank-2 tensor.
+The most commonly used representation is the four-(pseudo)vector $S^{\alpha}$, often referred to as the four-spin vector.
+It is defined such that, in the rest frame, its spatial components coincide with the components of $\vec{s}$, while its time-like component vanishes.
+When normalized by its invariant magnitude, $S^{\alpha}$ becomes the polarization four-vector.
+Being space-like, its spatial components remain non-zero in all inertial frames.
+[^60] An arguably more physically natural representation is the antisymmetric four-tensor.
+The motion of a spinning point particle with mass $m$ and charge $e$ in Minkowski spacetime is then described by its polarization tensor $D_{\mu\nu}$—an antisymmetric four-tensor that incorporates both intrinsic magnetic and electric dipole moments.
+It is defined as $D_{\mu\nu}=geS_{\mu\nu}/(2mc)$, where $S_{\mu\nu}$ is the intrinsic angular momentum tensor.
 Let the spin of the particle be represented in the rest frame by $\vec{s}$.
@@
 Their approach closely resembled the four-tensor equations already established in relativistic particle dynamics.
-However, Eq.
-(28) is more intricate than it might initially appear.
-It consists of a system of coupled differential equations, making direct solutions challenging—even in highly symmetric cases, such as a uniform magnetic field.
-The question of how to solve this four-tensor equation is particularly intriguing.
-In relativistic spin dynamics, a specific and highly convenient method is employed to tackle this problem.
-To apply Eq.(28) to specific problems it is convenient to introduce a three vector $\vec{s}~{}$ by the equation $\vec{s}=\vec{S}+S^{0}\vec{p}c/(\mathcal{E}+mc^{2})$.
+To apply Eq.
+(28) to concrete physical scenarios, it is necessary to express the standard three-dimensional spin vector $\vec{s}$~ in terms of the four-vector $S$.
+The relationship between $\vec{s}$ and $S$ follows directly from the Lorentz transformation: $\vec{s}=\vec{S}+S^{0}\vec{p}c/(\mathcal{E}+mc^{2})$.
 With the help of this relation, one can work out the equation of motion for $\vec{s}$~.
@@
 It uses a spin quantity defined in the proper frame but observed with respect to the lab frame axes.
-That means that we know the orientation of the proper spin with respect to the lab coordinate system which is moving with velocity $-\vec{v}$ and acceleration $-\gamma d\vec{v}/d\tau$ in the proper frame.
+That means that we know the orientation of the proper spin with respect to the lab coordinate system which is moving with velocity $\vec{v}$ and acceleration $\gamma d\vec{v}/d\tau$ in the proper frame.
 Let’s look at what the equation Eq.(29) says in a little more detail.
@@
 where $\vec{v}_{cov}=d\vec{x}_{cov}/dt$. Here, the covariant trajectory of the particle, as observed in the laboratory frame, results from a sequence of Lorentz transformations.
-[…]
+In this book, we emphasize the significant role of the Wigner rotation as a regulator of the velocity addition law.
+As previously discussed, the four-velocity of an accelerating particle moving along a curved path in the Lorentz lab frame cannot, in general, be decomposed as $u=(c\gamma,\vec{v}\gamma)$.
+[^61] Consider, for instance, a typical textbook treatment [72], which discusses the projection of an arbitrary world line onto the Lorentz lab frame basis: ”The particle, which is assumed to curry the charge $e$, creates the current density $j(x)=ec\int d\tau u(y)\delta^{4}(y-x(\tau))$.
+[…] Furthermore, in any frame of reference $K$, one recovers the expected expressions for the charge and current densities by integrating over $\tau$ by means of relation $d\tau=dt^{\prime}/\gamma$ between proper time and coordinate time and using the formula $\delta(y_{0}-x_{0}(\tau))=\delta(ct-ct^{\prime})=\delta(t-t^{\prime})/c$, $j_{0}(t,y)=ce\delta^{(3)}(y-x(t))\equiv c\rho(t,y)$, $j^{i}(t,y)=ev^{i}(t)\delta^{(3)}(y-x(t))$, i=1,2,3i=1,2,3”.
+The presentation of the time component simply as the relation $d\tau=dt^{\prime}/\gamma$ between proper time and coordinate time is based on the hidden assumption that the type of clock synchronization, which provides the time coordinate $t$ in the lab frame, is based on the use of the absolute time convention.
 19.1 The Dipole Approximation
@@
 When considering only the dipole component of radiation, all details of the electron trajectory are effectively ignored.
-Consequently, dipole radiation is completely insensitive to the distinction between covariant and non-covariant particle trajectories  [^59] Similarly, the ultrarelativistic asymptotic limit also simplifies covariant calculations.
+Consequently, dipole radiation is completely insensitive to the distinction between covariant and non-covariant particle trajectories.
+[^62] Similarly, the ultrarelativistic asymptotic limit also simplifies covariant calculations.
 This is due to the paraxial approximation, which naturally follows from the ultrarelativistic assumption.
-More details on this effect will be discussed in the next chapter..
-[…]
+More details on this effect will be discussed in the next chapter.
+We want now to solve electrodynamics equations mathematically in a general way and consider the radiation associated with the succeeding terms in (multi-pole) expansion of the field in powers of the ratio $v/c$.
+Radiation theory is naturally developed in the space-frequency domain, as one is usually interested in radiation properties at a given position in space and at a certain frequency.
+[^63] In this book we define the relation between temporal and frequency domain via the following definition of Fourier transform pair: $\bar{f}(\omega)=\int_{-\infty}^{\infty}dt~f(t)\exp(i\omega t)\leftrightarrow f(t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}d\omega\bar{f}(\omega)\exp(-i\omega t)$.
 Suppose we are interested in the radiation generated by an electron and observed far away from it.
@@
 We then obtain the following approximate expression for the the radiation field in the space-frequency domain:
-[…]
-where $\omega$ is the frequency, $(-e)$ is the negative electron charge and we make use of Gaussian units.
+$\vec{\bar{E}}(\vec{r}_{0},\omega)$ $=$ ${i\omega e\over{cr_{0}}}\exp\left[-\frac{i\omega}{c}\vec{n}\cdot\vec{r}_{0}\right]\int_{-\infty}^{\infty}dt~{\vec{n}\times\left[\vec{n}\times{\vec{\beta}(t)}\right]}\exp\left[i\omega\left(t+\frac{\vec{n}\cdot\vec{r}(t)}{c}\right)\right]$ (36)
+where $\omega$ is the frequency, (e)(-e) is the negative electron charge and we make use of Gaussian units.
 A different constant of proportionality in Eq.(LABEL:revwied) and the well-known textbooks is to be ascribed to the use of different units and the definition of the Fourier transform.
@@
 This implies that $S^{\prime}$ is related to $S$ by a Lorentz boost $L(\vec{v})$, with $\vec{v}$ aligned along the $z$ axis.
-The boost transforms a four-vector even $X$ in space-time according to $X^{\prime}=L(\vec{v})X$.
+The boost transforms a four-vector event $X$ in space-time according to $X^{\prime}=L(\vec{v})X$.
 Now, let us analyze the particle’s dynamics from the perspective of the $S^{\prime}$ frame.
-In this frame, the electron remains at rest, while the kicker moves toward it with velocity $-\vec{v}$.
+In this frame, the electron remains at rest, while the kicker moves toward it with velocity $\vec{v}$.
 Due to its motion, the kicker’s magnetic field generates an electric field perpendicular to it.
@@
 Within this second-order approximation, relativistic corrections to velocity composition—typically arising at third order—do not appear.
-[…]
+We begin by analyzing the textbook treatment of the composition of motions.
+Let SS" be a reference frame fixed with respect to a particle located downstream of the kicker.
+As is well known, non-collinear Lorentz boosts do not commute.
+However, in our second-order approximation, we can neglect the difference between $\gamma v_{y}/c$ and $\gamma_{z}v_{y}/c$, where $\gamma_{z}=1/\sqrt{1-v_{z}^{2}/c^{2}}$.
+Here $v_{z}=v(1-\theta_{k}^{2}/2)$ and $\theta_{k}=v_{y}/v=v_{y}/c$ in the ultrarelativistic regime of interest.
+Therefore, we can approximate the transformation from coordinates $X^{\prime}$ in frame $S^{\prime}$ to coordinates XX" in SS" using a sequence of two commuting non-collinear Lorentz boosts: $X"=L(\vec{e^{\prime}}_{y}v^{\prime}_{y})L(\vec{e^{\prime}}_{z}v^{\prime}_{z})X^{\prime}=L(\vec{e^{\prime}}_{z}v^{\prime}_{z})L(\vec{e^{\prime}}_{y}v^{\prime}_{y})X^{\prime}$, where $\vec{e^{\prime}}_{y}$ and $\vec{e^{\prime}}_{z}$ are unit vectors along the $x^{\prime}$ and $z^{\prime}$ axis respectively.
+It is important to note that, as seen by an observer in $S^{\prime}$, the axes of frameSS" remain parallel to those of $S^{\prime}$.
+The full transformation from the lab frame $S$ to the downstream particle frame SS" can thus be written as: $X"=L(\vec{e^{\prime}}_{y}v^{\prime}_{y})L(\vec{e^{\prime}}_{z}v^{\prime}_{z})L(\vec{e}_{z}v)X$.
+In the special case where the velocities are collinear, the composition of boosts simplifies: $L(\vec{e^{\prime}}_{z}v^{\prime}_{z})L(\vec{e}_{z}v)=L(\vec{e}_{z}v_{z})$, leading to the overall transformation: $X"=L(\vec{e^{\prime}}_{y}v^{\prime}_{y})L(\vec{e}_{z}v_{z})X$.
+Textbooks typically emphasize that, within the Lorentz framework, a magnetic field can only alter the direction of an electron’s motion, but not its speed.
 According to standard textbooks, in relativistic kinematics, the composition of perpendicular velocities is analyzed using Einstein’s velocity addition theorem.
@@
 This result highlights that the apparent contradiction is rooted in the presence of the Wigner rotation.
-As discussed in earlier chapters, the relation $dt^{\prime}=dt/\gamma$ is not valid when addressing the relativistic composition of perpendicular velocities within the Lorentz framework.
-It is important to highlight another effective approach to covariant electron tracking.
-We now revisit the motion of a relativistic electron accelerated by the kicker field, analyzing it from the perspective of an inertial lab observer without changing reference frames.
-Specifically, we consider an electron in the lab frame undergoing acceleration due to the kicker field—an example of an active boost.
-In this scenario, we track the evolution of the system over time within a single reference frame.
+Now let us revisit the kinematics of a relativistic electron accelerated by the kicker field, analyzing it from the perspective of an inertial lab observer without changing reference frames.
+Specifically, we consider an electron in the lab frame undergoing acceleration due to the kicker field.
 The simplest synchronization method involves maintaining the same set of uniformly synchronized clocks as in the pre-kick state, without modification.
@@
 The standard Galilean velocity addition rule is then applied to track the electron’s motion instant by instant as it moves through a constant magnetic field.
-In Chapter 3, we discussed how absolute time coordinatization can be transformed into Lorentz coordinatization.
 By combining Galilean transformations with variable changes, we ultimately derive the Lorentz transformation within the framework of absolute time coordinatization in the lab frame.
-To further illustrate this, we now examine how distant lab clocks resynchronize during the electron’s acceleration.
 This analysis establishes a direct link between the electron’s speed reduction after the kick (in Lorentz coordinates) and time dilation.
-As previously noted, the Lorentz coordinate system is purely a conceptual construct.
-Hypothetical clocks are required only for applying Maxwell’s equations in synchrotron radiation calculations.
 Suppose that, upstream of the kicker, we choose a Lorentz coordinate system in the lab frame.
@@
 Consequently, the post-kick motion can be described as a composition of two Lorentz boosts in the perpendicular $y$ and $z$ directions.
-The first boost imparts a velocity $v\theta_{k}\vec{e}_{y}$ along the $y$-axis, while the second applies an additional velocity $-(v\theta_{k}^{2}/2)\vec{e}_{z}$ along the $z$ axis.
+The first boost imparts a velocity $v\theta_{k}\vec{e}_{y}$ along the $y$-axis, while the second applies an additional velocity $(v\theta_{k}^{2}/2)\vec{e}_{z}$ along the $z$ axis.
 The second-order approximation ensures that the two boosts commute.
@@
 As a result, the total electron speed in the lab frame, after clock resynchronization downstream of the kicker, decreases from $v$ to $v(1-\theta_{k}^{2}/2)$.
-This analysis is performed in the limit of a small kick angle $\gamma v_{y}/c\ll 1$.
-Nevertheless, even in this simplified case and for a single electron, we are able to clearly demonstrate the distinction between covariant and non-covariant particle trajectories.
 19.3 Electron Motion Accelerated by a Kicker in a Bending Magnet
@@
 Here, we observe a second-order correction in $\theta_{k}^{2}$ that is, however, multiplied by a large factor $\gamma^{2}$.
-It is worth noting, however, that there exists an alternative and equally satisfactory explanation for the redshift.
-This interpretation employs a non-covariant framework, which considers non-covariant particle trajectories and applies Galilean transformations to the electromagnetic field equations.
-In this approach, the electron velocity remains unchanged according to non-covariant particle tracking.
-However, Maxwell’s equations are not invariant under Galilean transformations, leading to a change in the speed of light.
-Specifically, the velocity of light increases from $c$ (before the kick) to $c(1+\theta_{k}^{2}/2)$ (after the kick).
-This variation arises because, under the absolute time convention, clocks are not resynchronized after the kick.
-As a result, the speed of light differs from the electrodynamical constant $c$.
-Nevertheless, the ratio of the electron velocity to the speed of light remains invariant across different synchronization conventions—that is, it does not depend on the method of distant clock synchronization or the rate of the clocks.
-Our calculations demonstrate that both covariant and non-covariant treatments—when the correct coupling between fields and particles is used—yield the same prediction for the redshift.
-This prediction is convention-invariant and depends solely on the dimensionless parameter $v_{coord}/c_{coord}=v(1-\theta^{2}_{k}/2)/c=v/[c(1+\theta^{2}_{k}/2)]$, where $v_{coord}$ is the coordinate velocity of the electron, and $c_{coord}$ is the coordinate velocity of light.
 To validate the predictions of our synchrotron radiation theory, we propose an experimental test using third-generation synchrotron radiation sources.
@@
 In essence, we argue that, despite decades of experimental work, synchrotron radiation theory remains insufficiently confirmed, and a more precise test is both timely and necessary.
-18.5 Discussion
-We now wish to emphasize that when an electron is accelerated in the lab frame by a kicker, this acceleration is encoded in its covariant trajectory.
-A compelling question arises: Where, physically, is the information about the electron’s acceleration located when using absolute time coordinatization?
-[…]
-This apparent paradox—of an unchanged velocity despite acceleration—can be explained through dynamical reasoning involving the electron’s non-covariant trajectory.
-Without entering into a full derivation, we summarize the key insight as follows:
-For a rapidly moving electron, the transverse electric and magnetic fields are nearly equal in magnitude and orthogonal to each other—indistinguishable from the fields of an actual radiation beam.
-On a microscopic level, when the electron passes through a kicker, its electromagnetic fields are perturbed and now carry imprints of that acceleration.
-Under the framework of classical (Galilean) kinematics, the orientation of the virtual radiation phase front remains unchanged.
-However, since Maxwell’s equations are not invariant under Galilean transformations—as discussed throughout this article—the adoption of such kinematics necessitates the use of anisotropic field equations.
-Consequently, although the phase front remains planar, the direction of propagation is no longer perpendicular to it.
-In other words, the radiation beam’s direction of motion diverges from the normal to its phase front.
-Within the absolute time synchronization convention, electrodynamics predicts that the virtual radiation beam propagates along the direction of the kick, with a phase front tilt $\theta_{k}$.
-This tilt is the crucial clue to resolving the paradox.
-The key insight is this: The information about the electron’s acceleration is not lost—it is embedded in the perturbation of the electron’s self-fields.
-Mathematically, this information is encoded in the tilt of the phase front of the virtual radiation beam.
-Let us now briefly consider some additional, intriguing cases.
-Previously, we examined a simple scenario in which the kick angle is small.
-In such cases, when an electron passes through a weak kicker, it emits no appreciable synchrotron radiation.
-More precisely, the radiation is indistinguishable from the electron’s intrinsic electromagnetic fields.
-We can therefore assume that the electron remains in a “field-dressed” state — still accompanied by its usual virtual radiation fields — after passing through the weak kicker.
-Now, let us turn to a different situation.
-Suppose an ultrarelativistic electron in the laboratory frame is deflected by a dipole magnetic field with a kick angle much larger than $1/\gamma$, This setup corresponds to a typical synchrotron radiation scenario.
-At the exit of the strong kicker, the electron emerges effectively “naked” or “field-free,” meaning it is no longer accompanied by its usual virtual field structure.
-A process then begins in which the “field-dressed” state is gradually re-established — that is, the characteristic electromagnetic fields associated with a fast-moving charge begin to reform — over a certain formation length in the downstream straight section.
-At this point, information about the electron’s recent acceleration is carried away by the emitted synchrotron radiation.
-However, the electron’s original self-fields and its acceleration history are effectively erased in the strong kick.
-As a result, the electromagnetic fields surrounding the electron downstream of the kicker no longer encode this history.
-In other words, the electron’s covariant trajectory beyond the kicker contains no imprint of the preceding acceleration.
-We will revisit this topic in more detail in the next chapter.
 20 Synchrotron Radiation
@@
 (LABEL:revwied) as follows
-[…]
+$\vec{\bar{E}}(\vec{r}_{0},\omega)$ $=$ $-{i\omega e\over{cr_{0}}}\exp\left[-\frac{i\omega}{c}\vec{n}\cdot\vec{r}_{0}\right]\int_{-\infty}^{\infty}dt~[\vec{\beta}(t)-\vec{n}]\exp\left[i\omega\left(t+\frac{\vec{n}\cdot\vec{r}(t)}{c}\right)\right]$ (38)
 Eq. (LABEL:revwied) and Eq. (LABEL:rrevwied) are equivalent but include different integrands. This is no mistake, as different integrands can lead to the same integral (see Appendix A1).
@@
 The emission angle $\theta=\sqrt{\theta_{x}^{2}+\theta_{y}^{2}}$ is taking with respect to the $z$-axis.
-Here $\theta_{x}=x_{0}/z_{0}$ is the observation angle projected onto the $x-z$ plane, $\theta_{y}=y_{0}/z_{0}$ is the observation angle projected onto the $y-z$ plane.
+Here $\theta_{x}=x_{0}/z_{0}$ is the observation angle projected onto the xzx-z plane, $\theta_{y}=y_{0}/z_{0}$ is the observation angle projected onto the yzy-z plane.
 The components of the unit vector $\vec{n}$ can be approximated by $n_{z}=1-\theta_{x}^{2}/2-\theta_{y}^{2}/2$, $n_{x}=\theta_{x}$, $n_{y}=\theta_{y}$, so $\vec{n}\cdot(\vec{r}_{0}-\vec{r})=(z_{0}-z)(1-\theta_{x}^{2}/2-\theta_{y}^{2}/2)-x\theta_{x}-y\theta_{y}$.
@@
 (LABEL:rrevwied) in the far zone and paraxial approximation finally becomes
-[…]
+$\vec{\widetilde{{E}}}(z_{0},\vec{r}_{0},\omega)$ $=$ $-{i\omega e\over{c^{2}}z_{0}}\int_{-\infty}^{\infty}dz^{\prime}{\exp{\left[i\Phi_{T}\right]}}\left[\left({v_{x}(z^{\prime})\over{c}}-\theta_{x}\right){\vec{e_{x}}}+\left({v_{y}(z^{\prime})\over{c}}-\theta_{y}\right){\vec{e_{y}}}\right]$ (43)
 where the total phase $\Phi_{T}$ is
@@
 We write the undulator length as $L=N_{w}\lambda_{w}$, where $N_{w}$ is the number of undulator periods. With the help of Eq. (43) we obtain an expression, valid in the far zone:
-[…]
+${\vec{\widetilde{E}}}={i\omega e\over{c^{2}z_{0}}}\int_{-L/2}^{L/2}dz^{\prime}{\exp\left[i\Phi_{T}\right]\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]}\left[{K\over{\gamma}}\sin\left(k_{w}z^{\prime}\right)\vec{e}_{x}+\vec{\theta}\right]~.$ (50)
 Here
@@
 The resulting distribution for the slowly varying envelope of the electric field is given by:
-[…]
+${\vec{\widetilde{E}}}=-\frac{K\omega e}{2c^{2}z_{0}\gamma}A_{JJ}\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\int_{-L/2}^{L/2}dz^{\prime}\exp\left[i\left(C+{\omega\theta^{2}\over{2c}}\right)z^{\prime}\right]\vec{e}_{x}~,$ (55)
 Here $\omega=\omega_{r}+\Delta\omega$, $C=k_{w}\Delta\omega/\omega_{r}$ and
@@
 We analyze the radiation associated with the first-order term in the expansion of Eq.
-(147) in powers of the small parameter $[K\theta_{x}\omega/(\gamma k_{w}c)]$ (see Eq.
+(147) in powers of the small parameter $K\theta_{x}\omega/(\gamma k_{w}c)]$ (see Eq.
 (150)).
@@
 When performing this rotation, it is important to note that the phase factor in Eq.
-(61)—which represents a spherical wavefront emanating from $z=0$—remains invariant under rotation.
+(61)—which represents a spherical wavefront emanating from z=0z=0—remains invariant under rotation.
 In contrast, the argument of the $\mathrm{sinc}(\cdot)$ function in Eq.
@@
 Within this framework, introducing a kick leads to a non-zero red shift in the resonance frequency.
-This shift arises because, in Lorentz coordinates, the electron’s longitudinal velocity is reduced—from $v$ to $v-v\theta_{k}^{2}/2$while the speed of light remains constant at its electrodynamical value, $c$.
+This shift arises because, in Lorentz coordinates, the electron’s longitudinal velocity is reduced—from $v$ to $v-v\theta_{k}^{2}/2$ -while the speed of light remains constant at its electrodynamical value, $c$.
 Consequently, the expression given in Eq.
@@
 One way to highlight the incompatibility between the standard approach to relativistic electrodynamics, which typically involves Maxwell’s equations, and the particle trajectories derived from non-covariant particle tracking is through a direct laboratory test of synchrotron radiation theory.
-Essentially, we are asserting that, despite extensive measurements accumulated over several decades, synchrotron radiation theory has yet to be definitively confirmed by experimental evidence.
 Let us now explore the potential of synchrotron radiation sources to validate the predictions of a revised synchrotron radiation theory.
@@
 With Eq. (43) we obtain the radiation field amplitude in the far zone:
-[…]
+$\vec{\widetilde{E}}={i\omega e\over{c^{2}z_{0}}}\int_{-\infty}^{\infty}dz^{\prime}{e^{i\Phi_{T}}}\left({z^{\prime}+R\theta_{x}\over{R}}\vec{e_{x}}+\theta_{y}\vec{e_{y}}\right)$~ (77)
 where
@@
 With redefinition of $z^{\prime}$ as $z^{\prime}+R\theta_{x}$ under integral we obtain the final result:
-[…]
-×$\times\exp\left\{{i\omega\left[{z^{\prime}\over{2\gamma^{2}c}}\left(1+\gamma^{2}\theta_{y}^{2}\right)+{z^{\prime 3}\over{6R^{2}c}}\right]}\right\}~{},$ (83)
+$\vec{\widetilde{E}}={i\omega e\over{c^{2}z_{0}}}e^{i\Phi_{s}}e^{i\Phi_{0}}\int_{-\infty}^{\infty}dz^{\prime}\left({z^{\prime}\over{R}}\vec{e_{x}}+\theta_{y}\vec{e_{y}}\right)$ (82)
+$\times\exp\left\{{i\omega\left[{z^{\prime}\over{2\gamma^{2}c}}\left(1+\gamma^{2}\theta_{y}^{2}\right)+{z^{\prime 3}\over{6R^{2}c}}\right]}\right\}~,$ (83)
 where
@@
 Taking this into account, we define the modified angles $\bar{\theta}_{x}=\theta_{x}-l_{x}/z_{0}$ and $\bar{\theta}_{y}=\theta_{y}-l_{y}/z_{0}$ in order to obtain
-[…]
+$\vec{\widetilde{E}}={i\omega e\over{c^{2}z_{0}}}\int_{-\infty}^{\infty}dz^{\prime}{e^{i\Phi_{T}}}\left({z^{\prime}+R(\bar{\theta}_{x}-\eta_{x})\over{R}}\vec{e_{x}}+{(\bar{\theta}_{y}-\eta_{y})}\vec{e_{y}}\right)$~ (89)
 and
@@
 $\Phi_{T}=\omega\left({\bar{\theta}_{x}^{2}+\bar{\theta}_{y}^{2}\over{2c}}z_{0}\right)-{\omega R(\bar{\theta}_{x}-\eta_{x})\over{2c}}$ (92)
-×$\times\left({1\over{\gamma^{2}}}+(\bar{\theta}_{y}-\eta_{y})^{2}+{(\bar{\theta}_{x}-\eta_{x})^{2}\over{3}}\right)$ (93)
+$\times\left({1\over{\gamma^{2}}}+(\bar{\theta}_{y}-\eta_{y})^{2}+{(\bar{\theta}_{x}-\eta_{x})^{2}\over{3}}\right)$ (93)
 $+\left({{1\over{\gamma^{2}}}+(\bar{\theta}_{y}-\eta_{y})^{2}}\right){\omega\left(z^{\prime}+R(\bar{\theta}_{x}-\eta_{x})\right)\over{2c}}$ (94)
@@
 Redefinition of $z^{\prime}$ as $z^{\prime}+R(\bar{\theta}_{x}-\eta_{x})$ gives the result
-[…]
-×$\times\exp\left\{{i\omega\left[{z^{\prime}\over{2\gamma^{2}c}}\left(1+\gamma^{2}(\bar{\theta}_{y}-\eta_{y})^{2}\right)+{z^{\prime 3}\over{6R^{2}c}}\right]}\right\}~{},$ (97)
+$\vec{\widetilde{E}}={i\omega e\over{c^{2}z_{0}}}e^{i\Phi_{s}}e^{i\Phi_{0}}\int_{-\infty}^{\infty}dz^{\prime}\left({z^{\prime}\over{R}}\vec{e_{x}}+(\bar{\theta}_{y}-\eta_{y})\vec{e_{y}}\right)$ (96)
+$\times\exp\left\{{i\omega\left[{z^{\prime}\over{2\gamma^{2}c}}\left(1+\gamma^{2}(\bar{\theta}_{y}-\eta_{y})^{2}\right)+{z^{\prime 3}\over{6R^{2}c}}\right]}\right\}~,$ (97)
 where
@@
 In the far zone we can neglect terms in $l_{x}/z_{0}$ and $l_{y}/z_{0}$, which leads to
-[…]
-×$\times\exp\left\{{i\omega\left[{z^{\prime}\over{2\gamma^{2}c}}\left(1+\gamma^{2}\left(\theta_{y}-\eta_{y}\right)^{2}\right)+{z^{\prime 3}\over{6R^{2}c}}\right]}\right\}~{},$ (101)
+$\vec{\widetilde{E}}={i\omega e\over{c^{2}z_{0}}}e^{i\Phi_{s}}e^{i\Phi_{0}}\int_{-\infty}^{\infty}dz^{\prime}\left({z^{\prime}\over{R}}\vec{e_{x}}+\left(\theta_{y}-\eta_{y}\right)\vec{e_{y}}\right)$ (100)
+$\times\exp\left\{{i\omega\left[{z^{\prime}\over{2\gamma^{2}c}}\left(1+\gamma^{2}\left(\theta_{y}-\eta_{y}\right)^{2}\right)+{z^{\prime 3}\over{6R^{2}c}}\right]}\right\}~,$ (101)
 where
@@
 (LABEL:revwied).
-It is important to note that, in principle, the entire history of the electron - from $t^{\prime}=-\infty$ to $t^{\prime}=\infty$must be considered, since the integral in Eq.
+It is important to note that, in principle, the entire history of the electron - from $t^{\prime}=$-$\infty$ to $t^{\prime}=\infty$ -must be considered, since the integral in Eq.
 (LABEL:revwied) extends over this range.
@@
 In other words, we examine a standard synchrotron radiation setup, where an electron enters the system through a bending magnet, traverses a straight section, passes through an undulator, continues along another straight section, and exits via a second bending magnet.
-Although the integration in Eq.(43) is performed from $\infty$- ∞ to $\infty$, the only (edge) part of the trajectory into the bending magnets contributing to the integral is of the order of the radiation formation length $L_{f}$.
+Although the integration in Eq.(43) is performed from -$\infty$ to $\infty$, the only (edge) part of the trajectory into the bending magnets contributing to the integral is of the order of the radiation formation length $L_{f}$.
 Mathematically, this is reflected in the behavior of $\Phi_{T}(z^{\prime})$ in Eq.(43), which exhibits increasingly rapid oscillations as $z^{\prime}$ exceeds the formation length.
@@
 (101) can then be expressed in terms of modified Bessel functions:
-[…]
-[…]
+$\int_{0}^{\infty}x\sin[(3/2)\alpha(x+x^{3}/3)]dx=(1/\sqrt{3})K_{2/3}(\alpha)~,$ (106)
+$\int_{0}^{\infty}\cos[(3/2)\alpha(x+x^{3}/3)]dx=(1/\sqrt{3})K_{1/3}(\alpha)~.$ (107)
 Then, making the necessary variable changes, the formula reduces to Eq.( 105).
@@
 In the conventional framework, this wavefront tilt is treated as a physical phenomenon.
-That is, a transverse kick does not alter the orientation of the modulation wavefront, thereby reducing the efficiency of radiation in the direction of the beam motion  [^66] In standard XFEL operation, the electron beam is guided to remain closely aligned with the undulator axis.
+That is, a transverse kick does not alter the orientation of the modulation wavefront, thereby reducing the efficiency of radiation in the direction of the beam motion.
+[^68] In standard XFEL operation, the electron beam is guided to remain closely aligned with the undulator axis.
 Nonetheless, random perturbations in the focusing system can introduce angular deviations or ”kicks.” The resultant misalignment between beam trajectory and wavefront normal has been discussed in the literature.
@@
 Studies of trajectory errors in XFEL amplification have shown that undulator magnetic fields must meet stringent tolerances.
-Interestingly, these tolerances have proven to be more stringent than necessary according to the corrected, four-dimensional (4D) covariant XFEL theory—an insight that helps explain the remarkable success of XFEL technology in recent decades..
+Interestingly, these tolerances have proven to be more stringent than necessary according to the corrected, four-dimensional (4D) covariant XFEL theory—an insight that helps explain the remarkable success of XFEL technology in recent decades.
 Figure 61: Result of covariant particle tracking.
@@
 This raises an important question: how should synchronization be defined in the lab frame after this acceleration?
-Prior to the kicker setup, we chose a Lorentz coordinate system for the lab frame  [^67] Conventional XFEL theory adopts the absolute time convention to describe particle dynamics.
+Prior to the kicker setup, we chose a Lorentz coordinate system for the lab frame.
+[^69] Conventional XFEL theory adopts the absolute time convention to describe particle dynamics.
 Notably, employing the standard coupling of Maxwell’s equations with the corrected Newtonian equations to compute radiation from an ultrarelativistic electron beam in an undulator does not inherently introduce inaccuracies.
@@
 The most straightforward synchronization method employs the same set of synchronized clocks used in conventional particle tracking.
-This approach defines simultaneity based on light signals emitted from a stationary dipole source, assuming isotropic light propagation at speed $c$ in all directions..
+This approach defines simultaneity based on light signals emitted from a stationary dipole source, assuming isotropic light propagation at speed $c$ in all directions.
 After the beam acquires a small transverse velocity $v_{y}$, maintaining the original synchronization leads to complications in the electrodynamics of moving charges.
@@
 It is necessary to mention that in XFEL engineering we deal with very high-quality undulator systems, which have a sufficiently wide good-field-region, so that our studies, which refer to a simple model of undulator field nevertheless yield a correct quantitative description in a large variety of practical problems.
-The Lorentz force $\vec{F}=-e\vec{v}\times\vec{B}_{w}/c$ is used to derive the equation of motion of electrons with charge $-e$ and mass $m$ in the presence of magnetic field
+The Lorentz force $\vec{F}=-e\vec{v}\times\vec{B}_{w}/c$ is used to derive the equation of motion of electrons with charge e-e and mass $m$ in the presence of magnetic field
 $m\gamma\frac{dv_{x}}{dt}=\frac{e}{c}v_{z}B_{y}=-\frac{e}{c}v_{z}B_{w}\sin(k_{w}z)~,$ (109)
@@
 With the help of the identity
-×(×$\vec{\nabla}\times(\vec{\nabla}\times{\vec{E}})=\vec{\nabla}(\vec{\nabla}\cdot{\vec{E}})-\nabla^{2}\vec{E}$ (118)
+$\vec{\nabla}\times(\vec{\nabla}\times{\vec{E}})=\vec{\nabla}(\vec{\nabla}\cdot{\vec{E}})-\nabla^{2}\vec{E}$ (118)
 and Poisson equation
@@
 The field of electromagnetic wave has only transverse components, so the energy exchange between the electron and electromagnetic wave is due to a transverse component of the electron velocity.
-For effective energy exchange between the electron and the wave, the scalar product $-e\vec{v}_{\perp}\cdot\vec{E}$ should be kept nearly constant along the whole undulator length.
+For effective energy exchange between the electron and the wave, the scalar product $e\vec{v}_{\perp}\cdot\vec{E}$ should be kept nearly constant along the whole undulator length.
 We see that required synchronism $k_{w}+\omega/c-\omega/v_{z}=0$ takes place when the wave advances the electron beam by the wavelength at one undulator period $\lambda_{w}/v_{z}=\lambda/(c-v_{z})$, where $\lambda=2\pi/\omega$ is the radiation wavelength.
@@
 with $|\vec{r}-\vec{r^{\prime}}|=\sqrt{(x^{\prime}-x)^{2}+(y^{\prime}-y)^{2}+(z^{\prime}-z)}^{2}$. With the help of this Green function we can write a formal solution for the field equation as:
-[…]
+$\vec{\bar{E}}=\int d\vec{r^{\prime}}~G(\vec{r},\vec{r^{\prime}})\left[-4\pi i\frac{\omega}{c^{2}}\vec{\bar{j}_{\perp}}\right]~.$ (124)
 This is simply a mathematical description of the concept of Huygens’ secondary sources and wave propagation, which is, of course, well-known.
@@
 The diagonal path length $s$ exceeds the base by an amount $\Delta=s-z\simeq z\theta^{2}/2$.
-When $\Delta$ equals one radiation wavelength, destructive interference occurs, as the phase contributions from different oscillators become uniformly distributed over 00 to $2\pi$.
+When $\Delta$ equals one radiation wavelength, destructive interference occurs, as the phase contributions from different oscillators become uniformly distributed over 0 to $2\pi$.
 In the limit of a small electron beam size ($\sigma\to 0$), constructive interference occurs within an angle $\Delta\theta\backsimeq\sqrt{c/(\omega L_{w})}=1/(\sqrt{4\pi N_{w}}\gamma_{z})\ll 1/\gamma$, where $L_{w}=\lambda_{w}N_{w}$ is the undulator length.
@@
 This line of reasoning reinforces the conclusion that the results of the beam-splitting experiment in [79] validate our correction to the conventional theory of spontaneous undulator emission.
-20.7 Debunking the Modulation Wavefront Tilt Myth
-In conventional XFEL theory, wavefront tilt is typically regarded as a physically real effect.
-However, a common misconception in accelerator physics relates to the interpretation of this tilt.
-In the ultrarelativistic regime, wavefront tilt does not have an exact, objective meaning.
-Its value depends on the choice of clock synchronization procedure in the lab frame and, consequently, can be assigned arbitrarily within the interval $(0,\theta_{k})$.
+21.7 Wavefront Tilt and Degradation of Electron Beam Modulation
+In conventional XFEL theory, wavefront tilt is often treated as a physically meaningful and measurable quantity.
+However, a common misconception in accelerator physics concerns the interpretation of this tilt.
+In the ultrarelativistic regime, wavefront tilt lacks a unique, objective definition.
+Its value depends on the choice of clock synchronization convention in the laboratory frame and can therefore be assigned arbitrarily within the interval $(0,\theta_{k})$.
 For example, when the evolution of a modulated electron beam is described using Lorentz coordinates, the orientation of the modulation wavefront is always perpendicular to the electron beam velocity—i.e., $\theta_{tilt}=0$.
@@
 We start with the solution of Maxwell’s equation in the space-time domain, the well-known Lienard-Wiechert expression, and we subsequently apply a Fourier transformation.
-The Lienard-Wiechert expression for the electric field of a point charge $(-e)$ reads (see, e.g.
+The Lienard-Wiechert expression for the electric field of a point charge (e)(-e) reads (see, e.g.
 [41]):
@@
 Let us apply a Fourier transformation:
-[…]
-[…]
+$\vec{\bar{E}}(\vec{r}_{o},\omega)$ $=$ $-e\int_{-\infty}^{\infty}dt^{\prime}{\vec{n}-\vec{\beta}\over{\gamma^{2}(1-\vec{n}\cdot\vec{\beta})^{2}|\vec{r}_{o}-\vec{r^{\prime}}|^{2}}}\exp\left[i\omega\left(t^{\prime}+{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right]$ (127)
+$-{e\over{c}}\int_{-\infty}^{\infty}dt^{\prime}{\vec{n}\times[(\vec{n}-\vec{\beta})\times\dot{\vec{\beta}}]\over{(1-\vec{n}\cdot\vec{\beta})^{2}|\vec{r}_{o}-\vec{r^{\prime}}|}}\exp\left[i\omega\left(t^{\prime}+{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right].$
 As in Eq.
@@
 Eq. (127) can be written as
-$\vec{\bar{E}}$ (((( $\vec{r}_{o},\omega)=-{e}\int_{-\infty}^{\infty}dt^{\prime}{\vec{n}\over{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|^{2}}}\exp\left[i\omega\left(t^{\prime}+{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right]$ (130)
-[…]
+$\vec{\bar{E}}$ $($ $\vec{r}_{o},\omega)=-{e}\int_{-\infty}^{\infty}dt^{\prime}{\vec{n}\over{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|^{2}}}\exp\left[i\omega\left(t^{\prime}+{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right]$ (130)
+$+{e\over{c}}\int_{-\infty}^{\infty}dt^{\prime}{d\over{dt^{\prime}}}\left[{\vec{\beta}-\vec{n}\over{({1-\vec{n}\cdot\vec{\beta}})|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|}}\right]\exp\left[i\omega\left(t^{\prime}+{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right].$
 Eq. (LABEL:revtrasfbiss) may now be integrated by parts. When edge terms can be dropped one obtains [84]
-[…]
-×$\times\exp\left\{i\omega\left(t^{\prime}+{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right\}~{}.$
+$\vec{\bar{E}}(\vec{r_{o}},\omega)$ $=$ $-{i\omega e\over{c}}\int_{-\infty}^{\infty}dt^{\prime}\left[{\vec{\beta}-\vec{n}\over{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|}}-{ic\over{\omega}}{\vec{n}\over{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|^{2}}}\right]$ (133)
+$\times\exp\left\{i\omega\left(t^{\prime}+{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right\}~.$
 The only assumption made going from Eq.
@@
 (135) in the integration by parts, we obtain
-[…]
-×$\times\exp\left\{i\omega\left(t^{\prime}+{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right\}~{}.$
+$\vec{\bar{E}}(\vec{r_{o}},\omega)$ $=$ $-{i\omega e\over{c}}\int_{-\infty}^{\infty}dt^{\prime}\left[-{\vec{n}\times(\vec{n}\times{\vec{\beta}})\over{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|}}+{i{c}\over{\omega}}{\vec{\beta}-\vec{n}-2\vec{n}(\vec{n}\cdot\vec{\beta})\over{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|^{2}}}\right]$ (137)
+$\times\exp\left\{i\omega\left(t^{\prime}+{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right\}~.$
 Similarly as before, the edge terms have been dropped.
@@
 ${\vec{\widetilde{E}}}=\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\frac{i\omega e}{c^{2}z_{0}}$ (138)
-×$\times\int_{-L/2}^{L/2}dz^{\prime}\left\{\frac{K}{2i\gamma}\left[\exp\left(2ik_{w}z^{\prime}\right)-1\right]\vec{e}_{x}+\vec{\theta}\exp\left(ik_{w}z^{\prime}\right)\right\}$ (139)
-×$\times\exp\left[i\left(C+{\omega\theta^{2}\over{2c}}\right)z^{\prime}-{K\theta_{x}\over{\gamma}}{\omega\over{k_{w}c}}\cos(k_{w}z^{\prime})-{K^{2}\over{8\gamma^{2}}}{\omega\over{k_{w}c}}\sin(2k_{w}z^{\prime})\right]~{}.$ (140)
+$\times\int_{-L/2}^{L/2}dz^{\prime}\left\{\frac{K}{2i\gamma}\left[\exp\left(2ik_{w}z^{\prime}\right)-1\right]\vec{e}_{x}+\vec{\theta}\exp\left(ik_{w}z^{\prime}\right)\right\}$ (139)
+$\times\exp\left[i\left(C+{\omega\theta^{2}\over{2c}}\right)z^{\prime}-{K\theta_{x}\over{\gamma}}{\omega\over{k_{w}c}}\cos(k_{w}z^{\prime})-{K^{2}\over{8\gamma^{2}}}{\omega\over{k_{w}c}}\sin(2k_{w}z^{\prime})\right]~.$ (140)
 Here $\omega=\omega_{r}+\Delta\omega$, $C=k_{w}\Delta\omega/\omega_{r}$ and
@@
 ${\vec{\widetilde{E}}}=\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\frac{i\omega e}{c^{2}z_{0}}\sum_{m,n=-\infty}^{\infty}J_{m}(u)J_{n}(v)\exp\left[\frac{i\pi n}{2}\right]$ (144)
-×$\times\int_{-L/2}^{L/2}dz^{\prime}\exp\left[i\left(C+{\omega\theta^{2}\over{2c}}\right)z^{\prime}\right]$ (145)
-×$\times\left\{\frac{K}{2i\gamma}\left[\exp\left(2ik_{w}z^{\prime}\right)-1\right]\vec{e}_{x}+\vec{\theta}\exp\left(ik_{w}z^{\prime}\right)\right\}\exp\left[i(n+2m)k_{w}z^{\prime}\right]~{},$ (146)
+$\times\int_{-L/2}^{L/2}dz^{\prime}\exp\left[i\left(C+{\omega\theta^{2}\over{2c}}\right)z^{\prime}\right]$ (145)
+$\times\left\{\frac{K}{2i\gamma}\left[\exp\left(2ik_{w}z^{\prime}\right)-1\right]\vec{e}_{x}+\vec{\theta}\exp\left(ik_{w}z^{\prime}\right)\right\}\exp\left[i(n+2m)k_{w}z^{\prime}\right]~,$ (146)
 where [^70] Here the parameter $v$ should not be confused with the velocity.
@@
 $|v|={K|\theta_{x}|\over{\gamma}}{\omega\over{k_{w}c}}=\left(1+\frac{\Delta\omega}{\omega_{r}}\right)\frac{2\sqrt{2}K}{\sqrt{2+K^{2}}}\bar{\gamma}_{z}|\theta_{x}|\lesssim\bar{\gamma}_{z}|\theta_{x}|\ll 1~.$ (150)
-[…]
+This means that, independently of $K$, $|v|\ll 1$ and we may expand $J_{n}(v)$ in Eq.
+(147) according to $J_{n}(v)\simeq[2^{-n}/\Gamma(1+n)]~v^{n}$, $\Gamma(\cdot)$ being the Euler gamma function $\Gamma(z)=\int_{0}^{\infty}dt~t^{z-1}\exp[-t]$
 Similar reasonings can be done for frequencies around higher harmonics with a different definition of the detuning parameter $C$.
@@
 Then, since the integration in $dz^{\prime}$ is performed over a large number of undulator periods $N_{w}\gg 1$, all terms of the expansion in Eq.
-(147) but those for $m=-1$ and $m=0$ average to zero due to resonance approximation.
+(147) but those for m=1m=-1 and $m=0$ average to zero due to resonance approximation.
 Note that surviving contributions are proportional to $K/\gamma$, and can be traced back to the term in $\vec{e}_{x}$ only, while the term in $\vec{\theta}$ in Eq.
@@
 (147) is $v$ times the term with $n=0$ and is immediately negligible, regardless of the values of $m$.
-The term in $\vec{\theta}$ would survive averaging when $n=1,~{}m=-1$ and when $n=-1,~{}m=0$.
+The term in $\vec{\theta}$ would survive averaging when n=1,m=1n=1,~m=-1 and when n=1,m=0n=-1,~m=0.
 However, it scales as $\vec{\theta}v$.
@@
 All terms corresponding to larger values of $|n|$ are negligible.
-Summing up, all terms of the expansion in Eq. (143) but those for $n=0$ and $m=-1$ or $m=0$ give negligible contribution. After definition of
+Summing up, all terms of the expansion in Eq. (143) but those for $n=0$ and m=1m=-1 or $m=0$ give negligible contribution. After definition of
 $A_{JJ}=J_{0}\left(\frac{\omega K^{2}}{8k_{w}c\gamma^{2}}\right)-J_{1}\left(\frac{\omega K^{2}}{8k_{w}c\gamma^{2}}\right)~,$ (151)
 that can be calculated at $\omega=\omega_{r}$ since $|C|\ll k_{w}$, we have
-[…]
+${\vec{\widetilde{E}}}=-\frac{K\omega e}{2c^{2}z_{0}\gamma}A_{JJ}\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\int_{-L/2}^{L/2}dz^{\prime}\exp\left[i\left(C+{\omega\theta^{2}\over{2c}}\right)z^{\prime}\right]\vec{e}_{x}~.$ (152)
 22.3 A3. Approximating the Electron Path
@@
 The curvilinear abscissa on the path can then be written as
-[…]
-[…]
+$s(z)=\int_{0}^{z}dz^{\prime}\left[1+\left(\frac{dx}{dz^{\prime}}\right)^{2}+\left(\frac{dy}{dz^{\prime}}\right)^{2}\right]^{1/2}$ (156)
+$\simeq\int_{0}^{z}dz^{\prime}\left[1+\frac{1}{2}\left(\frac{dx_{r}}{dz^{\prime}}\right)^{2}+\frac{1}{2}\left(\frac{dy_{r}}{dz^{\prime}}\right)^{2}+\frac{1}{2}\left(\eta_{x}^{2}+\eta_{y}^{2}\right)+\eta_{x}\frac{dx_{r}}{dz^{\prime}}+\eta_{y}\frac{dy_{r}}{dz^{\prime}}\right]$ (157)
 $=s_{r}(z)+\frac{\eta^{2}z}{2}+x_{r}(z)\eta_{x}+y_{r}(z)\eta_{y}~,$ (158)
-[…]
+where we expanded the square root around unity in the first passage, we made use of Eq.
+(155), and of the fact that the curvilinear abscissa along the reference path is $s_{r}(z)\simeq z+\int_{0}^{z}dz^{\prime}[(dx_{r}/dz^{\prime})^{2}/2+(dy_{r}/dz^{\prime})^{2}/2]$.
 We now substitute Eq. (155) and Eq. (158) into Eq. (43) to obtain:
-[…]
-×$\times\left[\left({v_{x}(z^{\prime})\over{c}}-(\theta_{x}-\eta_{x})\right){\vec{e_{x}}}+\left({v_{y}(z^{\prime})\over{c}}-(\theta_{y}-\eta_{y})\right){\vec{e_{y}}}\right]~{},$ (160)
+$\vec{\widetilde{{E}}}(z_{0},\vec{r}_{0},\omega)=-{i\omega e\over{c^{2}}z_{0}}\int_{-\infty}^{\infty}dz^{\prime}{\exp{\left[i\Phi_{T}\right]}}$ (159)
+$\times\left[\left({v_{x}(z^{\prime})\over{c}}-(\theta_{x}-\eta_{x})\right){\vec{e_{x}}}+\left({v_{y}(z^{\prime})\over{c}}-(\theta_{y}-\eta_{y})\right){\vec{e_{y}}}\right]~,$ (160)
 where the total phase $\Phi_{T}$ is
@@
 Eq. (169) can be solved with the help of an appropriate Green’s function $G(z-z^{\prime},\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot})$ yielding
-[…]
+$\vec{{\bar{E}}}(z,\vec{r}_{\bot},\omega)$ $=$ $-4\pi\int_{-\infty}^{\infty}dz^{\prime}\int d\vec{r^{\prime}}_{\bot}\left(\frac{i\omega}{c^{2}}\vec{\bar{j}}-\nabla^{\prime}\bar{\rho}\right)G(z-z^{\prime},\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot})~,$ (170)
 the integration in $d\vec{r^{\prime}}_{\bot}$ being performed over the entire transverse plane. An explicit expression for the Green’s function to be used in Eq. (170) is given by
@@
 Since it is assumed that electrons are moving along the $z$-axis, we have $\vec{\bar{j}}_{\perp}=0$. Thus, after integration by parts, we obtain the solution
-[…]
-×$\times\frac{1}{z-z^{\prime}}{\rho_{\bot}}\left(\vec{r^{\prime}}_{\bot}\right)\left(\frac{\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}}{z-z^{\prime}}\right)~{}.$
+$\vec{\widetilde{E}}_{\bot}(z,\vec{r}_{\bot})$ $=$ $-\frac{i\omega}{c}\bar{f}(\omega)\int_{0}^{z}dz^{\prime}\int d\vec{r^{\prime}}_{\bot}\exp\left\{i\omega\left[\frac{\mid\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}\mid^{2}}{2c(z-z^{\prime})}\right]+i\frac{\omega z^{\prime}}{2c\gamma^{2}}\right\}$ (182)
+$\times\frac{1}{z-z^{\prime}}{\rho_{\bot}}\left(\vec{r^{\prime}}_{\bot}\right)\left(\frac{\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}}{z-z^{\prime}}\right)~.$
 Eq. (LABEL:ggeneralfin) describes the field at any position $z$.
@@
 $\vec{\widetilde{E}}_{\bot}(z,\vec{r}_{\bot})$ $=$ $-\frac{i\omega\bar{f}(\omega)}{c}\int d\vec{r^{\prime}}_{\bot}{\rho_{\bot}}\left(\vec{r^{\prime}}_{\bot}\right)\exp\left[\frac{i\omega z}{2c\gamma^{2}}\right]\Bigg{\{}\left[\frac{ic}{\omega}\frac{\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}}{\mid\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}\mid}\cdot\frac{d}{d\left[\mid\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}\mid\right]}\right]$ (185)
-×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\int_{0}^{\infty}\frac{d\xi}{\xi}\exp\left[+i\omega\frac{\mid\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}\mid^{2}}{2c\xi}-\frac{i\omega\xi}{2c\gamma^{2}}\right]\Bigg{\}}$
+$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\times\int_{0}^{\infty}\frac{d\xi}{\xi}\exp\left[+i\omega\frac{\mid\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}\mid^{2}}{2c\xi}-\frac{i\omega\xi}{2c\gamma^{2}}\right]\Bigg{\}}$
 We now use the fact that, for any real number $\alpha>0$:
-[…]
+$\int_{0}^{\infty}{d\xi}\exp\left[i\left(-\xi+{\alpha}/{\xi}\right)\right]/{\xi}=2K_{0}\left(2~\sqrt{\alpha}\right)~,$ (186)
 where $K_{0}$ is the zero order modified Bessel function of the second kind. Using Eq. (186) we can write Eq. (185) as
-[…]
+$\vec{\widetilde{E}}_{\bot}(z,\vec{r}_{\bot})=\frac{i\omega\bar{f}(\omega)}{c}\exp\left[\frac{i\omega z}{2c\gamma^{2}}\right]\int d\vec{r^{\prime}}_{\bot}{\rho_{\bot}}\left(\vec{r^{\prime}}_{\bot}\right)$ (187)
 […]